A Complete Bibliography of Scientific American: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

17 October 2018 Version 1.21

Title word cross-reference

# [3570]. #1 [1056].

$1.4 [2908]. $28.95 [2490]. $29.99 [2489]. 3 [775, 2967]. $30 [2491]. 3× [2670]. 4 [239]. _{2} [55, 1742]. G [1673].

-d [2967, 775, 239].

000-Mile [1631]. 000-Year-Old [2131]. 0425 [2437].

140-Character [124]. 170th [2471].

21st [2499, 1417]. 21st-century [2499].

5G [3747].

8 [589].

Abandoned [1780, 2425, 1429]. Ability [2280]. able [3752]. Abortion [1765, 3715]. Above [1519, 2995, 1395, 2811]. Absence [2052]. Absolute [498]. Absurd [2090]. Academia [2652]. Access [3096]. Accidental [2062]. Accidentally [900]. Achieve [1592]. Achilles [315, 3120]. Acid [3420, 2562]. Acoustic [144, 3047]. Acquisitions [206]. across [2063]. Act [974, 1864, 785, 950, 2233, 1872]. Action [358]. Actions [205]. Activities [613]. Actual [3420]. Actually [2004]. Actuary [760]. Acupuncture [2918]. Ad [2774]. Adams [2482]. Adapt [1456]. Adaptable [2870]. Add [1806, 2295, 2028]. Addiction [1616, 179]. Ado [1001]. Adolescent [3521]. Adoption [139]. Adrenaline [1416]. Adults [1089]. Advance [2275]. Advances [863, 3200, 3616, 1989, 802, 728, 842, 3183, 2036, 471, 618, 544, 1326, 2186, 2714, 2906, 2238, 2444, 2546, 2991, 2653, 2877, 3475, 1550, 983, 1248, 507, 1397, 3702, 2153, 3394, 3242, 1362, 1166, 1951, 2744, 3040, 1052, 581, 3560, 1636, 916, 3187, 1798, 2113, 3091, 3448, 1126, 2613, 2936, 3271, 3674, 2776, 1438, 2344, 3532, 688, 1210, 3735, 3157, 2581, 3420, 3772, 850, 2312, 3647, 3189, 3767, 766, 878, 1283, 2073, 1761, 1913, 2412, 2684, 276, 314, 397, 434, 1089, 1018, 1478, 1679, 1874, 2510, 2476, 350, 2272, 854, 3193, 3301, 3195, 1597, 653, 3329, 2807]. Advances [3197, 3504, 950, 859, 1717, 3589, 2379, 1836, 1515, 3358, 3130, 3214]. Advantage [745]. Adventures [2092]. Advertising [2984, 2392]. Advice [1368, 2245, 454]. Aeronautics [2180]. Aesop [2516]. Affair [170, 1161]. Affairs [2837]. Affordable [1743]. Afghanistan [755]. Afraid [2974]. Africa [904, 2036, 2706, 2463, 3357, 2801]. After [834, 421, 696, 2107, 3032, 1348, 2535]. Afterimages [2971]. Afterlife [50, 1919, 2699]. Again [2561, 3185, 3216, 2716, 3236]. Against [1747, 1902, 1792, 3229, 112, 3496, 445, 1509, 756]. Age [1415, 1893, 2053, 2286, 3273, 969, 1376, 190, 3195, 2940, 2762, 3020, 1592, 1569]. Ageless [655]. Agencies [3474]. Agenda [348, 312, 726, 579, 542, 432, 686, 469, 505, 395, 616, 651, 764, 800, 1087, 981, 1208, 1050, 914, 1246, 1016, 862, 1124, 1164, 876, 948, 1324, 1281, 1476, 1677, 1513, 1634, 1360, 1395, 1436, 1715, 1759, 2184, 2151, 1796, 1834, 1949, 2071, 1911, 1872, 1987, 2111, 2034, 2342, 2270, 2474, 2508, 2236, 2544, 2410, 2611, 2579, 2310, 2377, 2442, 2682, 3089, 2651, 2712, 2774, 2875, 3038, 2904, 2805, 2742, 2989, 2934, 3212, 3356, 3418, 3446, 3128, 3155, 3240, 3327, 3392, 3269, 3199, 3299, 3700, 3733]. Agenda [3765, 3587, 3502, 3558, 3672, 3473, 3530, 3645, 3614, 1392, 274, 1595, 1548]. Agent [2831]. Agents [139]. aggeration [593]. Aging [2455]. Ago [1671, 3, 27, 53, 79, 108, 134, 172, 199, 244, 273, 311, 347, 394, 431, 468, 504, 541, 578, 615, 650, 708, 725, 763, 799, 841, 875, 913, 947, 980, 1015, 1049, 1086, 1123, 1163, 1207, 1245, 1280, 1323, 1359, 1394, 1435, 1475, 1544, 1590, 1630, 1676, 1714, 1758, 1795, 1833, 1871, 1910, 1948, 1985, 2033, 2070, 2110, 2150, 2183, 2226, 2269, 2301, 2341, 2376, 2409, 2441, 2473, 2507, 2543, 2578, 2610, 2650, 2681, 2711, 2741, 2773, 2804, 2874, 2903, 2951, 2988, 3037, 3088]. Ago [3127, 3154, 3182, 3211, 3239, 3268, 3298, 3343, 3355, 3391, 3434, 3445, 3488, 3501, 3529, 3557, 3586, 3613, 3644, 3671, 3699, 3732, 3782]. Agony [2839]. Agree [1898, 1882]. Agriculture [573, 3351, 3554, 2900, 3123, 2604, 3204, 2144, 907]. Ahead [544, 291, 2345, 2584]. Ai [2825, 2825, 3293, 3685, 3196]. Aid [3713, 14]. Aided [3313]. Ails [1311]. Aim [1360]. Air [1264, 3144, 2019, 166, 720, 3188, 680, 1708]. Airplane [2911]. Aisle [1480]. Al [629]. Alabama [1837]. Alchemy [2007]. Alcohol [3300]. Algebra [9]. Alien [3771, 2496]. Aliens [462, 601]. Alignment [812]. Alike [522, 2267]. Alive [2129, 959]. All-There-Is [2089]. Allergy [2907, 3165]. Ally [756]. Almost [990]. Alone [3729]. Alpha [3201]. Alpinists [1268]. Also [1072, 1108, 1146, 1188, 1232]. Alter [2781]. Alvy [3517]. Always [1777, 2528]. Alzheimer [1732, 3229, 3701, 2402]. Alzheimer’s [164]. Am [3202, 2672, 3074]. Amazing [2432, 2870, 1617, 69]. Amazon [904, 424, 3301]. Ambition [2018, 1156]. America [1969, 1240]. American [2201, 2428, 2594, 3607, 1867, 1871, 1910, 1948, 1985, 2033, 2070, 2507, 3211, 3239, 3268, 3298, 3343, 1671, 2643, 3, 27, 53, 79, 108, 134, 172, 199, 244, 273, 311, 347, 394, 431, 468, 504, 541, 578, 615, 650, 708, 725, 763, 799, 841, 875, 913, 947, 980, 1015, 1049, 1086, 1123, 1163, 1207, 1245, 1280, 1323, 1359, 1394, 1435, 1475, 1544, 1590, 1630, 1676, 1714, 1758, 1795, 1833, 2110, 2150, 2183, 2226, 2269, 2301, 2341, 2376, 2409, 2441, 2473, 2543, 2578, 2610, 2650, 2681]. American [2711, 2741, 2773, 2804, 2874, 2903, 2951, 2988, 3037, 3088, 3127, 3154, 3182, 3355, 3391, 3434, 3445, 3488, 3501, 3529, 3557, 3586, 3613, 3644, 3671, 3699, 3732, 3782]. Americans [787]. Amoebas [2117]. among [1432]. Amps [2353]. Analgesics [249]. Analysis [268]. Anatomy [658, 909, 2738]. Ancestor [773]. Ancestry [1352]. Ancient [1114, 1536, 1839, 1916]. Ancillary [2169]. Angry [1577]. Anguish [1429]. Animal [1344, 1991, 1931, 3058, 1082, 3150, 865, 75, 538, 2435, 1534, 836, 228, 401, 1828, 2235, 1753]. Animals [2666, 2677, 1534, 900, 3166, 2121, 69, 3634]. Annotated [1303]. Announced [2747]. Anonymity [470]. Ansari [2490]. Answers [2090, 3492, 2946]. Ant [1442]. Antarctic [1080]. Antarctica [1145]. Anthrax [3236]. Anthropology [265, 3122]. Anti [188, 17, 124, 370, 43, 335, 150, 297, 68, 260, 216, 95, 455, 673, 785, 602, 491, 528, 565, 707, 637, 819, 749, 418, 1074, 1148, 968, 1036, 1190, 1110, 1002, 853, 1234, 1269, 935, 900, 1460, 1543, 1670, 1589, 1383, 1738, 1501, 1348, 1419, 1703, 1629, 1306, 1860, 1974, 1822, 2094, 2173, 2204, 1897, 1784, 1937, 2057, 2011, 2131, 2256, 2563, 2493, 2430, 2629, 2290, 2361, 2462, 2527, 2398, 2330, 2599, 2821, 2792, 2892, 3004, 3106, 2670, 2730, 2950, 3055, 2833, 2700, 2760, 2922, 3143, 3257, 3316, 3227, 3286, 3373, 3171, 3190, 3433, 3462, 3342, 3407, 3487]. Anti [3518, 3629, 3716, 3602, 3575, 3781, 3546, 3660, 3688, 3750]. Antibiotic [1888]. Antibiotics [3110]. Antidote [2569]. Antifreeze [3428]. Antimatter [107]. Antioxidants [1355]. Antiscience [3663]. Antiseptic [1750]. Ants [149, 2704, 836]. Anxiety [1554, 2028]. Anybody [1448, 3487, 1942]. Anything [842]. Anyway [3226]. Anywhere [3487]. Apocalypse [2422, 268, 3196]. App [1692, 547]. Appears [81]. Appendix [930]. Appetite [886]. Apple [188]. Apprentice [910]. Approach [8]. Approaches [3076]. Approximately [1497]. Apps [3626]. April [2745]. Archaeoastronomy [2574]. Archaeology [167, 683, 3470, 787, 497, 2026, 3259, 2221, 2601, 3294]. Architects [1755]. Architecture [1931, 2179]. Arctic [215, 129, 2063, 1247]. Aren’t [1103]. Arid [516]. Arm [2002, 837, 1216]. Armor [909, 739, 1509]. Arms [2579, 2821, 1286]. Arousal [2839]. Arrival [2128]. Arrives [2588]. Art [58, 2770, 1519, 47, 1276, 2980, 836, 3695, 3704, 1553, 1173]. Art-Eating [3704]. Artificial [1083, 3753, 534, 3548, 3768, 2354, 3042]. Artist [1807, 3515]. Ascent [2823]. Aside [2290]. Ask [258, 37, 2249]. Aspirin [3253]. Assassin [2655]. Assault [3356]. Assembled [1467]. Asteroid [492, 2930, 162]. Asteroids [2238]. Asthma [524]. Astro [3675]. Astrobiology [789]. Astronauts [789]. Astronomer [2668]. Astronomers [2006]. Astronomy [259, 2794, 96, 1787, 3631, 3786, 2433, 3119, 3559, 3605, 2258, 300, 2217, 1531, 1899, 2708, 1863, 2895, 3071, 3579, 2537, 2067, 1706, 2645, 44]. Astrophysics [751, 2293, 3260, 3230, 640, 2464, 2025, 1039, 125, 938, 2572, 463, 3691, 1754, 2824, 3073, 3520, 3349, 3495, 1112, 1386]. Atheist [1379]. Atheists [2699]. Atlantis [3301]. Atmospheric [1116]. Atom [752, 1712, 2353]. Atomic [2127, 2212, 1804]. Atomic-Scale [2212]. Atoms [2009, 1528]. Attack [750, 941, 900, 248]. Attacking [2014, 3039]. Attention [3092]. Audience [2163]. Auditory [2939]. Aunt [1073]. Aura [1681]. Aurora [2459]. Aussie [3330]. Autism [1241, 1210, 1661, 2866, 2939, 306]. Auto [3646]. Automating [3791]. Automotive [3572]. Autopilot [744]. Autopsy [1226]. Averse [1384]. Aviator [2973]. Avoid [3299, 1998]. Avoidable [3542]. Avoided [141]. awareness [158]. Awash [3335]. Away [2285, 885, 2205, 2156, 882, 84]. Awe [696, 1233, 1859]. Awful [3572]. Aziz [2490].

Ba [3684]. Ba-Boom [3684]. Babbling [2417]. Babel [2357]. Babes [673]. Babies [2677, 195, 2852, 3207, 2048]. Baby [3089, 3411, 1605, 638, 3317, 2417, 3361, 2606, 1962, 3632]. Back [2092, 3392, 1091, 3786, 723, 552, 735, 3569, 3693, 595, 6, 3403]. Backfire [3142]. Backseat [922]. Backup [3554]. Backyard [1395]. Bacon [1460]. Bacteria [85, 1721, 2630, 1453, 609, 957, 2240, 2259]. Bacteria-Style [1721]. Bacterial [47, 3535, 1681]. Bacteriology [1120]. Bad [1107, 2869, 1057, 1721, 647, 1453, 3447, 1976, 2113, 2871, 3421, 2563, 140]. Baddest [3550]. Bagel [1306]. Baggage [887]. Baghdad [437]. Bake [2426]. Balance [1641, 3739, 103]. Balancing [785, 2233]. Ball [2158, 216]. Balloon [1065]. Ballots [2657]. Baloney [2759]. Bambi [2116]. Ban [726, 2342]. Bananas [43]. Band [499]. Bandage [2367]. Bang [1653, 2590, 700, 225, 276, 929, 3437, 355, 2621, 2143]. Bank [3491]. Banking [2618]. Bankroll [1656]. Banning [1835]. Bans [1796]. Baobab [1646]. Bar [3740]. Barely [1417]. Baron [759]. Barrier [402, 1505]. Barriers [3642]. Based [1208]. Basic [2489]. Basics [2450, 3403]. Bat [2084]. Bathtub [61]. Bats [2417, 3133]. Batter [583, 440]. Batteries [2215, 3400, 3109]. Battery [2453, 1354, 3254]. Battery-Powered [1354]. Battle [2201, 2299, 3666, 1874, 2604, 3598]. Batty [2616, 3589]. Bayes [556]. Bazaar [714]. Be [304, 2442, 2934, 3733, 2722, 1154, 1554, 70, 2964, 3106, 3140, 3207, 3393, 3615, 2255, 3060, 3601, 1746]. Beacon [2143]. Beacons [1421]. Beaks [2916]. Bears [3302]. Beast [3577]. Beasts [627, 1705]. Beat [681]. Beaten [2715]. Beating [552]. Beauties [2671]. Beautiful [1736, 1414, 1600, 1031]. Beauty [627, 3440]. Became [2425]. Because [1970, 3419]. Beckons [1831]. Becoming [2142]. Bed [179]. Bee [3554, 1042, 1964, 3518]. Been [2358]. Beer [440]. Bees [2009, 2279, 1739]. Beetle [3468]. Beetles [1337, 1404, 391]. Beetroot [2276]. Before [536, 222, 1369, 1329, 1368, 3381]. Begin [3639]. Beginning [2059, 3325, 3319]. Begins [945]. Behave [2865]. Behavior [1106, 1082, 865, 75, 538, 1307, 2704, 1534, 836, 1828, 1753]. Behavioral [1976, 3203]. Behemoth [3540]. Behind [1625, 1249, 2729, 1952, 784]. Being [1974]. Beings [1417]. Beliefs [1669]. Believe [3560]. Believing [2963, 636]. Belong [263]. Below [1756]. Belt [2248, 3598]. Bend [2188]. Beneath [2325, 3040]. Benefactor [2683]. Benefits [2872]. Benevolence [1112]. Best [2421, 1267, 2594, 641, 687, 1194, 2172, 2139]. Bestiary [1417]. Better [710, 3212, 1403, 2915, 1868, 3351, 440, 1685, 2258, 654, 1090, 46, 720, 1010, 852, 130, 715, 36, 2615, 3254, 3400, 677, 2855, 325, 2528, 1117]. Betting [1584]. between [3642]. Beware [1324, 1759, 2682]. Beyond [1150, 1106, 1160, 1367, 1438, 2175, 1549, 3379, 3727, 561, 2868, 1406]. Bias [3791, 2580]. Big [536, 1653, 505, 2567, 2565, 2444, 1747, 3555, 12, 1569, 1287, 290, 3094, 1462, 1656, 915, 2256, 2792, 1673, 624, 3437, 3492, 549, 746, 2121, 30, 1772, 1873, 2621, 2143, 2064, 2021]. Bigfoot [1764]. Bigger [713, 36, 767, 2528]. Biggest [2747, 276, 3635]. Bilingual [3248, 3623]. Bill [102]. Billion [1665, 2706, 2908, 3204]. Billion-Dollar [3204]. Billions [2370]. Bind [1768]. Binds [2401]. Bingo [1543]. Biochemistry [1472, 1204, 3441]. Biodiversity [3523, 1119, 1168]. Bioengineering [1391, 910, 1788, 2768, 3080]. Bioethics [263]. Biofuel [85]. Biofuels [678]. Bioinformatics [2021]. Biological [200, 3236, 2271]. Biology [1241, 3058, 2023, 1114, 386, 2018, 535, 2466, 3292, 305, 3665, 2736, 607, 2030, 1790, 41, 1387, 2405, 644, 1670, 461, 3577, 2145, 3233, 3175, 760, 3324, 2061, 870, 833, 105, 2931, 2261, 19]. Biomechanics [2799, 2928]. Bionic [2915, 1317, 103, 1717]. Biophilia [2393]. Biotechnology [2959, 3079, 943, 793]. Bipedal [2899]. Bipolar [890]. Bird [3568, 1171, 1239, 2315, 3536, 1597, 1828, 325, 1340, 437]. Birds [2757, 1994, 1882, 2362, 1577]. Birdsong [854]. Birth [2394, 2436, 975]. Birthday [1025]. Births [3107, 172]. Bisexual [2842]. Bison [472, 2657, 2256]. Bitcoin [3492]. Bite [3550]. Bites [1439]. Bits [1546]. Bitter [2703]. Black [2059, 2597, 2425, 2293, 60, 3319, 867, 943, 3057, 1478, 1874, 3520, 350, 3569, 606, 2537, 1112]. Blackett [1345]. Blame [446]. Blast [2479]. Blind [2508, 3736, 1644, 3171, 3179, 120, 1532, 128]. Bling [42]. Blitzkrieg [2450]. Blizzard [2756]. Blizzards [2009]. Block [1868]. Blockchain [3493]. Blocking [941]. Blogger [3384]. Blood [2007, 2726, 3000, 385, 2891, 2940, 2026, 2228]. Bloodstream [3160]. Bloodthirsty [3677]. Bloody [2276]. Bloom [957]. Bloomsbury [2491]. Blows [1022]. Blowtorch [1488]. Blue [3555, 3655, 3052, 957, 459]. Blues [3052, 3482]. Blunders [1455]. BMI [1685]. Bodies [3058]. Body [3066, 920, 2030, 3682, 3739, 464]. Body-Snatching [920]. Bodyguards [2703]. Bold [1314]. Boldly [3180]. Bomb [1017, 2271]. Bombarded [674]. Bombs [3650]. Bond [176]. Bone [3702, 1102]. Bonobo [1379, 2841]. Bonobos [3561]. Boogie [3302]. Book [1144, 1417, 2597, 2595, 2491, 2489, 2490, 2627, 2596, 2626, 2488, 2594, 2625, 2558, 2668, 2667, 2726, 2758, 2697, 2757, 2698, 2666, 2727, 2728, 2756, 2696, 2665, 2695, 2725, 2755, 3629]. Books [2489, 1072]. Boom [1557, 1658, 30, 3684]. Boost [3623, 2995, 56]. Boosting [2894]. Borders [826]. Boring [1337]. Born [2794, 538, 1440, 1996, 2228, 3105]. Borrowing [379]. Boson [1129]. Boss [1463, 1373]. Bot [2384]. Botany [1008, 1508]. Both [2582]. Botox [169]. Bots [2554, 3160]. Bottom [2077, 1383, 1904]. Bottoms [2029]. Botulism [1768]. Bouillabaisse [1670]. Bounce [595]. Bound [973]. Boundaries [99]. Boundary [982, 2949]. Bounds [1121]. Boy [647]. Boys [842, 2844]. Bracing [549]. Brain [1231, 1105, 972, 908, 2894, 2910, 338, 3006, 3377, 1831, 456, 2635, 3521, 3582, 937, 2432, 2870, 366, 2707, 2871, 402, 1505, 1707, 1427, 1696, 1998, 2581, 3420, 1905, 1716, 2975, 2662, 3564, 3740, 1040, 2520, 2465, 1825, 823, 3761, 2734, 3203, 2623, 1350, 71, 1199, 3533, 636, 7, 412, 3458, 1752, 1125, 1331, 1862, 1836]. Brain-Boosting [2894]. Brains [2253, 1057, 1171, 1546, 2986, 786, 3666, 716, 3248, 3145, 809, 673, 595]. Brainy [3094, 702, 1828]. Brakes [3564]. Branches [2659]. Brave [1624]. Breaching [402]. Breadbox [2315]. Break [1222, 2936, 2647, 3760]. Breakdown [200]. Breaker [3177]. Breaking [1721, 1224, 1505, 3491, 101, 2563, 65]. Breakup [3536]. Breast [2996]. Breath [1453, 1527, 977, 680]. Breathe [1431]. Breathes [1605]. Breathing [3109]. Bred [1074]. Breed [325]. Breeding [131]. Brethren [213]. Brian [1972]. Bridge [2272]. Bridges [737]. Brief [2667, 2728, 2536]. Brigade [1836]. Bright [471, 1600, 984, 433, 3001, 411]. Brightest [687]. Brilliance [3382]. Brilliant [1455]. Bring [2185, 2745, 1148, 2383]. Bringing [2092, 3089, 3693]. Brink [1646, 1200]. Briny [2478]. Brittle [910]. Broadband [36]. Broken [3334, 39]. Brokers [3009]. Bronx [2256, 3133]. Brown [2478, 2050]. Brows [58]. Bubbles [700, 2025]. Buck [1097, 412]. Buckles [1371]. Budget [2554, 184]. Bug [1561, 1334, 2494, 229, 3204]. Bug-Eared [1334]. Bugs [1057, 318, 552, 3591, 985, 1720, 2750]. Build [2453, 1663, 2860, 677, 2855, 3265, 339]. Build-a-Battery [2453]. Building [3351, 3554, 3667, 2027, 2499, 3254, 1117]. Buildings [737, 716]. Built [3145]. Bulb [852]. Bullet [1354]. Bullets [2474, 144]. Bullies [3561]. Bum [2750]. Bummer [455]. Bump [2057]. Burden [633]. Burger [2746]. Burial [1839]. Buried [1612, 755]. Burn [2758]. Burning [471, 2363]. Business [1415, 1326, 3716]. Bust [2189]. Busting [3000]. Busy [1042]. Butchers [277]. Butt [2599]. Button [1183]. Butts [1340]. Buying [852]. Buzz [3193, 3668, 1479]. Buzzword [624]. Buzzworthy [3011]. bye [2999]. Bytes [2334].

C [1902, 2021]. Cable [2272]. Cable-Stayed [2272]. Caffeinated [1854]. Cage [1953]. Calamity [794, 1986]. Calculus [1907]. California [1726]. Calisthenics [2854]. Call [2196, 358, 2173, 2684, 3685, 217, 962]. Called [3712]. Calls [951]. Calories [1618]. Cam [3404]. Came [2097, 3016, 3498]. Camera [1528]. Cameraprints [1925]. Cameras [2640, 2184, 2216]. Campaign [2651, 3038]. Campus [624]. Can [3494, 1195, 3680, 1087, 1107, 842, 2722, 2293, 694, 791, 1152, 2912, 3641, 2481, 474, 1744, 533, 2936, 1554, 2822, 3357, 2050, 1155, 2615, 3615, 296, 2940, 3632, 2455, 3308, 3164, 1745]. Cancer [1858, 2013, 3253, 3076, 2765, 732, 1407, 903, 1926, 1826, 1709, 610, 2366, 2312, 2176, 2365, 3422, 692, 547, 398, 1875, 3206, 2121, 1529, 3578, 3571, 3777, 2946, 3075, 2764, 1940, 756]. Cancerous [2751]. Candy [178]. Cannabis [2936]. Can’t [1071, 3299, 802, 2858, 262, 1660]. Caps [187]. Captchas [931]. Captives [3470]. Captivity [1773]. Capture [2676, 2215, 3673]. Car [2776, 230]. Carbon [55, 2676, 1711, 439, 3673, 166, 3109]. Carbon-Breathing [3109]. Carbs [112]. Carcinogens [89]. Card [3262]. Cardio [112]. Cardiology [3147]. Care [2712, 641, 2266, 3491]. Career [2280]. Careful [17]. Cars [1884, 546, 829, 314, 2827, 3081]. Cartoon [1538]. Case [2427, 3148, 1026, 1792, 1866, 1276, 3547, 2343, 1611, 1153, 1835, 2152, 2437]. Cassava [131]. Cassettes [2947]. Cassini [3415]. Casting [392]. Castles [720]. Casualties [426]. Cat [2524]. Catalog [2469]. Catalysts [118]. Catastrophic [211]. Catch [54, 2791]. Catcher [1978, 1335]. Catching [3608, 2164]. Caterpillar [3679]. Cats [1855, 3094, 3395, 1190, 667]. Cause [1529]. Caution [2184]. Cautionary [2427]. Cave [2669]. Caveman [1257]. CDC [3357]. Celebrating [197, 1615]. Celebration [752, 2608, 607]. Celebrity [3124, 3012]. Cell [535, 805, 760, 56, 2788, 1788, 2145]. Cells [2208, 2210, 3665, 323, 2434, 233, 1350]. Cellular [2737, 2405]. Census [1994]. Centauri [3201]. Centennial [1986]. Central [2203]. Century [513, 1506, 1862, 2499, 1417]. Century-Old [513]. Chairs [2165]. Challenged [102]. Challenges [2565]. Chamber [3234]. Chameleon [111]. Champ [889]. Chance [2512, 44]. Change [215, 2502, 834, 25, 3086, 2336, 151, 1273, 3049, 1424, 642, 1906, 3363, 499, 3121, 3692, 3749, 2043, 129, 76, 426, 30]. Changed [3470]. Changes [1618]. Changing [372, 383, 821, 1740, 2631, 3108, 3606, 115, 2206, 1271, 2207, 1752].

Chaos [2794, 535]. Chappe [1110]. Character [124]. Charade [2223]. Charge [1097, 208, 1130]. Charging [2214, 445]. Charmers [75]. Charts [2514]. Chaser [1843]. Chasing [1266, 1219]. Chaste [3376]. Cheap [2714]. Cheaper [1403]. Check [2925, 2446, 2354]. Checking [1555]. Checks [737]. Cheese [2493, 220]. Cheeseburger [894]. Cheetahs [3744]. Chelyabinsk [1451]. Chemical [2640, 90, 336]. Chemically [1250]. Chemicals [3418]. Chemistry [752, 2007, 3551, 2770, 3264, 1535, 1507, 3345]. Chemo [2198]. Chemotherapy [3132]. Chernobyl [2743]. Chestnut [1867]. Chickens [846]. Chickpea [1135]. Chief [308]. Child [728, 2867, 3527, 2606, 1661, 2866, 1429, 2854]. Children [1780, 2869]. Chimp [726]. China [1195]. Chip [1747, 3276, 823, 478, 1100]. Chocolate [278, 907]. Choice [2843, 3348]. Cholesterol [781, 57, 1137]. Christmas [1252]. Chronic [358]. Chronicled [1671, 3, 27, 53, 79, 108, 134, 172, 199, 244, 273, 311, 347, 394, 431, 468, 504, 541, 578, 615, 650, 708, 725, 763, 799, 841, 875, 913, 947, 980, 1015, 1049, 1086, 1123, 1163, 1207, 1245, 1280, 1323, 1359, 1394, 1435, 1475, 1544, 1590, 1630, 1676, 1714, 1758, 1795, 1833, 1871, 1910, 1948, 1985, 2033, 2070, 2110, 2150, 2183, 2226, 2269, 2301, 2341, 2376, 2409, 2441, 2473, 2507, 2543, 2578, 2610, 2650, 2681, 2711, 2741, 2773, 2804, 2874, 2903, 2951, 2988, 3037, 3088]. Chronicled [3127, 3154, 3182, 3211, 3239, 3268, 3298, 3343, 3355, 3391, 3434, 3445, 3488, 3501, 3529, 3557, 3586, 3613, 3644, 3671, 3699, 3732, 3782]. Chronobiology [2294]. Chunk [3462]. Cigarettes [2905, 1929]. Circle [2318, 3085, 2731]. Circles [3696]. Circuit [2002, 805]. Circuits [1482, 97]. Circuses [1074]. Cities [710, 686, 713, 2186, 3320]. Citizen [458, 1356, 906, 1196]. Citizens [1511]. Citrus [3106]. City [717, 684, 3295, 719, 722, 3727, 2528]. Claims [580]. Clap [998]. Class [2758]. Classification [2721]. Classroom [1635, 1125]. Claw [1809]. Clean [579, 529, 2878, 193, 1140, 3062]. Cleaning [2535]. Cleanup [2714, 2942]. Clear [2712, 1684]. Clearing [734]. Clever [2892, 1297]. Click [1381, 2664, 882, 364]. Clicks [3753, 3449]. Cliffhanger [896]. Climate [2508, 2502, 116, 4, 944, 2714, 1711, 1159, 834, 1319, 2336, 1506, 1898, 1273, 2829, 642, 344, 1906, 3156, 3469, 2220, 2647, 1115, 1353, 718, 65, 499, 3121, 3692, 1635, 3262, 2043, 129, 2735, 3083, 76, 2099, 3019, 426]. Climategate [29]. Climbing [967]. ClimeApocalypse [2056]. Clinical [3700, 2016]. Clock [3567, 15]. Clocks [3665, 2294]. Clone [2396]. Clones [2992]. Cloning [152]. Close [1372]. Closer [1492]. Closing [293]. Clothes [3114]. Clots [3000]. Clotting [2230]. Cloud [1335, 943, 3469, 1478, 1815]. Clouds [1116]. Cloudy [2512, 2187, 44]. Club [954, 596]. Clue [2412]. Clues [1798, 3655, 1709]. Clusters [1386]. Clutch [3143]. Co [2751, 55, 1742]. Co-conspirators [2751]. Coal [471, 11]. Coast [55]. Coat [2597, 2887]. Cobalt [3482]. Cobb [2489]. Cocaine [701]. Cockroach [1024, 1876]. Cocktail [534]. Cocoon [343]. Code [3158, 1467, 379, 3383, 3795, 3796, 2489]. Codes [3740]. Coding [1176, 2926]. Coffee [958, 2347, 2144]. Cognition [2677]. Cognitive [2894, 460, 3524, 3328, 1982, 2762]. Cold [1402, 414, 111, 3139, 962, 975]. coli [660]. Collar [436]. Collecting [3674]. Collective [2704]. Colleges [2071]. Collide [3378]. Collider [2806]. Collision [908, 162]. Color [1011]. Colors [49, 2887, 2858, 2972, 3208]. Combinations [2556]. Come [959, 2677, 1943, 190, 860]. Comeback [1915]. Comes [270, 2777]. Comet [2708, 2046]. Comfort [3139, 1372]. Coming [579, 1557, 1319, 196, 849, 1754, 2323, 3746, 2896]. Command [2889]. Commercial [210]. Common [414]. Comparative [2018]. Comparatively [38]. Compartments [1305]. Compass [2737, 865]. Compassionate [1911]. Competition [274]. Competitor [2551]. Complexity [1582]. Component [2939]. Composites [1743]. Compromises [3398]. Computer [2, 2295, 2024, 427, 3338, 3082, 782, 3313]. Computer-Aided [3313]. Computers [825, 23, 786, 915]. Computing [1855, 3148, 794, 2403, 1314, 1559]. concentration [3491]. Conditions [982]. Condors [850]. Confederacy [1320]. Confessions [1783]. Confusion [1377, 1137]. Congress [3392, 3741, 3531]. Connect [2127]. Connected [1983]. Connection [2052, 366, 2718]. Connections [1317, 2881, 2798, 2470, 425]. Conquer [2371, 2656, 2909]. Consciousness [1525, 605, 3436, 3638]. Consensus [2628]. Consent [487]. Consequences [432]. Conservation [3014, 3410, 1276, 2898, 3095, 3468, 3526, 3346]. Consilience [2628]. Conspiracy [369, 1147, 2203]. conspirators [2751]. Constant [2052, 1673]. Contact [2697, 421]. Contagion [879]. Contagious [512]. Continent [939]. Continents [20]. Continue [2147]. Continuum [25, 2181]. Contradictions [1147]. Contrasts [2538]. Control [2091, 646, 208, 1849, 2047, 2658, 3077, 2313, 1263]. Controlled [2632, 76]. Controllers [1078]. Controlling [338]. Controls [90, 2776, 3493, 2243]. Conundrum [781, 2721, 3469]. Conversation [3741]. Conway [3798, 2491]. Cook [2588]. Cookbook [2094]. Cooking [668, 693, 485]. Cookout [1620]. Cool [1763, 1485, 1254, 3114]. Cooling [2236]. Copernicus [1792]. Coping [2956]. Cops [1686, 2184]. Copy [1822, 670]. Copycats [3051]. Coral [545, 1944, 1297]. Corals [3494, 2548]. Cord [3370]. Core [1443]. Cornboy [3204]. Corner [3747]. Corners [2641]. Correct [1497]. Correcting [2211]. Correlations [2488]. Cosmetics [3418]. Cosmic [16, 3065, 3475, 1763, 1761, 2067, 27]. Cosmigraphics [2251]. Cosmology [2059, 1653, 647, 189, 675, 337, 2602, 3319, 568, 867, 3174, 18, 2143, 2533, 269, 3288, 2733, 3442, 530, 2924, 3072]. Cosmos [640, 2464, 1679, 2895, 3071, 2540]. Cossack [2427]. Cost [3792, 709, 1307, 1428, 3117, 1430]. Costs [1672]. Couch [3624]. Coughing [1798]. Could [2806, 1914, 1745, 1749, 1746, 930, 847, 118, 3466, 2304, 1750, 3320, 269, 217, 3132, 3491]. Couldn’t [2254]. Count [1071]. Countdown [1626]. Counter [2583]. Countries [1194]. Country [186]. Couple [2847]. Couples [1457]. Coupling [2846]. Courtesy [1602]. Courtroom [532]. Cousin [3347]. Couture [1257]. Cover [1130]. Coverage [1765]. Cowboys [2475]. Cox [1972]. Coyotes [1636]. Crab [521]. Crack [2489, 1650, 3795, 3796]. Cracking [513]. Crackle [225]. Cracks [1102, 1507]. Cramming [917]. CRAPs [527]. Crash [1697, 2166, 2688]. Craving [1700, 2663, 292]. Crazy [2755, 1348]. Cream [1295]. Create [1084]. Created [3492]. Creation [2060, 2667, 156, 1357, 1282]. Creative [2852]. Creativity [575, 1198, 3313, 1385, 3029]. Creatures [1946, 1149, 2892]. Creep [1495]. Cresting [3107]. Crime [1373, 871, 681]. Criminal [3150, 2940]. Criminally [926]. Criminals [1686]. Crimp [3218]. Crisis [26, 838, 457, 1939, 2586, 255, 617, 3542]. Critical [1062, 15, 93, 67, 214, 122, 186, 41, 148, 257, 599]. Critters [2808]. Crops [1094, 3218, 516, 828]. Crossing [3662]. Crossroads [1655, 2012]. Crosswalk [2615]. Crosswalks [1636]. Crowd [181]. Crowds [688]. Crown [572]. Crucial [2876]. Cruel [1548, 2411]. Crunch [1289]. Crunchers [374]. Crushers [3702]. Cry [2279]. Crybaby [2196]. Cryogenic [668]. Cryptic [1260]. Crystal [592]. CSI [3150, 1500]. CT [1529]. Cu [3101]. Cull [1521]. Culprit [1875]. Cultivate [40]. Cultural [1068]. Culture [1198, 2389]. Cup [807]. Curb [1999]. Cure [1980, 557, 1166, 1440, 1425, 1311, 451, 2801, 306, 292]. Curing [414]. Curiosity [2886]. Curious [1106, 1592, 2054, 2491, 2852, 3798]. Currency [3492, 826]. Current [3004]. Curse [1815, 2402]. Cusp [2546]. Cut [2651, 935, 3370]. Cutting [3362]. Cybersecurity [765, 645, 2064]. Cyberwar [2805, 2372]. Cyborg [391]. Cycle [1508, 3276, 2785, 3480]. Cypress [1726].

d [2967, 775, 239]. DACA [3447]. Dad [3713, 2859]. Daddy [1689]. Dads [1773]. Daguerreotypes [1276]. Dam [2476]. Damp [175]. Dams [2445]. Dance [1024, 3324]. Dancing [2238, 602]. Danger [202, 1248, 1502, 2019, 67]. Dangerous [1184, 926, 1550, 3414, 1341]. Dangerously [257]. Dangers [866, 1628, 1890]. Dark [2595, 2991, 751, 1992, 60, 2339, 337, 2602, 3691, 93, 2412, 71, 3637, 2733, 3495, 1863]. Darkness [1869, 164]. Darwin [324, 3693, 3727]. Data [1677, 116, 1747, 1674, 1569, 1961, 750, 2999, 3143, 1662, 1033, 1356, 3512, 1772, 2134, 1873, 2064, 2021]. Data-Driven [1662]. Dataclysm [2055]. Dating [2850]. Daughters [3387, 3588]. Dave [3539]. Dawn [3237, 3448, 1531, 420, 72]. Day [3672, 421, 916, 188]. Daylight [3547]. Days [67]. Dazzling [1413, 833]. Dead [2092, 3614, 983, 2550, 533, 1214, 1148, 2324, 2574]. Deadliest [54, 3477]. Deadlock [2647]. Deadly [351, 1116, 1259, 163, 2556]. Deal [3177, 3342]. Dear [360]. Dearth [652]. Death [2008, 3410, 2674, 492, 3707, 278, 1613, 501, 66, 2820, 3424, 1247]. Deathly [3617]. Deaths [3408, 1750]. Deathwatch [2117]. Debate [903, 1673, 1891, 65, 530]. Decades [2445]. Deception [2598]. Deceptions [2968]. Deck [906]. Decline [748]. Decoded [1608]. Decoding [123, 3771]. Deed [420]. Deep [3706, 1971, 2884, 1420, 2589, 1517, 2962, 3177, 2189, 3610]. Deep-Sea [2589, 3610]. Deep-Space [3177]. Deepening [255]. Deer [2998]. Defanging [2613]. Defeat [24, 2050]. Defeating [1903]. Defense [3075, 2764]. Deficits [184]. Deflect [1745]. Defusing [2271]. Defying [348]. Degrees [901, 2612]. Delays [3222]. Deliver [651, 3143, 2017]. Deluge [834, 1369]. Delusion [1859]. Demand [1810]. Dementia [57, 1468]. Demise [3447]. Democracy [3241, 256]. Democrats [2582]. Demons [498]. Dengue [507, 2438]. Dentist [2550]. Department [871]. Depend [1208]. Depends [2054, 3493]. Depressed [2994]. Depression [2297]. Depths [3499, 925]. Dereliction [3765]. Descends [2386]. Desert [1264, 1003]. Deserted [2778]. Design [2126, 2188, 2754, 1223]. Designing [3295]. Designs [1179]. Desire [2054]. Desperate [306]. Dessert [2663]. Destiny [1324]. Destroyers [1084]. Destroying [3521]. Destruction [2404]. Detect [1926]. Detected [89]. Detectives [2006, 1182]. Detector [2603, 369]. Detects [880]. Deter [2121]. Deterrents [2084]. Developing [2020]. Development [3411, 2606, 1429, 146]. Developments [65, 255, 91, 39, 120, 184, 146, 212, 13]. Devices [3183]. Devil [610]. Diabetes [3525, 896, 3323, 744]. Diabolical [1068, 1536]. Diagnostics [796, 3117]. Diamond [3506, 1283]. Dice [2524]. Dictators [1382]. Did [167, 3639]. Didn’t [2743]. Diehards [1382]. Diet [1400, 3664, 3483, 3132]. Diets [2360]. Difference [2095, 3034, 2072]. Different [2267]. Differently [3723]. Digging [2202, 788]. Digit [1054]. Digital [1415, 1592, 1248, 3728, 1569, 2850, 992, 902, 3492, 2392, 3459, 3493]. Digital-Subscription [3459]. Digitizer [308]. Dilemma [1164, 1691, 2973, 3223, 2154, 3459]. Dimensions [153]. Din [2317]. Dining [602]. Dinner [1133, 2317]. Dinosaur [1252, 501, 759]. Dinosaurs [2595, 2644, 3604, 939, 1950, 1432, 3208]. Diplomacy [396]. Direction [2155, 2998]. Dirt [1622, 1032]. Dirty [1326, 191]. Disappearing [1276]. Disappointment [1413]. Disaster [2476]. Disasters [3762]. Disclosure [1949]. Discontent [1273]. Discovered [2006]. Discovery [1671, 2181, 3669, 3546, 1880, 3, 27, 53, 79, 108, 134, 172, 199, 244, 273, 311, 347, 394, 431, 468, 504, 541, 578, 615, 650, 708, 725, 763, 799, 841, 875, 913, 947, 980, 1015, 1049, 1086, 1123, 1163, 1207, 1245, 1280, 1323, 1359, 1394, 1435, 1475, 1544, 1590, 1630, 1676, 1714, 1758, 1795, 1833, 1871, 1910, 1948, 1985, 2033, 2070, 2110, 2150, 2183, 2226, 2269, 2301, 2341, 2376, 2409, 2441, 2473, 2507, 2543, 2578, 2610, 2650, 2681, 2711, 2741, 2773, 2804, 2874, 2903]. Discovery [2951, 2988, 3037, 3088, 3127, 3154, 3182, 3211, 3239, 3268, 3298, 3343, 3355, 3391, 3434, 3445, 3488, 3501, 3529, 3557, 3586, 3613, 3644, 3671, 3699, 3732, 3782]. Disease [3299, 3606, 422, 646, 2701, 1751, 1043, 2603, 2723, 3221, 130]. Diseases [358, 495, 24, 1389, 1471]. Disestimation [362]. Disguise [574]. Disgusting [1597]. Dish [1846, 495, 653]. Disobedience [3148]. Disorientation [3543]. Dispersants [251]. Dispirit [2379]. Displays [2211, 1413]. Dissent [3243]. Distance [2822]. Distant [1531]. Distraction [1553]. Dithering [2742]. Diverse [2309]. Diversions [1998]. Diversity [3700, 2135]. Divide [2036, 2656, 1054, 2909, 3662, 1647]. Divided [3092]. Diving [3534, 2264]. Divorce [2869]. DIY [1963]. DNA [2328, 1114, 1807, 2249, 190, 382, 1390, 3567, 2807, 109, 376]. Do [1499, 1281, 1476, 2151, 1970, 842, 713, 2190, 2677, 2843, 2249, 1155, 2615, 3452, 2947, 3283, 3484, 1529, 669, 3164, 2994]. Doc [1400]. Docs [2474]. Doctor [2477, 2416, 622, 968, 1480, 1266]. Does [1908, 675, 263, 436, 992, 2004, 2280, 258]. Doesn’t [2465, 1067, 1669]. Dog [2468, 3265]. Dogged [2235]. Dogs [3526]. Doing [1638, 121]. Dolfriends [1766]. Dolittle [1266]. Dollar [3204, 2125]. Dolphin [1391]. Dolphins [1471]. Dome [3245]. Donald [3038]. Donate [412]. Done [1366]. Donkey [2007]. Donor [628]. Donors [113]. Don’t [1104, 2270, 2508, 2875, 2037, 1524, 825, 3560, 2880, 828, 3257, 415, 1778, 3235, 1030, 1217, 3164]. Doomsday [3457, 15, 2791]. Door [2248]. Dose [79]. Doses [1587]. Double [1303, 3678, 1033]. Doubles [3728]. Doubt [2329]. Doubts [251]. Down [351, 2038, 1212, 2779, 1033, 3746, 3760, 2047]. Dozing [2352]. Dr [341, 542]. Dragnet [800, 1761]. Dragon [2007]. Dragonflies [2356]. Drain [2581, 2734]. Drawing [2786]. Drawn [1302]. Dream [506, 2201]. Dreams [1741, 205, 3304]. Dried [2075]. Drillers [196]. Drilling [258, 2897, 3316]. Drink [729, 3235]. Drinking [1229, 1763, 1603]. Drive [209, 2503]. Drivel [2759]. Driven [1974, 3510, 1662]. Driverless [2345]. Drivers [922]. Driveway [1309]. Driving [3778, 2827, 3081]. Drone [1309]. Drones [63, 3534, 2512, 860, 1708]. Drops [2412]. Drought [3099, 849]. Drug [2651, 2774, 1182, 2494, 1274, 2087, 2556, 1796]. Drug-Resistant [2494]. Druggists [3733]. Drugs [1124, 2151, 3734, 2365, 3660, 190]. Drunker [1191]. Dry [1003]. Drying [1363]. Dueling [1935]. Dumb [2754]. Dummies [1697, 2166]. Dummy [2688]. Dung [1252, 2007, 1404]. Dust [864, 2248, 2331, 1478, 27, 264, 1803]. Dusty [2352]. Duty [3765]. Dwarf [1863, 1848]. Dwindling [936]. Dying [2410, 298, 3419, 1597, 1737]. Dynamics [807, 1583, 1829].

E-Cigarettes [2905, 1929]. E-mail [862, 2324, 3140]. E-mails [3213]. E-Readers [332]. E.T. [951, 3742]. Each [937, 2977]. Eagle [3648]. Ear [1653, 2686, 1552]. Eared [1334]. Early [2391, 108]. Ears [103]. Earth [187, 1186, 2625, 2508, 2726, 3068, 448, 2877, 3185, 1680, 98, 2811, 2617, 2572, 2258, 168, 609, 2480, 1775, 1821, 20, 2122, 3084, 1865]. Earthquake [160]. Earthquakes [258, 880, 2897, 3754]. Easier [1605]. East [55]. Easy [3096, 38, 1771, 62]. Eat [591, 995, 298, 1615, 2985, 1478, 2547, 1881]. Eaters [954]. Eating [3425, 378, 3704, 3336, 1032]. Eats [743]. Ebola [2896, 2335, 2244, 2291]. Eccentrics [2006]. Echo [3194, 3234]. Eclipse [3349]. Eclipses [3350]. Ecology [462, 1431, 3438, 1042, 340, 424, 3696, 2797, 1867, 2767, 3634]. Economic [794, 2237, 3492]. Economics [1981, 2958]. Economy [1208, 3089, 2172, 390, 3789, 3491]. Ecosystems [1200, 30]. Eden [2797]. Edge [2089, 1510, 1156]. Edible [352]. Edit [2722, 1652]. Editing [2736]. Editor [1674, 1712, 1831, 1756, 1793, 25, 197, 170, 51, 1, 271, 242, 345, 106, 77, 132, 309, 723, 684, 392, 797, 648, 613, 761, 576, 502, 539, 466, 429, 1121, 1084, 1161, 1243, 911, 978, 1205, 945, 839, 873, 1047, 1013, 1546, 1278, 1433, 1321, 1392, 1632, 1357, 1511, 1593, 1473, 1946, 1983, 1869, 2181, 1908, 2068, 2148, 2031, 2108, 2299, 2339, 2267, 2505, 2471, 2608, 2224, 2576, 2407, 2374, 2541, 2439, 2648, 3086, 2932, 2709, 2986, 2771, 2739, 2872, 2802]. Editor [2901, 3035, 2679, 3325, 3180, 3295, 3237, 3209, 3443, 3416, 3266, 3353, 3152, 3125, 3389, 3555, 3642, 3527, 3499, 3584, 3730, 3471, 3669, 3611, 3697, 3763, 2851, 3064, 2963, 2832, 3015]. Editorial [2505]. Education [1107, 838, 138, 2065, 2814, 2926, 3693, 2343, 3348, 1577, 1117]. Effect [1802, 92, 2601]. Effects [3137, 2888]. Efficient [719]. Effort [1314, 2234]. Eggs [1771]. Eggshell [2814]. Einstein [2524, 2523, 2529, 2425, 2538, 2830, 2505, 2531, 2533, 2535]. Elections [3050, 3186]. Electric [1058, 2718, 230, 829, 314, 2562, 3197, 3279]. Electron [1025]. Electronic [2635, 221]. Electronics [2332]. Elegant [2678]. Element [1053]. Elemental [2691]. Elements [2200, 2127, 3797]. Elephant [3331]. Elephants [3193]. Elixir [1328]. Elusive [303, 3258]. Embrace [1589]. Embryo [778]. Embryos [2442]. Emergent [797]. Emerging [3606, 3464, 1389, 1471]. Emission [3332]. Empire [1664]. Emptiest [2924, 3072]. Enceladus [3007]. End [1228, 2008, 2052, 1796, 3356, 3587, 242, 1680, 3085, 3374, 2237, 67, 1387, 621, 3357, 3758, 261, 269, 2919, 3504]. End-of-Days [67]. Endangered [1872]. Ending [2411]. Ends [2572]. Enemy [905, 531]. Energy [567, 678, 1202, 2178, 2676, 940, 1664, 189, 3667, 423, 709, 3059, 1428, 193, 1672, 795, 72, 161, 606, 71, 307, 2733, 2467, 1789, 872, 1772]. Engine [847]. Engineer [1230]. Engineered [1623, 1313]. Engineering [2177, 1317, 1079, 3466, 1943, 643, 1354, 1581, 2307, 339]. Engineers [915]. Engines [712]. Enigma [513, 2763, 3533]. Enjoy [2262]. Enlightenment [1242]. Enough [2989, 3240, 429, 2985, 113]. Entertainment [528]. Entropy [498]. Environment [904, 1533, 864, 1238, 1755, 196, 3580, 533, 238, 2897, 166, 344, 2146, 267, 2647, 2956, 2801, 22, 48, 2296, 76, 1200, 426]. Environmental [100, 3792]. Epidemic [3558, 2936, 580, 2014, 3607, 2423, 1999]. Epidemiology [1978, 2925, 3608]. Epilepsy [3223, 1928]. Epoch [1288]. Equals [2670]. Equation [2133]. Equations [2678]. Erasing [1007]. Eric [2490]. Error [3517]. Errors [116, 290]. Escape [1723, 1293]. Essay [2530]. Eternal [3096, 261]. Ethic [2310]. Ethics [247, 1274, 3522]. Ethnic [745]. Ethology [2435]. Eureka [2782, 1912]. Europa [2478, 3096, 2378]. Even [2683]. Events [151, 3004]. Ever [2433, 3599]. Everests [3593]. Every [1309, 1506, 805]. Everybody [3353]. Everyone [3375, 2141]. Everything [1381, 2127, 2009, 1618, 151, 303, 384, 3156]. Everywhere [1270, 2518]. Evidence [2568, 2156, 819]. Evident [2934]. Evil [1623, 2905, 3373, 601, 1268]. Evolution [2126, 2427, 2823, 3149, 3604, 676, 3028, 2179, 194, 3550, 1889, 1470, 1352, 73, 45, 2769, 644, 420, 2495, 2373, 3597, 2468, 1076, 3173, 1654, 1385, 299, 3261, 3061, 3233, 1006, 3324, 3493, 1282, 3664, 1705, 970, 2096, 1900, 2292, 2732, 3498, 1582, 2929, 3122, 2128].

Evolutionary [1070, 1114, 3043]. Evolved [3718]. Evolves [1823]. Evolving [2107, 3032, 299]. Exam [895]. Exceptions [3780]. Excerpt [101]. Exchange [155]. Excluded [3780]. Exclusive [301, 101]. Execution [1911]. Exercise [1580, 3173]. Exist [675]. Exoplanet [3593]. Expected [1238]. Expensive [1374]. Experience [2008, 108]. Experiment [2953, 2503, 76, 1159]. Experimental [790, 789]. Experimentation [1574]. Experiments [2532, 790]. Expert [138, 91]. Experts [258, 1107, 37]. Explicator [1645]. Explode [1043]. Exploded [1513]. Exploiting [1444]. Exploration [390, 2263, 832, 2264, 1904, 1430]. Exports [1716]. Exposed [2137, 566]. Exposure [1166]. Extinction [1312, 83, 1817]. Extra [2024, 204, 153]. Extrasensory [564]. Extraterrestrial [154]. Extreme [259, 1857, 1819, 2625, 2373]. Extremely [1402]. Eye [1759, 2993, 205, 644, 2485, 1010, 3648, 178, 2632, 3359, 1237]. Eye-Candy [178]. Eye-Controlled [2632]. Eyeglasses [2816]. Eyes [125, 3560, 2978, 2493, 1375, 569].

F***ing [596]. Fables [2516]. Fabrication [2483]. Fabrics [1091]. Face [1952, 2977, 3373, 3329]. Facebook [3711, 2449]. Faced [83]. Faces [707, 2684, 1753]. Facilities [3645]. Fact [1054, 1099, 1555]. Fact-Checking [1555]. Factiness [3545]. Factor [1120, 1243, 313, 1495, 2104, 3025]. Factories [422]. Factors [2230]. Factory [1477, 3123]. Facts [1187, 3241, 3142, 3715, 1282, 3483]. Fail [512]. Fair [469]. Faired [2622]. Fairies [1220]. Fairness [686]. Fairy [3696]. Faith [1062, 214]. Faithful [2966]. Fake [169, 2586, 3168, 3514]. Fall [19]. Fallacy [2676, 3780]. Fallout [2298]. False [678, 580, 1906, 72]. Familiar [271]. Family [219, 1545, 2860, 1072, 2132, 135, 1486, 2098, 3017]. Famine [1584]. Fan [2591]. Fantasies [370]. Fantastic [1093]. Fantasy [2336]. Far [3152, 938]. Faraway [1787]. Farm [1621]. Farmers [1882]. Fascinations [261]. Fast [571, 2810, 956]. Faster [1238]. Fastest [1991, 1923]. Fasting [1298]. Fat [2573, 436, 2050, 1619, 248]. Fate [1313, 2431, 136, 2145]. Fatigue [628]. Fats [1835]. Fatter [1191, 2923]. Fatuous [370]. Faults [2810]. Faulty [97]. Faux [1481]. Favor [2950]. Fear [432, 2653, 2867, 1930, 2828]. Fears [1348]. Feast [1593, 1615]. Feathered [1296]. Federal [1726]. Feed [791, 131, 1652]. Feel [3467, 129]. Feeling [802, 476]. Fees [1033]. Feets [2629]. Felicity [2950]. Feline [2950]. Female [3377]. Fence [2377, 226]. Fence-sitters [2377]. Fends [57]. Fermenters [2306]. Ferocious [2449]. Fertilizers [133]. Fetal [909]. Fetch [2930]. Fever [507, 1938, 1856, 3608]. Few [2498, 2946]. Fewer [2016]. Fi [999]. Fiber [2306]. Fibs [2511]. Fickle [1220]. Fiction [990]. Fiddling [1634]. Field [2480]. FIFA [1722]. Fight [2083, 537, 1746, 1595, 3726, 1256, 91]. Fighter [1046, 808, 3115]. Fighting [3430]. Fights [3222]. Filming [239]. Films [3484]. Filter [377]. Filth [1861]. Filthy [193]. Final [1852, 245, 3659]. Finance [2661]. Financial [490, 3491]. Find [2877, 3331]. Finder [1054, 1099]. Finding [3793, 2252, 2648, 315]. Finish [320]. Fire [1231, 2288, 3337, 1019, 2434, 2363]. Fired [2210]. Firefighters [3273]. Fires [471]. Fireworks [2824]. First [2, 764, 1969, 231, 3411, 2452, 1899, 3520, 3492, 787, 1962, 277, 970, 1620]. Fish [3537, 1431, 1133, 3650, 702, 3164]. Fishing [3438, 2370, 1173]. Fishy [3540]. Fit [2277, 319]. Fittest [2128]. Five [1665, 866, 1588]. Fix [3755, 457, 617, 1836, 2788, 2231]. Fix-It [1836]. Fixing [3767, 39, 48]. Flags [1748]. Flame [1838]. Flames [603]. Flare [2366]. Flash [776]. Flashes [3579]. Flatland [971]. Fleet [1079]. Flexible [1749]. Flier [62]. Flies [920]. Flight [3128, 887, 3222, 286, 2180, 1388]. Flimflam [490]. Flip [2122]. Float [404]. Flock [447, 3738]. Flood [2938, 1104]. Flooding [544, 3769]. Floor [1069, 3227]. Flop [634]. Flow [361, 328, 1028]. Flower [2483]. Flowers [2349, 2383, 441]. Flu [422, 1718, 1777, 3503, 445, 3570]. Fluid [807, 1583, 1829]. Flunk [1968]. Flush [2430]. Fly [1884, 1067, 1847]. Flying [2419, 3426, 120, 3175]. Foam [873, 352]. Fog [3362, 2831]. Fold [253]. Folds [6]. Foliage [2044]. Follow [1728, 1404]. Following [1674]. Food [537, 1618, 1595, 1445, 1256, 2027, 1616, 1617, 1615, 2520, 2985, 1003, 964, 3286, 1614, 131, 2262, 767, 1621, 617, 2360, 459, 3455, 117, 872]. Foods [1623]. Foolish [1314]. Foolishness [214]. Fools [897]. Foot [1079]. Footprint [1916]. Foragers [3505]. Forbidden [49]. Forces [443]. Forecast [3247, 2187]. Forecaster [2748]. Forensic [1009, 2437, 3393, 2526]. Forensics [837, 757, 1390, 264]. Forest [1363, 1837, 2780, 1427, 2420]. Forestalling [164]. Forests [2504]. Forever [1734, 262, 525, 822]. Forged [2007]. Forget [2912, 1514, 2048, 917]. Forging [2329]. Forgotten [2253]. Format [3430]. Former [230]. Formula [794, 2492]. Fortress [2259]. Forum [1760, 1017, 2806, 1361, 2475, 2545, 687, 765, 2185, 1088, 1209, 1596, 1051, 3213, 1873, 2935, 3270, 3186, 3588, 3559, 1988, 1325, 2072, 949, 3241, 3673, 2237, 982, 3447, 849, 3734, 2509, 349, 433, 1797, 2990, 3419, 2905, 275, 727, 3328, 580, 3646, 1514, 652, 2378, 877, 1477, 1912, 2411, 506, 915, 1716, 313, 1549, 2271, 801, 3357, 2612, 3156, 3129, 2683, 2713, 2876, 3701, 2112, 3503, 2580, 2652, 1678, 2311, 2775, 3531, 3474, 1165, 3393, 3615, 1950, 1635, 617, 2343, 3300, 2035, 3766, 1282, 3090, 396, 1247, 3039, 1437, 2443, 1835, 543, 1125, 1396, 470, 2152, 2743]. Forward [132, 2932]. Fossil [1941, 3396, 1428, 2217, 3316, 3198]. Fossils [1481, 135]. Foundations [2967, 2965]. Four [2201, 1275, 776]. Fracking [764, 1557, 1691, 1603, 795]. Freaks [1036]. Free [1834, 2071, 3129, 2265, 3476, 1982, 1109]. Freediving [1971]. Freedom [2612, 808, 470]. Freeze [825]. French [630]. Frenzy [3537]. Fresh [981, 680]. Friction [965]. Frictionless [2661]. Friend [3678, 462]. Friends [1545, 3717, 2849, 475, 2511, 450]. Friendships [2044]. Fries [630]. Fright [3286]. Frogs [1773]. Frontier [1858, 1852, 1560, 1225]. Frontiers [2266]. Frothy [1523]. Frozen [3163, 3620, 1555]. Fruit [981]. Fruitless [3505]. Fruits [3163, 2027]. Frustration [2330]. Fry [2388]. Fudge [313]. Fuel [2210, 1618, 2172, 829]. Fueled [32]. Fuels [1428]. Fuggedaboutit [1636]. Fukushima [2298, 2717]. Fulfill [1051]. Full [1949, 1190]. Fundamental [2535]. Funding [1878, 3756]. Fungi [1751, 3704, 3620, 117]. Furry [3716]. Fuse [3199]. Fusion [1966, 2636, 1041, 3059, 3187, 72, 161]. Futile [3103]. Future [1208, 1415, 1699, 1423, 2054, 3067, 2890, 3299, 3606, 3490, 1464, 2337, 648, 978, 1278, 1433, 3209, 151, 2364, 938, 1901, 1520, 3187, 1462, 122, 3221, 491, 2283, 897, 1853, 2284, 3484, 1314, 1621, 907, 1005, 3231, 831, 3075, 2764]. Fuzzy [2730, 1880].

Gadget [488]. Gadgets [1814, 562, 634]. Gains [539]. Galactic [2350]. Galápagos [3466, 2191, 2767]. Galaxies [2009, 1600, 568, 1863]. Gambled [2425]. Gambling [1696]. Game [1045, 2630, 1341, 216, 1897, 373, 2424, 889]. Games [2894, 2872, 1827, 2527, 2147]. Gaming [35]. Gap [2140, 3603, 293, 3711, 3386, 1018, 3790, 2035]. Gaps [1461, 1542, 2921]. Garbage [562]. Garden [40]. Gardeners [1260]. Gardner [203]. Gas [2564, 381, 74]. Gasping [3188]. Gate [2466]. Gather [940]. Gatherer [2309]. Gathering [1535]. Gauging [2888]. Gave [2273, 3492]. Gays [2843]. Gazing [2977]. Gender [2140, 3375, 3376, 3711, 1461, 3378, 2652, 3090]. Gene [1586, 2682, 1209, 2229, 3775, 2573, 2207, 1864, 2305, 793, 2946]. Generation [2487, 2862]. Genes [3367, 2675, 3656]. Genesis [1936]. Genetic [800, 2489, 2906, 1204, 3466, 1905, 2362, 1867, 3398, 3795, 3796]. Genetics [3466, 114, 3291, 3265]. Genie [2229, 2207]. Genius [1624, 2428, 2491, 2757, 2539, 797, 1536, 2683, 3798, 1338, 2062, 2491]. Geniuses [3538]. Genome [461]. Genomics [3207, 2646]. Gentle [2368]. Geography [1905]. Geologic [1799]. Geologist [1104]. Geology [73, 2738, 3437, 792, 755, 2955]. Geometric [384]. Geometry [1736, 3787]. Geothermal [258, 1557]. Germ [2766, 1978, 830]. Germany [1193]. Germination [2197, 2389]. Germs [2197, 815, 320]. Gerrymandering [3787]. Get [1087, 862, 3418, 1044, 3062, 2808, 3193, 2693, 184, 2354, 3164]. Gets [1923, 3784, 3042]. Getting [289, 519, 1152, 323, 2220, 1383]. Ghost [2971]. Ghostly [1421]. Ghosts [2350, 3242]. Giant [2025, 3402, 792, 3397]. Giants [33, 177, 2769]. Gig.U [741]. Ginkgo [1496]. Giraffe [1627]. Girl [1539, 1923, 3383]. Girls [2844, 2391]. Give [809]. Glacial [3566]. Glad [3054]. Glamorous [329]. Glass [2314, 1629]. Glasses [775, 1520]. Glimpse [2000, 3492]. Glitches [1677]. Global [1436, 2020, 1238, 342, 170, 1161, 358, 2829, 642, 1536, 612, 2382, 2881, 714, 22, 3262, 48]. Globaloney [672]. Glocks [2474]. Gloom [3054]. Glory [611, 869]. Glow [2621]. Glue [2401, 1425]. Gluten [1816, 1812, 3476]. Gluten-Free [3476]. GM [1595]. GMOs [2637, 2027, 2360]. Go [3502, 558, 2444, 3180, 3165, 890, 1744, 2113, 2378, 1785, 2057, 929, 1028, 2943]. Goals [146]. Goat [2891]. God [1062, 317, 1549, 1737, 2221]. Gods [1542, 3406, 2026]. Goes [1569, 3174, 2038, 2312, 3684]. Going [2105, 3030, 1092, 1517, 1494, 484, 2672, 3074, 328, 1611, 1890, 662]. Gold [1715, 2430, 3316, 3479, 2087]. Goldbach [988]. Goldilocks [867]. Gone [2590, 3438, 1329, 1339, 3781]. Goo [1533]. Good [304, 1411, 266, 1107, 408, 1868, 2864, 3630, 1453, 2999, 1976, 2113, 1554, 1638, 2985, 81, 1110, 1055, 1773, 1298, 1753, 3514]. Good-bye [2999]. Google [1495, 1752]. Goose [2270]. Got [735, 2533]. Government [1390]. GPS [1941, 289]. Grab [3220]. Grade [2215]. Gramps [1814]. Grams [2930]. Grandma [3311]. Grandmother [1350]. Grandparents [676, 3028]. Graphene [1992, 2551, 3621]. Graphic [1075, 3783, 3463, 1823, 298, 336, 371, 638, 674, 529, 786, 492, 750, 566, 603, 709, 820, 419, 456, 1191, 901, 1111, 1037, 1420, 1502, 1545, 1591, 1631, 1307, 1704, 2205, 1898, 2012, 2058, 1938, 1975, 2172, 2630, 2331, 2564, 2399, 2257, 2291, 2600, 2463, 2731, 2701, 2923, 2793, 3056, 2893, 3005, 3317, 3374, 3172, 3408, 3344, 3287, 3435, 3144, 3228, 3630, 3751, 3489, 3661, 3717, 3576, 3603, 1149, 1349, 2494, 1739, 2822, 2362, 2431, 2761, 852, 936, 1003, 1270, 969, 1235, 1461, 1384, 1672, 1861, 1785, 2132, 3107, 2952]. Graphic [3258, 3192, 3547, 3519, 3689, 2528, 2671, 2095]. Grass [1599, 1029]. Gravitational [1673]. Gravity [232, 971, 188, 17, 124, 370, 43, 335, 150, 297, 68, 260, 216, 95, 455, 673, 785, 602, 491, 528, 565, 707, 637, 819, 749, 418, 1074, 1148, 968, 1036, 1190, 1110, 1002, 853, 1234, 1269, 935, 900, 1460, 1543, 1670, 1589, 1383, 1738, 1501, 1348, 1419, 1703, 1629, 1306, 1860, 1974, 1822, 2094, 2173, 2204, 1897, 1784, 1937, 2057, 2011, 2131, 2256, 2563, 2493, 2430, 2629, 2290, 2361, 2462, 2527, 2398, 2330, 2599, 2821, 2792, 2892, 3004, 3106, 2670, 2730, 2950, 3055, 2833, 2700, 2760, 2922, 3143, 3257, 3316, 3227, 3286, 3373, 3171, 3190, 3433, 3462, 3342]. Gravity [3407, 3487, 3518, 3629, 3716, 3602, 3575, 3781, 3546, 3660, 3688, 3750]. Great [1592, 2036, 3185, 336, 903, 1645, 642, 2581, 2968, 1944, 2629, 3349, 1891, 2954, 2483, 1159]. Greater [611]. Greatest [2128, 2489, 3795, 3796]. Green [717, 406, 3251]. Greener [710, 717, 546, 3511]. Greenhouse [1533, 2564, 43]. Greenland [3294]. Grid [3667]. Gridlock [433]. Grief [598, 1944]. Grit [493]. Grizzly [2245]. Ground [2658, 2582]. Grounds [260]. Groundwater [3439, 2618]. Group [2470]. Grow [1237, 669]. Growing [573, 1029, 1846]. Growth [198, 102, 2237, 101]. Guard [2123]. Guardian [683, 3526]. Guards [1747]. Guide [2236, 2362]. Guitar [1177]. Gulf [1353]. Gumming [886]. Gun [1459, 2817]. Gunk [2478]. Gunland [3412]. Guns [3257, 2552]. Gunshot [1545]. Gut [1957, 1746, 114, 2781, 2338, 2308, 2354, 2303, 1967]. Guts [435, 3189].

Habit [1946, 101]. Habits [1976]. Hack [2705, 581]. Hacker [2236, 240, 1100]. Hacker-Ready [1100]. Hackers [750, 2509, 1650, 275, 2593]. Hacking [2447, 645, 1708]. Hadrosaur [2273]. Haemophilia [2227, 2266]. Hail [2818]. Hair [1846]. Haitian [1773]. Half [1187]. Half-Life [1187]. Hallucination [1418]. Hallucinogens [387]. Halting [771, 342]. Hammer [2100]. Hand [883, 1419, 1219, 3246]. Hands [906, 3164]. Happened [3224]. Happy [2471, 2863, 2847, 1025]. Harassment [3530, 3758]. Hard [876, 2503]. Hardest [3721]. Hardware [3599, 240]. Harm [2151, 1591, 1366]. Harrisment [150]. Harvest [3351].

Hatch [1293]. Hated [2425]. Havoc [2279]. Hawking [2425, 317, 353]. Hazardous [1677]. Head [2532]. Headache [180]. Headed [1158]. Heading [3743]. Heads [2739, 2752, 3720]. Heal [616]. Healers [126]. Healing [848]. Health [1016, 1677, 1436, 3356, 1580, 641, 1066, 745, 974, 2335, 1285, 3253, 342, 500, 1612, 1732, 737, 1852, 2674, 703, 3581, 781, 487, 1102, 1182, 848, 932, 1341, 1453, 457, 3447, 2925, 2960, 293, 366, 866, 414, 524, 387, 905, 598, 2918, 2753, 1262, 1696, 3052, 2663, 3312, 2573, 610, 1494, 1816, 331, 2703, 1536, 2165, 3188, 2723, 1929, 2520, 2249, 2391, 3008, 3077, 3139, 3338, 2465, 715, 815, 1140, 896, 964, 998, 1226, 1376, 1651, 1565, 1412, 1777, 451, 2050, 3412, 3607, 2283, 2623, 2423, 2486, 3000, 3223, 2245, 2438, 3179, 3633, 428, 3289, 3790, 2308, 2831]. Health [3102, 3235, 3369, 2354, 2592, 2888, 561, 1005, 2817, 2063, 1032, 1118, 1298, 1529, 1890, 2124, 3668, 822, 2004, 2694, 1078, 1967, 2198, 2323, 3429, 3458, 3598, 3684, 3712, 3746, 3656, 3483, 3542, 3514, 3571, 3625, 3777, 633, 2087, 2556, 2946, 2455, 2788, 3282, 3167, 669, 2896, 3403]. Health-Wealth [3790]. Healthier [2299, 1295]. Healthy [2306, 99, 1126, 3452]. Hear [2481, 296, 3308]. Hearing [312, 2498, 2497, 3713, 34, 3684]. Heart [2605, 366, 1424, 2368, 3334, 3147, 3428]. Heart-Brain [366]. Heartbeat [3363]. Heartbeats [3679]. Heartburn [180, 2781]. Hearts [1009]. Heat [2215, 1420, 476, 2638, 129]. Heated [1443]. Heats [2447]. Heavens [1821]. Heavy [58]. Heavyweights [3551]. Heck [3407]. Heel [315, 3120]. Helix [1303]. Hell [1053, 2466]. Help [3680, 3759, 1988, 475, 689, 1661, 2866, 1076, 1773, 2397, 2004]. Helper [3311, 1889]. Helpers [2418]. Helping [1592, 3089]. Helpless [1223]. Hepatitis [1902]. Her [1253, 1176, 692, 280]. Here [1381, 1892, 2003, 1211, 2607, 2719, 2255, 860]. Hereditary [206]. Heretic [344]. Hiccupping [1106]. Hidden [1550, 2464, 866, 2497, 3137, 1672, 964, 835, 569, 1890, 105]. Hiding [2777]. Higgs [758, 1129, 1201]. High [1498, 2904, 3250, 58, 1578, 1099, 1402, 2082, 1003, 779, 3280, 3175, 3455, 3164]. High-Flying [3175]. High-Jump [1099]. High-Speed [1578]. Higher [86]. Hills [2752]. Him [1881]. Himalayan [2476]. Historical [2327]. Histories [2132]. History [1907, 1068, 2667, 2875, 1792, 465, 3414, 2536, 574, 2531, 1044, 1617, 611, 975, 3084, 2955]. Hit [944, 1762, 2058, 2365]. Hitchhikers [582]. Hits [2160, 2195, 2246, 2278, 2320, 2351, 2381, 2415, 2451, 2484, 2517, 2553, 2585, 2619, 2660, 2689, 2720, 2749, 2784, 2813, 2883, 2913, 2944, 2997, 3046, 3098, 3200, 3138, 3161, 3622, 3653, 3217, 3249, 3277, 3309, 3683, 3708, 3737, 3333, 3366, 3399, 3427, 3776, 3454, 3481, 3513, 3541, 3565, 3592]. Hitting [2077]. HIV [941, 1078, 554]. HIV’s [3120]. Hobbit [2162]. Hockey [944]. Hold [981, 2934, 1771, 1993, 775]. Holding [977]. Holds [151]. Hole [2059, 2425, 2293, 1478, 1874, 2537]. Holes [60, 3319, 867, 3057, 3520, 1112]. Holey [1642]. Hollers [1923]. Hollywood [1141, 1678]. Hollywood-Style [1678]. Home [2698, 2444, 3014, 2155, 1052, 3245, 1448, 1492, 3048, 217]. Homecoming [1876]. Homicide [3766]. Hominin [3724]. Homo [1070, 3358]. Honest [3703]. Honey [3751]. Honing [312]. Honors [613]. Hooey [2759]. Hope [1570, 3096, 3266, 5, 1906]. Hopeful [953]. Horizon [1308, 1160]. Horses [2575]. Horton [2491, 3798]. Hospital [2253]. Hospitals [3129]. Hot [1059, 3563, 883, 3307, 1183, 1217]. Hotspots [2058, 2793]. Houdini [454]. House [667]. Household [2319]. Hovers [2158]. Howdy [1134]. Hubble [3192]. Human [1070, 2008, 2427, 2953, 3058, 646, 676, 152, 1243, 3697, 709, 3293, 1470, 3022, 1352, 122, 3202, 1040, 374, 1385, 2338, 2961, 970, 2096, 1900, 2162, 2292, 2732, 3018, 3069]. Humanely [3159]. Humanitatis [2429]. Humanity [1735, 2068, 2925, 236, 3021, 1151, 1315, 3033, 3015]. Humans [1654, 83]. Hummer [2792]. Humvee [231]. Hundred [380]. Hunt [2901, 1132]. Hunter [2708, 2309]. Hunting [2217]. Huntington [2929]. Hurdle [2904]. Hurricanes [771]. Hurt [992]. Hurtfuls [1111]. Hybrids [1470, 3022, 191]. Hydrogen [3251]. Hydrothermal [78]. Hype [1570]. Hypothetical [2090]. Hysteria [3176].

Ice [187, 3446, 1224, 1799, 837, 2999, 3337, 3534, 3271, 985, 2146, 2080, 1115, 3195, 1295]. Ice-Diving [3534]. Iceman [322]. Ick [1589]. Icon [2191]. Icy [3162]. Idea [2425, 3646, 1477, 1656]. Ideas [372, 821, 1740, 2631, 3108, 3086, 2206, 2139, 1271, 2207, 2219, 147]. Identities [3493]. Ideology [80]. If [2090, 2100, 3491]. Ignites [1673]. III [2202]. I’ll [3685]. Illiteracy [3038]. Illness [914, 500, 3661, 653]. Illnesses [1746]. Illusion [165, 2222, 2980, 2965]. Illusions [2967, 2979, 2981, 2980, 2985, 2983]. Illustrated [1303]. I’m [1400, 329, 3539, 543]. Imagery [2406]. Imagine [3170]. Imagined [1417]. Imaging [532, 496, 239]. Immortality [967]. Immune [2361]. Immunity [1285, 3738, 2222]. Immunology [2232, 868, 2222, 3206]. Impact [3493]. Impending [2476, 2122]. Imperfect [1853]. Imperiled [94]. Implant [1127]. Implicates [2807]. Import [1129]. Importance [382]. Important [1950]. Impossible [3730, 1465, 2981, 2180]. Impracticality [894]. Impressions [2452]. Inches [2345]. Inclusion [2133]. Income [2010, 3766]. Inconvenient [2146]. Increase [118]. increased [3491]. Incredible [352]. Incredibly [1402]. Index [281]. India [1573]. Indicate [2940]. Indirect [1826]. Indisputable [1282]. Industrious [149]. Ineffective [3538]. Inequality [3788, 3792, 2958, 2010, 3766]. Infant [2707, 1750]. Infection [3656]. Infections [1746]. Infectious [422, 1751, 1043]. Infested [2359]. Infinite [1157]. Inflammation [3483]. Inflated [1913]. Inflation [530]. Infographics [2594]. Information [2053, 388, 942, 1866, 1827, 1510, 21, 308, 517, 1662]. Infrastructure [3270]. Infrequencies [2130]. Ingestible [3112]. Ingredient [168, 1928]. Inheritance [2061]. Injectable [2635]. Injection [3645]. Injury [3743]. Inner [575, 508, 2771, 2617, 126, 1236, 461, 1386, 174]. Innermost [494]. Innocent [2807]. Innovate [3056]. Innovation [372, 1659, 1671, 2631, 3108, 2608, 3237, 712, 1012, 1271, 2642, 3, 27, 53, 79, 108, 134, 172, 199, 244, 273, 311, 347, 394, 431, 468, 504, 541, 578, 615, 650, 708, 725, 763, 799, 841, 875, 913, 947, 980, 1015, 1049, 1086, 1123, 1163, 1207, 1245, 1280, 1323, 1359, 1394, 1435, 1475, 1544, 1590, 1630, 1676, 1714, 1758, 1795, 1833, 1871, 1910, 1948, 1985, 2033, 2070, 2110, 2150, 2183, 2226, 2269, 2301, 2341, 2376, 2409, 2441, 2473, 2507, 2543, 2578, 2610, 2650, 2681, 2711]. Innovation [2741, 2773, 2804, 2874, 2903, 2951, 2988, 3037, 3088, 3127, 3154, 3182, 3211, 3239, 3268, 3298, 3343, 3355, 3391, 3434, 3445, 3488, 3501, 3529, 3557, 3586, 3613, 3644, 3671, 3699, 3732, 3782]. Innovations [2302, 130, 2266]. Innovator [1296]. Insanity [1702]. Insect [2290]. Insects [3228, 769]. Insecure [2281]. Inspire [1744]. Instant [737, 777, 1806, 104]. Instincts [401]. Insulin [981]. Intelligence [1083, 3753, 534, 639, 3548, 154]. Interior [1204]. International [619]. Internet [629, 469, 1970, 1745, 274, 1988, 589, 1510, 2809, 2194, 80, 3001]. Internets [1325]. Interrupted [3568]. Interstellar [2286]. Intolerance [880]. Intricate [3318]. Introducing [3047]. Introduction [1538, 2529, 2327, 2964, 261, 1614]. Invaders [3330]. Invasion [63, 1442, 2600]. Invasive [1615, 2495]. Inventing [2642]. Invention [1655, 1632]. Inventors [2006]. Investigate [1360]. Investigates [1104]. Invisible [2127, 2395, 816, 239]. Invited [2108]. Inviting [2141]. iOS [3283]. IPCC [116]. iPhone [1650]. iPhones [483]. iResearch [2456]. Irrational [1189]. Irrationals [1071]. Island [2288, 3551, 186, 2336]. Isle [2942]. Isn’t [2528]. Issue [3375, 3353, 1614]. Itch [2346, 2771, 622, 2796]. Itself [1179, 2064]. I’ve [735]. iVote [2693].

Jaguars [3680]. Jam [289]. Jane [389]. Jar [2539]. Jargon [2040, 2157, 2115, 1990, 2086, 2192]. Jefferson [465]. Jellies [2690]. Jet [1259, 2220]. Jewel [572, 3198]. Jittery [1704]. Jobs [1208]. John [2491, 2190, 3798]. Johnny [2858]. Join [596]. Joint [2234]. Joke [2131, 782]. Journey [1473, 2068, 1204, 2885, 3412, 494, 1904, 3015]. Joy [1185, 2555]. Joyful [1113]. Judgment [1167]. Juice [1387]. Jump [1099, 390]. Jump-Starting [390]. Junctions [143]. June [589]. Jungle [2624, 389]. Junk [2328, 382]. Jupiter [2885]. Jupiters [2496]. Jurassic [3198]. Just [1970, 500, 282, 1806, 2295, 1177, 113, 1089, 2713, 3747, 882, 3380, 3164, 2849]. Justice [3150, 1936, 2397].

Kansas [848]. Kant [3601]. Kapitza [1169]. KEA [3681]. Keep [1195, 469, 2306, 1152, 3129]. Keeping [3209, 2204, 1628]. Kelvin [1639]. Kepler [1727, 1919]. Key [3063, 1574]. Kicks [2599]. Kid [411]. Kiddo [2421]. Kidney [113]. Kids [1124, 2853, 2615, 2845, 2501]. Kill [2637, 2270, 2465, 284, 931, 3668, 2152, 2743]. Killed [2644, 1141]. Killer [115, 538, 3489, 315, 1718, 2165, 801, 528, 377, 3656]. Killers [3206, 830]. Killing [1471]. Kilogram [2941]. Kind [2755, 265, 1997, 2061, 970]. Kindergarten [3435]. Kinetic [1202, 2980]. King [2202, 66, 1705]. Kings [1380]. Kitchen [2588]. Kite [1202]. Kits [2812]. Kitty [1471]. Klinenberg [2490]. Knee [2323]. Knocking [1845]. Knockoff [1748]. Knocks [876]. Know [2054, 2089, 1970, 2040, 2157, 1618, 2677, 519, 949, 3641, 2115, 921, 361, 1990, 3060, 2086, 2192, 375]. Know-It-All [375]. Knowledge [1793, 3266, 3270, 3597, 3615, 3060, 2949]. Known [2023]. Knows [2421]. Kool [14]. Kool-Aid [14]. Korea [3783].

Lab [608, 2811, 3145, 2878, 1748]. Lab-Built [3145]. Laboratory [256]. Labs [202, 926]. Lactose [880]. Lake [340]. Lakes [1115]. Land [1991, 2545, 2399, 516, 1642]. Landmarks [2043]. Landscaping [3044]. Language [2800, 460, 3063, 1199]. Large [3364, 11, 2261]. Larry [66]. Laser [326]. Last [1779, 2410, 2695, 974, 265, 642, 1225, 1527, 2684, 320, 1060, 1566, 3370, 1201, 322, 3724]. Lasting [173, 2079]. Later [212]. Laughed [1082]. Laughing [1106]. Laughs [1444]. Launch [2369]. Laws [2990]. Layers [3084, 2955]. Laying [268]. Laziness [3043]. Lead [1686, 1745, 1612, 290, 246, 2888]. Leaden [1841]. Leader [1728]. Leaders [2570, 3773]. Leaf [307, 2383]. Leaking [189]. Leanings [2841]. Leap [576]. Leaps [3450]. Learn [2826, 3158, 3763, 2912]. Learner [2499, 677, 2855]. Learning [1592, 2856, 1569, 1575, 3293, 3770, 2065, 1572, 3784, 2022]. Lecturing [2066]. Ledge [2744]. Left [1249, 267, 2729, 2940, 1347]. Legacy [2534, 1612, 2411, 3095, 964]. Legal [2039]. Legionnaire [2462]. Lego [1743, 1733]. Lego-like [1743]. Legos [2212]. Legs [1024]. Lending [1552]. Lens [144, 1790, 130, 2386]. Less [713, 1222, 700, 2568, 450, 872]. Lesser [2905]. Lesson [241]. Lessons [2728]. Let [2875, 3128, 558, 1408, 1431, 3443, 2279, 2147]. Letdown [2316]. Lethal [342, 79, 3571]. Letter [3155]. Letters [107, 133, 2, 198, 78, 171, 26, 243, 52, 272, 310, 346, 393, 430, 467, 503, 540, 577, 614, 649, 685, 724, 762, 798, 840, 874, 912, 946, 979, 1014, 1048, 1085, 1122, 1162, 1206, 1244, 1279, 1322, 1358, 1393, 1434, 1474, 1512, 1547, 1594, 1633, 1675, 1713, 1757, 1794, 1832, 1870, 1909, 1947, 1984, 2032, 2069, 2109, 2149, 2182, 2225, 2268, 2300, 2340, 2375, 2408, 2440, 2472, 2506, 2542, 2577, 2609, 2649, 2680, 2710, 2740, 2772, 2803, 2873, 2902, 2933, 2987, 3036, 3087, 3126, 3153, 3181, 3210, 3238, 3267, 3296]. Letters [3297, 3326, 3354, 3390, 3417, 3444, 3472, 3500, 3528, 3556, 3585, 3612, 3643, 3670, 3698, 3731, 3764]. Lettuce [1460]. Level [891, 3164]. Liar [3703]. Liberty [3328]. Lice [3282, 1030]. Lichen [3292]. Lichens [3219]. Lie [1104, 1030]. Lies [3753, 3040, 3305, 899]. Life [1187, 1667, 1343, 1972, 1857, 1895, 2489, 2254, 3068, 679, 2123, 156, 201, 2588, 1787, 3631, 2835, 1114, 386, 1045, 3795, 3796, 2318, 2935, 1, 930, 2179, 638, 1420, 3085, 2731, 3630, 3603, 2466, 722, 3381, 2780, 238, 263, 1790, 18, 2708, 1236, 50, 2768, 3080, 1503, 230, 1977, 2373, 95, 3688, 461, 1727, 3510, 2223, 373, 2492, 3517, 1386, 3639, 3676, 105, 3345, 2575, 389, 1318, 1582, 2489]. Lifelong [3770]. Lifesaver [74]. Lift [3219]. Lifting [943, 2402]. Liftoff [335]. Lifts [210]. Light [1667, 2665, 1745, 252, 338, 2433, 3201, 803, 3052, 621, 2319, 7]. Lightning [2793]. Lights [684, 1528, 2880, 645]. Like [167, 1059, 1029, 3287, 3561, 435, 823, 963, 1217, 1743]. Limb [1701]. Limbs [1717]. Limit [3214]. Limits [1150, 2442, 51, 639, 977, 3233, 3640]. Line [2959, 3079]. Lingering [2510]. Linguistics [2800, 3063, 3178]. Link [3272, 2305]. Linking [1285]. Links [820, 3261]. Lions [2744]. Lips [2837]. Liquid [829, 1829]. Liquids [890]. List [395, 1281, 3283]. Listening [3194]. Little [1914, 2636, 12, 847, 475, 1889, 915, 2969, 746, 1118]. Live [388, 1149, 533, 1154, 262, 331, 113, 1119, 3192, 1654, 172, 66]. Liver [1647]. Lives [3645, 2299, 2575]. Living [1894, 2545, 1788, 1278, 98, 1591, 2960, 3078, 1809, 263, 257, 3364, 3082, 604, 2261]. Lizard [3453]. Local [22, 718]. Lock [2167, 55]. Lockdown [1774]. Lockstep [1572]. Logging [3680]. Logic [1305]. Logic-Tight [1305]. Logician [1230]. Logjam [65]. Loko [776]. LOL [1400, 82]. Loneliness [3497]. Lonely [1848]. Long [388, 1966, 887, 1294, 1709, 113, 1917, 3192, 1654, 511, 1789]. Long-Term [1294]. Longer [331, 2079]. Longer-Lasting [2079]. Longest [2326, 1149]. Longevity [3078, 870]. Look [143, 3006, 3552, 436]. Looking [2055, 2932, 2709, 18, 1718, 3635]. Looks [509, 963]. Looming [3123]. Loops [1004]. Loose [1917]. Lose [1062]. Losing [2582]. Loss [2497, 3289]. Lost [1145, 2202, 2052, 2200, 2254, 33, 3290, 3797, 568, 939]. Lot [397, 3322]. Love [1666, 545, 820, 456, 521, 2692, 148, 1615, 2979, 2808, 2775, 1881, 3486, 2240, 1432, 2994]. Loved [2425, 3410]. Loving [1518]. Low [2215, 779, 1430]. Low-Cost [1430]. Low-Grade [2215]. Lump [11]. Lunar [246, 3044]. Lure [2880]. Lust [2835, 1401, 2834, 2836]. Lying [1896]. Lyme [2510]. Lysol [1401].

Ma [143]. Machine [1625, 1791, 1575, 3293, 158, 2768, 3080, 2403, 367, 1530, 831]. Machines [1083, 2034, 2826, 1744, 2632, 1157, 1466, 192, 2639]. Machismo [10]. Mad [1700]. Madagascar [1646]. Maddening [2796]. Made [2002, 879, 3157]. Magellanic [1397]. Magic [1760, 1580, 1805, 1454, 3339]. Magical [2626]. Magnetic [2480, 2122]. mail [862, 2324, 3140]. mails [3213]. Mains [2783]. Make [3185, 1743, 1644, 1462, 436, 3436, 2750, 3757, 3048, 1619, 1660]. Makeover [2941]. Makes [2101, 3023, 2864, 937, 1915, 1193]. Making [1433, 890, 1766, 3293, 2470, 88, 3061]. Malaria [3357]. Males [995]. Malformed [2253]. Malignancy [2013]. Malleable [2148]. Mammals [2823, 3751, 1251]. Mammograms [561]. Mammoth [2396, 1555]. Mammoths [472]. Man [437, 2597, 2254, 2559, 879, 3157, 2193, 633]. Man-Made [879, 3157]. Manhattan [1346]. Mania [3657]. Manipulated [1746]. Manipulating [3640]. Manipulation [181]. Manuals [3224]. Manufacturing [1462, 1678]. Manure [2240]. Many [2887, 2935, 2224, 3056, 1657]. Mapped [1069]. Maps [3244]. Marble [3555]. March [660, 2504, 3240, 2013, 1751]. Marijuana [2283, 3458]. Marine [3290, 1420, 3577, 1904]. Mark [2365]. Market [212]. Marketable [2148]. Mars [832, 960, 1469, 788, 2419, 81, 1977]. Marshmallow [2091]. Martian [893]. Martin [203]. Marvels [2031]. Mash [31]. Mash-Up [31]. Mask [1961]. Masks [3707]. Mass [1989, 1422, 3176, 15, 93, 67, 214, 122, 186, 41, 148, 257, 2738, 599]. Masses [1975, 715]. Massively [1571]. Mastering [2091]. Mastermind [1304]. Masters [574]. Material [590]. Math [171, 1666, 2758, 740, 3172, 682, 1002, 2730, 2470]. Mathematical [2280, 3798]. mathematician [3793, 2252]. Mathematics [1907, 570, 1945, 3752, 3787, 1830, 2678, 3583, 2147, 2469, 3695]. Matter [2595, 2991, 3691, 1262, 2412, 3637, 2311, 2242, 3348]. Matters [1932, 3375, 345, 2530, 93, 3701]. Matthew [2489]. May [2547, 2745, 2383]. Mayhem [10]. McKibben [102]. Me [1820, 694, 1524, 2481, 474, 436, 216, 296, 2066, 543, 3164]. Meaning [3292, 2492, 3517]. Means [2249, 2964, 1785, 3553]. Measles [554]. Measure [1346, 2686, 3774]. Measurement [1143, 1673]. Measuring [1422, 3440]. Meat [608, 3252, 1262]. Mechanical [2177]. Mechanics [2617]. Medal [1075]. media [3570]. Medical [1144, 1368, 2712, 796, 1852, 1317, 2936, 1376, 2283, 343, 3102, 1274, 2694, 2332]. Medicinal [2148]. Medicine [1038, 571, 1423, 1819, 3606, 641, 1980, 1009, 237, 978, 3525, 1560, 47, 2364, 3414, 903, 1901, 387, 302, 495, 126, 24, 770, 2753, 1826, 1709, 941, 2603, 262, 2675, 1864, 2176, 3595, 531, 2434, 190, 1355, 2580, 3179, 2266, 680, 3323, 103, 1005, 1532, 164, 2402, 2646, 2218, 3578, 2332, 3231, 3496, 74, 3075, 2764, 3553, 1940, 2259, 756]. Medieval [1398]. Meditator [2174]. Mediterranean [1484]. Meds [2583, 652, 2781]. Meet [3463, 773]. Meeting [1960]. Meets [2485]. Mega [849]. Megafloods [1319]. Meltdown [972, 159, 1080, 3580, 396]. Melting [2285]. Melts [2156]. Memories [1007]. Memorize [1566]. Memory [2129, 2295, 2344, 859, 3318]. Men [3252, 1205, 3376, 2256, 3380, 3777, 176]. Menace [573, 173, 1695]. Mending [3334]. Menstrual [3276]. Mental [914, 2253, 2815, 1285, 3581, 3661, 653, 2308]. Menu [2893]. Mercaptan [565]. mercury [3164, 1774]. Merge [1181]. Merger [3657]. Messenger [2322]. Messengers [3475, 3605]. Messy [3289, 3483]. Meta [722]. Metal [3183, 2899]. Metallic [1744]. Metals [1760]. Meteor [1132]. Meteorites [2745, 1351]. Meteorology [1979, 1010]. Meter [3436, 1652]. Methuselah [3003]. Metrology [3176]. Mexico [1660]. mHealth [1051]. Mice [1644, 2878]. Mickey [222]. Micro [3112, 1454]. Microbe [827, 2305, 3649]. Microbes [2306, 1746, 984, 885, 1362, 2302, 3175]. Microbial [1021, 1564]. Microbiology [609, 3785]. Microbiome [2304, 2338, 2309, 2996, 2307, 2303]. Microchips [23]. Micromachines [1813]. Microphone [137]. MicroRNA [806]. Microscopic [1989, 177]. Microwave [2633]. Microwaves [803]. Midterm [2112]. Might [810, 83]. Migrate [3228]. Migrations [1075]. Mild [2173]. Mile [1631]. Mileage [3646, 46]. Miles [2811]. Milestone [1966]. Military [3473]. Milk [1253, 2891, 1611]. Milky [751, 1723, 2025, 3729, 3091, 3619, 2217]. Millennium [146]. Millie [1073]. Million [1158, 2125]. Million-Dollar [2125]. Millions [1927]. Mind [312, 2053, 2491, 3066, 1786, 1241, 481, 1277, 911, 3209, 1827, 1098, 1707, 1113, 2969, 3386, 835, 464, 1155, 854, 1263, 1530, 2834, 2174, 666, 1073, 1628, 3749, 306, 950, 2566, 889, 2854, 1504, 3798]. Mind-Reading [1530, 666]. Mindful [2148]. Minds [1592, 2596, 1321, 2576, 1657, 2292, 3026]. Mine [2293]. Minerals [73, 806]. Miners [827]. Miniature [3183]. Miniatures [833]. Mining [3570, 3610]. Minuscule [1560]. Minutes [3184]. Miracle [1799, 343]. Mirror [1186, 2937, 1472]. Mislead [3102]. Misled [3519]. Misperception [2967, 2965]. Miss [3625]. Missing [3272, 1041, 1288, 3387, 168, 3742, 3591, 3261]. Mission [3201, 2930, 1384]. Mite [1991]. Mix [2871, 3197]. Mixology [1813]. Mobs [776]. Model [2911, 2382, 1251]. Models [2235]. Modern [2006, 2490, 3794, 1398, 2490]. Modification [3243]. Molasses [1583]. Mold [2519]. Molecules [2937, 1472, 3118]. Molten [2617]. Mom [3216, 884]. Moment [309]. Money [542, 3490, 419, 2273, 21, 2614]. Mongolian [3526]. Monikers [2242]. Monitor [822]. Monitoring [1376]. Monkey [2583]. Monster [3694, 3520, 2777]. Moon [3396, 973, 70, 2876, 3508, 3705, 3003, 2343]. Moonlet [1996]. Moons [1787, 1050]. Moose [465]. Moral [228, 2289]. Morality [3733, 3725]. Morbid [2886]. Morse [3158]. Mortal [2783]. Mosquito [3424]. Mosquitoes [3045]. Most [1068, 342, 1560, 2495]. Moth [3359]. Mother [638, 3532, 2547]. Motherhood [1018]. Mothering [3602]. Moths [2880]. Motion [2640, 1321, 1469, 1155]. Mount [967]. Mountains [3507, 2476]. Mourn [1534]. Mouse [3303, 31, 222]. Mouth [664, 1809]. Move [3244]. Movement [2260]. Movements [205]. Moves [2009, 3507]. Movies [992, 1454, 2789]. Moving [911, 3402]. Mr. [3778]. Much [2935, 429, 3641, 267, 517, 1001]. Muddying [3055]. Muir [2190]. Mules [1021]. Multimillion [1617]. Multimillion-Year [1617]. Multiverse [675, 18, 3288]. Murder [2669, 2569]. Murderer [1989, 2738]. Muscles [3767]. Museums [3652]. Mushroom [2736, 2746, 1805, 2193]. Music [312, 2129]. Must [3393]. Mutant [57]. Mutants [1031]. Mutating [1686]. My [862, 3678, 717, 1463, 2399, 581, 1400, 2708, 565, 1478, 3331, 854, 1227, 3685, 884, 511]. Myself [1820]. Mysterians [3659]. Mysteries [3066, 753, 3263]. Mysterious [1070, 3749]. Mystery [1397, 2464, 958, 619, 1564, 592, 1484, 3649, 3291, 3633, 3705, 2437, 2732, 3018, 3069]. Myth [1911, 2918, 1355, 601, 2010]. Myths [3376, 1588, 3122].

n [2664]. Nada [1973]. Naked [2018, 45, 3024, 35]. Name [2858, 1543]. Name-O [1543]. Nano [1466, 830]. Nano-Size [830]. Nanobacteria [19]. Nanobots [2369]. Nanoparticles [2216]. Nanotechnology [2017]. Naps [1333]. NASA [1050, 2508, 3128, 3222, 210, 81, 2321]. NASA’s [2747]. Nation [1191]. National [2043]. Nations [1415]. Natives [1621]. Natural [538, 3762, 574, 3610]. Natural-Born [538]. Nature [1070, 1933, 1011, 2227, 932, 3100, 3562, 1903, 2321, 379, 3640]. Navigating [3368]. Neandertal [213, 174, 2292, 3026]. Neandertals [167]. Near [3201, 1754]. Near-Light-Speed [3201]. Necessary [223, 2816, 102]. Ned [2620]. Need [2089, 3700, 1578, 12, 828, 2863, 3626, 13]. Needs [3389, 2509, 2683]. Negating [29]. Neglect [914]. Neglected [24]. Neighbor [1134]. Neighbors [3463]. Neither [2719]. Neo [3347]. Nerve [2840]. Nest [521]. Nestlé [2338]. Net [392, 1891]. Nether [1933]. Network [1038, 1180, 2038]. Networked [2435, 2106, 3031]. Networks [104, 3491]. Neurobiology [2796]. Neurology [3229]. Neurons [1023, 2028]. Neuroprosthetics [464]. Neuroscience [1007, 2327, 972, 3263, 2539, 1277, 338, 3006, 3666, 639, 3582, 3552, 937, 532, 2432, 2509, 2707, 97, 1505, 2028, 3690, 3145, 605, 3436, 1113, 1661, 1905, 2497, 3177, 2297, 1040, 2965, 2672, 3074, 3761, 2265, 2734, 835, 3062, 3203, 1942, 3784, 569, 1350, 71, 2174, 1320, 1237, 1199, 3318, 2571, 677, 493, 1567, 2062, 3632, 1468, 2931, 1862, 425, 128]. Neuroscientist [2762, 3020]. Neuroscientists [2943]. Neurosurgeons [1935]. Neutrino [1725, 1422, 125, 3409]. Neutrinos [810, 2572, 148]. Neutron [2763]. Never [1830, 2038, 3140, 2198]. Newborn [1066]. Newest [550, 701, 773]. News [200, 4, 173, 3260, 365, 400, 449, 473, 560, 632, 661, 697, 736, 245, 54, 80, 28, 3168, 3010, 109, 135]. Newtonians [2425]. Next [270, 1851, 1158, 2248, 3125, 23, 1287, 1462, 1656, 1451, 3437, 3657]. Nexus [1632, 711]. Nice [1724, 320]. Night [3367, 205, 3579, 2406, 3504]. Nightmare [765]. Nights [1275]. Nine [2942, 2795]. Niño [3013]. Nitrogen [48]. No [542, 2055, 2151, 2904, 143, 2816, 1725, 846, 1744, 887, 2027, 3312, 1572, 186, 2975, 1329, 655, 1249, 2760, 1366, 2713, 1720, 593, 2808, 3170, 3574, 1847]. No-Fly [1847]. No-Show [1725]. Noah [1104]. Nobel [1164, 607, 1535, 410, 1081, 707]. Noble [707]. Noir [2262]. Nondonors [113]. None [3171]. Noodles [235]. Noodling [235]. nor [2719]. North [3783, 2584]. Northwest [3258, 2809]. Nose [2620, 695]. Notable [1108, 1146, 1188, 1072, 1232]. Note [3085]. Nothing [397, 1001]. Novel [1745, 3076, 3172]. Nuclear [579, 2742, 3327, 550, 631, 2316, 1664, 3242, 2298, 606, 22, 1310, 1973, 155, 2743]. Nukes [3199]. Number [2011, 374]. Numbers [1071, 952, 1483, 1491, 1516, 1526, 1551, 1556, 1604, 1637, 1640, 1648, 1683, 1688, 1694, 1719, 1729, 1541, 1767, 1769, 1800, 1808, 1811, 1840, 1842, 1879, 1885, 1887, 1918, 1924, 1955, 1959, 1965, 1995, 2041, 2045, 2078, 2252, 2534, 2413, 570, 59, 988, 290, 1897, 2527, 3695, 2250, 3429, 1437, 3793, 510, 518]. Numerology [4]. Nurse [1488]. Nurture [3562, 2321]. Nurtures [932]. Nutrition [2338].

O [1543, 565]. Obama [1691, 3095]. Obamacare [1765]. Obesity [457, 866, 2050]. Object [241, 137]. Oblivion [68]. Observatory [300]. Obviously [1629]. OCD [2254]. Ocean [1069, 1971, 1779, 3694, 1367, 1126, 238]. Oceanography [3124, 2404, 1944]. Oceans [2986, 3499, 2333]. Octopus [1663, 3198]. Octopuses [1744]. Odd [1457]. Odds [268]. Ode [2322]. Odyssey [2252, 3793]. Off [3743, 2442, 3199, 591, 1730, 57, 2304, 2675, 210, 3193, 931, 1597, 1660, 2715, 1940, 2545, 3014, 1400]. Off-Limits [2442]. Officials [2904]. Offshore [142].

Oh [335, 2865]. Oil [26, 1246, 2714, 254, 3709]. Old [513, 1023, 3734, 2519, 2131, 398, 2198]. Oldest [1894, 3358, 1865]. Olives [2576, 2604]. Olympians [768]. Olympic [1111, 2904]. Olympics [2914]. Omelet [2670]. On/Off [2675]. Once [1699, 137, 1912, 1503, 1532]. One [536, 2055, 2201, 1711, 380, 1886, 1712, 2893, 3100, 3216, 113, 3492, 549, 702, 1619, 2103]. One-Stop [1711]. One-Way [1886]. One-World [2893]. Ones [1868]. Online [1267, 997]. Only [2560, 2458, 3741, 3491]. Op [2980]. Open [664, 452, 13, 1628]. Opening [2811]. Opens [1953]. Operation [3323]. Operations [3514]. Opinion [3327]. Opioid [3408]. Opioids [3419, 3746, 3625]. Opportunity [1573]. Opting [1502, 934]. Oracle [1945]. Oral [2232, 895, 3777]. Orange [1387, 2831]. Orangutan [3595]. Orbital [390]. Order [3492]. Orderly [535]. Organ [3411, 569, 2832]. Organelle [3093]. Organic [118]. Organisms [2209]. Organized [2053]. Organs [3058, 1426, 2079, 478, 1374]. Organs-on-a-Chip [478]. Orgasmic [2834]. Origami [300, 253]. Origin [2076, 2274]. Origins [218, 2252, 676, 1385, 3029, 20, 3725, 3259, 1582, 3793]. Orphaned [2750]. Orthopaedics [2234]. Other [1246, 2427, 2031, 3376, 2971, 2977, 767, 1197, 2124, 3549]. Our [505, 1699, 1786, 1746, 2418, 2337, 2648, 2771, 2739, 3584, 3186, 3728, 3588, 3447, 1273, 3665, 3187, 2990, 205, 1615, 2895, 3071, 2700, 1913, 3484, 569, 3609, 213, 3493, 3720, 135, 174, 970, 3347, 3164]. Ourselves [1971, 3584, 3352, 3609, 899, 3496]. Out-of-This-World [2347]. Outbreak [658]. Outbreaks [3299, 3570]. Outer [2771, 2795]. Outlawing [3461]. Outlook [2013, 2227]. Outrageous [2461]. Outshining [2467]. Outsmarting [507, 732]. Over-the-Counter [2583]. Overhaul [3093]. Overkill [1888]. Overlap [3661]. Overload [2053]. Overoptimism [934]. Overreaction [2592]. Overweight [119]. Owl [3367, 360]. Own [2860, 1237, 2996]. Owns [948, 3614]. Oxygen [1781].

P.U. [3101]. Pacemakers [2239]. Pacific [2627, 160]. Pages [2516]. Pain [3467, 2975, 2623, 2219, 2218]. Painful [1007, 3633]. Painkiller [2936, 1515]. Paleo [2360, 3664]. Paleoanthropology [3347]. Paleontology [1941, 2400, 2644, 759, 1239, 976, 939, 385, 501, 1432, 3208]. Pals [1834]. Pampering [1607]. Panacea [3312]. Pancreas [3042]. Pandemic [2727, 163]. Paper [1707, 3117]. Parable [1651]. Paradise [1302, 3322]. Paradox [2753, 3173, 2929]. Parasite [342, 224]. Parasites [1540]. Parent [439, 2864, 3303]. Parental [2243]. Parenting [2851]. Parents [2750]. Park [631]. Parking [3322]. Parkinson [3274]. Parsing [656]. Part [3718, 3727, 3724]. Parthenote [778]. participation [3491]. Particle [1228, 2401, 2763, 2621, 1421, 1236, 1549, 1939, 619, 1856, 3409, 1958, 3440]. Particles [447, 2241, 1679, 1064]. Particular [2136, 2555]. Partner [2587]. Parts [2201]. Party [534, 2124]. Pass [1488, 3257]. Passage [3258, 2809]. Passengers [2747]. Passing [2112]. Password [704, 3140]. Passwords [2919]. Past [948, 1825, 121, 3285]. Pastime [692]. Patent [1658, 1562, 955, 989, 1096, 1178, 1218, 1255, 1290, 1330, 1364, 1399, 1450, 1489, 1522, 2012, 182, 1687, 288, 327, 354, 409, 523, 555, 625, 690, 730, 856]. Path [870, 2715]. Pathogen [1291]. Pathogens [1608, 221]. Pathology [243]. Paths [761]. Patient [868]. Patients [2410]. Pay [2304, 2664, 1660]. Pays [1730]. Peace [1064, 2171]. Peacekeepers [2303]. Peak [3717]. Peaked [3559]. Pediatric [3369]. Pee [2037, 3190]. Peeling [885]. Peels [1601]. Peer [2861]. Peering [3416]. Pencil [246]. Pending [182]. Penguin [2490]. Penguins [3519]. Pennies [3117]. People [1415, 1540, 1013, 1898, 3419, 546, 3575, 1927, 2646, 1396]. Percent [1197, 2095, 3034]. Perception [49, 51, 2024, 2598]. Perchance [3775, 1567]. Perfect [1818, 1639, 3577]. Perfectly [2984, 3542]. Performers [2616]. Peril [1576]. Perils [1066, 2407, 670]. Periodic [640, 1507, 2200, 3797]. Peripheral [2168]. Perish [3502]. Permafrost [3121]. Permission [3332]. Perpetual [2171]. Persistence [2344]. Persona [2460, 3624]. Personal [2418, 2015, 1571]. Personally [2138]. Perspective [2967, 2231, 2022]. Perspectives [12, 90, 38, 211, 254, 64, 183, 145, 119]. Pertussis [1651]. Pervspicacity [1784, 2833]. Pest [2319]. Pests [1518]. PET [118]. Petrovich [1169]. Pets [3781, 1405]. Phage [1120]. Phantom [1563]. Pharoahs [683]. Phi [1105]. Philosopher [2525]. Philosophy [790, 682]. Phone [2477]. Phones [1733]. Phoning [1738]. Phosphorus [340]. Photicular [2558]. Photographic [859]. Photonics [1011]. Photos [1652]. Photosynthesis [843]. Physician [616]. Physicist [2525, 3741]. Physics [341, 1301, 2728, 2806, 1004, 1824, 1791, 647, 165, 971, 61, 2910, 1722, 978, 1908, 2401, 1099, 758, 3662, 2763, 303, 1421, 1568, 1236, 384, 1939, 1537, 3057, 1081, 1523, 853, 3082, 3409, 902, 572, 1242, 2607, 2540, 869, 1917, 2385, 2363, 2535, 498, 1201, 1316, 1272, 1444, 604, 3440, 1504]. Physiology [977]. Pick [2083, 1770]. Picky [954]. Picturing [2251]. Pieces [1041]. Pig [505, 2722]. Pigs [2713]. Pill [1210, 3008]. Pills [237, 434, 2713, 1748]. Pince [1524]. Pinning [351]. Pinot [2262]. Pinwheel [1170]. Piracy [992]. Pitfalls [698]. Pizzaiolo [3281]. Place [2648, 1743, 2895, 3071, 3106, 2924, 3072, 497]. Places [137, 466]. Plain [1101]. Plan [24, 210, 3227, 2817]. Planet [1699, 2236, 3584, 99, 791, 1077, 2503, 1313, 70, 494, 3388, 2296, 1607, 2679, 2795, 2702, 3070]. Planetary [2333, 2673, 2930, 70, 1469, 1977, 3415, 3007, 788, 3465, 3549]. Planets [1563, 2935, 1270, 1283, 1848, 217]. Planning [606]. Plant [631, 1008, 1387, 378]. Plant-Eating [378]. Plants [550, 1503, 350, 129, 3590, 3308]. Plastic [241, 1601, 3335, 1093, 655, 2080, 3509, 2284, 3336]. Plastic-Eating [3336]. Plastics [118, 2213]. Plate [3448]. Plates [767]. Play [2491, 728, 3527, 1724, 2863, 2277, 3798]. Player [2500]. Players [3785]. Pleasantly [1579]. Please [1596, 1524, 360]. Pliable [1744]. Plight [3124, 850, 3012]. Plowshares [1540]. Plug [191]. Plug-in [191]. Plugged [2779]. Pluto [2454, 1406, 2175, 3465]. Plutonium [2587]. Pocket [1749]. Poet [590]. Pogue [3778]. Point [2289]. Points [2136]. Poised [2454, 30].

Poison [1493, 1770, 2687, 3425, 2313]. Poison-Eating [3425]. Poisoned [1987, 28]. Poisoning [964]. Polar [2558, 159, 1080, 3271]. Pole [2584]. Police [3733]. Policy [39, 120, 3522]. Polio [974]. Political [395, 2666]. Politically [1189]. Politics [348, 3499, 488]. Poll [301]. Pollen [813]. Pollination [2725]. Pollster [2154]. Pollute [1951]. Polluters [3144]. Pollution [2611]. Polyester [813]. Polyglot [1658]. Pond [844]. Pontoon [1921]. Ponytails [1690]. Ponz [1981]. Pool [2037]. Poor [131, 633]. Pop [3174]. Popcorn [2385]. Poplars [3425]. Pops [1098]. Popular [450]. Population [2957]. Pornception [564]. Porpoise [3510]. Porpoise-Driven [3510]. Portrait [109]. Positive [698, 2861]. Positives [855]. Post [3473]. Postmodernism [3372]. Postpartum [3745]. Postponed [302]. Pot [2687, 2990, 1494, 1928]. Potato [3624]. Potions [1748]. Poultry [3433]. Pound [2832]. Poverty [2568, 2723, 91, 3113]. Poverty-Predicting [3113]. Power [579, 752, 1458, 2725, 3652, 2894, 142, 1664, 2882, 275, 2707, 2194, 283, 1998, 3615, 2274, 1657, 2743]. Powered [1354, 2779]. Powerhouse [2287]. Powers [2102, 2846, 3027]. Poxviruses [1389]. Practically [406]. Pre [871]. Pre-Crime [871]. Precipitation [1027]. Precision [2753]. Predator [1900]. Predicament [3369]. Predict [3484, 1314, 831, 2280, 3570]. Predicting [3113]. Prediction [3121]. Preferential [2111]. Prefers [1707]. Pregnancies [2544]. Prehistoric [2414, 1286]. Prejudice [877, 1509]. Prenatal [2682]. Prepare [3672]. Prescription [2816]. Presence [92]. Preserve [1872]. Presidential [1205, 150]. Press [2490, 3]. Pressure [2799, 2861]. pretty [3164]. Prevent [1750, 2817]. Prevented [704]. Preventing [3581, 2829, 1451]. Prevention [2019]. Price [1498, 2611, 3479]. Prices [2651, 553]. Prickly [1515]. Primate [2106, 3031]. Primates [1668, 3159]. Prime [988]. Primeval [1027]. Primitive [1351]. Primordial [1170]. Princess [319]. Print [1465, 3275]. Prion [1922]. Prism [3047]. Privacy [1710]. Prize [1164]. Prizes [986]. Probably [1497, 353, 450]. Probes [2635]. Probing [925, 3274]. Probiotics [3312]. Problem [505, 1246, 1824, 3721, 82, 534, 3752, 1908, 3575, 1240, 965, 1733, 3051, 3515, 2652, 2921, 48, 3090]. Problems [3433]. Process [39, 13]. Processed [1617]. Produce [1518]. Produce-Loving [1518]. Product [3599]. Prof [2927]. Professional [1012]. Programmable [2177]. Progress [746, 53]. Project [118, 1040]. Prolific [50]. Promiscuous [3376]. Promise [678, 1051, 41, 1576, 2017]. Promises [2407]. Prompt [116]. Proof [1934, 3730, 1812, 1418]. Proofing [2914]. Propeller [2153]. Prospecting [3394]. Prostate [903]. Protect [1016, 2498]. Protection [1726, 670]. Protector [3097]. Proteins [535, 2113]. Proton [1824, 282]. Protozoans [1690]. Provide [2946]. Prune [1567]. Pseudoscience [706, 2526]. Psychology [1786, 2371, 1868, 460, 3467, 1203, 3352, 195, 3663, 3609, 3497, 14, 493, 2762, 1752, 1509]. Psychopaths [1203]. Pterosaur [2042]. Puberty [2391]. Public [3502, 974, 2335, 500, 2674, 905, 3077, 3412, 3607, 2438, 428, 3668, 2896]. Publicly [2358]. Puke [360]. Pulse [2239]. Pumped [1742]. Pun [3092].

Pun-Divided [3092]. Punk [324]. Pupae [1921]. Pursuit [2139, 2235]. Pursuits [1081]. Push [1091]. Pushed [3005]. Pushing [2266]. Put [3392, 2037, 1749, 1775]. Puts [137]. Putting [179, 744, 497]. Puzzle [2128, 3409, 2733, 2296]. Puzzlers [1960]. Puzzles [1070]. Puzzling [1673]. Pyramid [2601]. Pyramids [2576]. Python [30].

Q [3151]. Qaeda [629]. Quack [2921]. Quakes [3157]. Quantum [1855, 1004, 1791, 971, 252, 137, 3450, 1160, 576, 2705, 3662, 1568, 3057, 891, 2798, 1242, 2540, 3288, 1316, 1272, 604, 1504, 3111]. Quarks [1205, 1236]. Quest [1, 1793, 498]. Question [1361, 294, 452, 1310]. Questions [2090, 1158, 3635, 3492]. Quick [2160, 2195, 2246, 2278, 2320, 2351, 2381, 2415, 2451, 2484, 2517, 2553, 2585, 2619, 2660, 2689, 2720, 2749, 2784, 2813, 2883, 2913, 2944, 2997, 3046, 3098, 3200, 3138, 3161, 3622, 3653, 3217, 3249, 3277, 3309, 3683, 3708, 3737, 3333, 3366, 3399, 3427, 3776, 3454, 3481, 3513, 3541, 3565, 3592, 617, 3164]. Quiet [838, 204, 1118]. Quirky [3621]. Quit [2218]. Quite [1339]. Quotable [279, 316, 321, 356, 359, 363, 399, 407, 413, 438, 444, 477, 479, 486, 514, 520, 559, 585, 597, 623, 663, 742].

Rabid [1068]. Rabies [1565, 3761]. Race [2489, 2579, 1680, 2821, 1447, 1286, 3795, 3796]. Racetrack [621]. Racial [2580]. Racists [3341]. Rack [1861]. Radiance [1893]. Radiation [3278]. Radical [567, 423, 3439]. Radioactive [428]. Radius [1824]. Rail [127]. Railroad [643]. Rain [1979, 1363, 1172, 1837, 2247, 2348, 499, 2420]. Rain-Forest [2420]. Rainbow [551, 3361, 233]. Raining [851]. Raise [3300]. Raises [3492]. Raising [2853, 1581]. Ran [1258]. Random [2540]. Range [3014, 1052]. Ranges [2247]. Rapid [1850, 205]. Rare [3229, 1726, 1773]. Rat [1082]. Rates [524]. ray [1791, 463]. Rays [1116, 2164]. Razor [1745]. Razor-thin [1745]. Reach [2534]. Reaction [2361]. Reactions [2640, 3264, 1967]. Reactors [3344]. Read [862, 2945, 2822]. Readers [332]. Reading [1077, 3116, 1530, 666, 1681]. Ready [1360, 591, 1127, 3221, 994, 1100]. Real [169, 995, 1487, 21, 3691, 41, 1568, 1764, 2586, 297, 3784, 1342, 3042, 784, 3664]. Realities [3642]. Reality [2531]. Really [2538, 223, 675, 102, 3521, 992, 3626]. Reanimate [1476, 1596]. Reason [3413]. Rebirth [1867]. Reboot [3531]. Recall [777, 1649]. Recently [2092]. Reckon [443]. Reckoning [3212]. Reclaiming [2052]. Recommendations [187, 215, 259, 416, 453, 563, 635, 747, 783, 817, 1184, 1264, 1228, 898, 933, 966, 1000, 1034, 1068, 1104, 1142, 1414, 1455, 1496, 1627, 1540, 1668, 1378, 1734, 1300, 1698, 1343, 1585, 1931, 1854, 2052, 2089, 1779, 2167, 2200, 1817, 1969, 2006, 2126, 1892, 2325, 2393, 2425, 2488, 2594, 2625, 2251, 2356, 2285, 2558, 2522, 2457, 2665, 2695, 2725, 2755, 2790, 2819, 2890, 2920, 2948, 3002, 3104, 3141, 3169, 3191, 3225, 3255, 3284, 3314, 3340, 3371, 3405, 3431, 3460, 3485, 3516, 3544, 3573, 3600]. Recommendations [3627, 3658, 3686, 3714, 3748, 3779, 3053, 295, 333, 368, 489, 526, 600, 671, 705, 861]. Recommended [16, 123, 149, 94, 42, 416, 453, 563, 635, 747, 783, 817, 898, 933, 966, 1000, 1034, 2790, 2819, 2890, 2920, 2948, 3002, 3104, 3141, 3169, 3191, 3225, 3255, 3284, 3314, 3340, 3371, 3405, 3431, 3460, 3485, 3516, 3544, 3573, 3600, 3627, 3658, 3686, 3714, 3748, 3779, 1108, 1146, 1188, 1232, 3053, 69, 295, 333, 368, 489, 526, 600, 671, 705, 861]. Recorder [3535]. Records [2694]. Recovering [1726]. Recovery [950]. Recyclable [2213]. Recycling [118]. Red [1378, 1715, 1077, 2959, 3079, 2976, 1607]. Redacted [3522]. Reduce [3300]. Reflections [1627, 502]. Reforms [120, 212]. Refugees [2735, 3083]. Regaining [103]. Regenerate [3456]. Regions [1933]. Reimagined [3582]. Reinvented [2531]. Reinventing [307]. Related [1446]. Relativity [2522, 2534]. reliant [2870]. Relics [2770]. Relief [3745]. Remains [916]. Remapped [3619]. Remarkable [1069, 3135]. Remedy [2786]. Remember [1456, 3248, 1922]. Remembering [2458, 2479]. Remembrance [1825]. Remote [1263]. Renaissance [3279]. Rendezvous [2046]. Renditions [2981]. Renegade [1971]. Renewable [1672]. Renewables [52]. Renewed [5]. Rent [1183]. Repair [2323]. Repeat [2346]. Replace [3132]. Replacing [1426]. Replanting [1427, 828]. Report [218, 821, 1423, 1592, 1655, 2565, 2825, 3490, 3788, 3464, 98, 1569, 2364, 1462, 3320, 2499, 2926, 3693, 1005, 3348, 3231, 2764]. Reproducible [3757]. Reproduction [3553]. Reprogram [3456]. Reprogrammable [2208]. Reptiles [3576]. Republicans [2582]. Reputation [3421]. Requiem [3346]. Require [3327]. Required [2027, 655]. Rescue [618, 844, 2186, 1213, 3045, 996]. Research [1796, 2980, 692, 3757, 2338, 1274]. Resemblance [2966]. Reshaping [1415]. Reshuffled [3344]. Resilient [2863]. Resistance [1046, 3103]. Resistant [2494]. Resourceful [1964]. Resources [3610]. Respect [819]. Response [1850]. Responsibilities [3232]. Responsible [2776]. Restart [270]. Restless [418]. Restore [3456]. Results [3184, 3164]. Resume [2420]. Resurrection [2906, 3468]. Rethink [3756]. Rethinking [506, 1565, 1353]. Retiarius [2384]. Retreat [3162]. Return [3387, 2153, 998, 2876, 1621]. Revealed [111, 3465]. Revelation [1804]. revelations [16]. Revenge [3282]. Review [2597, 2595, 2491, 2489, 2490, 2627, 2596, 2626, 2488, 2594, 2625, 2668, 2667, 2726, 2758, 2697, 2757, 2698, 2666, 2727, 2728, 2756, 2696, 2665, 2695, 2725, 2755, 116]. Reviews [215, 259, 416, 453, 563, 635, 747, 783, 817, 1184, 1264, 1228, 898, 933, 966, 1000, 1034, 1068, 1104, 1142, 1414, 1455, 1496, 1627, 1540, 1668, 1378, 1734, 1300, 1698, 1343, 1585, 1931, 1854, 2052, 2089, 1779, 2167, 2200, 1817, 1969, 2006, 2126, 1892, 2325, 2393, 2425, 2488, 2594, 2625, 2251, 2356, 2285, 2558, 2522, 2457, 2665, 2695, 2725, 2755, 2790, 2819, 2890, 2920, 2948, 3002, 3104, 3141, 3169, 3191, 3225, 3255, 3284, 3314, 3340, 3371, 3405, 3431, 3460, 3485, 3516, 3544, 3573, 3600, 3627]. Reviews [3658, 3686, 3714, 3748, 3779, 3053, 295, 333, 368, 489, 526, 600, 671, 705, 861]. Revolution [1142, 1901, 302, 1135, 2926, 459, 2323]. Revolutionary [127, 2118]. Reward [2836]. Rewired [1717]. Rewrite [3275]. Rewritten [2096]. rex [2273]. Rhetoric [779]. Rich [500]. Richard [2202]. Riches [755]. Rickety [2427]. Riddance [266]. Ride [1969, 581]. Rigged [3789]. Right [2410, 2755, 2774, 553, 3328, 323, 2782, 1044, 1177, 2390, 3008, 353, 335, 417, 3197, 3712]. Righteousness [1035]. Ring [1508, 1327, 3480]. Rings [2673, 2363]. Ripples [3073]. Rise [2825, 2400, 2374, 104, 2291, 427, 2272, 3107, 1466, 1789, 3571, 1900, 19]. Rising [1195, 2709, 529, 2058, 1349, 2557]. Risk [268, 1384, 1131]. Risk-Averse [1384]. Risks [539, 701, 1349, 987]. Risky [2579, 1307]. River [2561, 1258]. Rivers [3509]. RNA [1901]. Roaches [1213]. Road [2522, 1966, 2716, 637, 2752]. Roads [375]. Roam [667]. Roberts [2491]. Robo [3281, 427, 2818]. Robo-Taxi [2818]. Robobees [1388]. Robocrats [2185]. Robot [304, 2201, 3250, 1463, 3148, 3311, 1663, 237, 209, 3681, 378, 1221]. Robotic [3515]. Robotics [304, 1663, 391, 3202, 2899, 2961, 192]. Robots [2605, 1744, 2039, 3548, 1552, 3112, 2557, 2828]. Rock [55, 2077]. Rocket [1539, 2436, 2747]. Rocketry [2633]. Rockets [210, 3307]. Rocks [1104, 2739, 3563, 3674, 801, 175, 1865]. Rocky [3647]. Roe [3381]. Role [2382]. Rollback [3646, 1765]. Rolled [1749]. Rolled-up [1749]. Romance [2490, 3285, 3794, 2490]. Romania [1780]. Room [2560, 2480, 157]. Room-temperature [157]. Roots [1338, 3039]. Rope [1829]. Ropes [3413]. Rot [3430]. Rough [3506]. Round [3657]. Rover [1378, 81]. Rovers [893]. Royalty [1915]. Rubbery [2314]. Rudolph [654]. Ruled [1503]. Rules [2805, 556, 637, 1002]. Rumbles [3706]. Rumen [1501]. Rummaging [245]. Run [2273]. Running [1493]. Runs [2561, 1234]. Russia [1664]. Rust [2326].

Sabotaged [1436]. Sacred [818]. Safe [3645, 2870, 1929, 729, 3140, 1284]. Safer [1124, 3435, 2552]. Safety [764, 2114]. Saga [2096]. Said [2848]. Sailing [3401]. Salad [1631]. Sale [2694]. Salesman [1026]. Sally [1969]. Salmon [3507, 666]. Salt [2074]. Saltwater [2900]. Salubriousness [818]. Sam [2482]. Sampler [2589]. Sandy [1559]. sans [187]. Santa [360]. Sapiens [3719, 3358]. Sass [2257]. SAT [2514]. Satellite [2406]. Satellites [851, 458, 3111]. Saturn [2673, 3415]. Saturnalia [1703]. Saucers [2419]. Save [3494, 1381, 3645, 2806, 930, 3466, 3320, 412]. Saved [236, 3021, 3388]. Saving [2116, 2770, 2797, 3166, 3547, 2064, 2144, 3329, 3423]. Savvy [710]. Say [778, 1830, 3257, 2713]. Sayonara [1682]. Scale [2212]. Scaled [1212]. Scan [200, 4, 173, 365, 400, 449, 473, 560, 632, 661, 697, 736, 205, 245, 54, 80, 28, 109, 135]. Scanning [1732, 951]. Scans [1529]. Scare [1595]. Scaredy [3395]. Scaredy-Cats [3395]. Scatter [1456]. Scene [2940]. Scent [3617, 754, 646]. Scents [88]. Schedule [2390]. Schedules [3589]. Schizophrenia [3263]. School [3473, 202, 723, 1569, 2082, 1254]. Schooling [638]. Schools [1578]. Schrödinger [2524]. Science [301, 42, 348, 312, 281, 726, 579, 542, 432, 686, 469, 505, 395, 616, 651, 764, 800, 772, 774, 804, 863, 1267, 1075, 1087, 981, 1208, 1050, 914, 1246, 1016, 862, 1124, 1164, 876, 948, 986, 1107, 1324, 1281, 1476, 1677, 1513, 1634, 1360, 1395, 1436, 1715, 1655, 1971, 1759, 2184, 2151, 1796, 1834, 1949, 2071, 1911, 1872, 1987, 2111, 2034, 2626, 2567, 2342, 2270, 2474, 2508, 2236, 2544, 2410, 2611, 2579, 2310, 2377, 2442, 2565, 2682, 3089, 2651, 2712, 2774, 2875, 3038, 2904, 2805, 2742]. Science [2989, 2934, 3050, 3009, 3011, 3212, 3356, 3418, 3446, 3128, 3155, 3240, 3327, 3392, 3269, 3199, 3299, 3375, 3783, 3652, 3700, 3733, 3765, 3587, 3502, 3558, 3672, 3473, 3530, 3645, 3614, 3788, 3755, 679, 908, 2894, 973, 388, 116, 990, 2496, 3463, 2137, 1066, 745, 2119, 3253, 1612, 1732, 1823, 1852, 1196, 703, 1792, 2337, 197, 345, 106, 648, 502, 839, 1473, 2541, 3443, 3266, 3125, 3471, 2295, 781, 3694, 3290, 2024, 465, 2179, 1116, 64, 274, 1194, 1595, 1548, 1730, 3376, 949, 298, 336, 371, 638, 674, 529, 786]. Science [492, 750, 566, 603, 709, 820, 419, 456, 1191, 901, 1111, 1037, 1506, 1420, 1502, 1545, 1591, 1631, 1307, 1704, 2205, 1898, 2012, 2058, 1938, 1975, 2172, 2630, 2331, 2564, 2399, 2257, 2291, 2600, 2463, 2731, 2701, 2923, 2793, 3056, 2893, 3005, 3317, 3374, 3172, 3408, 3344, 3287, 3435, 3144, 3228, 3630, 3751, 3489, 3661, 3717, 3576, 3603, 3524, 3414, 1080, 487, 1102, 1182, 848, 932, 1341, 1453, 3064, 405, 293, 366, 1976, 414, 524, 1149, 1349, 2133, 2494, 168, 598, 2918, 2753, 1262, 1739, 1696, 3052, 2663, 3312, 2333]. Science [1044, 2065, 877, 3663, 2673, 1494, 1816, 427, 331, 2822, 1710, 611, 2930, 612, 2165, 1689, 2362, 50, 682, 3188, 70, 2723, 1929, 2431, 2520, 2249, 2391, 2761, 3139, 3338, 1503, 178, 852, 936, 1003, 1270, 969, 1235, 1461, 1384, 1672, 1861, 1785, 1469, 1825, 1977, 815, 1140, 896, 964, 998, 1226, 1376, 1651, 1565, 1412, 1777, 2373, 2563, 2922, 451, 2132, 308, 2050, 3635, 494, 131, 2223, 2283, 2623, 2423, 2486, 3000, 3223, 869, 2659, 1240, 3107, 2952, 3258, 3192, 3547, 3519, 3689, 3415, 3007, 3234, 2437, 3474, 1356, 3393, 3102, 3369, 3413, 362, 3010, 1192, 1982, 2354, 2592]. Science [2888, 121, 784, 417, 1035, 1500, 1459, 1347, 1669, 1896, 3372, 3486, 561, 2817, 20, 788, 1032, 3465, 1298, 1529, 1890, 2124, 975, 2004, 2694, 906, 3090, 1660, 1967, 2198, 2323, 3429, 3458, 3598, 3684, 3712, 3746, 3656, 3483, 3542, 3514, 3571, 3625, 3777, 633, 2087, 2556, 2946, 2455, 2788, 3282, 3167, 2528, 2575, 2671, 1125, 669, 1117, 3549, 389, 1318, 2095, 3403, 3084, 1865]. Science-Based [1208]. Sciences [2986]. Scientia [2429]. Scientific [2090, 2357, 1511, 1593, 3393, 2507, 3211, 3239, 3268, 3298, 3343, 1871, 1910, 1948, 1985, 2033, 2070, 1671, 2643, 3, 27, 53, 79, 108, 134, 172, 199, 244, 273, 311, 347, 394, 431, 468, 504, 541, 578, 615, 650, 708, 725, 763, 799, 841, 875, 913, 947, 980, 1015, 1049, 1086, 1123, 1163, 1207, 1245, 1280, 1323, 1359, 1394, 1435, 1475, 1544, 1590, 1630, 1676, 1714, 1758, 1795, 1833, 2110, 2150, 2183, 2226, 2269, 2301, 2341, 2376, 2409, 2441, 2473, 2543, 2578, 2610, 2650]. Scientific [2681, 2711, 2741, 2773, 2804, 2874, 2903, 2951, 2988, 3037, 3088, 3127, 3154, 3182, 3355, 3391, 3434, 3445, 3488, 3501, 3529, 3557, 3586, 3613, 3644, 3671, 3699, 3732, 3782]. Scientist [3213, 3124, 868, 3012, 963, 543]. Scientists [3759, 427, 185, 2501]. Scientology [784]. SciPhones [483]. Scoop [1032]. Scooting [68]. Scope [3710]. Score [2088]. Scorecard [1659, 1194]. Scorpions [1284]. Scourge [1536, 626]. Scrappy [1405].

Scratch [2346, 3110]. Screen [3220, 2487, 2862]. Screening [1066]. Screens [2682, 1749, 1299, 3572]. Scrivener [480]. Scrubs [227]. Sculpting [2981]. Scum [844]. Sea [1857, 2830, 2478, 873, 3290, 2589, 533, 3478, 236, 3021, 3610, 3007, 1904]. Seafood [1606]. Seals [3135]. Search [2202, 1004, 1787, 2496, 1610, 3611, 423, 2263, 2702, 3070, 1977, 2403, 1530]. Searching [3495, 3628]. Seas [371, 3489, 3618, 1503]. Seashells [3583]. Season [1777]. Seat [2398]. Second [764, 3327, 3264, 2780, 1864, 1727, 2896]. Seconds [536]. Secret [1907, 2489, 3631, 2853, 2840, 111, 3647, 2928, 3432, 2575, 3795, 3796]. Secrets [1422, 2851, 3305, 3057, 1351, 1078]. Secure [3186]. Security [1747, 169, 2372, 1988, 3404, 1708]. See [2119, 694, 474, 2416, 3338, 3452, 816, 1532, 1480, 1221]. See-Through [2119]. Seeds [1980, 424, 1468]. Seeing [49, 2602, 3736, 1172, 2641, 2963, 2976, 525]. Seem [205]. Seer [1012]. Sees [2319]. Seismology [536, 3754]. Self [2091, 2127, 2934, 2870, 158, 3548, 2397, 2827, 3081, 2004]. Self-awareness [158]. Self-Control [2091]. Self-Driving [2827, 3081]. Self-Evident [2934]. Self-Help [2397, 2004]. Self-reliant [2870]. Self-Taught [3548]. Semiconsciously [3158]. Send [360, 2992]. Sensation [2796]. Sensational [305, 3]. Sense [3669, 3690, 3682]. Sensed [92]. Sensed-Presence [92]. Senseless [2990]. Senses [3771, 3769, 1320]. Sensory [2024]. Sequenced [3207]. Sergei [1169]. Serious [2872, 2863, 184, 3601, 1444, 2090]. Session [741]. Set [1127, 1259]. Seven [2728, 2930]. Seven-Year [2930]. Seveneves [2457]. Seventh [3690]. Sex [3375, 115, 223, 2840, 3376, 1133, 2549, 3507, 3660, 3701, 3378, 147, 2832, 2841]. Sex-Changing [115]. Shades [2349, 598]. Shadow [2200, 2031, 2439, 942, 3797]. Shadows [2974, 3549]. Shaky [260]. Shamed [2358]. Shaming [3001]. Shape [351, 1943, 3583, 695, 3397]. Shape-Shifters [3397]. Shape-Shifting [1943]. Shaped [3597]. Shapes [460]. Shaping [1735]. Share [522, 2022]. Sharing [2005]. Shattered [1352]. Sheet [985]. Sheets [253]. Shell [1045]. Shield [1775]. Shields [110]. Shifter [351]. Shifters [3397]. Shifting [1943, 499]. Shipments [28]. Shock [1625, 696, 1412, 1233, 2332]. Shocked [2092]. Shocks [2099, 3019]. Shopping [314]. Short [3199]. Shot [2907, 1902, 1216, 3508, 3003]. Shots [2343]. Should [2442, 2904, 3327, 1710, 3207, 2828]. Shouldn’t [3733]. Show [1725, 1419, 2684, 2775]. Showers [2745, 2383, 250]. Shrinking [3634]. Shrunk [3751]. Shyness [2868]. Sick [3646, 1235]. Sickle [2788]. Sickness [2063]. Side [2200, 751, 1992, 60, 3797, 433, 3137, 2838, 2861, 3001]. Sides [2380, 3518]. Siege [3446]. Sight [557, 1101, 205, 128, 3116]. Sight-Reading [3116]. Signals [496]. Significance [1937, 2067]. Silent [371, 2423, 3574]. Silicon [857, 2467]. Silk [343, 2752]. Silos [3760]. Simple [2498, 136]. Simulating [1788]. Simulation [3082]. Sin [185]. Singing [2348]. Siobhan [2491]. Siri [1103, 2521]. sitters [2377]. Six [2706]. Sixth [1817]. Size [2463, 830, 1748]. Sizing [1606]. Skeptic [369, 40, 296, 256, 121, 14, 213, 66, 92, 334, 147, 185, 636, 748, 564, 490, 672, 454, 601, 784, 417, 818, 527, 706, 1268, 1073, 1147, 967, 1109, 899, 1001, 934, 1189, 1233, 1035, 858, 1500, 1382, 1628, 1588, 1737, 1542, 1459, 1305, 1347, 1418, 1702, 1669, 1859, 2056, 2203, 1783, 1936, 1821, 2130, 2010, 1973, 2171, 1896, 2093, 2628, 2562, 2329, 2526, 2255, 2492, 2289, 2461, 2360, 2598, 2429, 2397, 2699, 2949, 3105, 2791, 2820, 2759, 2729, 2669, 3003, 2921, 2891, 3054, 3196, 3170, 3461, 3372, 3285, 3406, 3341, 3142, 3432, 3226, 3256, 3315]. Skeptic [3517, 3715, 3545, 3780, 3659, 3486, 3749, 3574, 3628, 3687, 3601]. Skeptical [454]. Skeptics [3406]. Sketch [1807]. Skies [2281, 2333, 1531, 3511]. Skill [654]. Skills [2956]. Skin [2884, 824, 34]. Skinlike [3360]. Skull [1952]. Skulls [757]. Sky [2089, 1422, 3605, 768, 2993, 2769, 1375, 3406, 3754, 3214]. Skyscrapers [297]. Slayer [1153]. Sleep [408, 2541, 3775, 3784, 2571]. Sleeping [2082, 905, 792, 1153, 2671]. Sleepyheads [1598]. Sleuth [796]. Sleuths [2879, 1138]. Slime [3441]. Slinking [1017]. Slinky [1371]. Slip [2279]. Slippery [3735]. Slither [3498]. Slope [3735]. Slow [2640, 3651, 1789]. Slow-Motion [2640]. Small [1587, 2388, 282, 77, 2364, 290, 2405, 1750, 2968, 1089, 1318]. Smaller [1403]. Smallest [481, 582, 789]. Smart [2853, 3667, 2852, 1520, 3654, 2005, 1968]. Smarter [710, 1152, 2367, 1103, 711]. Smartest [609]. Smartphone [1926]. Smartphones [1745, 3521]. Smarts [3478]. Smasher [2353]. Smashup [3119]. Smell [441]. Smells [1008]. Smoke [734, 89, 428]. Smokers [689]. Smoky [3451]. Smut [2838]. Snake [254]. Snakebites [2613]. Snakes [3498, 30]. Snaking [2159]. Snap [1743, 1167, 111, 225]. Snatching [920]. Sneezing [1055]. Sniff [283]. Snooping [1747]. Snowball [2756, 448]. So-Called [3712]. Soaked [2654]. Soar [3408]. Soaring [524]. Sociability [32]. Sociable [2863]. Social [1038, 2704, 770, 3234, 711, 3493, 249, 3570]. Society [2337, 106, 1710, 1240, 1662, 2841]. Sociobiology [1045]. Sociology [681]. Sock [2114]. Soft [1482, 2635]. Software [3116, 3113]. Soil [3247, 2801, 53]. Solar [96, 2178, 2515, 3119, 3350, 1883, 3457, 178, 2879, 3349, 1789, 1810]. Sold [2519]. Soldiers [928]. Solid [2981, 2087]. Solid-Gold [2087]. Solitary [1548]. Solitude [1665]. Solution [1711, 423, 2448, 1426, 2900]. Solutions [100, 567, 3157, 2232]. solv [3752]. Solve [1164]. Solved [82]. Solving [2128, 534]. Some [1059, 1090, 1093, 2775, 2164, 1217]. something [3164]. Sometimes [1554]. Song [1610]. Sonic [144]. Soon [1754, 2304]. Sooner [64]. Soot [928]. Sooty [2584]. Sorry [3539]. Soul [1105, 3628]. Soul-Searching [3628]. Sound [1886, 2334, 144, 8, 2446, 2214]. Soundings [1069]. Sounds [3287, 3215, 88]. Sour [1611, 250]. Source [2297]. Sous [405]. Southward [2600]. Sovereign [1702]. Space [1969, 2523, 2251, 2579, 973, 2546, 2475, 2496, 318, 2937, 1852, 1441, 2690, 3499, 3394, 421, 1275, 390, 2962, 1077, 2654, 3674, 3273, 832, 2175, 2795, 2702, 3070, 3177, 1063, 2972, 175, 766, 969, 921, 2189, 3146, 902, 494, 458, 1351, 217, 1447, 3423, 1315, 1430, 2924, 3072, 1333, 572]. Space-Saving [3423].

Spaceflight [1368, 3201]. Spaces [466]. Spacetime [873, 3057, 3073, 3146, 3636]. Spare [2201, 3246]. Sparks [575]. Speak [757]. Speaking [838, 3449]. Special [218, 752, 821, 1423, 1592, 1655, 2101, 2565, 2825, 3023, 3490, 3788, 3464, 98, 1569, 2364, 1462, 3320, 2499, 2926, 3693, 1005, 3231, 2764, 3132]. Species [94, 1872, 2842, 2058, 1615, 1093, 2495, 1119, 3061]. Speciesist [1783]. Speed [1667, 9, 1578, 3201, 803, 2945, 2928]. Speed-Read [2945]. Spell [1139]. Sperm [1644]. Spherical [743]. Spice [1861]. Spider [1599, 2223]. Spiders [1065, 1720]. Spies [1395, 1436, 1759, 587]. Spike [3317]. Spin [485, 2785, 3010]. Spinning [2700]. Spiral [1247]. Spiraled [3077]. Spirit [81]. Spiritual [1907]. Spit [2210]. Spit-Fired [2210]. Split [3264, 2787]. Split-Second [3264]. Sport [2697]. Sports [1586, 1331]. Spots [411]. Spreaders [1165]. Spreads [2063]. Spring [846, 1938]. Springs [3096, 3345]. Sprint [3566]. Spurious [2488]. Squad [3271]. Squeeze [2701]. Squid [286]. Squirrels [3462]. Stack [901]. Staggering [3546]. Stairs [2159]. Stale [441]. Stalking [1873]. Stamp [3168]. Stampede [2767]. Stand [651, 322]. Standing [3724]. Standoff [1396]. Star [1700, 3260, 910, 2709, 1249, 1527, 1141, 1386, 2443]. Star-Craving [1700]. Stardust [1142]. Starfish [2655]. Starlight [2006, 1899]. Starry [1275]. Stars [938, 1404, 996, 480, 3635, 929, 2574, 44]. Starship [1315, 3033]. Start [242, 64]. Starting [390, 2391, 2289]. Stat [588, 594, 665, 699, 733, 738, 888, 892, 918, 991, 1175]. State [1655, 2253, 2565, 3009, 3755, 2502, 2133, 3413, 1192]. Static [87]. Statins [87]. Station [572]. Statistical [1937]. Statistics [1538, 2093]. Status [329]. Stay [3213, 2003]. Stayed [2272]. Staying [839]. Stays [854]. STD [3184]. STDs [3131]. Steady [2291]. Stealing [1168]. Stealth [1291]. Steel [551, 3365]. Steer [1684]. Stellar [2824]. Stem [323, 56]. Steppe [3526]. Steps [2498, 1962]. Stick [944, 612]. Still [2722, 2923, 1993, 182, 2107, 3032, 15, 1348, 2241, 2262, 1193]. Sting [2219]. Stinks [3629]. Stock [3676]. Stolen [1866]. Stomach [2494, 3167]. Stone [1742, 385, 2762, 3020, 2026]. Stonehenge [497]. Stop [800, 2254, 2544, 2742, 3558, 1711, 1607, 589, 801, 2455, 2066, 3491]. Stopper [2438]. Stopping [1332]. Storage [3439, 1294]. Storing [55]. Storm [1380, 1851, 1506, 1287, 3457, 1010, 496, 2221]. Stormy [2339]. Story [1071, 1069, 2129, 2254, 3016, 945, 1357, 2802, 3697, 3750, 2076, 220, 3795]. Storybook [893]. Straight [2053, 703]. Strange [2427, 3260, 1454, 1316]. Strangest [570, 1239]. Strawberries [2513]. Stray [2474]. Stream [1763, 2220, 1353]. Streams [1525]. Street [710, 1301, 1886, 721, 2039]. Street-Legal [2039]. Street-Savvy [710]. Stress [688]. Stretchable [1749]. Stretching [2230]. String [570]. Stringing [142]. Stringy [1316]. Structures [1743]. Stuck [81]. Studio [2163]. Study [2538, 124]. Stuff [1932, 1464, 1741, 3304, 1179, 3609]. Stunt [2125]. Style [1721, 3602, 1678]. Subject [2456]. Subscription [3459]. Substance [2257]. Substances [1745]. Subtract [2028]. Succeeding [1047]. Success [2802, 3229, 3432, 2280]. Successful [2851]. Successor [2073]. Sucker [305]. Sucks [693]. Sufferers [3165]. Suffering [22, 1927]. Suffers [3530]. Suffocated [3618]. Sugar [1957, 1689]. Suggests [2478]. Suicide [3581]. Suicides [3558, 2817]. Suits [2244]. Summer [692]. Summon [1979]. Sun [3631, 3230, 3611, 1843, 204]. Sunken [2263, 2513]. Sunny [2838]. Suns [3588, 1706]. Sunscreen [3360]. Sunshine [1408, 3401]. Sunspot [3689]. Super [1622, 1039, 1425, 1294, 2673, 2591, 2521, 1165, 1151, 3282, 3118]. Super-Molecules [3118]. Super-Spreaders [1165]. Superatoms [3118]. Superbug [1561]. Superbugs [1332, 3123]. Supercharge [3632]. Supercharging [3611]. Superconductors [157]. Superego [857]. Superfast [1745]. Supergiant [1682]. Supergrid [339]. Supermarket [254]. Supernova [1754]. Supernovae [1039]. Supersmart [2828]. Supersonic [61]. Superstorm [1559]. Superstrong [1743]. Supersymmetry [983, 1939]. Superweeds [573]. Supply [1621]. Support [1669]. Surf [241, 1956]. Surface [2325, 1756, 1367, 749]. Surge [1458, 3097]. Surgery [655]. Surprise [2564, 3689]. Surprised [1579]. Surprising [2788, 1582]. Survival [2962]. Survive [2372, 1287]. Surviving [66, 2093]. Suspect [772, 774, 804, 863]. Suspicious [2596]. Sustain [791]. Sustainability [3494, 567, 537, 98, 2706, 791, 3439, 2029, 2504, 459, 2370, 872, 2735, 3083]. Sustainable [1606, 3320, 65, 255, 91, 39, 120, 184, 146, 212, 13, 907]. Swagger [2414]. Swallow [3230]. Swallows [2298]. Swamp [1755]. Swan [606]. Swap [2074, 1133]. Swapping [815]. Sweaty [3594]. Sweep [1140]. Sweet [1426, 3245]. Sweeteners [2354]. Swell [3280]. Swept [904, 84]. Swim [1583]. Swimmers [2799]. Swimming [960]. Swindle [2515]. Switch [252, 2675, 1940]. Switches [2637, 835, 2152]. Switching [2380]. Sword [2169]. Symbiont [2170]. Symbol [1915]. Symptoms [1438]. Sync [891]. Syndrome [908]. Synthetic [41]. Syria [2735, 3083]. System [616, 96, 3119, 1412]. Systems [91].

T [2904, 2124]. T. [2273]. Table [640, 1507, 2200, 3797]. Tablet [1749]. Taboos [2360]. Tadpoles [1441]. Tag [326, 3134, 1235]. Taikonauts [878]. Tail [927, 1219]. Tailing [1449]. Tainted [1603]. Take [3128, 3199, 216, 3583]. Taken [761]. Takeoff [994, 2355, 199]. Takes [3099, 1088, 3189, 1922]. Taking [1368, 3149, 1762, 2998, 2138, 2219, 3676]. Tale [1935, 1391, 740, 1325, 2957, 2221]. Talent [2035]. Tales [2427, 2762, 3020]. Talk [703, 721, 2417, 1409, 2606, 2405, 367, 2685]. Talkers [956]. Talking [3352, 1445, 3722]. Talks [2614]. Tame [508, 1826]. Tan [2342]. Tangled [1644, 3146]. Tank [2422]. Tape [2686, 3535]. Tapping [3321]. Target [1165]. Taste [1644, 1090, 2703]. Taste-Blind [1644]. Tasteless [1460]. Tastier [2027]. Tattoos [1690]. Taught [3548]. Taxes [779, 3300]. Taxi [2818]. Taxonomy [3041]. Taxpayer [1602]. Teach [2927, 2639]. Teachable [309]. Teacher [1117]. Team [3134, 3785, 2500]. Tears [2938]. Tech [529, 2811, 2345, 1778, 2624, 3051, 3484, 2042]. Technical [1241]. Techno [3719]. TechnoFiles [294, 367, 332, 746, 599, 415, 562, 488, 816, 782, 452, 670, 704, 525, 634, 1033, 897, 1141, 1263, 1183, 857, 1103, 1067, 965, 931, 1227, 999, 1697, 1413, 1613, 1652, 1495, 1566, 1733, 1342, 1454, 1530, 1377, 1299, 1815, 2166, 1853, 1930, 1891, 2125, 2005, 1968, 2051, 1778, 2088, 2199, 2284, 2593, 2456, 2324, 2557, 2521, 2392, 2624, 2487, 2424, 2355, 2250, 3001, 2664, 2754, 2919, 2693, 2818, 2789, 3051, 3103, 2724, 2947, 2889, 3254, 3313, 3370, 3459, 3430, 3168, 3283, 3339, 3224, 3194, 3140, 3404, 3572, 3626, 3778, 3484, 3685, 3747, 3657, 3515, 3543, 3599, 3713]. Technologies [3464]. Technology [266, 821, 1308, 3067, 608, 1041, 2334, 942, 2260, 104, 3521, 2705, 2436, 2605, 1866, 1827, 1510, 21, 2027, 2798, 343, 1662, 1067, 965, 2051, 3339, 3493, 240, 871, 3259, 1388]. Tectonics [3448]. Teen [2544, 3099, 2432, 2870, 2871, 3458, 2379]. Teenagers [2201]. Telescope [2645]. Telescopic [3134]. Tell [1107, 2904, 349, 997, 899]. Tells [1971, 3164]. Telltale [1009, 3215]. Telomeres [511]. temperature [157]. Temperatures [2058]. Tense [2890]. Tension [3041, 749]. Tensions [878]. Tentative [1915]. Term [1294, 1377]. Terminate [183]. Terminators [183]. Termites [3696]. Territory [3310, 2514]. Terror [145]. Terrorism [1715, 1588, 2397]. Terrorists [1903, 3256]. Test [1324, 2091, 1102, 3439, 209, 605, 2166, 2537, 2562, 2688, 2548]. Test-Tube [2548]. Testing [726, 12, 547, 2499, 2857, 2500]. Testosterone [32]. Testosterone-Fueled [32]. Tests [380, 3205, 3102, 2946]. Texas [2253]. Text [1697, 2199]. Textbooks [959]. Thar [1022]. Theft [750]. Their [265, 475, 1062, 2365, 2684]. Them [944, 1514, 3724, 809]. Themselves [1083, 253, 2639]. Theorem [2394]. Theoretical [3057, 2241]. Theorist [1277]. Theory [1818, 2126, 570, 3035, 2536, 303, 384, 1327, 245, 276, 2470, 369]. Therapy [2015, 1864, 2176, 2685, 3147, 2229]. There [2089, 2559, 287, 3377, 2433, 2760, 547, 1942, 2719, 3684, 3164]. Therefore [1139]. These [2107, 3032, 1348, 3130, 3164]. They’re [418]. Thief [1227]. thin [1745]. Thing [1462, 3407, 1566, 3514]. Things [1592, 1894, 1943, 1825, 2057, 746, 3060, 1193, 2194]. Think [1083, 167, 2055, 2422, 132, 195, 2852, 1139, 1710]. Thinker [2085]. Thinkers [1062]. Thinking [1541, 2053, 211, 698, 3663, 3651, 2308]. Thinks [823]. Thirdhand [89]. Thirsty [516]. Thorny [226]. Those [287, 2947]. Thought [2254, 460, 1849, 1797, 2503, 2532, 790, 2985]. Thoughts [1868, 754]. Threat [169, 2958, 3123, 1389, 176]. Threatened [2043]. Threatening [238]. Threatens [1420, 3777]. Threats [100, 3221, 2420]. Three [2023, 2832]. Three-Pound [2832]. Thrive [1773]. Thrombosis [2233]. Tibetans [3195]. Ticket [1430]. Ticks [15]. Tiger [508, 2891]. Tigers [1734, 2898]. Tight [1305]. Timber [3596]. Time [2059, 2251, 1851, 165, 3786, 1712, 2536, 3319, 2015, 1289, 2962, 3773, 2378, 3236, 3722, 923, 976, 269, 3547, 294, 931, 3531, 2230]. Timed [2984]. Timeline [2264]. Times [2523, 1883, 2204]. Timing [3135]. Tinder [3744]. Tiniest [1439]. Tiny [2387, 1503]. Titan [1956]. Titans [1006]. To-Do [1281, 3283]. Today [3295, 2717]. Toddler [2812]. Together [659, 1047].

Toilets [2387]. Told [1970]. Toll [375]. Tolls [410]. Tomatoes [1090]. Tomorrow [3295, 1005, 2829]. Tongue [111, 3532, 1269, 364]. Too [1059, 429, 1254, 2878, 1372, 517, 512, 2961, 2198, 2501]. Tool [2812]. Toolbox [1748]. Toolbox-size [1748]. Tools [201, 3130]. Tooth [1138]. Toothless [2916]. Top [3464, 3144, 2047, 1176, 2748, 2551]. Top-down [2047]. Tortoise [811]. Tortoises [618]. Total [1649]. Touch [802, 3368, 1299, 2487, 2862, 3572]. Tourette [584]. Towering [1646]. Toxic [3418, 3497, 74]. Toxin [2856]. Toxins [727]. Trace [2970]. Traces [1407]. Track [571, 1227]. Tracking [2898, 1063, 997]. Traffic [1802, 141]. Trafficking [3596]. Trail [3013, 1709, 766, 224]. Trails [1674]. Trait [3043]. Traitors [1118]. Trans [1835]. Transformed [3091]. Transformers [3231]. Transfusion [2448]. Transgender [2845]. Transistor [376]. Translation [2330]. Transparency [3159]. Transparent [2209, 2337]. Transplant [1647]. Transportation [127, 46, 191, 3081]. Transylvania [759]. Trap [3382, 3303, 501]. Traps [3424]. Trash [766, 381, 3321, 280]. Trashing [1053]. Trauma [3166]. Travel [2536]. Traveler [976]. Traveling [1026]. Travels [2761]. Trawling [2634]. Treasure [679, 740, 2901, 2263, 1841, 280]. Treat [3159]. Treating [2856, 584, 2297]. Treatment [2111, 2219]. Treatments [2304]. Tree [591, 1329, 3578]. Trees [679, 1004, 1951, 1646, 3594]. Tresses [1917]. Triage [3278]. Trials [3700, 2016]. Tribalism [3524]. Tribes [2051]. Tribulations [2016]. Trick [1829]. Tricks [1011, 1023]. Trigger [258]. Trillion [425]. Trillions [3228]. Trip [2532]. Triumph [3604, 1006]. Trivers [1234]. Trolls [3384]. Tropical [24]. Trouble [3287, 1816, 739, 332, 999, 1299, 2566, 1845]. Troubled [3523]. True [1970, 2254, 1051, 1307, 2463, 3100, 1428, 3750, 3060, 493, 3208]. Truer [2982]. Truffles [105]. Trump [3038]. Trumps [3156, 3054]. Trust [301, 1107, 1513, 3203, 2311, 3493, 543, 1778]. Trustworthy [1274]. Truth [1587, 3009, 3035, 2844, 3524, 349, 45, 3024, 2982, 795, 191, 2392, 3289, 3226, 2827, 3081]. Try [3712]. Trying [3062]. Tsunami [1312]. Tsunamis [3215]. Tube [2548]. Tubes [1750]. Tug [232]. Tumors [669]. Tuna [1449]. Tunnel [2696]. Turbines [142]. Turing [3205]. Turn [2486]. Turning [1601, 1742]. Turns [144]. Turtle [1962]. Turtles [1063, 1684]. Tushi [2599]. Tusk [1803]. TV [411, 2563]. Tweaking [843]. Tweaks [1477, 1036]. Tweets [2934, 287, 2614]. Tweezers [1958]. Twice [1912]. Twilight [350]. Twin [3107]. Twist [2746, 136]. Twisted [2969]. Twisters [1269]. Twists [2431, 2145]. Twitching [1409]. Twitter [1953]. Twitterverse [656]. Two [2071, 137, 2102, 2846, 3027, 1325, 1506, 1883, 2957, 2612, 1997, 2051, 1706]. Two-Year [2071]. Tycoons [921]. Type [1747]. Tyrannosaurs [2400]. Tyrants [2374].

U. [1196]. U.S. [1087, 3221, 30, 1602]. UAPs [527]. UFOs [527]. Ugly [2201, 1730, 2808]. Ultimate [1038, 1791, 3493, 3115]. Ultrahard [2213]. Ultrasonic [630]. Unblock [2815]. Uncanny [128]. Uncharted [3310]. Unclear [2355]. Unconscious [1786]. Uncounted [715]. Uncover [2252, 3793]. Uncowed [853]. Underage [119]. Underground [3059, 1180, 643, 1742]. Undermines [2714, 3241]. Underwater [2264, 3692]. Undifferentiated [247]. Undocumented [2201]. Unease [2462]. Unfolding [1895]. Unfree [81]. Unhealthy [2310]. Unification [341]. Uninformed [487]. Unintelligible [2126]. Uninterrupted [142]. Unique [937]. Uniqueness [122, 3718]. Universal [812, 3503]. Universe [1228, 1300, 2006, 2127, 1641, 189, 2031, 2439, 3174, 1913, 2469, 3442, 2562, 3170]. University [926]. Unkindest [935]. Unknown [2653, 336, 2954]. Unknowns [2023]. Unlikely [3604]. Unlocking [3291, 2552]. Unnatural [3232]. Unpredictive [2199]. Unquantum [1272]. Unraveling [1397]. Unreality [1970]. Unschooled [876]. Unseen [386, 2000]. Unsettled [2088]. Unsolved [753]. Unsound [1250]. Untangled [2724]. Unusual [1548]. Unwanted [3772]. Unyielding [3263]. Update [329]. Upgrade [1613, 2424]. Upset [3167]. Urban [2983, 715, 229, 2952]. Urbanite [2880]. Urge [1181]. Urgency [2691]. Urine [441]. Use [210]. used [3491]. User [3224]. Using [1926].

v [3787]. Vaccination [703]. Vaccine [3076, 2765, 2418, 3422, 1777, 3503, 975]. Vaccines [571, 3120]. Vacuum [2638]. Vanished [2970, 3285, 3294]. Vanishing [1101]. Vanquishing [3525]. Vaquita [3346]. vector [2229]. Veggies [2027, 3136]. Vents [78]. Venus [2775]. Versus [2021]. Vertical [199]. Very [3697]. Vial [145]. Vibrations [1411, 2282]. Vicissitudes [2549]. Victory [194]. Vide [405]. Video [2894, 2216, 1141, 3543, 889]. Videos [383]. Videotape [3753]. View [2712, 2136, 1501, 284]. Views [3068, 2986, 3597]. Vikings [3294]. Violence [748, 3300, 3039]. Violent [20]. Viral [1092, 2312, 1785, 662]. Virology [1046]. Virtual [1970, 21, 1226]. Virtually [2118]. Virtues [1226]. Virus [1068, 2176, 3115]. Viruses [1686, 2634, 140]. Visible [137, 2142]. Vision [463, 953, 2211, 2499, 2857, 3136]. Vision-Correcting [2211]. Visions [2696]. Visual [2967, 2981, 2980, 2985]. Vital [3503]. Vitamin [56]. Vitamins [806]. Vive [2072]. Volcano [2695]. Volcanoes [1914, 3426]. Voting [1209]. Voyage [1105, 1306]. Voyager [1698]. vs [3253, 317, 1561, 472, 539, 786, 2257, 3204, 2321, 3372, 554].

Wag [3453]. Waiting [1725, 758, 1043]. Wall [1301, 1974]. Wallet [919, 824]. Walls [3269, 1537, 3654]. Wandering [378]. Waning [3306]. Want [1252, 2917]. Wanted [1406, 1988, 1689]. Wants [949, 1390]. War [2326, 3327, 3587, 59, 2546, 2335, 3241, 2279, 1878, 2922, 836, 80, 22, 3474, 3441, 1347, 3461, 192, 975, 3496, 1345, 1796]. Warfare [2766, 47]. Warming [1238, 2829, 642, 612, 3262]. Warms [1634]. Warning [2020, 544, 2275]. Warp [9]. Warp-Speed [9]. Warrant [862]. Warrior [3151]. Wars [171, 2800, 2178, 2829, 2645, 2443]. Was [2007, 448, 1579, 324, 2433, 2718, 353, 1543, 2538]. Washing [166, 1219]. Washington [1281, 3155]. Wasn’t [2559]. Waste [2742, 2215, 2917].

Wasteland [969]. Wasting [3767]. Watch [1562, 955, 989, 1096, 1178, 1218, 1255, 1290, 1330, 1364, 1399, 1450, 1489, 1522, 1687, 288, 354, 409, 2406, 523, 555, 625, 690, 730, 856, 2482, 327]. Watches [919, 1968]. Watchful [3590]. Watching [3783, 415, 3404]. Water [1229, 1094, 1585, 3672, 3096, 1806, 2674, 1037, 1603, 2882, 193, 1537, 1214, 1019, 377, 3280, 3235, 2518]. Watermarks [86]. Waters [2778, 3523, 3055]. Watson [509]. Waves [3694, 2404, 2214]. Way [1886, 3762, 1591, 1826, 1044, 1176, 832, 968, 767, 3105, 751, 1723, 2025, 3729, 3091, 3619, 2217]. Ways [3056]. Weakest [2480]. Weakness [2259]. Wealth [2952, 3790, 3491]. Weapon [3477]. Weapons [3236, 3129]. Weapons-Free [3129]. Wear [2349, 1514]. Wearable [3512]. Weather [123, 3013, 3667, 2993]. Weather-Smart [3667]. Web [388, 942, 936, 1863, 3318]. Weed [3587, 115, 2586, 3257]. Weedy [626]. Week [188]. Weighing [987, 855]. Weight [3289]. Weird [1343, 2220]. Weirdest [3599]. Weirdness [1806, 1504]. Welcome [3353, 2098, 3017]. Well [1987, 2622, 3484, 3497]. We’re [1158, 1446]. West [1585, 3165]. We’ve [2052]. Whale [1414]. Whales [50]. Wheat [3476]. Wheel [2776, 2593]. Wheels [2700]. Where [2097, 2105, 3016, 3030, 1158, 2679, 2677, 3576, 1882, 1744, 2134, 1739, 1329, 2672, 3074, 2607, 3048, 2859, 2943, 667, 2672, 3074]. Which [2155, 3216, 1149, 434, 1119, 1619]. While [1634]. Whispers [2060, 2073]. Whistled [3178]. White [2597]. Who [948, 1069, 1436, 2055, 2254, 2559, 2538, 3614, 2826, 2172, 1923, 2776, 1656, 1056, 997, 415, 1151, 3388, 3315, 3493]. Whole [397, 2469]. Whom [410]. Whorl [2662]. Whose [3130]. Wi [999]. Wi-Fi [999]. Wide [392, 664]. Wider [3603]. Wild [1610, 1739, 626, 3602, 3781, 1521, 3262]. Wildfire [3616]. Wildlife [3151]. Will [2034, 2722, 3763, 151, 589, 1154, 2416, 1656, 3338, 2684, 2265, 404, 2262, 3140, 1982, 2943, 1619, 1480, 1221, 1119]. Willy [1834]. Wimp [810]. WIMPs [2684]. Win [2916]. Wind [1265, 940, 1789]. Windfall [1782]. Winding [1966]. Window [314]. Winds [3049]. Wine [3709, 3451]. Wing [3149]. Winged [194]. Wingman [884]. Wings [1356]. Winner [1088]. Wins [2692]. Winter [446]. Winters [1273]. Wipe [3275]. Wipeout [793]. Wired [1797, 3723]. Wireless [104, 2214]. Wisdom [287, 1203, 2704, 1598, 1437]. Wise [1344]. Wish [395, 17, 2820]. Wishes [893]. Witch [2668]. Witches [3256]. Withdraw [176]. Within [1300, 865, 2299, 3416, 3552, 2302, 531, 557, 3665, 2294]. Without [2299, 2970, 1982, 1585, 3269, 2590, 826]. Witness [1080]. Wizard [1265]. Wobble [1136]. Wolf [2468]. Woman [1069, 1969, 2560, 3388]. Women [1016, 2904, 3356, 3375, 3376, 3773, 331, 3385, 176]. Wonder [2087]. Wonders [1972, 3065, 96, 2364, 95, 1318]. Won’t [1109, 2218, 3625]. Wood [2787, 3768, 3365]. Woodland [1695]. Woods [3306]. Wooing [2377]. Word [3178]. Words [2458, 1866, 2760]. Work [1184, 2034, 2310, 2306, 434, 509, 3385]. Workouts [3152]. Works [1580, 2030, 682, 2135]. World [372, 383, 821, 1068, 1668, 1634, 1655, 1740, 1894, 2052, 2631, 2570, 3108, 3009, 3606, 3755, 2371, 2020, 3470, 2177, 342, 77, 1983, 3086, 3389, 2260, 1194, 2850, 2893, 2206, 791, 1487, 2133, 1271, 2642, 2207, 2347, 1638, 2748, 3492, 3413, 1192, 1982, 604, 2565]. Worlds [1406, 2224, 2771, 337, 21, 2957, 1706, 3549]. Worm [75, 1328, 1060]. Wormholes [3057]. Worms [1744, 3675, 3336]. Worry [3236, 415]. Worse [2486]. Worst [1930]. Worth [2458, 1131, 3610]. Worts [207]. Would [2190, 70, 831]. Wrapped [592]. Wraps [140]. Wreck [1581]. Wretched [1975]. wring [3164]. Wrinkle [923]. Writing [1267]. Written [1422]. Wrong [1618, 1044, 2533, 2011, 417]. WTF [2126].

X [463, 1185, 1791, 2679, 2702, 3070, 593]. X-aggeration [593]. X-ray [463, 1791]. XX [3379]. XY [3379].

Yawn [811]. Yawning [1106]. Year [752, 2071, 3064, 1617, 257, 809, 2930, 1750, 2131, 858]. Years [1665, 1671, 2643, 1158, 23, 1506, 3350, 2107, 3032, 212, 3, 27, 53, 79, 108, 134, 172, 199, 244, 273, 311, 347, 394, 431, 468, 504, 541, 578, 615, 650, 708, 725, 763, 799, 841, 875, 913, 947, 980, 1015, 1049, 1086, 1123, 1163, 1207, 1245, 1280, 1323, 1359, 1394, 1435, 1475, 1544, 1590, 1630, 1676, 1714, 1758, 1795, 1833, 1871, 1910, 1948, 1985, 2033, 2070, 2110, 2150, 2183, 2226, 2269, 2301, 2341, 2376, 2409, 2441, 2473, 2507, 2543, 2578, 2610, 2650, 2681, 2711, 2741]. Years [2773, 2804, 2874, 2903, 2951, 2988, 3037, 3088, 3127, 3154, 3182, 3211, 3239, 3268, 3298, 3343, 3355, 3391, 3434, 3445, 3488, 3501, 3529, 3557, 3586, 3613, 3644, 3671, 3699, 3732, 3782]. Yeast [1963, 3423]. Yes [1720, 1446]. Yogurt [995]. Yorick [2971]. York [681]. You’ll [1566]. Young [3759, 839, 2576, 418]. You’re [2108, 1235, 450]. Yourself [2371]. YouTube [451]. You’ve [2358].

Zero [2252, 3672, 234, 102, 1812, 498, 3793]. Zika [2761, 2881, 2914, 3077]. Zika-Proofing [2914]. Zinc [1365]. Zombie [769, 2931]. Zombies [2255]. Zone [1847]. Zones [350]. Zoology [2830]. Zoom [2009]. Zoomable [3442].

References

DiChristina:2010:ELQ

[1] Mariette DiChristina. From the Editor: Life quest. Scientific American, 302(1):6, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

      scientificamerican/journal/v302/n1/full/scientificamerican0110-
      6.html;   http://www.nature.com/scientificamerican/journal/
      v302/n1/pdf/scientificamerican0110-6.pdf.

Anonymous:2010:LFC

[2] Anonymous. Letters: The first computer. Scientific American, 302(1):8– 10, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n1/full/scientificamerican0110-8.html;   http:
//www.nature.com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
8.pdf.

Schlenoff:2010:YAIa

[3] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: Sensational press. Scientific American, 302(1):12, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-
12.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-12.pdf.

Biello:2010:NSC

[4] David Biello. News scan: Climate numerology. Scientific American, 302(1):14–15, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-
14.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-14.pdf.

Harmon:2010:RH

[5] Katherine Harmon. Renewed hope. Scientific American, 302(1):15–16, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n1/full/scientificamerican0110-15.html;   http:
//www.nature.com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
15.pdf.

Woodbury:2010:BF

[6] M. A. Woodbury. Back into the folds. Scientific American, 302(1):16–18, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n1/full/scientificamerican0110-16.html;   http:
//www.nature.com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
16.pdf.

Stix:2010:LB

[7] Gary Stix. A light in the brain. Scientific American, 302(1):18–20, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n1/full/scientificamerican0110-18.html;   http:
//www.nature.com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
18.pdf.

Greenemeier:2010:SA

[8] Larry Greenemeier. Sound approach. Scientific American, 302(1):20–22, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n1/full/scientificamerican0110-20.html;   http:
//www.nature.com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
20.pdf.

Castelvecchi:2010:WSA

[9] Davide Castelvecchi. Warp-speed algebra. Scientific American, 302(1):22–24, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-
22.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-22.pdf.

Choi:2010:MM

[10] Charles Q. Choi. Machismo mayhem. Scientific American, 302(1):24, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n1/full/scientificamerican0110-24.html;   http:
//www.nature.com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
24.pdf.

Musser:2010:LLC

[11] George Musser. A large lump of coal. Scientific American, 302(1):26, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n1/full/scientificamerican0110-26.html;   http:
//www.nature.com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
26.pdf.

Editors:2010:PBN

[12] The Editors. Perspectives: Big need for a little testing. Scientific American, 302(1):28, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-

28.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-28.pdf.

Sachs:2010:SDN

[13] Jeffrey D. Sachs. Sustainable developments: The need for open process. Scientific American, 302(1):30, January 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n1/full/scientificamerican0110-30.html;   http://www.nature.
com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
30.pdf.

Shermer:2010:SKA

[14] Michael Shermer. Skeptic: Kool-Aid psychology. Scientific American, 302(1):39, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-
39.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-39.pdf.

Krauss:2010:CMDa

[15] Lawrence M. Krauss. Critical mass: The Doomsday Clock still ticks. Scientific American, 302(1):40, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v302/n1/full/scientificamerican0110-
40.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-40.pdf.

Anonymous:2010:RCR

[16] Anonymous. Recommended: Cosmic revelations. Scientific American, 302(1):98, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-
98.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-98.pdf.

Mirsky:2010:AGCa

[17] Steve Mirsky. Anti gravity: Careful what you wish for. Scientific American, 302(1):100, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-
100.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-100.pdf.

Jenkins:2010:CLL

[18] Alejandro Jenkins and Gilad Perez. Cosmology: Looking for life in the multiverse. Scientific American, 302(1):42–49, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n1/full/scientificamerican0110-42.html;   http://www.nature.
com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
42.pdf.

Young:2010:BRF

[19] John D. Young and Jan Martel. Biology: The rise and fall of nanobacteria. Scientific American, 302(1):52–59, January 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n1/full/scientificamerican0110-52.html;   http://www.nature.
com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
52.pdf.

Simpson:2010:ESV

[20] Sarah Simpson. Earth science: Violent origins of continents. Scientific American, 302(1):60–67, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-
60.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-60.pdf.

Heeks:2010:ITR

[21] Richard Heeks. Information technology: Real money from virtual worlds. Scientific American, 302(1):68–73, January 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n1/full/scientificamerican0110-68.html;   http://www.nature.
com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
68.pdf.

Robock:2010:ELN

[22] Alan Robock and Owen Brian Toon. Environment: Local nuclear war, global suffering. Scientific American, 302(1):74–81, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n1/full/scientificamerican0110-74.html;   http://www.nature.
com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
74.pdf.

Editors:2010:CNY

[23] The Editors. Computers: The next 20 years of microchips. Scientific American, 302(1):82–89, January 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n1/full/scientificamerican0110-
82.html;   http://www.nature.com/scientificamerican/journal/
v302/n1/pdf/scientificamerican0110-82.pdf.

Hotez:2010:MPD

[24] Peter Jay Hotez. Medicine: A plan to defeat neglected tropical diseases. Scientific American, 302(1):90–96, January 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n1/full/scientificamerican0110-90.html;   http://www.nature.
com/scientificamerican/journal/v302/n1/pdf/scientificamerican0110-
90.pdf.

DiChristina:2010:ECC

[25] Mariette DiChristina. From the Editor: Continuum of change. Scientific American, 302(2):6, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
6.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-6.pdf.

Anonymous:2010:LOC

[26] Anonymous. Letters: The oil crisis. Scientific American, 302(2):8–10, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n2/full/scientificamerican0210-8.html;   http:
//www.nature.com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
8.pdf.

Schlenoff:2010:YAIb

[27] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: Cosmic dust. Scientific American, 302(2):12, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
12.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-12.pdf.

Mukerjee:2010:NSP

[28] Madhusree Mukerjee. News scan: Poisoned shipments. Scientific American, 302(2):14–15, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
14.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-14.pdf.

Biello:2010:NC

[29] David Biello. Negating “Climategate”. Scientific American, 302(2):16, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n2/full/scientificamerican0210-16a.html;   http:
//www.nature.com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
16a.pdf.

Tennesen:2010:PBB

[30] Michael Tennesen. Python boom: Big snakes poised to change U.S. ecosystems. Scientific American, 302(2):16–18, February 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n2/full/scientificamerican0210-16b.html;   http://www.nature.
com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
16b.pdf.

Scudellari:2010:MM

[31] Megan Scudellari. Mouse mash-up. Scientific American, 302(2):18–21, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n2/full/scientificamerican0210-18.html;   http:
//www.nature.com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
18.pdf.

Harmon:2010:TFS

[32] Katherine Harmon. Testosterone-fueled sociability. Scientific American, 302(2):20, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
20.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-20.pdf.

Choi:2010:LG

[33] Charles Q. Choi. Lost giants. Scientific American, 302(2):21–22, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n2/full/scientificamerican0210-21.html;   http:
//www.nature.com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
21.pdf.

Storrs:2010:HS

[34] Carina Storrs. Hearing with skin. Scientific American, 302(2):22, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n2/full/scientificamerican0210-22.html;   http:
//www.nature.com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
22.pdf.

Kuchinskas:2010:NG

[35] Susan Kuchinskas. Naked gaming. Scientific American, 302(2):24, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n2/full/scientificamerican0210-24.html;   http:
//www.nature.com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
24.pdf.

Moyer:2010:BBB

[36] Michael Moyer. Bigger, better broadband. Scientific American, 302(2):26, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
26.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-26.pdf.

Little:2010:AE

[37] Crispin T. S. Little. Ask the experts. Scientific American, 302(2):86, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n2/full/scientificamerican0210-86.html;   http:
//www.nature.com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
86.pdf.

Editors:2010:PCE

[38] The Editors. Perspectives: Comparatively easy. Scientific American, 302(2):26, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
26A.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-26A.pdf.

Sachs:2010:SDFa

[39] Jeffrey D. Sachs. Sustainable developments: Fixing the broken policy process. Scientific American, 302(2):28, February 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n2/full/scientificamerican0210-28.html;   http://www.nature.
com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
28.pdf.

Shermer:2010:SCYa

[40] Michael Shermer. Skeptic: Cultivate your garden. Scientific American, 302(2):30, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
30.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-30.pdf.

Krauss:2010:CMR

[41] Lawrence M. Krauss. Critical mass: The real promise of synthetic biology. Scientific American, 302(2):32, February 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n2/full/scientificamerican0210-32.html;   http://www.nature.
com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
32.pdf.

Anonymous:2010:RSB

[42] Anonymous. Recommended: The science of bling. Scientific American, 302(2):87, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
87.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-87.pdf.

Mirsky:2010:AGG

[43] Steve Mirsky. Anti gravity: Greenhouse bananas. Scientific American, 302(2):88, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
88.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-88.pdf.

Young:2010:ACC

[44] Erick T. Young. Astronomy: Cloudy with a chance of stars. Scientific American, 302(2):34–41, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
34.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-34.pdf.

Jablonski:2010:ENT

[45] Nina G. Jablonski. Evolution: The naked truth. Scientific American, 302(2):42–49, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
42.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-42.pdf.

Knight:2010:TBM

[46] Ben Knight. Transportation: Better mileage now. Scientific American, 302(2):50–55, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
50.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-50.pdf.

Finlay:2010:MAB

[47] B. Brett Finlay. Medicine: The art of bacterial warfare. Scientific American, 302(2):56–63, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
56.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-56.pdf.

Townsend:2010:EFG

[48] Alan R. Townsend and Robert W. Howarth. Environment: Fixing the global nitrogen problem. Scientific American, 302(2):64–71, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n2/full/scientificamerican0210-64.html;   http:

//www.nature.com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
64.pdf.

Billock:2010:PSF

[49] Vincent A. Billock and Brian H. Tsou. Perception: Seeing forbidden colors. Scientific American, 302(2):72–77, February 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n2/full/scientificamerican0210-72.html;   http://www.nature.
com/scientificamerican/journal/v302/n2/pdf/scientificamerican0210-
72.pdf.

Little:2010:LSP

[50] Crispin T. S. Little. Life science: The prolific afterlife of whales. Scientific American, 302(2):78–84, February 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n2/full/scientificamerican0210-
78.html;   http://www.nature.com/scientificamerican/journal/
v302/n2/pdf/scientificamerican0210-78.pdf.

DiChristina:2010:ELP

[51] Mariette DiChristina. From the Editor: Limits of perception. Scientific American, 302(3):6, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
6.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-6.pdf.

Anonymous:2010:LR

[52] Anonymous. Letters: Renewables. Scientific American, 302(3):8–10, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n3/full/scientificamerican0310-8.html;   http:
//www.nature.com/scientificamerican/journal/v302/n3/pdf/scientificamerican0310-
8.pdf.

Schlenoff:2010:YAIc

[53] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: Soil progress. Scientific American, 302(3):12, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-

12.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-12.pdf.

Moyer:2010:NSD

[54] Michael Moyer. News scan: The deadliest catch. Scientific American, 302(3):14–17, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
14.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-14.pdf.

Biello:2010:CRL

[55] David Biello. Carbon rock lock: Storing CO_{2} on the East Coast. Scientific American, 302(3):17, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
17a.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-17a.pdf.

Storrs:2010:SCV

[56] Carina Storrs. Stem cell vitamin boost. Scientific American, 302(3):17– 18, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n3/full/scientificamerican0310-17b.html;   http:
//www.nature.com/scientificamerican/journal/v302/n3/pdf/scientificamerican0310-
17b.pdf.

Harmon:2010:MCF

[57] Katherine Harmon. Mutant cholesterol fends off dementia. Scientific American, 302(3):18, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
18a.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-18a.pdf.

Choi:2010:HBH

[58] Charles Q. Choi. Heavy brows, high art. Scientific American, 302(3):18– 19, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n3/full/scientificamerican0310-18b.html;   http:
//www.nature.com/scientificamerican/journal/v302/n3/pdf/scientificamerican0310-
18b.pdf.

Baker:2010:NW

[59] Linda Baker. Numbers war. Scientific American, 302(3):20–21, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n3/full/scientificamerican0310-20.html;   http:
//www.nature.com/scientificamerican/journal/v302/n3/pdf/scientificamerican0310-
20.pdf.

Choi:2010:DSB

[60] Charles Q. Choi. Dark side of black holes. Scientific American, 302(3):22–24, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
22.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-22.pdf.

Castelvecchi:2010:SBP

[61] Davide Castelvecchi. Supersonic bathtub physics. Scientific American, 302(3):23, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
23.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-23.pdf.

Nash:2010:EF

[62] Jim Nash. Easy flier. Scientific American, 302(3):24–25, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n3/full/scientificamerican0310-24.html;   http://www.nature.
com/scientificamerican/journal/v302/n3/pdf/scientificamerican0310-
24.pdf.

Castelvecchi:2010:ID

[63] Davide Castelvecchi. Invasion of the drones. Scientific American, 302(3):25–27, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
25.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-25.pdf.

Editors:2010:PSS

[64] The Editors. Perspectives: Start science sooner. Scientific American, 302(3):28, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
28.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-28.pdf.

Sachs:2010:SDB

[65] Jeffrey D. Sachs. Sustainable developments: Breaking the climate debate logjam. Scientific American, 302(3):30, March 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n3/full/scientificamerican0310-30.html;   http://www.nature.
com/scientificamerican/journal/v302/n3/pdf/scientificamerican0310-
30.pdf.

Shermer:2010:SSD

[66] Michael Shermer. Skeptic: Surviving death on Larry King Live. Scientific American, 302(3):32, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
32.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-32.pdf.

Krauss:2010:CME

[67] Lawrence M. Krauss. Critical mass: End-of-days danger. Scientific American, 302(3):34, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
34.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-34.pdf.

Mirsky:2010:AGSa

[68] Steve Mirsky. Anti gravity: Scooting toward oblivion. Scientific American, 302(3):84, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
84.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-84.pdf.

Wong:2010:RAA

[69] Kate Wong. Recommended: Amazing animals. Scientific American, 302(3):84, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-

84A.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-84A.pdf.

Lorenz:2010:PSM

[70] Ralph Lorenz and Christophe Sotin. Planetary science: The Moon that would be a planet. Scientific American, 302(3):36–43, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n3/full/scientificamerican0310-36.html;   http://www.nature.
com/scientificamerican/journal/v302/n3/pdf/scientificamerican0310-
36.pdf.

Raichle:2010:NBD

[71] Marcus E. Raichle. Neuroscience: The brain’s dark energy. Scientific American, 302(3):44–49, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
44.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-44.pdf.

Moyer:2010:EFF

[72] Michael Moyer. Energy: Fusion’s false dawn. Scientific American, 302(3):50–57, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
50.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-50.pdf.

Hazen:2010:GEM

[73] Robert M. Hazen. Geology: Evolution of minerals. Scientific American, 302(3):58–65, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
58.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-58.pdf.

Wang:2010:MTG

[74] Rui Wang. Medicine: Toxic gas, lifesaver. Scientific American, 302(3):66–71, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
66.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-66.pdf.

Catania:2010:ABW

[75] Kenneth Catania. Animal behavior: Worm charmers. Scientific American, 302(3):72–76, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n3/full/scientificamerican0310-
72.html;   http://www.nature.com/scientificamerican/journal/
v302/n3/pdf/scientificamerican0310-72.pdf.

Wullschleger:2010:ECC

[76] Stan D. Wullschleger and Maya Strahl. Environment: Climate change: A controlled experiment. Scientific American, 302(3):78–83, March 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n3/full/scientificamerican0310-78.html;   http://www.nature.
com/scientificamerican/journal/v302/n3/pdf/scientificamerican0310-
78.pdf.

DiChristina:2010:ESW

[77] Mariette DiChristina. From the Editor: Small world. Scientific American, 302(4):6, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
6.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-6.pdf.

Anonymous:2010:LHV

[78] Anonymous. Letters: Hydrothermal vents. Scientific American, 302(4): 8–10, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n4/full/scientificamerican0410-8.html;   http:
//www.nature.com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
8.pdf.

Schlenoff:2010:YAId

[79] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: Lethal dose. Scientific American, 302(4):12, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
12.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-12.pdf.

Moyer:2010:NSI

[80] Michael Moyer. News scan: Internet ideology war. Scientific American, 302(4):14–16, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
14.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-14.pdf.

Matson:2010:USN

[81] John Matson. Unfree spirit: NASA’s Mars Rover appears stuck for good. Scientific American, 302(4):16, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v302/n4/full/scientificamerican0410-
16a.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-16a.pdf.

Castelvecchi:2010:PSL

[82] Davide Castelvecchi. Problem solved, LOL. Scientific American, 302(4):16–18, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
16b.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-16b.pdf.

Storrs:2010:HMF

[83] Carina Storrs. Humans might have faced extinction. Scientific American, 302(4):18, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
18.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-18.pdf.

Wong:2010:S

[84] Kate Wong. Swept away. Scientific American, 302(4):20, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n4/full/scientificamerican0410-20.html;   http://www.nature.
com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
20.pdf.

Biello:2010:BB

[85] David Biello. Biofuel from bacteria. Scientific American, 302(4):21–22, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n4/full/scientificamerican0410-21.html;   http:
//www.nature.com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
21.pdf.

Biello:2010:HW

[86] David Biello. Higher watermarks. Scientific American, 302(4):22–24, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n4/full/scientificamerican0410-22.html;   http:
//www.nature.com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
22.pdf.

Moyer:2010:SS

[87] Melinda Wenner Moyer. Static over statins. Scientific American, 302(4):26–28, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
26.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-26.pdf.

Peeples:2010:MSS

[88] Lynne Peeples. Making scents of sounds. Scientific American, 302(4):28– 29, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n4/full/scientificamerican0410-28.html;   http:
//www.nature.com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
28.pdf.

Harmon:2010:CDT

[89] Katherine Harmon. Carcinogens detected from “thirdhand smoke”. Scientific American, 302(4):29, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
29.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-29.pdf.

Editors:2010:PCC

[90] The Editors. Perspectives: Chemical controls. Scientific American, 302(4):30, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-

30.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-30.pdf.

Sachs:2010:SDE

[91] Jeffrey D. Sachs. Sustainable developments: Expert systems fight poverty. Scientific American, 302(4):32, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v302/n4/full/scientificamerican0410-
32.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-32.pdf.

Shermer:2010:SSP

[92] Michael Shermer. Skeptic: The sensed-presence effect. Scientific American, 302(4):34, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
34.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-34.pdf.

Krauss:2010:CMDb

[93] Lawrence M. Krauss. Critical mass: Dark matters. Scientific American, 302(4):35, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
35A.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-35A.pdf.

Anonymous:2010:RIS

[94] Anonymous. Recommended: Imperiled species. Scientific American, 302(4):86, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
86.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-86.pdf.

Mirsky:2010:AGW

[95] Steve Mirsky. Anti gravity: Wonders of life. Scientific American, 302(4): 88, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n4/full/scientificamerican0410-88.html;   http:
//www.nature.com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
88.pdf.

Bell:2010:AWS

[96] Edward Bell. Astronomy: 8 wonders of the solar system. Scientific American, 302(4):36–43, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
36.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-36.pdf.

Insel:2010:NFC

[97] Thomas R. Insel. Neuroscience: Faulty circuits. Scientific American, 302(4):44–51, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
44.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-44.pdf.

Editors:2010:SRS

[98] The Editors. Special report: Sustainability: Living on a new Earth. Scientific American, 302(4):53, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
53.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-53.pdf.

Foley:2010:BHP

[99] Jonathan Foley. Boundaries for a healthy planet. Scientific American, 302(4):54–57, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
54.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-54.pdf.

Anonymous:2010:SET

[100] Anonymous. Solutions to environmental threats. Scientific American, 302(4):58–60, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
58.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-58.pdf.

McKibben:2010:EEB

[101] Bill McKibben. Exclusive excerpt: Breaking the growth habit. Scientific American, 302(4):61–65, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n4/full/scientificamerican0410-
61.html;   http://www.nature.com/scientificamerican/journal/
v302/n4/pdf/scientificamerican0410-61.pdf. See rebuttal [102].

Fischetti:2010:BMC

[102] Mark Fischetti. Bill McKibben, challenged: Is zero growth really necessary? Scientific American, 302(4):66–67, April 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n4/full/scientificamerican0410-66.html;   http://www.nature.
com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-

66.pdf. See [101].

Santina:2010:MRB

[103] Charles C. Della Santina. Medicine: Regaining balance with bionic ears. Scientific American, 302(4):68–71, April 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n4/full/scientificamerican0410-68.html;   http://www.nature.
com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
68.pdf.

Effros:2010:TRI

[104] Michelle Effros, Andrea Goldsmith, and Muriel Médard. Technology: The rise of instant wireless networks. Scientific American, 302(4):72–77, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n4/full/scientificamerican0410-72.html;   http:
//www.nature.com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
72.pdf.

Trappe:2010:BHL

[105] James M. Trappe and Andrew W. Claridge. Biology: The hidden life of truffles. Scientific American, 302(4):78–84, April 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n4/full/scientificamerican0410-78.html;   http://www.nature.
com/scientificamerican/journal/v302/n4/pdf/scientificamerican0410-
78.pdf.

DiChristina:2010:ESS

[106] Mariette DiChristina. From the Editor: Society and science. Scientific American, 302(5):6, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
6.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-6.pdf.

Anonymous:2010:LA

[107] Anonymous. Letters: Antimatter. Scientific American, 302(5):8–10, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n5/full/scientificamerican0510-8.html;   http:
//www.nature.com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
8.pdf.

Schlenoff:2010:YAIe

[108] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: Early experience. Scientific American, 302(5):12, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
12.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-12.pdf.

Soares:2010:NSP

[109] Christine Soares. News scan: Portrait in DNA. Scientific American, 302(5):14–17, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
14.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-14.pdf.

Matson:2010:S

[110] John Matson. Shields up. Scientific American, 302(5):17–18, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n5/full/scientificamerican0510-17.html;   http://www.nature.
com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
17.pdf.

Harmon:2010:CSS

[111] Katherine Harmon. Cold snap: A secret of the chameleon’s tongue revealed. Scientific American, 302(5):18, May 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n5/full/scientificamerican0510-18.html;   http://www.nature.
com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
18.pdf.

Moyer:2010:CAC

[112] Melinda Wenner Moyer. Carbs against cardio. Scientific American, 302(5):19–21, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
19.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-19.pdf.

Moisse:2010:OEK

[113] Kate Moisse. One’s enough: Kidney donors live just as long as nondonors. Scientific American, 302(5):20, May 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n5/full/scientificamerican0510-20.html;   http://www.nature.
com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
20.pdf.

Harmon:2010:GG

[114] Katherine Harmon. Genetics in the gut. Scientific American, 302(5):22– 24, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n5/full/scientificamerican0510-22.html;   http:
//www.nature.com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
22.pdf.

Biello:2010:SCW

[115] David Biello. Sex-changing weed killer. Scientific American, 302(5):24– 27, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n5/full/scientificamerican0510-24.html;   http:
//www.nature.com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
24.pdf.

Biello:2010:IEP

[116] David Biello. IPCC errors prompt a review on climate science data. Scientific American, 302(5):26–27, May 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n5/full/scientificamerican0510-26.html;   http://www.nature.
com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
26.pdf.

Tennesen:2010:MFF

[117] Michael Tennesen. More food from fungi? Scientific American, 302(5): 27–28, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n5/full/scientificamerican0510-27.html;   http:
//www.nature.com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
27.pdf.

Greenmeier:2010:PPO

[118] Larry Greenmeier. PET project: Organic catalysts could increase plastics recycling. Scientific American, 302(5):28, May 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n5/full/scientificamerican0510-28.html;   http://www.nature.
com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
28.pdf.

Editors:2010:PUO

[119] The Editors. Perspectives: Underage, overweight. Scientific American, 302(5):30, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
30.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-30.pdf.

Sachs:2010:SDFb

[120] Jeffrey D. Sachs. Sustainable developments: Flying blind in policy reforms. Scientific American, 302(5):32, May 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n5/full/scientificamerican0510-32.html;   http://www.nature.
com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
32.pdf.

Shermer:2010:SDS

[121] Michael Shermer. Skeptic: Doing science in the past. Scientific American, 302(5):34, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
34.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-34.pdf.

Krauss:2010:CMH

[122] Lawrence M. Krauss. Critical mass: Human uniqueness and the future. Scientific American, 302(5):36, May 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n5/full/scientificamerican0510-36.html;   http://www.nature.
com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
36.pdf.

Anonymous:2010:RDW

[123] Anonymous. Recommended: Decoding weather. Scientific American, 302(5):86, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
86.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-86.pdf.

Mirsky:2010:AGCb

[124] Steve Mirsky. Anti gravity: 140-character study. Scientific American, 302 (5):88, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n5/full/scientificamerican0510-88.html;   http:
//www.nature.com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
88.pdf.

Gelmini:2010:ATN

[125] Graciela B. Gelmini, Alexander Kusenko, and Thomas J. Weiler. Astrophysics: Through neutrino eyes. Scientific American, 302(5):38–45, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n5/full/scientificamerican0510-38.html;   http:
//www.nature.com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
38.pdf.

Hochedlinger:2010:MYI

[126] Konrad Hochedlinger. Medicine: Your inner healers. Scientific American, 302(5):46–53, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
46.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-46.pdf.

Brown:2010:TRR

[127] Stuart F. Brown. Transportation: Revolutionary rail. Scientific American, 302(5):54–59, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
54.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-54.pdf.

deGelder:2010:NUS

[128] Beatrice de Gelder. Neuroscience: Uncanny sight in the blind. Scientific American, 302(5):60–65, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
60.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-60.pdf.

Sturm:2010:CCA

[129] Matthew Sturm. Climate change: Arctic plants feel the heat. Scientific American, 302(5):66–73, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
66.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-66.pdf.

May:2010:IBL

[130] Mike May. Innovations: A better lens on disease. Scientific American, 302(5):74–77, May 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n5/full/scientificamerican0510-
74.html;   http://www.nature.com/scientificamerican/journal/
v302/n5/pdf/scientificamerican0510-74.pdf.

Nassar:2010:FSB

[131] Nagib Nassar and Rodomiro Ortiz. Food science: Breeding cassava to feed the poor. Scientific American, 302(5):78–84, May 2010.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n5/full/scientificamerican0510-78.html;   http://www.nature.
com/scientificamerican/journal/v302/n5/pdf/scientificamerican0510-
78.pdf.

DiChristina:2010:ETF

[132] Mariette DiChristina. From the Editor: Think forward. Scientific American, 302(6):4, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
4.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-4.pdf.

Anonymous:2010:LF

[133] Anonymous. Letters: Fertilizers. Scientific American, 302(6):6–8, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-6.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
6.pdf.

Schlenoff:2010:YAIf

[134] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 302(6):10, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-10.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
10.pdf.

Wong:2010:NSF

[135] Kate Wong. News scan: Fossils of our family. Scientific American, 302(6):12–14, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
12.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-12.pdf.

Musser:2010:STF

[136] George Musser. A simple twist of fate. Scientific American, 302(6):14– 17, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-14.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
14.pdf.

Castelvecchi:2010:QMP

[137] Davide Castelvecchi. “quantum microphone” puts visible object in two places at once. Scientific American, 302(6):16, June 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n6/full/scientificamerican0610-16.html;   http://www.nature.
com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
16.pdf.

Choi:2010:EE

[138] Charles Q. Choi. Expert education. Scientific American, 302(6):17–18, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-17.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
17.pdf.

Wapner:2010:AA

[139] Jessica Wapner. Adoption agents. Scientific American, 302(6):19–20, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-19.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
19.pdf.

Roehr:2010:BWV

[140] Bob Roehr. Bad wraps on viruses. Scientific American, 302(6):21, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-21.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
21.pdf.

Fischetti:2010:TA

[141] Mark Fischetti. Traffic avoided. Scientific American, 302(6):22–23, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-22.html;   http:

//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
22.pdf.

Biello:2010:SOT

[142] David Biello. Stringing offshore turbines for uninterrupted power. Scientific American, 302(6):23, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
23.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-23.pdf.

Castelvecchi:2010:LMN

[143] Davide Castelvecchi. Look, Ma — no junctions! Scientific American, 302(6):26, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
26a.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-26a.pdf.

Greenemeier:2010:ALT

[144] Larry Greenemeier. Acoustic lens turns sound into sonic bullets. Scientific American, 302(6):26, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
26b.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-26b.pdf.

Editors:2010:PTV

[145] The Editors. Perspectives: Terror in a vial. Scientific American, 302(6): 28, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-28.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
28.pdf.

Sachs:2010:SDMa

[146] Jeffrey D. Sachs. Sustainable developments: Millennium development goals at 10. Scientific American, 302(6):30, June 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n6/full/scientificamerican0610-30.html;   http://www.nature.
com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
30.pdf.

Shermer:2010:SWI

[147] Michael Shermer. Skeptic: When ideas have sex. Scientific American, 302(6):32, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
32.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-32.pdf.

Krauss:2010:CMW

[148] Lawrence M. Krauss. Critical mass: Why I love neutrinos. Scientific American, 302(6):34, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
34.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-34.pdf.

Anonymous:2010:RIA

[149] Anonymous. Recommended: Industrious ants. Scientific American, 302(6):90, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
90.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-90.pdf.

Mirsky:2010:AGP

[150] Steve Mirsky. Anti gravity: Presidential harrisment. Scientific American, 302(6):92, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
92.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-92.pdf.

Editors:2010:WFH

[151] The Editors. What the future holds: 12 events that will change everything. Scientific American, 302(6):36, June 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n6/full/scientificamerican0610-36a.html;   http://www.nature.
com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
36a.pdf.

Choi:2010:CH

[152] Charles Q. Choi. Cloning of a human. Scientific American, 302(6):36– 38, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-36b.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
36b.pdf.

Musser:2010:ED

[153] George Musser. Extra dimensions. Scientific American, 302(6):39, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-39.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
39.pdf.

Matson:2010:EI

[154] John Matson. Extraterrestrial intelligence. Scientific American, 302(6): 40, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-40a.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
40a.pdf.

Yam:2010:NE

[155] Philip Yam. Nuclear exchange. Scientific American, 302(6):40, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-40b.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
40b.pdf.

Biello:2010:CL

[156] David Biello. Creation of life. Scientific American, 302(6):42, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n6/full/scientificamerican0610-42.html;   http://www.nature.
com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
42.pdf.

Moyer:2010:RTS

[157] Michael Moyer. Room-temperature superconductors. Scientific American, 302(6):43, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
43.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-43.pdf.

Greenemeier:2010:MSA

[158] Larry Greenemeier. Machine self-awareness. Scientific American, 302(6):44–45, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
44.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-44.pdf.

Biello:2010:PM

[159] David Biello. Polar meltdown. Scientific American, 302(6):45, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n6/full/scientificamerican0610-45.html;   http://www.nature.
com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
45.pdf.

Harmon:2010:PE

[160] Katherine Harmon. Pacific earthquake. Scientific American, 302(6):46– 47, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-46.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
46.pdf.

Moyer:2010:FE

[161] Michael Moyer. Fusion energy. Scientific American, 302(6):47, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n6/full/scientificamerican0610-47.html;   http://www.nature.
com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
47.pdf.

Lloyd:2010:AC

[162] Robin Lloyd. Asteroid collision. Scientific American, 302(6):48, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-48a.html;   http:

//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
48a.pdf.

Harmon:2010:DP

[163] Katherine Harmon. Deadly pandemic. Scientific American, 302(6):48, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v302/n6/full/scientificamerican0610-48b.html;   http:
//www.nature.com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
48b.pdf.

Stix:2010:MAF

[164] Gary Stix. Medicine: Alzheimer’s: Forestalling the darkness. Scientific American, 302(6):50–57, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
50.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-50.pdf.

Callender:2010:PTI

[165] Craig Callender. Physics: Is time an illusion? Scientific American, 302(6):58–65, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
58.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-58.pdf.

Lackner:2010:EWC

[166] Klaus S. Lackner. Environment: Washing carbon out of the air. Scientific American, 302(6):66–71, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
66.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-66.pdf.

Anonymous:2010:ADN

[167] Anonymous. Archaeology: Did Neandertals think like us? Scientific American, 302(6):72–75, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
72.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-72.pdf.

Hirose:2010:ESE

[168] Kei Hirose. Earth science: The Earth’s missing ingredient. Scientific American, 302(6):76–83, June 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v302/n6/full/scientificamerican0610-
76.html;   http://www.nature.com/scientificamerican/journal/
v302/n6/pdf/scientificamerican0610-76.pdf.

Coleman:2010:SFB

[169] Ken Coleman and Raymond A. Zilinskas. Security: Fake botox, real threat. Scientific American, 302(6):84–89, June 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v302/
n6/full/scientificamerican0610-84.html;   http://www.nature.
com/scientificamerican/journal/v302/n6/pdf/scientificamerican0610-
84.pdf.

DiChristina:2010:EGA

[170] Mariette DiChristina. From the Editor: A global affair. Scientific American, 303(1):8, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
8.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-8.pdf.

Anonymous:2010:LMW

[171] Anonymous. Letters: Math wars. Scientific American, 303(1):10–12, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-10.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
10.pdf.

Schlenoff:2010:YAIg

[172] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: Live births. Scientific American, 303(1):14, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
14.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-14.pdf.

Biello:2010:NSL

[173] David Biello. News scan: Lasting menace. Scientific American, 303(1): 16–18, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-16.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
16.pdf.

Wong:2010:OIN

[174] Kate Wong. Our inner Neandertal. Scientific American, 303(1):18– 20, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-18.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
18.pdf.

Matson:2010:DRS

[175] John Matson. Damp rocks from space. Scientific American, 303(1): 20, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-20a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
20a.pdf.

Wickelgren:2010:UTW

[176] Ingrid Wickelgren. Under threat, women bond and men withdraw. Scientific American, 303(1):20, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
20b.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-20b.pdf.

Harmon:2010:MG

[177] Katherine Harmon. Microscopic giants. Scientific American, 303(1): 22, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-22a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
22a.pdf.

Matson:2010:ECS

[178] John Matson. Eye-candy solar science. Scientific American, 303(1): 22, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-22b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
22b.pdf.

Soares:2010:PAB

[179] Christine Soares. Putting addiction to bed. Scientific American, 303(1): 23–24, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-23.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
23.pdf.

Moyer:2010:HH

[180] Melinda Wenner Moyer. Heartburn headache. Scientific American, 303(1):24–26, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
24.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-24.pdf.

Moyer:2010:MC

[181] Michael Moyer. Manipulation of the crowd. Scientific American, 303(1): 26–28, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-26.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
26.pdf.

Greenemeier:2010:PSP

[182] Larry Greenemeier. Patent still pending. Scientific American, 303(1): 28, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-28.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
28.pdf.

Editors:2010:PTT

[183] The Editors. Perspectives: Terminate the terminators. Scientific American, 303(1):30, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-

30.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-30.pdf.

Sachs:2010:SDG

[184] Jeffrey D. Sachs. Sustainable developments: Get serious about budget deficits. Scientific American, 303(1):32, July 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n1/full/scientificamerican0710-32.html;   http://www.nature.
com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
32.pdf.

Shermer:2010:SWS

[185] Michael Shermer. Skeptic: When scientists sin. Scientific American, 303 (1):34, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-34.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
34.pdf.

Krauss:2010:CMN

[186] Lawrence M. Krauss. Critical mass: No country is an island. Scientific American, 303(1):36, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
36.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-36.pdf.

Anonymous:2010:RES

[187] Anonymous. Recommendations: Earth sans ice caps. Scientific American, 303(1):86, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
86.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-86.pdf.

Mirsky:2010:AGA

[188] Steve Mirsky. Anti gravity: An apple a day for a week. Scientific American, 303(1):88, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
88.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-88.pdf.

Davis:2010:CUL

[189] Tamara M. Davis. Cosmology: Is the universe leaking energy? Scientific American, 303(1):38–47, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
38.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-38.pdf.

Morrow:2010:MDD

[190] Matthew P. Morrow and David B. Weiner. Medicine: DNA drugs come of age. Scientific American, 303(1):48–53, July 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n1/full/scientificamerican0710-48.html;   http://www.nature.
com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
48.pdf.

Moyer:2010:TDT

[191] Michael Moyer. Transportation: The dirty truth about plugin hybrids. Scientific American, 303(1):54–55, July 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n1/full/scientificamerican0710-54.html;   http://www.nature.
com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
54.pdf.

Singer:2010:RWM

[192] P. W. Singer. Robotics: War of the machines. Scientific American, 303(1):56–63, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
56.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-56.pdf.

Little:2010:ECE

[193] Jane Braxton Little. Energy: Clean energy from filthy water. Scientific American, 303(1):64–69, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
64.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-64.pdf.

Dyke:2010:EWV

[194] Gareth Dyke. Evolution: Winged victory. Scientific American, 303(1): 70–75, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n1/full/scientificamerican0710-70.html;   http:
//www.nature.com/scientificamerican/journal/v303/n1/pdf/scientificamerican0710-
70.pdf.

Gopnik:2010:PHB

[195] Alison Gopnik. Psychology: How babies think. Scientific American, 303(1):76–81, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
76.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-76.pdf.

Fischetti:2010:EDC

[196] Mark Fischetti. Environment: The drillers are coming. Scientific American, 303(1):82–85, July 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n1/full/scientificamerican0710-
82.html;   http://www.nature.com/scientificamerican/journal/
v303/n1/pdf/scientificamerican0710-82.pdf.

DiChristina:2010:ECS

[197] Mariette DiChristina. From the Editor: Celebrating science. Scientific American, 303(2):6, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
6.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-6.pdf.

Anonymous:2010:LG

[198] Anonymous. Letters: Growth. Scientific American, 303(2):8–10, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-8.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
8.pdf.

Schlenoff:2010:YAIh

[199] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: Vertical takeoff. Scientific American, 303(2):12–13, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
12.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-12.pdf.

Biello:2010:NSB

[200] David Biello. News scan: Biological breakdown. Scientific American, 303(2):14–17, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
14.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-14.pdf.

Biello:2010:TL

[201] David Biello and Katherine Harmon. Tools for life. Scientific American, 303(2):17–18, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
17.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-17.pdf.

Benderly:2010:DSL

[202] Beryl Lieff Benderly. Danger in school labs. Scientific American, 303(2):18–20, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
18.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-18.pdf.

Anonymous:2010:MG

[203] Anonymous. Martin Gardner, 1914–2010. Scientific American, 303(2): 20, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-20a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
20a.pdf.

Matson:2010:EQS

[204] John Matson. An extra quiet Sun. Scientific American, 303(2):20–21, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-20b.html;   http:

//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
20b.pdf.

Jabr:2010:NSR

[205] Ferris Jabr. Night sight: Rapid eye movements seem to scan the actions in our dreams. Scientific American, 303(2):22, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n2/full/scientificamerican0810-22a.html;   http://www.nature.
com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
22a.pdf.

Minkel:2010:HA

[206] JR Minkel. Hereditary acquisitions. Scientific American, 303(2):22–24, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-22b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
22b.pdf.

Harmon:2010:WA

[207] Katherine Harmon. Worts and all. Scientific American, 303(2):24–26, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-24.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
24.pdf.

Fischetti:2010:CUC

[208] Mark Fischetti. Charge under control. Scientific American, 303(2):26– 27, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-26.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
26.pdf.

Greenemeier:2010:RTD

[209] Larry Greenemeier. Robot test drive. Scientific American, 303(2):28, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-28a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
28a.pdf.

Matson:2010:NPU

[210] John Matson. NASA’s plan to use commercial rockets lifts off. Scientific American, 303(2):28, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
28b.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-28b.pdf.

Editors:2010:PCT

[211] The Editors. Perspectives: Catastrophic thinking. Scientific American, 303(2):30, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
30.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-30.pdf.

Sachs:2010:SDMb

[212] Jeffrey D. Sachs. Sustainable developments: Market reforms, 20 years later. Scientific American, 303(2):32, August 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n2/full/scientificamerican0810-32.html;   http://www.nature.
com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
32.pdf.

Shermer:2010:SON

[213] Michael Shermer. Skeptic: Our Neandertal brethren. Scientific American, 303(2):34, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
34.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-34.pdf.

Krauss:2010:CMF

[214] Lawrence M. Krauss. Critical mass: Faith and foolishness. Scientific American, 303(2):36, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
36.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-36.pdf.

Anonymous:2010:RRA

[215] Anonymous. Reviews and recommendations: Arctic change. Scientific American, 303(2):90, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
90.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-90.pdf.

Mirsky:2010:AGT

[216] Steve Mirsky. Anti gravity: Take me out of the ball game. Scientific American, 303(2):92, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
92.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-92.pdf.

Sasselov:2010:SPW

[217] Dimitar D. Sasselov and Diana Valencia. Space: Planets we could call home. Scientific American, 303(2):38–45, August 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n2/full/scientificamerican0810-38.html;   http://www.nature.
com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
38.pdf.

Anonymous:2010:SRO

[218] Anonymous. Special report: Origins. Scientific American, 303(2):46, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-46.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
46.pdf.

Borrell:2010:AF

[219] Brendan Borrell. All in the family. Scientific American, 303(2):47, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-47a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
47a.pdf.

Moyer:2010:CS

[220] Melinda Wenner Moyer. Cheese story. Scientific American, 303(2):47, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-47b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
47b.pdf.

May:2010:EP

[221] Mike May. Electronic pathogens. Scientific American, 303(2):47, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-47c.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
47c.pdf.

Borrell:2010:BMM

[222] Brendan Borrell. Before Mickey Mouse. Scientific American, 303(2):48, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-48a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
48a.pdf.

Borrell:2010:SRN

[223] Brendan Borrell. Is sex really necessary? Scientific American, 303(2):48– 49, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-48b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
48b.pdf.

May:2010:PT

[224] Mike May. On the parasite’s trail. Scientific American, 303(2):48–49, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-48c.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
48c.pdf.

May:2010:SCB

[225] Mike May. Snap, crackle, bang. Scientific American, 303(2):49, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-49.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
49.pdf.

May:2010:TF

[226] Mike May. Thorny fence. Scientific American, 303(2):50, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n2/full/scientificamerican0810-50a.html;   http://www.nature.
com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
50a.pdf.

May:2010:S

[227] Mike May. Scrubs. Scientific American, 303(2):50, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n2/full/scientificamerican0810-50b.html;   http://www.nature.
com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
50b.pdf.

Moyer:2010:MA

[228] Melinda Wenner Moyer. Moral animal. Scientific American, 303(2):51, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-51a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
51a.pdf.

Moyer:2010:UB

[229] Melinda Wenner Moyer. Urban bug. Scientific American, 303(2):51, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-51b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
51b.pdf.

May:2010:FLE

[230] Mike May. Former life of the electric car. Scientific American, 303(2):51, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-51c.html;   http:

//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
51c.pdf.

Borrell:2010:FH

[231] Brendan Borrell. The first Humvee. Scientific American, 303(2):52, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-52a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
52a.pdf.

Borrell:2010:GT

[232] Brendan Borrell. Gravity’s tug. Scientific American, 303(2):52–53, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-52b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
52b.pdf.

Moyer:2010:RC

[233] Melinda Wenner Moyer. Rainbow cells. Scientific American, 303(2):53, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-53a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
53a.pdf.

Borrell:2010:Z

[234] Brendan Borrell. Zero. Scientific American, 303(2):53, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n2/full/scientificamerican0810-53b.html;   http://www.nature.
com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
53b.pdf.

May:2010:NN

[235] Mike May. Noodling the noodles. Scientific American, 303(2):53, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n2/full/scientificamerican0810-53c.html;   http:
//www.nature.com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
53c.pdf.

Marean:2010:WSS

[236] Curtis W. Marean. When the sea saved humanity. Scientific American, 303(2):54–61, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
54.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-54.pdf.

Dario:2010:MRP

[237] Paolo Dario and Arianna Menciassi. Medicine: Robot pills. Scientific American, 303(2):62–65, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
62.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-62.pdf.

Hardt:2010:ETO

[238] Marah J. Hardt and Carl Safina. Environment: Threatening ocean life from the inside out. Scientific American, 303(2):66–73, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n2/full/scientificamerican0810-66.html;   http://www.nature.
com/scientificamerican/journal/v303/n2/pdf/scientificamerican0810-
66.pdf.

Zewail:2010:IFI

[239] Ahmed H. Zewail. Imaging: Filming the invisible in 4-D. Scientific American, 303(2):74–81, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
74.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-74.pdf.

Villasenor:2010:THY

[240] John Villasenor. Technology: The hacker in your hardware. Scientific American, 303(2):82–88, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
82.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-82.pdf.

Ackerman:2010:OLP

[241] Jennifer Ackerman. Object lesson: Plastic surf. Scientific American, 303(2):88–89, August 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n2/full/scientificamerican0810-
88.html;   http://www.nature.com/scientificamerican/journal/
v303/n2/pdf/scientificamerican0810-88.pdf.

DiChristina:2010:ESE

[242] Mariette DiChristina. From the Editor: Start of the end. Scientific American, 303(3):6, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
6.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-6.pdf.

Anonymous:2010:LP

[243] Anonymous. Letters: Pathology. Scientific American, 303(3):8–10, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n3/full/scientificamerican0910-8.html;   http:
//www.nature.com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
8.pdf.

Schlenoff:2010:YAIi

[244] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 303(3):12, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n3/full/scientificamerican0910-12.html;   http:
//www.nature.com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
12.pdf.

Merali:2010:NSR

[245] Zeeya Merali. News scan: Rummaging for a final theory. Scientific American, 303(3):14–17, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
14.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-14.pdf.

Matson:2010:LPL

[246] John Matson. Lunar pencil lead. Scientific American, 303(3):17, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n3/full/scientificamerican0910-17.html;   http:
//www.nature.com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
17.pdf.

Lehrman:2010:UE

[247] Sally Lehrman. Undifferentiated ethics. Scientific American, 303 (3):18–20, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
18.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-18.pdf.

Westly:2010:FA

[248] Erica Westly. Fat attack. Scientific American, 303(3):20–22, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n3/full/scientificamerican0910-20.html;   http:
//www.nature.com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
20.pdf.

Stix:2010:SA

[249] Gary Stix. Social analgesics. Scientific American, 303(3):22–23, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n3/full/scientificamerican0910-22.html;   http:
//www.nature.com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
22.pdf.

Tennesen:2010:SS

[250] Michael Tennesen. Sour showers. Scientific American, 303(3):23–24, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n3/full/scientificamerican0910-23.html;   http:
//www.nature.com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
23.pdf.

Biello:2010:DD

[251] David Biello. Doubts on dispersants. Scientific American, 303(3):24–26, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n3/full/scientificamerican0910-24.html;   http:
//www.nature.com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
24.pdf.

Castelvecchi:2010:QLS

[252] Davide Castelvecchi. Quantum light switch. Scientific American, 303(3):26–28, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
26.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-26.pdf.

Matson:2010:OSF

[253] John Matson. Origami sheets that fold themselves. Scientific American, 303(3):28, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
28.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-28.pdf.

Editors:2010:PSO

[254] The Editors. Perspectives: Snake oil in the supermarket. Scientific American, 303(3):30, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
30.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-30.pdf.

Sachs:2010:SDD

[255] Jeffrey D. Sachs. Sustainable developments: The deepening crisis. Scientific American, 303(3):32, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v303/n3/full/scientificamerican0910-
32.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-32.pdf.

Shermer:2010:SDL

[256] Michael Shermer. Skeptic: Democracy’s laboratory. Scientific American, 303(3):34, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-

34.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-34.pdf.

Krauss:2010:CMY

[257] Lawrence M. Krauss. Critical mass: A year of living dangerously. Scientific American, 303(3):36, September 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n3/full/scientificamerican0910-36.html;   http://www.nature.
com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
36.pdf.

Anonymous:2010:AEH

[258] Anonymous. Ask the experts: How does geothermal drilling trigger earthquakes? Scientific American, 303(3):100, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n3/full/scientificamerican0910-100.html;   http://www.nature.
com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
100.pdf.

Anonymous:2010:RRE

[259] Anonymous. Reviews and recommendations: Extreme astronomy. Scientific American, 303(3):102, September 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n3/full/scientificamerican0910-102.html;   http://www.nature.
com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
102.pdf.

Mirsky:2010:AGSb

[260] Steve Mirsky. Anti gravity: Shaky grounds. Scientific American, 303(3):104, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
104.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-104.pdf.

Moyer:2010:IEF

[261] Michael Moyer. Introduction: Eternal fascinations with the end. Scientific American, 303(3):38–41, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v303/n3/full/scientificamerican0910-

38.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-38.pdf.

Kirkwood:2010:MWC

[262] Thomas Kirkwood. Medicine: Why can’t we live forever? Scientific American, 303(3):42–49, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v303/n3/full/scientificamerican0910-
42.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-42.pdf.

Henig:2010:BWD

[263] Robin Marantz Henig. Bioethics: When does life belong to the living? Scientific American, 303(3):50–55, September 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n3/full/scientificamerican0910-50.html;   http://www.nature.
com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
50.pdf.

Vass:2010:FDD

[264] Arpad A. Vass. Forensics: Dust to dust. Scientific American, 303(3):56–59, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
56.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-56.pdf.

Davis:2010:ALT

[265] Wade Davis. Anthropology: Last of their kind. Scientific American, 303(3):60–67, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
60.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-60.pdf.

Anonymous:2010:TGR

[266] Anonymous. Technology: Good riddance. Scientific American, 303(3):68–73, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
68.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-68.pdf.

Moyer:2010:EHM

[267] Michael Moyer. Environment: How much is left? Scientific American, 303(3):74–81, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
74.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-74.pdf.

Matson:2010:RAL

[268] John Matson and John Pavlus. Risk analysis: Laying odds on the apocalypse. Scientific American, 303(3):82–83, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n3/full/scientificamerican0910-82.html;   http://www.nature.
com/scientificamerican/journal/v303/n3/pdf/scientificamerican0910-
82.pdf.

Musser:2010:CCT

[269] George Musser. Cosmology: Could time end? Scientific American, 303(3):84–91, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
84.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-84.pdf.

Anonymous:2010:RWC

[270] Anonymous. Restart: What comes next. Scientific American, 303(3):92–97, September 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n3/full/scientificamerican0910-
92.html;   http://www.nature.com/scientificamerican/journal/
v303/n3/pdf/scientificamerican0910-92.pdf.

DiChristina:2010:ENY

[271] Mariette DiChristina. From the Editor: New yet familiar. Scientific American, 303(4):8, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
8.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-8.pdf.

Anonymous:2010:Lb

[272] Anonymous. Letters. Scientific American, 303(4):10–12, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n4/full/scientificamerican1010-10.html;   http://www.nature.
com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
10.pdf.

Schlenoff:2010:YAIj

[273] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 303(4):102, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-102.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
102.pdf.

Editors:2010:SAC

[274] The Editors. Science agenda: Competition and the Internet. Scientific American, 303(4):14, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
14.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-14.pdf.

Hathaway:2010:FPH

[275] Melissa Hathaway. Forum: Power hackers. Scientific American, 303(4): 16, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-16.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
16.pdf.

Moyer:2010:ABB

[276] Michael Moyer. Advances: The biggest bang theory. Scientific American, 303(4):18, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
18.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-18.pdf.

Wong:2010:FB

[277] Kate Wong. The first butchers. Scientific American, 303(4):21, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-21a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
21a.pdf.

Moyer:2010:DC

[278] Michael Moyer. Death and chocolate. Scientific American, 303(4):21, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-21b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
21b.pdf.

Anonymous:2010:Qa

[279] Anonymous. Quotable. Scientific American, 303(4):21, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n4/full/scientificamerican1010-21c.html;   http://www.nature.
com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
21c.pdf.

Zeliadt:2010:THT

[280] Nicholette Zeliadt. Trash is her treasure. Scientific American, 303(4):22, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-22a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
22a.pdf.

Anonymous:2010:SI

[281] Anonymous. Science index. Scientific American, 303(4):22, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-22b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
22b.pdf.

Castelvecchi:2010:JHS

[282] Davide Castelvecchi. Just how small is the proton? Scientific American, 303(4):24, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
24a.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-24a.pdf.

Jabr:2010:PS

[283] Ferris Jabr. The power of sniff. Scientific American, 303(4):24, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-24b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
24b.pdf.

Moyer:2010:VK

[284] Melinda Wenner Moyer. A view to a kill. Scientific American, 303(4):26, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-26.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
26.pdf.

Kuchment:2010:WIa

[285] Anna Kuchment. What is it? Scientific American, 303(4):27, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-27.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
27.pdf.

Jabr:2010:FS

[286] Ferris Jabr. Flight of the squid. Scientific American, 303(4):28, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-28.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
28.pdf.

Choi:2010:TWT

[287] Charles Q. Choi. There’s wisdom in those tweets. Scientific American, 303(4):29, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-

29a.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-29a.pdf.

Kuchment:2010:PWa

[288] Anna Kuchment. Patent watch. Scientific American, 303(4):29, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-29b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
29b.pdf.

Choi:2010:GGJ

[289] Charles Q. Choi. Getting GPS out of a jam. Scientific American, 303(4): 30, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-30.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
30.pdf.

Gelman:2010:WSN

[290] Andrew Gelman. When small numbers lead to big errors. Scientific American, 303(4):31, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
31a.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-31a.pdf.

Greenemeier:2010:WA

[291] Larry Greenemeier. What’s ahead. Scientific American, 303(4):31, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-31b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
31b.pdf.

Stix:2010:CC

[292] Gary Stix. Craving a cure. Scientific American, 303(4):32, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n4/full/scientificamerican1010-32.html;   http://www.nature.
com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
32.pdf.

Gorman:2010:SHC

[293] Christine Gorman. The science of health: Closing the health gap. Scientific American, 303(4):34–36, October 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n4/full/scientificamerican1010-34.html;   http://www.nature.
com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
34.pdf.

Pogue:2010:TQT

[294] David Pogue. TechnoFiles: Question time. Scientific American, 303(4): 38, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n4/full/scientificamerican1010-38.html;   http:
//www.nature.com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
38.pdf.

Wong:2010:RRRa

[295] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 303(4):96–97, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
96.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-96.pdf.

Shermer:2010:SCYb

[296] Michael Shermer. Skeptic: Can you hear me now? Scientific American, 303(4):98, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
98.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-98.pdf.

Mirsky:2010:AGR

[297] Steve Mirsky. Anti gravity: Real skyscrapers. Scientific American, 303(4):100, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
100.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-100.pdf.

Fischetti:2010:GSD

[298] Mark Fischetti. Graphic science: Dying to eat. Scientific American, 303(4):104, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
104.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-104.pdf.

Pritchard:2010:EHW

[299] Jonathan K. Pritchard. Evolution: How we are evolving. Scientific American, 303(4):40–47, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
40.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-40.pdf.

Irion:2010:AOO

[300] Robert Irion. Astronomy: Origami observatory. Scientific American, 303(4):48–55, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
48.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-48.pdf.

Anonymous:2010:EPS

[301] Anonymous. Exclusive poll: In science we trust. Scientific American, 303(4):56–59, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
56.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-56.pdf.

Hall:2010:MRP

[302] Stephen S. Hall. Medicine: Revolution postponed. Scientific American, 303(4):60–67, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
60.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-60.pdf.

Hawking:2010:PET

[303] Stephen Hawking and Leonard Mlodinow. Physics: The (elusive) theory of everything. Scientific American, 303(4):68–71, October 2010.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n4/full/scientificamerican1010-68.html;   http://www.nature.
com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
68.pdf.

Anderson:2010:RRG

[304] Michael Anderson and Susan Leigh Anderson. Robotics: Robot be good. Scientific American, 303(4):72–77, October 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n4/full/scientificamerican1010-72.html;   http://www.nature.
com/scientificamerican/journal/v303/n4/pdf/scientificamerican1010-
72.pdf.

Grasso:2010:BSS

[305] Frank W. Grasso. Biology: Sensational sucker. Scientific American, 303(4):78–79, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
78.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-78.pdf.

Shute:2010:MDA

[306] Nancy Shute. Mind: Desperate for an autism cure. Scientific American, 303(4):80–85, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
80.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-80.pdf.

Regalado:2010:ERL

[307] Antonio Regalado. Energy: Reinventing the leaf. Scientific American, 303(4):86–89, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
86.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-86.pdf.

Moyer:2010:ISD

[308] Michael Moyer. Information science: Digitizer in chief. Scientific American, 303(4):90–94, October 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n4/full/scientificamerican1010-
90.html;   http://www.nature.com/scientificamerican/journal/
v303/n4/pdf/scientificamerican1010-90.pdf.

DiChristina:2010:ETM

[309] Mariette DiChristina. From the Editor: Teachable moment. Scientific American, 303(5):10, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
10.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-10.pdf.

Anonymous:2010:Lc

[310] Anonymous. Letters. Scientific American, 303(5):12–14, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n5/full/scientificamerican1110-12.html;   http://www.nature.
com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
12.pdf.

Schlenoff:2010:YAIk

[311] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 303(5):90, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-90.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
90.pdf.

Anonymous:2010:SAH

[312] Anonymous. Science agenda: Hearing the music, honing the mind. Scientific American, 303(5):16, November 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n5/full/scientificamerican1110-16.html;   http://www.nature.
com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
16.pdf.

Lilienfeld:2010:FFF

[313] Scott O. Lilienfeld. Forum: Fudge factor. Scientific American, 303(5):18, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-18.html;   http:

//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
18.pdf.

Moyer:2010:AWS

[314] Michael Moyer. Advances: Window shopping for electric cars. Scientific American, 303(5):20, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
20.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-20.pdf.

Interlandi:2010:FKA

[315] Jeneen Interlandi. Finding a killer’s Achilles’ heel. Scientific American, 303(5):21, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
21a.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-21a.pdf.

Anonymous:2010:Qb

[316] Anonymous. Quotable. Scientific American, 303(5):21, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n5/full/scientificamerican1110-21b.html;   http://www.nature.
com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
21b.pdf.

Castelvecchi:2010:HVG

[317] Davide Castelvecchi. Hawking vs. God. Scientific American, 303(5):21, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-21c.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
21c.pdf.

Choi:2010:BS

[318] Charles Q. Choi. Bugs in space. Scientific American, 303(5):22, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-22.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
22.pdf.

Piore:2010:FP

[319] Adam Piore. Fit for a princess. Scientific American, 303(5):24, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-24.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
24.pdf.

Moyer:2010:NGF

[320] Melinda Wenner Moyer. Nice germs finish last. Scientific American, 303(5):25, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
25a.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-25a.pdf.

Anonymous:2010:Qc

[321] Anonymous. Quotable. Scientific American, 303(5):25, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n5/full/scientificamerican1110-25b.html;   http://www.nature.
com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
25b.pdf.

Wong:2010:ILS

[322] Kate Wong. The Iceman’s last stand. Scientific American, 303(5):26, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-26.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
26.pdf.

Interlandi:2010:GIR

[323] Jeneen Interlandi. Getting it right on stem cells. Scientific American, 303(5):27, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
27.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-27.pdf.

Biello:2010:DWP

[324] David Biello. Darwin was a punk. Scientific American, 303(5):28, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-28a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
28a.pdf.

Tucker:2010:BBB

[325] John H. Tucker. To breed a better bird. Scientific American, 303(5):28, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-28b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
28b.pdf.

Choi:2010:LT

[326] Charles Q. Choi. Laser tag. Scientific American, 303(5):30, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-30a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
30a.pdf.

Kuchment:2010:PWb

[327] Anna Kuchment. Patent watch. Scientific American, 303(5):30, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-30b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
30b.pdf.

Musser:2010:GF

[328] George Musser. Going with the flow. Scientific American, 303(5):31, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-31.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
31.pdf.

Tucker:2010:SUI

[329] John H. Tucker. Status update: “i’m so glamorous”. Scientific American, 303(5):32, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-

32.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-32.pdf.

Kuchment:2010:WIb

[330] Anna Kuchment. What is it? Scientific American, 303(5):33, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n5/full/scientificamerican1110-33.html;   http:
//www.nature.com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
33.pdf.

Kirkwood:2010:SHW

[331] Thomas Kirkwood. The science of health: Why women live longer. Scientific American, 303(5):34–35, November 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n5/full/scientificamerican1110-34.html;   http://www.nature.
com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
34.pdf.

Pogue:2010:TTR

[332] David Pogue. TechnoFiles: The trouble with e-readers. Scientific American, 303(5):36, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
36.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-36.pdf.

Wong:2010:RRRb

[333] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 303(5):84, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
84.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-84.pdf.

Shermer:2010:SSS

[334] Michael Shermer. Skeptic: The skeptic’s skeptic. Scientific American, 303(5):86, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
86.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-86.pdf.

Mirsky:2010:AGO

[335] Steve Mirsky. Anti gravity: Oh, we have liftoff all right. Scientific American, 303(5):88, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
88.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-88.pdf.

Fischetti:2010:GSG

[336] Mark Fischetti. Graphic science: The great chemical unknown. Scientific American, 303(5):92, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
92.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-92.pdf.

Feng:2010:CDW

[337] Jonathan Feng and Mark Trodden. Cosmology: Dark worlds. Scientific American, 303(5):38–45, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v303/n5/full/scientificamerican1110-
38.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-38.pdf.

Deisseroth:2010:NCB

[338] Karl Deisseroth. Neuroscience: Controlling the brain with light. Scientific American, 303(5):48–55, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v303/n5/full/scientificamerican1110-
48.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-48.pdf.

Wald:2010:EHB

[339] Matthew L. Wald. Engineering: How to build the Supergrid. Scientific American, 303(5):56–61, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v303/n5/full/scientificamerican1110-
56.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-56.pdf.

Fischetti:2010:EPL

[340] Mark Fischetti. Ecology: Phosphorus lake. Scientific American, 303(5):62–63, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
62.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-62.pdf.

Aczel:2010:PDU

[341] Amir D. Aczel. Physics: Dr. unification. Scientific American, 303(5):64–67, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
64.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-64.pdf.

Carmichael:2010:GHH

[342] Mary Carmichael. Global health: Halting the world’s most lethal parasite. Scientific American, 303(5):68–75, November 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n5/full/scientificamerican1110-68.html;   http://www.nature.
com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
68.pdf.

Omenetto:2010:TSC

[343] Fiorenzo Omenetto and David Kaplan. Technology: From silk cocoon to medical miracle. Scientific American, 303(5):76–77, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n5/full/scientificamerican1110-76.html;   http://www.nature.
com/scientificamerican/journal/v303/n5/pdf/scientificamerican1110-
76.pdf.

Lemonick:2010:ECH

[344] Michael D. Lemonick. Environment: Climate heretic. Scientific American, 303(5):78–83, November 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n5/full/scientificamerican1110-
78.html;   http://www.nature.com/scientificamerican/journal/
v303/n5/pdf/scientificamerican1110-78.pdf.

DiChristina:2010:ESM

[345] Mariette DiChristina. From the Editor: Science that matters. Scientific American, 303(6):10, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
10.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-10.pdf.

Anonymous:2010:Ld

[346] Anonymous. Letters. Scientific American, 303(6):12–14, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-12.html;   http://www.nature.
com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
12.pdf.

Schlenoff:2010:YAIl

[347] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 303(6):106, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-106.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
106.pdf.

Anonymous:2010:SAD

[348] Anonymous. Science agenda: Defying politics. Scientific American, 303(6):16, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
16.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-16.pdf.

Greely:2010:FTT

[349] Hank Greely. Forum: To tell the truth. Scientific American, 303(6):18, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-18.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
18.pdf.

Nelson:2010:ABP

[350] Bryn Nelson. Advances: Black plants and twilight zones. Scientific American, 303(6):20, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-

20.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-20.pdf.

Carmichael:2010:PDS

[351] Mary Carmichael. Pinning down a deadly shape shifter. Scientific American, 303(6):21, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
21.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-21.pdf.

Gibbs:2010:IEF

[352] W. Wayt Gibbs and Nathan Myhrvold. The incredible, edible foam. Scientific American, 303(6):22, December 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-22.html;   http://www.nature.
com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
22.pdf.

Matson:2010:HWR

[353] John Matson. Hawking was right (probably). Scientific American, 303(6):24, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
24.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-24.pdf.

Kuchment:2010:PWc

[354] Anna Kuchment. Patent watch. Scientific American, 303(6):25, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-25.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
25.pdf.

Wong:2010:B

[355] Kate Wong. Out with a bang. Scientific American, 303(6):26, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-26a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
26a.pdf.

Anonymous:2010:Qd

[356] Anonymous. Quotable. Scientific American, 303(6):26, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-26b.html;   http://www.nature.
com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
26b.pdf.

Kuchment:2010:WIc

[357] Anna Kuchment. What is it? Scientific American, 303(6):27, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-27.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
27.pdf.

Gorman:2010:GCA

[358] Christine Gorman. A global call to action on chronic diseases. Scientific American, 303(6):28, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
28a.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-28a.pdf.

Anonymous:2010:Qe

[359] Anonymous. Quotable. Scientific American, 303(6):28, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-28b.html;   http://www.nature.
com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
28b.pdf.

Kuchment:2010:DSP

[360] Anna Kuchment. Dear Santa: Please send owl puke. Scientific American, 303(6):29, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
29.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-29.pdf.

Moyer:2010:KF

[361] Michael Moyer. Know the flow. Scientific American, 303(6):30, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-30.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
30.pdf.

Seife:2010:SD

[362] Charles Seife. The science of “disestimation”. Scientific American, 303(6):31, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
31a.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-31a.pdf.

Anonymous:2010:Qf

[363] Anonymous. Quotable. Scientific American, 303(6):31, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-31b.html;   http://www.nature.
com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
31b.pdf.

Song:2010:CT

[364] Lisa Song. A click of the tongue. Scientific American, 303(6):32, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-32a.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
32a.pdf.

Hackett:2010:NS

[365] George Hackett. News scan. Scientific American, 303(6):32, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-32b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
32b.pdf.

Gorman:2010:SHH

[366] Christine Gorman. The science of health: The heart-brain connection. Scientific American, 303(6):36–38, December 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-36.html;   http://www.nature.

com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
36.pdf.

Pogue:2010:TTM

[367] David Pogue. TechnoFiles: Talk to the machine. Scientific American, 303(6):40, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
40.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-40.pdf.

Wong:2010:RRRc

[368] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 303(6):100, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
100.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-100.pdf.

Shermer:2010:SCT

[369] Michael Shermer. Skeptic: The conspiracy theory detector. Scientific American, 303(6):102, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
102.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-102.pdf.

Mirsky:2010:AGF

[370] Steve Mirsky. Anti gravity: Fatuous fantasies. Scientific American, 303(6):104, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
104.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-104.pdf.

Fischetti:2010:GSS

[371] Mark Fischetti. Graphic science: Silent seas. Scientific American, 303(6):108, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
108.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-108.pdf.

Anonymous:2010:IWC

[372] Anonymous. Innovation: World changing ideas. Scientific American, 303(6):42, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
42.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-42.pdf.

Pavlus:2010:GL

[373] John Pavlus. The game of life. Scientific American, 303(6):43–44, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-43.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
43.pdf.

Pavlus:2010:HNC

[374] John Pavlus. Human number crunchers. Scientific American, 303(6):44, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-44.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
44.pdf.

Vanderbilt:2010:KIA

[375] Tom Vanderbilt. Know-it-all toll roads. Scientific American, 303(6):45, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-45.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
45.pdf.

Svoboda:2010:DT

[376] Elizabeth Svoboda. The DNA transistor. Scientific American, 303(6):46, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-46.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
46.pdf.

Moyer:2010:KWF

[377] Melinda Wenner Moyer. A killer water filter. Scientific American, 303(6):48, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
48a.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-48a.pdf.

Pavlus:2010:WPE

[378] John Pavlus. A wandering, plant-eating robot. Scientific American, 303(6):48, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
48b.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-48b.pdf.

Pavlus:2010:BNC

[379] John Pavlus. Borrowing nature’s code. Scientific American, 303(6):49, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-49.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
49.pdf.

Carmichael:2010:OHT

[380] Mary Carmichael. One hundred tests. Scientific American, 303(6):50, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-50.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
50.pdf.

Wald:2010:GT

[381] Matthew L. Wald. Gas from trash. Scientific American, 303(6):52, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-52.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
52.pdf.

Moyer:2010:IJD

[382] Melinda Wenner Moyer. The importance of junk DNA. Scientific American, 303(6):53, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-

53a.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-53a.pdf.

Anonymous:2010:WCV

[383] Anonymous. World changing videos. Scientific American, 303(6):53, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v303/n6/full/scientificamerican1210-53b.html;   http:
//www.nature.com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
53b.pdf.

Lisi:2010:PGT

[384] A. Garrett Lisi and James Owen Weatherall. Physics: A geometric theory of everything. Scientific American, 303(6):54–61, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-54.html;   http://www.nature.
com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
54.pdf.

Schweitzer:2010:PBS

[385] Mary H. Schweitzer. Paleontology: Blood from stone. Scientific American, 303(6):62–69, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
62.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-62.pdf.

Castelvecchi:2010:BLU

[386] Davide Castelvecchi. Biology: Life unseen. Scientific American, 303(6):70–75, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
70.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-70.pdf.

Griffiths:2010:HHM

[387] Roland R. Griffiths and Charles S. Grob. Health: Hallucinogens as medicine. Scientific American, 303(6):76–79, December 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-76.html;   http://www.nature.

com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
76.pdf.

Berners-Lee:2010:ISL

[388] Tim Berners-Lee. Information science: Long live the web. Scientific American, 303(6):80–85, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v303/n6/full/scientificamerican1210-
80.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-80.pdf.

Wong:2010:LSJ

[389] Kate Wong. Life science: Jane of the jungle. Scientific American, 303(6):86–87, December 2010. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v303/n6/full/scientificamerican1210-
86.html;   http://www.nature.com/scientificamerican/journal/
v303/n6/pdf/scientificamerican1210-86.pdf.

Freedman:2010:SEJ

[390] David H. Freedman. Space exploration: Jump-starting the orbital economy. Scientific American, 303(6):88–93, December 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-88.html;   http://www.nature.
com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
88.pdf.

Maharbiz:2010:RCB

[391] Michel M. Maharbiz and Hirotaka Sato. Robotics: Cyborg beetles. Scientific American, 303(6):94–99, December 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v303/
n6/full/scientificamerican1210-94.html;   http://www.nature.
com/scientificamerican/journal/v303/n6/pdf/scientificamerican1210-
94.pdf.

DiChristina:2011:ECW

[392] Mariette DiChristina. From the Editor: Casting a wide net. Scientific American, 304(1):8, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-

8.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-8.pdf.

Anonymous:2011:La

[393] Anonymous. Letters. Scientific American, 304(1):10–11, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n1/full/scientificamerican0111-10.html;   http://www.nature.
com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
10.pdf.

Schlenoff:2011:YAIa

[394] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 304(1):86, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-86.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
86.pdf.

Anonymous:2011:SAPa

[395] Anonymous. Science agenda: A political wish list. Scientific American, 304(1):12, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
12.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-12.pdf.

Victor:2011:FDM

[396] David G. Victor. Forum: Diplomacy’s meltdown. Scientific American, 304(1):14, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
14.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-14.pdf.

Moyer:2011:AWL

[397] Michael Moyer. Advances: A whole lot of nothing. Scientific American, 304(1):16, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
16.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-16.pdf.

Moyer:2011:HOY

[398] Melinda Wenner Moyer. How old is your cancer? Scientific American, 304(1):17, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
17a.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-17a.pdf.

Anonymous:2011:Qa

[399] Anonymous. Quotable. Scientific American, 304(1):17, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n1/full/scientificamerican0111-17b.html;   http://www.nature.
com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
17b.pdf.

Hackett:2011:NSa

[400] George Hackett. News scan. Scientific American, 304(1):17, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-17c.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
17c.pdf.

Paulos:2011:AI

[401] John Allen Paulos. Animal instincts. Scientific American, 304(1):18, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-18.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
18.pdf.

Interlandi:2011:BBBa

[402] Jeneen Interlandi. Breaching the brain barrier. Scientific American, 304(1):20, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
20a.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-20a.pdf.

Chin:2011:WIa

[403] Ann Chin. What is it? Scientific American, 304(1):20, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n1/full/scientificamerican0111-20b.html;   http://www.nature.
com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
20b.pdf.

Naish:2011:WIF

[404] Darren Naish. Will it float? Scientific American, 304(1):22, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-22.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
22.pdf.

Gibbs:2011:SSV

[405] W. Wayt Gibbs and Nathan Myhrvold. The science of sous vide. Scientific American, 304(1):24, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
24.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-24.pdf.

Kuchment:2011:PG

[406] Anna Kuchment. Practically green. Scientific American, 304(1):25, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-25a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
25a.pdf.

Anonymous:2011:Qb

[407] Anonymous. Quotable. Scientific American, 304(1):25, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n1/full/scientificamerican0111-25b.html;   http://www.nature.
com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
25b.pdf.

Arnold:2011:WSG

[408] Carrie Arnold. Why sleep is good for you. Scientific American, 304(1):26, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-26a.html;   http:

//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
26a.pdf.

Kuchment:2011:PW

[409] Anna Kuchment. Patent watch. Scientific American, 304(1):26, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-26b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
26b.pdf.

Kuchment:2011:WNT

[410] Anna Kuchment. For whom the Nobel tolls. Scientific American, 304(1): 27, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-27.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
27.pdf.

Rubenzahl:2011:BSK

[411] Lauren Rubenzahl. The bright spots of kids’ tv. Scientific American, 304(1):28, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
28.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-28.pdf.

Stix:2011:DYB

[412] Gary Stix. Donate your brain, save a buck. Scientific American, 304(1): 29, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-29a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
29a.pdf.

Anonymous:2011:Qc

[413] Anonymous. Quotable. Scientific American, 304(1):29, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n1/full/scientificamerican0111-29b.html;   http://www.nature.
com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
29b.pdf.

Greenwood:2011:SHC

[414] Veronique Greenwood. The science of health: Curing the common cold. Scientific American, 304(1):30–31, January 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n1/full/scientificamerican0111-30.html;   http://www.nature.
com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
30.pdf.

Pogue:2011:TDW

[415] David Pogue. TechnoFiles: Don’t worry about who’s watching. Scientific American, 304(1):32, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
32.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-32.pdf.

Anonymous:2011:RRRa

[416] Anonymous. Reviews and recommendations: Recommended. Scientific American, 304(1):82, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
82.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-82.pdf.

Shermer:2011:SSR

[417] Michael Shermer. Skeptic: The science of right and wrong. Scientific American, 304(1):83, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
83.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-83.pdf.

Mirsky:2011:AGT

[418] Steve Mirsky. Anti gravity: They’re young, and they’re restless. Scientific American, 304(1):84, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
84.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-84.pdf.

Fischetti:2011:GSM

[419] Mark Fischetti. Graphic science: Money for science. Scientific American, 304(1):88, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
88.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-88.pdf.

Long:2011:EDD

[420] John A. Long. Evolution: Dawn of the deed. Scientific American, 304(1):34–39, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
34.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-34.pdf.

Folger:2011:SCD

[421] Tim Folger. Space: Contact: The day after. Scientific American, 304(1):40–45, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
40.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-40.pdf.

Branswell:2011:IDF

[422] Helen Branswell. Infectious disease: Flu factories. Scientific American, 304(1):46–51, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
46.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-46.pdf.

Fischetti:2011:ESR

[423] Mark Fischetti. Energy: In search of the radical solution. Scientific American, 304(1):52–55, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
52.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-52.pdf.

Kuchment:2011:ESA

[424] Anna Kuchment. Ecology: Seeds of the Amazon. Scientific American, 304(1):56–57, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
56.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-56.pdf.

Zimmer:2011:NTC

[425] Carl Zimmer. Neuroscience: 100 trillion connections. Scientific American, 304(1):58–63, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
58.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-58.pdf.

deSherbinin:2011:ECC

[426] Alex de Sherbinin, Koko Warner, and Charles Ehrhart. Environment: Casualties of climate change. Scientific American, 304(1):64–71, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n1/full/scientificamerican0111-64.html;   http:
//www.nature.com/scientificamerican/journal/v304/n1/pdf/scientificamerican0111-
64.pdf.

King:2011:CSR

[427] Ross D. King. Computer science: Rise of the robo scientists. Scientific American, 304(1):72–77, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
72.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-72.pdf.

Rego:2011:PHR

[428] Brianna Rego. Public health: Radioactive smoke. Scientific American, 304(1):78–81, January 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n1/full/scientificamerican0111-
78.html;   http://www.nature.com/scientificamerican/journal/
v304/n1/pdf/scientificamerican0111-78.pdf.

DiChristina:2011:ETM

[429] Mariette DiChristina. From the Editor: Too much and not enough. Scientific American, 304(2):6, February 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-6.html;   http://www.nature.
com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
6.pdf.

Anonymous:2011:Lb

[430] Anonymous. Letters. Scientific American, 304(2):8–10, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-8.html;   http://www.nature.
com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
8.pdf.

Schlenoff:2011:YAIb

[431] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 304(2):91, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n2/full/scientificamerican0211-91.html;   http:
//www.nature.com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
91.pdf.

Anonymous:2011:SAFa

[432] Anonymous. Science agenda: Fear and its consequences. Scientific American, 304(2):14, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
14.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-14.pdf.

Grifo:2011:FBS

[433] Francesca Grifo. Forum: The bright side of gridlock. Scientific American, 304(2):16, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
16.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-16.pdf.

Moyer:2011:AWP

[434] Melinda Wenner Moyer. Advances: Which pills work? Scientific American, 304(2):20, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-

20.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-20.pdf.

Jabr:2011:TLY

[435] Ferris Jabr. They like your guts. Scientific American, 304(2):22, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n2/full/scientificamerican0211-22a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
22a.pdf.

Katsnelson:2011:DCM

[436] Alla Katsnelson. Does this collar make me look fat? Scientific American, 304(2):22, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
22b.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-22b.pdf.

Anonymous:2011:BMB

[437] Anonymous. The Bird Man of Baghdad. Scientific American, 304(2):24, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n2/full/scientificamerican0211-24a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
24a.pdf.

Anonymous:2011:Qd

[438] Anonymous. Quotable. Scientific American, 304(2):24, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-24b.html;   http://www.nature.
com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
24b.pdf.

Castelvecchi:2011:YPC

[439] Davide Castelvecchi. Not your parents’ carbon. Scientific American, 304(2):25, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
25a.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-25a.pdf.

Gibbs:2011:BBB

[440] W. Wayt Gibbs and Nathan Myhrvold. Beer batter is better. Scientific American, 304(2):25, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
25b.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-25b.pdf.

Spinney:2011:YSF

[441] Laura Spinney. You smell flowers, I smell stale urine. Scientific American, 304(2):26, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
26a.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-26a.pdf.

Chin:2011:WIb

[442] Ann Chin. What is it? Scientific American, 304(2):26, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-26b.html;   http://www.nature.
com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
26b.pdf.

Musser:2011:FR

[443] George Musser. Forces to reckon with. Scientific American, 304(2):27, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n2/full/scientificamerican0211-27a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
27a.pdf.

Anonymous:2011:Qe

[444] Anonymous. Quotable. Scientific American, 304(2):27, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-27b.html;   http://www.nature.
com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
27b.pdf.

Wapner:2011:CAF

[445] Jessica Wapner. Charging against the flu. Scientific American, 304(2):28, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n2/full/scientificamerican0211-28a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
28a.pdf.

Moyer:2011:BIW

[446] Melinda Wenner Moyer. Blame it on winter. Scientific American, 304(2):28, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
28b.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-28b.pdf.

Aczel:2011:PF

[447] Amir D. Aczel. Particles that flock. Scientific American, 304(2):30, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n2/full/scientificamerican0211-30.html;   http:
//www.nature.com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
30.pdf.

Arnold:2011:WEW

[448] Carrie Arnold. When Earth was a snowball. Scientific American, 304(2):32, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
32a.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-32a.pdf.

Hackett:2011:NSb

[449] George Hackett. News scan. Scientific American, 304(2):32, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n2/full/scientificamerican0211-32b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
32b.pdf.

Paulos:2011:WYP

[450] John Allen Paulos. Why you’re probably less popular than your friends. Scientific American, 304(2):33, February 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-33.html;   http://www.nature.

com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
33.pdf.

Moisse:2011:SHY

[451] Katie Moisse. The science of health: The YouTube cure. Scientific American, 304(2):34–37, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
34.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-34.pdf.

Pogue:2011:TOQ

[452] David Pogue. TechnoFiles: An open question. Scientific American, 304(2):38, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
38.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-38.pdf.

Anonymous:2011:RRRb

[453] Anonymous. Reviews and recommendations: Recommended. Scientific American, 304(2):88, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
88.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-88.pdf.

Shermer:2011:SHS

[454] Michael Shermer. Skeptic: Houdini’s skeptical advice. Scientific American, 304(2):89, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
89.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-89.pdf.

Mirsky:2011:AGBa

[455] Steve Mirsky. Anti gravity: La bummer. Scientific American, 304(2):89, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n2/full/scientificamerican0211-90.html;   http:
//www.nature.com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
90.pdf.

Fischetti:2011:GSY

[456] Mark Fischetti. Graphic science: Your brain in love. Scientific American, 304(2):92, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
92.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-92.pdf.

Freedman:2011:HHF

[457] David H. Freedman. Health: How to fix the obesity crisis. Scientific American, 304(2):40–47, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
40.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-40.pdf.

Pang:2011:SCS

[458] Alex Soojung-Kim Pang and Bob Twiggs. Space: Citizen satellites. Scientific American, 304(2):48–53, February 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-48.html;   http://www.nature.
com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
48.pdf.

Simpson:2011:SBF

[459] Sarah Simpson. Sustainability: The blue food revolution. Scientific American, 304(2):54–61, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
54.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-54.pdf.

Boroditsky:2011:CPH

[460] Lera Boroditsky. Cognitive psychology: How language shapes thought. Scientific American, 304(2):62–65, February 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-62.html;   http://www.nature.
com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
62.pdf.

Misteli:2011:BIL

[461] Tom Misteli. Biology: The inner life of the genome. Scientific American, 304(2):66–73, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
66.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-66.pdf.

Borrell:2011:EFA

[462] Brendan Borrell. Ecology: A friend to aliens. Scientific American, 304(2):74–77, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
74.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-74.pdf.

Harrison:2011:AXR

[463] Fiona Harrison and Charles J. Hailey. Astrophysics: X-ray vision. Scientific American, 304(2):78–79, February 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n2/full/scientificamerican0211-78.html;   http://www.nature.
com/scientificamerican/journal/v304/n2/pdf/scientificamerican0211-
78.pdf.

Nicolelis:2011:NMB

[464] Miguel A. L. Nicolelis. Neuroprosthetics: Mind out of body. Scientific American, 304(2):80–83, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
80.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-80.pdf.

Dugatkin:2011:HSJ

[465] Lee Dugatkin. History of science: Jefferson’s moose. Scientific American, 304(2):84–87, February 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n2/full/scientificamerican0211-
84.html;   http://www.nature.com/scientificamerican/journal/
v304/n2/pdf/scientificamerican0211-84.pdf.

DiChristina:2011:ESP

[466] Mariette DiChristina. From the Editor: Spaces and places. Scientific American, 304(3):6, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
6.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-6.pdf.

Anonymous:2011:Lc

[467] Anonymous. Letters. Scientific American, 304(3):8–10, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n3/full/scientificamerican0311-8.html;   http://www.nature.
com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
8.pdf.

Schlenoff:2011:YAIc

[468] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 304(3):79, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-79.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
79.pdf.

Anonymous:2011:SAK

[469] Anonymous. Science agenda: Keep the Internet fair. Scientific American, 304(3):12, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
12.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-12.pdf.

Zittrain:2011:FFA

[470] Jonathan Zittrain. Forum: Freedom and anonymity. Scientific American, 304(3):13, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
13.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-13.pdf.

Biello:2011:ACF

[471] David Biello. Advances: Coal fires burning bright. Scientific American, 304(3):14, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
14.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-14.pdf.

Coffey:2011:BVM

[472] Rebecca Coffey. Bison vs. mammoths. Scientific American, 304(3):15, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-15a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
15a.pdf.

Hackett:2011:NSc

[473] George Hackett. News scan. Scientific American, 304(3):15, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n3/full/scientificamerican0311-15b.html;   http://www.nature.
com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
15b.pdf.

Greenemeier:2011:CYS

[474] Larry Greenemeier. Can you see me now? Scientific American, 304(3): 16, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-16.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
16.pdf.

Ghose:2011:LHT

[475] Tia Ghose. A little help from their friends. Scientific American, 304(3): 17, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-17.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
17.pdf.

Matson:2011:FH

[476] John Matson. Feeling the heat. Scientific American, 304(3):18, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-18a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
18a.pdf.

Anonymous:2011:Qf

[477] Anonymous. Quotable. Scientific American, 304(3):18, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n3/full/scientificamerican0311-18b.html;   http://www.nature.
com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
18b.pdf.

Moyer:2011:OC

[478] Melinda Wenner Moyer. Organs-on-a-chip. Scientific American, 304(3): 19, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-19a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
19a.pdf.

Anonymous:2011:Qg

[479] Anonymous. Quotable. Scientific American, 304(3):19, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n3/full/scientificamerican0311-19b.html;   http://www.nature.
com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
19b.pdf.

Matson:2011:SS

[480] John Matson. Scrivener to the stars. Scientific American, 304(3):20, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-20.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
20.pdf.

Castelvecchi:2011:SM

[481] Davide Castelvecchi. The smallest mind. Scientific American, 304(3):21, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-21a.html;   http:

//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
21a.pdf.

Chin:2011:WIc

[482] Ann Chin. What is it? Scientific American, 304(3):21, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n3/full/scientificamerican0311-21b.html;   http://www.nature.
com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
21b.pdf.

Choi:2011:IS

[483] Charles Q. Choi. From iPhones to SciPhones. Scientific American, 304(3):22, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
22a.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-22a.pdf.

McCook:2011:SG

[484] Alison McCook. She’s 11, going on 2,500. Scientific American, 304(3): 22, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-22b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
22b.pdf.

Gibbs:2011:NSC

[485] W. Wayt Gibbs and Nathan Myhrvold. A new spin on cooking. Scientific American, 304(3):23, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
23a.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-23a.pdf.

Anonymous:2011:Qh

[486] Anonymous. Quotable. Scientific American, 304(3):23, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n3/full/scientificamerican0311-23b.html;   http://www.nature.
com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
23b.pdf.

Franklin:2011:SHU

[487] Deborah Franklin. The science of health: Uninformed consent. Scientific American, 304(3):24–25, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
24.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-24.pdf.

Pogue:2011:TGP

[488] David Pogue. TechnoFiles: Gadget politics. Scientific American, 304(3): 26, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n3/full/scientificamerican0311-26.html;   http:
//www.nature.com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
26.pdf.

Wong:2011:RRRa

[489] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 304(3):76, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
76.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-76.pdf.

Shermer:2011:SFF

[490] Michael Shermer. Skeptic: Financial flimflam. Scientific American, 304(3):77, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
77.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-77.pdf.

Mirsky:2011:AGF

[491] Steve Mirsky. Anti gravity: The future is now. Scientific American, 304(3):78, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
78.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-78.pdf.

Fischetti:2011:GSDa

[492] Mark Fischetti. Graphic science: Death by asteroid. Scientific American, 304(3):80, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
80.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-80.pdf.

Stix:2011:PNT

[493] Gary Stix. Psychology: The neuroscience of true grit. Scientific American, 304(3):28–33, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
28.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-28.pdf.

Murchie:2011:SSJ

[494] Scott L. Murchie, Ronald J. Vervack, Jr., and Brian J. Anderson. Space science: Journey to the innermost planet. Scientific American, 304(3):34–39, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
34.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-34.pdf.

Hall:2011:MDD

[495] Stephen S. Hall. Medicine: Diseases in a dish. Scientific American, 304(3):40–45, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
40.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-40.pdf.

Schoonover:2011:ISS

[496] Carl Schoonover. Imaging: Signals in a storm. Scientific American, 304(3):46–47, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
46.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-46.pdf.

Underhill:2011:APS

[497] William Underhill. Archaeology: Putting Stonehenge in its place. Scientific American, 304(3):48–53, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v304/n3/full/scientificamerican0311-

48.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-48.pdf.

Raizen:2011:PDE

[498] Mark G. Raizen. Physics: Demons, entropy and the quest for absolute zero. Scientific American, 304(3):54–59, March 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n3/full/scientificamerican0311-54.html;   http://www.nature.
com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
54.pdf.

Sachs:2011:CCS

[499] Julian P. Sachs and Conor L. Myhrvold. Climate change: A shifting band of rain. Scientific American, 304(3):60–65, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n3/full/scientificamerican0311-60.html;   http://www.nature.
com/scientificamerican/journal/v304/n3/pdf/scientificamerican0311-
60.pdf.

Carmichael:2011:PHJ

[500] Mary Carmichael. Public health: Not just an illness of the rich. Scientific American, 304(3):66–69, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
66.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-66.pdf.

Sereno:2011:PDD

[501] Paul C. Sereno. Paleontology: Dinosaur death trap. Scientific American, 304(3):70–75, March 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n3/full/scientificamerican0311-
70.html;   http://www.nature.com/scientificamerican/journal/
v304/n3/pdf/scientificamerican0311-70.pdf.

DiChristina:2011:ERS

[502] Mariette DiChristina. From the Editor: Reflections from science. Scientific American, 304(4):6, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-

6.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-6.pdf.

Anonymous:2011:Ld

[503] Anonymous. Letters. Scientific American, 304(4):8–10, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-8.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
8.pdf.

Schlenoff:2011:YAId

[504] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 304(4):94, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-94.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
94.pdf.

Anonymous:2011:SAO

[505] Anonymous. Science agenda: Our big pig problem. Scientific American, 304(4):12, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
12.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-12.pdf.

Krauss:2011:FRD

[506] Lawrence M. Krauss. Forum: Rethinking the dream. Scientific American, 304(4):14, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
14.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-14.pdf.

Coffey:2011:AOD

[507] Rebecca Coffey. Advances: Outsmarting dengue fever. Scientific American, 304(4):16, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
16.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-16.pdf.

Choi:2011:TYI

[508] Charles Q. Choi. Tame your inner tiger. Scientific American, 304(4):19, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-19a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
19a.pdf.

Moyer:2011:WLW

[509] Michael Moyer. Watson looks for work. Scientific American, 304(4):19, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-19b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
19b.pdf.

Anonymous:2011:Na

[510] Anonymous. By the numbers. Scientific American, 304(4):19, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-19c.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
19c.pdf.

Singer:2011:MWL

[511] Thea Singer. My, what long telomeres you have. Scientific American, 304(4):22, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
22a.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-22a.pdf.

Power:2011:TCF

[512] Carla Power. Too contagious to fail. Scientific American, 304(4):22, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-22b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
22b.pdf.

Castelvecchi:2011:CCO

[513] Davide Castelvecchi. Cracking a century-old enigma. Scientific American, 304(4):24, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
24a.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-24a.pdf.

Anonymous:2011:Qi

[514] Anonymous. Quotable. Scientific American, 304(4):24, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-24b.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
24b.pdf.

Chin:2011:WId

[515] Ann Chin. What is it? Scientific American, 304(4):25, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-25.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
25.pdf.

Mazumdar:2011:ALT

[516] Sudip Mazumdar. Arid land, thirsty crops. Scientific American, 304(4): 26, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-26.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
26.pdf.

Moyer:2011:TMI

[517] Melinda Wenner Moyer. Too much information? Scientific American, 304(4):27, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
27a.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-27a.pdf.

Anonymous:2011:Nb

[518] Anonymous. By the numbers. Scientific American, 304(4):27, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-27b.html;   http://www.nature.

com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
27b.pdf.

Easter:2011:GKY

[519] Michael Easter. Getting to know you. Scientific American, 304(4):28, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-28a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
28a.pdf.

Anonymous:2011:Qj

[520] Anonymous. Quotable. Scientific American, 304(4):28, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-28b.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
28b.pdf.

Funkhouser:2011:CLN

[521] David Funkhouser. Crab love nest. Scientific American, 304(4):29, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-29.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
29.pdf.

Arnold:2011:SSA

[522] Carrie Arnold. To share and share alike. Scientific American, 304(4):30– 31, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-30.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
30.pdf.

Piore:2011:PWa

[523] Adam Piore. Patent watch. Scientific American, 304(4):31, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-31.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
31.pdf.

Greenwood:2011:SHW

[524] Veronique Greenwood. The science of health: Why are asthma rates soaring? Scientific American, 304(4):32–33, April 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-32.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
32.pdf.

Pogue:2011:TSF

[525] David Pogue. TechnoFiles: Seeing forever. Scientific American, 304(4): 34, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n4/full/scientificamerican0411-34.html;   http:
//www.nature.com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
34.pdf.

Wong:2011:RRRb

[526] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 304(4):88, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
88.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-88.pdf.

Shermer:2011:SUU

[527] Michael Shermer. Skeptic: UFOs, UAPs and CRAPs. Scientific American, 304(4):90, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
90.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-90.pdf.

Mirsky:2011:AGK

[528] Steve Mirsky. Anti gravity: Killer entertainment. Scientific American, 304(4):92, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
92.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-92.pdf.

Fischetti:2011:GSCa

[529] Mark Fischetti. Graphic science: Clean tech rising. Scientific American, 304(4):96, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
96.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-96.pdf.

Steinhardt:2011:CID

[530] Paul J. Steinhardt. Cosmology: The inflation debate. Scientific American, 304(4):36–43, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
36.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-36.pdf.

McKenna:2011:MEW

[531] Maryn McKenna. Medicine: The enemy within. Scientific American, 304(4):46–53, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
46.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-46.pdf.

Gazzaniga:2011:INC

[532] Michael S. Gazzaniga. Imaging: Neuroscience in the courtroom. Scientific American, 304(4):54–59, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
54.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-54.pdf.

Haddok:2011:ECD

[533] Eitan Haddok. Environment: Can the dead sea live? Scientific American, 304(4):60–65, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
60.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-60.pdf.

Collins:2011:AIS

[534] Graham P. Collins. Artificial intelligence: Solving the cocktail party problem. Scientific American, 304(4):66–67, April 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-66.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
66.pdf.

Dunker:2011:CBO

[535] A. Keith Dunker and Richard W. Kriwacki. Cell biology: The orderly chaos of proteins. Scientific American, 304(4):68–73, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n4/full/scientificamerican0411-68.html;   http://www.nature.
com/scientificamerican/journal/v304/n4/pdf/scientificamerican0411-
68.pdf.

Allen:2011:SSB

[536] Richard Allen. Seismology: Seconds before the big one. Scientific American, 304(4):74–79, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
74.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-74.pdf.

Borrell:2011:SFF

[537] Brendan Borrell. Sustainability: Food fight. Scientific American, 304(4):80–83, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
80.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-80.pdf.

Catania:2011:ABN

[538] Kenneth C. Catania. Animal behavior: Natural-born killer. Scientific American, 304(4):84–87, April 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n4/full/scientificamerican0411-
84.html;   http://www.nature.com/scientificamerican/journal/
v304/n4/pdf/scientificamerican0411-84.pdf.

DiChristina:2011:ERV

[539] Mariette DiChristina. From the Editor: Risks vs. gains. Scientific American, 304(5):4, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
4.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-4.pdf.

Anonymous:2011:Le

[540] Anonymous. Letters. Scientific American, 304(5):6–8, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-6.html;   http://www.nature.
com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
6.pdf.

Schlenoff:2011:YAIe

[541] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 304(5): 94, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-94.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
94.pdf.

Anonymous:2011:SAD

[542] Anonymous. Science agenda: Dr. no money. Scientific American, 304(5): 10, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-10.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
10.pdf.

Willingham:2011:FTM

[543] Daniel T. Willingham. Forum: Trust me, I’m a scientist. Scientific American, 304(5):12, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
12.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-12.pdf.

Biello:2011:AWF

[544] David Biello. Advances: Warning: Flooding ahead. Scientific American, 304(5):16, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-

16.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-16.pdf.

Coffey:2011:CL

[545] Rebecca Coffey. Coral in love. Scientific American, 304(5):18, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-18a.html;   http://www.nature.
com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
18a.pdf.

Kuchment:2011:WCG

[546] Anna Kuchment. When cars are greener than people. Scientific American, 304(5):18, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
18b.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-18b.pdf.

Moyer:2011:CTT

[547] Melinda Wenner Moyer. Cancer testing? there’s an app for that. Scientific American, 304(5):20, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
20a.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-20a.pdf.

Chin:2011:WIe

[548] Ann Chin. What is it? Scientific American, 304(5):20, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-20b.html;   http://www.nature.
com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
20b.pdf.

Perkins:2011:BBO

[549] Sid Perkins. Bracing for the big one. Scientific American, 304(5):22, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-22a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
22a.pdf.

Biello:2011:NNP

[550] David Biello. The newest nuclear plants. Scientific American, 304(5): 22, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-22b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
22b.pdf.

Choi:2011:SR

[551] Charles Q. Choi. Steel rainbow. Scientific American, 304(5):23, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-23.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
23.pdf.

Interlandi:2011:BBBb

[552] Jeneen Interlandi. Beating back the bugs. Scientific American, 304(5): 24, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-24a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
24a.pdf.

Easter:2011:PR

[553] Michael Easter. The prices are right. Scientific American, 304(5):24, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-24b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
24b.pdf.

Westly:2011:HVM

[554] Erica Westly. HIV vs. measles. Scientific American, 304(5):26, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-26a.html;   http://www.nature.
com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
26a.pdf.

Piore:2011:PWb

[555] Adam Piore. Patent watch. Scientific American, 304(5):26, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-26b.html;   http://www.nature.
com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
26b.pdf.

McGrayne:2011:WBR

[556] Sharon Bertsch McGrayne. Why Bayes rules. Scientific American, 304 (5):28, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-28a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
28a.pdf.

Collins:2011:CWS

[557] Sonya Collins. A cure within sight? Scientific American, 304(5):28, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-28b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
28b.pdf.

Biello:2011:LGI

[558] David Biello. “let’s go for it”. Scientific American, 304(5):30, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-30a.html;   http://www.nature.
com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
30a.pdf.

Anonymous:2011:Qk

[559] Anonymous. Quotable. Scientific American, 304(5):30, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-30b.html;   http://www.nature.
com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
30b.pdf.

Hackett:2011:NSd

[560] George Hackett. News scan. Scientific American, 304(5):31, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-31.html;   http://www.nature.

com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
31.pdf.

Shute:2011:SHB

[561] Nancy Shute. The science of health: Beyond mammograms. Scientific American, 304(5):32–34, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
32.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-32.pdf.

Pogue:2011:TGG

[562] David Pogue. TechnoFiles: Gadgets are garbage. Scientific American, 304(5):36, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
36.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-36.pdf.

Anonymous:2011:RRRc

[563] Anonymous. Reviews and recommendations: Recommended. Scientific American, 304(5):88, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
88.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-88.pdf.

Shermer:2011:SEP

[564] Michael Shermer. Skeptic: Extrasensory pornception. Scientific American, 304(5):90, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
90.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-90.pdf.

Mirsky:2011:AGM

[565] Steve Mirsky. Anti gravity: O mercaptan, my mercaptan. Scientific American, 304(5):92, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
92.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-92.pdf.

Fischetti:2011:GSE

[566] Mark Fischetti. Graphic science: Exposed. Scientific American, 304(5): 96, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-96.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
96.pdf.

Anonymous:2011:SRE

[567] Anonymous. Sustainability: 7 radical energy solutions. Scientific American, 304(5):38–45, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
38.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-38.pdf.

Geach:2011:CLG

[568] James E. Geach. Cosmology: The lost galaxies. Scientific American, 304(5):46–53, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
46.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-46.pdf.

Provencio:2011:NHO

[569] Ignacio Provencio. Neuroscience: The hidden organ in our eyes. Scientific American, 304(5):54–59, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
54.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-54.pdf.

Baez:2011:MSN

[570] John C. Baez and John Huerta. Mathematics: The strangest numbers in string theory. Scientific American, 304(5):60–65, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n5/full/scientificamerican0511-60.html;   http://www.nature.
com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
60.pdf.

Aderem:2011:MFT

[571] Alan Aderem. Medicine: Fast track to vaccines. Scientific American, 304(5):66–71, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
66.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-66.pdf.

Musser:2011:PSS

[572] George Musser. Physics: The Space Station’s crown jewel. Scientific American, 304(5):72–73, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
72.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-72.pdf.

Adler:2011:AGM

[573] Jerry Adler. Agriculture: The growing menace from superweeds. Scientific American, 304(5):74–79, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
74.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-74.pdf.

Forbes:2011:NHM

[574] Peter Forbes. Natural history: Masters of disguise. Scientific American, 304(5):80–83, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n5/full/scientificamerican0511-
80.html;   http://www.nature.com/scientificamerican/journal/
v304/n5/pdf/scientificamerican0511-80.pdf.

Anstead:2011:CIS

[575] Alicia Anstead. Creativity: Inner sparks. Scientific American, 304(5): 84–87, May 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n5/full/scientificamerican0511-84.html;   http:
//www.nature.com/scientificamerican/journal/v304/n5/pdf/scientificamerican0511-
84.pdf.

DiChristina:2011:EQL

[576] Mariette DiChristina. From the Editor: Quantum leap. Scientific American, 304(6):6, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
6.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-6.pdf.

Anonymous:2011:Lf

[577] Anonymous. Letters. Scientific American, 304(6):10–12, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n6/full/scientificamerican0611-10.html;   http://www.nature.
com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
10.pdf.

Schlenoff:2011:YAIf

[578] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 304(6):91, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-91.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
91.pdf.

Anonymous:2011:SAC

[579] Anonymous. Science agenda: Coming clean about nuclear power. Scientific American, 304(6):14, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
14.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-14.pdf.

Ioannidis:2011:FEF

[580] John P. A. Ioannidis. Forum: An epidemic of false claims. Scientific American, 304(6):16, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
16.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-16.pdf.

Greenemeier:2011:AHM

[581] Larry Greenemeier. Advances: Hack my ride. Scientific American, 304(6):18, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-

18.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-18.pdf.

Martinez:2011:SH

[582] Amanda Rose Martinez. The smallest hitchhikers. Scientific American, 304(6):21, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
21a.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-21a.pdf.

Easter:2011:BB

[583] Michael Easter. A batter for a batter. Scientific American, 304(6):21, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-21b.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
21b.pdf.

Collins:2011:TT

[584] Sonya Collins. Treating Tourette’s. Scientific American, 304(6):22, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-22a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
22a.pdf.

Anonymous:2011:Ql

[585] Anonymous. Quotable. Scientific American, 304(6):22, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n6/full/scientificamerican0611-22b.html;   http://www.nature.
com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
22b.pdf.

Chin:2011:WIf

[586] Ann Chin. What is it? Scientific American, 304(6):23, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n6/full/scientificamerican0611-23.html;   http://www.nature.
com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
23.pdf.

Moyer:2011:SI

[587] Melinda Wenner Moyer. Spies inside. Scientific American, 304(6):24, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-24a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
24a.pdf.

Anonymous:2011:Sa

[588] Anonymous. Stat. Scientific American, 304(6):24, June 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n6/full/scientificamerican0611-24b.html;   http://www.nature.
com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
24b.pdf.

Fischetti:2011:WIS

[589] Mark Fischetti. Will the Internet stop on June 8? Scientific American, 304(6):25, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
25.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-25.pdf.

Gibbs:2011:MP

[590] W. Wayt Gibbs. Material poet. Scientific American, 304(6):26, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-26.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
26.pdf.

Arnold:2011:TRE

[591] Carrie Arnold. Off the tree, ready to eat. Scientific American, 304(6): 28, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-28a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
28a.pdf.

Musser:2011:MWC

[592] George Musser. A mystery wrapped in a crystal. Scientific American, 304(6):28, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
28b.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-28b.pdf.

Paulos:2011:NXA

[593] John Allen Paulos. No X-aggeration. Scientific American, 304(6):30, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-30a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
30a.pdf.

Anonymous:2011:Sb

[594] Anonymous. Stat. Scientific American, 304(6):30, June 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n6/full/scientificamerican0611-30b.html;   http://www.nature.
com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
30b.pdf.

Requarth:2011:HBB

[595] Tim Requarth and Meehan Crist. How brains bounce back. Scientific American, 304(6):31, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
31.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-31.pdf.

Raymond:2011:JFC

[596] Joan Raymond. Join the f***ing club. Scientific American, 304(6):32, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-32a.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
32a.pdf.

Anonymous:2011:Qm

[597] Anonymous. Quotable. Scientific American, 304(6):32, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n6/full/scientificamerican0611-32b.html;   http://www.nature.

com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
32b.pdf.

Hughes:2011:SHS

[598] Virginia Hughes. The science of health: Shades of grief. Scientific American, 304(6):34–35, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
34.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-34.pdf.

Pogue:2011:TCM

[599] David Pogue. TechnoFiles: Critical mass. Scientific American, 304(6): 36, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v304/n6/full/scientificamerican0611-36.html;   http:
//www.nature.com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
36.pdf.

Wong:2011:RRRc

[600] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 304(6):88, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
88.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-88.pdf.

Shermer:2011:SME

[601] Michael Shermer. Skeptic: The myth of the evil aliens. Scientific American, 304(6):89, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
89.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-89.pdf.

Mirsky:2011:AGD

[602] Steve Mirsky. Anti gravity: Dining and dancing. Scientific American, 304(6):90, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
90.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-90.pdf.

Fischetti:2011:GSF

[603] Mark Fischetti. Graphic science: Up in flames. Scientific American, 304(6):92, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
92.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-92.pdf.

Vedral:2011:PLQ

[604] Vlatko Vedral. Physics: Living in a quantum world. Scientific American, 304(6):38–43, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
38.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-38.pdf.

Koch:2011:NTC

[605] Christof Koch and Giulio Tononi. Neuroscience: A test for consciousness. Scientific American, 304(6):44–47, June 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n6/full/scientificamerican0611-44.html;   http://www.nature.
com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
44.pdf.

Piore:2011:NEP

[606] Adam Piore. Nuclear energy: Planning for the black swan. Scientific American, 304(6):48–53, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
48.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-48.pdf.

Jabr:2011:BNC

[607] Ferris Jabr. Biology: A Nobel celebration. Scientific American, 304(6):54–63, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
54.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-54.pdf.

Bartholet:2011:TIM

[608] Jeffrey Bartholet. Technology: Inside the meat lab. Scientific American, 304(6):64–69, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
64.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-64.pdf.

Kuchment:2011:MSB

[609] Anna Kuchment and staff editor. Microbiology: The smartest bacteria on Earth. Scientific American, 304(6):70–71, June 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v304/
n6/full/scientificamerican0611-70.html;   http://www.nature.
com/scientificamerican/journal/v304/n6/pdf/scientificamerican0611-
70.pdf.

Jones:2011:HDC

[610] Menna E. Jones and Hamish McCallum. Health: The devil’s cancer. Scientific American, 304(6):72–77, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v304/n6/full/scientificamerican0611-
72.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-72.pdf.

Larson:2011:HSG

[611] Edward J. Larson. History of science: Greater glory. Scientific American, 304(6):78–83, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
78.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-78.pdf.

Lemonick:2011:GWS

[612] Michael D. Lemonick. Global warming: “i stick to the science”. Scientific American, 304(6):84–87, June 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v304/n6/full/scientificamerican0611-
84.html;   http://www.nature.com/scientificamerican/journal/
v304/n6/pdf/scientificamerican0611-84.pdf.

DiChristina:2011:EHA

[613] Mariette DiChristina. From the Editor: Honors and activities. Scientific American, 305(1):6, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
6.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-6.pdf.

Anonymous:2011:Lg

[614] Anonymous. Letters. Scientific American, 305(1):8–10, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-8.html;   http://www.nature.
com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
8.pdf.

Schlenoff:2011:YAIg

[615] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 305(1): 87, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-87.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
87.pdf.

Anonymous:2011:SAPb

[616] Anonymous. Science agenda: Physician, heal the system. Scientific American, 305(1):12–13, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
12.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-12.pdf.

Searchinger:2011:FQF

[617] Timothy Searchinger. Forum: A quick fix to the food crisis. Scientific American, 305(1):14, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
14.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-14.pdf.

Biello:2011:ATR

[618] David Biello. Advances: Tortoises to the rescue. Scientific American, 305 (1):16, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-16.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
16.pdf.

Matson:2011:IPM

[619] John Matson. International particle of mystery. Scientific American, 305 (1):19, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-19a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
19a.pdf.

Chin:2011:WIg

[620] Ann Chin. What is it? Scientific American, 305(1):19, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-19b.html;   http://www.nature.
com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
19b.pdf.

Lewinski:2011:LER

[621] John Scott Lewinski. Light at the end of the racetrack. Scientific American, 305(1):20, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
20.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-20.pdf.

Kuchment:2011:ID

[622] Anna Kuchment. Itch doctor. Scientific American, 305(1):21, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-21a.html;   http://www.nature.
com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
21a.pdf.

Anonymous:2011:Qn

[623] Anonymous. Quotable. Scientific American, 305(1):21, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-21b.html;   http://www.nature.
com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
21b.pdf.

Nelson:2011:BBC

[624] Bryn Nelson. Big buzzword on campus. Scientific American, 305(1): 22, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-22a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
22a.pdf.

Piore:2011:PWc

[625] Adam Piore. Patent watch. Scientific American, 305(1):22, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-22b.html;   http://www.nature.
com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
22b.pdf.

Madren:2011:WWS

[626] Carrie Madren. A wild, weedy scourge. Scientific American, 305(1): 24, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-24a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
24a.pdf.

Coffey:2011:BB

[627] Rebecca Coffey. Beauty and the beasts. Scientific American, 305(1): 24, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-24b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
24b.pdf.

Bai:2011:DF

[628] Nina Bai. Donor fatigue. Scientific American, 305(1):26, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-26a.html;   http://www.nature.

com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
26a.pdf.

Anonymous:2011:AQI

[629] Anonymous. Al Qaeda and the Internet. Scientific American, 305(1): 26, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-26b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
26b.pdf.

Gibbs:2011:UFF

[630] W. Wayt Gibbs and Nathan Myhrvold. Ultrasonic French fries. Scientific American, 305(1):27, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
27.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-27.pdf.

Biello:2011:NPN

[631] David Biello. From nuclear plant to nuclear park? Scientific American, 305(1):28, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
28a.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-28a.pdf.

Hackett:2011:NSe

[632] George Hackett. News scan. Scientific American, 305(1):28, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-28b.html;   http://www.nature.
com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
28b.pdf.

Wapner:2011:SHP

[633] Jessica Wapner. The science of health: Poor man’s burden. Scientific American, 305(1):30–32, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
30.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-30.pdf.

Pogue:2011:TWG

[634] David Pogue. TechnoFiles: Why gadgets flop. Scientific American, 305 (1):34, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-34.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
34.pdf.

Anonymous:2011:RRRd

[635] Anonymous. Reviews and recommendations: Recommended. Scientific American, 305(1):84, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
84.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-84.pdf.

Shermer:2011:SBB

[636] Michael Shermer. Skeptic: The believing brain. Scientific American, 305 (1):85, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-85.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
85.pdf.

Mirsky:2011:AGRa

[637] Steve Mirsky. Anti gravity: Rules of the road. Scientific American, 305 (1):86, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n1/full/scientificamerican0711-86.html;   http:
//www.nature.com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
86.pdf.

Fischetti:2011:GSBa

[638] Mark Fischetti. Graphic science: Baby’s life, mother’s schooling. Scientific American, 305(1):88, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
88.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-88.pdf.

Fox:2011:NLI

[639] Douglas Fox. Neuroscience: The limits of intelligence. Scientific American, 305(1):36–43, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
36.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-36.pdf.

Croswell:2011:APT

[640] Ken Croswell. Astrophysics: The periodic table of the cosmos. Scientific American, 305(1):44–49, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
44.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-44.pdf.

Begley:2011:HCB

[641] Sharon Begley. Health care: The best medicine. Scientific American, 305(1):50–55, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
50.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-50.pdf.

Kump:2011:CCL

[642] Lee R. Kump. Climate change: The last great global warming. Scientific American, 305(1):56–61, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
56.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-56.pdf.

Kuchment:2011:EUR

[643] Anna Kuchment and staff editor. Engineering: Underground railroad. Scientific American, 305(1):62–63, July 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-62.html;   http://www.nature.
com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
62.pdf.

Lamb:2011:BEE

[644] Trevor D. Lamb. Biology: Evolution of the eye. Scientific American, 305(1):64–69, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-

64.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-64.pdf.

Nicol:2011:CHL

[645] David M. Nicol. Cybersecurity: Hacking the lights out. Scientific American, 305(1):70–75, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
70.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-70.pdf.

Carlson:2011:DCS

[646] John R. Carlson and Allison F. Carey. Disease control: Scent of a human. Scientific American, 305(1):76–79, July 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n1/full/scientificamerican0711-76.html;   http://www.nature.
com/scientificamerican/journal/v305/n1/pdf/scientificamerican0711-
76.pdf.

Byrne:2011:CBB

[647] Peter Byrne. Cosmology: Bad boy of physics. Scientific American, 305(1):80–83, July 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n1/full/scientificamerican0711-
80.html;   http://www.nature.com/scientificamerican/journal/
v305/n1/pdf/scientificamerican0711-80.pdf.

DiChristina:2011:EFS

[648] Mariette DiChristina. From the Editor: A future with science. Scientific American, 305(2):8, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
8.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-8.pdf.

Anonymous:2011:Lh

[649] Anonymous. Letters. Scientific American, 305(2):10–12, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-10.html;   http://www.nature.
com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
10.pdf.

Schlenoff:2011:YAIh

[650] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 305(2):94, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-94.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
94.pdf.

Anonymous:2011:SASa

[651] Anonymous. Science agenda: Stand and deliver. Scientific American, 305(2):14, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
14.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-14.pdf.

Kaitin:2011:FDN

[652] Kenneth I. Kaitin and Christopher P. Milne. Forum: A dearth of new meds. Scientific American, 305(2):16, August 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-16.html;   http://www.nature.
com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
16.pdf.

Requarth:2011:AMI

[653] Tim Requarth. Advances: Mental illness in a dish. Scientific American, 305(2):18, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
18.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-18.pdf.

Hodge:2011:SBT

[654] Anne-Marie Hodge. A skill better than Rudolph’s. Scientific American, 305(2):20, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
20a.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-20a.pdf.

Matson:2011:ANP

[655] John Matson. Ageless. no plastic surgery required. Scientific American, 305(2):20, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
20b.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-20b.pdf.

Diep:2011:PT

[656] Francie Diep. Parsing the twitterverse. Scientific American, 305(2):22, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-22a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
22a.pdf.

Chin:2011:WIh

[657] Ann Chin. What is it? Scientific American, 305(2):22, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-22b.html;   http://www.nature.
com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
22b.pdf.

Arnold:2011:AO

[658] Carrie Arnold. Anatomy of an outbreak. Scientific American, 305(2):24, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-24a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
24a.pdf.

Carmichael:2011:ATN

[659] Mary Carmichael. All together now. Scientific American, 305(2):24, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-24b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
24b.pdf.

Gorman:2011:CM

[660] Christine Gorman. E. coli on the march. Scientific American, 305(2):26, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-26a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
26a.pdf.

Hackett:2011:NSf

[661] George Hackett. News scan. Scientific American, 305(2):26, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-26b.html;   http://www.nature.
com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
26b.pdf.

Wapner:2011:GV

[662] Jessica Wapner. Going viral. Scientific American, 305(2):28, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-28a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
28a.pdf.

Anonymous:2011:Qo

[663] Anonymous. Quotable. Scientific American, 305(2):28, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-28b.html;   http://www.nature.
com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
28b.pdf.

Diep:2011:MWO

[664] Francie Diep. Mouth wide open. Scientific American, 305(2):29, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-29a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
29a.pdf.

Anonymous:2011:Sc

[665] Anonymous. Stat. Scientific American, 305(2):29, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-29b.html;   http://www.nature.

com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
29b.pdf.

Seife:2011:MRS

[666] Charles Seife. The mind-reading salmon. Scientific American, 305(2):30, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-30a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
30a.pdf.

Venkataramanan:2011:WHC

[667] Madhumita Venkataramanan. Where house cats roam. Scientific American, 305(2):30, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
30b.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-30b.pdf.

Gibbs:2011:CC

[668] W. Wayt Gibbs and Nathan Myhrvold. Cryogenic cooking. Scientific American, 305(2):31, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
31.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-31.pdf.

Willyard:2011:SHH

[669] Cassandra Willyard. The science of health: How do tumors grow? Scientific American, 305(2):32–34, August 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-32.html;   http://www.nature.
com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
32.pdf.

Pogue:2011:TPC

[670] David Pogue. TechnoFiles: The perils of copy protection. Scientific American, 305(2):36–37, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
36.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-36.pdf.

Wong:2011:RRRd

[671] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 305(2):88, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
88.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-88.pdf.

Shermer:2011:SG

[672] Michael Shermer. Skeptic: Globaloney. Scientific American, 305(2):90, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n2/full/scientificamerican0811-90.html;   http:
//www.nature.com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
90.pdf.

Mirsky:2011:AGBb

[673] Steve Mirsky. Anti gravity: Out of the brains of babes. Scientific American, 305(2):92, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
92.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-92.pdf.

Fischetti:2011:GSBb

[674] Mark Fischetti. Graphic science: Bombarded. Scientific American, 305(2):96, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
96.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-96.pdf.

Ellis:2011:CDM

[675] George F. R. Ellis. Cosmology: Does the multiverse really exist? Scientific American, 305(2):38–43, August 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-38.html;   http://www.nature.
com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
38.pdf.

Caspari:2011:HOE

[676] Rachel Caspari. Human origins: The evolution of grandparents. Scientific American, 305(2):44–49, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n2/full/scientificamerican0811-
44.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-44.pdf.

Stix:2011:NHB

[677] Gary Stix. Neuroscience: How to build a better learner. Scientific American, 305(2):50–57, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
50.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-50.pdf.

Biello:2011:EFP

[678] David Biello. Energy: The false promise of biofuels. Scientific American, 305(2):58–65, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
58.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-58.pdf.

Bai:2011:LST

[679] Nina Bai. Life science: Treasure in the trees. Scientific American, 305(2):66–67, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
66.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-66.pdf.

Rowe:2011:MBF

[680] Steven M. Rowe, J. P. Clancy, and Eric J. Sorscher. Medicine: A breath of fresh air. Scientific American, 305(2):68–73, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n2/full/scientificamerican0811-68.html;   http://www.nature.
com/scientificamerican/journal/v305/n2/pdf/scientificamerican0811-
68.pdf.

Zimring:2011:SHN

[681] Franklin E. Zimring. Sociology: How New York beat crime. Scientific American, 305(2):74–79, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
74.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-74.pdf.

Livio:2011:PSW

[682] Mario Livio. Philosophy of science: Why math works. Scientific American, 305(2):80–83, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
80.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-80.pdf.

Bartholet:2011:AGP

[683] Jeffrey Bartholet. Archaeology: Guardian of the pharoahs. Scientific American, 305(2):84–87, August 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n2/full/scientificamerican0811-
84.html;   http://www.nature.com/scientificamerican/journal/
v305/n2/pdf/scientificamerican0811-84.pdf.

DiChristina:2011:ECL

[684] Mariette DiChristina. From the Editor: City lights. Scientific American, 305(3):6, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
6.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-6.pdf.

Anonymous:2011:Li

[685] Anonymous. Letters. Scientific American, 305(3):8–10, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n3/full/scientificamerican0911-8.html;   http://www.nature.
com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
8.pdf.

Anonymous:2011:SAFb

[686] Anonymous. Science agenda: In fairness to cities. Scientific American, 305(3):14, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
14.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-14.pdf.

Bloomberg:2011:FBB

[687] Michael R. Bloomberg. Forum: The best and the brightest. Scientific American, 305(3):16, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
16.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-16.pdf.

Katsnelson:2011:ASC

[688] Alla Katsnelson. Advances: The stress of crowds. Scientific American, 305(3):18, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
18.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-18.pdf.

Interlandi:2011:NHS

[689] Jeneen Interlandi. New help for smokers. Scientific American, 305(3):20, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-20a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
20a.pdf.

Piore:2011:PWd

[690] Adam Piore. Patent watch. Scientific American, 305(3):20, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-20b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
20b.pdf.

Matson:2011:WI

[691] John Matson. What is it? Scientific American, 305(3):21, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-21.html;   http:

//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
21.pdf.

Matson:2011:HSP

[692] John Matson. Her summer pastime? cancer research. Scientific American, 305(3):22, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
22.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-22.pdf.

Gibbs:2011:CS

[693] W. Wayt Gibbs and Nathan Myhrvold. Cooking that sucks. Scientific American, 305(3):23, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
23.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-23.pdf.

Choi:2011:CYS

[694] Charles Q. Choi. Can you see me now? Scientific American, 305(3):24, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-24a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
24a.pdf.

Raymond:2011:SN

[695] Joan Raymond. The shape of a nose. Scientific American, 305(3):24, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-24b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
24b.pdf.

Greenemeier:2011:ASA

[696] Larry Greenemeier. After shock and awe. Scientific American, 305(3):26, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-26a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
26a.pdf.

Hackett:2011:NSg

[697] George Hackett. News scan. Scientific American, 305(3):26, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-26b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
26b.pdf.

Katsnelson:2011:PPT

[698] Alla Katsnelson. The pitfalls of positive thinking. Scientific American, 305(3):28, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
28a.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-28a.pdf.

Anonymous:2011:Sd

[699] Anonymous. Stat. Scientific American, 305(3):28, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n3/full/scientificamerican0911-28b.html;   http://www.nature.
com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
28b.pdf.

Choi:2011:LBM

[700] Charles Q. Choi. Less bang, more bubbles. Scientific American, 305(3):28, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
28c.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-28c.pdf.

Diep:2011:CNR

[701] Francie Diep. Cocaine’s newest risks. Scientific American, 305(3):30, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-30.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
30.pdf.

Requarth:2011:OBF

[702] Tim Requarth. One brainy fish. Scientific American, 305(3):31, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-31.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
31.pdf.

Daley:2011:SHS

[703] Matthew F. Daley and Jason M. Glanz. The science of health: Straight talk about vaccination. Scientific American, 305(3):32–34, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-32.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
32.pdf.

Pogue:2011:TPP

[704] David Pogue. TechnoFiles: Password prevented. Scientific American, 305(3):36–37, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
36.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-36.pdf.

Wong:2011:RRRe

[705] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 305(3):90, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
90.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-90.pdf.

Shermer:2011:SWP

[706] Michael Shermer. Skeptic: What is pseudoscience? Scientific American, 305(3):92, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
92.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-92.pdf.

Mirsky:2011:AGN

[707] Steve Mirsky. Anti gravity: Noble Nobel faces. Scientific American, 305(3):94, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-

94.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-94.pdf.

Schlenoff:2011:YAIi

[708] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 305(3):95, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-95.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
95.pdf.

Fischetti:2011:GSH

[709] Mark Fischetti. Graphic science: The human cost of energy. Scientific American, 305(3):96, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
96.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-96.pdf.

Anonymous:2011:CSG

[710] Anonymous. Cities: Smarter, greener, better: Street-savvy. Scientific American, 305(3):38–41, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n3/full/scientificamerican0911-
38.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-38.pdf.

Ratti:2011:SSN

[711] Carlo Ratti and Anthony Townsend. Smarter: The social nexus. Scientific American, 305(3):42–48, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n3/full/scientificamerican0911-
42.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-42.pdf.

Glaeser:2011:EI

[712] Edward Glaeser. Engines of innovation. Scientific American, 305(3):50–55, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
50.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-50.pdf.

Bettencourt:2011:BCD

[713] Luís M. A. Bettencourt and Geoffrey B. West. Bigger cities do more with less. Scientific American, 305(3):52–53, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n3/full/scientificamerican0911-52.html;   http://www.nature.
com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
52.pdf.

Neuwirth:2011:GB

[714] Robert Neuwirth. Global bazaar. Scientific American, 305(3):56–63, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-56.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
56.pdf.

McGranahan:2011:BHU

[715] Gordon McGranahan and David Satterthwaite. Better health for the uncounted urban masses. Scientific American, 305(3):62, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n3/full/scientificamerican0911-62.html;   http://www.nature.
com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
62.pdf.

Glaeser:2011:BB

[716] Edward Glaeser. Brains over buildings. Scientific American, 305(3):64, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-64.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
64.pdf.

Biello:2011:GHG

[717] David Biello. Greener: How green is my city. Scientific American, 305(3):66–69, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
66.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-66.pdf.

Rosenzweig:2011:ACL

[718] Cynthia Rosenzweig. All climate is local. Scientific American, 305(3):70–73, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
70.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-70.pdf.

Fischetti:2011:EC

[719] Mark Fischetti. The efficient city. Scientific American, 305(3):74–75, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-74.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
74.pdf.

Lamster:2011:BCA

[720] Mark Lamster. Better: Castles in the air. Scientific American, 305(3):76–83, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
76.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-76.pdf.

Easter:2011:ST

[721] Michael Easter and Gary Stix. Street talk. Scientific American, 305(3):84–87, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n3/full/scientificamerican0911-
84.html;   http://www.nature.com/scientificamerican/journal/
v305/n3/pdf/scientificamerican0911-84.pdf.

Gibson:2011:LMC

[722] William Gibson. Life in the meta city. Scientific American, 305(3):88–89, September 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n3/full/scientificamerican0911-88.html;   http:
//www.nature.com/scientificamerican/journal/v305/n3/pdf/scientificamerican0911-
88.pdf.

DiChristina:2011:EBS

[723] Mariette DiChristina. From the Editor: Back to school. Scientific American, 305(4):6, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
6.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-6.pdf.

Anonymous:2011:Lj

[724] Anonymous. Letters. Scientific American, 305(4):8–10, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n4/full/scientificamerican1011-8.html;   http://www.nature.
com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
8.pdf.

Schlenoff:2011:YAIj

[725] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 305(4):94, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-94.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
94.pdf.

Anonymous:2011:SAB

[726] Anonymous. Science agenda: Ban chimp testing. Scientific American, 305(4):12, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
12.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-12.pdf.

Hunt:2011:FTA

[727] Patricia Hunt. Forum: Toxins all around us. Scientific American, 305(4): 14, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-14.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
14.pdf.

Begley:2011:AMT

[728] Sharon Begley. Advances: More than child’s play. Scientific American, 305(4):16, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-

16.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-16.pdf.

Moyer:2011:ISD

[729] Melinda Wenner Moyer. Is it safe to drink? Scientific American, 305(4): 18, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-18a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
18a.pdf.

Piore:2011:PWe

[730] Adam Piore. Patent watch. Scientific American, 305(4):18, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-18b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
18b.pdf.

Chin:2011:WIi

[731] Ann Chin. What is it? Scientific American, 305(4):19, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n4/full/scientificamerican1011-19.html;   http://www.nature.
com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
19.pdf.

Diep:2011:OC

[732] Francie Diep. Outsmarting cancer. Scientific American, 305(4):20, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-20a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
20a.pdf.

Anonymous:2011:Se

[733] Anonymous. Stat. Scientific American, 305(4):20, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n4/full/scientificamerican1011-20b.html;   http://www.nature.
com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
20b.pdf.

Diep:2011:CS

[734] Francie Diep. Clearing the smoke. Scientific American, 305(4):21, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-21.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
21.pdf.

Johnson:2011:IGY

[735] Eric Michael Johnson. “i’ve got your back”. Scientific American, 305(4): 22, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-22a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
22a.pdf.

Hackett:2011:NSh

[736] George Hackett. News scan. Scientific American, 305(4):22, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-22b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
22b.pdf.

Choi:2011:IHC

[737] Charles Q. Choi. Instant health checks for buildings and bridges. Scientific American, 305(4):24, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
24a.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-24a.pdf.

Anonymous:2011:Sf

[738] Anonymous. Stat. Scientific American, 305(4):24, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n4/full/scientificamerican1011-24b.html;   http://www.nature.
com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
24b.pdf.

Moyer:2011:TA

[739] Michael Moyer. The trouble with armor. Scientific American, 305(4):25, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-25.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
25.pdf.

Castelvecchi:2011:TMT

[740] Davide Castelvecchi. A tale of math treasure. Scientific American, 305(4):26, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
26.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-26.pdf.

Greenemeier:2011:GUN

[741] Larry Greenemeier. Gig.u is now in session. Scientific American, 305(4): 27, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-27a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
27a.pdf.

Anonymous:2011:Qp

[742] Anonymous. Quotable. Scientific American, 305(4):27, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n4/full/scientificamerican1011-27b.html;   http://www.nature.
com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
27b.pdf.

Gibbs:2011:SE

[743] W. Wayt Gibbs and Nathan Myhrvold. Spherical eats. Scientific American, 305(4):28, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
28a.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-28a.pdf.

Svoboda:2011:PDA

[744] Elizabeth Svoboda. Putting diabetes on autopilot. Scientific American, 305(4):28, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-

28b.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-28b.pdf.

Blue:2011:SHE

[745] Laura Blue. The science of health: The ethnic health advantage. Scientific American, 305(4):30–32, October 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n4/full/scientificamerican1011-30.html;   http://www.nature.
com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
30.pdf.

Pogue:2011:TBP

[746] David Pogue. TechnoFiles: Big progress on the little things. Scientific American, 305(4):34–35, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
34.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-34.pdf.

Anonymous:2011:RRRe

[747] Anonymous. Reviews and recommendations: Recommended. Scientific American, 305(4):88, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
88.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-88.pdf.

Shermer:2011:SDV

[748] Michael Shermer. Skeptic: The decline of violence. Scientific American, 305(4):90, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
90.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-90.pdf.

Mirsky:2011:AGS

[749] Steve Mirsky. Anti gravity: Surface tension. Scientific American, 305(4): 92, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-92.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
92.pdf.

Fischetti:2011:GSDb

[750] Mark Fischetti. Graphic science: Data theft: Hackers attack. Scientific American, 305(4):100, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
100.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-100.pdf.

Blitz:2011:ADS

[751] Leo Blitz. Astrophysics: The dark side of the Milky Way. Scientific American, 305(4):36–45, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
36.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-36.pdf.

Anonymous:2011:SYC

[752] Anonymous. Special year of chemistry celebration: Atom power. Scientific American, 305(4):46–47, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n4/full/scientificamerican1011-
46.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-46.pdf.

Ball:2011:UM

[753] Philip Ball. 10 unsolved mysteries. Scientific American, 305(4):48–53, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n4/full/scientificamerican1011-48.html;   http:
//www.nature.com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
48.pdf.

Blum:2011:SYT

[754] Deborah Blum. The scent of your thoughts. Scientific American, 305(4):54–57, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
54.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-54.pdf.

Simpson:2011:GAB

[755] Sarah Simpson. Geology: Afghanistan’s buried riches. Scientific American, 305(4):58–65, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
58.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-58.pdf.

vonHofe:2011:MNA

[756] Eric von Hofe. Medicine: A new ally against cancer. Scientific American, 305(4):66–71, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
66.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-66.pdf.

Kuchment:2011:FHS

[757] Anna Kuchment and staff editor. Forensics: How skulls speak. Scientific American, 305(4):72–73, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
72.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-72.pdf.

Folger:2011:PWH

[758] Tim Folger. Physics: Waiting for the Higgs. Scientific American, 305(4):74–79, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-
74.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-74.pdf.

Dyke:2011:PDB

[759] Gareth Dyke. Paleontology: The Dinosaur Baron of Transylvania. Scientific American, 305(4):80–83, October 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n4/full/scientificamerican1011-80.html;   http://www.nature.
com/scientificamerican/journal/v305/n4/pdf/scientificamerican1011-
80.pdf.

Singer:2011:BAC

[760] Thea Singer. Biology: Actuary of the cell. Scientific American, 305(4):84–87, October 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n4/full/scientificamerican1011-

84.html;   http://www.nature.com/scientificamerican/journal/
v305/n4/pdf/scientificamerican1011-84.pdf.

DiChristina:2011:EPT

[761] Mariette DiChristina. From the Editor: Paths taken. Scientific American, 305(5):6, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
6.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-6.pdf.

Anonymous:2011:Lk

[762] Anonymous. Letters. Scientific American, 305(5):8–10, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n5/full/scientificamerican1111-8.html;   http://www.nature.
com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
8.pdf.

Schlenoff:2011:YAIk

[763] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 305(5):98, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-98.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
98.pdf.

Anonymous:2011:SASb

[764] Anonymous. Science agenda: Safety first, fracking second. Scientific American, 305(5):12, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
12.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-12.pdf.

Borg:2011:FCN

[765] Scott Borg. Forum: A cybersecurity nightmare. Scientific American, 305(5):14, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
14.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-14.pdf.

Matson:2011:ATS

[766] John Matson. Advances: On the trail of space trash. Scientific American, 305(5):18, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
18.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-18.pdf.

Paulos:2011:BPM

[767] John Allen Paulos. Bigger plates, more food — or is it the other way around? Scientific American, 305(5):20, November 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n5/full/scientificamerican1111-20.html;   http://www.nature.
com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
20.pdf.

Godkin:2011:OS

[768] David Godkin. Olympians of the sky. Scientific American, 305(5):21, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-21.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
21.pdf.

Kuchment:2011:ZI

[769] Anna Kuchment. Zombie insects. Scientific American, 305(5):22, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-22.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
22.pdf.

Interlandi:2011:SM

[770] Jeneen Interlandi. Social medicine. Scientific American, 305(5):23, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-23.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
23.pdf.

Biello:2011:HH

[771] David Biello. Halting hurricanes. Scientific American, 305(5):24, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-24a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
24a.pdf.

Anonymous:2011:SSa

[772] Anonymous. Suspect science. Scientific American, 305(5):24, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-24b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
24b.pdf.

Hodge:2011:MYN

[773] Anne-Marie Hodge. Meet your newest ancestor. Scientific American, 305(5):26, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
26a.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-26a.pdf.

Anonymous:2011:SSb

[774] Anonymous. Suspect science. Scientific American, 305(5):26, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-26b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
26b.pdf.

Greenemeier:2011:DHG

[775] Larry Greenemeier. 3-D, hold the glasses. Scientific American, 305(5):27, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-27.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
27.pdf.

Goldman:2011:FMF

[776] Jason G. Goldman. Of flash mobs and Four Loko. Scientific American, 305(5):28, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
28a.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-28a.pdf.

Choi:2011:IR

[777] Charles Q. Choi. Instant recall. Scientific American, 305(5):28, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-28b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
28b.pdf.

Galef:2011:YSE

[778] Julia Galef. You say embryo, I say parthenote. Scientific American, 305(5):30, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
30a.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-30a.pdf.

Nelson:2011:LTH

[779] Bryn Nelson. Low taxes, high rhetoric. Scientific American, 305(5):30, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-30b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
30b.pdf.

Chin:2011:WIj

[780] Ann Chin. What is it? Scientific American, 305(5):31, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n5/full/scientificamerican1111-31.html;   http://www.nature.
com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
31.pdf.

Diep:2011:SHC

[781] Francie Diep. The science of health: Cholesterol conundrum. Scientific American, 305(5):32–33, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n5/full/scientificamerican1111-

32.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-32.pdf.

Pogue:2011:TJY

[782] David Pogue. TechnoFiles: The joke’s on your computer. Scientific American, 305(5):34–35, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n5/full/scientificamerican1111-
34.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-34.pdf.

Anonymous:2011:RRRf

[783] Anonymous. Reviews and recommendations: Recommended. Scientific American, 305(5):92, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
92.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-92.pdf.

Shermer:2011:SRS

[784] Michael Shermer. Skeptic: The real science behind scientology. Scientific American, 305(5):94, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
94.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-94.pdf.

Mirsky:2011:AGBc

[785] Steve Mirsky. Anti gravity: Balancing act. Scientific American, 305(5):96, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
96.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-96.pdf.

Fischetti:2011:GSCb

[786] Mark Fischetti. Graphic science: Computers vs. brains. Scientific American, 305(5):104, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
104.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-104.pdf.

Pringle:2011:AFA

[787] Heather Pringle. Archaeology: The first Americans. Scientific American, 305(5):36–45, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
36.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-36.pdf.

Smith:2011:PSD

[788] Peter H. Smith. Planetary science: Digging mars. Scientific American, 305(5):46–54, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
46.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-46.pdf.

Warmflash:2011:EAS

[789] David Warmflash. Experimental astrobiology: The smallest astronauts. Scientific American, 305(5):54–55, November 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n5/full/scientificamerican1111-54.html;   http://www.nature.
com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
54.pdf.

Knobe:2011:EPT

[790] Joshua Knobe. Experimental philosophy: Thought experiments. Scientific American, 305(5):56–59, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n5/full/scientificamerican1111-
56.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-56.pdf.

Foley:2011:SCW

[791] Jonathan A. Foley. Sustainability: Can we feed the world & sustain the planet? Scientific American, 305(5):60–65, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n5/full/scientificamerican1111-60.html;   http://www.nature.
com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
60.pdf.

Perkins:2011:GSG

[792] Sid Perkins. Geology: Sleeping giant. Scientific American, 305(5):66–67, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n5/full/scientificamerican1111-66.html;   http:
//www.nature.com/scientificamerican/journal/v305/n5/pdf/scientificamerican1111-
66.pdf.

Trivedi:2011:BWG

[793] Bijal P. Trivedi. Biotechnology: The wipeout gene. Scientific American, 305(5):68–75, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
68.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-68.pdf.

Freedman:2011:CFE

[794] David H. Freedman. Computing: A formula for economic calamity. Scientific American, 305(5):76–79, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v305/n5/full/scientificamerican1111-
76.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-76.pdf.

Mooney:2011:ETA

[795] Chris Mooney. Energy: The truth about fracking. Scientific American, 305(5):80–85, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
80.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-80.pdf.

Borrell:2011:DMS

[796] Brendan Borrell. Diagnostics: The medical sleuth. Scientific American, 305(5):86–90, November 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n5/full/scientificamerican1111-
86.html;   http://www.nature.com/scientificamerican/journal/
v305/n5/pdf/scientificamerican1111-86.pdf.

DiChristina:2011:EEG

[797] Mariette DiChristina. From the Editor: Emergent genius. Scientific American, 305(6):6, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
6.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-6.pdf.

Anonymous:2011:Ll

[798] Anonymous. Letters. Scientific American, 305(6):8–10, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n6/full/scientificamerican1211-8.html;   http://www.nature.
com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
8.pdf.

Schlenoff:2011:YAIl

[799] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 305(6):106, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-106.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
106.pdf.

Anonymous:2011:SASc

[800] Anonymous. Science agenda: Stop the genetic dragnet. Scientific American, 305(6):14, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
14.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-14.pdf.

Lu:2011:FSK

[801] Edward T. Lu. Forum: Stop the killer rocks. Scientific American, 305(6):16, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
16.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-16.pdf.

Begley:2011:ACT

[802] Sharon Begley. Advances: Can’t touch this feeling. Scientific American, 305(6):18, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-

18.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-18.pdf.

Gibbs:2011:MSL

[803] W. Wayt Gibbs and Nathan Myhrvold. Microwaves and the speed of light. Scientific American, 305(6):20, December 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n6/full/scientificamerican1211-20a.html;   http://www.nature.
com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
20a.pdf.

Anonymous:2011:SSc

[804] Anonymous. Suspect science. Scientific American, 305(6):20, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-20b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
20b.pdf.

Requarth:2011:CEC

[805] Tim Requarth and Greg Wayne. A circuit in every cell. Scientific American, 305(6):21, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
21.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-21.pdf.

Hodge:2011:VMM

[806] Anne-Marie C. Hodge. Vitamins, minerals and MicroRNA. Scientific American, 305(6):22, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
22a.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-22a.pdf.

Choi:2011:FDC

[807] Charles Q. Choi. Fluid dynamics in a cup. Scientific American, 305(6):22, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-22b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
22b.pdf.

Zittrain:2011:FF

[808] Jonathan Zittrain. Freedom fighter. Scientific American, 305(6):24, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-24.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
24.pdf.

Kuchment:2011:YGT

[809] Anna Kuchment. This year, give them brains. Scientific American, 305(6):26–27, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
26.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-26.pdf.

Castelvecchi:2011:WNM

[810] Davide Castelvecchi. Why neutrinos might wimp out. Scientific American, 305(6):28, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
28a.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-28a.pdf.

Anonymous:2011:YT

[811] Anonymous. Yawn of the tortoise. Scientific American, 305(6):28, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-28b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
28b.pdf.

Moyer:2011:UA

[812] Michael Moyer. Universal alignment. Scientific American, 305(6):30, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-30.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
30.pdf.

Eveleth:2011:PP

[813] Rose Eveleth. From pollen to polyester. Scientific American, 305(6):32, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-32.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
32.pdf.

Biello:2011:WI

[814] David Biello. What is it? Scientific American, 305(6):33, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n6/full/scientificamerican1211-33.html;   http://www.nature.
com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
33.pdf.

McKenna:2011:SHS

[815] Maryn McKenna. The science of health: Swapping germs. Scientific American, 305(6):34–36, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n6/full/scientificamerican1211-
34.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-34.pdf.

Pogue:2011:THS

[816] David Pogue. TechnoFiles: How to see the invisible. Scientific American, 305(6):38–39, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
38.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-38.pdf.

Anonymous:2011:RRRg

[817] Anonymous. Reviews and recommendations: Recommended. Scientific American, 305(6):100, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
100.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-100.pdf.

Shermer:2011:SSS

[818] Michael Shermer. Skeptic: Sacred salubriousness. Scientific American, 305(6):102, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-

102.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-102.pdf.

Mirsky:2011:AGRb

[819] Steve Mirsky. Anti gravity: Respect for evidence. Scientific American, 305(6):104, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
104.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-104.pdf.

Fischetti:2011:GSL

[820] Mark Fischetti. Graphic science: The links we love. Scientific American, 305(6):110, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
110.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-110.pdf.

Anonymous:2011:TSR

[821] Anonymous. Technology special report: World changing ideas. Scientific American, 305(6):40–41, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n6/full/scientificamerican1211-
40.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-40.pdf.

Svoboda:2011:FHM

[822] Elizabeth Svoboda. The forever health monitor. Scientific American, 305(6):42, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
42.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-42.pdf.

Mims:2011:CTL

[823] Christopher Mims. A chip that thinks like a brain. Scientific American, 305(6):43, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
43.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-43.pdf.

Mims:2011:WYS

[824] Christopher Mims. The wallet in your skin. Scientific American, 305(6):44, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
44.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-44.pdf.

Diep:2011:CDF

[825] Francie Diep. Computers that don’t freeze up. Scientific American, 305(6):45, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
45.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-45.pdf.

Peck:2011:CB

[826] Morgen Peck. Currency without borders. Scientific American, 305(6):46, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-46a.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
46a.pdf.

Fecht:2011:MM

[827] Sarah Fecht. Microbe miners. Scientific American, 305(6):46, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v305/n6/full/scientificamerican1211-46b.html;   http:
//www.nature.com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
46b.pdf.

Mims:2011:CDN

[828] Christopher Mims. Crops that don’t need replanting. Scientific American, 305(6):48, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
48.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-48.pdf.

Mims:2011:LFE

[829] Christopher Mims. Liquid fuel for electric cars. Scientific American, 305(6):50–51, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
50.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-50.pdf.

Svoboda:2011:NSG

[830] Elizabeth Svoboda. Nano-size germ killers. Scientific American, 305(6):51, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
51.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-51.pdf.

Weinberger:2011:MWP

[831] David Weinberger. The machine that would predict the future. Scientific American, 305(6):52–57, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n6/full/scientificamerican1211-
52.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-52.pdf.

Landau:2011:SEW

[832] Damon Landau and Nathan J. Strange. Space exploration: This way to mars. Scientific American, 305(6):58–65, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n6/full/scientificamerican1211-58.html;   http://www.nature.
com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
58.pdf.

Stix:2011:BDM

[833] Gary Stix and staff writer. Biology: Dazzling miniatures. Scientific American, 305(6):66–71, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n6/full/scientificamerican1211-
66.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-66.pdf.

Carey:2011:CCA

[834] John A. Carey. Climate change: After the deluge. Scientific American, 305(6):72–75, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-

72.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-72.pdf.

Nestler:2011:NHS

[835] Eric J. Nestler. Neuroscience: Hidden switches in the mind. Scientific American, 305(6):76–83, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n6/full/scientificamerican1211-
76.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-76.pdf.

Moffett:2011:ABA

[836] Mark W. Moffett. Animal behavior: Ants & the art of war. Scientific American, 305(6):84–89, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v305/n6/full/scientificamerican1211-
84.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-84.pdf.

Fitzpatrick:2011:FAI

[837] Colleen Fitzpatrick. Forensics: Arm in the ice. Scientific American, 305(6):90–93, December 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v305/n6/full/scientificamerican1211-
90.html;   http://www.nature.com/scientificamerican/journal/
v305/n6/pdf/scientificamerican1211-90.pdf.

Borrell:2011:ESQ

[838] Brendan Borrell. Education: Speaking out on the “quiet crisis”. Scientific American, 305(6):94–99, December 2011. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v305/
n6/full/scientificamerican1211-94.html;   http://www.nature.
com/scientificamerican/journal/v305/n6/pdf/scientificamerican1211-
94.pdf.

DiChristina:2012:ESSa

[839] Mariette DiChristina. From the Editor: The science of staying young. Scientific American, 306(1):6, January 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n1/full/scientificamerican0112-6.html;   http://www.nature.

com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
6.pdf.

Anonymous:2012:Lb

[840] Anonymous. Letters. Scientific American, 306(1):8–10, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n1/full/scientificamerican0112-8.html;   http://www.nature.
com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
8.pdf.

Schlenoff:2012:YAIa

[841] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 306(1):87, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-87.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
87.pdf.

Begley:2012:AAB

[842] Sharon Begley. Advances: Anything boys can do …. Scientific American, 306(1):25, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
25.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-25.pdf.

Biello:2012:TP

[843] David Biello. Tweaking photosynthesis. Scientific American, 306(1):21, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-21.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
21.pdf.

Biello:2012:PSR

[844] David Biello. Pond scum to the rescue. Scientific American, 306(1):22, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-22a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
22a.pdf.

Chin:2012:WIa

[845] Ann Chin. What is it? Scientific American, 306(1):22, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n1/full/scientificamerican0112-22b.html;   http://www.nature.
com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
22b.pdf.

Corbyn:2012:NSC

[846] Zoë Corbyn. No spring chickens. Scientific American, 306(1):19, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-19.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
19.pdf.

Eveleth:2012:LEC

[847] Rose Eveleth. The little engine that could. Scientific American, 306(1): 24, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-24b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
24b.pdf.

Franklin:2012:SHH

[848] Deborah Franklin. The science of health: Healing Kansas. Scientific American, 306(1):28–30, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
28.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-28.pdf.

Gleick:2012:FCM

[849] Peter H. Gleick and Matthew Heberger. Forum: The coming mega drought. Scientific American, 306(1):14, January 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n1/full/scientificamerican0112-14.html;   http://www.nature.
com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
14.pdf.

Little:2012:APC

[850] Jane Braxton Little. Advances: Plight of the condors. Scientific American, 306(1):18, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
18a.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-18a.pdf.

Matson:2012:WRS

[851] John Matson. Why it’s raining satellites. Scientific American, 306(1):26, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-26a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
26a.pdf.

Matson:2012:GSB

[852] John Matson. Graphic science: Buying a better bulb. Scientific American, 306(1):92, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
92.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-92.pdf.

Mirsky:2012:AGP

[853] Steve Mirsky. Anti gravity: Physics uncowed. Scientific American, 306(1):86, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
86.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-86.pdf.

Ouellette:2012:AYB

[854] Jennifer Ouellette. Advances: Your birdsong stays on my mind. Scientific American, 306(1):26, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
26b.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-26b.pdf.

Paulos:2012:WP

[855] John Allen Paulos. Weighing the positives. Scientific American, 306(1): 20, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-20.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
20.pdf.

Piore:2012:PW

[856] Adam Piore. Patent watch. Scientific American, 306(1):23, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-23.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
23.pdf.

Pogue:2012:TSS

[857] David Pogue. TechnoFiles: Silicon superego. Scientific American, 306(1): 31, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-31.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
31.pdf.

Shermer:2012:SY

[858] Michael Shermer. Skeptic: In the year 9595. Scientific American, 306(1): 85, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-85.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
85.pdf.

Stix:2012:APM

[859] Gary Stix. Advances: Photographic memory. Scientific American, 306(1):24, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
24a.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-24a.pdf.

Villasenor:2012:HCD

[860] John Villasenor. Here come the drones. Scientific American, 306(1):16, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-16.html;   http:

//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
16.pdf.

Wong:2012:RRR

[861] Kate Wong. Reviews and recommendations: Recommended. Scientific American, 306(1):84, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
84.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-84.pdf.

Anonymous:2012:SAR

[862] Anonymous. Science agenda: Read my e-mail? get a warrant. Scientific American, 306(1):12, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
12.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-12.pdf.

Anonymous:2012:ASS

[863] Anonymous. Advances: Suspect science. Scientific American, 306(1):18, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-18b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
18b.pdf.

Borrell:2012:ED

[864] Brendan Borrell. Environment: Dust up. Scientific American, 306(1):80– 83, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n1/full/scientificamerican0112-80.html;   http:
//www.nature.com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
80.pdf.

Castelvecchi:2012:ABC

[865] Davide Castelvecchi. Animal behavior: The compass within. Scientific American, 306(1):48–53, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
48.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-48.pdf.

Gorman:2012:HFH

[866] Christine Gorman. Health: Five hidden dangers of … obesity. Scientific American, 306(1):60–61, January 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n1/full/scientificamerican0112-60.html;   http://www.nature.
com/scientificamerican/journal/v306/n1/pdf/scientificamerican0112-
60.pdf.

Greene:2012:CGB

[867] Jenny E. Greene. Cosmology: Goldilocks black holes. Scientific American, 306(1):40–47, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
40.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-40.pdf.

Harmon:2012:IPS

[868] Katherine Harmon. Immunology: The patient scientist. Scientific American, 306(1):54–59, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
54.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-54.pdf.

Nussenzveig:2012:PSG

[869] H. Moysés Nussenzveig. Physics: The science of the glory. Scientific American, 306(1):68–73, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
68.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-68.pdf.

Stipp:2012:BNP

[870] David Stipp. Biology: A new path to longevity. Scientific American, 306(1):32–39, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
32.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-32.pdf.

Vlahos:2012:TDP

[871] James Vlahos. Technology: The department of pre-crime. Scientific American, 306(1):62–67, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
62.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-62.pdf.

Webber:2012:SMF

[872] Michael E. Webber. Sustainability: More food, less energy. Scientific American, 306(1):74–79, January 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n1/full/scientificamerican0112-
74.html;   http://www.nature.com/scientificamerican/journal/
v306/n1/pdf/scientificamerican0112-74.pdf.

DiChristina:2012:ESSb

[873] Mariette DiChristina. From the Editor: A sea of spacetime foam? Scientific American, 306(2):4, February 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n2/full/scientificamerican0212-4.html;   http://www.nature.
com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
4.pdf.

Anonymous:2012:Lc

[874] Anonymous. Letters. Scientific American, 306(2):6–8, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n2/full/scientificamerican0212-6.html;   http://www.nature.
com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
6.pdf.

Schlenoff:2012:YAIb

[875] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 306(2):87, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n2/full/scientificamerican0212-87.html;   http:
//www.nature.com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
87.pdf.

Anonymous:2012:SAU

[876] Anonymous. Science agenda: Unschooled in hard knocks. Scientific American, 306(2):10, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
10.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-10.pdf.

Kaplan:2012:FSP

[877] David Kaplan. Forum: Science and prejudice. Scientific American, 306(2):11, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
11.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-11.pdf.

Matson:2012:ATT

[878] John Matson. Advances: Tensions over Taikonauts. Scientific American, 306(2):12, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
12.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-12.pdf.

Interlandi:2012:MMC

[879] Jeneen Interlandi. A man-made contagion. Scientific American, 306(2):14, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
14a.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-14a.pdf.

Gaskill:2012:IDE

[880] Melissa Gaskill. It detects earthquakes and lactose intolerance. Scientific American, 306(2):14, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
14b.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-14b.pdf.

Chin:2012:WIb

[881] Ann Chin. What is it? Scientific American, 306(2):15, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n2/full/scientificamerican0212-15.html;   http://www.nature.
com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
15.pdf.

Pycha:2012:JC

[882] Anne Pycha. Just a click away. Scientific American, 306(2):16, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n2/full/scientificamerican0212-16a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
16a.pdf.

Matson:2012:HH

[883] John Matson. The not so hot hand. Scientific American, 306(2):16, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n2/full/scientificamerican0212-16b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
16b.pdf.

Raymond:2012:MMW

[884] Joan Raymond. Mom is my wingman. Scientific American, 306(2):18, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n2/full/scientificamerican0212-18a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
18a.pdf.

Eveleth:2012:PM

[885] Rose Eveleth. Peeling away microbes. Scientific American, 306(2):18, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n2/full/scientificamerican0212-18b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
18b.pdf.

Jabr:2012:GA

[886] Ferris Jabr. Gumming up appetite. Scientific American, 306(2):20, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n2/full/scientificamerican0212-20a.html;   http:

//www.nature.com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
20a.pdf.

Harmon:2012:LFN

[887] Katherine Harmon. A long flight but no baggage. Scientific American, 306(2):20, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
20b.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-20b.pdf.

Anonymous:2012:Sa

[888] Anonymous. Stat. Scientific American, 306(2):20, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n2/full/scientificamerican0212-20c.html;   http://www.nature.
com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
20c.pdf.

Upson:2012:IMV

[889] Sandra Upson. Inside the mind of a video game champ. Scientific American, 306(2):21, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
21.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-21.pdf.

Gibbs:2012:MLG

[890] W. Wayt Gibbs and Nathan Myhrvold. Making liquids go bipolar. Scientific American, 306(2):22, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v306/n2/full/scientificamerican0212-
22a.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-22a.pdf.

Matson:2012:SQL

[891] John Matson. In sync, on a quantum level. Scientific American, 306(2):22, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
22b.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-22b.pdf.

Anonymous:2012:Sb

[892] Anonymous. Stat. Scientific American, 306(2):22, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n2/full/scientificamerican0212-22c.html;   http://www.nature.
com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
22c.pdf.

Mellow:2012:SWM

[893] Glendon Mellow. Storybook wishes for Martian Rovers. Scientific American, 306(2):24, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
24a.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-24a.pdf.

Wogan:2012:IC

[894] David Wogan. The impracticality of a cheeseburger. Scientific American, 306(2):24, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
24b.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-24b.pdf.

Harmon:2012:OE

[895] Katherine Harmon. Oral exam. Scientific American, 306(2):24, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n2/full/scientificamerican0212-24c.html;   http:
//www.nature.com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
24c.pdf.

McKenna:2012:SHD

[896] Maryn McKenna. The science of health: A diabetes cliffhanger. Scientific American, 306(2):26–28, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
26.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-26.pdf.

Pogue:2012:TFF

[897] David Pogue. TechnoFiles: The future is for fools. Scientific American, 306(2):29, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
29.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-29.pdf.

Anonymous:2012:RRRa

[898] Anonymous. Reviews and recommendations: Recommended. Scientific American, 306(2):82–83, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
82.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-82.pdf.

Shermer:2012:SLW

[899] Michael Shermer. Skeptic: Lies we tell ourselves. Scientific American, 306(2):84, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
84.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-84.pdf.

Mirsky:2012:AGW

[900] Steve Mirsky. Anti gravity: When animals accidentally attack. Scientific American, 306(2):86, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
86.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-86.pdf.

Fischetti:2012:GSH

[901] Mark Fischetti. Graphic science: How science degrees stack up. Scientific American, 306(2):88, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
88.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-88.pdf.

Moyer:2012:PSD

[902] Michael Moyer. Physics: Is space digital? Scientific American, 306(2):30–37, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-

30.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-30.pdf.

Garnick:2012:MGP

[903] Marc B. Garnick. Medicine: The great prostate cancer debate. Scientific American, 306(2):38–43, February 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n2/full/scientificamerican0212-38.html;   http://www.nature.
com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
38.pdf.

Bartholet:2012:ESA

[904] Jeffrey Bartholet. Environment: Swept from Africa to the Amazon. Scientific American, 306(2):44–49, February 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n2/full/scientificamerican0212-44.html;   http://www.nature.
com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
44.pdf.

Haynes:2012:PHS

[905] Kenneth F. Haynes. Public health: Sleeping with the enemy. Scientific American, 306(2):50–55, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
50.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-50.pdf.

Thompson:2012:CSA

[906] Kalee Thompson. Citizen science: All hands on deck. Scientific American, 306(2):56–59, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
56.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-56.pdf.

Schmitz:2012:SAF

[907] Harold Schmitz and Howard-Yana Shapiro. Sustainable agriculture: The future of chocolate. Scientific American, 306(2):60–65, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n2/full/scientificamerican0212-60.html;   http://www.nature.

com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
60.pdf.

Bartholet:2012:BSC

[908] Jeffrey Bartholet. Brain science: The collision syndrome. Scientific American, 306(2):66–71, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
66.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-66.pdf.

Kalb:2012:AFA

[909] Claudia Kalb. Anatomy: Fetal armor. Scientific American, 306(2):72–73, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n2/full/scientificamerican0212-72.html;   http:
//www.nature.com/scientificamerican/journal/v306/n2/pdf/scientificamerican0212-
72.pdf.

Cook:2012:BBS

[910] Gareth Cook. Bioengineering: The brittle star’s apprentice. Scientific American, 306(2):74–77, February 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n2/full/scientificamerican0212-
74.html;   http://www.nature.com/scientificamerican/journal/
v306/n2/pdf/scientificamerican0212-74.pdf.

DiChristina:2012:EMM

[911] Mariette DiChristina. From the Editor: The moving mind. Scientific American, 306(3):4, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
4.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-4.pdf.

Anonymous:2012:Ld

[912] Anonymous. Letters. Scientific American, 306(3):6–7, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n3/full/scientificamerican0312-6.html;   http://www.nature.
com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
6.pdf.

Schlenoff:2012:YAIc

[913] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 306(3):82–83, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-82.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
82.pdf.

Anonymous:2012:SAN

[914] Anonymous. Science agenda: A neglect of mental illness. Scientific American, 306(3):8, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
8.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-8.pdf.

Lamb:2012:FBC

[915] Donald Q. Lamb. Forum: Big computers for little engineers. Scientific American, 306(3):10, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
10.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-10.pdf.

Grossman:2012:ARD

[916] Elizabeth Grossman. Advances: Remains of the day. Scientific American, 306(3):11, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
11.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-11.pdf.

Stix:2012:FC

[917] Gary Stix. Forget cramming. Scientific American, 306(3):12, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-12a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
12a.pdf.

Anonymous:2012:Sc

[918] Anonymous. Stat. Scientific American, 306(3):12, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n3/full/scientificamerican0312-12b.html;   http://www.nature.
com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
12b.pdf.

Choi:2012:IWY

[919] Charles Q. Choi. It watches your wallet. Scientific American, 306(3):13, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-13.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
13.pdf.

Harmon:2012:BSF

[920] Katherine Harmon. Body-snatching flies. Scientific American, 306(3): 14, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-14a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
14a.pdf.

Matson:2012:KYS

[921] John Matson. Know your space tycoons. Scientific American, 306(3):14, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-14b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
14b.pdf.

Arnold:2012:BD

[922] Carrie Arnold. Backseat drivers. Scientific American, 306(3):15, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-15.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
15.pdf.

Matson:2012:NWT

[923] John Matson. A new wrinkle in time. Scientific American, 306(3):16, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-16.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
16.pdf.

Chin:2012:WIc

[924] Ann Chin. What is it? Scientific American, 306(3):17, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n3/full/scientificamerican0312-17.html;   http://www.nature.
com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
17.pdf.

Eveleth:2012:PD

[925] Rose Eveleth. Probing the depths. Scientific American, 306(3):18, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-18.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
18.pdf.

Benderly:2012:ULC

[926] Beryl Lieff Benderly. Are university labs criminally dangerous? Scientific American, 306(3):19, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
19.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-19.pdf.

Greenmeier:2012:AT

[927] Larry Greenmeier. It’s all in the tail. Scientific American, 306(3):20, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-20.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
20.pdf.

Biello:2012:SS

[928] David Biello. Soot soldiers. Scientific American, 306(3):21, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n3/full/scientificamerican0312-21.html;   http://www.nature.

com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
21.pdf.

Oakes:2012:SGB

[929] Kelly Oakes. Stars that go out with a bang. Scientific American, 306(3): 22, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n3/full/scientificamerican0312-22a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
22a.pdf.

Dunn:2012:YAC

[930] Rob Dunn. Your appendix could save your life. Scientific American, 306(3):22, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
22b.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-22b.pdf.

Pogue:2012:TTK

[931] David Pogue. TechnoFiles: Time to kill off captchas. Scientific American, 306(3):23, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
23.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-23.pdf.

Franklin:2012:SHN

[932] Deborah Franklin. The science of health: Nature that nurtures. Scientific American, 306(3):24–25, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
24.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-24.pdf.

Anonymous:2012:RRRb

[933] Anonymous. Reviews and recommendations: Recommended. Scientific American, 306(3):76–77, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
76.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-76.pdf.

Shermer:2012:SOO

[934] Michael Shermer. Skeptic: Opting out of overoptimism. Scientific American, 306(3):78, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
78.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-78.pdf.

Mirsky:2012:AGU

[935] Steve Mirsky. Anti gravity: The unkindest cut. Scientific American, 306(3):80, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
80.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-80.pdf.

Matson:2012:GSD

[936] John Matson. Graphic science: The dwindling web. Scientific American, 306(3):88, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
88.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-88.pdf.

Gage:2012:NWM

[937] Fred H. Gage and Alysson R. Muotri. Neuroscience: What makes each brain unique. Scientific American, 306(3):26–31, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n3/full/scientificamerican0312-26.html;   http://www.nature.
com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
26.pdf.

Goldsmith:2012:AFF

[938] Donald Goldsmith. Astrophysics: The far, far future of stars. Scientific American, 306(3):32–39, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
32.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-32.pdf.

Sampson:2012:PDL

[939] Scott D. Sampson. Paleontology: Dinosaurs of the lost continent. Scientific American, 306(3):40–47, March 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n3/full/scientificamerican0312-40.html;   http://www.nature.
com/scientificamerican/journal/v306/n3/pdf/scientificamerican0312-
40.pdf.

Castelvecchi:2012:EGW

[940] Davide Castelvecchi. Energy: Gather the wind. Scientific American, 306(3):48–53, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
48.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-48.pdf.

June:2012:MBH

[941] Carl June and Bruce Levine. Medicine: Blocking HIV’s attack. Scientific American, 306(3):54–59, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
54.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-54.pdf.

Dibbell:2012:ITS

[942] Julian Dibbell. Information technology: The shadow web. Scientific American, 306(3):60–65, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
60.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-60.pdf.

Henig:2012:BLB

[943] Robin Marantz Henig. Biotechnology: Lifting the black cloud. Scientific American, 306(3):66–71, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
66.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-66.pdf.

Biello:2012:CHT

[944] David Biello. Climate: Hit them with the hockey stick. Scientific American, 306(3):72–75, March 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n3/full/scientificamerican0312-
72.html;   http://www.nature.com/scientificamerican/journal/
v306/n3/pdf/scientificamerican0312-72.pdf.

DiChristina:2012:ESB

[945] Mariette DiChristina. From the Editor: The story begins. Scientific American, 306(4):4, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
4.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-4.pdf.

Anonymous:2012:Lf

[946] Anonymous. Letters. Scientific American, 306(4):6–8, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-6.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
6.pdf.

Schlenoff:2012:YAId

[947] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 306(4):86–87, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n4/full/scientificamerican0412-86.html;   http:
//www.nature.com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
86.pdf.

Anonymous:2012:SAW

[948] Anonymous. Science agenda: Who owns the past? Scientific American, 306(4):9, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
9.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-9.pdf.

Firestein:2012:FWS

[949] Stuart Firestein. Forum: What science wants to know. Scientific American, 306(4):10, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
10.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-10.pdf.

Stix:2012:AMR

[950] Gary Stix. Advances: The mind recovery act. Scientific American, 306(4):12, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
12.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-12.pdf.

Choi:2012:SC

[951] Charles Q. Choi. Scanning for E.T.’s calls. Scientific American, 306(4): 14, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n4/full/scientificamerican0412-14a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
14a.pdf.

Anonymous:2012:N

[952] Anonymous. By the numbers. Scientific American, 306(4):14, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-14b.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
14b.pdf.

Jabr:2012:HV

[953] Ferris Jabr. Hopeful vision. Scientific American, 306(4):15, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-15.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
15.pdf.

Madren:2012:PEC

[954] Carrie Madren. Picky eaters club. Scientific American, 306(4):16, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n4/full/scientificamerican0412-16a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
16a.pdf.

Eveleth:2012:PWa

[955] Rose Eveleth. Patent watch. Scientific American, 306(4):16, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-16b.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
16b.pdf.

Pycha:2012:FT

[956] Anne Pycha. Fast talkers. Scientific American, 306(4):18, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-18.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
18.pdf.

Laursen:2012:BBB

[957] Lucas Laursen. Blue bacteria in bloom. Scientific American, 306(4):19, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n4/full/scientificamerican0412-19.html;   http:
//www.nature.com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
19.pdf.

Eveleth:2012:CM

[958] Rose Eveleth. Coffee mystery. Scientific American, 306(4):20, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-20a.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
20a.pdf.

Choi:2012:TCA

[959] Charles Q. Choi. Textbooks come alive. Scientific American, 306(4):20, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n4/full/scientificamerican0412-20b.html;   http:

//www.nature.com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
20b.pdf.

Matson:2012:SM

[960] John Matson. Swimming on mars. Scientific American, 306(4):22, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n4/full/scientificamerican0412-22.html;   http:
//www.nature.com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
22.pdf.

Harmon:2012:WI

[961] Katherine Harmon. What is it? Scientific American, 306(4):23, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n4/full/scientificamerican0412-23.html;   http:
//www.nature.com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
23.pdf.

Scharf:2012:CC

[962] Caleb A. Scharf. Cold call. Scientific American, 306(4):24, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-24a.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
24a.pdf.

Wilcox:2012:WSL

[963] Christie Wilcox. This is what a scientist looks like. Scientific American, 306(4):24, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
24b.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-24b.pdf.

McKenna:2012:SHF

[964] Maryn McKenna. The science of health: Food poisoning’s hidden legacy. Scientific American, 306(4):26–27, April 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-26.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
26.pdf.

Pogue:2012:TTF

[965] David Pogue. TechnoFiles: Technology’s friction problem. Scientific American, 306(4):28, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
28.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-28.pdf.

Anonymous:2012:RRRc

[966] Anonymous. Reviews and recommendations: Recommended. Scientific American, 306(4):80, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
80.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-80.pdf.

Shermer:2012:SCM

[967] Michael Shermer. Skeptic: Climbing mount immortality. Scientific American, 306(4):82, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
82.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-82.pdf.

Mirsky:2012:AGD

[968] Steve Mirsky. Anti gravity: The doctor is way out. Scientific American, 306(4):84, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
84.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-84.pdf.

Matson:2012:GSS

[969] John Matson. Graphic science: Space age wasteland. Scientific American, 306(4):88, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
88.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-88.pdf.

Wong:2012:HEF

[970] Kate Wong. Human evolution: First of our kind. Scientific American, 306(4):30–39, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
30.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-30.pdf.

Carlip:2012:PQG

[971] Steven Carlip. Physics: Quantum gravity in Flatland. Scientific American, 306(4):40–47, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
40.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-40.pdf.

Arnsten:2012:NYB

[972] Amy Arnsten, Carolyn M. Mazure, and Rajita Sinha. Neuroscience: This is your brain in meltdown. Scientific American, 306(4):48–53, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n4/full/scientificamerican0412-48.html;   http://www.nature.
com/scientificamerican/journal/v306/n4/pdf/scientificamerican0412-
48.pdf.

Belfiore:2012:SSB

[973] Michael Belfiore. Space science: Bound for the Moon. Scientific American, 306(4):54–59, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
54.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-54.pdf.

Branswell:2012:PHP

[974] Helen Branswell. Public health: Polio’s last act. Scientific American, 306(4):60–65, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
60.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-60.pdf.

Swanson:2012:HSB

[975] William Swanson. History of science: Birth of a cold war vaccine. Scientific American, 306(4):66–69, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-

66.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-66.pdf.

Milner:2012:PTT

[976] Richard Milner. Paleontology: Time traveler. Scientific American, 306(4):70–73, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
70.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-70.pdf.

Parkes:2012:PLB

[977] Michael J. Parkes. Physiology: The limits of breath holding. Scientific American, 306(4):74–79, April 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n4/full/scientificamerican0412-
74.html;   http://www.nature.com/scientificamerican/journal/
v306/n4/pdf/scientificamerican0412-74.pdf.

DiChristina:2012:ENP

[978] Mariette DiChristina. From the Editor: New physics and future medicine. Scientific American, 306(5):6, May 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n5/full/scientificamerican0512-6.html;   http://www.nature.
com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
6.pdf.

Anonymous:2012:Lg

[979] Anonymous. Letters. Scientific American, 306(5):8–10, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n5/full/scientificamerican0512-8.html;   http://www.nature.
com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
8.pdf.

Schlenoff:2012:YAIe

[980] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 306(5): 94, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-94.html;   http:

//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
94.pdf.

Anonymous:2012:SAFa

[981] Anonymous. Science agenda: Fresh fruit, hold the insulin. Scientific American, 306(5):12, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
12.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-12.pdf.

Gast:2012:FBC

[982] Alice P. Gast. Forum: Boundary conditions. Scientific American, 306(5): 14, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-14.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
14.pdf.

Castelvecchi:2012:ASD

[983] Davide Castelvecchi. Advances: Is supersymmetry dead? Scientific American, 306(5):16–18, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
16.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-16.pdf.

Dybas:2012:BM

[984] Cheryl Lyn Dybas. Bright microbes. Scientific American, 306(5):19, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-19.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
19.pdf.

Katz:2012:BIS

[985] Cheryl Katz. Bugs in the ice sheet. Scientific American, 306(5):20, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-20a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
20a.pdf.

Anonymous:2012:SP

[986] Anonymous. Science prizes. Scientific American, 306(5):20, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n5/full/scientificamerican0512-20b.html;   http://www.nature.
com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
20b.pdf.

Moyer:2012:WR

[987] Melinda Wenner Moyer. Weighing the risks. Scientific American, 306(5): 22, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-22.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
22.pdf.

Castelvecchi:2012:GPN

[988] Davide Castelvecchi. Goldbach’s prime numbers. Scientific American, 306(5):23, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
23a.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-23a.pdf.

Eveleth:2012:PWb

[989] Rose Eveleth. Patent watch. Scientific American, 306(5):23, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n5/full/scientificamerican0512-23b.html;   http://www.nature.
com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
23b.pdf.

Biello:2012:ASF

[990] David Biello. “it’s almost science fiction”. Scientific American, 306(5): 24, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-24a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
24a.pdf.

Anonymous:2012:Sd

[991] Anonymous. Stat. Scientific American, 306(5):24, May 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n5/full/scientificamerican0512-24b.html;   http://www.nature.
com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
24b.pdf.

Moyer:2012:DDP

[992] Michael Moyer. Does digital piracy really hurt movies? Scientific American, 306(5):25, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
25a.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-25a.pdf.

Eveleth:2012:WI

[993] Rose Eveleth. What is it? Scientific American, 306(5):25, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n5/full/scientificamerican0512-25b.html;   http://www.nature.
com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
25b.pdf.

Oberg:2012:RT

[994] James E. Oberg. Not ready for takeoff. Scientific American, 306(5): 26, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-26a.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
26a.pdf.

Dolgin:2012:RME

[995] Elie Dolgin. Real males eat yogurt. Scientific American, 306(5):26, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-26b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
26b.pdf.

Kuchner:2012:SR

[996] Marc Kuchner. Stars to the rescue. Scientific American, 306(5):28, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-28a.html;   http:

//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
28a.pdf.

Moyer:2012:HTW

[997] Michael Moyer. How to tell who’s tracking you online. Scientific American, 306(5):28, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
28b.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-28b.pdf.

McKenna:2012:SHR

[998] Maryn McKenna. The science of health: Return of the clap. Scientific American, 306(5):30–31, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
30.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-30.pdf.

Pogue:2012:TTW

[999] David Pogue. TechnoFiles: The trouble with wi-fi. Scientific American, 306(5):32, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
32.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-32.pdf.

Anonymous:2012:RRRd

[1000] Anonymous. Reviews and recommendations: Recommended. Scientific American, 306(5):84–85, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
84.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-84.pdf.

Shermer:2012:SMA

[1001] Michael Shermer. Skeptic: Much ado about nothing. Scientific American, 306(5):86, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
86.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-86.pdf.

Mirsky:2012:AGM

[1002] Steve Mirsky. Anti gravity: Math rules. Scientific American, 306(5): 92, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-92.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
92.pdf.

Matson:2012:GSH

[1003] John Matson. Graphic science: High and dry in the food desert. Scientific American, 306(5):96, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
96.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-96.pdf.

Bern:2012:QPL

[1004] Zvi Bern, Lance J. Dixon, and David A. Kosower. Quantum physics: Loops, trees and the search for new physics. Scientific American, 306(5): 34–41, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-34.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
34.pdf.

Shute:2012:FHS

[1005] Nancy Shute, Ferris Jabr, and Katherine Harmon. Future health special report: Tomorrow’s medicine. Scientific American, 306(5):42–47, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-42.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
42.pdf.

Rogers:2012:ETT

[1006] Kristina A. Curry Rogers and Michael D. D’Emic. Evolution: Triumph of the titans. Scientific American, 306(5):48–55, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n5/full/scientificamerican0512-48.html;   http://www.nature.
com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
48.pdf.

Adler:2012:NEP

[1007] Jerry Adler. Neuroscience: Erasing painful memories. Scientific American, 306(5):56–61, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
56.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-56.pdf.

Chamovitz:2012:BWP

[1008] Daniel Chamovitz. Botany: What a plant smells. Scientific American, 306(5):62–65, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
62.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-62.pdf.

Chin:2012:FMT

[1009] Ann Chin. Forensic medicine: Telltale hearts. Scientific American, 306(5):66–67, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
66.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-66.pdf.

Lubchenco:2012:MBE

[1010] Jane Lubchenco and Jack Hayes. Meteorology: A better eye on the storm. Scientific American, 306(5):68–73, May 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n5/full/scientificamerican0512-68.html;   http://www.nature.
com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
68.pdf.

Ball:2012:PNC

[1011] Philip Ball. Photonics: Nature’s color tricks. Scientific American, 306(5): 74–79, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n5/full/scientificamerican0512-74.html;   http:
//www.nature.com/scientificamerican/journal/v306/n5/pdf/scientificamerican0512-
74.pdf.

Greenemeier:2012:IPS

[1012] Larry Greenemeier. Innovation: Professional seer. Scientific American, 306(5):80–83, May 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n5/full/scientificamerican0512-
80.html;   http://www.nature.com/scientificamerican/journal/
v306/n5/pdf/scientificamerican0512-80.pdf.

DiChristina:2012:EWP

[1013] Mariette DiChristina. From the Editor: We the people. Scientific American, 306(6):6, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
6.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-6.pdf.

Anonymous:2012:Lh

[1014] Anonymous. Letters. Scientific American, 306(6):8–10, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n6/full/scientificamerican0612-8.html;   http://www.nature.
com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
8.pdf.

Schlenoff:2012:YAIf

[1015] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 306(6):94–95, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-94.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
94.pdf.

Anonymous:2012:SAP

[1016] Anonymous. Science agenda: Protect women’s health. Scientific American, 306(6):12, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
12.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-12.pdf.

Allison:2012:FST

[1017] Graham T. Allison. Forum: Slinking toward the bomb. Scientific American, 306(6):14, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
14.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-14.pdf.

Moyer:2012:AMG

[1018] Melinda Wenner Moyer. Advances: The motherhood gap. Scientific American, 306(6):16, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
16.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-16.pdf.

Matson:2012:FW

[1019] John Matson. Fire and water. Scientific American, 306(6):18, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n6/full/scientificamerican0612-18a.html;   http://www.nature.
com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
18a.pdf.

Castelvecchi:2012:WI

[1020] Davide Castelvecchi. What is it? Scientific American, 306(6):18, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-18b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
18b.pdf.

Jabr:2012:MM

[1021] Ferris Jabr. Microbial mules. Scientific American, 306(6):20, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n6/full/scientificamerican0612-20.html;   http://www.nature.
com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
20.pdf.

Choi:2012:TSB

[1022] Charles Q. Choi. Thar she blows! Scientific American, 306(6):21, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-21.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
21.pdf.

Crist:2012:ONN

[1023] Meehan Crist. Old neurons, new tricks. Scientific American, 306(6):22, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-22.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
22.pdf.

Eveleth:2012:WCL

[1024] Rose Eveleth. When cockroach legs dance. Scientific American, 306(6): 23, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-23.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
23.pdf.

Wilczek:2012:HBE

[1025] Frank Wilczek. Happy birthday, electron. Scientific American, 306(6): 24, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-24.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
24.pdf.

Cook:2012:CTS

[1026] William J. Cook. The case of the traveling salesman. Scientific American, 306(6):26, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
26a.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-26a.pdf.

Biello:2012:PP

[1027] David Biello. Primeval precipitation. Scientific American, 306(6):26, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-26b.html;   http:

//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
26b.pdf.

Perkins:2012:GF

[1028] Sid Perkins. Go with the flow. Scientific American, 306(6):27, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n6/full/scientificamerican0612-27a.html;   http://www.nature.
com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
27a.pdf.

Eveleth:2012:LGG

[1029] Rose Eveleth. It’s not “like growing grass”. Scientific American, 306(6): 27, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-27b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
27b.pdf.

Wilcox:2012:LDL

[1030] Christie Wilcox. Lice don’t lie. Scientific American, 306(6):28, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n6/full/scientificamerican0612-28a.html;   http://www.nature.
com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
28a.pdf.

Jabr:2012:BM

[1031] Ferris Jabr. Beautiful mutants. Scientific American, 306(6):28, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-28b.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
28b.pdf.

Starks:2012:SHS

[1032] Philip T. B. Starks and Brittany L. Slabach. The science of health: The scoop on eating dirt. Scientific American, 306(6):30–32, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n6/full/scientificamerican0612-30.html;   http://www.nature.
com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
30.pdf.

Pogue:2012:TDD

[1033] David Pogue. TechnoFiles: Down with double data fees! Scientific American, 306(6):34, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
34.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-34.pdf.

Anonymous:2012:RRRe

[1034] Anonymous. Reviews and recommendations: Recommended. Scientific American, 306(6):84–85, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
84.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-84.pdf.

Shermer:2012:SSR

[1035] Michael Shermer. Skeptic: The science of righteousness. Scientific American, 306(6):86, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
86.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-86.pdf.

Mirsky:2012:AGFa

[1036] Steve Mirsky. Anti gravity: Freaks and tweaks. Scientific American, 306(6):88, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
88.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-88.pdf.

Fischetti:2012:GSW

[1037] Mark Fischetti. Graphic science: Water in, water out. Scientific American, 306(6):96, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
96.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-96.pdf.

Ackerman:2012:MUS

[1038] Jennifer Ackerman. Medicine: The ultimate social network. Scientific American, 306(6):36–43, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
36.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-36.pdf.

Gal-Yam:2012:ASS

[1039] Avishay Gal-Yam. Astrophysics: Super supernovae. Scientific American, 306(6):44–49, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
44.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-44.pdf.

Markram:2012:NHB

[1040] Henry Markram. Neuroscience: The human brain project. Scientific American, 306(6):50–55, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
50.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-50.pdf.

Brumfiel:2012:TFM

[1041] Geoff Brumfiel. Technology: Fusion’s missing pieces. Scientific American, 306(6):56–61, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
56.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-56.pdf.

Eveleth:2012:EBB

[1042] Rose Eveleth. Ecology: Busy bee. Scientific American, 306(6):62–63, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v306/n6/full/scientificamerican0612-62.html;   http:
//www.nature.com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
62.pdf.

Guterl:2012:IDW

[1043] Fred Guterl. Infectious disease: Waiting to explode. Scientific American, 306(6):64–69, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-

64.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-64.pdf.

Kaiser:2012:HSR

[1044] David Kaiser and Angela N. H. Creager. History of science: The right way to get it wrong. Scientific American, 306(6):70–75, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v306/
n6/full/scientificamerican0612-70.html;   http://www.nature.
com/scientificamerican/journal/v306/n6/pdf/scientificamerican0612-
70.pdf.

Chase:2012:SLS

[1045] Ivan Chase. Sociobiology: Life is a shell game. Scientific American, 306(6):76–79, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
76.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-76.pdf.

Borrell:2012:VRF

[1046] Brendan Borrell. Virology: Resistance fighter. Scientific American, 306(6):80–83, June 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v306/n6/full/scientificamerican0612-
80.html;   http://www.nature.com/scientificamerican/journal/
v306/n6/pdf/scientificamerican0612-80.pdf.

DiChristina:2012:EST

[1047] Mariette DiChristina. From the Editor: Succeeding together. Scientific American, 307(1):8, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
8.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-8.pdf.

Anonymous:2012:Li

[1048] Anonymous. Letters. Scientific American, 307(1):10–12, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n1/full/scientificamerican0712-10.html;   http://www.nature.
com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
10.pdf.

Schlenoff:2012:YAIg

[1049] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 307(1):86– 87, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-86.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
86.pdf.

Anonymous:2012:SAM

[1050] Anonymous. Science agenda: To the Moons, NASA. Scientific American, 307(1):14, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
14.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-14.pdf.

Collins:2012:FHF

[1051] Francis Collins. Forum: How to fulfill the true promise of “mHealth”. Scientific American, 307(1):16, July 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n1/full/scientificamerican0712-16.html;   http://www.nature.
com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
16.pdf.

Greenberg:2012:AHR

[1052] Susan H. Greenberg. Advances: A home on the range. Scientific American, 307(1):18–19, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
18.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-18.pdf.

Biello:2012:TEH

[1053] David Biello. Trashing the “element from hell”. Scientific American, 307 (1):19, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-19a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
19a.pdf.

Coffey:2012:FFD

[1054] Rebecca Coffey. Fact finder: Digit divide. Scientific American, 307(1): 19, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-19b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
19b.pdf.

Moyer:2012:WSG

[1055] Melinda Wenner Moyer. Why sneezing is good for you. Scientific American, 307(1):20, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
20.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-20.pdf.

Langville:2012:W

[1056] Amy N. Langville and Carl D. Meyer. Who’s #1? Scientific American, 307(1):21, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
21.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-21.pdf.

Biello:2012:BBB

[1057] David Biello. Bad for bugs and brains? Scientific American, 307(1): 22, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-22a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
22a.pdf.

Eveleth:2012:E

[1058] Rose Eveleth. It’s electric. Scientific American, 307(1):22, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n1/full/scientificamerican0712-22b.html;   http://www.nature.
com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
22b.pdf.

Coffey:2012:SLI

[1059] Rebecca Coffey. Some like it too hot. Scientific American, 307(1): 23, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-23.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
23.pdf.

Nelson:2012:LW

[1060] Roxanne Nelson. The last worm. Scientific American, 307(1):24, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-24a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
24a.pdf.

Chin:2012:WId

[1061] Ann Chin. What is it? Scientific American, 307(1):24, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n1/full/scientificamerican0712-24b.html;   http://www.nature.
com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
24b.pdf.

Grewal:2012:HCT

[1062] Daisy Grewal. How critical thinkers lose their faith in god. Scientific American, 307(1):26, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
26.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-26.pdf.

Madren:2012:TTS

[1063] Carrie Madren. Tracking turtles from space. Scientific American, 307(1): 27, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-27.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
27.pdf.

Musser:2012:PP

[1064] George Musser. Particles for peace. Scientific American, 307(1):28, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-28a.html;   http:

//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
28a.pdf.

Kuchment:2012:HSB

[1065] Anna Kuchment. How spiders “balloon”. Scientific American, 307(1): 28, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-28b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
28b.pdf.

Bleicher:2012:SHP

[1066] Ariel Bleicher. The science of health: Perils of newborn screening. Scientific American, 307(1):30–31, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
30.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-30.pdf.

Pogue:2012:TTD

[1067] David Pogue. TechnoFiles: Technology that doesn’t fly. Scientific American, 307(1):32, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
32.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-32.pdf.

Anonymous:2012:RRRf

[1068] Anonymous. Reviews and recommendations: Rabid: A cultural history of the world’s most diabolical virus. Scientific American, 307(1): 82, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-82a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
82a.pdf.

Anonymous:2012:SSR

[1069] Anonymous. Soundings: The story of the remarkable woman who mapped the ocean floor. Scientific American, 307(1):82, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n1/full/scientificamerican0712-82b.html;   http://www.nature.

com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
82b.pdf.

Anonymous:2012:HME

[1070] Anonymous. Homo mysterious: Evolutionary puzzles of human nature. Scientific American, 307(1):82, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n1/full/scientificamerican0712-
82c.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-82c.pdf.

Anonymous:2012:ISN

[1071] Anonymous. The irrationals: A story of the numbers you can’t count on. Scientific American, 307(1):82, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n1/full/scientificamerican0712-
82d.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-82d.pdf.

Kuchment:2012:NFB

[1072] Anna Kuchment. Also notable: Family books. Scientific American, 307 (1):82, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-82e.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
82e.pdf.

Shermer:2012:SAM

[1073] Michael Shermer. Skeptic: Aunt Millie’s mind. Scientific American, 307 (1):84, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-84.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
84.pdf.

Mirsky:2012:AGBa

[1074] Steve Mirsky. Anti gravity: Bred and circuses. Scientific American, 307 (1):85, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-85.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
85.pdf.

Anonymous:2012:GSM

[1075] Anonymous. Graphic science: Medal migrations. Scientific American, 307(1):92, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
92.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-92.pdf.

Nowak:2012:EWW

[1076] Martin A. Nowak. Evolution: Why we help. Scientific American, 307(1): 34–39, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n1/full/scientificamerican0712-34.html;   http:
//www.nature.com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
34.pdf.

Grotzinger:2012:SRR

[1077] John P. Grotzinger and Ashwin Vasavada. Space: Reading the red planet. Scientific American, 307(1):40–43, July 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n1/full/scientificamerican0712-40.html;   http://www.nature.
com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
40.pdf.

Walker:2012:HSH

[1078] Bruce D. Walker. Health: Secrets of the HIV controllers. Scientific American, 307(1):44–51, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
44.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-44.pdf.

Greenemeier:2012:EFF

[1079] Larry Greenemeier. Engineering: Fleet of foot. Scientific American, 307(1):52–53, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
52.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-52.pdf.

Fox:2012:PSW

[1080] Douglas Fox. Polar science: Witness to an Antarctic meltdown. Scientific American, 307(1):54–61, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
54.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-54.pdf.

Matson:2012:PNP

[1081] John Matson and Ferris Jabr. Physics: Nobel pursuits. Scientific American, 307(1):62–73, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
62.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-62.pdf.

Bering:2012:ABR

[1082] Jesse Bering. Animal behavior: The rat that laughed. Scientific American, 307(1):74–77, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n1/full/scientificamerican0712-
74.html;   http://www.nature.com/scientificamerican/journal/
v307/n1/pdf/scientificamerican0712-74.pdf.

Abu-Mostafa:2012:AIM

[1083] Yaser S. Abu-Mostafa. Artificial intelligence: Machines that think for themselves. Scientific American, 307(1):78–81, July 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n1/full/scientificamerican0712-78.html;   http://www.nature.
com/scientificamerican/journal/v307/n1/pdf/scientificamerican0712-
78.pdf.

DiChristina:2012:EDC

[1084] Mariette DiChristina. From the Editor: Destroyers that create. Scientific American, 307(2):6, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
6.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-6.pdf.

Anonymous:2012:Lj

[1085] Anonymous. Letters. Scientific American, 307(2):8–9, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-8.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
8.pdf.

Schlenoff:2012:YAIh

[1086] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 307(2):90, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-90.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
90.pdf.

Anonymous:2012:SAC

[1087] Anonymous. Science agenda: Can the U.S. get an “A” in science? Scientific American, 307(2):12, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
12.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-12.pdf.

Casadevall:2012:FWT

[1088] Arturo Casadevall and Ferric C. Fang. Forum: Winner takes all. Scientific American, 307(2):13, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
13.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-13.pdf.

Moyer:2012:AJS

[1089] Melinda Wenner Moyer. Advances: Not just “small adults”. Scientific American, 307(2):18, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
18.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-18.pdf.

Jabr:2012:WST

[1090] Ferris Jabr. Why some tomatoes taste better. Scientific American, 307(2):20, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
20.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-20.pdf.

Castelvecchi:2012:FPB

[1091] Davide Castelvecchi. Fabrics that push back. Scientific American, 307(2):21, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
21.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-21.pdf.

Arnold:2012:GV

[1092] Carrie Arnold. Going viral. Scientific American, 307(2):22, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-22.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
22.pdf.

Madren:2012:SSP

[1093] Carrie Madren. For some species, plastic is fantastic. Scientific American, 307(2):23, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
23.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-23.pdf.

Anonymous:2012:WC

[1094] Anonymous. Water for crops. Scientific American, 307(2):24, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-24.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
24.pdf.

Chin:2012:WIe

[1095] Ann Chin. What is it? Scientific American, 307(2):26, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-26a.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
26a.pdf.

Eveleth:2012:PWc

[1096] Rose Eveleth. Patent watch. Scientific American, 307(2):26, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-26b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
26b.pdf.

Biello:2012:MCB

[1097] David Biello. More charge for the buck. Scientific American, 307(2):27, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-27.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
27.pdf.

Jabr:2012:MP

[1098] Ferris Jabr. Mind pops. Scientific American, 307(2):28, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-28a.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
28a.pdf.

Eveleth:2012:FFH

[1099] Rose Eveleth. Fact finder: High-jump physics. Scientific American, 307(2):28, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
28b.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-28b.pdf.

Villasenor:2012:HRC

[1100] John Villasenor. A hacker-ready chip. Scientific American, 307(2):29, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-29a.html;   http:

//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
29a.pdf.

Harmon:2012:VPS

[1101] Katherine Harmon. Vanishing in plain sight. Scientific American, 307(2): 29, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-29b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
29b.pdf.

Franklin:2012:SHC

[1102] Deborah Franklin. The science of health: Cracks in the bone test. Scientific American, 307(2):30–31, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n2/full/scientificamerican0812-
30.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-30.pdf.

Pogue:2012:TSW

[1103] David Pogue. TechnoFiles: Siri, why aren’t you smarter? Scientific American, 307(2):33, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
33.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-33.pdf.

Anonymous:2012:RRRg

[1104] Anonymous. Reviews and recommendations: The rocks don’t lie: A geologist investigates Noah’s Flood. Scientific American, 307(2):84, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-84a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
84a.pdf.

Anonymous:2012:PVB

[1105] Anonymous. Phi: A voyage from the brain to the soul. Scientific American, 307(2):84, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
84b.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-84b.pdf.

Anonymous:2012:CBY

[1106] Anonymous. Curious behavior: Yawning, laughing, hiccupping, and beyond. Scientific American, 307(2):84, August 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-84c.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
84c.pdf.

Anonymous:2012:WCY

[1107] Anonymous. When can you trust the experts? how to tell good science from bad in education. Scientific American, 307(2):84, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-84d.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
84d.pdf.

Kuchment:2012:RNa

[1108] Anna Kuchment. Recommended: Also notable. Scientific American, 307(2):84, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
84e.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-84e.pdf.

Shermer:2012:SFW

[1109] Michael Shermer. Skeptic: Free won’t. Scientific American, 307(2):86, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-86.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
86.pdf.

Mirsky:2012:AGG

[1110] Steve Mirsky. Anti gravity: Good chappe. Scientific American, 307(2): 88, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-88.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
88.pdf.

Fischetti:2012:GSO

[1111] Mark Fischetti. Graphic science: Olympic hurtfuls. Scientific American, 307(2):92, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
92.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-92.pdf.

Scharf:2012:ABB

[1112] Caleb Scharf. Astrophysics: The benevolence of black holes. Scientific American, 307(2):34–39, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
34.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-34.pdf.

Kringelbach:2012:NJM

[1113] Morten L. Kringelbach and Kent C. Berridge. Neuroscience: The joyful mind. Scientific American, 307(2):40–45, August 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-40.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
40.pdf.

Campbell:2012:EBN

[1114] Kevin L. Campbell and Michael Hofreiter. Evolutionary biology: New life for ancient DNA. Scientific American, 307(2):46–51, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-46.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
46.pdf.

Perkins:2012:CLI

[1115] Sid Perkins. Climate: Lakes on ice. Scientific American, 307(2):52–53, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n2/full/scientificamerican0812-52.html;   http:
//www.nature.com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
52.pdf.

Dwyer:2012:ASD

[1116] Joseph R. Dwyer and David M. Smith. Atmospheric science: Deadly rays from clouds. Scientific American, 307(2):54–59, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n2/full/scientificamerican0812-54.html;   http://www.nature.
com/scientificamerican/journal/v307/n2/pdf/scientificamerican0812-
54.pdf.

Wingert:2012:EBB

[1117] Pat Wingert. Education: Building a better science teacher. Scientific American, 307(2):60–67, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
60.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-60.pdf.

Stipp:2012:HQL

[1118] David Stipp. Health: Quiet little traitors. Scientific American, 307(2):68–73, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
68.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-68.pdf.

Nijhuis:2012:BWS

[1119] Michelle Nijhuis. Biodiversity: Which species will live? Scientific American, 307(2):74–79, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
74.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-74.pdf.

Borrell:2012:BPF

[1120] Brendan Borrell. Bacteriology: Phage factor. Scientific American, 307(2):80–83, August 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n2/full/scientificamerican0812-
80.html;   http://www.nature.com/scientificamerican/journal/
v307/n2/pdf/scientificamerican0812-80.pdf.

DiChristina:2012:EB

[1121] Mariette DiChristina. From the Editor: Out of bounds. Scientific American, 307(3):6, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
6.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-6.pdf.

Anonymous:2012:Lk

[1122] Anonymous. Letters. Scientific American, 307(3):8–9, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n3/full/scientificamerican0912-8.html;   http://www.nature.
com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
8.pdf.

Schlenoff:2012:YAIi

[1123] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 307(3):94–95, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-94.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
94.pdf.

Anonymous:2012:SASa

[1124] Anonymous. Science agenda: Safer drugs for kids. Scientific American, 307(3):12, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
12.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-12.pdf.

Willingham:2012:FBS

[1125] Daniel T. Willingham. Forum: Brain science in the classroom. Scientific American, 307(3):14, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
14.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-14.pdf.

Halpern:2012:AHH

[1126] Benjamin S. Halpern. Advances: How healthy is your ocean? Scientific American, 307(3):16–17, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n3/full/scientificamerican0912-
16.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-16.pdf.

Berlanga:2012:RSI

[1127] Oscar Berlanga. Ready, set, implant. Scientific American, 307(3):18, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-18a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
18a.pdf.

Chin:2012:WIf

[1128] Ann Chin. What is it? Scientific American, 307(3):18, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n3/full/scientificamerican0912-18b.html;   http://www.nature.
com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
18b.pdf.

Garisto:2012:IHB

[1129] Robert Garisto and Abhishek Agarwal. The import of the Higgs boson. Scientific American, 307(3):20–21, September 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n3/full/scientificamerican0912-20a.html;   http://www.nature.
com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
20a.pdf.

Lamb:2012:CC

[1130] Evelyn Lamb. Cover charge. Scientific American, 307(3):20, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-20b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
20b.pdf.

Mosher:2012:WR

[1131] Dave Mosher. Not worth the risk? Scientific American, 307(3):22, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-22.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
22.pdf.

Fessenden:2012:MH

[1132] Marissa Fessenden. Meteor hunt. Scientific American, 307(3):23, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-23.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
23.pdf.

Harmon:2012:FSD

[1133] Katherine Harmon. Fish swap dinner for sex. Scientific American, 307(3):24, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
24a.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-24a.pdf.

Platt:2012:HN

[1134] John R. Platt. Howdy, neighbor. Scientific American, 307(3):24, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-24b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
24b.pdf.

Nukala:2012:CR

[1135] Aishwarya Nukala. Chickpea revolution. Scientific American, 307(3):26, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-26.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
26.pdf.

Matson:2012:WW

[1136] John Matson. Wobble, wobble. Scientific American, 307(3):27, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-27a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
27a.pdf.

Singer:2012:CC

[1137] Thea Singer. Cholesterol confusion. Scientific American, 307(3):27, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-27b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
27b.pdf.

Wong:2012:TS

[1138] Kate Wong. Tooth sleuths. Scientific American, 307(3):28, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-28a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
28a.pdf.

Jabr:2012:TTS

[1139] Ferris Jabr. I think, therefore I spell. Scientific American, 307(3):28, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-28b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
28b.pdf.

McKenna:2012:SHC

[1140] Maryn McKenna. The science of health: Clean sweep. Scientific American, 307(3):30–31, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
30.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-30.pdf.

Pogue:2012:THK

[1141] David Pogue. TechnoFiles: Hollywood killed the video star. Scientific American, 307(3):36, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-

36.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-36.pdf.

Anonymous:2012:RRS

[1142] Anonymous. Reviews and recommendations: Stardust revolution. Scientific American, 307(3):90, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n3/full/scientificamerican0912-
90a.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-90a.pdf.

Anonymous:2012:M

[1143] Anonymous. Measurement. Scientific American, 307(3):90, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-90b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
90b.pdf.

Anonymous:2012:MB

[1144] Anonymous. The medical book. Scientific American, 307(3):90, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-90c.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
90c.pdf.

Anonymous:2012:LA

[1145] Anonymous. Lost Antarctica. Scientific American, 307(3):90, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-90d.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
90d.pdf.

Kuchment:2012:RNb

[1146] Anna Kuchment. Recommended: Also notable. Scientific American, 307(3):90, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
90e.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-90e.pdf.

Shermer:2012:SCC

[1147] Michael Shermer. Skeptic: Conspiracy contradictions. Scientific American, 307(3):91, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
91.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-91.pdf.

Mirsky:2012:AGBb

[1148] Steve Mirsky. Anti gravity: Bring out your dead. Scientific American, 307(3):92, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
92.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-92.pdf.

Guterl:2012:GSW

[1149] Fred Guterl. Graphic science: Which creatures live the longest? Scientific American, 307(3):96, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n3/full/scientificamerican0912-
96.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-96.pdf.

Anonymous:2012:BL

[1150] Anonymous. Beyond limits. Scientific American, 307(3):38–39, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-38.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
38.pdf.

Sapolsky:2012:WWS

[1151] Robert M. Sapolsky. Who we are: Super humanity. Scientific American, 307(3):40–43, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
40.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-40.pdf.

Folger:2012:CWK

[1152] Tim Folger. Can we keep getting smarter? Scientific American, 307(3):44–47, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
44.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-44.pdf.

Vlahos:2012:CSS

[1153] James Vlahos. The case of the sleeping slayer. Scientific American, 307(3):48–53, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
48.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-48.pdf.

Harmon:2012:HWA

[1154] Katherine Harmon. How we all will live to be 100. Scientific American, 307(3):54–57, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
54.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-54.pdf.

Nicolelis:2012:WWC

[1155] Miguel A. L. Nicolelis. What we can do: Mind in motion. Scientific American, 307(3):58–63, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n3/full/scientificamerican0912-
58.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-58.pdf.

Mosher:2012:EA

[1156] Dave Mosher. The edge of ambition. Scientific American, 307(3):64–65, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-64.html;   http:
//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
64.pdf.

Pavlus:2012:MI

[1157] John Pavlus. Machines of the infinite. Scientific American, 307(3):66–71, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n3/full/scientificamerican0912-66.html;   http:

//www.nature.com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
66.pdf.

Castelvecchi:2012:WWH

[1158] Davide Castelvecchi. Where we’re headed: Questions for the next million years. Scientific American, 307(3):72–77, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n3/full/scientificamerican0912-72.html;   http://www.nature.
com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
72.pdf.

Caldeira:2012:GCE

[1159] Ken Caldeira. The Great Climate Experiment. Scientific American, 307(3):78–83, September 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n3/full/scientificamerican0912-
78.html;   http://www.nature.com/scientificamerican/journal/
v307/n3/pdf/scientificamerican0912-78.pdf.

Deutsch:2012:BQH

[1160] David Deutsch and Artur Ekert. Beyond the quantum horizon. Scientific American, 307(3):84–89, September 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n3/full/scientificamerican0912-84.html;   http://www.nature.
com/scientificamerican/journal/v307/n3/pdf/scientificamerican0912-
84.pdf.

DiChristina:2012:EGA

[1161] Mariette DiChristina. From the Editor: A global affair. Scientific American, 307(4):6, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
6.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-6.pdf.

Anonymous:2012:Ll

[1162] Anonymous. Letters. Scientific American, 307(4):8–10, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n4/full/scientificamerican1012-8.html;   http://www.nature.

com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
8.pdf.

Schlenoff:2012:YAIj

[1163] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 307(4):98–99, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-98.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
98.pdf.

Anonymous:2012:SASb

[1164] Anonymous. Science agenda: Solve the Nobel Prize dilemma. Scientific American, 307(4):12, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
12.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-12.pdf.

Ryan:2012:FTS

[1165] Kathleen A. Ryan. Forum: Target the super-spreaders. Scientific American, 307(4):14, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
14.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-14.pdf.

Fessenden:2012:AEC

[1166] Marissa Fessenden. Advances: The exposure cure. Scientific American, 307(4):16, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
16.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-16.pdf.

Greenemeier:2012:SJ

[1167] Larry Greenemeier. Snap judgment. Scientific American, 307(4):18, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-18a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
18a.pdf.

Yuhas:2012:SB

[1168] Daisy Yuhas. Stealing for biodiversity. Scientific American, 307(4):18, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-18b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
18b.pdf.

Greenemeier:2012:SPK

[1169] Larry Greenemeier. Sergei Petrovich Kapitza. Scientific American, 307(4):19, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
19.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-19.pdf.

Matson:2012:PP

[1170] John Matson. Primordial pinwheel. Scientific American, 307(4):20, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-20a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
20a.pdf.

Bushwick:2012:BB

[1171] Sophie Bushwick. Bird brains. Scientific American, 307(4):20, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-20b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
20b.pdf.

Greenemeier:2012:SR

[1172] Larry Greenemeier. Seeing in the rain. Scientific American, 307(4):21, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-21.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
21.pdf.

Wagner:2012:AF

[1173] Eric Wagner. The art of fishing. Scientific American, 307(4):22, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-22.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
22.pdf.

Crew:2012:WI

[1174] Becky Crew. What is it? Scientific American, 307(4):23, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n4/full/scientificamerican1012-23a.html;   http://www.nature.
com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
23a.pdf.

Greenemeier:2012:S

[1175] Larry Greenemeier. Stat. Scientific American, 307(4):23, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n4/full/scientificamerican1012-23b.html;   http://www.nature.
com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
23b.pdf.

Kuchment:2012:CHW

[1176] Anna Kuchment. Coding her way to the top. Scientific American, 307(4): 24, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-24.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
24.pdf.

Lamb:2012:JRG

[1177] Evelyn Lamb. A “just right” guitar. Scientific American, 307(4):26, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-26a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
26a.pdf.

Fessenden:2012:PWa

[1178] Marissa Fessenden. Patent watch. Scientific American, 307(4):26, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-26b.html;   http:

//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
26b.pdf.

Matson:2012:SDI

[1179] John Matson. Stuff that designs itself. Scientific American, 307(4):27, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-27.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
27.pdf.

Gould:2012:UN

[1180] S. E. Gould. Underground network. Scientific American, 307(4):28, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-28a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
28a.pdf.

Matson:2012:UM

[1181] John Matson. Urge to merge. Scientific American, 307(4):28, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-28b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
28b.pdf.

Franklin:2012:SHD

[1182] Deborah Franklin. The science of health: Drug detectives. Scientific American, 307(4):30–32, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
30.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-30.pdf.

Pogue:2012:TRY

[1183] David Pogue. TechnoFiles: Rent out your hot button. Scientific American, 307(4):34, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
34.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-34.pdf.

Anonymous:2012:RRDa

[1184] Anonymous. Reviews and recommendations: “dangerous work”. Scientific American, 307(4):90–91, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n4/full/scientificamerican1012-
90a.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-90a.pdf.

Anonymous:2012:JX

[1185] Anonymous. The joy of X. Scientific American, 307(4):90–91, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-90b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
90b.pdf.

Anonymous:2012:ME

[1186] Anonymous. Mirror Earth. Scientific American, 307(4):90–91, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-90c.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
90c.pdf.

Anonymous:2012:HLF

[1187] Anonymous. The half-life of facts. Scientific American, 307(4):90–91, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-90d.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
90d.pdf.

Kuchment:2012:RNc

[1188] Anna Kuchment. Recommended: Also notable. Scientific American, 307(4):90–91, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
90e.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-90e.pdf.

Shermer:2012:SPI

[1189] Michael Shermer. Skeptic: Politically irrational. Scientific American, 307(4):92, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
92.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-92.pdf.

Mirsky:2012:AGFb

[1190] Steve Mirsky. Anti gravity: It’s full of cats. Scientific American, 307(4): 96, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-96.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
96.pdf.

Fischetti:2012:GSF

[1191] Mark Fischetti. Graphic science: Fatter, drunker nation. Scientific American, 307(4):100, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
100.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-100.pdf.

Sexton:2012:SWS

[1192] John Sexton. State of the world’s science: State of the world’s science. Scientific American, 307(4):36–40, October 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n4/full/scientificamerican1012-36.html;   http://www.nature.
com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
36.pdf.

Theil:2012:WGS

[1193] Stefan Theil. Why Germany still makes things. Scientific American, 307(4):41–43, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
41.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-41.pdf.

Editors:2012:SWB

[1194] The Editors. Scorecard: The world’s best countries in science. Scientific American, 307(4):44–45, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-

44.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-44.pdf.

Altbach:2012:CCK

[1195] Philip G. Altbach and Qi Wang. Can China keep rising? Scientific American, 307(4):46–47, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
46.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-46.pdf.

Crow:2012:CSU

[1196] Michael M. Crow. Citizen Science U. Scientific American, 307(4):48–49, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-48.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
48.pdf.

Stephan:2012:OP

[1197] Paula Stephan. The other 1 percent. Scientific American, 307(4):50–51, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-50.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
50.pdf.

Guterl:2012:CC

[1198] Fred Guterl. Culture of creativity. Scientific American, 307(4):52–53, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-52.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
52.pdf.

Sejnowski:2012:NLB

[1199] Terry Sejnowski and Tobi Delbruck. Neuroscience: The language of the brain. Scientific American, 307(4):54–59, October 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n4/full/scientificamerican1012-54.html;   http://www.nature.
com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
54.pdf.

Zimmer:2012:EEB

[1200] Carl Zimmer. Environment: Ecosystems on the brink. Scientific American, 307(4):60–65, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
60.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-60.pdf.

Riordan:2012:PHL

[1201] Michael Riordan, Guido Tonelli, and Sau Lan Wu. Physics: The Higgs at last. Scientific American, 307(4):66–73, October 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n4/full/scientificamerican1012-66.html;   http://www.nature.
com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
66.pdf.

Biello:2012:EKK

[1202] David Biello. Energy: Kinetic kite. Scientific American, 307(4):74–75, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n4/full/scientificamerican1012-74.html;   http:
//www.nature.com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
74.pdf.

Dutton:2012:PWP

[1203] Kevin Dutton. Psychology: The wisdom of psychopaths. Scientific American, 307(4):76–79, October 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n4/full/scientificamerican1012-
76.html;   http://www.nature.com/scientificamerican/journal/
v307/n4/pdf/scientificamerican1012-76.pdf.

Hall:2012:BJG

[1204] Stephen S. Hall. Biochemistry: Journey to the genetic interior. Scientific American, 307(4):80–84, October 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n4/full/scientificamerican1012-80.html;   http://www.nature.
com/scientificamerican/journal/v307/n4/pdf/scientificamerican1012-
80.pdf.

DiChristina:2012:EQP

[1205] Mariette DiChristina. From the Editor: Of quarks and (presidential) men. Scientific American, 307(5):6, November 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-6.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
6.pdf.

Anonymous:2012:Lm

[1206] Anonymous. Letters. Scientific American, 307(5):8–10, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-8.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
8.pdf.

Schlenoff:2012:YAIk

[1207] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 307(5):90–91, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-90.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
90.pdf.

Anonymous:2012:SAFb

[1208] Anonymous. Science agenda: Future jobs depend on a sciencebased economy. Scientific American, 307(5):12, November 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-12.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
12.pdf.

Charney:2012:FVG

[1209] Evan Charney and William English. Forum: The voting gene. Scientific American, 307(5):14, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
14.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-14.pdf.

Katsnelson:2012:AAP

[1210] Alla Katsnelson. Advances: The autism pill. Scientific American, 307(5):16, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
16.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-16.pdf.

Matson:2012:YH

[1211] John Matson. You are here. Scientific American, 307(5):18, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-18a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
18a.pdf.

Matson:2012:S

[1212] John Matson. Scaled down. Scientific American, 307(5):18, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-18b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
18b.pdf.

Greenemeier:2012:RR

[1213] Larry Greenemeier. Roaches to the rescue. Scientific American, 307(5):19, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
19.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-19.pdf.

Madren:2012:DW

[1214] Carrie Madren. Dead in the water. Scientific American, 307(5):20, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-20.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
20.pdf.

Chin:2012:WIg

[1215] Ann Chin. What is it? Scientific American, 307(5):22, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-22.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
22.pdf.

Moyer:2012:SA

[1216] Melinda Wenner Moyer. A shot in the arm. Scientific American, 307(5):23, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
23.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-23.pdf.

Yuhas:2012:SDL

[1217] Daisy Yuhas. Some don’t like it hot. Scientific American, 307(5):24, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-24a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
24a.pdf.

Fessenden:2012:PWb

[1218] Marissa Fessenden. Patent watch. Scientific American, 307(5):24, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-24b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
24b.pdf.

Nuwer:2012:TCH

[1219] Rachel Nuwer. From tail chasing to hand washing. Scientific American, 307(5):25, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
25.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-25.pdf.

Fessenden:2012:FF

[1220] Marissa Fessenden. Fickle fairies. Scientific American, 307(5):26, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-26a.html;   http:

//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
26a.pdf.

Zeldovich:2012:RWS

[1221] Lina Zeldovich. The robot will see you now. Scientific American, 307(5):26, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
26b.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-26b.pdf.

Castelvecchi:2012:BMB

[1222] Davide Castelvecchi. Break more to break less. Scientific American, 307(5):27, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
27.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-27.pdf.

Wong:2012:HD

[1223] Kate Wong. Helpless by design? Scientific American, 307(5):28, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-28a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
28a.pdf.

Conant:2012:BI

[1224] Eve Conant. Breaking the ice. Scientific American, 307(5):28, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-28b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
28b.pdf.

Matson:2012:LF

[1225] John Matson. The last frontier. Scientific American, 307(5):29, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-29.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
29.pdf.

McKenna:2012:SHV

[1226] Maryn McKenna. The science of health: Virtues of the virtual autopsy. Scientific American, 307(5):30–32, November 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-30.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
30.pdf.

Pogue:2012:TTM

[1227] David Pogue. TechnoFiles: To track my thief. Scientific American, 307(5):34, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
34.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-34.pdf.

Anonymous:2012:RRP

[1228] Anonymous. Reviews and recommendations: The particle at the end of the universe. Scientific American, 307(5):84, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-84a.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
84a.pdf.

Anonymous:2012:DW

[1229] Anonymous. Drinking water. Scientific American, 307(5):84, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-84b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
84b.pdf.

Anonymous:2012:LE

[1230] Anonymous. The logician and the engineer. Scientific American, 307(5):84, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
84c.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-84c.pdf.

Anonymous:2012:BF

[1231] Anonymous. Brain on fire. Scientific American, 307(5):84, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n5/full/scientificamerican1112-84d.html;   http:
//www.nature.com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
84d.pdf.

Kuchment:2012:RNd

[1232] Anna Kuchment. Recommended: Also notable. Scientific American, 307(5):84, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
84e.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-84e.pdf.

Shermer:2012:SSA

[1233] Michael Shermer. Skeptic: Shock and awe. Scientific American, 307(5):86, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
86.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-86.pdf.

Mirsky:2012:AGTa

[1234] Steve Mirsky. Anti gravity: A trivers runs through it. Scientific American, 307(5):88, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
88.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-88.pdf.

Matson:2012:GST

[1235] John Matson. Graphic science: Tag — you’re sick. Scientific American, 307(5):92, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
92.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-92.pdf.

Lincoln:2012:PPI

[1236] Don Lincoln. Particle physics: The inner life of quarks. Scientific American, 307(5):36–43, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
36.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-36.pdf.

Sasai:2012:NGY

[1237] Yoshiki Sasai. Neuroscience: Grow your own eye. Scientific American, 307(5):44–49, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n5/full/scientificamerican1112-
44.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-44.pdf.

Carey:2012:EGW

[1238] John Carey. Environment: Global warming: Faster than expected? Scientific American, 307(5):50–55, November 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-50.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
50.pdf.

Fordyce:2012:PSB

[1239] R. Ewan Fordyce and Daniel T. Ksepka. Paleontology: The strangest bird. Scientific American, 307(5):56–61, November 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-56.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
56.pdf.

Otto:2012:SSA

[1240] Shawn Lawrence Otto. Science and society: America’s science problem. Scientific American, 307(5):62–71, November 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n5/full/scientificamerican1112-62.html;   http://www.nature.
com/scientificamerican/journal/v307/n5/pdf/scientificamerican1112-
62.pdf.

Baron-Cohen:2012:BAT

[1241] Simon Baron-Cohen. Biology: Autism and the technical mind. Scientific American, 307(5):72–75, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n5/full/scientificamerican1112-
72.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-72.pdf.

Musser:2012:QPN

[1242] George Musser. Quantum physics: A new enlightenment. Scientific American, 307(5):76–81, November 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n5/full/scientificamerican1112-
76.html;   http://www.nature.com/scientificamerican/journal/
v307/n5/pdf/scientificamerican1112-76.pdf.

DiChristina:2012:EHF

[1243] Mariette DiChristina. From the Editor: The human factor. Scientific American, 307(6):4, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
4.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-4.pdf.

Anonymous:2012:Ln

[1244] Anonymous. Letters. Scientific American, 307(6):6–8, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n6/full/scientificamerican1212-6.html;   http://www.nature.
com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
6.pdf.

Schlenoff:2012:YAIl

[1245] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 307(6):88–89, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-88.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
88.pdf.

Anonymous:2012:SAO

[1246] Anonymous. Science agenda: The other oil problem. Scientific American, 307(6):10, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-

10.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-10.pdf.

Wadhams:2012:FAD

[1247] Peter Wadhams. Forum: The Arctic “death spiral”. Scientific American, 307(6):12, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
12.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-12.pdf.

Choi:2012:ADD

[1248] Charles Q. Choi. Advances: Digital danger. Scientific American, 307(6):14, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
14.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-14.pdf.

Matson:2012:NSL

[1249] John Matson. No star left behind. Scientific American, 307(6):16, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-16a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
16a.pdf.

Moyer:2012:CU

[1250] Melinda Wenner Moyer. Chemically unsound. Scientific American, 307(6):16, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
16b.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-16b.pdf.

Yuhas:2012:MM

[1251] Daisy Yuhas. Model mammals. Scientific American, 307(6):17, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-17.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
17.pdf.

Anonymous:2012:AWC

[1252] Anonymous. All I want for Christmas is dinosaur dung. Scientific American, 307(6):18–19, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
18.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-18.pdf.

Fessenden:2012:HMH

[1253] Marissa Fessenden. His milk, her milk. Scientific American, 307(6):20, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-20.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
20.pdf.

Lamb:2012:TCS

[1254] Evelyn Lamb. Too cool for school. Scientific American, 307(6):22, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-22a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
22a.pdf.

Fessenden:2012:PWc

[1255] Marissa Fessenden. Patent watch. Scientific American, 307(6):22, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-22b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
22b.pdf.

Jabr:2012:FF

[1256] Ferris Jabr. Food fight. Scientific American, 307(6):23, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n6/full/scientificamerican1212-23.html;   http://www.nature.
com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
23.pdf.

Wong:2012:CC

[1257] Kate Wong. Caveman couture. Scientific American, 307(6):24, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-24a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
24a.pdf.

Scharf:2012:RRT

[1258] Caleb A. Scharf. A river ran through it. Scientific American, 307(6):24, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-24b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
24b.pdf.

Greenemeier:2012:DJS

[1259] Larry Greenemeier. A deadly jet set. Scientific American, 307(6):26, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-26a.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
26a.pdf.

Hodge:2012:CG

[1260] Anne-Marie Hodge. Cryptic gardeners. Scientific American, 307(6):26, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-26b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
26b.pdf.

Chin:2012:WIh

[1261] Ann Chin. What is it? Scientific American, 307(6):27, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n6/full/scientificamerican1212-27.html;   http://www.nature.
com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
27.pdf.

Jabr:2012:SHM

[1262] Ferris Jabr. The science of health: Meat of the matter. Scientific American, 307(6):28–30, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-

28.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-28.pdf.

Pogue:2012:TRC

[1263] David Pogue. TechnoFiles: The remote control in your mind. Scientific American, 307(6):32, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
32.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-32.pdf.

Anonymous:2012:RRDb

[1264] Anonymous. Reviews and recommendations: Desert air. Scientific American, 307(6):82, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
82a.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-82a.pdf.

Anonymous:2012:WW

[1265] Anonymous. Wind wizard. Scientific American, 307(6):82, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-82b.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
82b.pdf.

Anonymous:2012:CDD

[1266] Anonymous. Chasing Doctor Dolittle. Scientific American, 307(6):82, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v307/n6/full/scientificamerican1212-82c.html;   http:
//www.nature.com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
82c.pdf.

Anonymous:2012:BSW

[1267] Anonymous. The best science writing online 2012. Scientific American, 307(6):82, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
82d.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-82d.pdf.

Shermer:2012:SAE

[1268] Michael Shermer. Skeptic: The alpinists of evil. Scientific American, 307(6):84, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
84.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-84.pdf.

Mirsky:2012:AGTb

[1269] Steve Mirsky. Anti gravity: Tongue twisters. Scientific American, 307(6):86, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
86.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-86.pdf.

Matson:2012:GSP

[1270] John Matson. Graphic science: Planets everywhere. Scientific American, 307(6):92, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
92.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-92.pdf.

Jabr:2012:IWC

[1271] Ferris Jabr, Katherine Harmon, Emily Laber-Warren, David Biello, Adam Piore, Daisy Yuhas, Christopher Mims, Marissa Fessenden, and John Carey. Innovation: World changing ideas. Scientific American, 307(6):34–35, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
34.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-34.pdf.

Tong:2012:PUQ

[1272] David Tong. Physics: The unquantum quantum. Scientific American, 307(6):46–49, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
46.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-46.pdf.

Greene:2012:CCW

[1273] Charles H. Greene. Climate change: The winters of our discontent. Scientific American, 307(6):50–55, December 2012. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n6/full/scientificamerican1212-50.html;   http://www.nature.
com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
50.pdf.

Seife:2012:MED

[1274] Charles Seife. Medical ethics: Is drug research trustworthy? Scientific American, 307(6):56–63, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v307/n6/full/scientificamerican1212-
56.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-56.pdf.

Frebel:2012:SFS

[1275] Anna Frebel. Space: Four starry nights. Scientific American, 307(6):64–69, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
64.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-64.pdf.

Grushkin:2012:ACC

[1276] Daniel Grushkin. Art conservation: The case of the disappearing Daguerreotypes. Scientific American, 307(6):70–73, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v307/
n6/full/scientificamerican1212-70.html;   http://www.nature.
com/scientificamerican/journal/v307/n6/pdf/scientificamerican1212-
70.pdf.

Cook:2012:NMT

[1277] Gareth Cook. Neuroscience: Mind theorist. Scientific American, 307(6):74–77, December 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v307/n6/full/scientificamerican1212-
74.html;   http://www.nature.com/scientificamerican/journal/
v307/n6/pdf/scientificamerican1212-74.pdf.

DiChristina:2013:ELF

[1278] Mariette DiChristina. From the Editor: Living in the future. Scientific American, 308(1):4, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
4.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-4.pdf.

Anonymous:2013:La

[1279] Anonymous. Letters. Scientific American, 308(1):6–8, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n1/full/scientificamerican0113-6.html;   http://www.nature.
com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
6.pdf.

Schlenoff:2013:YAIa

[1280] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 308(1):79, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-79.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
79.pdf.

Anonymous:2013:SADa

[1281] Anonymous. Science agenda: A to-do list for Washington. Scientific American, 308(1):10, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
10.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-10.pdf.

Tanenbaum:2013:FCE

[1282] Jacob Tanenbaum. Forum: Creation, evolution and indisputable facts. Scientific American, 308(1):11, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v308/n1/full/scientificamerican0113-
11.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-11.pdf.

Matson:2013:ADP

[1283] John Matson. Advances: Diamond planets. Scientific American, 308(1): 12, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-12.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
12.pdf.

Vance:2013:SS

[1284] Erik Vance. Safe from scorpions. Scientific American, 308(1):14, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-14a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
14a.pdf.

Cahalan:2013:LIM

[1285] Susannah Cahalan. Linking immunity and mental health. Scientific American, 308(1):14, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
14b.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-14b.pdf.

Wong:2013:PAR

[1286] Kate Wong. A prehistoric arms race. Scientific American, 308(1):16, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-16a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
16a.pdf.

Fischetti:2013:HSN

[1287] Mark Fischetti. How to survive the next big storm. Scientific American, 308(1):16, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
16b.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-16b.pdf.

Castelvecchi:2013:ME

[1288] Davide Castelvecchi. The missing epoch. Scientific American, 308(1):17, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-17.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
17.pdf.

Greenemeier:2013:CT

[1289] Larry Greenemeier. Crunch time. Scientific American, 308(1):18, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-18a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
18a.pdf.

Fessenden:2013:PWa

[1290] Marissa Fessenden. Patent watch. Scientific American, 308(1):18, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-18b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
18b.pdf.

Fessenden:2013:SP

[1291] Marissa Fessenden. Stealth pathogen. Scientific American, 308(1):19, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-19.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
19.pdf.

Chin:2013:WIa

[1292] Ann Chin. What is it? Scientific American, 308(1):20, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n1/full/scientificamerican0113-20a.html;   http://www.nature.
com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
20a.pdf.

Wagner:2013:EH

[1293] Eric Wagner. The escape hatch. Scientific American, 308(1):20, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-20b.html;   http:

//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
20b.pdf.

Hornyak:2013:SLT

[1294] Timothy Hornyak. Super long-term storage. Scientific American, 308(1): 21, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-21a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
21a.pdf.

Vega:2013:HIC

[1295] Cesar Vega. Healthier ice cream? Scientific American, 308(1):21, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-21b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
21b.pdf.

Goldman:2013:FI

[1296] Jason G. Goldman. A feathered innovator. Scientific American, 308(1): 22, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-22a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
22a.pdf.

Wilcox:2013:CC

[1297] Christie Wilcox. Clever coral. Scientific American, 308(1):22, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-22b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
22b.pdf.

Stipp:2013:SHF

[1298] David Stipp. The science of health: Is fasting good for you? Scientific American, 308(1):23–24, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
23.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-23.pdf.

Pogue:2013:TTT

[1299] David Pogue. TechnoFiles: The trouble with touch screens. Scientific American, 308(1):25, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
25.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-25.pdf.

Anonymous:2013:RRU

[1300] Anonymous. Reviews and recommendations: The universe within. Scientific American, 308(1):76, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
76a.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-76a.pdf.

Anonymous:2013:PWS

[1301] Anonymous. The physics of wall street. Scientific American, 308(1):76, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-76b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
76b.pdf.

Anonymous:2013:DP

[1302] Anonymous. Drawn from paradise. Scientific American, 308(1):76, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-76c.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
76c.pdf.

Anonymous:2013:AID

[1303] Anonymous. The annotated and illustrated double helix. Scientific American, 308(1):76, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
76d.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-76d.pdf.

Anonymous:2013:M

[1304] Anonymous. Mastermind. Scientific American, 308(1):76, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n1/full/scientificamerican0113-76e.html;   http://www.nature.
com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
76e.pdf.

Shermer:2013:SLT

[1305] Michael Shermer. Skeptic: Logic-tight compartments. Scientific American, 308(1):77, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
77.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-77.pdf.

Mirsky:2013:AGV

[1306] Steve Mirsky. Anti gravity: Voyage of the bagel. Scientific American, 308(1):78, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
78.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-78.pdf.

Fischetti:2013:GST

[1307] Mark Fischetti. Graphic science: The true cost of risky behavior. Scientific American, 308(1):80, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
80.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-80.pdf.

Anonymous:2013:TH

[1308] Anonymous. Technology: Over the horizon. Scientific American, 308(1):26–28, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
26.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-26.pdf.

Cummings:2013:DED

[1309] Mary Cummings. A drone in every driveway. Scientific American, 308(1):28–29, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-

28.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-28.pdf.

Rosenbaum:2013:NQ

[1310] Ron Rosenbaum. The nuclear question. Scientific American, 308(1):30– 31, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-30.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
30.pdf.

Lewis:2013:CWA

[1311] Ricki Lewis. A cure for what ails you. Scientific American, 308(1):32, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-32.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
32.pdf.

Lovejoy:2013:TE

[1312] Thomas Lovejoy. A tsunami of extinction. Scientific American, 308(1):33–34, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
33.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-33.pdf.

Keith:2013:FEP

[1313] David W. Keith and Andy Parker. The fate of an engineered planet. Scientific American, 308(1):34–36, January 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n1/full/scientificamerican0113-34.html;   http://www.nature.
com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
34.pdf.

Regis:2013:BFE

[1314] Ed Regis. A bold and foolish effort to predict the future of computing. Scientific American, 308(1):36–37, January 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n1/full/scientificamerican0113-36.html;   http://www.nature.

com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
36.pdf.

Smith:2013:SSH

[1315] Cameron M. Smith. Space: Starship humanity. Scientific American, 308(1):38–43, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
38.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-38.pdf.

Sachdev:2013:QPS

[1316] Subir Sachdev. Quantum physics: Strange and stringy. Scientific American, 308(1):44–51, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n1/full/scientificamerican0113-
44.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-44.pdf.

Cullen:2013:MEB

[1317] D. Kacy Cullen and Douglas H. Smith. Medical engineering: Bionic connections. Scientific American, 308(1):52–57, January 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n1/full/scientificamerican0113-52.html;   http://www.nature.
com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
52.pdf.

Wong:2013:LSS

[1318] Kate Wong. Life science: Small wonders. Scientific American, 308(1):58– 63, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n1/full/scientificamerican0113-58.html;   http:
//www.nature.com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
58.pdf.

Dettinger:2013:CCM

[1319] Michael D. Dettinger and B. Lynn Ingram. Climate: The coming megafloods. Scientific American, 308(1):64–71, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n1/full/scientificamerican0113-64.html;   http://www.nature.

com/scientificamerican/journal/v308/n1/pdf/scientificamerican0113-
64.pdf.

Rosenblum:2013:NCS

[1320] Lawrence D. Rosenblum. Neuroscience: A confederacy of senses. Scientific American, 308(1):72–75, January 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v308/n1/full/scientificamerican0113-
72.html;   http://www.nature.com/scientificamerican/journal/
v308/n1/pdf/scientificamerican0113-72.pdf.

DiChristina:2013:EMM

[1321] Mariette DiChristina. From the Editor: Minds in motion. Scientific American, 308(2):6, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
6.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-6.pdf.

Anonymous:2013:Lb

[1322] Anonymous. Letters. Scientific American, 308(2):8–10, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n2/full/scientificamerican0213-8.html;   http://www.nature.
com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
8.pdf.

Schlenoff:2013:YAIb

[1323] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 308(2):80, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-80.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
80.pdf.

Anonymous:2013:SAB

[1324] Anonymous. Science agenda: Beware the destiny test. Scientific American, 308(2):12, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
12.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-12.pdf.

Fertik:2013:FTT

[1325] Michael Fertik. Forum: A tale of two Internets. Scientific American, 308(2):13, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
13.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-13.pdf.

Biello:2013:ADB

[1326] David Biello. Advances: A dirty business. Scientific American, 308(2):14, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-14.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
14.pdf.

Matson:2013:RT

[1327] John Matson. Ring theory. Scientific American, 308(2):15, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-15a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
15a.pdf.

Harmon:2013:WE

[1328] Katherine Harmon. Worm elixir. Scientific American, 308(2):15, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-15b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
15b.pdf.

Madren:2013:WNT

[1329] Carrie Madren. Where no tree has gone before. Scientific American, 308(2):16, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
16.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-16.pdf.

Fessenden:2013:PWb

[1330] Marissa Fessenden. Patent watch. Scientific American, 308(2):17, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-17a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
17a.pdf.

Yuhas:2013:YBS

[1331] Daisy Yuhas. Your brain on sports. Scientific American, 308(2):17, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-17b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
17b.pdf.

Harmon:2013:SS

[1332] Katherine Harmon. Stopping superbugs. Scientific American, 308(2):18, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-18a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
18a.pdf.

Worth:2013:NS

[1333] Katie Worth. Naps in space. Scientific American, 308(2):18, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-18b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
18b.pdf.

Fessenden:2013:BE

[1334] Marissa Fessenden. Bug-eared. Scientific American, 308(2):19, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-19.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
19.pdf.

Fessenden:2013:CC

[1335] Marissa Fessenden. Cloud catcher. Scientific American, 308(2):20, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-20.html;   http:

//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
20.pdf.

Wild:2013:WIa

[1336] Alex Wild. What is it? Scientific American, 308(2):21, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n2/full/scientificamerican0213-21.html;   http://www.nature.
com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
21.pdf.

Harmon:2013:BB

[1337] Katherine Harmon. Boring beetles. Scientific American, 308(2):22, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-22.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
22.pdf.

Stix:2013:RG

[1338] Gary Stix. Roots of genius. Scientific American, 308(2):23, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-23.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
23.pdf.

Matson:2013:QG

[1339] John Matson. Not quite gone. Scientific American, 308(2):24, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-24a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
24a.pdf.

Waters:2013:BB

[1340] Hannah Waters. Bird butts. Scientific American, 308(2):24, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-24b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
24b.pdf.

Franklin:2013:SHD

[1341] Deborah Franklin. The science of health: A dangerous game. Scientific American, 308(2):27–28, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
27.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-27.pdf.

Pogue:2013:TR

[1342] David Pogue. TechnoFiles: Out with the real. Scientific American, 308(2):29, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
29.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-29.pdf.

Anonymous:2013:RRWa

[1343] Anonymous. Reviews and recommendations: Weird life. Scientific American, 308(2):74, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
74a.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-74a.pdf.

Anonymous:2013:AW

[1344] Anonymous. Animal wise. Scientific American, 308(2):74, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-74b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
74b.pdf.

Anonymous:2013:BW

[1345] Anonymous. Blackett’s War. Scientific American, 308(2):74, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-74c.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
74c.pdf.

Anonymous:2013:MM

[1346] Anonymous. The measure of Manhattan. Scientific American, 308(2):74, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-74d.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
74d.pdf.

Shermer:2013:SLW

[1347] Michael Shermer. Skeptic: The Left’s war on science. Scientific American, 308(2):76, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
76.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-76.pdf.

Mirsky:2013:AGSa

[1348] Steve Mirsky. Anti gravity: Still crazy after all these fears. Scientific American, 308(2):78, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
78.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-78.pdf.

Guterl:2013:GSR

[1349] Fred Guterl. Graphic science: Rising risks. Scientific American, 308(2):82, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
82.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-82.pdf.

Quiroga:2013:NBC

[1350] Rodrigo Quian Quiroga, Itzhak Fried, and Christof Koch. Neuroscience: Brain cells for grandmother. Scientific American, 308(2):30–35, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n2/full/scientificamerican0213-30.html;   http:
//www.nature.com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
30.pdf.

Rubin:2013:SSP

[1351] Alan E. Rubin. Space: Secrets of primitive meteorites. Scientific American, 308(2):36–41, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-

36.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-36.pdf.

Harmon:2013:HES

[1352] Katherine Harmon. Human evolution: Shattered ancestry. Scientific American, 308(2):42–49, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
42.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-42.pdf.

Riser:2013:CRG

[1353] Stephen C. Riser and M. Susan Lozier. Climate: Rethinking the Gulf Stream. Scientific American, 308(2):50–55, February 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n2/full/scientificamerican0213-50.html;   http://www.nature.
com/scientificamerican/journal/v308/n2/pdf/scientificamerican0213-
50.pdf.

Mone:2013:EBP

[1354] Gregory Mone. Engineering: The battery-powered bullet. Scientific American, 308(2):56–61, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
56.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-56.pdf.

Moyer:2013:MMA

[1355] Melinda Wenner Moyer. Medicine: The myth of antioxidants. Scientific American, 308(2):62–67, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
62.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-62.pdf.

Rosner:2013:CSD

[1356] Hillary Rosner. Citizen science: Data on wings. Scientific American, 308(2):68–73, February 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n2/full/scientificamerican0213-
68.html;   http://www.nature.com/scientificamerican/journal/
v308/n2/pdf/scientificamerican0213-68.pdf.

DiChristina:2013:ESCa

[1357] Mariette DiChristina. From the Editor: The story of creation. Scientific American, 308(3):4, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
4.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-4.pdf.

Anonymous:2013:Lc

[1358] Anonymous. Letters. Scientific American, 308(3):6–8, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n3/full/scientificamerican0313-6.html;   http://www.nature.
com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
6.pdf.

Schlenoff:2013:YAIc

[1359] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 308(3):86–87, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-86.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
86.pdf.

Anonymous:2013:SARa

[1360] Anonymous. Science agenda: Ready. aim. investigate. Scientific American, 308(3):10, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
10.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-10.pdf.

Bartels:2013:FWY

[1361] Dennis M. Bartels. Forum: What is your question? Scientific American, 308(3):12, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
12.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-12.pdf.

Eveleth:2013:AM

[1362] Rose Eveleth. Advances: Up with microbes. Scientific American, 308(3): 14, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-14.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
14.pdf.

Fraser:2013:DRF

[1363] Barbara Fraser. A drying rain forest. Scientific American, 308(3):16, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-16.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
16.pdf.

Fessenden:2013:PWc

[1364] Marissa Fessenden. Patent watch. Scientific American, 308(3):17, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-17.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
17.pdf.

Underwood:2013:Z

[1365] Anne Underwood. From a to zinc. Scientific American, 308(3):18, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-18.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
18.pdf.

Moyer:2013:NHD

[1366] Melinda Wenner Moyer. No harm done? Scientific American, 308(3):19, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-19.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
19.pdf.

Frenkel:2013:BOS

[1367] Karen A. Frenkel. Beyond the ocean’s surface. Scientific American, 308(3):24, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
24a.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-24a.pdf.

Anonymous:2013:MAB

[1368] Anonymous. Medical advice before taking a spaceflight. Scientific American, 308(3):24, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
24b.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-24b.pdf.

Fischetti:2013:BD

[1369] Mark Fischetti. Before the deluge. Scientific American, 308(3):26, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-26.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
26.pdf.

Chin:2013:WIb

[1370] Ann Chin. What is it? Scientific American, 308(3):27, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n3/full/scientificamerican0313-27.html;   http://www.nature.
com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
27.pdf.

Lamb:2013:HSB

[1371] Evelyn Lamb. How the slinky buckles. Scientific American, 308(3):28, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-28.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
28.pdf.

Matson:2013:TCC

[1372] John Matson. Not too close for comfort. Scientific American, 308(3):30, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-30a.html;   http:

//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
30a.pdf.

Fessenden:2013:CBa

[1373] Marissa Fessenden. Crime boss. Scientific American, 308(3):30, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-30b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
30b.pdf.

Wilcox:2013:EO

[1374] Christie Wilcox. Expensive organs. Scientific American, 308(3):32, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-32a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
32a.pdf.

Platt:2013:ES

[1375] John R. Platt. Eyes in the sky. Scientific American, 308(3):32, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-32b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
32b.pdf.

McKenna:2013:SHN

[1376] Maryn McKenna. The science of health: The new age of medical monitoring. Scientific American, 308(3):33–34, March 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n3/full/scientificamerican0313-33.html;   http://www.nature.
com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
33.pdf.

Pogue:2013:TTC

[1377] David Pogue. TechnoFiles: Term of confusion. Scientific American, 308(3):35, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
35.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-35.pdf.

Anonymous:2013:RRR

[1378] Anonymous. Reviews and recommendations: Red rover. Scientific American, 308(3):82, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
82a.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-82a.pdf.

Anonymous:2013:BA

[1379] Anonymous. The bonobo and the atheist. Scientific American, 308(3): 82, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-82b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
82b.pdf.

Anonymous:2013:SK

[1380] Anonymous. Storm kings. Scientific American, 308(3):82, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n3/full/scientificamerican0313-82c.html;   http://www.nature.
com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
82c.pdf.

Anonymous:2013:SEC

[1381] Anonymous. To save everything click here. Scientific American, 308(3): 82, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n3/full/scientificamerican0313-82d.html;   http:
//www.nature.com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
82d.pdf.

Shermer:2013:SDD

[1382] Michael Shermer. Skeptic: Dictators and diehards. Scientific American, 308(3):84, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
84.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-84.pdf.

Mirsky:2013:AGG

[1383] Steve Mirsky. Anti gravity: Getting to the bottom. Scientific American, 308(3):85, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
85.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-85.pdf.

Matson:2013:GSM

[1384] John Matson. Graphic science: Mission: Risk-averse. Scientific American, 308(3):88, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
88.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-88.pdf.

Pringle:2013:HEO

[1385] Heather Pringle. Human evolution: The origins of creativity. Scientific American, 308(3):36–43, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
36.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-36.pdf.

Stahler:2013:AIL

[1386] Steven W. Stahler. Astrophysics: The inner life of star clusters. Scientific American, 308(3):44–51, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
44.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-44.pdf.

Kuchment:2013:PBE

[1387] Anna Kuchment. Plant biology: The end of orange juice. Scientific American, 308(3):44–51, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
52.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-52.pdf.

Wood:2013:TFR

[1388] Robert Wood, Radhika Nagpal, and Gu-Yeon Wei. Technology: Flight of the Robobees. Scientific American, 308(3):60–65, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n3/full/scientificamerican0313-60.html;   http://www.nature.

com/scientificamerican/journal/v308/n3/pdf/scientificamerican0313-
60.pdf.

Shah:2013:EDN

[1389] Sonia Shah. Emerging diseases: New threat from poxviruses. Scientific American, 308(3):66–71, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
66.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-66.pdf.

Murphy:2013:FGW

[1390] Erin Murphy. Forensics: The government wants your DNA. Scientific American, 308(3):72–77, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
72.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-72.pdf.

Anthes:2013:BDT

[1391] Emily Anthes. Bioengineering: A dolphin’s tale. Scientific American, 308(3):78–81, March 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n3/full/scientificamerican0313-
78.html;   http://www.nature.com/scientificamerican/journal/
v308/n3/pdf/scientificamerican0313-78.pdf.

DiChristina:2013:ENA

[1392] Mariette DiChristina. From the Editor: A new agenda. Scientific American, 308(4):6, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
6.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-6.pdf.

Anonymous:2013:Ld

[1393] Anonymous. Letters. Scientific American, 308(4):8–10, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n4/full/scientificamerican0413-8.html;   http://www.nature.
com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
8.pdf.

Schlenoff:2013:YAId

[1394] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 308(4):90–91, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-90.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
90.pdf.

Anonymous:2013:SASa

[1395] Anonymous. Science agenda: The spies above your backyard. Scientific American, 308(4):12, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
12.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-12.pdf.

Yang:2013:FPS

[1396] Guobin Yang. Forum: The people’s standoff. Scientific American, 308(4): 14, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-14.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
14.pdf.

Croswell:2013:AUM

[1397] Ken Croswell. Advances: Unraveling a Magellanic mystery. Scientific American, 308(4):16, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
16.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-16.pdf.

Arnold:2013:MM

[1398] Carrie Arnold. Medieval and modern. Scientific American, 308(4):19, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-19a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
19a.pdf.

Fessenden:2013:PWd

[1399] Marissa Fessenden. Patent watch. Scientific American, 308(4):19, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-19b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
19b.pdf.

Laursen:2013:DIM

[1400] Lucas Laursen. “Doc, I’m off My Diet. LOL”. Scientific American, 308(4):20, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
20a.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-20a.pdf.

Coffey:2013:LL

[1401] Rebecca Coffey. Of lust and Lysol. Scientific American, 308(4):20, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-20b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
20b.pdf.

Fessenden:2013:EHI

[1402] Marissa Fessenden. Extremely high and incredibly cold. Scientific American, 308(4):22, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
22a.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-22a.pdf.

Biello:2013:BCS

[1403] David Biello. Better, cheaper, smaller. Scientific American, 308(4):22, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-22b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
22b.pdf.

Harmon:2013:DBF

[1404] Katherine Harmon. Dung beetles follow the stars. Scientific American, 308(4):24, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
24a.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-24a.pdf.

Wong:2013:SP

[1405] Kate Wong. Scrappy pets. Scientific American, 308(4):24, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n4/full/scientificamerican0413-24b.html;   http://www.nature.
com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
24b.pdf.

Croswell:2013:WNW

[1406] Ken Croswell. Wanted: New worlds beyond Pluto. Scientific American, 308(4):25, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
25a.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-25a.pdf.

Ferguson:2013:TC

[1407] William Ferguson. Traces of cancer. Scientific American, 308(4):25, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-25b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
25b.pdf.

Biello:2013:LS

[1408] David Biello. Let the sunshine in. Scientific American, 308(4):26, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-26.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
26.pdf.

Greenemeier:2013:TT

[1409] Larry Greenemeier. Twitching to talk. Scientific American, 308(4):27, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-27.html;   http:

//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
27.pdf.

Crew:2013:WI

[1410] Becky Crew. What is it? Scientific American, 308(4):28, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n4/full/scientificamerican0413-28.html;   http://www.nature.
com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
28.pdf.

Anderson:2013:GV

[1411] Mark Anderson. Good vibrations. Scientific American, 308(4):29, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-29.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
29.pdf.

McKenna:2013:SHS

[1412] Maryn McKenna. The science of health: Shock to the system. Scientific American, 308(4):30–32, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
30.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-30.pdf.

Pogue:2013:TDD

[1413] David Pogue. TechnoFiles: Disappointment of dazzling displays. Scientific American, 308(4):34, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
34.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-34.pdf.

Anonymous:2013:RRBa

[1414] Anonymous. Reviews and recommendations: Beautiful whale. Scientific American, 308(4):84, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
84a.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-84a.pdf.

Anonymous:2013:NDA

[1415] Anonymous. The new digital age: Reshaping the future of people, nations and business. Scientific American, 308(4):84, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n4/full/scientificamerican0413-84b.html;   http://www.nature.
com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
84b.pdf.

Anonymous:2013:A

[1416] Anonymous. Adrenaline. Scientific American, 308(4):84, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n4/full/scientificamerican0413-84c.html;   http://www.nature.
com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
84c.pdf.

Anonymous:2013:BBI

[1417] Anonymous. The book of barely imagined beings: A 21st Century bestiary. Scientific American, 308(4):84, April 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n4/full/scientificamerican0413-84d.html;   http://www.nature.
com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
84d.pdf.

Shermer:2013:SPH

[1418] Michael Shermer. Skeptic: Proof of hallucination. Scientific American, 308(4):86, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
86.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-86.pdf.

Mirsky:2013:AGSb

[1419] Steve Mirsky. Anti gravity: Show your hand. Scientific American, 308(4): 88, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-88.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
88.pdf.

Fischetti:2013:GSDa

[1420] Mark Fischetti. Graphic science: Deep heat threatens marine life. Scientific American, 308(4):92, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
92.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-92.pdf.

Hirsch:2013:PPG

[1421] Martin Hirsch, Heinrich Päs, and Werner Porod. Particle physics: Ghostly beacons of new physics. Scientific American, 308(4):40–47, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-40.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
40.pdf.

Das:2013:MMN

[1422] Sudeep Das and Tristan L. Smith. Measuring mass: The neutrino’s secrets, written on the sky. Scientific American, 308(4):46, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n4/full/scientificamerican0413-46.html;   http://www.nature.
com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
46.pdf.

Anonymous:2013:SRF

[1423] Anonymous. Special report: The future of medicine: The future of medicine. Scientific American, 308(4):48–49, April 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n4/full/scientificamerican0413-48.html;   http://www.nature.
com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
48.pdf.

Jabr:2013:CH

[1424] Ferris Jabr. A change of heart. Scientific American, 308(4):50–52, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-50.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
50.pdf.

Gorman:2013:SGC

[1425] Christine Gorman. The super glue cure. Scientific American, 308(4):52– 53, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n4/full/scientificamerican0413-52.html;   http:
//www.nature.com/scientificamerican/journal/v308/n4/pdf/scientificamerican0413-
52.pdf.

Harmon:2013:SSR

[1426] Katherine Harmon. A sweet solution for replacing organs. Scientific American, 308(4):54–55, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
54.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-54.pdf.

Jabr:2013:RBF

[1427] Ferris Jabr. Replanting the brain’s forest. Scientific American, 308(4):55–57, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
55.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-55.pdf.

Inman:2013:ETC

[1428] Mason Inman. Energy: The true cost of fossil fuels. Scientific American, 308(4):58–61, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
58.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-58.pdf.

Nelson:2013:CDA

[1429] Charles A. Nelson III, Nathan A. Fox, and Charles H. Zeanah, Jr. Child development: Anguish of the abandoned child. Scientific American, 308(4):62–67, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
62.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-62.pdf.

Stern:2013:SEL

[1430] S. Alan Stern. Space exploration: The low-cost ticket to space. Scientific American, 308(4):68–73, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
68.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-68.pdf.

Borrell:2013:ELF

[1431] Brendan Borrell. Ecology: Let the fish breathe. Scientific American, 308(4):74–79, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
74.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-74.pdf.

Switek:2013:PLA

[1432] Brian Switek. Paleontology: Love among the dinosaurs. Scientific American, 308(4):80–83, April 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n4/full/scientificamerican0413-
80.html;   http://www.nature.com/scientificamerican/journal/
v308/n4/pdf/scientificamerican0413-80.pdf.

DiChristina:2013:EMF

[1433] Mariette DiChristina. From the Editor: Making the future. Scientific American, 308(5):6, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
6.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-6.pdf.

Anonymous:2013:Le

[1434] Anonymous. Letters. Scientific American, 308(5):8–10, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-8.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
8.pdf.

Schlenoff:2013:YAIe

[1435] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 308(5):86– 87, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-86.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
86.pdf.

Anonymous:2013:SASb

[1436] Anonymous. Science agenda: The spies who sabotaged global health. Scientific American, 308(5):12, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
12.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-12.pdf.

West:2013:FWN

[1437] Geoffrey West. Forum: Wisdom in numbers. Scientific American, 308(5): 14, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-14.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
14.pdf.

Jabr:2013:ABS

[1438] Ferris Jabr. Advances: Beyond symptoms. Scientific American, 308(5): 17, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-17.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
17.pdf.

Biello:2013:TB

[1439] David Biello. The tiniest bites. Scientific American, 308(5):18, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-18a.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
18a.pdf.

Fessenden:2013:CBb

[1440] Marissa Fessenden. A cure is born. Scientific American, 308(5):18, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-18b.html;   http:

//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
18b.pdf.

Croswell:2013:TS

[1441] Ken Croswell. Tadpoles in space. Scientific American, 308(5):19, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-19.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
19.pdf.

Fessenden:2013:AI

[1442] Marissa Fessenden. Ant invasion! Scientific American, 308(5):20, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-20.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
20.pdf.

Ferguson:2013:HC

[1443] William Ferguson. Heated to the core. Scientific American, 308(5): 21, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-21.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
21.pdf.

Vance:2013:EPS

[1444] Erik Vance. Exploiting physics for serious laughs. Scientific American, 308(5):22, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
22a.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-22a.pdf.

Fessenden:2013:YFT

[1445] Marissa Fessenden. It’s your food talking. Scientific American, 308(5): 22, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-22b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
22b.pdf.

Wong:2013:YWR

[1446] Kate Wong. Yes, we’re related. Scientific American, 308(5):24, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-24a.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
24a.pdf.

Scharf:2013:SR

[1447] Caleb A. Scharf. Space race. Scientific American, 308(5):24, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-24b.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
24b.pdf.

Matson:2013:AH

[1448] John Matson. Anybody home? Scientific American, 308(5):26, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-26.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
26.pdf.

Fessenden:2013:TT

[1449] Marissa Fessenden. Tailing tuna. Scientific American, 308(5):27, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-27a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
27a.pdf.

Fessenden:2013:PWe

[1450] Marissa Fessenden. Patent watch. Scientific American, 308(5):27, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-27b.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
27b.pdf.

Matson:2013:PNC

[1451] John Matson. Preventing the next Chelyabinsk. Scientific American, 308 (5):28, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-28.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
28.pdf.

Wild:2013:WIb

[1452] Alex Wild. What is it? Scientific American, 308(5):29, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-29.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
29.pdf.

Franklin:2013:SHG

[1453] Deborah Franklin. The science of health: Good bacteria for bad breath. Scientific American, 308(5):30–32, May 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-30.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
30.pdf.

Pogue:2013:TSMa

[1454] David Pogue. TechnoFiles: The strange magic of micro movies. Scientific American, 308(5):34, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
34.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-34.pdf.

Anonymous:2013:RRBb

[1455] Anonymous. Reviews and recommendations: Brilliant blunders. Scientific American, 308(5):82, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
82a.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-82a.pdf.

Anonymous:2013:SARb

[1456] Anonymous. Scatter, adapt, and remember. Scientific American, 308(5): 82, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-82b.html;   http:

//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
82b.pdf.

Anonymous:2013:OC

[1457] Anonymous. Odd couples. Scientific American, 308(5):82, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-82c.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
82c.pdf.

Anonymous:2013:PS

[1458] Anonymous. The power surge. Scientific American, 308(5):82, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-82d.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
82d.pdf.

Shermer:2013:SGS

[1459] Michael Shermer. Skeptic: Gun science. Scientific American, 308(5): 83, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-83.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
83.pdf.

Mirsky:2013:AGBa

[1460] Steve Mirsky. Anti gravity: Bacon, lettuce and tasteless. Scientific American, 308(5):84, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
84.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-84.pdf.

Matson:2013:GSG

[1461] John Matson. Graphic science: Gender gaps. Scientific American, 308 (5):88, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-88.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
88.pdf.

Hausmann:2013:SRF

[1462] Ricardo Hausmann. Special report: The future of manufacturing: How to make the next big thing. Scientific American, 308(5):36–37, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-36.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
36.pdf.

Bourne:2013:MBR

[1463] David Bourne. My boss the robot. Scientific American, 308(5):38– 41, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-38.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
38.pdf.

Ashley:2013:FS

[1464] Steven Ashley. Future stuff. Scientific American, 308(5):42–43, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-42.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
42.pdf.

Greenemeier:2013:PI

[1465] Larry Greenemeier. To print the impossible. Scientific American, 308(5): 44–47, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-44.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
44.pdf.

Roco:2013:RNM

[1466] Mihail C. Roco. Rise of the nano machines. Scientific American, 308(5): 48–49, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-48.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
48.pdf.

Myers:2013:AC

[1467] James D. Myers. Assembled in code. Scientific American, 308(5):50– 51, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n5/full/scientificamerican0513-50.html;   http:
//www.nature.com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
50.pdf.

Walker:2013:NSD

[1468] Lary C. Walker and Mathias Jucker. Neuroscience: Seeds of dementia. Scientific American, 308(5):52–57, May 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-52.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
52.pdf.

McEwen:2013:PSM

[1469] Alfred S. McEwen. Planetary science: Mars in motion. Scientific American, 308(5):58–65, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
58.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-58.pdf.

Hammer:2013:EHH

[1470] Michael F. Hammer. Evolution: Human hybrids. Scientific American, 308(5):66–71, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
66.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-66.pdf.

Solomon:2013:EDH

[1471] Christopher Solomon. Emerging diseases: How kitty is killing the dolphins. Scientific American, 308(5):72–77, May 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n5/full/scientificamerican0513-72.html;   http://www.nature.
com/scientificamerican/journal/v308/n5/pdf/scientificamerican0513-
72.pdf.

Everts:2013:BMM

[1472] Sarah Everts. Biochemistry: Mirror molecules. Scientific American, 308(5):78–81, May 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n5/full/scientificamerican0513-
78.html;   http://www.nature.com/scientificamerican/journal/
v308/n5/pdf/scientificamerican0513-78.pdf.

DiChristina:2013:ESJ

[1473] Mariette DiChristina. From the Editor: A science journey. Scientific American, 308(6):6, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
6.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-6.pdf.

Anonymous:2013:Lf

[1474] Anonymous. Letters. Scientific American, 308(6):8–10, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-8.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
8.pdf.

Schlenoff:2013:YAIf

[1475] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 308(6):94–95, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n6/full/scientificamerican0613-94.html;   http:
//www.nature.com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
94.pdf.

Anonymous:2013:SADb

[1476] Anonymous. Science agenda: Do not reanimate. Scientific American, 308(6):12, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
12.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-12.pdf.

Kappos:2013:FTI

[1477] David J. Kappos. Forum: Tweaks to the idea factory. Scientific American, 308(6):14, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
14.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-14.pdf.

Moyer:2013:ACB

[1478] Michael Moyer. Advances: Cloud to black hole: Eat my dust. Scientific American, 308(6):16, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
16.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-16.pdf.

Underhill:2013:WB

[1479] William Underhill. What’s the buzz? Scientific American, 308(6):19, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n6/full/scientificamerican0613-19a.html;   http:
//www.nature.com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
19a.pdf.

Westly:2013:DWS

[1480] Erica Westly. The doctor will see you in aisle 3. Scientific American, 308(6):19, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
19b.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-19b.pdf.

Laursen:2013:FF

[1481] Lucas Laursen. Faux fossils. Scientific American, 308(6):20, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-20.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
20.pdf.

Choi:2013:SC

[1482] Charles Q. Choi. Soft circuits. Scientific American, 308(6):21, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-21a.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
21a.pdf.

Anonymous:2013:Na

[1483] Anonymous. By the numbers. Scientific American, 308(6):21, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-21b.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
21b.pdf.

Nuwer:2013:MM

[1484] Rachel Nuwer. Mediterranean mystery. Scientific American, 308(6):22, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n6/full/scientificamerican0613-22.html;   http:
//www.nature.com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
22.pdf.

Greenemeier:2013:CI

[1485] Larry Greenemeier. Cool it. Scientific American, 308(6):23, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-23.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
23.pdf.

Wong:2013:AF

[1486] Kate Wong. All in the family. Scientific American, 308(6):24, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-24a.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
24a.pdf.

Greenemeier:2013:RW

[1487] Larry Greenemeier. Real world. Scientific American, 308(6):24, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n6/full/scientificamerican0613-24b.html;   http:

//www.nature.com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
24b.pdf.

Choi:2013:NPB

[1488] Charles Q. Choi. “nurse, pass the blowtorch”. Scientific American, 308(6):26, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
26.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-26.pdf.

Fessenden:2013:PWf

[1489] Marissa Fessenden. Patent watch. Scientific American, 308(6):27, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n6/full/scientificamerican0613-27.html;   http:
//www.nature.com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
27.pdf.

Bushwick:2013:WI

[1490] Sophie Bushwick. What is it? Scientific American, 308(6):28, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-28a.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
28a.pdf.

Anonymous:2013:Nb

[1491] Anonymous. By the numbers. Scientific American, 308(6):28, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-28b.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
28b.pdf.

Matson:2013:CH

[1492] John Matson. Closer to home. Scientific American, 308(6):29, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-29.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
29.pdf.

Biello:2013:RP

[1493] David Biello. Running on poison. Scientific American, 308(6):30, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n6/full/scientificamerican0613-30.html;   http:
//www.nature.com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
30.pdf.

Khamsi:2013:SHG

[1494] Roxanne Khamsi. The science of health: Going to pot. Scientific American, 308(6):34–36, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
34.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-34.pdf.

Pogue:2013:TGC

[1495] David Pogue. TechnoFiles: Google’s creep factor. Scientific American, 308(6):38, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
38.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-38.pdf.

Anonymous:2013:RRGa

[1496] Anonymous. Reviews and recommendations: Ginkgo. Scientific American, 308(6):88, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
88a.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-88a.pdf.

Anonymous:2013:PAC

[1497] Anonymous. Probably approximately correct. Scientific American, 308(6):88, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
88b.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-88b.pdf.

Anonymous:2013:HP

[1498] Anonymous. High price. Scientific American, 308(6):88, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-88c.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
88c.pdf.

Anonymous:2013:HWD

[1499] Anonymous. How we do it. Scientific American, 308(6):88, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-88d.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
88d.pdf.

Shermer:2013:SCS

[1500] Michael Shermer. Skeptic: CSI, science. Scientific American, 308(6):90, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n6/full/scientificamerican0613-90.html;   http:
//www.nature.com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
90.pdf.

Mirsky:2013:AGR

[1501] Steve Mirsky. Anti gravity: Rumen with a view. Scientific American, 308(6):92, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
92.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-92.pdf.

Fischetti:2013:GSDb

[1502] Mark Fischetti. Graphic science: The danger of opting out. Scientific American, 308(6):96, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
96.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-96.pdf.

Martin:2013:LST

[1503] Ronald Martin and Antonietta Quigg. Life science: Tiny plants that once ruled the seas. Scientific American, 308(6):40–45, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-40.html;   http://www.nature.

com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
40.pdf.

vonBaeyer:2013:PQW

[1504] Hans Christian von Baeyer. Physics: Quantum weirdness? It’s all in your mind. Scientific American, 308(6):46–51, June 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-46.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
46.pdf.

Interlandi:2013:NBB

[1505] Jeneen Interlandi. Neuroscience: Breaking the brain barrier. Scientific American, 308(6):52–57, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
52.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-52.pdf.

Fischetti:2013:CSS

[1506] Mark Fischetti. Climate science: Storm of the century every two years. Scientific American, 308(6):58–67, June 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v308/
n6/full/scientificamerican0613-58.html;   http://www.nature.
com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
58.pdf.

Scerri:2013:CCP

[1507] Eric Scerri. Chemistry: Cracks in the periodic table. Scientific American, 308(6):68–73, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
68.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-68.pdf.

Fessenden:2013:BRC

[1508] Marissa Fessenden. Botany: Ring cycle. Scientific American, 308(6):74– 75, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v308/n6/full/scientificamerican0613-74.html;   http:

//www.nature.com/scientificamerican/journal/v308/n6/pdf/scientificamerican0613-
74.pdf.

Yong:2013:PAA

[1509] Ed Yong. Psychology: Armor against prejudice. Scientific American, 308(6):76–80, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
76.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-76.pdf.

Greenemeier:2013:ITE

[1510] Larry Greenemeier. Information technology: Edge of the Internet. Scientific American, 308(6):82–86, June 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v308/n6/full/scientificamerican0613-
82.html;   http://www.nature.com/scientificamerican/journal/
v308/n6/pdf/scientificamerican0613-82.pdf.

DiChristina:2013:ESCb

[1511] Mariette DiChristina. From the Editor: Scientific citizens. Scientific American, 309(1):4, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
4.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-4.pdf.

Anonymous:2013:Lg

[1512] Anonymous. Letters. Scientific American, 309(1):6–8, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n1/full/scientificamerican0713-6.html;   http://www.nature.
com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
6.pdf.

Anonymous:2013:SAE

[1513] Anonymous. Science agenda: Exploded trust. Scientific American, 309 (1):10, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-10.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
10.pdf.

Kahn:2013:FWT

[1514] Philippe Kahn. Forum: Wear them, forget them. Scientific American, 309(1):12, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
12.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-12.pdf.

Weintraub:2013:APP

[1515] Arlene Weintraub. Advances: Prickly painkiller. Scientific American, 309 (1):14, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-14.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
14.pdf.

Anonymous:2013:Nc

[1516] Anonymous. By the numbers. Scientific American, 309(1):17, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n1/full/scientificamerican0713-17b.html;   http://www.nature.
com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
17b.pdf.

Greenemeier:2013:GD

[1517] Larry Greenemeier. Going deep. Scientific American, 309(1):17, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-17a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
17a.pdf.

Kuchment:2013:PLP

[1518] Anna Kuchment. Produce-loving pests. Scientific American, 309(1): 18, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-18.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
18.pdf.

Fessenden:2013:AA

[1519] Marissa Fessenden. Art from above. Scientific American, 309(1):20, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-20.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
20.pdf.

Greenemeier:2013:FSG

[1520] Larry Greenemeier. The future of smart glasses. Scientific American, 309 (1):21, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-21.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
21.pdf.

Ravindran:2013:CW

[1521] Shruti Ravindran. Cull of the wild. Scientific American, 309(1):22, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-22.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
22.pdf.

Fessenden:2013:PWg

[1522] Marissa Fessenden. Patent watch. Scientific American, 309(1):23, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-23b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
23b.pdf.

Matson:2013:FP

[1523] John Matson. Frothy physics. Scientific American, 309(1):23, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n1/full/scientificamerican0713-23a.html;   http://www.nature.
com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
23a.pdf.

Crew:2013:DPM

[1524] Becky Crew. Don’t pince me, please. Scientific American, 309(1):24, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-24a.html;   http:

//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
24a.pdf.

Jabr:2013:SC

[1525] Ferris Jabr. Streams of consciousness. Scientific American, 309(1):24, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-24b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
24b.pdf.

Anonymous:2013:Nd

[1526] Anonymous. By the numbers. Scientific American, 309(1):25, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n1/full/scientificamerican0713-25b.html;   http://www.nature.
com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
25b.pdf.

Matson:2013:SLB

[1527] John Matson. A star’s last breath. Scientific American, 309(1):25, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-25a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
25a.pdf.

Greenemeier:2013:LCA

[1528] Larry Greenemeier. Lights, camera, atoms. Scientific American, 309(1): 26, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-26.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
26.pdf.

Storrs:2013:SHD

[1529] Carina Storrs. The science of health: Do CT scans cause cancer? Scientific American, 309(1):30–32, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
30.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-30.pdf.

Pogue:2013:TSMb

[1530] David Pogue. TechnoFiles: In search of a mind-reading machine. Scientific American, 309(1):34, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
34.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-34.pdf.

Lemonick:2013:ADD

[1531] Michael D. Lemonick. Astronomy: The dawn of distant skies. Scientific American, 309(1):40–47, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
40.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-40.pdf.

Sinha:2013:MOB

[1532] Pawan Sinha. Medicine: Once blind and now they see. Scientific American, 309(1):48–55, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
48.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-48.pdf.

Biello:2013:EGG

[1533] David Biello. Environment: Greenhouse goo. Scientific American, 309(1):56–61, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
56.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-56.pdf.

King:2013:ABW

[1534] Barbara J. King. Animal behavior: When animals mourn. Scientific American, 309(1):62–67, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
62.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-62.pdf.

Jabr:2013:CNG

[1535] Ferris Jabr. Chemistry: a Nobel gathering. Scientific American, 309(1): 68–79, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-68.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
68.pdf.

Lehrman:2013:GHD

[1536] Sally Lehrman. Global health: The diabolical genius of an ancient scourge. Scientific American, 309(1):80–85, July 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n1/full/scientificamerican0713-80.html;   http://www.nature.
com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
80.pdf.

Mackenzie:2013:PWW

[1537] Dana Mackenzie. Physics: Walls of water. Scientific American, 309(1): 86–89, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-86.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
86.pdf.

Anonymous:2013:CIS

[1538] Anonymous. The cartoon introduction to statistics. Scientific American, 309(1):90, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
90b.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-90b.pdf.

Anonymous:2013:RG

[1539] Anonymous. Rocket girl. Scientific American, 309(1):90, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n1/full/scientificamerican0713-90d.html;   http://www.nature.
com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
90d.pdf.

Anonymous:2013:RRPa

[1540] Anonymous. Reviews and recommendations: People, parasites, and plowshares. Scientific American, 309(1):90, July 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n1/full/scientificamerican0713-90a.html;   http://www.nature.
com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
90a.pdf.

Anonymous:2013:TN

[1541] Anonymous. Thinking in numbers. Scientific American, 309(1):90, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-90c.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
90c.pdf.

Shermer:2013:SGG

[1542] Michael Shermer. Skeptic: Gods of the gaps. Scientific American, 309(1): 92, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-92.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
92.pdf.

Mirsky:2013:AGBb

[1543] Steve Mirsky. Anti gravity: Bingo was not his name-o. Scientific American, 309(1):94, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-
94.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-94.pdf.

Schlenoff:2013:YAIg

[1544] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 309(1):96– 97, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n1/full/scientificamerican0713-96.html;   http:
//www.nature.com/scientificamerican/journal/v309/n1/pdf/scientificamerican0713-
96.pdf.

Fischetti:2013:GSF

[1545] Mark Fischetti. Graphic science: Family, friends and gunshot. Scientific American, 309(1):98, July 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n1/full/scientificamerican0713-

98.html;   http://www.nature.com/scientificamerican/journal/
v309/n1/pdf/scientificamerican0713-98.pdf.

DiChristina:2013:EBB

[1546] Mariette DiChristina. From the Editor: Bits for brains. Scientific American, 309(2):6, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
6.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-6.pdf.

Anonymous:2013:Lh

[1547] Anonymous. Letters. Scientific American, 309(2):8–9, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n2/full/scientificamerican0813-8.html;   http://www.nature.
com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
8.pdf.

Editors:2013:SAS

[1548] The Editors. Science agenda: Solitary is cruel and unusual. Scientific American, 309(2):10, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
10.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-10.pdf.

Lincoln:2013:FBG

[1549] Don Lincoln. Forum: Beyond the God particle. Scientific American, 309(2):12, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
12.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-12.pdf.

Branswell:2013:AHD

[1550] Helen Branswell. Advances: Hidden and dangerous. Scientific American, 309(2):14–15, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
14.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-14.pdf.

Anonymous:2013:Ne

[1551] Anonymous. By the numbers. Scientific American, 309(2):15, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-15b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
15b.pdf.

Mone:2013:LRE

[1552] Gregory Mone. Lending robots an ear. Scientific American, 309(2):15, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-15a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
15a.pdf.

Vance:2013:AD

[1553] Erik Vance. The art of distraction. Scientific American, 309(2):16, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-16.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
16.pdf.

Jabr:2013:ACG

[1554] Ferris Jabr. Anxiety can be good for you — sometimes. Scientific American, 309(2):19, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
19a.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-19a.pdf.

Wong:2013:FCF

[1555] Kate Wong. Fact-checking a frozen mammoth. Scientific American, 309(2):19, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
19b.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-19b.pdf.

Anonymous:2013:Nf

[1556] Anonymous. By the numbers. Scientific American, 309(2):20, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-20b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
20b.pdf.

Biello:2013:CBG

[1557] David Biello. The coming boom in geothermal fracking. Scientific American, 309(2):20, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
20a.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-20a.pdf.

Gravotta:2013:WI

[1558] Luciana Gravotta. What is it? Scientific American, 309(2):21, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-21.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
21.pdf.

Roulstone:2013:CSS

[1559] Ian Roulstone and John Norbury. Computing Superstorm Sandy. Scientific American, 309(2):22, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
22.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-22.pdf.

Fessenden:2013:MMM

[1560] Marissa Fessenden. Medicine’s most minuscule frontier. Scientific American, 309(2):23, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
23.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-23.pdf.

Choi:2013:BVS

[1561] Charles Q. Choi. Bug vs. superbug. Scientific American, 309(2):24, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-24a.html;   http:

//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
24a.pdf.

Duhaime-Ross:2013:PW

[1562] Arielle Duhaime-Ross. Patent watch. Scientific American, 309(2):24, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-24b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
24b.pdf.

Cowen:2013:PP

[1563] Ron Cowen. Phantom planets. Scientific American, 309(2):25, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-25.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
25.pdf.

Moyer:2013:MM

[1564] Melinda Wenner Moyer. A microbial mystery. Scientific American, 309(2):26, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
26.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-26.pdf.

McKenna:2013:SHR

[1565] Maryn McKenna. The science of health: Rethinking rabies. Scientific American, 309(2):28–30, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
28.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-28.pdf.

Pogue:2013:TLT

[1566] David Pogue. TechnoFiles: The last thing you’ll memorize. Scientific American, 309(2):32, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
32.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-32.pdf.

Tononi:2013:NPP

[1567] Giulio Tononi and Chiara Cirelli. Neuroscience: Perchance to prune. Scientific American, 309(2):34–39, August 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n2/full/scientificamerican0813-34.html;   http://www.nature.
com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
34.pdf.

Kuhlmann:2013:QPW

[1568] Meinard Kuhlmann. Quantum physics: What is real? Scientific American, 309(2):40–47, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
40.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-40.pdf.

Editors:2013:SRL

[1569] The Editors. Special report: Learning in The Digital Age: Big data goes to school. Scientific American, 309(2):48–52, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n2/full/scientificamerican0813-48.html;   http://www.nature.
com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
48.pdf.

Bartholet:2013:HH

[1570] Jeffrey Bartholet. Hype and hope. Scientific American, 309(2):53–61, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-53.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
53.pdf.

Norvig:2013:MP

[1571] Peter Norvig. Massively personal. Scientific American, 309(2):55, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-55.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
55.pdf.

Khan:2013:NML

[1572] Salman Khan. No more lockstep learning. Scientific American, 309(2): 57, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-57.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
57.pdf.

Agarwal:2013:OI

[1573] Pawan Agarwal. An opportunity for India. Scientific American, 309(2):58–59, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
58.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-58.pdf.

Lue:2013:EK

[1574] Robert A. Lue. “Experimentation is key”. Scientific American, 309(2):60–61, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
60.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-60.pdf.

Fletcher:2013:ML

[1575] Seth Fletcher. Machine learning. Scientific American, 309(2):62–68, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-62.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
62.pdf.

Ravitch:2013:PP

[1576] Diane Ravitch. Promise and peril. Scientific American, 309(2):66–67, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-66.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
66.pdf.

Vesterbacka:2013:EAB

[1577] Peter Vesterbacka. Education is for the (angry) birds. Scientific American, 309(2):68, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
68.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-68.pdf.

Duncan:2013:WWN

[1578] Arne Duncan. Why we need high-speed schools. Scientific American, 309(2):69–71, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
69.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-69.pdf.

Bartholet:2013:WPS

[1579] Jeffrey Bartholet. “I was pleasantly surprised”. Scientific American, 309(2):72–73, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
72.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-72.pdf.

Bassuk:2013:HWE

[1580] Shari S. Bassuk, Timothy S. Church, and JoAnn E. Manson. Health: Why exercise works magic. Scientific American, 309(2):74–79, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-74.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
74.pdf.

Nadeau:2013:ERW

[1581] Barbie Latza Nadeau. Engineering: Raising the wreck. Scientific American, 309(2):80–83, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
80.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-80.pdf.

Zimmer:2013:ESO

[1582] Carl Zimmer. Evolution: The surprising origins of life’s complexity. Scientific American, 309(2):84–89, August 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n2/full/scientificamerican0813-84.html;   http://www.nature.
com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
84.pdf.

Jabr:2013:FDH

[1583] Ferris Jabr. Fluid dynamics: How to swim in molasses. Scientific American, 309(2):90–95, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
90.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-90.pdf.

Anonymous:2013:BF

[1584] Anonymous. Betting on famine. Scientific American, 309(2):96, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-96b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
96b.pdf.

Anonymous:2013:RRWb

[1585] Anonymous. Reviews and recommendations: The West without water. Scientific American, 309(2):96, August 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n2/full/scientificamerican0813-96a.html;   http://www.nature.
com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
96a.pdf.

Anonymous:2013:SG

[1586] Anonymous. The sports gene. Scientific American, 309(2):96, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-96d.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
96d.pdf.

Anonymous:2013:TSD

[1587] Anonymous. The truth in small doses. Scientific American, 309(2):96, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-96c.html;   http:

//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
96c.pdf.

Shermer:2013:SFM

[1588] Michael Shermer. Skeptic: Five myths of terrorism. Scientific American, 309(2):98, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
98.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-98.pdf.

Mirsky:2013:AGE

[1589] Steve Mirsky. Anti gravity: Embrace the ick. Scientific American, 309(2): 99, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-99.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
99.pdf.

Schlenoff:2013:YAIh

[1590] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 309(2):100, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-100.html;   http:
//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
100.pdf.

Fischetti:2013:GSL

[1591] Mark Fischetti. Graphic science: Living in harm’s way. Scientific American, 309(2):102, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n2/full/scientificamerican0813-
102.html;   http://www.nature.com/scientificamerican/journal/
v309/n2/pdf/scientificamerican0813-102.pdf.

Anonymous:2013:SRL

[1592] Anonymous. Special report: Learning in The Digital Age: Helping curious minds achieve great things. Scientific American, 309(2):103–106, August 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n2/full/scientificamerican0813-103A.html;   http:

//www.nature.com/scientificamerican/journal/v309/n2/pdf/scientificamerican0813-
103A.pdf.

DiChristina:2013:ESF

[1593] Mariette DiChristina. From the Editor: a scientific feast. Scientific American, 309(3):4, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
4.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-4.pdf.

Anonymous:2013:Li

[1594] Anonymous. Letters. Scientific American, 309(3):6–7, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n3/full/scientificamerican0913-6.html;   http://www.nature.
com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
6.pdf.

Editors:2013:SAF

[1595] The Editors. Science agenda: Fight the GM food scare. Scientific American, 309(3):10, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
10.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-10.pdf.

Church:2013:FPR

[1596] George Church. Forum: Please reanimate. Scientific American, 309(3):12, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
12.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-12.pdf.

Ravindran:2013:ADB

[1597] Shruti Ravindran. Advances: “disgusting bird” is dying off. Scientific American, 309(3):14, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
14.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-14.pdf.

Moyer:2013:WS

[1598] Melinda Wenner Moyer. Wisdom of the sleepyheads. Scientific American, 309(3):15, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
15.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-15.pdf.

Duhaime-Ross:2013:SG

[1599] Arielle Duhaime-Ross. The spider in the grass. Scientific American, 309(3):16, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
16.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-16.pdf.

Croswell:2013:GBB

[1600] Ken Croswell. Galaxies bright and beautiful. Scientific American, 309(3):19, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
19a.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-19a.pdf.

DiChristina:2013:TPP

[1601] Mariette DiChristina and Rachel Scheer. Turning peels into plastic. Scientific American, 309(3):19, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n3/full/scientificamerican0913-
19b.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-19b.pdf.

Worth:2013:CUT

[1602] Katie Worth. Courtesy of the U.S. taxpayer. Scientific American, 309(3):20, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
20.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-20.pdf.

Fischetti:2013:FTD

[1603] Mark Fischetti. Fracking and tainted drinking water. Scientific American, 309(3):21, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
21.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-21.pdf.

Anonymous:2013:Ng

[1604] Anonymous. By the numbers. Scientific American, 309(3):22, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-22b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
22b.pdf.

Fessenden:2013:BBE

[1605] Marissa Fessenden. A baby breathes easier. Scientific American, 309(3):22, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
22a.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-22a.pdf.

Hodge:2013:SSS

[1606] Anne-Marie C. Hodge. Sizing up sustainable seafood. Scientific American, 309(3):23, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
23.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-23.pdf.

Collins:2013:SPR

[1607] Nathan Collins. Stop pampering the Red Planet. Scientific American, 309(3):24, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
24a.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-24a.pdf.

Harmon:2013:PD

[1608] Katherine Harmon. Pathogens decoded. Scientific American, 309(3):24, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-24b.html;   http:

//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
24b.pdf.

Guterl:2013:WI

[1609] Sophie Guterl. What is it? Scientific American, 309(3):25, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-25.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
25.pdf.

Crew:2013:SWS

[1610] Becky Crew. In search of a wild song. Scientific American, 309(3):26, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-26b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
26b.pdf.

Stix:2013:CMG

[1611] Gary Stix. The case for milk is going sour. Scientific American, 309(3):26, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-26a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
26a.pdf.

Chen:2013:SHL

[1612] Ingfei Chen. The science of health: Lead’s buried legacy. Scientific American, 309(3):28–30, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
28.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-28.pdf.

Pogue:2013:TDU

[1613] David Pogue. TechnoFiles: Death to the upgrade. Scientific American, 309(3):32, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
32.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-32.pdf.

Moyer:2013:IFI

[1614] Michael Moyer. Introduction: The food issue. Scientific American, 309(3):34–39, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
34.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-34.pdf.

Lai:2013:FCO

[1615] Bun Lai. Feast: Celebrating our love of food: How (and why) to eat invasive species. Scientific American, 309(3):40–43, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n3/full/scientificamerican0913-40.html;   http://www.nature.
com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
40.pdf.

Kenny:2013:FA

[1616] Paul J. Kenny. The food addiction. Scientific American, 309(3):44–49, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-44.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
44.pdf.

Kim:2013:AMY

[1617] Evelyn Kim. The amazing multimillion-year history of processed food. Scientific American, 309(3):50–55, September 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n3/full/scientificamerican0913-50.html;   http://www.nature.
com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
50.pdf.

Dunn:2013:FHF

[1618] Rob Dunn. Fuel: How food changes us: Everything you know about calories is wrong. Scientific American, 309(3):56–59, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n3/full/scientificamerican0913-56.html;   http://www.nature.
com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
56.pdf.

Taubes:2013:WOW

[1619] Gary Taubes. Which one will make you fat? Scientific American, 309(3):60–65, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
60.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-60.pdf.

Wong:2013:FC

[1620] Kate Wong. The first cookout. Scientific American, 309(3):66–69, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-66.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
66.pdf.

Rosner:2013:FFF

[1621] Hillary Rosner. Farm: The future of the food supply: Return of the natives. Scientific American, 309(3):70–75, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n3/full/scientificamerican0913-70.html;   http://www.nature.
com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
70.pdf.

Conniff:2013:SD

[1622] Richard Conniff. Super dirt. Scientific American, 309(3):76–79, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-76.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
76.pdf.

Freedman:2013:EFE

[1623] David H. Freedman. Are engineered foods evil? Scientific American, 309(3):80–85, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
80.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-80.pdf.

Anonymous:2013:BGa

[1624] Anonymous. Brave genius. Scientific American, 309(3):90, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-90d.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
90d.pdf.

Anonymous:2013:BSM

[1625] Anonymous. Behind the shock machine. Scientific American, 309(3):90, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-90c.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
90c.pdf.

Anonymous:2013:C

[1626] Anonymous. Countdown. Scientific American, 309(3):90, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-90b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
90b.pdf.

Anonymous:2013:RRGb

[1627] Anonymous. Reviews and recommendations: Giraffe reflections. Scientific American, 309(3):90, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n3/full/scientificamerican0913-
90a.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-90a.pdf.

Shermer:2013:SDK

[1628] Michael Shermer. Skeptic: The dangers of keeping an open mind. Scientific American, 309(3):92, September 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n3/full/scientificamerican0913-92.html;   http://www.nature.
com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
92.pdf.

Mirsky:2013:AGT

[1629] Steve Mirsky. Anti gravity: Through a glass, obviously. Scientific American, 309(3):93, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
93.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-93.pdf.

Schlenoff:2013:YAIi

[1630] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 309(3):94–95, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n3/full/scientificamerican0913-94.html;   http:
//www.nature.com/scientificamerican/journal/v309/n3/pdf/scientificamerican0913-
94.pdf.

Fischetti:2013:GSM

[1631] Mark Fischetti. Graphic science: The 5,000-mile salad. Scientific American, 309(3):96, September 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n3/full/scientificamerican0913-
96.html;   http://www.nature.com/scientificamerican/journal/
v309/n3/pdf/scientificamerican0913-96.pdf.

DiChristina:2013:ENI

[1632] Mariette DiChristina. From the Editor: Nexus of invention. Scientific American, 309(4):6, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
6.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-6.pdf.

Anonymous:2013:Lj

[1633] Anonymous. Letters. Scientific American, 309(4):8–10, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-8.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
8.pdf.

Anonymous:2013:SAF

[1634] Anonymous. Science agenda: Fiddling while the world warms. Scientific American, 309(4):12, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
12.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-12.pdf.

Scott:2013:FCC

[1635] Eugenie C. Scott and Minda Berbeco. Forum: Climate in the classroom. Scientific American, 309(4):14, October 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-14.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
14.pdf.

Greenspan:2013:ACC

[1636] Jesse Greenspan. Advances: Coyotes in the crosswalks? Fuggedaboutit! Scientific American, 309(4):17, October 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-17.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
17.pdf.

Anonymous:2013:Nh

[1637] Anonymous. By the numbers. Scientific American, 309(4):18, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-18b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
18b.pdf.

Maron:2013:DWG

[1638] Dina Fine Maron. Doing a world of good. Scientific American, 309(4):18, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-18a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
18a.pdf.

Billings:2013:PK

[1639] Lee Billings. The perfect Kelvin. Scientific American, 309(4):19, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-19.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
19.pdf.

Anonymous:2013:Ni

[1640] Anonymous. By the numbers. Scientific American, 309(4):20, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-20b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
20b.pdf.

Choi:2013:UB

[1641] Charles Q. Choi. Universe out of balance. Scientific American, 309(4):20, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-20a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
20a.pdf.

Duhaime-Ross:2013:HL

[1642] Arielle Duhaime-Ross. The Holey Land. Scientific American, 309(4):21, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-21.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
21.pdf.

Matson:2013:WI

[1643] John Matson. What is it? Scientific American, 309(4):22, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-22.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
22.pdf.

Greenwood:2013:TBM

[1644] Veronique Greenwood. Taste-blind mice make tangled sperm. Scientific American, 309(4):23, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
23.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-23.pdf.

Hayes:2013:GE

[1645] Brian Hayes. The great explicator. Scientific American, 309(4):24, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-24.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
24.pdf.

Platt:2013:MTB

[1646] John R. Platt. Madagascar’s towering baobab trees on the brink. Scientific American, 309(4):26, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
26.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-26.pdf.

Maron:2013:LTD

[1647] Dina Fine Maron. The liver transplant divide. Scientific American, 309(4):28, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
28.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-28.pdf.

Anonymous:2013:Nj

[1648] Anonymous. By the numbers. Scientific American, 309(4):29, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-29c.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
29c.pdf.

Duhaime-Ross:2013:TR

[1649] Arielle Duhaime-Ross. Total recall. Scientific American, 309(4):29, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-29a.html;   http:

//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
29a.pdf.

Greenemeier:2013:HCI

[1650] Larry Greenemeier. Hackers crack the iPhone. Scientific American, 309(4):29, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
29b.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-29b.pdf.

McKenna:2013:SHP

[1651] Maryn McKenna. The science of health: The pertussis parable. Scientific American, 309(4):34–36, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
34.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-34.pdf.

Pogue:2013:TEY

[1652] David Pogue. TechnoFiles: Edit your photos? Feed the meter. Scientific American, 309(4):38, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
38.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-38.pdf.

Andersen:2013:CEB

[1653] Ross D. Andersen. Cosmology: An ear to the big bang. Scientific American, 309(4):40–47, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
40.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-40.pdf.

Pringle:2013:ELL

[1654] Heather Pringle. Evolution: Long live the humans. Scientific American, 309(4):48–55, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
48.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-48.pdf.

Anonymous:2013:SRS

[1655] Anonymous. Special report: State of the world’s science 2013: Crossroads of invention. Scientific American, 309(4):56–57, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-56.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
56.pdf.

Kappos:2013:WWB

[1656] David J. Kappos. Who will bankroll the next big idea? Scientific American, 309(4):58–61, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
58.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-58.pdf.

Suresh:2013:PMM

[1657] Subra Suresh. The power of many minds. Scientific American, 309(4):60, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-60.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
60.pdf.

Branstetter:2013:PPB

[1658] Lee Branstetter, Guangwei Li, and Francisco Veloso. The polyglot patent boom. Scientific American, 309(4):62–63, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-62.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
62.pdf.

Anonymous:2013:IS

[1659] Anonymous. Innovation scorecard. Scientific American, 309(4):64–65, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-64.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
64.pdf.

Vance:2013:WCM

[1660] Erik Vance. Why can’t Mexico make science pay off? Scientific American, 309(4):66–71, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
66.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-66.pdf.

Lange:2013:NHC

[1661] Nicholas Lange and Christopher J. McDougle. Neuroscience: Help for the child with autism. Scientific American, 309(4):72–77, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-72.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
72.pdf.

Pentland:2013:ITD

[1662] Alex ‘Sandy’ Pentland. Information technology: The data-driven society. Scientific American, 309(4):78–83, October 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-78.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
78.pdf.

Courage:2013:RHB

[1663] Katherine Harmon Courage. Robotics: How to build a robot octopus. Scientific American, 309(4):84–87, October 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n4/full/scientificamerican1013-84.html;   http://www.nature.
com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
84.pdf.

Conant:2013:ERN

[1664] Eve Conant. Energy: Russia’s new empire: Nuclear power. Scientific American, 309(4):88–93, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
88.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-88.pdf.

Anonymous:2013:FBY

[1665] Anonymous. Five billion years of solitude. Scientific American, 309(4): 94, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-94c.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
94c.pdf.

Anonymous:2013:LM

[1666] Anonymous. Love and math. Scientific American, 309(4):94, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-94d.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
94d.pdf.

Anonymous:2013:LSL

[1667] Anonymous. Life at the speed of light. Scientific American, 309(4):94, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-94b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
94b.pdf.

Anonymous:2013:RRPb

[1668] Anonymous. Reviews and recommendations: Primates of the world. Scientific American, 309(4):94, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n4/full/scientificamerican1013-
94a.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-94a.pdf.

Shermer:2013:SWS

[1669] Michael Shermer. Skeptic: When science doesn’t support beliefs. Scientific American, 309(4):95, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
95.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-95.pdf.

Mirsky:2013:AGBc

[1670] Steve Mirsky. Anti gravity: Bouillabaisse biology. Scientific American, 309(4):96, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n4/full/scientificamerican1013-
96.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-96.pdf.

Anonymous:2013:YAI

[1671] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 309(4):98–99, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n4/full/scientificamerican1013-98.html;   http:
//www.nature.com/scientificamerican/journal/v309/n4/pdf/scientificamerican1013-
98.pdf.

Matson:2013:GSR

[1672] John Matson. Graphic science: Renewable energy’s hidden costs. Scientific American, 309(4):100, October 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n4/full/scientificamerican1013-
100.html;   http://www.nature.com/scientificamerican/journal/
v309/n4/pdf/scientificamerican1013-100.pdf.

Moskowitz:2013:PMB

[1673] Clara Moskowitz. Puzzling measurement of “Big G” gravitational constant ignites debate. Scientific American, ??(??):??, September 18, 2013. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.scientificamerican.com/
article.cfm?id=puzzling-measurement-of-big-g-gravitational-
constant-ignites-debate-slide-show.

DiChristina:2013:EFD

[1674] Mariette DiChristina. From the editor: Following data trails. Scientific American, 309(5):4, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
4.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-4.pdf.

Anonymous:2013:Lk

[1675] Anonymous. Letters. Scientific American, 309(5):6–7, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n5/full/scientificamerican1113-6.html;   http://www.nature.

com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
6.pdf.

Schlenoff:2013:YAIj

[1676] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 309(5):81, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-81.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
81.pdf.

Anonymous:2013:SADc

[1677] Anonymous. Science agenda: Data glitches are hazardous to your health. Scientific American, 309(5):10, November 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n5/full/scientificamerican1113-10.html;   http://www.nature.
com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
10.pdf.

Rangaswami:2013:FMH

[1678] J. P. Rangaswami. Forum: Manufacturing, Hollywood-style. Scientific American, 309(5):12, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
12.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-12.pdf.

Moyer:2013:ACP

[1679] Michael Moyer. Advances: a cosmos of particles. Scientific American, 309(5):15, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
15.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-15.pdf.

Duhaime-Ross:2013:NRE

[1680] Arielle Duhaime-Ross. A new race to Earth’s end. Scientific American, 309(5):16, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-

16a.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-16a.pdf.

Smith:2013:RYB

[1681] Peter Andrey Smith. Reading your bacterial aura. Scientific American, 309(5):16, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
16b.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-16b.pdf.

Matson:2013:SS

[1682] John Matson. Sayonara, supergiant. Scientific American, 309(5):18–19, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-18a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
18a.pdf.

Anonymous:2013:Nk

[1683] Anonymous. By the numbers. Scientific American, 309(5):18, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-18b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
18b.pdf.

Nuwer:2013:SCT

[1684] Rachel Nuwer. Steer clear, turtles. Scientific American, 309(5):20, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-20.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
20.pdf.

Gorman:2013:WBT

[1685] Christine Gorman. What’s better than BMI? Scientific American, 309(5):21, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
21.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-21.pdf.

Arnold:2013:MVL

[1686] Carrie Arnold. Mutating viruses lead cops to the criminals. Scientific American, 309(5):22–23, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n5/full/scientificamerican1113-
22a.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-22a.pdf.

Guterl:2013:PW

[1687] Sophie Guterl. Patent watch. Scientific American, 309(5):22, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-22b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
22b.pdf.

Anonymous:2013:Nl

[1688] Anonymous. By the numbers. Scientific American, 309(5):23, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-23a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
23a.pdf.

Levitan:2013:WSS

[1689] Dave Levitan. Wanted: Science sugar daddy. Scientific American, 309(5):23, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
23b.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-23b.pdf.

LeCompte:2013:TPP

[1690] Celeste LeCompte. Tattoos, ponytails and protozoans. Scientific American, 309(5):24, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
24.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-24.pdf.

Biello:2013:OFD

[1691] David Biello. Obama’s fracking dilemma. Scientific American, 309(5):25, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-25.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
25.pdf.

Jabr:2013:A

[1692] Ferris Jabr. A is for app. Scientific American, 309(5):26, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n5/full/scientificamerican1113-26.html;   http://www.nature.
com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
26.pdf.

Feltman:2013:WIa

[1693] Rachel Feltman. What is it? Scientific American, 309(5):27, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-27a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
27a.pdf.

Anonymous:2013:Nm

[1694] Anonymous. By the numbers. Scientific American, 309(5):27, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-27b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
27b.pdf.

Chakradhar:2013:WM

[1695] Shraddha Chakradhar. Woodland menace. Scientific American, 309(5):27, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
27c.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-27c.pdf.

Jabr:2013:SHG

[1696] Ferris Jabr. The science of health: Gambling on the brain. Scientific American, 309(5):28–30, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n5/full/scientificamerican1113-

28.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-28.pdf.

Pogue:2013:TCT

[1697] David Pogue. TechnoFiles: Crash text dummies. Scientific American, 309(5):32, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
32.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-32.pdf.

Anonymous:2013:RRV

[1698] Anonymous. Reviews and recommendations: Voyager. Scientific American, 309(5):78, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
78a.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-78a.pdf.

Anonymous:2013:OOF

[1699] Anonymous. Our once and future planet. Scientific American, 309(5):78, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-78b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
78b.pdf.

Anonymous:2013:SCM

[1700] Anonymous. Star-craving mad. Scientific American, 309(5):78, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-78c.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
78c.pdf.

Anonymous:2013:Ll

[1701] Anonymous. Out on a limb. Scientific American, 309(5):78, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-78d.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
78d.pdf.

Shermer:2013:SSI

[1702] Michael Shermer. Skeptic: Sovereign insanity. Scientific American, 309(5):79, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
79.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-79.pdf.

Mirsky:2013:AGSc

[1703] Steve Mirsky. Anti gravity: Saturnalia. Scientific American, 309(5):80, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n5/full/scientificamerican1113-80.html;   http:
//www.nature.com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
80.pdf.

Fischetti:2013:GSW

[1704] Mark Fischetti. Graphic science: Why so jittery? Scientific American, 309(5):82, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
82.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-82.pdf.

Werdelin:2013:EKB

[1705] Lars Werdelin. Evolution: King of beasts. Scientific American, 309(5):34–39, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
34.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-34.pdf.

Welsh:2013:AWT

[1706] William F. Welsh and Laurance R. Doyle. Astronomy: Worlds with two suns. Scientific American, 309(5):40–47, November 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n5/full/scientificamerican1113-40.html;   http://www.nature.
com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
40.pdf.

Jabr:2013:MWB

[1707] Ferris Jabr. Mind: Why the brain prefers paper. Scientific American, 309(5):48–53, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n5/full/scientificamerican1113-
48.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-48.pdf.

Wesson:2013:ASH

[1708] Kyle Wesson and Todd Humphreys. Air security: Hacking drones. Scientific American, 309(5):54–59, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n5/full/scientificamerican1113-
54.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-54.pdf.

Johnson:2013:MLT

[1709] George Johnson. Medicine: The long trail of cancer’s clues. Scientific American, 309(5):60–63, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n5/full/scientificamerican1113-
60.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-60.pdf.

Lanier:2013:SSH

[1710] Jaron Lanier. Science and society: How should we think about privacy? Scientific American, 309(5):64–71, November 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n5/full/scientificamerican1113-64.html;   http://www.nature.
com/scientificamerican/journal/v309/n5/pdf/scientificamerican1113-
64.pdf.

Bryant:2013:COS

[1711] Steven L. Bryant. Climate: The one-stop carbon solution. Scientific American, 309(5):72–77, November 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n5/full/scientificamerican1113-
72.html;   http://www.nature.com/scientificamerican/journal/
v309/n5/pdf/scientificamerican1113-72.pdf.

DiChristina:2013:EOA

[1712] Mariette DiChristina. From the editor: One atom at a time. Scientific American, 309(6):6, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
6.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-6.pdf.

Anonymous:2013:Ln

[1713] Anonymous. Letters. Scientific American, 309(6):8–9, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-8.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
8.pdf.

Schlenoff:2013:YAIk

[1714] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 309(6):86, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-86.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
86.pdf.

Anonymous:2013:SAT

[1715] Anonymous. Science agenda: Terrorism’s red gold. Scientific American, 309(6):10, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
10.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-10.pdf.

Levy:2013:FBE

[1716] Harold O. Levy. Forum: Brain exports. Scientific American, 309(6):13, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-13.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
13.pdf.

Stix:2013:ABL

[1717] Gary Stix. Advances: Bionic limbs, rewired. Scientific American, 309(6):14, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
14.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-14.pdf.

Kuchment:2013:LFK

[1718] Anna Kuchment. Looking for a flu killer. Scientific American, 309(6):16, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-16a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
16a.pdf.

Anonymous:2013:Nn

[1719] Anonymous. By the numbers. Scientific American, 309(6):16, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-16b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
16b.pdf.

Nuwer:2013:BYS

[1720] Rachel Nuwer. Bugs, yes. spiders, no. Scientific American, 309(6):17, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-17.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
17.pdf.

Braun:2013:BBB

[1721] Robyn Braun. Breaking bad, bacteria-style. Scientific American, 309(6):18, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
18.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-18.pdf.

Chiaet:2013:FP

[1722] Julianne Chiaet. FIFA physics. Scientific American, 309(6):19, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-19.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
19.pdf.

Collins:2013:EMW

[1723] Nathan Collins. Escape from the Milky Way. Scientific American, 309(6):20, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
20.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-20.pdf.

Goldman:2013:PN

[1724] Jason G. Goldman. Play nice. Scientific American, 309(6):21, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-21.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
21.pdf.

Cofield:2013:WNN

[1725] Calla Cofield. Waiting for a neutrino no-show. Scientific American, 309(6):22, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
22.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-22.pdf.

Platt:2013:RCC

[1726] John R. Platt. Rare California cypress is recovering under Federal protection. Scientific American, 309(6):23, December 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-23.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
23.pdf.

Moskowitz:2013:KSL

[1727] Clara Moskowitz. Kepler’s second life. Scientific American, 309(6):24, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-24.html;   http:

//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
24.pdf.

Arnold:2013:FL

[1728] Carrie Arnold. Follow the leader. Scientific American, 309(6):26, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-26a.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
26a.pdf.

Anonymous:2013:No

[1729] Anonymous. By the numbers. Scientific American, 309(6):26, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-26b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
26b.pdf.

Feltman:2013:USP

[1730] Rachel Feltman. Ugly science pays off. Scientific American, 309(6):27, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-27.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
27.pdf.

Feltman:2013:WIb

[1731] Rachel Feltman. What is it? Scientific American, 309(6):28, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-28.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
28.pdf.

Chen:2013:SHS

[1732] Ingfei Chen. The science of health: Scanning for Alzheimer’s. Scientific American, 309(6):30–31, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n6/full/scientificamerican1213-
30.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-30.pdf.

Pogue:2013:TPL

[1733] David Pogue. TechnoFiles: The problem with Lego phones. Scientific American, 309(6):32, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
32.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-32.pdf.

Anonymous:2013:RRT

[1734] Anonymous. Reviews and recommendations: Tigers forever. Scientific American, 309(6):80, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
80a.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-80a.pdf.

Anonymous:2013:SH

[1735] Anonymous. Shaping humanity. Scientific American, 309(6):80, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-80b.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
80b.pdf.

Anonymous:2013:BGb

[1736] Anonymous. Beautiful geometry. Scientific American, 309(6):80, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-80c.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
80c.pdf.

Shermer:2013:SGD

[1737] Michael Shermer. Skeptic: Is God dying? Scientific American, 309(6):82, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-82.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
82.pdf.

Mirsky:2013:AGP

[1738] Steve Mirsky. Anti gravity: Phoning it in. Scientific American, 309(6):84, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-84.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
84.pdf.

Jabr:2013:GSW

[1739] Ferris Jabr. Graphic science: Where the wild bees are. Scientific American, 309(6):88, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
88.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-88.pdf.

Anonymous:2013:WCI

[1740] Anonymous. World changing ideas. Scientific American, 309(6):34–35, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-34.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
34.pdf.

Ceder:2013:SD

[1741] Gerbrand Ceder and Kristin Persson. The stuff of dreams. Scientific American, 309(6):36–40, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n6/full/scientificamerican1213-
36.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-36.pdf.

Levitan:2013:TCP

[1742] Dave Levitan. Turning CO_{2}, pumped underground, into stone. Scientific American, 309(6):41–42, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v309/n6/full/scientificamerican1213-
41.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-41.pdf.

Fessenden:2013:LLC

[1743] Marissa Fessenden. Lego-like composites snap into place to make superstrong, affordable structures. Scientific American, 309(6):42–43, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-42.html;   http:

//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
42.pdf.

Greenemeier:2013:WOI

[1744] Larry Greenemeier. Worms and octopuses inspire pliable machines that go where no metallic robots can. Scientific American, 309(6):43–44, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-43.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
43.pdf.

Billings:2013:NSC

[1745] Lee Billings. Novel substances that can deflect light could lead to superfast Internet and razor-thin Smartphones. Scientific American, 309(6):44, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
44.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-44.pdf.

Courage:2013:MOG

[1746] Katherine Harmon Courage. Microbes in our gut could be manipulated to fight illnesses and infections. Scientific American, 309(6):45–46, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v309/n6/full/scientificamerican1213-45.html;   http:
//www.nature.com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
45.pdf.

Choi:2013:NTS

[1747] Charles Q. Choi. A new type of security chip guards against big data snooping. Scientific American, 309(6):46–47, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-46.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
46.pdf.

Yuhas:2013:TSL

[1748] Daisy Yuhas. A toolbox-size lab flags knockoff pills and potions. Scientific American, 309(6):47–48, December 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-47.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
47.pdf.

Choi:2013:FSS

[1749] Charles Q. Choi. Flexible, stretchable screens could put a rolled-up tablet in your pocket. Scientific American, 309(6):48, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-48a.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
48a.pdf.

Maron:2013:STA

[1750] Dina Fine Maron. Small tubes of antiseptic could prevent 500,000 infant deaths a year. Scientific American, 309(6):48–49, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-48b.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
48b.pdf.

Frazer:2013:IDF

[1751] Jennifer Frazer. Infectious disease: Fungi on the March. Scientific American, 309(6):50–57, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
50.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-50.pdf.

Wegner:2013:PHG

[1752] Daniel M. Wegner and Adrian F. Ward. Psychology: How Google is changing your brain. Scientific American, 309(6):58–61, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-58.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
58.pdf.

Tibbetts:2013:ABG

[1753] Elizabeth A. Tibbetts and Adrian G. Dyer. Animal behavior: Good with faces. Scientific American, 309(6):62–67, December 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-62.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
62.pdf.

Jayawardhana:2013:ACS

[1754] Ray Jayawardhana. Astrophysics: Coming soon: a supernova near you. Scientific American, 309(6):68–73, December 2013. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v309/
n6/full/scientificamerican1213-68.html;   http://www.nature.
com/scientificamerican/journal/v309/n6/pdf/scientificamerican1213-
68.pdf.

Carey:2013:EAS

[1755] John Carey. Environment: Architects of the swamp. Scientific American, 309(6):74–79, December 2013. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v309/n6/full/scientificamerican1213-
74.html;   http://www.nature.com/scientificamerican/journal/
v309/n6/pdf/scientificamerican1213-74.pdf.

DiChristina:2014:EBS

[1756] Mariette DiChristina. From the editor: Below the surface. Scientific American, 310(1):4, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
4.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-4.pdf.

Anonymous:2014:La

[1757] Anonymous. Letters. Scientific American, 310(1):6–7, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n1/full/scientificamerican0114-6.html;   http://www.nature.
com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
6.pdf.

Schlenoff:2014:YAIa

[1758] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 310(1):82, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-82.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
82.pdf.

Anonymous:2014:SABa

[1759] Anonymous. Science agenda: Beware the eye spies. Scientific American, 310(1):10, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
10.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-10.pdf.

Ali:2014:FMM

[1760] Saleem H. Ali. Forum: Magic metals. Scientific American, 310(1):12, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-12.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
12.pdf.

Moskowitz:2014:ACD

[1761] Clara Moskowitz. Advances: Cosmic dragnet. Scientific American, 310(1):13, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
13.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-13.pdf.

Fecht:2014:TH

[1762] Sarah Fecht. Taking the hit. Scientific American, 310(1):14, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-14.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
14.pdf.

Cowen:2014:DCC

[1763] Ron Cowen. Drinking from the cool cosmic stream. Scientific American, 310(1):17, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-

17a.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-17a.pdf.

Laursen:2014:RB

[1764] Lucas Laursen. The real Bigfoot. Scientific American, 310(1):17, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-17b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
17b.pdf.

Maron:2014:UOR

[1765] Dina Fine Maron. Under obamacare, a rollback of abortion coverage. Scientific American, 310(1):18, January 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n1/full/scientificamerican0114-18.html;   http://www.nature.
com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
18.pdf.

Goldman:2014:MD

[1766] Jason G. Goldman. Making dolfriends. Scientific American, 310(1):19, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-19a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
19a.pdf.

Anonymous:2014:Na

[1767] Anonymous. By the numbers. Scientific American, 310(1):19, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-19b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
19b.pdf.

Branswell:2014:BB

[1768] Helen Branswell. A botulism bind. Scientific American, 310(1):20, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-20a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
20a.pdf.

Anonymous:2014:Nb

[1769] Anonymous. By the numbers. Scientific American, 310(1):20, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-20b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
20b.pdf.

Biello:2014:PYP

[1770] David Biello. Pick your poison. Scientific American, 310(1):21, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-21.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
21.pdf.

Feltman:2014:EHE

[1771] Rachel Feltman. Over easy, hold the eggs. Scientific American, 310(1):22–23, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
22a.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-22a.pdf.

Wogan:2014:BDB

[1772] David Wogan. Big data, big energy. Scientific American, 310(1):22, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-22b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
22b.pdf.

Platt:2014:GDH

[1773] John R. Platt. Good dads help rare Haitian frogs thrive in captivity. Scientific American, 310(1):23, January 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n1/full/scientificamerican0114-23.html;   http://www.nature.
com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
23.pdf.

Madren:2014:ML

[1774] Carrie Madren. Mercury lockdown. Scientific American, 310(1):24, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-24a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
24a.pdf.

Moskowitz:2014:PES

[1775] Clara Moskowitz. Put up the Earth shield. Scientific American, 310(1): 24, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-24b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
24b.pdf.

Nuwer:2014:WI

[1776] Rachel Nuwer. What is it? Scientific American, 310(1):25, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n1/full/scientificamerican0114-25.html;   http://www.nature.
com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
25.pdf.

McKenna:2014:SHF

[1777] Maryn McKenna. The science of health: a flu vaccine That’s always in season. Scientific American, 310(1):26–27, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n1/full/scientificamerican0114-26.html;   http://www.nature.
com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
26.pdf.

Pogue:2014:TTW

[1778] David Pogue. TechnoFiles: In tech we don’t Trust. Scientific American, 310(1):28, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
28.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-28.pdf.

Anonymous:2014:RRLa

[1779] Anonymous. Reviews and recommendations: The last ocean. Scientific American, 310(1):78, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
78a.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-78a.pdf.

Anonymous:2014:RAC

[1780] Anonymous. Romania’s abandoned children. Scientific American, 310(1):78, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
78b.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-78b.pdf.

Anonymous:2014:O

[1781] Anonymous. Oxygen. Scientific American, 310(1):78, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n1/full/scientificamerican0114-78c.html;   http://www.nature.
com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
78c.pdf.

Anonymous:2014:W

[1782] Anonymous. Windfall. Scientific American, 310(1):78, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n1/full/scientificamerican0114-78d.html;   http://www.nature.
com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
78d.pdf.

Shermer:2014:SCS

[1783] Michael Shermer. Skeptic: Confessions of a speciesist. Scientific American, 310(1):79, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
79.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-79.pdf.

Mirsky:2014:AGP

[1784] Steve Mirsky. Anti gravity: Pervspicacity. Scientific American, 310(1): 80, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-80.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
80.pdf.

Matson:2014:GSW

[1785] John Matson. Graphic science: What it means to go viral. Scientific American, 310(1):84, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
84.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-84.pdf.

Bargh:2014:POU

[1786] John A. Bargh. Psychology: Our unconscious mind. Scientific American, 310(1):30–37, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
30.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-30.pdf.

Billings:2014:ASL

[1787] Lee Billings. Astronomy: The search for life on faraway moons. Scientific American, 310(1):38–43, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
38.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-38.pdf.

Covert:2014:BSL

[1788] Markus W. Covert. Bioengineering: Simulating a living Cell. Scientific American, 310(1):44–51, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-
44.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-44.pdf.

Smil:2014:ELS

[1789] Vaclav Smil. Energy: The long slow rise of solar and wind. Scientific American, 310(1):52–57, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n1/full/scientificamerican0114-

52.html;   http://www.nature.com/scientificamerican/journal/
v310/n1/pdf/scientificamerican0114-52.pdf.

Jabr:2014:BLU

[1790] Ferris Jabr. Biology: Life under the lens. Scientific American, 310(1):58– 63, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-58.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
58.pdf.

Berrah:2014:QPU

[1791] Nora Berrah and Philip H. Bucksbaum. Quantum physics: The ultimate X-ray machine. Scientific American, 310(1):64–71, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n1/full/scientificamerican0114-64.html;   http://www.nature.
com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
64.pdf.

Danielson:2014:HSC

[1792] Dennis Danielson and Christopher M. Graney. History of science: The case against Copernicus. Scientific American, 310(1):72–77, January 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n1/full/scientificamerican0114-72.html;   http:
//www.nature.com/scientificamerican/journal/v310/n1/pdf/scientificamerican0114-
72.pdf.

DiChristina:2014:EKQ

[1793] Mariette DiChristina. From the editor: Knowledge quest. Scientific American, 310(2):4, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
4.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-4.pdf.

Anonymous:2014:Lb

[1794] Anonymous. Letters. Scientific American, 310(2):6–7, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n2/full/scientificamerican0214-6.html;   http://www.nature.

com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
6.pdf.

Schlenoff:2014:YAIb

[1795] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 310(2):81, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-81.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
81.pdf.

Anonymous:2014:SAE

[1796] Anonymous. Science agenda: End the Drug War’s research bans. Scientific American, 310(2):10, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v310/n2/full/scientificamerican0214-
10.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-10.pdf.

Hamilton:2014:FWT

[1797] Roy H. Hamilton and Jihad Zreik. Forum: Wired for thought. Scientific American, 310(2):12, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
12.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-12.pdf.

Haelle:2014:ACC

[1798] Tara Haelle. Advances: Coughing up clues. Scientific American, 310(2):13, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
13.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-13.pdf.

Easter:2014:GMI

[1799] Michael Easter. Geologic miracle on ice. Scientific American, 310(2):14, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-14a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
14a.pdf.

Anonymous:2014:Nc

[1800] Anonymous. By the numbers. Scientific American, 310(2):14, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-14b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
14b.pdf.

Feltman:2014:WI

[1801] Rachel Feltman. What is it? Scientific American, 310(2):17, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-17a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
17a.pdf.

Fecht:2014:TE

[1802] Sarah Fecht. The traffic effect. Scientific American, 310(2):17, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-17b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
17b.pdf.

Wong:2014:TD

[1803] Kate Wong. Tusk to dust. Scientific American, 310(2):18, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n2/full/scientificamerican0214-18.html;   http://www.nature.
com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
18.pdf.

Moskowitz:2014:AR

[1804] Clara Moskowitz. Atomic revelation. Scientific American, 310(2):19, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-19.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
19.pdf.

Nuwer:2014:MM

[1805] Rachel Nuwer. Mushroom magic. Scientific American, 310(2):20, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-20.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
20.pdf.

Collins:2014:IWJ

[1806] Nathan Collins. Instant weirdness — just add water. Scientific American, 310(2):21, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
21.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-21.pdf.

Feltman:2014:DSA

[1807] Rachel Feltman. The DNA sketch artist. Scientific American, 310(2):22, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-22a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
22a.pdf.

Anonymous:2014:Nd

[1808] Anonymous. By the numbers. Scientific American, 310(2):22, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-22b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
22b.pdf.

Goldman:2014:LCM

[1809] Jason G. Goldman. Living claw to mouth. Scientific American, 310(2):23, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
23.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-23.pdf.

Wogan:2014:SD

[1810] David Wogan. Solar on demand. Scientific American, 310(2):24, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-24a.html;   http:

//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
24a.pdf.

Anonymous:2014:Ne

[1811] Anonymous. By the numbers. Scientific American, 310(2):24, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-24b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
24b.pdf.

Minnick:2014:PZG

[1812] Fred Minnick. 80 proof, zero gluten. Scientific American, 310(2):25, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-25.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
25.pdf.

Feltman:2014:MM

[1813] Rachel Feltman. Mixology micromachines. Scientific American, 310(2):26, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
26.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-26.pdf.

Bruzek:2014:GG

[1814] Alison Bruzek. Gadgets for gramps. Scientific American, 310(2):27, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-27.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
27.pdf.

Pogue:2014:TCC

[1815] David Pogue. TechnoFiles: The curse of the cloud. Scientific American, 310(2):28, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
28.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-28.pdf.

Khamsi:2014:SHT

[1816] Roxanne Khamsi. The science of health: The trouble with gluten. Scientific American, 310(2):30–32, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v310/n2/full/scientificamerican0214-
30.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-30.pdf.

Anonymous:2014:RRSa

[1817] Anonymous. Reviews and recommendations: The Sixth Extinction. Scientific American, 310(2):76, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v310/n2/full/scientificamerican0214-
76a.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-76a.pdf.

Anonymous:2014:PT

[1818] Anonymous. The perfect theory. Scientific American, 310(2):76, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-76b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
76b.pdf.

Anonymous:2014:EM

[1819] Anonymous. Extreme medicine. Scientific American, 310(2):76, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-76c.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
76c.pdf.

Anonymous:2014:MMW

[1820] Anonymous. Me, myself, and why. Scientific American, 310(2):76, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-76d.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
76d.pdf.

Shermer:2014:SHE

[1821] Michael Shermer. Skeptic: Heavens on Earth. Scientific American, 310(2):77, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
77.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-77.pdf.

Mirsky:2014:AGCa

[1822] Steve Mirsky. Anti gravity: Copy that. Scientific American, 310(2):78, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n2/full/scientificamerican0214-78.html;   http:
//www.nature.com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
78.pdf.

Christiansen:2014:GSH

[1823] Jen Christiansen and Mark Fischetti. Graphic science: How science evolves. Scientific American, 310(2):82, February 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n2/full/scientificamerican0214-82.html;   http://www.nature.
com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
82.pdf.

Bernauer:2014:PPR

[1824] Jan C. Bernauer and Randolf Pohl. Physics: The proton radius problem. Scientific American, 310(2):32–39, February 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n2/full/scientificamerican0214-32.html;   http://www.nature.
com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
32.pdf.

McGaugh:2014:BSR

[1825] James L. McGaugh and Aurora LePort. Brain science: Remembrance of all things past. Scientific American, 310(2):40–45, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n2/full/scientificamerican0214-40.html;   http://www.nature.
com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
40.pdf.

Jain:2014:MIW

[1826] Rakesh K. Jain. Medicine: An indirect way to tame cancer. Scientific American, 310(2):46–53, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
46.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-46.pdf.

Gershenfeld:2014:ITM

[1827] Alan Gershenfeld. Information technology: Mind games. Scientific American, 310(2):54–59, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
54.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-54.pdf.

Smith:2014:ABB

[1828] Carolynn “K lynn” L. Smith and Sarah L. Zielinski. Animal behavior: Brainy bird. Scientific American, 310(2):60–65, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n2/full/scientificamerican0214-60.html;   http://www.nature.
com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
60.pdf.

Ribe:2014:FDL

[1829] Neil M. Ribe, Mehdi Habibi, and Daniel Bonn. Fluid dynamics: The liquid rope trick. Scientific American, 310(2):66–71, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n2/full/scientificamerican0214-66.html;   http://www.nature.
com/scientificamerican/journal/v310/n2/pdf/scientificamerican0214-
66.pdf.

Hand:2014:MNS

[1830] David J. Hand. Mathematics: Never say never. Scientific American, 310(2):72–75, February 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n2/full/scientificamerican0214-
72.html;   http://www.nature.com/scientificamerican/journal/
v310/n2/pdf/scientificamerican0214-72.pdf.

DiChristina:2014:EBB

[1831] Mariette DiChristina. From the editor: The brain beckons. Scientific American, 310(3):4, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
4.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-4.pdf.

Anonymous:2014:Lc

[1832] Anonymous. Letters. Scientific American, 310(3):6–8, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n3/full/scientificamerican0314-6.html;   http://www.nature.
com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
6.pdf.

Schlenoff:2014:YAIc

[1833] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 310(3):86, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-86.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
86.pdf.

Anonymous:2014:SAFa

[1834] Anonymous. Science agenda: Free Willy — and all his pals. Scientific American, 310(3):10, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
10.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-10.pdf.

Willett:2014:FCB

[1835] Walter Willett. Forum: The case for banning trans fats. Scientific American, 310(3):13, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
13.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-13.pdf.

Weiner:2014:ABF

[1836] Debra Weiner. Advances: The Brain’s fix-it brigade. Scientific American, 310(3):16, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-

16.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-16.pdf.

Greenspan:2014:RFA

[1837] Jesse Greenspan. The rain forest of Alabama. Scientific American, 310(3):18, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
18a.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-18a.pdf.

Maron:2014:F

[1838] Dina Fine Maron. Flame out. Scientific American, 310(3):18, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-18b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
18b.pdf.

Wong:2014:AB

[1839] Kate Wong. Ancient burial. Scientific American, 310(3):19, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n3/full/scientificamerican0314-19a.html;   http://www.nature.
com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
19a.pdf.

Anonymous:2014:Nf

[1840] Anonymous. By the numbers. Scientific American, 310(3):19, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-19b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
19b.pdf.

Moskowitz:2014:LT

[1841] Clara Moskowitz. Leaden treasure. Scientific American, 310(3):20, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-20a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
20a.pdf.

Anonymous:2014:Ng

[1842] Anonymous. By the numbers. Scientific American, 310(3):20, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-20b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
20b.pdf.

Greenemeier:2014:SC

[1843] Larry Greenemeier. Sun chaser. Scientific American, 310(3):23, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-23a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
23a.pdf.

Sneed:2014:WIa

[1844] Annie Sneed. What is it? Scientific American, 310(3):23, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n3/full/scientificamerican0314-23b.html;   http://www.nature.
com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
23b.pdf.

Maron:2014:TK

[1845] Dina Fine Maron. Trouble knocking. Scientific American, 310(3):24, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-24.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
24.pdf.

Guenard:2014:GHD

[1846] Rebecca Guenard. Growing hair in a dish. Scientific American, 310(3): 25, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-25.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
25.pdf.

Smith:2014:NFZ

[1847] Peter Andrey Smith. No-fly zone. Scientific American, 310(3):26, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-26.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
26.pdf.

Matson:2014:LDP

[1848] John Matson. Lonely dwarf planets. Scientific American, 310(3):27, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-27.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
27.pdf.

Goldman:2014:TCa

[1849] Jason G. Goldman. Thought control. Scientific American, 310(3):28, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-28.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
28.pdf.

Feltman:2014:RR

[1850] Rachel Feltman. Rapid response. Scientific American, 310(3):29, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-29.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
29.pdf.

Biello:2014:SNT

[1851] David Biello. The storm next time. Scientific American, 310(3):30, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-30.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
30.pdf.

Courage:2014:SHS

[1852] Katherine Harmon Courage. The science of health: Space: The final medical frontier. Scientific American, 310(3):32–34, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n3/full/scientificamerican0314-32.html;   http://www.nature.

com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
32.pdf.

Pogue:2014:TFI

[1853] David Pogue. TechnoFiles: Future imperfect. Scientific American, 310(3):36, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
36.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-36.pdf.

Anonymous:2014:RRC

[1854] Anonymous. Reviews and recommendations: Caffeinated. Scientific American, 310(3):80, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
80a.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-80a.pdf.

Anonymous:2014:CQC

[1855] Anonymous. Computing with quantum cats. Scientific American, 310(3): 80, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-80b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
80b.pdf.

Moskowitz:2014:PF

[1856] Clara Moskowitz. Particle fever. Scientific American, 310(3):80, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-80c.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
80c.pdf.

Anonymous:2014:ELS

[1857] Anonymous. The extreme life of the sea. Scientific American, 310(3):80, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-80d.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
80d.pdf.

Anonymous:2014:CF

[1858] Anonymous. On the cancer frontier. Scientific American, 310(3):80, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-80e.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
80e.pdf.

Shermer:2014:SAD

[1859] Michael Shermer. Skeptic: The awe delusion. Scientific American, 310(3):82, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
82.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-82.pdf.

Mirsky:2014:AG

[1860] Steve Mirsky. Anti gravity: 007 and 7. Scientific American, 310(3):84, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n3/full/scientificamerican0314-84.html;   http:
//www.nature.com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
84.pdf.

Matson:2014:GSF

[1861] John Matson. Graphic science: Filth in the spice rack. Scientific American, 310(3):88, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
88.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-88.pdf.

Yuste:2014:NNC

[1862] Rafael Yuste and George M. Church. Neuroscience: The new century of the brain. Scientific American, 310(3):38–45, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n3/full/scientificamerican0314-38.html;   http://www.nature.
com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
38.pdf.

Libeskind:2014:ADG

[1863] Noam I. Libeskind. Astronomy: Dwarf galaxies and the Dark Web. Scientific American, 310(3):46–51, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v310/n3/full/scientificamerican0314-
46.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-46.pdf.

Lewis:2014:MGT

[1864] Ricki Lewis. Medicine: Gene therapy’s second act. Scientific American, 310(3):52–57, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
52.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-52.pdf.

Zimmer:2014:ESO

[1865] Carl Zimmer. Earth science: The oldest rocks on Earth. Scientific American, 310(3):58–63, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n3/full/scientificamerican0314-
58.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-58.pdf.

Garner:2014:ITC

[1866] Harold “Skip” Garner. Information technology: The case of the stolen words. Scientific American, 310(3):64–67, March 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n3/full/scientificamerican0314-64.html;   http://www.nature.
com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
64.pdf.

Powell:2014:EAC

[1867] William Powell. Ecology: The American chestnut’s genetic rebirth. Scientific American, 310(3):68–73, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v310/n3/full/scientificamerican0314-
68.html;   http://www.nature.com/scientificamerican/journal/
v310/n3/pdf/scientificamerican0314-68.pdf.

Bilalic:2014:PWG

[1868] Merim Bilalić and Peter McLeod. Psychology: Why good thoughts block better ones. Scientific American, 310(3):74–79, March 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n3/full/scientificamerican0314-74.html;   http://www.nature.
com/scientificamerican/journal/v310/n3/pdf/scientificamerican0314-
74.pdf.

DiChristina:2014:ED

[1869] Mariette DiChristina. From the Editor: Out of darkness. Scientific American, 310(4):6, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
6.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-6.pdf.

Anonymous:2014:Ld

[1870] Anonymous. Letters. Scientific American, 310(4):8–10, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-8.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
8.pdf.

Schlenoff:2014:YAId

[1871] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 310(4):90–91, April 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-90.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
90.pdf.

Anonymous:2014:SAPa

[1872] Anonymous. Science agenda: Preserve the Endangered Species Act. Scientific American, 310(4):12, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
12.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-12.pdf.

Crawford:2014:FBD

[1873] Kate Crawford. Forum: Big Data stalking. Scientific American, 310(4): 14, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-14.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
14.pdf.

Moyer:2014:ANB

[1874] Michael Moyer. Advances: The new black hole battle. Scientific American, 310(4):16, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
16.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-16.pdf.

Moyer:2014:CC

[1875] Melinda Wenner Moyer. Cancer culprit? Scientific American, 310(4):18, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-18.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
18.pdf.

Nuwer:2014:CH

[1876] Rachel Nuwer. Cockroach homecoming. Scientific American, 310(4):21, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-21a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
21a.pdf.

Sneed:2014:WIb

[1877] Annie Sneed. What is it? Scientific American, 310(4):21, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-21b.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
21b.pdf.

Greenemeier:2014:WFI

[1878] Larry Greenemeier. War funding inside? Scientific American, 310(4):22, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-22a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
22a.pdf.

Anonymous:2014:Nh

[1879] Anonymous. By the numbers. Scientific American, 310(4):22, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-22b.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
22b.pdf.

Nuwer:2014:FD

[1880] Rachel Nuwer. Fuzzy discovery. Scientific American, 310(4):23, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-23.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
23.pdf.

Rogers:2014:LHE

[1881] Nala Rogers. Love him or eat him? Scientific American, 310(4):24, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-24.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
24.pdf.

Goldman:2014:WFB

[1882] Jason G. Goldman. Where farmers and birds agree. Scientific American, 310(4):26, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
26.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-26.pdf.

Giller:2014:STT

[1883] Geoffrey Giller. Solar times two. Scientific American, 310(4):27, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-27.html;   http:

//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
27.pdf.

Hsu:2014:WCF

[1884] Jeremy Hsu. When cars fly. Scientific American, 310(4):28, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-28a.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
28a.pdf.

Anonymous:2014:Ni

[1885] Anonymous. By the numbers. Scientific American, 310(4):28, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-28b.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
28b.pdf.

Choi:2014:OWS

[1886] Charles Q. Choi. One-way street for sound. Scientific American, 310(4): 31, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-31a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
31a.pdf.

Anonymous:2014:Nj

[1887] Anonymous. By the numbers. Scientific American, 310(4):31, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-31b.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
31b.pdf.

Maron:2014:AO

[1888] Dina Fine Maron. Antibiotic overkill. Scientific American, 310(4):32, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-32.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
32.pdf.

Greenwood:2014:ELH

[1889] Veronique Greenwood. Evolution’s little helper. Scientific American, 310(4):33, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
33.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-33.pdf.

Storrs:2014:SHH

[1890] Carina Storrs. The science of health: Hidden dangers of going under. Scientific American, 310(4):34–35, April 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-34.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
34.pdf.

Pogue:2014:TGN

[1891] David Pogue. TechnoFiles: The great net debate. Scientific American, 310(4):36, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
36.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-36.pdf.

Anonymous:2014:RRY

[1892] Anonymous. Reviews and recommendations: You are here. Scientific American, 310(4):86, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
86a.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-86a.pdf.

Anonymous:2014:AR

[1893] Anonymous. The age of radiance. Scientific American, 310(4):86, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-86b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
86b.pdf.

Anonymous:2014:OLT

[1894] Anonymous. The oldest living things in the world. Scientific American, 310(4):86, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
86c.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-86c.pdf.

Anonymous:2014:LU

[1895] Anonymous. Life unfolding. Scientific American, 310(4):86, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-86d.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
86d.pdf.

Shermer:2014:SSL

[1896] Michael Shermer. Skeptic: The science of lying. Scientific American, 310(4):87, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
87.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-87.pdf.

Mirsky:2014:AGN

[1897] Steve Mirsky. Anti gravity: Numbers game. Scientific American, 310(4): 88, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-88.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
88.pdf.

Fischetti:2014:GSC

[1898] Mark Fischetti. Graphic science: On climate, the people agree. Scientific American, 310(4):92, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
92.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-92.pdf.

Lemonick:2014:AFS

[1899] Michael D. Lemonick. Astronomy: The first starlight. Scientific American, 310(4):38–45, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
38.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-38.pdf.

Wong:2014:ERH

[1900] Kate Wong. Evolution: Rise of the human predator. Scientific American, 310(4):46–51, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
46.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-46.pdf.

Gorman:2014:FMR

[1901] Christine Gorman and Dina Fine Maron. The future of medicine 2014: The RNA revolution. Scientific American, 310(4):52–59, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-52.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
52.pdf.

Gorman:2014:NSA

[1902] Christine Gorman. A new shot against Hepatitis C. Scientific American, 310(4):56–57, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
56.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-56.pdf.

Jabr:2014:DNT

[1903] Ferris Jabr. Defeating nature’s terrorists. Scientific American, 310(4): 58, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n4/full/scientificamerican0414-58.html;   http:
//www.nature.com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
58.pdf.

Schrope:2014:MEJ

[1904] Mark Schrope. Marine exploration: Journey to the bottom of the sea. Scientific American, 310(4):60–69, April 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-60.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
60.pdf.

Lein:2014:NGG

[1905] Ed Lein and Mike Hawrylycz. Neuroscience: The genetic geography of the brain. Scientific American, 310(4):70–77, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-70.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
70.pdf.

Mann:2014:CCF

[1906] Michael E. Mann. Climate change: False hope. Scientific American, 310(4):78–81, April 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n4/full/scientificamerican0414-
78.html;   http://www.nature.com/scientificamerican/journal/
v310/n4/pdf/scientificamerican0414-78.pdf.

Alexander:2014:MSS

[1907] Amir Alexander. Mathematics: The secret spiritual history of calculus. Scientific American, 310(4):82–85, April 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n4/full/scientificamerican0414-82.html;   http://www.nature.
com/scientificamerican/journal/v310/n4/pdf/scientificamerican0414-
82.pdf.

DiChristina:2014:EDP

[1908] Mariette DiChristina. From the Editor: Does physics have a problem? Scientific American, 310(5):6, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
6.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-6.pdf.

Anonymous:2014:Le

[1909] Anonymous. Letters. Scientific American, 310(5):8–9, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-8.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
8.pdf.

Schlenoff:2014:YAIe

[1910] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 310(5):82, May 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-82.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
82.pdf.

Anonymous:2014:SAM

[1911] Anonymous. Science agenda: The myth of the compassionate execution. Scientific American, 310(5):12, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v310/n5/full/scientificamerican0514-
12.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-12.pdf.

Kiermer:2014:FEO

[1912] Veronique Kiermer. Forum: Eureka once, Eureka twice. Scientific American, 310(5):13, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
13.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-13.pdf.

Moskowitz:2014:AOI

[1913] Clara Moskowitz. Advances: Our inflated universe. Scientific American, 310(5):14, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
14.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-14.pdf.

Biello:2014:LVC

[1914] David Biello. The little volcanoes that could. Scientific American, 310 (5):16, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-16.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
16.pdf.

Platt:2014:SRM

[1915] John R. Platt. A symbol of royalty makes a tentative comeback. Scientific American, 310(5):17, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
17.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-17.pdf.

Wong:2014:AF

[1916] Kate Wong. Ancient footprint. Scientific American, 310(5):18, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-18a.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
18a.pdf.

Nuwer:2014:PLL

[1917] Rachel Nuwer. The physics of long, loose tresses. Scientific American, 310 (5):18, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-18b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
18b.pdf.

Anonymous:2014:Nk

[1918] Anonymous. By the numbers. Scientific American, 310(5):20, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-20a.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
20a.pdf.

Moskowitz:2014:KA

[1919] Clara Moskowitz. Kepler’s afterlife. Scientific American, 310(5):20, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-20b.html;   http:

//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
20b.pdf.

Sneed:2014:WIc

[1920] Annie Sneed. What is it? Scientific American, 310(5):21, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-21.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
21.pdf.

Sneed:2014:PP

[1921] Annie Sneed. A pontoon of pupae. Scientific American, 310(5):22, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-22.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
22.pdf.

Skwarecki:2014:ITP

[1922] Beth Skwarecki. It takes a prion to remember. Scientific American, 310 (5):23, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-23.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
23.pdf.

Goldman:2014:WHF

[1923] Jason G. Goldman. He who hollers fastest gets the girl. Scientific American, 310(5):26, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
26a.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-26a.pdf.

Anonymous:2014:Nl

[1924] Anonymous. By the numbers. Scientific American, 310(5):26, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-26b.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
26b.pdf.

Grossman:2014:C

[1925] Wendy M. Grossman. Cameraprints. Scientific American, 310(5):27, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-27.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
27.pdf.

Giller:2014:USD

[1926] Geoffrey Giller. Using a Smartphone to detect cancer. Scientific American, 310(5):28, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
28a.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-28a.pdf.

Stix:2014:MPS

[1927] Gary Stix. “millions of people are suffering”. Scientific American, 310(5): 28, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-28b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
28b.pdf.

Sneed:2014:PIE

[1928] Annie Sneed. Pot ingredient for epilepsy. Scientific American, 310(5): 29, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-29.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
29.pdf.

Maron:2014:SHC

[1929] Dina Fine Maron. The science of health: Are e-cigarettes safe? Scientific American, 310(5):31–32, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
31.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-31.pdf.

Pogue:2014:TFW

[1930] David Pogue. TechnoFiles: Fear the worst. Scientific American, 310(5): 33, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-33.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
33.pdf.

Anonymous:2014:RRA

[1931] Anonymous. Reviews and recommendations: Animal architecture. Scientific American, 310(5):76, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
76a.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-76a.pdf.

Anonymous:2014:SM

[1932] Anonymous. Stuff matters. Scientific American, 310(5):76, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-76b.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
76b.pdf.

Anonymous:2014:NNR

[1933] Anonymous. Nature’s nether regions. Scientific American, 310(5):76, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-76c.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
76c.pdf.

Anonymous:2014:Pa

[1934] Anonymous. Proof. Scientific American, 310(5):76, May 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-76d.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
76d.pdf.

Anonymous:2014:TDN

[1935] Anonymous. The tale of the dueling neurosurgeons. Scientific American, 310(5):76, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-

76e.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-76e.pdf.

Shermer:2014:SGJ

[1936] Michael Shermer. Skeptic: The genesis of justice. Scientific American, 310(5):78, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
78.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-78.pdf.

Mirsky:2014:AGS

[1937] Steve Mirsky. Anti gravity: Statistical significance. Scientific American, 310(5):80, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
80.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-80.pdf.

Fischetti:2014:GSS

[1938] Mark Fischetti. Graphic science: Spring fever. Scientific American, 310 (5):84, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-84.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
84.pdf.

Lykken:2014:PPS

[1939] Joseph Lykken and Maria Spiropulu. Particle physics: Supersymmetry and the crisis in physics. Scientific American, 310(5):34–39, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-34.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
34.pdf.

Wolchok:2014:MCS

[1940] Jedd D. Wolchok. Medicine: Cancer’s off switch. Scientific American, 310(5):40–45, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
40.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-40.pdf.

Anemone:2014:PFG

[1941] Robert L. Anemone and Charles W. Emerson. Paleontology: Fossil GPS. Scientific American, 310(5):46–51, May 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n5/full/scientificamerican0514-46.html;   http://www.nature.
com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
46.pdf.

Owen:2014:NAT

[1942] Adrian M. Owen. Neuroscience: Is anybody in there? Scientific American, 310(5):52–57, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
52.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-52.pdf.

Kota:2014:ESS

[1943] Sridhar Kota. Engineering: Shape-shifting things to come. Scientific American, 310(5):58–65, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
58.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-58.pdf.

McCalman:2014:OGC

[1944] Iain McCalman. Oceanography: The great coral grief. Scientific American, 310(5):66–69, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n5/full/scientificamerican0514-
66.html;   http://www.nature.com/scientificamerican/journal/
v310/n5/pdf/scientificamerican0514-66.pdf.

Bleicher:2014:MO

[1945] Ariel Bleicher. Mathematics: The oracle. Scientific American, 310(5): 70–75, May 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n5/full/scientificamerican0514-70.html;   http:
//www.nature.com/scientificamerican/journal/v310/n5/pdf/scientificamerican0514-
70.pdf.

DiChristina:2014:ECH

[1946] Mariette DiChristina. From the Editor: Creatures of habit. Scientific American, 310(6):4, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
4.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-4.pdf.

Anonymous:2014:Lf

[1947] Anonymous. Letters. Scientific American, 310(6):6, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-6.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
6.pdf.

Schlenoff:2014:YAIf

[1948] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 310(6):84, June 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-84.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
84.pdf.

Anonymous:2014:SAFb

[1949] Anonymous. Science agenda: Full disclosure. Scientific American, 310(6):10, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
10.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-10.pdf.

Schweitzer:2014:FDI

[1950] Mary H. Schweitzer. Forum: Dinosaurs are important. Scientific American, 310(6):12, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
12.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-12.pdf.

Fischetti:2014:ATP

[1951] Mark Fischetti. Advances: Trees that pollute. Scientific American, 310(6):14, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
14.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-14.pdf.

Haelle:2014:FBS

[1952] Tara Haelle. The face behind the skull. Scientific American, 310(6):15, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-15.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
15.pdf.

Moyer:2014:TOC

[1953] Melinda Wenner Moyer. Twitter opens its cage. Scientific American, 310(6):16, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
16.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-16.pdf.

Sneed:2014:WId

[1954] Annie Sneed. What is it? Scientific American, 310(6):19, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-19a.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
19a.pdf.

Anonymous:2014:Nm

[1955] Anonymous. By the numbers. Scientific American, 310(6):19, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-19b.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
19b.pdf.

Moskowitz:2014:ST

[1956] Clara Moskowitz. Surf’s up on Titan. Scientific American, 310(6):20, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-20.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
20.pdf.

Biba:2014:SG

[1957] Erin Biba. Sugar gut. Scientific American, 310(6):21, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-21.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
21.pdf.

Nuwer:2014:PT

[1958] Rachel Nuwer. Particle tweezers. Scientific American, 310(6):22, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-22a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
22a.pdf.

Anonymous:2014:Nn

[1959] Anonymous. By the numbers. Scientific American, 310(6):22, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-22b.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
22b.pdf.

Richards:2014:MP

[1960] Dana Richards. Meeting of the puzzlers. Scientific American, 310(6):23, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-23.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
23.pdf.

Emspak:2014:DM

[1961] Jesse Emspak. Data mask. Scientific American, 310(6):24, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-24a.html;   http://www.nature.

com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
24a.pdf.

Skwarecki:2014:TBF

[1962] Beth Skwarecki. Turtle baby’s first steps. Scientific American, 310(6): 24, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-24b.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
24b.pdf.

Biello:2014:DY

[1963] David Biello. DIY yeast. Scientific American, 310(6):25, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-25.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
25.pdf.

Goldman:2014:BR

[1964] Jason G. Goldman. Bee resourceful. Scientific American, 310(6):26, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-26a.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
26a.pdf.

Anonymous:2014:No

[1965] Anonymous. By the numbers. Scientific American, 310(6):26, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-26b.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
26b.pdf.

Biello:2014:MLW

[1966] David Biello. A milestone on the long and winding road to fusion. Scientific American, 310(6):27, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
27.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-27.pdf.

Wallis:2014:SHG

[1967] Claudia Wallis. The science of health: Gut reactions. Scientific American, 310(6):30–33, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
30.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-30.pdf.

Pogue:2014:TSW

[1968] David Pogue. TechnoFiles: Smart watches flunk out. Scientific American, 310(6):37, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
37.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-37.pdf.

Anonymous:2014:RRSb

[1969] Anonymous. Reviews and recommendations: Sally ride: America’s first woman in space. Scientific American, 310(6):80, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-80a.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
80a.pdf.

Anonymous:2014:VUJ

[1970] Anonymous. Virtual unreality: Just because the Internet told you, how do you know it’s true? Scientific American, 310(6):80, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-80b.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
80b.pdf.

Anonymous:2014:DFR

[1971] Anonymous. Deep: Freediving, renegade science, and what the ocean tells us about ourselves. Scientific American, 310(6):80, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-80c.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
80c.pdf.

Anonymous:2014:BCW

[1972] Anonymous. Brian Cox’s wonders of life. Scientific American, 310(6): 80, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-80d.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
80d.pdf.

Shermer:2014:SNN

[1973] Michael Shermer. Skeptic: Nuclear nada. Scientific American, 310(6): 81, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-81.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
81.pdf.

Mirsky:2014:AGB

[1974] Steve Mirsky. Anti gravity: Being driven up the wall. Scientific American, 310(6):82, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
82.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-82.pdf.

Fischetti:2014:GSWa

[1975] Mark Fischetti. Graphic science: The not so wretched masses. Scientific American, 310(6):86, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
86.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-86.pdf.

Graybiel:2014:BSG

[1976] Ann M. Graybiel and Kyle S. Smith. Behavioral science: Good habits, bad habits. Scientific American, 310(6):38–43, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-38.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
38.pdf.

McKay:2014:PSH

[1977] Christopher P. McKay and Victor Parro García. Planetary science: How to search for life on Mars. Scientific American, 310(6):44–49, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-44.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
44.pdf.

Ecker:2014:EGC

[1978] David J. Ecker. Epidemiology: Germ catcher. Scientific American, 310(6):50–55, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
50.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-50.pdf.

Baum:2014:MSR

[1979] Dan Baum. Meteorology: Summon the rain. Scientific American, 310(6):56–63, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
56.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-56.pdf.

Borrell:2014:MSC

[1980] Brendan Borrell. Medicine: Seeds of a cure. Scientific American, 310(6):64–69, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v310/n6/full/scientificamerican0614-
64.html;   http://www.nature.com/scientificamerican/journal/
v310/n6/pdf/scientificamerican0614-64.pdf.

Basu:2014:EP

[1981] Kaushik Basu. Economics: The ponz. Scientific American, 310(6):70– 75, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v310/n6/full/scientificamerican0614-70.html;   http:
//www.nature.com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
70.pdf.

Shariff:2014:CSW

[1982] Azim F. Shariff and Kathleen D. Vohs. Cognitive science: The world without free will. Scientific American, 310(6):76–79, June 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v310/
n6/full/scientificamerican0614-76.html;   http://www.nature.
com/scientificamerican/journal/v310/n6/pdf/scientificamerican0614-
76.pdf.

DiChristina:2014:ECW

[1983] Mariette DiChristina. From the Editor: A connected world. Scientific American, 311(1):6, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
6.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-6.pdf.

Anonymous:2014:Lg

[1984] Anonymous. Letters. Scientific American, 311(1):8–9, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-8.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
8.pdf.

Schlenoff:2014:YAIg

[1985] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 311(1):90, July 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-90.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
90.pdf.

Schlenoff:2014:CC

[1986] Daniel C. Schlenoff. Centennial of a calamity. Scientific American, 311(1):92–93, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
92.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-92.pdf.

Anonymous:2014:SAPb

[1987] Anonymous. Science agenda: Poisoned well. Scientific American, 311(1): 12, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-12.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
12.pdf.

Felten:2014:FHW

[1988] Edward W. Felten and Joshua A. Kroll. Forum: Help wanted on Internet security. Scientific American, 311(1):14, July 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-14.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
14.pdf.

Arnold:2014:AMM

[1989] Carrie Arnold. Advances: Microscopic mass murderer. Scientific American, 311(1):16–17, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
16.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-16.pdf.

Nuwer:2014:KJ

[1990] Rachel Nuwer. Know the jargon. Scientific American, 311(1):18, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-18a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
18a.pdf.

Anonymous:2014:FAL

[1991] Anonymous. Fastest animal on land? A mite. Scientific American, 311 (1):18, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-18b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
18b.pdf.

Bourzac:2014:GDS

[1992] Katherine Bourzac. Graphene’s dark side. Scientific American, 311(1): 19, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-19a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
19a.pdf.

Giller:2014:HS

[1993] Geoffrey Giller. Hold still. Scientific American, 311(1):19, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-19b.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
19b.pdf.

Goldman:2014:CB

[1994] Jason G. Goldman. Census for the birds. Scientific American, 311(1): 20, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-20a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
20a.pdf.

Anonymous:2014:Np

[1995] Anonymous. By the numbers. Scientific American, 311(1):20, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-20b.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
20b.pdf.

Croswell:2014:MB

[1996] Ken Croswell. A moonlet is Born. Scientific American, 311(1):21, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-21.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
21.pdf.

Marshall:2014:TK

[1997] Lisa Marshall. Two of a kind. Scientific American, 311(1):22, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-22.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
22.pdf.

Jabr:2014:BPA

[1998] Ferris Jabr. The brain’s power to avoid diversions. Scientific American, 311(1):23, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
23.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-23.pdf.

Sneed:2014:HCE

[1999] Annie Sneed. How to curb an epidemic. Scientific American, 311(1): 24, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-24.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
24.pdf.

Moskowitz:2014:GU

[2000] Clara Moskowitz. A glimpse at the unseen. Scientific American, 311(1): 25, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-25.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
25.pdf.

Moskowitz:2014:WI

[2001] Clara Moskowitz. What is it? Scientific American, 311(1):26, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-26a.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
26a.pdf.

Bennington-Castro:2014:CMY

[2002] Joseph Bennington-Castro. The circuit made for your arm. Scientific American, 311(1):26, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-

26b.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-26b.pdf.

Fischman:2014:HS

[2003] Josh Fischman. Here to stay. Scientific American, 311(1):27, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-27.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
27.pdf.

Szalavitz:2014:SHW

[2004] Maia Szalavitz. The science of health: When does self-help actually help? Scientific American, 311(1):30–31, July 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-30.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
30.pdf.

Pogue:2014:TSS

[2005] David Pogue. TechnoFiles: Smart sharing. Scientific American, 311(1): 33, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-33.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
33.pdf.

Anonymous:2014:RRSc

[2006] Anonymous. Reviews and recommendations: Starlight detectives: How astronomers, inventors, and eccentrics discovered the modern universe. Scientific American, 311(1):84, July 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-84a.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
84a.pdf.

Anonymous:2014:CAD

[2007] Anonymous. The chemistry of alchemy: From dragon’s blood to donkey dung, how chemistry was forged. Scientific American, 311(1):84, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-84b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
84b.pdf.

Anonymous:2014:EHE

[2008] Anonymous. The end: The human experience of death. Scientific American, 311(1):84, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
84c.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-84c.pdf.

Anonymous:2014:ZHE

[2009] Anonymous. Zoom: How everything moves, from atoms and galaxies to blizzards and bees. Scientific American, 311(1):84, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-84d.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
84d.pdf.

Shermer:2014:SMI

[2010] Michael Shermer. Skeptic: The myth of income inequality. Scientific American, 311(1):86, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
86.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-86.pdf.

Mirsky:2014:AGW

[2011] Steve Mirsky. Anti gravity: Wrong number. Scientific American, 311(1): 88, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-88.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
88.pdf.

Fischetti:2014:GSP

[2012] Mark Fischetti. Graphic science: Patent crossroads. Scientific American, 311(1):96, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-

96.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-96.pdf.

Brody:2014:OCM

[2013] Herb Brody. Outlook: Cancer: The March on malignancy. Scientific American, 311(1):103, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S3.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S3.pdf.

May:2014:AE

[2014] Mike May. Attacking an epidemic. Scientific American, 311(1):104– 105, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-S4.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
S4.pdf.

Gravitz:2014:TTP

[2015] Lauren Gravitz. Therapy: This time it’s personal. Scientific American, 311(1):106–108, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S6.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S6.pdf.

Eisenstein:2014:CTM

[2016] Michael Eisenstein. Clinical trials: More trials, fewer tribulations. Scientific American, 311(1):109–111, July 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-S9.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
S9.pdf.

Wright:2014:NDP

[2017] Jessica Wright. Nanotechnology: Deliver on a promise. Scientific American, 311(1):112–113, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S12.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S12.pdf.

Deweerdt:2014:CBN

[2018] Sarah Deweerdt. Comparative biology: Naked ambition. Scientific American, 311(1):114–115, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S14.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S14.pdf.

Kessler:2014:PAD

[2019] Rebecca Kessler. Prevention: Air of danger. Scientific American, 311(1):116–117, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S16.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S16.pdf.

Bender:2014:DWG

[2020] Eric Bender. Developing world: Global warning. Scientific American, 311(1):118–119, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S18.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S18.pdf.

Savage:2014:BBD

[2021] Neil Savage. Bioinformatics: Big Data versus the Big C. Scientific American, 311(1):120–121, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S20.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S20.pdf.

Quackenbush:2014:PLS

[2022] John Quackenbush. Perspective: Learning to share. Scientific American, 311(1):122, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S22.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S22.pdf.

Bourzac:2014:BTK

[2023] Katherine Bourzac. Biology: Three known unknowns. Scientific American, 311(1):123–125, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
S23.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-S23.pdf.

Dublon:2014:CSE

[2024] Gershon Dublon and Joseph A. Paradiso. Computer science: Extra sensory perception. Scientific American, 311(1):36–41, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-36.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
36.pdf.

Finkbeiner:2014:AGB

[2025] Douglas Finkbeiner, Meng Su, and Dmitry Malyshev. Astrophysics: Giant bubbles of the Milky Way. Scientific American, 311(1):42–47, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-42.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
42.pdf.

Vance:2014:AGB

[2026] Erik Vance. Archaeology: Gods of blood & stone. Scientific American, 311(1):48–55, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
48.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-48.pdf.

Jabr:2014:FTB

[2027] Ferris Jabr. Food technology: Building tastier fruits & veggies (no GMOs required). Scientific American, 311(1):56–61, July 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-56.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
56.pdf.

Kheirbek:2014:NAN

[2028] Mazen A. Kheirbek and René Hen. Neuroscience: Add neurons, subtract anxiety. Scientific American, 311(1):62–67, July 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n1/full/scientificamerican0714-62.html;   http://www.nature.
com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
62.pdf.

Heffernan:2014:SB

[2029] Olive Heffernan. Sustainability: Bottoms up. Scientific American, 311(1):68–75, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n1/full/scientificamerican0714-
68.html;   http://www.nature.com/scientificamerican/journal/
v311/n1/pdf/scientificamerican0714-68.pdf.

Jabr:2014:BBW

[2030] Ferris Jabr. Biology: Body works. Scientific American, 311(1):76– 83, July 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n1/full/scientificamerican0714-76.html;   http:
//www.nature.com/scientificamerican/journal/v311/n1/pdf/scientificamerican0714-
76.pdf.

DiChristina:2014:ESU

[2031] Mariette DiChristina. From the Editor: The shadow universe and other marvels. Scientific American, 311(2):4, August 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-4.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
4.pdf.

Anonymous:2014:Lh

[2032] Anonymous. Letters. Scientific American, 311(2):6–7, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-6.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
6.pdf.

Schlenoff:2014:YAIh

[2033] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 311(2):82, August 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-82.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
82.pdf.

Anonymous:2014:SAW

[2034] Anonymous. Science agenda: Will work for machines. Scientific American, 311(2):10, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
10.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-10.pdf.

Subotnik:2014:FTG

[2035] Rena F. Subotnik, Paula Olszewski-Kubilius, and Frank C. Worrell. Forum: The talent gap. Scientific American, 311(2):13, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-13.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
13.pdf.

Biba:2014:AAG

[2036] Erin Biba. Advances: Africa’s great divide. Scientific American, 311(2):17–18, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
17.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-17.pdf.

Arnold:2014:DPP

[2037] Carrie Arnold. Don’t put the “pee” in pool. Scientific American, 311(2): 18, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-18.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
18.pdf.

Iozzio:2014:NNG

[2038] Corinne Iozzio. A network that never goes down. Scientific American, 311(2):19, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
19.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-19.pdf.

Iozzio:2014:SLR

[2039] Corinne Iozzio. Street-legal robots. Scientific American, 311(2):20, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-20.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
20.pdf.

Biello:2014:KJ

[2040] David Biello. Know the jargon. Scientific American, 311(2):21, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-21a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
21a.pdf.

Anonymous:2014:Nq

[2041] Anonymous. By the numbers. Scientific American, 311(2):21, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-21b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
21b.pdf.

Sneed:2014:PT

[2042] Annie Sneed. Pterosaur tech. Scientific American, 311(2):22–23, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-22a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
22a.pdf.

Sneed:2014:NLT

[2043] Annie Sneed. 5 national landmarks threatened by climate change. Scientific American, 311(2):22–23, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v311/n2/full/scientificamerican0814-

22b.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-22b.pdf.

Smith:2014:FF

[2044] Peter Andrey Smith. Foliage friendships. Scientific American, 311(2):24– 25, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-24.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
24.pdf.

Anonymous:2014:Nr

[2045] Anonymous. By the numbers. Scientific American, 311(2):25, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-25.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
25.pdf.

Moskowitz:2014:RC

[2046] Clara Moskowitz. Rendezvous with a comet. Scientific American, 311(2): 26, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-26.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
26.pdf.

Goldman:2014:TCb

[2047] Jason G. Goldman. Top-down control. Scientific American, 311(2):27, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-27.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
27.pdf.

Sneed:2014:WBF

[2048] Annie Sneed. Why babies forget. Scientific American, 311(2):28, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-28a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
28a.pdf.

Williams:2014:WI

[2049] Amber Williams. What is it? Scientific American, 311(2):28, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-28b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
28b.pdf.

Moyer:2014:SHC

[2050] Melinda Wenner Moyer. The science of health: Can brown fat defeat obesity? Scientific American, 311(2):30–32, August 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-30.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
30.pdf.

Pogue:2014:TTT

[2051] David Pogue. TechnoFiles: The two tribes of technology. Scientific American, 311(2):34, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
34.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-34.pdf.

Anonymous:2014:RREa

[2052] Anonymous. Reviews and recommendations: The end of absence: Reclaiming what we’ve lost in a world of constant connection. Scientific American, 311(2):78, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
78a.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-78a.pdf.

Anonymous:2014:OMT

[2053] Anonymous. The organized mind: Thinking straight in the age of information overload. Scientific American, 311(2):78, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-78b.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
78b.pdf.

Anonymous:2014:CDK

[2054] Anonymous. Curious: The desire to know and why your future depends on it. Scientific American, 311(2):78, August 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-78c.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
78c.pdf.

Anonymous:2014:DWW

[2055] Anonymous. Dataclysm: Who we are (when we think no one’s looking). Scientific American, 311(2):78, August 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-78d.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
78d.pdf.

Shermer:2014:SC

[2056] Michael Shermer. Skeptic: ClimeApocalypse! Scientific American, 311(2):79, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
79.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-79.pdf.

Mirsky:2014:AGT

[2057] Steve Mirsky. Anti gravity: Things that go bump. Scientific American, 311(2):80, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
80.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-80.pdf.

Fischetti:2014:GSR

[2058] Mark Fischetti. Graphic science: Rising temperatures hit species hotspots. Scientific American, 311(2):84, August 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-84.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
84.pdf.

Afshordi:2014:CBH

[2059] Niayesh Afshordi, Robert B. Mann, and Razieh Pourhasan. Cosmology: The black hole at the beginning of time. Scientific American, 311(2):36– 43, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-36.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
36.pdf.

Afshordi:2014:WC

[2060] Niayesh Afshordi, Robert B. Mann, and Razieh Pourhasan. Whispers from creation. Scientific American, 311(2):42, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-42.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
42.pdf.

Skinner:2014:BNK

[2061] Michael K. Skinner. Biology: A new kind of inheritance. Scientific American, 311(2):44–51, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
44.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-44.pdf.

Treffert:2014:NAG

[2062] Darold A. Treffert. Neuroscience: Accidental genius. Scientific American, 311(2):52–57, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
52.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-52.pdf.

Solomon:2014:HSS

[2063] Christopher Solomon. Health: Sickness spreads across the Arctic. Scientific American, 311(2):58–63, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v311/n2/full/scientificamerican0814-
58.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-58.pdf.

Pentland:2014:CSB

[2064] Alex ‘Sandy’ Pentland. Cybersecurity: Saving Big Data from itself. Scientific American, 311(2):64–67, August 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n2/full/scientificamerican0814-64.html;   http://www.nature.
com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
64.pdf.

Kantrowitz:2014:ESL

[2065] Barbara Kantrowitz. Education: The science of learning. Scientific American, 311(2):68–73, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
68.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-68.pdf.

Wieman:2014:SLM

[2066] Carl Wieman. Stop lecturing me. Scientific American, 311(2):70–71, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n2/full/scientificamerican0814-70.html;   http:
//www.nature.com/scientificamerican/journal/v311/n2/pdf/scientificamerican0814-
70.pdf.

Scharf:2014:ACS

[2067] Caleb Scharf. Astronomy: Cosmic (in)significance. Scientific American, 311(2):74–77, August 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n2/full/scientificamerican0814-
74.html;   http://www.nature.com/scientificamerican/journal/
v311/n2/pdf/scientificamerican0814-74.pdf.

DiChristina:2014:EHJ

[2068] Mariette DiChristina. From the Editor: Humanity’s journey. Scientific American, 311(3):4, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
4.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-4.pdf.

Anonymous:2014:Li

[2069] Anonymous. Letters. Scientific American, 311(3):6–8, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n3/full/scientificamerican0914-6.html;   http://www.nature.
com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
6.pdf.

Schlenoff:2014:YAIi

[2070] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 311(3):98, September 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n3/full/scientificamerican0914-98.html;   http://www.nature.
com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
98.pdf.

Anonymous:2014:SAFc

[2071] Anonymous. Science agenda: Free up the two-year colleges. Scientific American, 311(3):12, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
12.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-12.pdf.

Fields:2014:FVD

[2072] R. Douglas Fields. Forum: Vive la différence. Scientific American, 311(3):14, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
14.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-14.pdf.

McKee:2014:AWS

[2073] Maggie McKee. Advances: Whispers of a successor. Scientific American, 311(3):19–20, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
19.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-19.pdf.

Bennington-Castro:2014:SS

[2074] Joseph Bennington-Castro. Salt swap. Scientific American, 311(3):20, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-20a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
20a.pdf.

Maron:2014:D

[2075] Dina Fine Maron. Dried up. Scientific American, 311(3):20, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-20b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
20b.pdf.

Moskowitz:2014:OS

[2076] Clara Moskowitz. An origin story. Scientific American, 311(3):21, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-21.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
21.pdf.

Laursen:2014:HRB

[2077] Lucas Laursen. Hitting rock bottom. Scientific American, 311(3):22, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-22.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
22.pdf.

Anonymous:2014:Ns

[2078] Anonymous. By the numbers. Scientific American, 311(3):23, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-23.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
23.pdf.

Maron:2014:LLO

[2079] Dina Fine Maron. Longer-lasting organs. Scientific American, 311(3):24, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-24.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
24.pdf.

Nuwer:2014:PI

[2080] Rachel Nuwer. Plastic on ice. Scientific American, 311(3):25, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-25.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
25.pdf.

Sneed:2014:WIe

[2081] Annie Sneed. What is it? Scientific American, 311(3):26, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-26.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
26.pdf.

Fischetti:2014:STH

[2082] Mark Fischetti. Sleeping through high school. Scientific American, 311(3):27, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
27.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-27.pdf.

Adams:2014:PF

[2083] Jill U. Adams. Pick a fight. Scientific American, 311(3):28, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-28.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
28.pdf.

Drouin:2014:BD

[2084] Roger Drouin. Bat deterrents. Scientific American, 311(3):29, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-29.html;   http:

//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
29.pdf.

Goldman:2014:T

[2085] Jason G. Goldman. The thinker. Scientific American, 311(3):30, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-30a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
30a.pdf.

Sneed:2014:KJa

[2086] Annie Sneed. Know the jargon. Scientific American, 311(3):30, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-30b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
30b.pdf.

Wapner:2014:SHS

[2087] Jessica Wapner. The science of health: The solid-gold wonder drug. Scientific American, 311(3):32–33, September 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n3/full/scientificamerican0914-32.html;   http://www.nature.
com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
32.pdf.

Pogue:2014:TUS

[2088] David Pogue. TechnoFiles: Unsettled score. Scientific American, 311(3):35, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
35.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-35.pdf.

Anonymous:2014:RREb

[2089] Anonymous. Reviews and recommendations: The edge of the sky: All you need to know about the all-there-is. Scientific American, 311(3):92, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-92a.html;   http:

//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
92a.pdf.

Anonymous:2014:WIS

[2090] Anonymous. What if? Serious scientific answers to absurd hypothetical questions. Scientific American, 311(3):92, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n3/full/scientificamerican0914-92b.html;   http://www.nature.
com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
92b.pdf.

Anonymous:2014:MTM

[2091] Anonymous. The marshmallow test: Mastering self-control. Scientific American, 311(3):92, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
92c.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-92c.pdf.

Anonymous:2014:SABb

[2092] Anonymous. Shocked: Adventures in bringing back the recently dead. Scientific American, 311(3):92, September 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n3/full/scientificamerican0914-92d.html;   http://www.nature.
com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
92d.pdf.

Shermer:2014:SSS

[2093] Michael Shermer. Skeptic: Surviving statistics. Scientific American, 311(3):94, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
94.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-94.pdf.

Mirsky:2014:AGCb

[2094] Steve Mirsky. Anti gravity: It’s a cookbook! Scientific American, 311(3):96, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-

96.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-96.pdf.

Wong:2014:GSP

[2095] Kate Wong. Graphic science: The 1 percent difference. Scientific American, 311(3):100, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
100.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-100.pdf.

Wong:2014:EHS

[2096] Kate Wong. Evolution: The human saga: Evolution rewritten. Scientific American, 311(3):36–39, September 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n3/full/scientificamerican0914-36.html;   http://www.nature.
com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
36.pdf.

Anonymous:2014:WWC

[2097] Anonymous. Where we came from. Scientific American, 311(3):40–41, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-40.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
40.pdf.

Wood:2014:WF

[2098] Bernard Wood. Welcome to the family. Scientific American, 311(3):42–47, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
42.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-42.pdf.

deMenocal:2014:CS

[2099] Peter B. deMenocal. Climate shocks. Scientific American, 311(3):48–53, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-48.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
48.pdf.

Tattersall:2014:IH

[2100] Ian Tattersall. If I had a hammer. Scientific American, 311(3):54–59, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-54.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
54.pdf.

Anonymous:2014:WMU

[2101] Anonymous. What makes us special. Scientific American, 311(3):60–61, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-60.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
60.pdf.

Edgar:2014:PT

[2102] Blake Edgar. Powers of two. Scientific American, 311(3):62–67, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-62.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
62.pdf.

deWaal:2014:OA

[2103] Frans de Waal. One for all. Scientific American, 311(3):68–71, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-68.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
68.pdf.

Stix:2014:IF

[2104] Gary Stix. The “it” factor. Scientific American, 311(3):72–79, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-72.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
72.pdf.

Anonymous:2014:WWG

[2105] Anonymous. Where we are going. Scientific American, 311(3):80–81, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n3/full/scientificamerican0914-80.html;   http:
//www.nature.com/scientificamerican/journal/v311/n3/pdf/scientificamerican0914-
80.pdf.

Fischetti:2014:NP

[2106] Mark Fischetti. The networked primate. Scientific American, 311(3):82–85, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
82.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-82.pdf.

Hawks:2014:SEA

[2107] John Hawks. Still evolving (after all these years). Scientific American, 311(3):86–91, September 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n3/full/scientificamerican0914-
86.html;   http://www.nature.com/scientificamerican/journal/
v311/n3/pdf/scientificamerican0914-86.pdf.

DiChristina:2014:EYI

[2108] Mariette DiChristina. From the Editor: You’re invited. Scientific American, 311(4):6, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
6.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-6.pdf.

Anonymous:2014:Lj

[2109] Anonymous. Letters. Scientific American, 311(4):8–10, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-8.html;   http://www.nature.
com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
8.pdf.

Schlenoff:2014:YAIj

[2110] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 311(4):99, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-99.html;   http:

//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
99.pdf.

Anonymous:2014:SAPc

[2111] Anonymous. Science agenda: Preferential treatment. Scientific American, 311(4):12, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
12.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-12.pdf.

Nelson:2014:FPM

[2112] Ben Nelson. Forum: Passing the midterm. Scientific American, 311(4): 15, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-15.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
15.pdf.

Haelle:2014:AWG

[2113] Tara Haelle. Advances: When good proteins go bad. Scientific American, 311(4):19–20, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
19.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-19.pdf.

Scheer:2014:SS

[2114] Rachel Scheer and Annie Sneed. Safety in a sock. Scientific American, 311(4):20, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
20.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-20.pdf.

Goldman:2014:KJ

[2115] Jason G. Goldman. Know the jargon. Scientific American, 311(4):21, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-21.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
21.pdf.

Drouin:2014:SB

[2116] Roger Drouin. Saving “Bambi”. Scientific American, 311(4):22–23, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-22.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
22.pdf.

Frazer:2014:AD

[2117] Jennifer Frazer. Amoebas on deathwatch. Scientific American, 311(4):24–25, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
24.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-24.pdf.

Iozzio:2014:VR

[2118] Corinne Iozzio. Virtually revolutionary. Scientific American, 311(4):26, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-26.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
26.pdf.

Calderone:2014:STS

[2119] Julia Calderone. See-through science. Scientific American, 311(4):27, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-27a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
27a.pdf.

Schultz:2014:WI

[2120] Kevin Schultz. What is it? Scientific American, 311(4):27, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-27b.html;   http://www.nature.
com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
27b.pdf.

Sneed:2014:HBA

[2121] Annie Sneed. How big animals deter cancer. Scientific American, 311(4): 28, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-28.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
28.pdf.

Sneed:2014:EIM

[2122] Annie Sneed. Earth’s impending magnetic flip. Scientific American, 311(4):29, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
29.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-29.pdf.

Biba:2014:LG

[2123] Erin Biba. Life guard. Scientific American, 311(4):30, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-30.html;   http://www.nature.
com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
30.pdf.

Storrs:2014:SHO

[2124] Carina Storrs. The science of health: The other T party. Scientific American, 311(4):32–34, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
32.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-32.pdf.

Pogue:2014:TMD

[2125] David Pogue. TechnoFiles: Million-dollar stunt. Scientific American, 311(4):36, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
36.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-36.pdf.

Anonymous:2014:RRW

[2126] Anonymous. Reviews and recommendations: WTF, evolution?! A theory of unintelligible design. Scientific American, 311(4):96, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-96a.html;   http://www.nature.

com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
96a.pdf.

Anonymous:2014:YAS

[2127] Anonymous. Your atomic self: The invisible elements that connect you to everything in the universe. Scientific American, 311(4):96, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-96b.html;   http://www.nature.
com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
96b.pdf.

Anonymous:2014:AFS

[2128] Anonymous. Arrival of the fittest: Solving Evolution’s greatest puzzle. Scientific American, 311(4):96, October 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-96c.html;   http://www.nature.
com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
96c.pdf.

Anonymous:2014:AIS

[2129] Anonymous. Alive inside: A story of music & memory. Scientific American, 311(4):96, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
96d.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-96d.pdf.

Shermer:2014:SI

[2130] Michael Shermer. Skeptic: Infrequencies. Scientific American, 311(4):97, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-97.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
97.pdf.

Mirsky:2014:AGY

[2131] Steve Mirsky. Anti gravity: The 2,000-year-old joke. Scientific American, 311(4):98, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-

98.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-98.pdf.

Moskowitz:2014:GSF

[2132] Clara Moskowitz. Graphic science: Family histories. Scientific American, 311(4):100, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
100.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-100.pdf.

Guterl:2014:SWS

[2133] Fred Guterl. The state of the world’s science 2014: The inclusion equation. Scientific American, 311(4):38–40, October 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-38.html;   http://www.nature.
com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
38.pdf.

Guterl:2014:WD

[2134] Fred Guterl. Where are the data? Scientific American, 311(4):40–41, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-40.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
40.pdf.

Phillips:2014:HDW

[2135] Katherine W. Phillips. How diversity works. Scientific American, 311(4):42–47, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
42.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-42.pdf.

Medin:2014:PPV

[2136] Douglas Medin, Carol D. Lee, and Megan Bang. Particular points of view. Scientific American, 311(4):44–45, October 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-44.html;   http://www.nature.

com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
44.pdf.

Bishop:2014:SE

[2137] Steven Bishop. Science exposed. Scientific American, 311(4):46, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-46.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
46.pdf.

Lee:2014:TIP

[2138] D. N. Lee. Taking it personally. Scientific American, 311(4):47, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-47.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
47.pdf.

Hill:2014:PBI

[2139] Stephanie C. Hill. In pursuit of the best ideas. Scientific American, 311(4):48–49, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
48.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-48.pdf.

Anonymous:2014:GG

[2140] Anonymous. Gender gap. Scientific American, 311(4):50–51, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-50.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
50.pdf.

Plaut:2014:IE

[2141] Victoria Plaut. Inviting everyone in. Scientific American, 311(4):52–57, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n4/full/scientificamerican1014-52.html;   http:
//www.nature.com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
52.pdf.

Welle:2014:BV

[2142] Brian Welle and Megan Smith. Becoming visible. Scientific American, 311(4):54–55, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
54.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-54.pdf.

Krauss:2014:CBB

[2143] Lawrence M. Krauss. Cosmology: A beacon from the Big Bang. Scientific American, 311(4):58–67, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
58.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-58.pdf.

Rosner:2014:ASC

[2144] Hillary Rosner. Agriculture: Saving coffee. Scientific American, 311(4):68–73, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
68.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-68.pdf.

Piccolo:2014:CBT

[2145] Stefano Piccolo. Cell biology: Twists of fate. Scientific American, 311(4):74–81, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
74.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-74.pdf.

Margonelli:2014:EII

[2146] Lisa Margonelli. Environment: An inconvenient ice. Scientific American, 311(4):82–89, October 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n4/full/scientificamerican1014-
82.html;   http://www.nature.com/scientificamerican/journal/
v311/n4/pdf/scientificamerican1014-82.pdf.

Mulcahy:2014:MLG

[2147] Colm Mulcahy and Dana Richards. Mathematics: Let the games continue. Scientific American, 311(4):90–95, October 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n4/full/scientificamerican1014-90.html;   http://www.nature.
com/scientificamerican/journal/v311/n4/pdf/scientificamerican1014-
90.pdf.

DiChristina:2014:EMM

[2148] Mariette DiChristina. From the Editor: Mindful, medicinal, malleable and marketable. Scientific American, 311(5):6, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n5/full/scientificamerican1114-6.html;   http://www.nature.
com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
6.pdf.

Anonymous:2014:Lk

[2149] Anonymous. Letters. Scientific American, 311(5):8–9, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n5/full/scientificamerican1114-8.html;   http://www.nature.
com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
8.pdf.

Schlenoff:2014:YAIk

[2150] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 311(5):89, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-89.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
89.pdf.

Anonymous:2014:SAD

[2151] Anonymous. Science agenda: Do no harm — and no drugs. Scientific American, 311(5):12, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
12.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-12.pdf.

Zittrain:2014:FCK

[2152] Jonathan Zittrain. Forum: The case for kill switches. Scientific American, 311(5):14, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
14.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-14.pdf.

Dillow:2014:ARP

[2153] Clay Dillow. Advances: Return of the propeller. Scientific American, 311(5):17–21, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
17.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-17.pdf.

Nuzzo:2014:PD

[2154] Regina Nuzzo. Pollsters’ dilemma. Scientific American, 311(5):21, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-21.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
21.pdf.

Goldman:2014:WDH

[2155] Jason G. Goldman. Which direction home? Scientific American, 311(5):22, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
22.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-22.pdf.

Laursen:2014:WEM

[2156] Lucas Laursen. When evidence melts away. Scientific American, 311(5):23, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
23.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-23.pdf.

Crow:2014:KJ

[2157] Diana Crow. Know the jargon. Scientific American, 311(5):24, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-24.html;   http:

//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
24.pdf.

Greenemeier:2014:BH

[2158] Larry Greenemeier. The ball that hovers. Scientific American, 311(5):25, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-25.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
25.pdf.

Baggaley:2014:SS

[2159] Kate Baggaley. Snaking stairs. Scientific American, 311(5):26, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-26.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
26.pdf.

Anonymous:2014:QHa

[2160] Anonymous. Quick hits. Scientific American, 311(5):27, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n5/full/scientificamerican1114-27a.html;   http://www.nature.
com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
27a.pdf.

Christiansen:2014:WI

[2161] Jen Christiansen. What is it? Scientific American, 311(5):27, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-27b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
27b.pdf.

Wong:2014:HH

[2162] Kate Wong. Human or hobbit? Scientific American, 311(5):28–29, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-28.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
28.pdf.

Gammon:2014:IAS

[2163] Katharine Gammon. Inside the audience studio. Scientific American, 311(5):30–31, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
30.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-30.pdf.

Weiner:2014:CSR

[2164] Debra Weiner. Catching some rays. Scientific American, 311(5):32, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-32.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
32.pdf.

Levine:2014:SHK

[2165] James Levine. The science of health: Killer chairs. Scientific American, 311(5):34–35, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
34.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-34.pdf.

Pogue:2014:TCT

[2166] David Pogue. TechnoFiles: Crash test dummies. Scientific American, 311(5):36, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
36.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-36.pdf.

Anonymous:2014:RRLb

[2167] Anonymous. Reviews and recommendations: Lock in. Scientific American, 311(5):86, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
86a.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-86a.pdf.

Anonymous:2014:Pb

[2168] Anonymous. The peripheral. Scientific American, 311(5):86, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-86b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
86b.pdf.

Anonymous:2014:AS

[2169] Anonymous. Ancillary sword. Scientific American, 311(5):86, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-86c.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
86c.pdf.

Anonymous:2014:S

[2170] Anonymous. Symbiont. Scientific American, 311(5):86, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n5/full/scientificamerican1114-86d.html;   http://www.nature.
com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
86d.pdf.

Shermer:2014:SPP

[2171] Michael Shermer. Skeptic: Perpetual peace. Scientific American, 311(5):88, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
88.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-88.pdf.

Fischetti:2014:GSWb

[2172] Mark Fischetti and Kevin Schultz. Graphic science: Who has the best fuel economy? Scientific American, 311(5):90, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n5/full/scientificamerican1114-90.html;   http://www.nature.
com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
90.pdf.

Mirsky:2014:AGCc

[2173] Steve Mirsky. Anti gravity: The call of the mild. Scientific American, 311(5):92, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-

92.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-92.pdf.

Ricard:2014:NMM

[2174] Matthieu Ricard, Antoine Lutz, and Richard J. Davidson. Neuroscience: Mind of the meditator. Scientific American, 311(5):38–45, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-38.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
38.pdf.

Lemonick:2014:SPB

[2175] Michael D. Lemonick. Space: Pluto and beyond. Scientific American, 311(5):46–53, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
46.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-46.pdf.

Mahoney:2014:MVT

[2176] Douglas J. Mahoney, David F. Stojdl, and Gordon Laird. Medicine: Virus therapy for cancer. Scientific American, 311(5):54–59, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-54.html;   http:
//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
54.pdf.

Campbell:2014:MEP

[2177] Thomas A. Campbell, Skylar Tibbits, and Banning Garrett. Mechanical engineering: The programmable world. Scientific American, 311(5):60–65, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
60.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-60.pdf.

Biello:2014:ESW

[2178] David Biello. Energy: Solar wars. Scientific American, 311(5):66–71, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n5/full/scientificamerican1114-66.html;   http:

//www.nature.com/scientificamerican/journal/v311/n5/pdf/scientificamerican1114-
66.pdf.

Dunn:2014:LSE

[2179] Rob Dunn. Life science: The evolution of architecture. Scientific American, 311(5):72–77, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
72.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-72.pdf.

Noonan:2014:AIF

[2180] David Noonan. Aeronautics: Impossible flight. Scientific American, 311(5):78–83, November 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n5/full/scientificamerican1114-
78.html;   http://www.nature.com/scientificamerican/journal/
v311/n5/pdf/scientificamerican1114-78.pdf.

DiChristina:2014:EDC

[2181] Mariette DiChristina. From the Editor: The discovery continuum. Scientific American, 311(6):4, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
4.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-4.pdf.

Anonymous:2014:Ll

[2182] Anonymous. Letters. Scientific American, 311(6):6–8, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n6/full/scientificamerican1214-6.html;   http://www.nature.
com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
6.pdf.

Schlenoff:2014:YAIl

[2183] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 311(6):98, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-98.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
98.pdf.

Anonymous:2014:SAC

[2184] Anonymous. Science agenda: Caution: Cops with cameras. Scientific American, 311(6):10, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
10.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-10.pdf.

Calo:2014:FBR

[2185] Ryan Calo. Forum: Bring on the robocrats. Scientific American, 311(6):12, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
12.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-12.pdf.

Biello:2014:ACR

[2186] David Biello. Advances: Cities to the rescue. Scientific American, 311(6):15–19, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
15.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-15.pdf.

Nordrum:2014:FC

[2187] Amy Nordrum. Forecast: Cloudy. Scientific American, 311(6):19, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-19a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
19a.pdf.

Bourzac:2014:BD

[2188] Katherine Bourzac. Bend by design. Scientific American, 311(6):19, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-19b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
19b.pdf.

Moskowitz:2014:DSB

[2189] Clara Moskowitz. Deep space or bust. Scientific American, 311(6):20, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-20.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
20.pdf.

Drouin:2014:WWJ

[2190] Roger Drouin. What would John Muir do now? Scientific American, 311(6):22, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
22.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-22.pdf.

StFleur:2014:GI

[2191] Nicholas St. Fleur. A Galápagos icon. Scientific American, 311(6):23, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-23.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
23.pdf.

Sneed:2014:KJb

[2192] Annie Sneed. Know the jargon. Scientific American, 311(6):24, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-24.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
24.pdf.

Smith:2014:MM

[2193] Peter Andrey Smith. Mushroom man. Scientific American, 311(6):28–29, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-28.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
28.pdf.

Iozzio:2014:PIT

[2194] Corinne Iozzio. Power to the Internet of Things. Scientific American, 311(6):30, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-

30.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-30.pdf.

Anonymous:2014:QHb

[2195] Anonymous. Quick hits. Scientific American, 311(6):31, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n6/full/scientificamerican1214-31.html;   http://www.nature.
com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
31.pdf.

Goldman:2014:CC

[2196] Jason G. Goldman. Call of the crybaby. Scientific American, 311(6):32, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-32.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
32.pdf.

Kuchment:2014:GG

[2197] Anna Kuchment. Of germs and germination. Scientific American, 311(6):33, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
33.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-33.pdf.

Wallis:2014:SHN

[2198] Claudia Wallis. The science of health: Never too old for chemo. Scientific American, 311(6):34–35, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v311/n6/full/scientificamerican1214-
34.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-34.pdf.

Pogue:2014:TUT

[2199] David Pogue. TechnoFiles: Unpredictive text. Scientific American, 311(6):38, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
38.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-38.pdf.

Anonymous:2014:RRLc

[2200] Anonymous. Reviews and recommendations: The lost elements: The Periodic Table’s shadow side. Scientific American, 311(6):92, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-92a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
92a.pdf.

Anonymous:2014:SPF

[2201] Anonymous. Spare parts: Four undocumented teenagers, one ugly robot, and the battle for the American Dream. Scientific American, 311(6):92, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-92b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
92b.pdf.

Anonymous:2014:DRI

[2202] Anonymous. Digging for Richard III: The search for the lost king. Scientific American, 311(6):92, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v311/n6/full/scientificamerican1214-
92c.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-92c.pdf.

Shermer:2014:SCC

[2203] Michael Shermer. Skeptic: Conspiracy central. Scientific American, 311(6):94, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
94.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-94.pdf.

Mirsky:2014:AGK

[2204] Steve Mirsky. Anti gravity: Keeping up with the times. Scientific American, 311(6):96, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
96.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-96.pdf.

Fischetti:2014:GS

[2205] Mark Fischetti. Graphic science: Up, up and away. Scientific American, 311(6):100, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
100.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-100.pdf.

Fletcher:2014:WCI

[2206] Seth Fletcher. World changing ideas 2014. Scientific American, 311(6):40–41, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
40.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-40.pdf.

Knox:2014:WCI

[2207] Margaret Knox. World changing ideas 2014: The gene Genie. Scientific American, 311(6):42–46, December 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n6/full/scientificamerican1214-42.html;   http://www.nature.
com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
42.pdf.

Bradley:2014:RC

[2208] Ryan Bradley. Reprogrammable cells. Scientific American, 311(6):47, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-47.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
47.pdf.

Bradley:2014:TO

[2209] Ryan Bradley. Transparent organisms. Scientific American, 311(6):48, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-48a.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
48a.pdf.

Bradley:2014:SFF

[2210] Ryan Bradley. Spit-fired fuel cells. Scientific American, 311(6):48, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-48b.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
48b.pdf.

Nuwer:2014:VCD

[2211] Rachel Nuwer. Vision-correcting displays. Scientific American, 311(6):49, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
49.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-49.pdf.

Geim:2014:ASL

[2212] Andre K. Geim. Atomic-scale legos. Scientific American, 311(6):50–51, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v311/n6/full/scientificamerican1214-50.html;   http:
//www.nature.com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
50.pdf.

Nuwer:2014:URP

[2213] Rachel Nuwer. Ultrahard recyclable plastics. Scientific American, 311(6):52, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
52a.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-52a.pdf.

Nuwer:2014:WCS

[2214] Rachel Nuwer. Wireless charging with sound waves. Scientific American, 311(6):52, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
52b.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-52b.pdf.

Bradley:2014:BCL

[2215] Ryan Bradley. Batteries that capture low-grade waste heat. Scientific American, 311(6):53, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
53a.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-53a.pdf.

Fogelson:2014:VCN

[2216] Ben Fogelson. Video cameras for nanoparticles. Scientific American, 311(6):53, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
53b.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-53b.pdf.

Johnston:2014:AFH

[2217] Kathryn V. Johnston. Astronomy: Fossil hunting in the Milky Way. Scientific American, 311(6):54–59, December 2014. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v311/
n6/full/scientificamerican1214-54.html;   http://www.nature.
com/scientificamerican/journal/v311/n6/pdf/scientificamerican1214-
54.pdf.

Sutherland:2014:MPW

[2218] Stephani Sutherland. Medicine: Pain that won’t quit. Scientific American, 311(6):60–67, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
60.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-60.pdf.

Peplow:2014:MTI

[2219] Mark Peplow. More treatment ideas: Taking the sting out of pain. Scientific American, 311(6):66, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v311/n6/full/scientificamerican1214-
66.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-66.pdf.

Masters:2014:CJS

[2220] Jeff Masters. Climate: The jet stream is getting weird. Scientific American, 311(6):68–75, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-

68.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-68.pdf.

Zorich:2014:ASG

[2221] Zach Zorich. Archaeology: The Storm God’s tale. Scientific American, 311(6):76–79, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
76.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-76.pdf.

Kucharski:2014:III

[2222] Adam J. Kucharski. Immunology: Immunity’s illusion. Scientific American, 311(6):80–85, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
80.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-80.pdf.

Nelson:2014:LSS

[2223] Ximena Nelson. Life science: The spider’s charade. Scientific American, 311(6):86–89, December 2014. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v311/n6/full/scientificamerican1214-
86.html;   http://www.nature.com/scientificamerican/journal/
v311/n6/pdf/scientificamerican1214-86.pdf.

DiChristina:2015:EMW

[2224] Mariette DiChristina. From the Editor: Many worlds. Scientific American, 312(1):4, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
4.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-4.pdf.

Anonymous:2015:La

[2225] Anonymous. Letters. Scientific American, 312(1):6–8, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n1/full/scientificamerican0115-6.html;   http://www.nature.
com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
6.pdf.

Schlenoff:2015:YAIa

[2226] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 312(1):85, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-85.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
85.pdf.

Brody:2015:ONO

[2227] Herb Brody. Outlook: Nature outlook: Haemophilia. Scientific American, 312(1):S2–S3, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
S2.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-S2.pdf.

Savage:2015:BB

[2228] Neil Savage. Born in the blood. Scientific American, 312(1):S4–S5, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-S4.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
S4.pdf.

Gould:2015:GTG

[2229] Julie Gould. Gene therapy: Genie in a vector. Scientific American, 312(1):S6–S7, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
S6.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-S6.pdf.

Savage:2015:CFS

[2230] Neil Savage. Clotting factors: Stretching time. Scientific American, 312(1):S8–S10, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
S8.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-S8.pdf.

Pemberton:2015:PF

[2231] Stephen Pemberton. Perspective: The fix is in. Scientific American, 312(1):S11, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
S11.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-S11.pdf.

Dolgin:2015:IOS

[2232] Elie Dolgin. Immunology: Oral solutions. Scientific American, 312(1):S12–S13, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
S12.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-S12.pdf.

Willyard:2015:TBA

[2233] Cassandra Willyard. Thrombosis: Balancing act. Scientific American, 312(1):S14–S15, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
S14.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-S14.pdf.

Gammon:2015:OJE

[2234] Katharine Gammon. Orthopaedics: Joint effort. Scientific American, 312(1):S16–S17, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
S16.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-S16.pdf.

Sohn:2015:AMD

[2235] Emily Sohn. Animal models: Dogged pursuit. Scientific American, 312(1):S18–S19, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
S18.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-S18.pdf.

Anonymous:2015:SAHa

[2236] Anonymous. Science agenda: A hacker’s guide to planet cooling. Scientific American, 312(1):10, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
10.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-10.pdf.

Frey:2015:FEE

[2237] Carl Benedikt Frey. Forum: The end of economic growth? Scientific American, 312(1):12, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
12.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-12.pdf.

Billings:2015:ADA

[2238] Lee Billings. Advances: Dancing with the asteroids. Scientific American, 312(1):13–14, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
13.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-13.pdf.

Patel:2015:PP

[2239] Prachi Patel. The pulse of pacemakers. Scientific American, 312(1):14, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-14.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
14.pdf.

Smith:2015:BTL

[2240] Peter Andrey Smith. Bacteria? They love all manure. Scientific American, 312(1):15, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
15.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-15.pdf.

Moskowitz:2015:TPS

[2241] Clara Moskowitz. Theoretical particles, still theoretical. Scientific American, 312(1):16, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-

16.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-16.pdf.

Shiffman:2015:MM

[2242] David Shiffman. Monikers matter. Scientific American, 312(1):17, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-17.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
17.pdf.

Yuhas:2015:PC

[2243] Daisy Yuhas. Parental controls. Scientific American, 312(1):18, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-18.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
18.pdf.

Calderone:2015:ES

[2244] Julia Calderone. Ebola suits. Scientific American, 312(1):19, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-19.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
19.pdf.

Nordrum:2015:HAG

[2245] Amy Nordrum. Health advice from a grizzly. Scientific American, 312(1): 20, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-20.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
20.pdf.

Anonymous:2015:QHa

[2246] Anonymous. Quick hits. Scientific American, 312(1):21, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n1/full/scientificamerican0115-21.html;   http://www.nature.
com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
21.pdf.

Johnson:2015:RR

[2247] Scott K. Johnson. From rain to ranges. Scientific American, 312(1):22– 23, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-22a.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
22a.pdf.

Croswell:2015:DBN

[2248] Ken Croswell. The dust belt next door. Scientific American, 312(1):24, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-24.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
24.pdf.

Maron:2015:SHWa

[2249] Dina Fine Maron. The science of health: When DNA means “do not ask”. Scientific American, 312(1):28–30, January 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n1/full/scientificamerican0115-28.html;   http://www.nature.
com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
28.pdf.

Pogue:2015:TYN

[2250] David Pogue. TechnoFiles: You: By the numbers. Scientific American, 312(1):31, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
31.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-31.pdf.

Anonymous:2015:RRC

[2251] Anonymous. Reviews and recommendations: Cosmigraphics: Picturing space through time. Scientific American, 312(1):82, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n1/full/scientificamerican0115-82a.html;   http://www.nature.
com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
82a.pdf.

Anonymous:2015:FZM

[2252] Anonymous. Finding zero: A Mathematician’s odyssey to uncover the origins of numbers. Scientific American, 312(1):82, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n1/full/scientificamerican0115-82b.html;   http://www.nature.
com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
82b.pdf.

Anonymous:2015:MFB

[2253] Anonymous. Malformed: Forgotten brains of the Texas State Mental Hospital. Scientific American, 312(1):82, January 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n1/full/scientificamerican0115-82c.html;   http://www.nature.
com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
82c.pdf.

Anonymous:2015:MWC

[2254] Anonymous. The man who couldn’t stop: OCD and the true story of a life lost in thought. Scientific American, 312(1):82, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n1/full/scientificamerican0115-82d.html;   http://www.nature.
com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
82d.pdf.

Shermer:2015:SHZ

[2255] Michael Shermer. Skeptic: Here be zombies. Scientific American, 312(1): 83, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-83.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
83.pdf.

Mirsky:2015:AGBa

[2256] Steve Mirsky. Anti gravity: From big men to Bronx bison. Scientific American, 312(1):84, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
84.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-84.pdf.

Fischetti:2015:GSSa

[2257] Mark Fischetti. Graphic science: Sass vs. substance. Scientific American, 312(1):86, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
86.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-86.pdf.

Heller:2015:ABT

[2258] René Heller. Astronomy: Better than Earth. Scientific American, 312(1):32–39, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
32.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-32.pdf.

Zimmer:2015:MWB

[2259] Carl Zimmer. Medicine: A weakness in Bacteria’s fortress. Scientific American, 312(1):40–45, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
40.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-40.pdf.

Durand:2015:TWM

[2260] Frédo Durand, William T. Freeman, and Michael Rubinstein. Technology: A world — of — movement. Scientific American, 312(1):46–51, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-46.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
46.pdf.

Wong:2015:BLL

[2261] Kate Wong. Biology: Living large. Scientific American, 312(1):52–59, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n1/full/scientificamerican0115-52.html;   http:
//www.nature.com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
52.pdf.

Nicholas:2015:FWW

[2262] Kimberly A. Nicholas. Food: Will we still enjoy pinot noir? Scientific American, 312(1):60–67, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
60.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-60.pdf.

Hilts:2015:ESS

[2263] Philip J. Hilts. Exploration: In search of sunken treasure. Scientific American, 312(1):68–75, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
68.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-68.pdf.

Nordrum:2015:TDU

[2264] Amy Nordrum. Timeline: Diving into underwater exploration. Scientific American, 312(1):74–75, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
74.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-74.pdf.

Nahmias:2015:NWW

[2265] Eddy Nahmias. Neuroscience: Why we have free will. Scientific American, 312(1):76–79, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n1/full/scientificamerican0115-
76.html;   http://www.nature.com/scientificamerican/journal/
v312/n1/pdf/scientificamerican0115-76.pdf.

Reininger:2015:PFM

[2266] Armin J. Reininger, Alessandro Gringeri, Leonard A. Valentino, Gerald Spotts, Vadim Romanov, Robert Numerof, Friedrich Scheiflinger, and Bruce M. Ewenstein. Pushing the frontiers of medicine: Innovations in haemophilia care: Pushing the frontiers of medicine: innovations in haemophilia care. Scientific American, 312(1):W1–W4, January 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n1/full/scientificamerican0115-87.html;   http://www.nature.
com/scientificamerican/journal/v312/n1/pdf/scientificamerican0115-
87.pdf.

DiChristina:2015:EDY

[2267] Mariette DiChristina. From the Editor: Different yet alike. Scientific American, 312(2):6, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
6.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-6.pdf.

Anonymous:2015:Lb

[2268] Anonymous. Letters. Scientific American, 312(2):8–10, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n2/full/scientificamerican0215-8.html;   http://www.nature.
com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
8.pdf.

Schlenoff:2015:YAIb

[2269] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 312(2):86, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-86.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
86.pdf.

Anonymous:2015:SADa

[2270] Anonymous. Science agenda: Don’t kill the goose. Scientific American, 312(2):12, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
12.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-12.pdf.

Lipsitch:2015:FDB

[2271] Marc Lipsitch. Forum: Defusing a biological Bomb. Scientific American, 312(2):14, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
14.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-14.pdf.

Nordrum:2015:ARC

[2272] Amy Nordrum. Advances: The rise of the cable-stayed bridge. Scientific American, 312(2):15–19, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
15.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-15.pdf.

Godkin:2015:HGR

[2273] David Godkin. Hadrosaur gave T. rex a run for its money. Scientific American, 312(2):19, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
19.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-19.pdf.

Sneed:2015:OP

[2274] Annie Sneed. The origin of power. Scientific American, 312(2):20–21, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-20.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
20.pdf.

Landhuis:2015:AW

[2275] Esther Landhuis. Advance warning. Scientific American, 312(2):22, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-22.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
22.pdf.

Nordrum:2015:BB

[2276] Amy Nordrum. Bloody beetroot. Scientific American, 312(2):23, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-23.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
23.pdf.

Nuwer:2015:FP

[2277] Rachel Nuwer. Fit for play. Scientific American, 312(2):24, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-24.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
24.pdf.

Anonymous:2015:QHb

[2278] Anonymous. Quick hits. Scientific American, 312(2):25, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n2/full/scientificamerican0215-25.html;   http://www.nature.
com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
25.pdf.

Goldman:2015:CHL

[2279] Jason G. Goldman. Cry “havoc” and let slip the bees of war. Scientific American, 312(2):26, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
26.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-26.pdf.

Williams:2015:DMA

[2280] Amber Williams. Does mathematical ability predict career success? Scientific American, 312(2):28, February 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n2/full/scientificamerican0215-28a.html;   http://www.nature.
com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
28a.pdf.

Hemenway:2015:IS

[2281] Brett Hemenway and Bill Welser. Insecure skies. Scientific American, 312(2):28–29, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
28b.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-28b.pdf.

Nordrum:2015:NV

[2282] Amy Nordrum. New vibrations. Scientific American, 312(2):30, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-30.html;   http:

//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
30.pdf.

Noonan:2015:SHM

[2283] David Noonan. The science of health: Marijuana’s medical future. Scientific American, 312(2):32–34, February 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n2/full/scientificamerican0215-32.html;   http://www.nature.
com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
32.pdf.

Pogue:2015:TFP

[2284] David Pogue. TechnoFiles: The future is plastic. Scientific American, 312(2):35, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
35.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-35.pdf.

Anonymous:2015:RRM

[2285] Anonymous. Reviews and recommendations: Melting away. Scientific American, 312(2):82, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
82a.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-82a.pdf.

Anonymous:2015:IA

[2286] Anonymous. The interstellar age. Scientific American, 312(2):82, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-82b.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
82b.pdf.

Anonymous:2015:Pa

[2287] Anonymous. The powerhouse. Scientific American, 312(2):82, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-82c.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
82c.pdf.

Anonymous:2015:IF

[2288] Anonymous. Island on fire. Scientific American, 312(2):82, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-82d.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
82d.pdf.

Shermer:2015:SMS

[2289] Michael Shermer. Skeptic: A moral starting point. Scientific American, 312(2):84, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
84.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-84.pdf.

Mirsky:2015:AGIa

[2290] Steve Mirsky. Anti gravity: Insect aside. Scientific American, 312(2):85, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n2/full/scientificamerican0215-85.html;   http:
//www.nature.com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
85.pdf.

Fischetti:2015:GSSb

[2291] Mark Fischetti. Graphic science: The steady rise of Ebola. Scientific American, 312(2):88, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
88.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-88.pdf.

Wong:2015:HEN

[2292] Kate Wong. Human evolution: Neandertal minds. Scientific American, 312(2):36–43, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
36.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-36.pdf.

Brown:2015:ACW

[2293] Adam Brown. Astrophysics: Can we mine a black hole? Scientific American, 312(2):44–49, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
44.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-44.pdf.

Summa:2015:CCW

[2294] Keith C. Summa and Fred W. Turek. Chronobiology: The clocks within us. Scientific American, 312(2):50–55, February 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n2/full/scientificamerican0215-50.html;   http://www.nature.
com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
50.pdf.

DiVentra:2015:CSJ

[2295] Massimiliano Di Ventra and Yuriy V. Pershin. Computer science: Just add memory. Scientific American, 312(2):56–61, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n2/full/scientificamerican0215-56.html;   http://www.nature.
com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
56.pdf.

Webber:2015:EPP

[2296] Michael E. Webber. Environment: A puzzle for the planet. Scientific American, 312(2):62–67, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n2/full/scientificamerican0215-
62.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-62.pdf.

Lozano:2015:NTD

[2297] Andres M. Lozano and Helen S. Mayberg. Neuroscience: Treating depression at the source. Scientific American, 312(2):68–73, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n2/full/scientificamerican0215-68.html;   http://www.nature.
com/scientificamerican/journal/v312/n2/pdf/scientificamerican0215-
68.pdf.

Featherstone:2015:NFS

[2298] Steven Featherstone. Nuclear fallout: The swallows of Fukushima. Scientific American, 312(2):74–81, February 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v312/n2/full/scientificamerican0215-
74.html;   http://www.nature.com/scientificamerican/journal/
v312/n2/pdf/scientificamerican0215-74.pdf.

DiChristina:2015:EBW

[2299] Mariette DiChristina. From the Editor: The battle within and without for healthier lives. Scientific American, 312(3):3, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n3/full/scientificamerican0315-3.html;   http://www.nature.
com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
3.pdf.

Anonymous:2015:Lc

[2300] Anonymous. Letters. Scientific American, 312(3):4–5, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n3/full/scientificamerican0315-4.html;   http://www.nature.
com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
4.pdf.

Schlenoff:2015:YAIc

[2301] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 312(3):77, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-77.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
77.pdf.

Grogan:2015:IMW

[2302] David Grogan. Innovations in: The microbes within. Scientific American, 312(3):S1–S2, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
S1.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-S1.pdf.

Velasquez-Manoff:2015:GMP

[2303] Moises Velasquez-Manoff. Gut microbiome: The peacekeepers. Scientific American, 312(3):S3–S11, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
S3.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-S3.pdf.

Knight:2015:WMT

[2304] Rob Knight. Why microbiome treatments could pay off soon. Scientific American, 312(3):S5, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
S5.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-S5.pdf.

Ley:2015:GML

[2305] Ruth E. Ley. The gene–microbe link. Scientific American, 312(3):S7, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-S7.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
S7.pdf.

Anonymous:2015:YMW

[2306] Anonymous. Your microbes at work: Fiber fermenters keep us healthy. Scientific American, 312(3):S9, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v312/n3/full/scientificamerican0315-
S9.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-S9.pdf.

Sonnenburg:2015:ME

[2307] Justin L. Sonnenburg. Microbiome engineering. Scientific American, 312(3):S10, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
S10.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-S10.pdf.

Schmidt:2015:MHT

[2308] Charles Schmidt. Mental health: Thinking from the gut. Scientific American, 312(3):S12–S15, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
S12.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-S12.pdf.

Schnorr:2015:DMH

[2309] Stephanie L. Schnorr. The diverse microbiome of the hunter– gatherer. Scientific American, 312(3):S14–S15, March 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n3/full/scientificamerican0315-S14.html;   http://www.nature.
com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
S14.pdf.

Anonymous:2015:SAU

[2310] Anonymous. Science agenda: An unhealthy work ethic. Scientific American, 312(3):6, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
6.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-6.pdf.

Rauscher:2015:FMT

[2311] Karl Frederick Rauscher. Forum: A matter of trust. Scientific American, 312(3):8, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
8.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-8.pdf.

Lloyd:2015:ACG

[2312] Robin Lloyd. Advances: Cancer goes viral. Scientific American, 312(3):10–12, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
10.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-10.pdf.

Nordrum:2015:PC

[2313] Amy Nordrum. Poison control. Scientific American, 312(3):12, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-12a.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
12a.pdf.

Bourzac:2015:RG

[2314] Katherine Bourzac. Rubbery glass. Scientific American, 312(3):12, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-12b.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
12b.pdf.

Goldman:2015:BB

[2315] Jason G. Goldman. The bird with a breadbox. Scientific American, 312(3):14–15, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
14.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-14.pdf.

Biello:2015:NL

[2316] David Biello. Nuclear letdown. Scientific American, 312(3):16, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-16.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
16.pdf.

Williams:2015:DD

[2317] Amber Williams. Din for dinner. Scientific American, 312(3):17, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-17.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
17.pdf.

Croswell:2015:CL

[2318] Ken Croswell. Circle of life. Scientific American, 312(3):18–19, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-18.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
18.pdf.

Nuwer:2015:HPS

[2319] Rachel Nuwer. Household pest sees the light. Scientific American, 312(3): 20, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-20.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
20.pdf.

Anonymous:2015:QHc

[2320] Anonymous. Quick hits. Scientific American, 312(3):21, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n3/full/scientificamerican0315-21.html;   http://www.nature.
com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
21.pdf.

Nordrum:2015:NVN

[2321] Amy Nordrum. Nature vs. nurture vs. NASA. Scientific American, 312(3):22–23, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
22.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-22.pdf.

Moskowitz:2015:OM

[2322] Clara Moskowitz. Ode to the messenger. Scientific American, 312(3):24, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-24a.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
24a.pdf.

Wallis:2015:SHC

[2323] Claudia Wallis. The science of health: The coming revolution in knee repair. Scientific American, 312(3):25–26, March 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n3/full/scientificamerican0315-25.html;   http://www.nature.
com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
25.pdf.

Pogue:2015:TMD

[2324] David Pogue. TechnoFiles: Is e-mail dead? Scientific American, 312(3): 27, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-27.html;   http:

//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
27.pdf.

Anonymous:2015:RRBa

[2325] Anonymous. Reviews and recommendations: Beneath the surface. Scientific American, 312(3):72, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
72a.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-72a.pdf.

Anonymous:2015:RLW

[2326] Anonymous. Rust: The longest war. Scientific American, 312(3):72, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-72b.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
72b.pdf.

Anonymous:2015:NHI

[2327] Anonymous. Neuroscience: A historical introduction. Scientific American, 312(3):72, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
72c.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-72c.pdf.

Anonymous:2015:JD

[2328] Anonymous. Junk DNA. Scientific American, 312(3):72, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n3/full/scientificamerican0315-72d.html;   http://www.nature.
com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
72d.pdf.

Shermer:2015:SFD

[2329] Michael Shermer. Skeptic: Forging doubt. Scientific American, 312(3): 74, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-74.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
74.pdf.

Mirsky:2015:AGTa

[2330] Steve Mirsky. Anti gravity: Translation frustration. Scientific American, 312(3):76, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
76.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-76.pdf.

Fischetti:2015:GSD

[2331] Mark Fischetti. Graphic science: Dust up. Scientific American, 312(3): 78, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-78.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
78.pdf.

Tracey:2015:MES

[2332] Kevin J. Tracey. Medical electronics: Shock medicine. Scientific American, 312(3):28–35, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
28.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-28.pdf.

Jewitt:2015:PSO

[2333] David Jewitt and Edward D. Young. Planetary science: Oceans from the skies. Scientific American, 312(3):36–43, March 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n3/full/scientificamerican0315-36.html;   http://www.nature.
com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
36.pdf.

Cowen:2015:TSB

[2334] Ron Cowen. Technology: Sound bytes. Scientific American, 312(3):44– 47, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-44.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
44.pdf.

Branswell:2015:PHE

[2335] Helen Branswell. Public health: Ebola war. Scientific American, 312(3):48–55, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
48.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-48.pdf.

Donner:2015:CCF

[2336] Simon D. Donner. Climate change: Fantasy Island. Scientific American, 312(3):56–63, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n3/full/scientificamerican0315-
56.html;   http://www.nature.com/scientificamerican/journal/
v312/n3/pdf/scientificamerican0315-56.pdf.

Dennett:2015:SSO

[2337] Daniel C. Dennett and Deb Roy. Science and society: Our transparent future. Scientific American, 312(3):64–69, March 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n3/full/scientificamerican0315-64.html;   http://www.nature.
com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
64.pdf.

Rezzonico:2015:NRN

[2338] Enea Rezzonico, Annick Mercenier, Ed Baetge, Scott Parkinson, Thomas Beck, Johannes le Coutre, and Harald Brüssow. Nestlé’s research on nutrition and the human gut microbiome. Scientific American, 312(3):79– 85, March 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n3/full/scientificamerican0315-79.html;   http:
//www.nature.com/scientificamerican/journal/v312/n3/pdf/scientificamerican0315-
79.pdf.

DiChristina:2015:EDS

[2339] Mariette DiChristina. From the Editor: Dark and stormy. Scientific American, 312(4):4, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
4.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-4.pdf.

Anonymous:2015:Ld

[2340] Anonymous. Letters. Scientific American, 312(4):6–8, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n4/full/scientificamerican0415-6.html;   http://www.nature.
com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
6.pdf.

Schlenoff:2015:YAId

[2341] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 312(4):82, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-82.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
82.pdf.

Anonymous:2015:SAB

[2342] Anonymous. Science agenda: Ban the tan. Scientific American, 312(4): 10, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-10.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
10.pdf.

Shilling:2015:FCE

[2343] Russell Shilling. Forum: The case for education Moon shots. Scientific American, 312(4):12, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
12.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-12.pdf.

Jacobson:2015:APM

[2344] Roni Jacobson. Advances: The persistence of memory. Scientific American, 312(4):14–16, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
14.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-14.pdf.

Iozzio:2015:DTI

[2345] Corinne Iozzio. Driverless tech inches ahead. Scientific American, 312(4): 16, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-16.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
16.pdf.

Alfano:2015:ISR

[2346] Andrea Alfano. Itch, scratch, repeat. Scientific American, 312(4):17, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-17.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
17.pdf.

Lufkin:2015:WC

[2347] Bryan Lufkin. Out-of-this-world coffee. Scientific American, 312(4):18, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-18.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
18.pdf.

Lewin:2015:SR

[2348] Sarah Lewin. Singing in the rain. Scientific American, 312(4):19, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-19.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
19.pdf.

Goldman:2015:WFW

[2349] Jason G. Goldman. Why flowers wear shades. Scientific American, 312(4):20–21, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
20.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-20.pdf.

Croswell:2015:GG

[2350] Ken Croswell. Galactic ghosts. Scientific American, 312(4):22–23, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-22.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
22.pdf.

Anonymous:2015:QHd

[2351] Anonymous. Quick hits. Scientific American, 312(4):24, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n4/full/scientificamerican0415-24.html;   http://www.nature.
com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
24.pdf.

Smith:2015:DD

[2352] Peter Andrey Smith. Dusty dozing. Scientific American, 312(4):26, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-26.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
26.pdf.

Moskowitz:2015:ASA

[2353] Clara Moskowitz. Atom smasher amps up. Scientific American, 312(4): 30, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-30a.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
30a.pdf.

Shell:2015:SHA

[2354] Ellen Ruppel Shell. The science of health: Artificial sweeteners get a gut check. Scientific American, 312(4):32–34, April 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n4/full/scientificamerican0415-32.html;   http://www.nature.
com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
32.pdf.

Pogue:2015:TUT

[2355] David Pogue. TechnoFiles: Unclear for takeoff. Scientific American, 312(4):35, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-

35.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-35.pdf.

Anonymous:2015:RRD

[2356] Anonymous. Reviews and recommendations: Dragonflies. Scientific American, 312(4):76, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
76a.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-76a.pdf.

Anonymous:2015:SB

[2357] Anonymous. Scientific Babel. Scientific American, 312(4):76, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n4/full/scientificamerican0415-76b.html;   http://www.nature.
com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
76b.pdf.

Anonymous:2015:YBP

[2358] Anonymous. So you’ve been publicly shamed. Scientific American, 312(4):76, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
76c.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-76c.pdf.

Anonymous:2015:Ib

[2359] Anonymous. Infested. Scientific American, 312(4):76, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n4/full/scientificamerican0415-76d.html;   http://www.nature.
com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
76d.pdf.

Shermer:2015:SPDa

[2360] Michael Shermer. Skeptic: Paleo diets, GMOs and food taboos. Scientific American, 312(4):78, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
78.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-78.pdf.

Mirsky:2015:AGIb

[2361] Steve Mirsky. Anti gravity: Immune reaction. Scientific American, 312(4):80, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
80.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-80.pdf.

Lewin:2015:GSG

[2362] Sarah Lewin. Graphic science: A genetic guide to birds. Scientific American, 312(4):84, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
84.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-84.pdf.

Polchinski:2015:PBR

[2363] Joseph Polchinski. Physics: Burning rings of fire. Scientific American, 312(4):36–41, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
36.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-36.pdf.

Fischman:2015:SRF

[2364] Josh Fischman. Special report: Future of medicine 2015: Small wonders. Scientific American, 312(4):42–43, April 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n4/full/scientificamerican0415-42.html;   http://www.nature.
com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
42.pdf.

Maron:2015:CDH

[2365] Dina Fine Maron. Cancer drugs hit their mark. Scientific American, 312(4):44–46, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
44.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-44.pdf.

Krisch:2015:FC

[2366] Joshua A. Krisch. A flare for cancer. Scientific American, 312(4):46, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-46.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
46.pdf.

Peplow:2015:SB

[2367] Mark Peplow. A smarter bandage. Scientific American, 312(4):47–49, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-47.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
47.pdf.

Krisch:2015:GH

[2368] Joshua A. Krisch. Gentle on the heart. Scientific American, 312(4):49, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n4/full/scientificamerican0415-49.html;   http:
//www.nature.com/scientificamerican/journal/v312/n4/pdf/scientificamerican0415-
49.pdf.

Greenemeier:2015:LN

[2369] Larry Greenemeier. Launch the nanobots! Scientific American, 312(4):50–51, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
50.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-50.pdf.

Vance:2015:SFB

[2370] Erik Vance. Sustainability: Fishing for billions. Scientific American, 312(4):52–59, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
52.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-52.pdf.

Baumeister:2015:PCY

[2371] Roy F. Baumeister. Psychology: Conquer yourself, conquer the world. Scientific American, 312(4):60–65, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v312/n4/full/scientificamerican0415-
60.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-60.pdf.

Elazari:2015:SHS

[2372] Keren Elazari. Security: How to survive cyberwar. Scientific American, 312(4):66–69, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
66.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-66.pdf.

Meyer:2015:LSE

[2373] Axel Meyer. Life science: Extreme evolution. Scientific American, 312(4):70–75, April 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n4/full/scientificamerican0415-
70.html;   http://www.nature.com/scientificamerican/journal/
v312/n4/pdf/scientificamerican0415-70.pdf.

DiChristina:2015:ERT

[2374] Mariette DiChristina. From the Editor: Rise of the tyrants. Scientific American, 312(5):4, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
4.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-4.pdf.

Anonymous:2015:Le

[2375] Anonymous. Letters. Scientific American, 312(5):6–8, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n5/full/scientificamerican0515-6.html;   http://www.nature.
com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
6.pdf.

Schlenoff:2015:YAIe

[2376] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 312(5): 88, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-88.html;   http:

//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
88.pdf.

Anonymous:2015:SAWa

[2377] Anonymous. Science agenda: Wooing the fence-sitters. Scientific American, 312(5):10, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
10.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-10.pdf.

Kane:2015:FTG

[2378] Van R. Kane. Forum: Time to go to Europa. Scientific American, 312(5): 12, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-12.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
12.pdf.

Wapner:2015:ATD

[2379] Jessica Wapner. Advances: Teen dispirit. Scientific American, 312(5): 15–16, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-15.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
15.pdf.

Harrington:2015:SS

[2380] Rebecca Harrington. Switching sides. Scientific American, 312(5):16, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-16.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
16.pdf.

Anonymous:2015:QHe

[2381] Anonymous. Quick hits. Scientific American, 312(5):17, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n5/full/scientificamerican0515-17.html;   http://www.nature.
com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
17.pdf.

Loewenberg:2015:GRM

[2382] Sam Loewenberg. Global role model. Scientific American, 312(5):18– 25, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-18.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
18.pdf.

Sneed:2015:MFB

[2383] Annie Sneed. May flowers bring leaf showers. Scientific American, 312 (5):21, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-21a.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
21a.pdf.

Lufkin:2015:RB

[2384] Bryan Lufkin. The Retiarius bot. Scientific American, 312(5):21, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-21b.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
21b.pdf.

Nuwer:2015:PP

[2385] Rachel Nuwer. Popcorn physics 101. Scientific American, 312(5):22, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-22a.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
22a.pdf.

Patel:2015:LD

[2386] Prachi Patel. The lens descends. Scientific American, 312(5):22, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-22b.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
22b.pdf.

Goldman:2015:TT

[2387] Jason G. Goldman. Tiny toilets. Scientific American, 312(5):23, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-23.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
23.pdf.

Billings:2015:SF

[2388] Lee Billings. Small fry. Scientific American, 312(5):24, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n5/full/scientificamerican0515-24.html;   http://www.nature.
com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
24.pdf.

Lewin:2015:CG

[2389] Sarah Lewin. The culture of germination. Scientific American, 312(5): 26, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-26.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
26.pdf.

Lewin:2015:RS

[2390] Sarah Lewin. Right on schedule. Scientific American, 312(5):27, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-27.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
27.pdf.

Maron:2015:SHWb

[2391] Dina Fine Maron. The science of health: Why girls are starting puberty early. Scientific American, 312(5):28–30, May 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n5/full/scientificamerican0515-28.html;   http://www.nature.
com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
28.pdf.

Pogue:2015:TTD

[2392] David Pogue. TechnoFiles: Truth in digital advertising. Scientific American, 312(5):32, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-

32.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-32.pdf.

Anonymous:2015:RRBb

[2393] Anonymous. Reviews and recommendations: Biophilia. Scientific American, 312(5):82, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
82a.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-82a.pdf.

Anonymous:2015:BT

[2394] Anonymous. Birth of a theorem. Scientific American, 312(5):82, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-82b.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
82b.pdf.

Anonymous:2015:Ic

[2395] Anonymous. Invisible. Scientific American, 312(5):82, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n5/full/scientificamerican0515-82c.html;   http://www.nature.
com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
82c.pdf.

Anonymous:2015:HCM

[2396] Anonymous. How to clone a mammoth. Scientific American, 312(5): 82, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-82d.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
82d.pdf.

Shermer:2015:STS

[2397] Michael Shermer. Skeptic: Terrorism as self-help justice. Scientific American, 312(5):84, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
84.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-84.pdf.

Mirsky:2015:AGS

[2398] Steve Mirsky. Anti gravity: Have a seat. Scientific American, 312(5): 86, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n5/full/scientificamerican0515-86.html;   http:
//www.nature.com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
86.pdf.

Fischetti:2015:GSL

[2399] Mark Fischetti. Graphic science: This land is my land. Scientific American, 312(5):90, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
90.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-90.pdf.

Brusatte:2015:PRT

[2400] Stephen Brusatte. Paleontology: Rise of the tyrannosaurs. Scientific American, 312(5):34–41, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
34.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-34.pdf.

Ent:2015:PPG

[2401] Rolf Ent, Thomas Ullrich, and Raju Venugopalan. Particle physics: The glue that binds us. Scientific American, 312(5):42–49, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n5/full/scientificamerican0515-42.html;   http://www.nature.
com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
42.pdf.

Stix:2015:MLC

[2402] Gary Stix. Medicine: Lifting the curse of Alzheimer’s. Scientific American, 312(5):50–57, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
50.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-50.pdf.

Pavlus:2015:CSN

[2403] John Pavlus. Computing: The search for a new machine. Scientific American, 312(5):58–63, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
58.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-58.pdf.

Harris:2015:OWD

[2404] Mark Harris. Oceanography: Waves of destruction. Scientific American, 312(5):64–69, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
64.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-64.pdf.

Laird:2015:BCS

[2405] Dale W. Laird, Paul D. Lampe, and Ross G. Johnson. Biology: Cellular small talk. Scientific American, 312(5):70–77, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n5/full/scientificamerican0515-70.html;   http://www.nature.
com/scientificamerican/journal/v312/n5/pdf/scientificamerican0515-
70.pdf.

Miller:2015:SIN

[2406] Steven D. Miller. Satellite imagery: Night watch. Scientific American, 312(5):78–81, May 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n5/full/scientificamerican0515-
78.html;   http://www.nature.com/scientificamerican/journal/
v312/n5/pdf/scientificamerican0515-78.pdf.

DiChristina:2015:EPP

[2407] Mariette DiChristina. From the Editor: Promises and perils. Scientific American, 312(6):4, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
4.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-4.pdf.

Anonymous:2015:Lf

[2408] Anonymous. Letters. Scientific American, 312(6):6–8, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n6/full/scientificamerican0615-6.html;   http://www.nature.
com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
6.pdf.

Schlenoff:2015:YAIf

[2409] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 312(6):82, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-82.html;   http:
//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
82.pdf.

Anonymous:2015:SAL

[2410] Anonymous. Science agenda: A last right for dying patients. Scientific American, 312(6):10, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
10.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-10.pdf.

King:2015:FEC

[2411] Barbara J. King. Forum: Ending a cruel legacy. Scientific American, 312(6):12, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
12.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-12.pdf.

Moskowitz:2015:ADM

[2412] Clara Moskowitz. Advances: Dark matter drops a clue. Scientific American, 312(6):15–17, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
15.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-15.pdf.

Anonymous:2015:N

[2413] Anonymous. By the numbers. Scientific American, 312(6):17, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n6/full/scientificamerican0615-17.html;   http://www.nature.
com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
17.pdf.

Lewin:2015:PS

[2414] Sarah Lewin. Prehistoric swagger. Scientific American, 312(6):18, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-18.html;   http:
//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
18.pdf.

Anonymous:2015:QHf

[2415] Anonymous. Quick hits. Scientific American, 312(6):19, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n6/full/scientificamerican0615-19.html;   http://www.nature.
com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
19.pdf.

Jabr:2015:DWS

[2416] Ferris Jabr. The doctor will see you now. Scientific American, 312(6): 20, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-20.html;   http:
//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
20.pdf.

Goldman:2015:BBB

[2417] Jason G. Goldman. From babbling bats to baby talk. Scientific American, 312(6):21, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
21.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-21.pdf.

Courage:2015:OPV

[2418] Katherine Harmon Courage. Our personal vaccine helpers. Scientific American, 312(6):22, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
22.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-22.pdf.

Billings:2015:FSM

[2419] Lee Billings. Flying saucers for Mars. Scientific American, 312(6):23, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-23.html;   http:
//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
23.pdf.

Schiffman:2015:RFT

[2420] Richard Schiffman. Rain-forest threats resume. Scientific American, 312(6):24, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
24.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-24.pdf.

Alfano:2015:KKB

[2421] Andrea Alfano. Kiddo knows best. Scientific American, 312(6):25, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-25.html;   http:
//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
25.pdf.

Biba:2015:ATT

[2422] Erin Biba. An apocalypse think tank. Scientific American, 312(6):26, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-26.html;   http:
//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
26.pdf.

Noonan:2015:SHS

[2423] David Noonan. The science of health: The not so silent epidemic. Scientific American, 312(6):27–28, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-

27.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-27.pdf.

Pogue:2015:TUG

[2424] David Pogue. TechnoFiles: The upgrade game. Scientific American, 312(6):29, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
29.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-29.pdf.

Anonymous:2015:RRBc

[2425] Anonymous. Reviews and recommendations: Black hole: How an idea abandoned by Newtonians, hated by Einstein, and gambled on by Hawking became loved. Scientific American, 312(6):78, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n6/full/scientificamerican0615-78a.html;   http://www.nature.
com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
78a.pdf.

Anonymous:2015:HB

[2426] Anonymous. How to bake. Scientific American, 312(6):78, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n6/full/scientificamerican0615-78b.html;   http://www.nature.
com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
78b.pdf.

Anonymous:2015:SCR

[2427] Anonymous. The strange case of the rickety Cossack: And other cautionary tales from human evolution. Scientific American, 312(6):78, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-78c.html;   http:
//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
78c.pdf.

Anonymous:2015:AG

[2428] Anonymous. American genius. Scientific American, 312(6):78, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-78d.html;   http:

//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
78d.pdf.

Shermer:2015:SSH

[2429] Michael Shermer. Skeptic: Scientia humanitatis. Scientific American, 312(6):80, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
80.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-80.pdf.

Mirsky:2015:AGGa

[2430] Steve Mirsky. Anti gravity: Gold flush. Scientific American, 312(6):81, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v312/n6/full/scientificamerican0615-81.html;   http:
//www.nature.com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
81.pdf.

Maron:2015:GST

[2431] Dina Fine Maron. Graphic science: Twists of fate. Scientific American, 312(6):84, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
84.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-84.pdf.

Giedd:2015:NAT

[2432] Jay N. Giedd. Neuroscience: The amazing teen brain. Scientific American, 312(6):32–37, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
32.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-32.pdf.

Dominguez:2015:AAL

[2433] Alberto Domínguez, Joel R. Primack, and Trudy E. Bell. Astronomy: All the light there ever was. Scientific American, 312(6):38–43, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n6/full/scientificamerican0615-38.html;   http://www.nature.
com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
38.pdf.

Mehal:2015:MCF

[2434] Wajahat Z. Mehal. Medicine: Cells on fire. Scientific American, 312(6):44–49, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
44.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-44.pdf.

Dugatkin:2015:ENA

[2435] Lee Alan Dugatkin and Matthew Hasenjager. Ethology: The networked animal. Scientific American, 312(6):50–55, June 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v312/
n6/full/scientificamerican0615-50.html;   http://www.nature.
com/scientificamerican/journal/v312/n6/pdf/scientificamerican0615-
50.pdf.

Freedman:2015:TBR

[2436] David H. Freedman. Technology: Birth of a rocket. Scientific American, 312(6):56–65, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
56.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-56.pdf.

Rose:2015:FSM

[2437] Ananda Rose. Forensic science: The mystery of Case 0425. Scientific American, 312(6):66–71, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
66.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-66.pdf.

ONeill:2015:PHD

[2438] Scott O’Neill. Public health: The dengue stopper. Scientific American, 312(6):72–77, June 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v312/n6/full/scientificamerican0615-
72.html;   http://www.nature.com/scientificamerican/journal/
v312/n6/pdf/scientificamerican0615-72.pdf.

DiChristina:2015:ESU

[2439] Mariette DiChristina. From the Editor: The shadow universe. Scientific American, 313(1):4, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
4.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-4.pdf.

Anonymous:2015:Lg

[2440] Anonymous. Letters. Scientific American, 313(1):6–8, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-6.html;   http://www.nature.
com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
6.pdf.

Schlenoff:2015:YAIg

[2441] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 313(1): 79, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-79.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
79.pdf.

Anonymous:2015:SAWb

[2442] Anonymous. Science agenda: Why embryos should not be off-limits. Scientific American, 313(1):10, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
10.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-10.pdf.

West:2015:FSW

[2443] Michael West. Forum: Star Wars. Scientific American, 313(1):12, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-12.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
12.pdf.

Billings:2015:AGB

[2444] Lee Billings. Advances: Go big or go home. Scientific American, 313(1): 14–16, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-14.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
14.pdf.

Lewin:2015:DD

[2445] Sarah Lewin. Dams over the decades. Scientific American, 313(1): 16, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-16.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
16.pdf.

Lewin:2015:SC

[2446] Sarah Lewin. Sound check. Scientific American, 313(1):17, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-17.html;   http://www.nature.
com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
17.pdf.

Emspak:2015:HH

[2447] Jesse Emspak. Hacking heats up. Scientific American, 313(1):18–19, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-18.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
18.pdf.

Haelle:2015:TS

[2448] Tara Haelle. Transfusion solution. Scientific American, 313(1):20, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-20.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
20.pdf.

Kerr:2015:FF

[2449] Millie Kerr. Facebook for the ferocious. Scientific American, 313(1): 21, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-21.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
21.pdf.

Palucka:2015:BB

[2450] Tim Palucka. Blitzkrieg basics. Scientific American, 313(1):22, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-22.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
22.pdf.

Anonymous:2015:QHg

[2451] Anonymous. Quick hits. Scientific American, 313(1):23, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-23.html;   http://www.nature.
com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
23.pdf.

Jabr:2015:FI

[2452] Ferris Jabr. First impressions. Scientific American, 313(1):24, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-24.html;   http://www.nature.
com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
24.pdf.

Bourzac:2015:BB

[2453] Katherine Bourzac. Build-a-battery. Scientific American, 313(1):25, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-25.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
25.pdf.

Anonymous:2015:PPa

[2454] Anonymous. Poised for Pluto. Scientific American, 313(1):26, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-26.html;   http://www.nature.

com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
26.pdf.

Weintraub:2015:SHC

[2455] Karen Weintraub. The science of health: Can we stop aging? Scientific American, 313(1):28–29, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
28.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-28.pdf.

Pogue:2015:TIS

[2456] David Pogue. TechnoFiles: iResearch subject. Scientific American, 313 (1):30, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-30.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
30.pdf.

Anonymous:2015:RRS

[2457] Anonymous. Reviews and recommendations: Seveneves. Scientific American, 313(1):76, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
76a.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-76a.pdf.

Anonymous:2015:OWW

[2458] Anonymous. The only words that are worth remembering. Scientific American, 313(1):76, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
76b.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-76b.pdf.

Anonymous:2015:A

[2459] Anonymous. Aurora. Scientific American, 313(1):76, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-76c.html;   http://www.nature.
com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
76c.pdf.

Anonymous:2015:Pb

[2460] Anonymous. Persona. Scientific American, 313(1):76, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-76d.html;   http://www.nature.
com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
76d.pdf.

Shermer:2015:SO

[2461] Michael Shermer. Skeptic: Outrageous. Scientific American, 313(1): 77, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-77.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
77.pdf.

Mirsky:2015:AGL

[2462] Steve Mirsky. Anti gravity: Legionnaires’ unease. Scientific American, 313(1):78, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
78.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-78.pdf.

Fischetti:2015:GST

[2463] Mark Fischetti. Graphic science: True size of Africa. Scientific American, 313(1):80, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
80.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-80.pdf.

Dobrescu:2015:AMH

[2464] Bogdan A. Dobrescu and Don Lincoln. Astrophysics: Mystery of the hidden cosmos. Scientific American, 313(1):32–39, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-32.html;   http://www.nature.
com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
32.pdf.

Mattson:2015:BHW

[2465] Mark P. Mattson. Brain health: What doesn’t kill you …. Scientific American, 313(1):40–45, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
40.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-40.pdf.

Fox:2015:BLH

[2466] Douglas Fox. Biology: Life at Hell’s gate. Scientific American, 313(1): 46–53, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n1/full/scientificamerican0715-46.html;   http:
//www.nature.com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
46.pdf.

Sivaram:2015:EOS

[2467] Varun Sivaram, Samuel D. Stranks, and Henry J. Snaith. Energy: Outshining silicon. Scientific American, 313(1):54–59, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n1/full/scientificamerican0715-54.html;   http://www.nature.
com/scientificamerican/journal/v313/n1/pdf/scientificamerican0715-
54.pdf.

Morell:2015:EWD

[2468] Virginia Morell. Evolution: From wolf to dog. Scientific American, 313(1):60–67, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
60.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-60.pdf.

Ornes:2015:MWU

[2469] Stephen Ornes. Mathematics: The whole universe catalog. Scientific American, 313(1):68–75, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
68.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-68.pdf.

Ornes:2015:GTM

[2470] Stephen Ornes. Group theory: The math of making connections. Scientific American, 313(1):74, July 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n1/full/scientificamerican0715-
74.html;   http://www.nature.com/scientificamerican/journal/
v313/n1/pdf/scientificamerican0715-74.pdf.

DiChristina:2015:EH

[2471] Mariette DiChristina. From the Editor: Happy 170th! Scientific American, 313(2):4, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
4.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-4.pdf.

Anonymous:2015:Lh

[2472] Anonymous. Letters. Scientific American, 313(2):6–8, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-6.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
6.pdf.

Schlenoff:2015:YAIh

[2473] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 313(2):86, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n2/full/scientificamerican0815-86.html;   http:
//www.nature.com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
86.pdf.

Anonymous:2015:SADb

[2474] Anonymous. Science agenda: Docs, glocks and stray bullets. Scientific American, 313(2):10, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
10.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-10.pdf.

Billings:2015:FSC

[2475] Linda Billings. Forum: Space cowboys. Scientific American, 313(2):12, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n2/full/scientificamerican0815-12.html;   http:
//www.nature.com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
12.pdf.

Mukerjee:2015:AID

[2476] Madhusree Mukerjee. Advances: The impending dam disaster in the Himalayan mountains. Scientific American, 313(2):14–16, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-14.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
14.pdf.

Iozzio:2015:DYP

[2477] Corinne Iozzio. The doctor is in (your phone). Scientific American, 313(2):18, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
18.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-18.pdf.

Billings:2015:EBG

[2478] Lee Billings. Europa’s “brown gunk” suggests a briny sea. Scientific American, 313(2):19, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
19.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-19.pdf.

Moskowitz:2015:RB

[2479] Clara Moskowitz. Remembering the blast. Scientific American, 313(2): 20, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n2/full/scientificamerican0815-20.html;   http:
//www.nature.com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
20.pdf.

Lewin:2015:RWM

[2480] Sarah Lewin. The room with the weakest magnetic field on Earth. Scientific American, 313(2):21, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
21.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-21.pdf.

Goldman:2015:CYH

[2481] Jason G. Goldman. Can you hear me now? Scientific American, 313(2): 22, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n2/full/scientificamerican0815-22.html;   http:
//www.nature.com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
22.pdf.

Smith:2015:WSA

[2482] Peter Andrey Smith. Watch out, Sam Adams. Scientific American, 313(2):23, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
23.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-23.pdf.

Weiner:2015:GFF

[2483] Debra Weiner. The great flower fabrication. Scientific American, 313(2): 24, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n2/full/scientificamerican0815-24.html;   http:
//www.nature.com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
24.pdf.

Anonymous:2015:QHh

[2484] Anonymous. Quick hits. Scientific American, 313(2):25, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-25.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
25.pdf.

Long:2015:WME

[2485] Kat Long. What meets the eye. Scientific American, 313(2):26, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n2/full/scientificamerican0815-26.html;   http:
//www.nature.com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
26.pdf.

Noonan:2015:SHT

[2486] David Noonan. The science of health: A turn for the worse. Scientific American, 313(2):27–28, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
27.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-27.pdf.

Pogue:2015:TTS

[2487] David Pogue. TechnoFiles: The touch screen generation. Scientific American, 313(2):29, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
29.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-29.pdf.

Anonymous:2015:RRBd

[2488] Anonymous. Reviews and recommendations: Book review: Spurious correlations. Scientific American, 313(2):82, August 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-82a.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
82a.pdf.

Anonymous:2015:BRL

[2489] Anonymous. Book review: Life’s greatest secret: Life’s Greatest Secret: The Race to Crack the Genetic Code, by Matthew Cobb, Basic Books: 2015 ($29.99). Scientific American, 313(2):82, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-82b.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
82b.pdf.

Anonymous:2015:BRM

[2490] Anonymous. Book review: Modern romance: Modern Romance, by Aziz Ansari, with and Eric Klinenberg, Penguin Press: 2015 ($28.95). Scientific American, 313(2):82, August 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-82c.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
82c.pdf.

Anonymous:2015:BRG

[2491] Anonymous. Book review: Genius at play: Genius at Play: The Curious Mind of John Horton Conway, by Siobhan Roberts, Bloomsbury: 2015 ($30). Scientific American, 313(2):82, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-82d.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
82d.pdf.

Shermer:2015:SML

[2492] Michael Shermer. Skeptic: The meaning of life in a formula. Scientific American, 313(2):83, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
83.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-83.pdf.

Mirsky:2015:AGC

[2493] Steve Mirsky. Anti gravity: The cheese eyes have it. Scientific American, 313(2):84, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
84.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-84.pdf.

Harrington:2015:GSD

[2494] Rebecca Harrington. Graphic science: Drug-resistant stomach bug. Scientific American, 313(2):88, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v313/n2/full/scientificamerican0815-
88.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-88.pdf.

Marean:2015:EMI

[2495] Curtis W. Marean. Evolution: The most invasive species of all. Scientific American, 313(2):32–39, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
32.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-32.pdf.

Billings:2015:SSS

[2496] Lee Billings. Space science: In search of alien Jupiters. Scientific American, 313(2):40–47, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
40.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-40.pdf.

Liberman:2015:NHH

[2497] M. Charles Liberman. Neuroscience: Hidden hearing loss. Scientific American, 313(2):48–53, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
48.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-48.pdf.

Liberman:2015:FSS

[2498] M. Charles Liberman. A few simple steps: How to protect your hearing. Scientific American, 313(2):52, August 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-52.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
52.pdf.

Paul:2015:SRB

[2499] Annie Murphy Paul. Special report: Building the 21st-century learner: A new vision for testing. Scientific American, 313(2):54–61, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n2/full/scientificamerican0815-54.html;   http://www.nature.
com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
54.pdf.

Tyre:2015:TTP

[2500] Peg Tyre. Testing the team player. Scientific American, 313(2):58–59, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n2/full/scientificamerican0815-58.html;   http:
//www.nature.com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
58.pdf.

Wysession:2015:KST

[2501] Michael Wysession. Kids are scientists, too. Scientific American, 313(2):62–63, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
62.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-62.pdf.

Baum:2015:CCS

[2502] Dan Baum. Climate: Change of state. Scientific American, 313(2):64– 71, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n2/full/scientificamerican0815-64.html;   http:
//www.nature.com/scientificamerican/journal/v313/n2/pdf/scientificamerican0815-
64.pdf.

Hidalgo:2015:TEP

[2503] César A. Hidalgo. Thought experiment: Planet hard drive. Scientific American, 313(2):72–75, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
72.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-72.pdf.

Rosner:2015:SFM

[2504] Hillary Rosner. Sustainability: Forests on the march. Scientific American, 313(2):76–81, August 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n2/full/scientificamerican0815-
76.html;   http://www.nature.com/scientificamerican/journal/
v313/n2/pdf/scientificamerican0815-76.pdf.

DiChristina:2015:EEE

[2505] Mariette DiChristina. From the Editor: Editorial and Einstein. Scientific American, 313(3):6, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
6.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-6.pdf.

Anonymous:2015:Li

[2506] Anonymous. Letters. Scientific American, 313(3):8–10, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-8.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
8.pdf.

Schlenoff:2015:YAIi

[2507] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in scientific American: 50, 100 & 150 years ago. Scientific American, 313(3):98, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-98.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
98.pdf.

Anonymous:2015:SADc

[2508] Anonymous. Science agenda: Don’t blind NASA to Earth’s climate. Scientific American, 313(3):12, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-12.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
12.pdf.

Goodwin:2015:FNN

[2509] Daniel Goodwin. Forum: Neuroscience needs hackers. Scientific American, 313(3):14, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
14.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-14.pdf.

Moyer:2015:ALL

[2510] Melinda Wenner Moyer. Advances: Lingering Lyme. Scientific American, 313(3):16–18, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
16.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-16.pdf.

Long:2015:FF

[2511] Kat Long. Fibs with friends. Scientific American, 313(3):19, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-19.html;   http:
//www.nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
19.pdf.

Hsu:2015:CCD

[2512] Jeremy Hsu. Cloudy with a chance of drones. Scientific American, 313(3):20, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
20.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-20.pdf.

Imbler:2015:SS

[2513] Sabrina Imbler. Sunken strawberries. Scientific American, 313(3):21, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-21.html;   http:
//www.nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
21.pdf.

Nuwer:2015:SCN

[2514] Rachel Nuwer. SAT charts new territory. Scientific American, 313(3):22, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-22.html;   http:
//www.nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
22.pdf.

Croswell:2015:SS

[2515] Ken Croswell. Solar swindle. Scientific American, 313(3):23, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-23.html;   http:

//www.nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
23.pdf.

Goldman:2015:PAF

[2516] Jason G. Goldman. From the pages of Aesop’s fables. Scientific American, 313(3):24, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
24.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-24.pdf.

Anonymous:2015:QHi

[2517] Anonymous. Quick hits. Scientific American, 313(3):25, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-25.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
25.pdf.

Temming:2015:WWE

[2518] Maria Temming. Water, water everywhere. Scientific American, 313(3):26, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
26.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-26.pdf.

Long:2015:OMS

[2519] Kat Long. Old mold sold. Scientific American, 313(3):27, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-27.html;   http:
//www.nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
27.pdf.

Maron:2015:SHB

[2520] Dina Fine Maron. The science of health: Brain food. Scientific American, 313(3):29–30, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
29.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-29.pdf.

Pogue:2015:TSS

[2521] David Pogue. TechnoFiles: Super Siri. Scientific American, 313(3):31, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-31.html;   http:
//www.nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
31.pdf.

Anonymous:2015:RRR

[2522] Anonymous. Reviews and recommendations: The road to relativity. Scientific American, 313(3):94, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-94a.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
94a.pdf.

Anonymous:2015:EHS

[2523] Anonymous. Einstein: His space and times. Scientific American, 313(3):94, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
94b.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-94b.pdf.

Anonymous:2015:EDS

[2524] Anonymous. Einstein’s dice and Schrödinger’s cat. Scientific American, 313(3):94, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
94c.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-94c.pdf.

Anonymous:2015:PPb

[2525] Anonymous. The physicist and the philosopher. Scientific American, 313(3):94, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
94d.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-94d.pdf.

Shermer:2015:SFP

[2526] Michael Shermer. Skeptic: Forensic pseudoscience. Scientific American, 313(3):95, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
95.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-95.pdf.

Mirsky:2015:AGN

[2527] Steve Mirsky. Anti gravity: Numbers games. Scientific American, 313(3):96, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
96.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-96.pdf.

Williams:2015:GSB

[2528] Amber Williams. Graphic science: A bigger city isn’t always better. Scientific American, 313(3):100, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-100.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
100.pdf.

Anonymous:2015:IE

[2529] Anonymous. Introduction: Einstein. Scientific American, 313(3):32–34, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n3/full/scientificamerican0915-32.html;   http:
//www.nature.com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
32.pdf.

Greene:2015:EWM

[2530] Brian Greene. Essay: Why He matters. Scientific American, 313(3):34–37, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
34.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-34.pdf.

Isaacson:2015:HHE

[2531] Walter Isaacson. History: How Einstein reinvented reality. Scientific American, 313(3):38–45, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v313/n3/full/scientificamerican0915-
38.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-38.pdf.

Hossenfelder:2015:TEH

[2532] Sabine Hossenfelder. Thought experiments: Head trip. Scientific American, 313(3):46–49, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
46.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-46.pdf.

Krauss:2015:CWE

[2533] Lawrence M. Krauss. Cosmology: What Einstein got wrong. Scientific American, 313(3):50–55, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v313/n3/full/scientificamerican0915-
50.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-50.pdf.

Anonymous:2015:LNR

[2534] Anonymous. A legacy in numbers: Relativity’s reach. Scientific American, 313(3):56–59, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
56.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-56.pdf.

Powell:2015:FPC

[2535] Corey S. Powell. Fundamental physics: Cleaning up after Einstein. Scientific American, 313(3):60–67, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-60.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
60.pdf.

Folger:2015:TBH

[2536] Tim Folger. Theory: A brief history of time travel. Scientific American, 313(3):68–73, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
68.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-68.pdf.

Psaltis:2015:ABH

[2537] Dimitrios Psaltis and Sheperd S. Doeleman. Astronomy: The black hole test. Scientific American, 313(3):74–79, September 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n3/full/scientificamerican0915-74.html;   http://www.nature.
com/scientificamerican/journal/v313/n3/pdf/scientificamerican0915-
74.pdf.

Anonymous:2015:SCW

[2538] Anonymous. A study in contrasts: Who was Einstein, really? Scientific American, 313(3):80–81, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v313/n3/full/scientificamerican0915-
80.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-80.pdf.

Burrell:2015:NGJ

[2539] Brian D. Burrell. Neuroscience: Genius in a jar. Scientific American, 313(3):82–87, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n3/full/scientificamerican0915-
82.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-82.pdf.

Musser:2015:QPC

[2540] George Musser. Quantum physics: Is the Cosmos random? Scientific American, 313(3):88–93, September 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v313/n3/full/scientificamerican0915-
88.html;   http://www.nature.com/scientificamerican/journal/
v313/n3/pdf/scientificamerican0915-88.pdf.

DiChristina:2015:ESS

[2541] Mariette DiChristina. From the Editor: Sleep science. Scientific American, 313(4):4, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
4.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-4.pdf.

Anonymous:2015:Lj

[2542] Anonymous. Letters. Scientific American, 313(4):6–8, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-6.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
6.pdf.

Schlenoff:2015:YAIj

[2543] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 313(4):86, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-86.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
86.pdf.

Anonymous:2015:SAHb

[2544] Anonymous. Science agenda: How to stop teen pregnancies. Scientific American, 313(4):10, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
10.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-10.pdf.

Blomqvist:2015:FLL

[2545] Linus Blomqvist, Ted Nordhaus, and Michael Shellenberger. Forum: Living off the land. Scientific American, 313(4):12, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-12.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
12.pdf.

Billings:2015:AWC

[2546] Lee Billings. Advances: Are we on the cusp of war — in space? Scientific American, 313(4):14–18, October 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-14.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
14.pdf.

Nuwer:2015:MML

[2547] Rachel Nuwer. Mother, may I … eat you? Scientific American, 313(4): 19, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-19.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
19.pdf.

Sneed:2015:TTC

[2548] Annie Sneed. Test-tube corals. Scientific American, 313(4):20, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-20.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
20.pdf.

Long:2015:VS

[2549] Kat Long. The vicissitudes of sex. Scientific American, 313(4):21, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-21.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
21.pdf.

Gannon:2015:DD

[2550] Megan Gannon. Dentist to the dead. Scientific American, 313(4):22, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-22.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
22.pdf.

Ossola:2015:GTC

[2551] Alexandra Ossola. Graphene’s top competitor. Scientific American, 313(4):23, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
23.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-23.pdf.

Schechter:2015:USG

[2552] Erik Schechter. Unlocking safer guns. Scientific American, 313(4):24, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-24.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
24.pdf.

Anonymous:2015:QHj

[2553] Anonymous. Quick hits. Scientific American, 313(4):25, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-25.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
25.pdf.

Iozzio:2015:BB

[2554] Corinne Iozzio. Bots on a budget. Scientific American, 313(4):26–27, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-26.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
26.pdf.

Temming:2015:PJ

[2555] Maria Temming. Particular joy. Scientific American, 313(4):28, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-28.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
28.pdf.

Wapner:2015:SHD

[2556] Jessica Wapner. The science of health: Deadly drug combinations. Scientific American, 313(4):29–30, October 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-29.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
29.pdf.

Pogue:2015:TRR

[2557] David Pogue. TechnoFiles: Robots rising. Scientific American, 313(4):32, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-32.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
32.pdf.

Anonymous:2015:RRP

[2558] Anonymous. Reviews and recommendations: Polar: A photicular book. Scientific American, 313(4):80, October 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-80a.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
80a.pdf.

Anonymous:2015:MWW

[2559] Anonymous. The man who wasn’t there. Scientific American, 313(4):80, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-80b.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
80b.pdf.

Anonymous:2015:OWR

[2560] Anonymous. The only woman in the room. Scientific American, 313(4): 80, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-80c.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
80c.pdf.

Anonymous:2015:RRA

[2561] Anonymous. A river runs again. Scientific American, 313(4):80, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-80d.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
80d.pdf.

Shermer:2015:SEU

[2562] Michael Shermer. Skeptic: The electric universe acid test. Scientific American, 313(4):82, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
82.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-82.pdf.

Mirsky:2015:AGBb

[2563] Steve Mirsky. Anti gravity: Breaking bad TV science. Scientific American, 313(4):84, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
84.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-84.pdf.

Fischetti:2015:GSG

[2564] Mark Fischetti. Graphic science: A greenhouse gas surprise. Scientific American, 313(4):88, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
88.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-88.pdf.

Anonymous:2015:SRS

[2565] Anonymous. Special report: State of the World’s science 2015: Big science, big challenges. Scientific American, 313(4):34–35, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-34.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
34.pdf.

Theil:2015:TM

[2566] Stefan Theil. Trouble in mind. Scientific American, 313(4):36–42, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-36.html;   http:

//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
36.pdf.

Anonymous:2015:HBS

[2567] Anonymous. How big is science? Scientific American, 313(4):40–41, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-40.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
40.pdf.

Karlan:2015:MEL

[2568] Dean Karlan. More evidence, less poverty. Scientific American, 313(4):43–45, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
43.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-43.pdf.

Velasco:2015:AM

[2569] Rodrigo Guerrero Velasco. An antidote to murder. Scientific American, 313(4):46–50, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
46.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-46.pdf.

Anonymous:2015:WL

[2570] Anonymous. World leaders. Scientific American, 313(4):51, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n4/full/scientificamerican1015-51.html;   http:
//www.nature.com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
51.pdf.

Stickgold:2015:NSI

[2571] Robert Stickgold. Neuroscience: Sleep on it! Scientific American, 313(4):52–57, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
52.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-52.pdf.

Halzen:2015:ANE

[2572] Francis Halzen. Astrophysics: Neutrinos at the ends of the Earth. Scientific American, 313(4):58–63, October 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-58.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
58.pdf.

Johnson:2015:HFG

[2573] Richard J. Johnson and Peter Andrews. Health: The fat gene. Scientific American, 313(4):64–69, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
64.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-64.pdf.

Symons:2015:ASD

[2574] Sarah Symons and Elizabeth Tasker. Archaeoastronomy: Stars of the dead. Scientific American, 313(4):70–75, October 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n4/full/scientificamerican1015-70.html;   http://www.nature.
com/scientificamerican/journal/v313/n4/pdf/scientificamerican1015-
70.pdf.

Williams:2015:LSS

[2575] Wendy Williams. Life science: The secret lives of horses. Scientific American, 313(4):76–79, October 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n4/full/scientificamerican1015-
76.html;   http://www.nature.com/scientificamerican/journal/
v313/n4/pdf/scientificamerican1015-76.pdf.

DiChristina:2015:EPO

[2576] Mariette DiChristina. From the Editor: Pyramids, olives and young minds. Scientific American, 313(5):6, November 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n5/full/scientificamerican1115-6.html;   http://www.nature.
com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
6.pdf.

Anonymous:2015:Lk

[2577] Anonymous. Letters. Scientific American, 313(5):8–10, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n5/full/scientificamerican1115-8.html;   http://www.nature.
com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
8.pdf.

Schlenoff:2015:YAIk

[2578] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 313(5):79, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n5/full/scientificamerican1115-79.html;   http:
//www.nature.com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
79.pdf.

Anonymous:2015:SAR

[2579] Anonymous. Science agenda: A risky arms race in space. Scientific American, 313(5):12, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
12.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-12.pdf.

Pearson:2015:FRB

[2580] Rachel Pearson. Forum: Racial bias in medicine. Scientific American, 313(5):14, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
14.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-14.pdf.

Kwon:2015:AGB

[2581] Diana Kwon. Advances: The great brain drain. Scientific American, 313(5):17–18, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
17.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-17.pdf.

Long:2015:BDR

[2582] Kat Long. Both Democrats and Republicans losing ground. Scientific American, 313(5):18, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
18.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-18.pdf.

Goldman:2015:CMM

[2583] Jason G. Goldman. Over-the-counter monkey meds. Scientific American, 313(5):19, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
19.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-19.pdf.

Laursen:2015:SNP

[2584] Lucas Laursen. A sooty North Pole ahead. Scientific American, 313(5):20–21, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
20.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-20.pdf.

Anonymous:2015:QHk

[2585] Anonymous. Quick hits. Scientific American, 313(5):22, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n5/full/scientificamerican1115-22.html;   http://www.nature.
com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
22.pdf.

Maron:2015:FWR

[2586] Dina Fine Maron. Fake weed, real crisis. Scientific American, 313(5):23, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n5/full/scientificamerican1115-23.html;   http:
//www.nature.com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
23.pdf.

Hsu:2015:PP

[2587] Jeremy Hsu. A partner for plutonium. Scientific American, 313(5):24, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n5/full/scientificamerican1115-24.html;   http:
//www.nature.com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
24.pdf.

Biello:2015:NCA

[2588] David Biello. New cook arrives in life’s kitchen. Scientific American, 313(5):25, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
25.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-25.pdf.

Fischetti:2015:DSS

[2589] Mark Fischetti. Deep-sea sampler. Scientific American, 313(5):26, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n5/full/scientificamerican1115-26a.html;   http:
//www.nature.com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
26a.pdf.

Billings:2015:GB

[2590] Lee Billings. Gone without a bang. Scientific American, 313(5):26–27, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n5/full/scientificamerican1115-26b.html;   http:
//www.nature.com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
26b.pdf.

Long:2015:SF

[2591] Kat Long. Super fan. Scientific American, 313(5):27, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n5/full/scientificamerican1115-27.html;   http://www.nature.
com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
27.pdf.

Shell:2015:SHO

[2592] Ellen Ruppel Shell. The science of health: Overreaction. Scientific American, 313(5):28–29, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-

28.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-28.pdf.

Pogue:2015:THW

[2593] David Pogue. TechnoFiles: Hackers at the wheel. Scientific American, 313(5):30, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
30.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-30.pdf.

Anonymous:2015:RRBe

[2594] Anonymous. Reviews and recommendations: Book review: The best American infographics 2015. Scientific American, 313(5):74, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n5/full/scientificamerican1115-74a.html;   http:
//www.nature.com/scientificamerican/journal/v313/n5/pdf/scientificamerican1115-
74a.pdf.

Anonymous:2015:BRD

[2595] Anonymous. Book review: Dark matter and the dinosaurs. Scientific American, 313(5):74, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
74b.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-74b.pdf.

Anonymous:2015:BRSa

[2596] Anonymous. Book review: Suspicious minds. Scientific American, 313(5):74, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
74c.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-74c.pdf.

Anonymous:2015:BRB

[2597] Anonymous. Book review: Black man in a white coat. Scientific American, 313(5):74, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
74d.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-74d.pdf.

Shermer:2015:SPDb

[2598] Michael Shermer. Skeptic: Perception deception. Scientific American, 313(5):75, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
75.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-75.pdf.

Mirsky:2015:AGTb

[2599] Steve Mirsky. Anti gravity: Tushi kicks butt. Scientific American, 313(5):76, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
76.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-76.pdf.

Fischetti:2015:GSSc

[2600] Mark Fischetti. Graphic science: A southward invasion. Scientific American, 313(5):80, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
80.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-80.pdf.

Zorich:2015:APE

[2601] Zach Zorich. Archaeology: The pyramid effect. Scientific American, 313(5):32–39, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
32.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-32.pdf.

Frieman:2015:CSD

[2602] Joshua Frieman. Cosmology: Seeing in the dark. Scientific American, 313(5):40–47, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
40.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-40.pdf.

Kelley:2015:MDD

[2603] Shana O. Kelley. Medicine: Disease detector. Scientific American, 313(5):48–51, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
48.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-48.pdf.

Nadeau:2015:ABO

[2604] Barbie Latza Nadeau. Agriculture: The battle of olives. Scientific American, 313(5):52–59, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
52.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-52.pdf.

Fung:2015:TRH

[2605] Pascale Fung. Technology: Robots with heart. Scientific American, 313(5):60–63, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
60.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-60.pdf.

Kuhl:2015:CDB

[2606] Patricia K. Kuhl. Child development: Baby talk. Scientific American, 313(5):64–69, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
64.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-64.pdf.

Musser:2015:PWH

[2607] George Musser. Physics: Where is here? Scientific American, 313(5):70–73, November 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n5/full/scientificamerican1115-
70.html;   http://www.nature.com/scientificamerican/journal/
v313/n5/pdf/scientificamerican1115-70.pdf.

DiChristina:2015:EIC

[2608] Mariette DiChristina. From the Editor: Innovation celebration. Scientific American, 313(6):4, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-

4.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-4.pdf.

Anonymous:2015:Ll

[2609] Anonymous. Letters. Scientific American, 313(6):6–8, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n6/full/scientificamerican1215-6.html;   http://www.nature.
com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
6.pdf.

Schlenoff:2015:YAIl

[2610] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 313(6):83, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-83.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
83.pdf.

Anonymous:2015:SAP

[2611] Anonymous. Science agenda: The price of pollution. Scientific American, 313(6):10, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
10.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-10.pdf.

Mann:2015:FTD

[2612] Michael E. Mann. Forum: Two degrees of freedom. Scientific American, 313(6):12, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
12.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-12.pdf.

Hsu:2015:ADS

[2613] Jeremy Hsu. Advances: Defanging snakebites. Scientific American, 313(6):14–16, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
14.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-14.pdf.

Nuwer:2015:MTT

[2614] Rachel Nuwer. Money talks — and tweets. Scientific American, 313(6):17, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
17.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-17.pdf.

Nuwer:2015:KCD

[2615] Rachel Nuwer. Kids can do better in the crosswalk. Scientific American, 313(6):18–19, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
18.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-18.pdf.

Godkin:2015:BP

[2616] David Godkin. Batty performers. Scientific American, 313(6):20, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-20.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
20.pdf.

Hall:2015:MMI

[2617] Shannon Hall. The molten mechanics of inner Earth. Scientific American, 313(6):21, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
21.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-21.pdf.

Little:2015:BG

[2618] Jane Braxton Little. Banking on groundwater. Scientific American, 313(6):22, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
22.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-22.pdf.

Anonymous:2015:QHl

[2619] Anonymous. Quick hits. Scientific American, 313(6):23, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n6/full/scientificamerican1215-23.html;   http://www.nature.
com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
23.pdf.

Gannon:2015:NN

[2620] Megan Gannon. Ned the nose. Scientific American, 313(6):24, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-24.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
24.pdf.

Hall:2015:BBP

[2621] Shannon Hall. The Big Bang’s particle glow. Scientific American, 313(6):25, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
25.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-25.pdf.

Kuchment:2015:SFW

[2622] Anna Kuchment. She faired well. Scientific American, 313(6):26, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-26.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
26.pdf.

Noonan:2015:SHP

[2623] David Noonan. The science of health: A pain in the brain. Scientific American, 313(6):27–28, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v313/n6/full/scientificamerican1215-
27.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-27.pdf.

Pogue:2015:TTJ

[2624] David Pogue. TechnoFiles: The tech jungle. Scientific American, 313(6):29, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-

29.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-29.pdf.

Anonymous:2015:RRBf

[2625] Anonymous. Reviews and recommendations: Book review: Extreme Earth. Scientific American, 313(6):80, December 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n6/full/scientificamerican1215-80a.html;   http://www.nature.
com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
80a.pdf.

Anonymous:2015:BRSb

[2626] Anonymous. Book review: Science of the magical. Scientific American, 313(6):80, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
80b.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-80b.pdf.

Anonymous:2015:BRP

[2627] Anonymous. Book review: Pacific. Scientific American, 313(6):80, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-80c.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
80c.pdf.

Shermer:2015:SCC

[2628] Michael Shermer. Skeptic: Consilience and consensus. Scientific American, 313(6):81, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
81.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-81.pdf.

Mirsky:2015:AGGb

[2629] Steve Mirsky. Anti gravity: Great feets. Scientific American, 313(6):82, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-82.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
82.pdf.

Fischetti:2015:GSB

[2630] Mark Fischetti. Graphic science: The bacteria game. Scientific American, 313(6):84, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
84.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-84.pdf.

Anonymous:2015:WCI

[2631] Anonymous. World changing ideas 2015: Innovation. Scientific American, 313(6):30–31, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
30.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-30.pdf.

Nuwer:2015:ECM

[2632] Rachel Nuwer. Eye-controlled machines. Scientific American, 313(6):32, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-32.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
32.pdf.

Billings:2015:MR

[2633] Lee Billings. Microwave rocketry. Scientific American, 313(6):33, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-33.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
33.pdf.

Nuwer:2015:TV

[2634] Rachel Nuwer. Trawling for viruses. Scientific American, 313(6):34, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-34a.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
34a.pdf.

Fletcher:2015:SIE

[2635] Seth Fletcher. Soft, injectable electronic probes for the brain. Scientific American, 313(6):34, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
34b.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-34b.pdf.

Biello:2015:LF

[2636] David Biello. Little fusion. Scientific American, 313(6):35, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-35.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
35.pdf.

Abbasi:2015:KSG

[2637] Jennifer Abbasi. Kill switches for GMOs. Scientific American, 313(6):36, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-36.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
36.pdf.

Nuwer:2015:HV

[2638] Rachel Nuwer. The heat vacuum. Scientific American, 313(6):37, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v313/n6/full/scientificamerican1215-37.html;   http:
//www.nature.com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
37.pdf.

Stix:2015:MTT

[2639] Gary Stix. Machines that teach themselves. Scientific American, 313(6):38, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
38.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-38.pdf.

Abbasi:2015:SMC

[2640] Jennifer Abbasi. Slow-motion cameras for chemical reactions. Scientific American, 313(6):39, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-

39a.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-39a.pdf.

Greenemeier:2015:SAC

[2641] Larry Greenemeier. Seeing around corners. Scientific American, 313(6):39, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
39b.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-39b.pdf.

Kevles:2015:IIW

[2642] Daniel J. Kevles. Innovation: Inventing the world 1845. Scientific American, 313(6):40–53, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
40.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-40.pdf.

Anonymous:2015:YSA

[2643] Anonymous. 170 years of Scientific American. Scientific American, 313(6):42–53, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
42.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-42.pdf.

Brusatte:2015:PWK

[2644] Stephen Brusatte. Paleontology: What killed the dinosaurs. Scientific American, 313(6):54–59, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v313/n6/full/scientificamerican1215-
54.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-54.pdf.

Worth:2015:ATW

[2645] Katie Worth. Astronomy: Telescope wars. Scientific American, 313(6):60–65, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
60.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-60.pdf.

Strauss:2015:MGP

[2646] Kevin A. Strauss. Medicine: Genomics for the people. Scientific American, 313(6):66–73, December 2015. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v313/n6/full/scientificamerican1215-
66.html;   http://www.nature.com/scientificamerican/journal/
v313/n6/pdf/scientificamerican1215-66.pdf.

Oreskes:2015:EHB

[2647] Naomi Oreskes. Environment: How to break the climate deadlock. Scientific American, 313(6):74–79, December 2015. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v313/
n6/full/scientificamerican1215-74.html;   http://www.nature.
com/scientificamerican/journal/v313/n6/pdf/scientificamerican1215-
74.pdf.

DiChristina:2016:EFO

[2648] Mariette DiChristina. From the Editor: Finding our place. Scientific American, 314(1):4, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
4.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-4.pdf.

Anonymous:2016:La

[2649] Anonymous. Letters. Scientific American, 314(1):6–8, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n1/full/scientificamerican0116-6.html;   http://www.nature.
com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
6.pdf.

Schlenoff:2016:YAIa

[2650] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 314(1):78, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-78.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
78.pdf.

Anonymous:2016:SACa

[2651] Anonymous. Science agenda: A campaign to cut drug prices. Scientific American, 314(1):10, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
10.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-10.pdf.

Pomeroy:2016:FAG

[2652] Claire Pomeroy. Forum: Academia’s gender problem. Scientific American, 314(1):11, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
11.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-11.pdf.

Billings:2016:AFU

[2653] Lee Billings. Advances: Fear of the unknown. Scientific American, 314(1):12–15, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
12.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-12.pdf.

Hall:2016:SS

[2654] Shannon Hall. Soaked in space. Scientific American, 314(1):14–15, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-14.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
14.pdf.

Platt:2016:SA

[2655] John R. Platt. The starfish assassin. Scientific American, 314(1):16, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-16.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
16.pdf.

Callier:2016:DC

[2656] Viviane Callier. Divide or conquer. Scientific American, 314(1):17, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-17.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
17.pdf.

Goldman:2016:BB

[2657] Jason G. Goldman. Ballots for bison. Scientific American, 314(1):18, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-18a.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
18a.pdf.

Iozzio:2016:GC

[2658] Corinne Iozzio. Ground control. Scientific American, 314(1):18, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-18b.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
18b.pdf.

Nuwer:2016:BS

[2659] Rachel Nuwer. Branches of science. Scientific American, 314(1):19, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-19.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
19.pdf.

Anonymous:2016:QHa

[2660] Anonymous. Quick hits. Scientific American, 314(1):20, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n1/full/scientificamerican0116-20.html;   http://www.nature.
com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
20.pdf.

Hall:2016:FF

[2661] Shannon Hall. Frictionless finance. Scientific American, 314(1):21, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-21.html;   http:

//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
21.pdf.

Makin:2016:BW

[2662] Simon Makin. The brain’s whorl. Scientific American, 314(1):22, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-22.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
22.pdf.

Jabr:2016:SHC

[2663] Ferris Jabr. The science of health: That craving for dessert. Scientific American, 314(1):23–24, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
23.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-23.pdf.

Pogue:2016:TCP

[2664] David Pogue. TechnoFiles: Click ‘n’ pay. Scientific American, 314(1):25, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-25.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
25.pdf.

Anonymous:2016:RRBa

[2665] Anonymous. Reviews and recommendations: Book review: Light. Scientific American, 314(1):74, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
74a.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-74a.pdf.

Anonymous:2016:BRPa

[2666] Anonymous. Book review: Political animals. Scientific American, 314(1): 74, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n1/full/scientificamerican0116-74b.html;   http:
//www.nature.com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
74b.pdf.

Anonymous:2016:BRBa

[2667] Anonymous. Book review: A brief history of creation. Scientific American, 314(1):74, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
74c.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-74c.pdf.

Anonymous:2016:BRA

[2668] Anonymous. Book review: The astronomer and the witch. Scientific American, 314(1):74, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
74d.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-74d.pdf.

Shermer:2016:SMC

[2669] Michael Shermer. Skeptic: Murder in the cave. Scientific American, 314(1):75, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
75.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-75.pdf.

Mirsky:2016:AGE

[2670] Steve Mirsky. Anti gravity: 3× equals an omelet. Scientific American, 314(1):76, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
76.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-76.pdf.

Williams:2016:GSS

[2671] Amber Williams. Graphic science: Sleeping beauties of science. Scientific American, 314(1):80, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
80.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-80.pdf.

Moser:2016:NWWa

[2672] May-Britt Moser and Edvard I. Moser. Neuroscience: Where am I? Where am I going? Scientific American, 314(1):26–33, January 2016.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n1/full/scientificamerican0116-26.html;   http://www.nature.
com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
26.pdf.

Kenworthy:2016:PSR

[2673] Matthew Kenworthy. Planetary science: Rings of a super Saturn. Scientific American, 314(1):34–41, January 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n1/full/scientificamerican0116-34.html;   http://www.nature.
com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
34.pdf.

Daigle:2016:PHD

[2674] Katy Daigle. Public health: Death in the water. Scientific American, 314(1):42–51, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
42.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-42.pdf.

Kozubek:2016:MSG

[2675] Jim Kozubek. Medicine: An on/off switch for genes. Scientific American, 314(1):52–57, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
52.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-52.pdf.

Biello:2016:ECC

[2676] David Biello. Energy: The carbon capture fallacy. Scientific American, 314(1):58–65, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
58.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-58.pdf.

Dunsworth:2016:CDA

[2677] Holly Dunsworth. Cognition: Do animals know where babies come from? Scientific American, 314(1):66–69, January 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n1/full/scientificamerican0116-66.html;   http://www.nature.
com/scientificamerican/journal/v314/n1/pdf/scientificamerican0116-
66.pdf.

Moskowitz:2016:MEE

[2678] Clara Moskowitz. Mathematics: Elegant equations. Scientific American, 314(1):70–73, January 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n1/full/scientificamerican0116-
70.html;   http://www.nature.com/scientificamerican/journal/
v314/n1/pdf/scientificamerican0116-70.pdf.

DiChristina:2016:EWP

[2679] Mariette DiChristina. From the Editor: Where is Planet X? Scientific American, 314(2):4, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
4.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-4.pdf.

Anonymous:2016:Lb

[2680] Anonymous. Letters. Scientific American, 314(2):6–8, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n2/full/scientificamerican0216-6.html;   http://www.nature.
com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
6.pdf.

Schlenoff:2016:YAIb

[2681] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 314(2):75, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n2/full/scientificamerican0216-75.html;   http:
//www.nature.com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
75.pdf.

Anonymous:2016:SABa

[2682] Anonymous. Science agenda: Beware prenatal gene screens. Scientific American, 314(2):10, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-

10.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-10.pdf.

Myhrvold:2016:FEG

[2683] Nathan Myhrvold. Forum: Even genius needs a benefactor. Scientific American, 314(2):11, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
11.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-11.pdf.

Moskowitz:2016:ALC

[2684] Clara Moskowitz. Advances: Last call: Will WIMPs show their faces? Scientific American, 314(2):12–14, February 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n2/full/scientificamerican0216-12.html;   http://www.nature.
com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
12.pdf.

Pycha:2016:TT

[2685] Anne Pycha. Talk therapy. Scientific American, 314(2):14–15, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n2/full/scientificamerican0216-14.html;   http:
//www.nature.com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
14.pdf.

Giller:2016:ETM

[2686] Geoffrey Giller. The ear as tape measure. Scientific American, 314(2):16, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n2/full/scientificamerican0216-16a.html;   http:
//www.nature.com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
16a.pdf.

Greenspan:2016:PP

[2687] Jesse Greenspan. Poison pot. Scientific American, 314(2):17, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n2/full/scientificamerican0216-17.html;   http:
//www.nature.com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
17.pdf.

Smith:2016:NCT

[2688] Peter Andrey Smith. The new crash test dummy. Scientific American, 314(2):18–19, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
18.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-18.pdf.

Anonymous:2016:QHb

[2689] Anonymous. Quick hits. Scientific American, 314(2):20, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n2/full/scientificamerican0216-20.html;   http://www.nature.
com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
20.pdf.

Croswell:2016:SJ

[2690] Ken Croswell. Space jellies. Scientific American, 314(2):21, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n2/full/scientificamerican0216-21.html;   http:
//www.nature.com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
21.pdf.

Hackett:2016:EU

[2691] Jennifer Hackett. Elemental urgency. Scientific American, 314(2):22, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n2/full/scientificamerican0216-22.html;   http:
//www.nature.com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
22.pdf.

Goldman:2016:WLW

[2692] Jason G. Goldman. When love wins. Scientific American, 314(2):23, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n2/full/scientificamerican0216-23.html;   http:
//www.nature.com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
23.pdf.

Pogue:2016:TGI

[2693] David Pogue. TechnoFiles: Get out the iVote. Scientific American, 314(2):24, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
24.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-24.pdf.

Tanner:2016:SHS

[2694] Adam Tanner. The science of health: For sale: Your medical records. Scientific American, 314(2):26–27, February 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n2/full/scientificamerican0216-26.html;   http://www.nature.
com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
26.pdf.

Anonymous:2016:RRBb

[2695] Anonymous. Reviews and recommendations: Book review: The last volcano. Scientific American, 314(2):72, February 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n2/full/scientificamerican0216-72a.html;   http://www.nature.
com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
72a.pdf.

Anonymous:2016:BRT

[2696] Anonymous. Book review: Tunnel visions. Scientific American, 314(2):72, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
72b.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-72b.pdf.

Anonymous:2016:BRC

[2697] Anonymous. Book review: Contact sport. Scientific American, 314(2):72, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
72c.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-72c.pdf.

Anonymous:2016:BRH

[2698] Anonymous. Book review: Home. Scientific American, 314(2):72, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n2/full/scientificamerican0216-72d.html;   http:
//www.nature.com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
72d.pdf.

Shermer:2016:SAA

[2699] Michael Shermer. Skeptic: Afterlife for atheists. Scientific American, 314(2):73, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
73.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-73.pdf.

Mirsky:2016:AGS

[2700] Steve Mirsky. Anti gravity: Spinning our wheels. Scientific American, 314(2):74, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
74.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-74.pdf.

Fischetti:2016:GSD

[2701] Mark Fischetti. Graphic science: Disease squeeze. Scientific American, 314(2):76, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
76.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-76.pdf.

Lemonick:2016:SSPa

[2702] Michael D. Lemonick. Space: The search for Planet X. Scientific American, 314(2):30–37, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
30.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-30.pdf.

Lee:2016:HBT

[2703] Robert J. Lee and Noam A. Cohen. Health: Bitter taste bodyguards. Scientific American, 314(2):38–43, February 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n2/full/scientificamerican0216-38.html;   http://www.nature.

com/scientificamerican/journal/v314/n2/pdf/scientificamerican0216-
38.pdf.

Gordon:2016:SBC

[2704] Deborah M. Gordon. Social behavior: Collective wisdom of ants. Scientific American, 314(2):44–47, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v314/n2/full/scientificamerican0216-
44.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-44.pdf.

Folger:2016:TQH

[2705] Tim Folger. Technology: The quantum hack. Scientific American, 314(2):48–55, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
48.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-48.pdf.

Engelman:2016:SSB

[2706] Robert Engelman. Sustainability: Six billion in Africa. Scientific American, 314(2):56–63, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
56.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-56.pdf.

Hensch:2016:NPI

[2707] Takao K. Hensch. Neuroscience: The power of the infant brain. Scientific American, 314(2):64–69, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
64.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-64.pdf.

Levy:2016:AML

[2708] David H. Levy. Astronomy: My life as a comet hunter. Scientific American, 314(2):70–71, February 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n2/full/scientificamerican0216-
70.html;   http://www.nature.com/scientificamerican/journal/
v314/n2/pdf/scientificamerican0216-70.pdf.

DiChristina:2016:ELR

[2709] Mariette DiChristina. From the Editor: Looking to the rising star. Scientific American, 314(3):4, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
4.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-4.pdf.

Anonymous:2016:Lc

[2710] Anonymous. Letters. Scientific American, 314(3):6–8, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n3/full/scientificamerican0316-6.html;   http://www.nature.
com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
6.pdf.

Schlenoff:2016:YAIc

[2711] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 314(3):75, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-75.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
75.pdf.

Anonymous:2016:SACb

[2712] Anonymous. Science agenda: A clear view of medical care. Scientific American, 314(3):10, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
10.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-10.pdf.

Nachman:2016:FPP

[2713] Keeve Nachman. Forum: Pills for pigs: Just say no. Scientific American, 314(3):11, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
11.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-11.pdf.

Biello:2016:ACO

[2714] David Biello. Advances: Cheap oil undermines climate cleanup. Scientific American, 314(3):12–15, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
12.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-12.pdf.

Wapner:2016:BP

[2715] Jessica Wapner. Off the beaten path. Scientific American, 314(3):16, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-16a.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
16a.pdf.

Hsu:2016:RA

[2716] Jeremy Hsu. On the road again. Scientific American, 314(3):16–17, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-16b.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
16b.pdf.

Mukerjee:2016:FT

[2717] Madhusree Mukerjee. Fukushima today. Scientific American, 314(3):18– 19, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-18.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
18.pdf.

Kwon:2016:CWE

[2718] Diana Kwon. The connection was electric. Scientific American, 314(3): 20, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-20a.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
20a.pdf.

Schmidt:2016:NHN

[2719] Charles Schmidt. Neither here nor there. Scientific American, 314(3):20, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-20b.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
20b.pdf.

Anonymous:2016:QHc

[2720] Anonymous. Quick hits. Scientific American, 314(3):21, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n3/full/scientificamerican0316-21.html;   http://www.nature.
com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
21.pdf.

Hackett:2016:CC

[2721] Jennifer Hackett. Classification conundrum. Scientific American, 314(3): 22, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-22a.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
22a.pdf.

Brouillette:2016:YCE

[2722] Monique Brouillette. You can edit a pig, but it will still be a pig. Scientific American, 314(3):22, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
22b.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-22b.pdf.

Marmot:2016:SHD

[2723] Michael Marmot. The science of health: The disease of poverty. Scientific American, 314(3):23–24, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
23.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-23.pdf.

Pogue:2016:TU

[2724] David Pogue. TechnoFiles: Untangled. Scientific American, 314(3):26, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-26.html;   http:

//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
26.pdf.

Anonymous:2016:RRBc

[2725] Anonymous. Reviews and recommendations: Book review: Pollination power. Scientific American, 314(3):72, March 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n3/full/scientificamerican0316-72a.html;   http://www.nature.
com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
72a.pdf.

Anonymous:2016:BRBb

[2726] Anonymous. Book review: Blood and Earth. Scientific American, 314(3): 72, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-72b.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
72b.pdf.

Anonymous:2016:BRPb

[2727] Anonymous. Book review: Pandemic. Scientific American, 314(3):72, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-72c.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
72c.pdf.

Anonymous:2016:BRSa

[2728] Anonymous. Book review: Seven brief lessons on physics. Scientific American, 314(3):72, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
72d.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-72d.pdf.

Shermer:2016:SLB

[2729] Michael Shermer. Skeptic: Left behind. Scientific American, 314(3):73, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-73.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
73.pdf.

Mirsky:2016:AGFa

[2730] Steve Mirsky. Anti gravity: Fuzzy math. Scientific American, 314(3):74, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-74.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
74.pdf.

Fischetti:2016:GSC

[2731] Mark Fischetti. Graphic science: The circle of life. Scientific American, 314(3):76, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
76.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-76.pdf.

Wong:2016:EMH

[2732] Kate Wong. Evolution: Mystery human. Scientific American, 314(3):28– 37, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n3/full/scientificamerican0316-28.html;   http:
//www.nature.com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
28.pdf.

Riess:2016:CPD

[2733] Adam G. Riess and Mario Livio. Cosmology: The puzzle of dark energy. Scientific American, 314(3):38–43, March 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n3/full/scientificamerican0316-38.html;   http://www.nature.
com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
38.pdf.

Nedergaard:2016:NBD

[2734] Maiken Nedergaard and Steven A. Goldman. Neuroscience: Brain drain. Scientific American, 314(3):44–49, March 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n3/full/scientificamerican0316-44.html;   http://www.nature.
com/scientificamerican/journal/v314/n3/pdf/scientificamerican0316-
44.pdf.

Wendle:2016:SSCa

[2735] John Wendle. Sustainability: Syria’s climate refugees. Scientific American, 314(3):50–55, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
50.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-50.pdf.

Hall:2016:BEM

[2736] Stephen S. Hall. Biology: Editing the mushroom. Scientific American, 314(3):56–63, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
56.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-56.pdf.

Adler:2016:CC

[2737] Paul N. Adler and Jeremy Nathans. The cellular compass. Scientific American, 314(3):67–71, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
67.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-67.pdf.

Lee:2016:GAM

[2738] Howard Lee. Geology: Anatomy of a mass murderer. Scientific American, 314(3):64–66, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n3/full/scientificamerican0316-
64.html;   http://www.nature.com/scientificamerican/journal/
v314/n3/pdf/scientificamerican0316-64.pdf.

DiChristina:2016:ERO

[2739] Mariette DiChristina. From the Editor: Rocks in our heads. Scientific American, 314(4):4, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
4.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-4.pdf.

Anonymous:2016:Ld

[2740] Anonymous. Letters. Scientific American, 314(4):6–8, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n4/full/scientificamerican0416-6.html;   http://www.nature.
com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
6.pdf.

Schlenoff:2016:YAId

[2741] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 314(4):82, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n4/full/scientificamerican0416-82.html;   http:
//www.nature.com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
82.pdf.

Anonymous:2016:SASa

[2742] Anonymous. Science agenda: Stop dithering on nuclear waste. Scientific American, 314(4):10, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
10.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-10.pdf.

vonHippel:2016:FCD

[2743] Frank von Hippel. Forum: Chernobyl didn’t kill nuclear power. Scientific American, 314(4):11, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
11.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-11.pdf.

Goldman:2016:ALL

[2744] Jason G. Goldman. Advances: Lions on the ledge. Scientific American, 314(4):12–14, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
12.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-12.pdf.

Hackett:2016:ASB

[2745] Jennifer Hackett. April showers bring May meteorites. Scientific American, 314(4):15, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-

15.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-15.pdf.

Jacewicz:2016:TMB

[2746] Natalie Jacewicz. A twist on the mushroom burger. Scientific American, 314(4):16, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
16.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-16.pdf.

Hackett:2016:PAN

[2747] Jennifer Hackett. Passengers announced for NASA’s biggest rocket. Scientific American, 314(4):17, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
17.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-17.pdf.

Nuwer:2016:FTW

[2748] Rachel Nuwer. Forecaster for the top of the world. Scientific American, 314(4):18–19, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
18.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-18.pdf.

Anonymous:2016:QHd

[2749] Anonymous. Quick hits. Scientific American, 314(4):20, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n4/full/scientificamerican0416-20.html;   http://www.nature.
com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
20.pdf.

Nuwer:2016:OBM

[2750] Rachel Nuwer. Orphaned bugs make bum parents. Scientific American, 314(4):21, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
21.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-21.pdf.

Callier:2016:CCC

[2751] Viviane Callier. Cancerous co-conspirators. Scientific American, 314(4): 22, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n4/full/scientificamerican0416-22.html;   http:
//www.nature.com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
22.pdf.

Qiu:2016:SRH

[2752] Jane Qiu. Silk Road heads for the hills. Scientific American, 314(4):23, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n4/full/scientificamerican0416-23.html;   http:
//www.nature.com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
23.pdf.

Interlandi:2016:SHP

[2753] Jeneen Interlandi. The science of health: The paradox of precision medicine. Scientific American, 314(4):24–25, April 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n4/full/scientificamerican0416-24.html;   http://www.nature.
com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
24.pdf.

Pogue:2016:TDD

[2754] David Pogue. TechnoFiles: Dumb design. Scientific American, 314(4): 26, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n4/full/scientificamerican0416-26.html;   http:
//www.nature.com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
26.pdf.

Anonymous:2016:RRBd

[2755] Anonymous. Reviews and recommendations: Book review: The right kind of crazy. Scientific American, 314(4):78, April 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n4/full/scientificamerican0416-78a.html;   http://www.nature.
com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
78a.pdf.

Anonymous:2016:BRSb

[2756] Anonymous. Book review: Snowball in a blizzard. Scientific American, 314(4):78, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
78b.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-78b.pdf.

Anonymous:2016:BRG

[2757] Anonymous. Book review: The genius of birds. Scientific American, 314(4):78, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
78c.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-78c.pdf.

Anonymous:2016:BRBc

[2758] Anonymous. Book review: Burn math class. Scientific American, 314(4): 78, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n4/full/scientificamerican0416-78d.html;   http:
//www.nature.com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
78d.pdf.

Shermer:2016:SHD

[2759] Michael Shermer. Skeptic: Hooey. drivel. baloney …. Scientific American, 314(4):79, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
79.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-79.pdf.

Mirsky:2016:AGT

[2760] Steve Mirsky. Anti gravity: There are no words. Scientific American, 314(4):80, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
80.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-80.pdf.

Maron:2016:GST

[2761] Dina Fine Maron. Graphic science: Travels with zika. Scientific American, 314(4):84, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
84.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-84.pdf.

Stout:2016:CPT

[2762] Dietrich Stout. Cognitive psychology: Tales of a stone age neuroscientist. Scientific American, 314(4):28–35, April 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n4/full/scientificamerican0416-28.html;   http://www.nature.
com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
28.pdf.

Greene:2016:PPN

[2763] Geoffrey L. Greene and Peter Geltenbort. Particle physics: The neutron enigma. Scientific American, 314(4):36–41, April 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n4/full/scientificamerican0416-36.html;   http://www.nature.
com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
36.pdf.

Weintraub:2016:SRF

[2764] Karen Weintraub. Special report: 2016 future of medicine: The cancer defense. Scientific American, 314(4):42–51, April 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n4/full/scientificamerican0416-42.html;   http://www.nature.
com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
42.pdf.

Carreno:2016:VC

[2765] Beatriz M. Carreno and Elaine R. Mardis. A vaccine for cancer? Scientific American, 314(4):46, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
46.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-46.pdf.

Alegre:2016:GW

[2766] Maria-Luisa Alegre and Thomas F. Gajewski. Germ warfare. Scientific American, 314(4):50, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
50.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-50.pdf.

Tullis:2016:EGS

[2767] Paul Tullis. Ecology: Galápagos stampede. Scientific American, 314(4):52–57, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
52.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-52.pdf.

Lu:2016:BMLa

[2768] Timothy K. Lu and Oliver Purcell. Bioengineering: Machine life. Scientific American, 314(4):58–63, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
58.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-58.pdf.

Ksepka:2016:EGS

[2769] Daniel T. Ksepka and Michael Habib. Evolution: Giants of the sky. Scientific American, 314(4):64–71, April 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n4/full/scientificamerican0416-64.html;   http://www.nature.
com/scientificamerican/journal/v314/n4/pdf/scientificamerican0416-
64.pdf.

Everts:2016:CAS

[2770] Sarah Everts. Chemistry: The art of saving relics. Scientific American, 314(4):72–77, April 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n4/full/scientificamerican0416-
72.html;   http://www.nature.com/scientificamerican/journal/
v314/n4/pdf/scientificamerican0416-72.pdf.

DiChristina:2016:EOW

[2771] Mariette DiChristina. From the Editor: Outer worlds and our inner itch. Scientific American, 314(5):4, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n5/full/scientificamerican0516-

4.html;   http://www.nature.com/scientificamerican/journal/
v314/n5/pdf/scientificamerican0516-4.pdf.

Anonymous:2016:Le

[2772] Anonymous. Letters. Scientific American, 314(5):6–8, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-6.html;   http://www.nature.
com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
6.pdf.

Schlenoff:2016:YAIe

[2773] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 314(5): 74, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-74.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
74.pdf.

Anonymous:2016:SADa

[2774] Anonymous. Science agenda: This drug ad is not right for you. Scientific American, 314(5):10, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n5/full/scientificamerican0516-
10.html;   http://www.nature.com/scientificamerican/journal/
v314/n5/pdf/scientificamerican0516-10.pdf.

Rodin:2016:FSV

[2775] Alexander Rodin. Forum: Show Venus some love. Scientific American, 314(5):11, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n5/full/scientificamerican0516-
11.html;   http://www.nature.com/scientificamerican/journal/
v314/n5/pdf/scientificamerican0516-11.pdf.

Iozzio:2016:AWR

[2776] Corinne Iozzio. Advances: Who’s responsible when a car controls the wheel? Scientific American, 314(5):12–13, May 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-12.html;   http://www.nature.

com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
12.pdf.

Nuwer:2016:MCH

[2777] Rachel Nuwer. Monster comes out of hiding. Scientific American, 314(5): 13, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-13.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
13.pdf.

Hall:2016:DW

[2778] Shannon Hall. Deserted waters. Scientific American, 314(5):14, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-14.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
14.pdf.

Nuwer:2016:PP

[2779] Rachel Nuwer. Plugged in or powered down? Scientific American, 314 (5):15, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-15.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
15.pdf.

Greenspan:2016:FSL

[2780] Jesse Greenspan. A forest’s second life. Scientific American, 314(5): 16, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-16.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
16.pdf.

Moyer:2016:HMA

[2781] Melinda Wenner Moyer. Heartburn meds alter the gut. Scientific American, 314(5):17, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n5/full/scientificamerican0516-
17.html;   http://www.nature.com/scientificamerican/journal/
v314/n5/pdf/scientificamerican0516-17.pdf.

Jacobson:2016:ER

[2782] Roni Jacobson. Is Eureka right? Scientific American, 314(5):18, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-18.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
18.pdf.

Lloyd:2016:MM

[2783] Robin Lloyd. Mortal mains. Scientific American, 314(5):19, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-19.html;   http://www.nature.
com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
19.pdf.

Anonymous:2016:QHe

[2784] Anonymous. Quick hits. Scientific American, 314(5):20, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-20.html;   http://www.nature.
com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
20.pdf.

Scharf:2016:SC

[2785] Caleb Scharf. Spin cycle. Scientific American, 314(5):21, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-21.html;   http://www.nature.
com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
21.pdf.

Wapner:2016:DR

[2786] Jessica Wapner. Drawing for a remedy. Scientific American, 314(5): 22, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-22.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
22.pdf.

Palucka:2016:HSW

[2787] Tim Palucka. How to split wood. Scientific American, 314(5):23, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-23.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
23.pdf.

Weintraub:2016:SHS

[2788] Karen Weintraub. The science of health: A surprising fix for sickle cell. Scientific American, 314(5):24–26, May 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-24.html;   http://www.nature.
com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
24.pdf.

Pogue:2016:TM

[2789] David Pogue. TechnoFiles: In the movies. Scientific American, 314(5): 27, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-27.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
27.pdf.

Anonymous:2016:RRRa

[2790] Anonymous. Reviews and recommendations: Recommended. Scientific American, 314(5):70, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n5/full/scientificamerican0516-
70.html;   http://www.nature.com/scientificamerican/journal/
v314/n5/pdf/scientificamerican0516-70.pdf.

Shermer:2016:SDC

[2791] Michael Shermer. Skeptic: Doomsday catch. Scientific American, 314(5): 72, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-72.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
72.pdf.

Mirsky:2016:AGB

[2792] Steve Mirsky. Anti gravity: Big hummer. Scientific American, 314(5): 73, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-73.html;   http:

//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
73.pdf.

Fischetti:2016:GSL

[2793] Mark Fischetti. Graphic science: Lightning hotspots. Scientific American, 314(5):76, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n5/full/scientificamerican0516-
76.html;   http://www.nature.com/scientificamerican/journal/
v314/n5/pdf/scientificamerican0516-76.pdf.

Batygin:2016:ABC

[2794] Konstantin Batygin, Gregory Laughlin, and Alessandro Morbidelli. Astronomy: Born of chaos. Scientific American, 314(5):28–37, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-28.html;   http://www.nature.
com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
28.pdf.

Lemonick:2016:PNO

[2795] Michael D. Lemonick. Planet Nine from outer space. Scientific American, 314(5):36, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n5/full/scientificamerican0516-
36.html;   http://www.nature.com/scientificamerican/journal/
v314/n5/pdf/scientificamerican0516-36.pdf.

Sutherland:2016:NMS

[2796] Stephani Sutherland. Neurobiology: The maddening sensation of itch. Scientific American, 314(5):38–43, May 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-38.html;   http://www.nature.
com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
38.pdf.

Nuwer:2016:ESE

[2797] Rachel Nuwer. Ecology: Saving Eden. Scientific American, 314(5):44– 49, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-44.html;   http:

//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
44.pdf.

Monroe:2016:TQC

[2798] Christopher R. Monroe, Robert J. Schoelkopf, and Mikhail D. Lukin. Technology: Quantum connections. Scientific American, 314(5):50–57, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-50.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
50.pdf.

Fischman:2016:BSU

[2799] Josh Fischman. Biomechanics: Swimmers under pressure. Scientific American, 314(5):58–59, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n5/full/scientificamerican0516-
58.html;   http://www.nature.com/scientificamerican/journal/
v314/n5/pdf/scientificamerican0516-58.pdf.

Balter:2016:LLW

[2800] Michael Balter. Linguistics: Language wars. Scientific American, 314(5): 60–65, May 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n5/full/scientificamerican0516-60.html;   http:
//www.nature.com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
60.pdf.

Reganold:2016:ECA

[2801] John P. Reganold and Jerry D. Glover. Environment: A cure for Africa’s soil. Scientific American, 314(5):66–69, May 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n5/full/scientificamerican0516-66.html;   http://www.nature.
com/scientificamerican/journal/v314/n5/pdf/scientificamerican0516-
66.pdf.

DiChristina:2016:ESS

[2802] Mariette DiChristina. From the Editor: A success story. Scientific American, 314(6):4, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-

4.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-4.pdf.

Anonymous:2016:Lf

[2803] Anonymous. Letters. Scientific American, 314(6):5–6, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n6/full/scientificamerican0616-5.html;   http://www.nature.
com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
5.pdf.

Schlenoff:2016:YAIf

[2804] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American. Scientific American, 314(6):79, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-79.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
79.pdf.

Anonymous:2016:SAR

[2805] Anonymous. Science agenda: Rules for cyberwar. Scientific American, 314(6):7, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
7.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-7.pdf.

Baer:2016:FCC

[2806] Howard Baer, Vernon D. Barger, and Jenny List. Forum: The collider that could save physics. Scientific American, 314(6):8, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n6/full/scientificamerican0616-8.html;   http://www.nature.
com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
8.pdf.

Smith:2016:AWD

[2807] Peter Andrey Smith. Advances: When DNA implicates the innocent. Scientific American, 314(6):11–12, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v314/n6/full/scientificamerican0616-

11.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-11.pdf.

Platt:2016:UCG

[2808] John R. Platt. Ugly critters get no love. Scientific American, 314(6):12– 13, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-12.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
12.pdf.

Hsu:2016:NPI

[2809] Jeremy Hsu. A Northwest Passage for the Internet. Scientific American, 314(6):14–15, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
14.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-14.pdf.

Hackett:2016:FF

[2810] Jennifer Hackett. Fast faults. Scientific American, 314(6):16, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n6/full/scientificamerican0616-16.html;   http://www.nature.
com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
16.pdf.

Hackett:2016:LTO

[2811] Jennifer Hackett. Lab tech opening: 249 miles above Earth. Scientific American, 314(6):17, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
17.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-17.pdf.

Goldman:2016:TTK

[2812] Jason G. Goldman. Toddlers’ tool kits. Scientific American, 314(6):18, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-18.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
18.pdf.

Anonymous:2016:QHf

[2813] Anonymous. Quick hits. Scientific American, 314(6):19, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n6/full/scientificamerican0616-19.html;   http://www.nature.
com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
19.pdf.

Nuwer:2016:EE

[2814] Rachel Nuwer. Eggshell education. Scientific American, 314(6):20, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-20.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
20.pdf.

Brouillette:2016:MU

[2815] Monique Brouillette. A mental unblock. Scientific American, 314(6):21, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-21.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
21.pdf.

Cepelewicz:2016:ENP

[2816] Jordana Cepelewicz. Eyeglasses, no prescription necessary. Scientific American, 314(6):22, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
22.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-22.pdf.

Shute:2016:SHP

[2817] Nancy Shute. The science of health: A plan to prevent gun suicides. Scientific American, 314(6):25–26, June 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n6/full/scientificamerican0616-25.html;   http://www.nature.
com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
25.pdf.

Pogue:2016:THR

[2818] David Pogue. TechnoFiles: Hail, robo-taxi. Scientific American, 314(6): 27, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-27.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
27.pdf.

Anonymous:2016:RRRb

[2819] Anonymous. Reviews and recommendations: Recommended. Scientific American, 314(6):76, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
76.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-76.pdf.

Shermer:2016:SDW

[2820] Michael Shermer. Skeptic: Death wish. Scientific American, 314(6):77, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-77.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
77.pdf.

Mirsky:2016:AGA

[2821] Steve Mirsky. Anti gravity: Arms race. Scientific American, 314(6):78, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-78.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
78.pdf.

Kwon:2016:GSC

[2822] Diana Kwon. Graphic science: Can you read this from a distance? Scientific American, 314(6):80, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
80.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-80.pdf.

Brusatte:2016:EAM

[2823] Stephen Brusatte and Zhe-Xi Luo. Evolution: Ascent of the mammals. Scientific American, 314(6):28–35, June 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n6/full/scientificamerican0616-28.html;   http://www.nature.
com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
28.pdf.

Kasen:2016:ASF

[2824] Daniel Kasen. Astrophysics: Stellar fireworks. Scientific American, 314(6):36–43, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
36.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-36.pdf.

Anonymous:2016:ASR

[2825] Anonymous. Ai special report: The rise of AI. Scientific American, 314(6):44–45, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
44.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-44.pdf.

Bengio:2016:MWL

[2826] Yoshua Bengio. Machines who learn. Scientific American, 314(6):46–51, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v314/n6/full/scientificamerican0616-46.html;   http:
//www.nature.com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
46.pdf.

Shladover:2016:TAS

[2827] Steven E. Shladover. The truth about “self-driving” cars. Scientific American, 314(6):52–57, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
52.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-52.pdf.

Russell:2016:SWF

[2828] Stuart Russell. Should we fear supersmart robots? Scientific American, 314(6):58–59, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-

58.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-58.pdf.

Holland:2016:GWP

[2829] Andrew Holland. Global warming: Preventing Tomorrow’s climate wars. Scientific American, 314(6):60–65, June 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v314/
n6/full/scientificamerican0616-60.html;   http://www.nature.
com/scientificamerican/journal/v314/n6/pdf/scientificamerican0616-
60.pdf.

Balcombe:2016:ZES

[2830] Jonathan Balcombe. Zoology: Einstein of the sea. Scientific American, 314(6):66–69, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
66.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-66.pdf.

Schmidt:2016:HFA

[2831] Charles Schmidt. Health: The fog of Agent Orange. Scientific American, 314(6):70–75, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v314/n6/full/scientificamerican0616-
70.html;   http://www.nature.com/scientificamerican/journal/
v314/n6/pdf/scientificamerican0616-70.pdf.

Stover:2016:EYT

[2832] Dawn Stover. From the Editor: Your three-pound sex organ. Scientific American, 25(1s):1, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1s/full/scientificamericansex0316-
1.html;   http://www.nature.com/scientificamerican/journal/
v25/n1s/pdf/scientificamericansex0316-1.pdf.

Mirsky:2016:AGP

[2833] Steve Mirsky. Anti gravity: Pervspicacity. Scientific American, 25(1s): 112, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-112.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
112.pdf.

Portner:2016:LOM

[2834] Martin Portner. Lust: The orgasmic mind. Scientific American, 25(1s): 4–9, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-4.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
4.pdf.

Cacioppo:2016:LL

[2835] Stephanie Cacioppo and John T. Cacioppo. Lust for life. Scientific American, 25(1s):10–14, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1s/full/scientificamericansex0316-
10.html;   http://www.nature.com/scientificamerican/journal/
v25/n1s/pdf/scientificamericansex0316-10.pdf.

Sukel:2016:LR

[2836] Kayt Sukel. Lust’s reward. Scientific American, 25(1s):14–17, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-14.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
14.pdf.

Walter:2016:AL

[2837] Chip Walter. Affairs of the lips. Scientific American, 25(1s):18–23, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-18.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
18.pdf.

Moyer:2016:SSS

[2838] Melinda Wenner Moyer. The sunny side of smut. Scientific American, 25(1s):24–25, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1s/full/scientificamericansex0316-
24.html;   http://www.nature.com/scientificamerican/journal/
v25/n1s/pdf/scientificamericansex0316-24.pdf.

Bohannon:2016:WAA

[2839] Cat Bohannon. When arousal is agony. Scientific American, 25(1s):26– 33, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-26.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
26.pdf.

Fields:2016:SSN

[2840] R. Douglas Fields. Sex and the secret nerve. Scientific American, 25(1s):34–41, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1s/full/scientificamericansex0316-
34.html;   http://www.nature.com/scientificamerican/journal/
v25/n1s/pdf/scientificamericansex0316-34.pdf.

deWaal:2016:LBS

[2841] Frans B. M. de Waal. Leanings: Bonobo sex and society. Scientific American, 25(1s):42–49, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1s/full/scientificamericansex0316-
42.html;   http://www.nature.com/scientificamerican/journal/
v25/n1s/pdf/scientificamericansex0316-42.pdf.

Driscoll:2016:BS

[2842] Emily V. Driscoll. Bisexual species. Scientific American, 25(1s):50–55, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-50.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
50.pdf.

Epstein:2016:DGC

[2843] Robert Epstein. Do gays have a choice? Scientific American, 25(1s):56– 63, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-56.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
56.pdf.

Eliot:2016:TAB

[2844] Lise Eliot. The truth about boys and girls. Scientific American, 25(1s):64–71, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1s/full/scientificamericansex0316-

64.html;   http://www.nature.com/scientificamerican/journal/
v25/n1s/pdf/scientificamericansex0316-64.pdf.

Russo:2016:TK

[2845] Francine Russo. Transgender kids. Scientific American, 25(1s):72–79, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-72.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
72.pdf.

Edgar:2016:CPT

[2846] Blake Edgar. Coupling: Powers of two. Scientific American, 25(1s):80– 85, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-80.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
80.pdf.

Pawelski:2016:HC

[2847] Suzann Pileggi Pawelski. The happy couple. Scientific American, 25(1s):86–91, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1s/full/scientificamericansex0316-
86.html;   http://www.nature.com/scientificamerican/journal/
v25/n1s/pdf/scientificamericansex0316-86.pdf.

Tannen:2016:SSS

[2848] Deborah Tannen. He said, she said. Scientific American, 25(1s):92–97, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-92.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
92.pdf.

Flora:2016:JF

[2849] Carlin Flora. “Just friends”. Scientific American, 25(1s):98–103, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1s/full/scientificamericansex0316-98.html;   http:
//www.nature.com/scientificamerican/journal/v25/n1s/pdf/scientificamericansex0316-
98.pdf.

Finkel:2016:DDW

[2850] Eli J. Finkel, Paul W. Eastwick, Benjamin R. Karney, Harry T. Reis, and Susan Sprecher. Dating in a digital world. Scientific American, 25(1s):104–111, March 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1s/full/scientificamericansex0316-
104.html;   http://www.nature.com/scientificamerican/journal/
v25/n1s/pdf/scientificamericansex0316-104.pdf.

Gawrylewski:2016:ESS

[2851] Andrea Gawrylewski. From the Editor: Secrets of successful parenting. Scientific American, 25(2s):1, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
1.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-1.pdf.

Gopnik:2016:SCC

[2852] Alison Gopnik. Smart, curious, creative: How babies think. Scientific American, 25(2s):4–9, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
4.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-4.pdf.

Dweck:2016:SRS

[2853] Carol S. Dweck. The secret to raising smart kids. Scientific American, 25(2s):10–17, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
10.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-10.pdf.

Wickelgren:2016:CCM

[2854] Ingrid Wickelgren. Calisthenics for a child’s mind. Scientific American, 25(2s):18–25, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
18.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-18.pdf.

Stix:2016:HBB

[2855] Gary Stix. How to build a better learner. Scientific American, 25(2s):26– 33, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2s/full/scientificamericankids0616-26.html;   http:
//www.nature.com/scientificamerican/journal/v25/n2s/pdf/scientificamericankids0616-
26.pdf.

Blair:2016:TTL

[2856] Clancy Blair. Treating a toxin to learning. Scientific American, 25(2s):34–37, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
34.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-34.pdf.

Paul:2016:NVT

[2857] Annie Murphy Paul. A new vision for testing. Scientific American, 25(2s):38–45, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
38.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-38.pdf.

Dye:2016:WJC

[2858] Melody Dye. Why Johnny can’t name his colors. Scientific American, 25(2s):46–49, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
46.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-46.pdf.

Raeburn:2016:WD

[2859] Paul Raeburn. Where’s Dad? Scientific American, 25(2s):102–109, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2s/full/scientificamericankids0616-102.html;
http://www.nature.com/scientificamerican/journal/v25/n2s/pdf/
scientificamericankids0616-102.pdf.

Jacobson:2016:BYO

[2860] Roni Jacobson. Build your own family. Scientific American, 25(2s):106– 107, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2s/full/scientificamericankids0616-106.html;
http://www.nature.com/scientificamerican/journal/v25/n2s/pdf/
scientificamericankids0616-106.pdf.

Paul:2016:PSP

[2861] Annie Murphy Paul. The positive side of peer pressure. Scientific American, 25(2s):110–111, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
110.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-110.pdf.

Pogue:2016:TSG

[2862] David Pogue. The touch screen generation. Scientific American, 25(2s): 112, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2s/full/scientificamericankids0616-112.html;
http://www.nature.com/scientificamerican/journal/v25/n2s/pdf/
scientificamericankids0616-112.pdf.

Moyer:2016:HSR

[2863] Melinda Wenner Moyer. Happy, sociable, resilient: The serious need for play. Scientific American, 25(2s):50–57, June 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n2s/full/scientificamericankids0616-50.html;   http://www.nature.
com/scientificamerican/journal/v25/n2s/pdf/scientificamericankids0616-
50.pdf.

Epstein:2016:WMG

[2864] Robert Epstein. What makes a good parent? Scientific American, 25(2s):58–63, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
58.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-58.pdf.

Wickelgren:2016:OB

[2865] Ingrid Wickelgren. Oh, behave! Scientific American, 25(2s):64–69, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2s/full/scientificamericankids0616-64.html;   http:

//www.nature.com/scientificamerican/journal/v25/n2s/pdf/scientificamericankids0616-
64.pdf.

Lange:2016:HCA

[2866] Nicholas Lange and Christopher J. McDougle. Help for the child with autism. Scientific American, 25(2s):70–75, June 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n2s/full/scientificamericankids0616-70.html;   http://www.nature.
com/scientificamerican/journal/v25/n2s/pdf/scientificamericankids0616-
70.pdf.

Bubrick:2016:FC

[2867] Jerry Bubrick. Fear not, child. Scientific American, 25(2s):76–83, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2s/full/scientificamericankids0616-76.html;   http:
//www.nature.com/scientificamerican/journal/v25/n2s/pdf/scientificamericankids0616-
76.pdf.

Wallis:2016:BS

[2868] Claudia Wallis. Beyond shyness. Scientific American, 25(2s):84–89, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2s/full/scientificamericankids0616-84.html;   http:
//www.nature.com/scientificamerican/journal/v25/n2s/pdf/scientificamericankids0616-
84.pdf.

Arkowitz:2016:DBC

[2869] Hal Arkowitz and Scott O. Lilienfeld. Is divorce bad for children? Scientific American, 25(2s):90–91, June 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n2s/full/scientificamericankids0616-90.html;   http://www.nature.
com/scientificamerican/journal/v25/n2s/pdf/scientificamericankids0616-
90.pdf.

Giedd:2016:ASS

[2870] Jay N. Giedd. Adaptable, safe, self-reliant: The amazing teen brain. Scientific American, 25(2s):92–97, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n2s/full/scientificamericankids0616-

92.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-92.pdf.

Hopson:2016:BMT

[2871] Janet Hopson. Bad mix for the teen brain. Scientific American, 25(2s):98–101, June 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2s/full/scientificamericankids0616-
98.html;   http://www.nature.com/scientificamerican/journal/
v25/n2s/pdf/scientificamericankids0616-98.pdf.

DiChristina:2016:ESB

[2872] Mariette DiChristina. From the Editor: Serious benefits from games. Scientific American, 315(1):4, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
4.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-4.pdf.

Anonymous:2016:Lg

[2873] Anonymous. Letters. Scientific American, 315(1):5–6, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n1/full/scientificamerican0716-5.html;   http://www.nature.
com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
5.pdf.

Schlenoff:2016:YAIg

[2874] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 315(1):75, July 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n1/full/scientificamerican0716-75.html;   http://www.nature.
com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
75.pdf.

Anonymous:2016:SADb

[2875] Anonymous. Science agenda: Don’t let history die. Scientific American, 315(1):7, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-

7.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-7.pdf.

Neal:2016:FRM

[2876] Clive R. Neal. Forum: A return to the Moon is crucial. Scientific American, 315(1):8, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
8.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-8.pdf.

Billings:2016:AHF

[2877] Lee Billings. Advances: How to find another Earth. Scientific American, 315(1):11–13, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
11.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-11.pdf.

Landhuis:2016:LMT

[2878] Esther Landhuis. Lab mice are too clean. Scientific American, 315(1): 12–13, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-12.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
12.pdf.

Nuwer:2016:SS

[2879] Rachel Nuwer. Solar sleuths. Scientific American, 315(1):14, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n1/full/scientificamerican0716-14.html;   http://www.nature.
com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
14.pdf.

Hackett:2016:LDL

[2880] Jennifer Hackett. Lights don’t lure urbanite moths. Scientific American, 315(1):15, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
15.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-15.pdf.

Maron:2016:ZGC

[2881] Dina Fine Maron. Zika’s global connections. Scientific American, 315(1): 16, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-16.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
16.pdf.

Hackett:2016:WP

[2882] Jennifer Hackett. Water power. Scientific American, 315(1):17, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-17.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
17.pdf.

Anonymous:2016:QHg

[2883] Anonymous. Quick hits. Scientific American, 315(1):18, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n1/full/scientificamerican0716-18.html;   http://www.nature.
com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
18.pdf.

Bourzac:2016:SD

[2884] Katherine Bourzac. Skin deep. Scientific American, 315(1):19, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n1/full/scientificamerican0716-19.html;   http://www.nature.
com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
19.pdf.

Lufkin:2016:JJ

[2885] Bryan Lufkin. Journey to Jupiter. Scientific American, 315(1):20, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-20.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
20.pdf.

Jacobson:2016:MC

[2886] Roni Jacobson. Morbid curiosity. Scientific American, 315(1):21, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-21.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
21.pdf.

Chain:2016:CMC

[2887] Lydia Chain. A coat of many colors. Scientific American, 315(1):22, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-22.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
22.pdf.

Shell:2016:SHG

[2888] Ellen Ruppel Shell. The science of health: Gauging the effects of lead. Scientific American, 315(1):23–24, July 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n1/full/scientificamerican0716-23.html;   http://www.nature.
com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
23.pdf.

Pogue:2016:TYC

[2889] David Pogue. TechnoFiles: At your command. Scientific American, 315 (1):25, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-25.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
25.pdf.

Anonymous:2016:RRRc

[2890] Anonymous. Reviews and recommendations: Recommended: Future tense. Scientific American, 315(1):72, July 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n1/full/scientificamerican0716-72.html;   http://www.nature.
com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
72.pdf.

Shermer:2016:STB

[2891] Michael Shermer. Skeptic: Tiger blood and goat milk. Scientific American, 315(1):73, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
73.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-73.pdf.

Mirsky:2016:AGCa

[2892] Steve Mirsky. Anti gravity: Clever creatures. Scientific American, 315 (1):74, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-74.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
74.pdf.

Fischetti:2016:GSO

[2893] Mark Fischetti. Graphic science: One-world menu. Scientific American, 315(1):76, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
76.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-76.pdf.

Bavelier:2016:CSB

[2894] Daphne Bavelier and C. Shawn Green. Cognitive science: The brainboosting power of video games. Scientific American, 315(1):26–31, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-26.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
26.pdf.

Libeskind:2016:AOPa

[2895] Noam I. Libeskind and R. Brent Tully. Astronomy: Our place in the Cosmos. Scientific American, 315(1):32–39, July 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n1/full/scientificamerican0716-32.html;   http://www.nature.
com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
32.pdf.

Yasmin:2016:PHE

[2896] Seema Yasmin. Public health: Ebola’s second coming. Scientific American, 315(1):40–45, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-

40.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-40.pdf.

Kuchment:2016:EDE

[2897] Anna Kuchment. Environment: Drilling for earthquakes. Scientific American, 315(1):46–53, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
46.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-46.pdf.

Karanth:2016:CTT

[2898] K. Ullas Karanth. Conservation: Tracking tigers. Scientific American, 315(1):54–59, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
54.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-54.pdf.

Pavlus:2016:RBM

[2899] John Pavlus. Robotics: Bipedal metal. Scientific American, 315(1):60– 67, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n1/full/scientificamerican0716-60.html;   http:
//www.nature.com/scientificamerican/journal/v315/n1/pdf/scientificamerican0716-
60.pdf.

Harris:2016:ASS

[2900] Mark Harris. Agriculture: Saltwater solution. Scientific American, 315(1):68–71, July 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n1/full/scientificamerican0716-
68.html;   http://www.nature.com/scientificamerican/journal/
v315/n1/pdf/scientificamerican0716-68.pdf.

DiChristina:2016:ETH

[2901] Mariette DiChristina. From the Editor: The treasure hunt is on. Scientific American, 315(2):4, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
4.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-4.pdf.

Anonymous:2016:Lh

[2902] Anonymous. Letters. Scientific American, 315(2):6–7, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n2/full/scientificamerican0816-6.html;   http://www.nature.
com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
6.pdf.

Schlenoff:2016:YAIh

[2903] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 315(2):78, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n2/full/scientificamerican0816-78.html;   http://www.nature.
com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
78.pdf.

Anonymous:2016:SAO

[2904] Anonymous. Science agenda: Olympic officials should tell women “high T” is no hurdle. Scientific American, 315(2):8, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n2/full/scientificamerican0816-8.html;   http://www.nature.
com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
8.pdf.

Hartung:2016:FLE

[2905] Thomas Hartung. Forum: The lesser evil of e-cigarettes. Scientific American, 315(2):9, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
9.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-9.pdf.

Biello:2016:AGR

[2906] David Biello. Advances: Genetic resurrection. Scientific American, 315(2):10–13, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
10.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-10.pdf.

Brouillette:2016:NAS

[2907] Monique Brouillette. The new allergy shot. Scientific American, 315(2):12–13, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
12.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-12.pdf.

Gorman:2016:B

[2908] Christine Gorman. $1.4 billion. Scientific American, 315(2):14, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-14a.html;   http:
//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
14a.pdf.

Crow:2016:DC

[2909] Diana Crow. Divide and conquer. Scientific American, 315(2):14–15, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-14b.html;   http:
//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
14b.pdf.

Cepelewicz:2016:YBP

[2910] Jordana Cepelewicz. Your brain on physics. Scientific American, 315(2): 15, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-15.html;   http:
//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
15.pdf.

Hackett:2016:MA

[2911] Jennifer Hackett. Model airplane. Scientific American, 315(2):16, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-16a.html;   http:
//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
16a.pdf.

Gholipour:2016:CWL

[2912] Bahar Gholipour. Can we learn how to forget? Scientific American, 315(2):16–17, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
16b.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-16b.pdf.

Anonymous:2016:QHh

[2913] Anonymous. Quick hits. Scientific American, 315(2):18, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n2/full/scientificamerican0816-18.html;   http://www.nature.
com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
18.pdf.

Sheikh:2016:ZPO

[2914] Knvul Sheikh. Zika-proofing the Olympics. Scientific American, 315(2): 19, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-19.html;   http:
//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
19.pdf.

Biello:2016:BB

[2915] David Biello. Bionic is better. Scientific American, 315(2):20, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-20.html;   http:
//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
20.pdf.

Godkin:2016:TBW

[2916] David Godkin. Toothless beaks for the win. Scientific American, 315(2): 21, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-21.html;   http:
//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
21.pdf.

Patel:2016:WW

[2917] Prachi Patel. Waste not, want not. Scientific American, 315(2):22, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-22.html;   http:

//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
22.pdf.

Interlandi:2016:SHA

[2918] Jeneen Interlandi. The science of health: The acupuncture myth. Scientific American, 315(2):24–25, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v315/n2/full/scientificamerican0816-
24.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-24.pdf.

Pogue:2016:TEP

[2919] David Pogue. TechnoFiles: The end of passwords. Scientific American, 315(2):26, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
26.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-26.pdf.

Anonymous:2016:RRRd

[2920] Anonymous. Reviews and recommendations: Recommended. Scientific American, 315(2):74, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
74.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-74.pdf.

Shermer:2016:SQG

[2921] Michael Shermer. Skeptic: The quack of the gaps problem. Scientific American, 315(2):75, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
75.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-75.pdf.

Mirsky:2016:AGW

[2922] Steve Mirsky. Anti gravity: The war by science. Scientific American, 315(2):76, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
76.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-76.pdf.

Fischetti:2016:GSF

[2923] Mark Fischetti. Graphic science: Fatter still. Scientific American, 315(2): 80, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n2/full/scientificamerican0816-80.html;   http:
//www.nature.com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
80.pdf.

Szapudi:2016:CEPa

[2924] István Szapudi. Cosmology: The emptiest place in space. Scientific American, 315(2):28–35, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
28.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-28.pdf.

Gibbs:2016:EHC

[2925] W. Wayt Gibbs. Epidemiology: Health check for humanity. Scientific American, 315(2):36–41, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
36.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-36.pdf.

Paul:2016:SRE

[2926] Annie Murphy Paul. Special report on education: The coding revolution. Scientific American, 315(2):42–49, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n2/full/scientificamerican0816-42.html;   http://www.nature.
com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
42.pdf.

Frederick:2016:TPH

[2927] Jennifer Frederick. Teach the Prof how to teach. Scientific American, 315(2):46–47, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
46.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-46.pdf.

Maron:2016:BSS

[2928] Dina Fine Maron. Biomechanics: The secret to speed. Scientific American, 315(2):50–55, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
50.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-50.pdf.

Zuccato:2016:EHP

[2929] Chiara Zuccato and Elena Cattaneo. Evolution: The Huntington’s paradox. Scientific American, 315(2):56–61, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n2/full/scientificamerican0816-56.html;   http://www.nature.
com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
56.pdf.

Lauretta:2016:PSS

[2930] Dante S. Lauretta. Planetary science: The seven-year mission to fetch 60 grams of asteroid. Scientific American, 315(2):62–69, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n2/full/scientificamerican0816-62.html;   http://www.nature.
com/scientificamerican/journal/v315/n2/pdf/scientificamerican0816-
62.pdf.

Wilcox:2016:BZN

[2931] Christie Wilcox. Biology: Zombie neuroscience. Scientific American, 315(2):70–73, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n2/full/scientificamerican0816-
70.html;   http://www.nature.com/scientificamerican/journal/
v315/n2/pdf/scientificamerican0816-70.pdf.

DiChristina:2016:ELF

[2932] Mariette DiChristina. From the Editor: Looking forward. Scientific American, 315(3):4, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
4.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-4.pdf.

Anonymous:2016:Li

[2933] Anonymous. Letters. Scientific American, 315(3):5–6, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n3/full/scientificamerican0916-5.html;   http://www.nature.
com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
5.pdf.

Anonymous:2016:SAT

[2934] Anonymous. Science agenda: The tweets we hold to be selfevident. Scientific American, 315(3):7, September 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n3/full/scientificamerican0916-7.html;   http://www.nature.
com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
7.pdf.

Davies:2016:FMP

[2935] Paul Davies. Forum: Many planets, not much life. Scientific American, 315(3):8, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
8.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-8.pdf.

Hsu:2016:ACM

[2936] Jeremy Hsu. Advances: Can medical cannabis break the painkiller epidemic? Scientific American, 315(3):10–12, September 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n3/full/scientificamerican0916-10.html;   http://www.nature.
com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
10.pdf.

Collins:2016:MMS

[2937] Nathan Collins. Mirror molecules in space. Scientific American, 315(3):14, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
14.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-14.pdf.

Caldwell:2016:FT

[2938] Noah Caldwell. A flood of tears. Scientific American, 315(3):15, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n3/full/scientificamerican0916-15.html;   http:
//www.nature.com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
15.pdf.

Pycha:2016:ACA

[2939] Anne Pycha. An auditory component to autism. Scientific American, 315(3):16, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
16a.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-16a.pdf.

Smith:2016:BLS

[2940] Peter Andrey Smith. Blood left at scene can indicate criminal age. Scientific American, 315(3):16, September 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n3/full/scientificamerican0916-16b.html;   http://www.nature.
com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
16b.pdf.

Sheikh:2016:KM

[2941] Knvul Sheikh. The kilogram makeover. Scientific American, 315 (3):18–19, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
18.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-18.pdf.

Greenspan:2016:CIN

[2942] Jesse Greenspan. Cleanup on Isle Nine. Scientific American, 315(3):20, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n3/full/scientificamerican0916-20.html;   http:
//www.nature.com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
20.pdf.

Stix:2016:WWA

[2943] Gary Stix. Where will all the new neuroscientists go? Scientific American, 315(3):21, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
21.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-21.pdf.

Anonymous:2016:QHi

[2944] Anonymous. Quick hits. Scientific American, 315(3):22, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n3/full/scientificamerican0916-22.html;   http://www.nature.
com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
22.pdf.

Jacobson:2016:WAS

[2945] Roni Jacobson. We all speed-read. Scientific American, 315(3):23, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n3/full/scientificamerican0916-23.html;   http:
//www.nature.com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
23.pdf.

Wapner:2016:SHC

[2946] Jessica Wapner. The science of health: Cancer gene tests provide few answers. Scientific American, 315(3):24–25, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n3/full/scientificamerican0916-24.html;   http://www.nature.
com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
24.pdf.

Pogue:2016:TWD

[2947] David Pogue. TechnoFiles: What to do with all those cassettes. Scientific American, 315(3):27, September 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n3/full/scientificamerican0916-27.html;   http://www.nature.
com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
27.pdf.

Anonymous:2016:RRRe

[2948] Anonymous. Reviews and recommendations: Recommended. Scientific American, 315(3):86, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
86.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-86.pdf.

Shermer:2016:SBK

[2949] Michael Shermer. Skeptic: At the boundary of knowledge. Scientific American, 315(3):88, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
88.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-88.pdf.

Mirsky:2016:AGFb

[2950] Steve Mirsky. Anti gravity: In favor of feline felicity. Scientific American, 315(3):90, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
90.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-90.pdf.

Schlenoff:2016:YAIi

[2951] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 315(3):91, September 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n3/full/scientificamerican0916-91.html;   http://www.nature.
com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
91.pdf.

Peek:2016:GSU

[2952] Katie Peek. Graphic science: Urban wealth. Scientific American, 315(3):92, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
92.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-92.pdf.

Anonymous:2016:HE

[2953] Anonymous. The human experiment. Scientific American, 315(3):28–29, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n3/full/scientificamerican0916-28.html;   http:
//www.nature.com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
28.pdf.

Robinson:2016:GU

[2954] Kim Stanley Robinson. The great unknown. Scientific American, 315(3):80–85, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
80.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-80.pdf.

Zalasiewicz:2016:GHL

[2955] Jan Zalasiewicz. Geology: A history in layers. Scientific American, 315(3):30–37, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
30.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-30.pdf.

Peek:2016:ECS

[2956] Katie Peek. Environment: Coping skills. Scientific American, 315(3):38–41, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
38.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-38.pdf.

Hvistendahl:2016:PTT

[2957] Mara Hvistendahl. Population: A tale of two worlds. Scientific American, 315(3):42–47, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
42.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-42.pdf.

Deaton:2016:ETI

[2958] Angus Deaton. Economics: The threat of inequality. Scientific American, 315(3):48–53, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
48.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-48.pdf.

Hall:2016:BRLa

[2959] Stephen S. Hall. Biotechnology: The red line. Scientific American, 315(3):54–61, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
54.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-54.pdf.

Gifford:2016:HL

[2960] Bill Gifford. Health: Living to 120. Scientific American, 315(3):62–69, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n3/full/scientificamerican0916-62.html;   http:
//www.nature.com/scientificamerican/journal/v315/n3/pdf/scientificamerican0916-
62.pdf.

Rosner:2016:RAT

[2961] Hillary Rosner. Robotics: All too human. Scientific American, 315(3):70–75, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
70.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-70.pdf.

Grinspoon:2016:SDT

[2962] David Grinspoon. Space: Deep time deep survival. Scientific American, 315(3):76–79, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n3/full/scientificamerican0916-
76.html;   http://www.nature.com/scientificamerican/journal/
v315/n3/pdf/scientificamerican0916-76.pdf.

Guynup:2016:ESB

[2963] Sharon Guynup. From the Editor: Is seeing believing? Scientific American, 25(3s):1, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3s/full/scientificamericanillusions0816-

1.html;   http://www.nature.com/scientificamerican/journal/
v25/n3s/pdf/scientificamericanillusions0816-1.pdf.

Martinez-Conde:2016:IWI

[2964] Susana Martinez-Conde and Stephen L. Macknik. Introduction: What it means to be you. Scientific American, 25(3s):4–5, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-4.html;   http://www.
nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
4.pdf.

Martinez-Conde:2016:FMN

[2965] Susana Martinez-Conde and Stephen L. Macknik. Foundations of misperception: The neuroscience of illusion. Scientific American, 25(3s):6–9, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3s/full/scientificamericanillusions0816-6.html;
 http://www.nature.com/scientificamerican/journal/v25/n3s/pdf/
scientificamericanillusions0816-6.pdf.

Martinez-Conde:2016:FR

[2966] Susana Martinez-Conde and Stephen L. Macknik. A faithful resemblance. Scientific American, 25(3s):18–21, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-18.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
18.pdf.

Macknik:2016:FMP

[2967] Stephen L. Macknik and Susana Martinez-Conde. Foundations of misperception: A perspective on 3-d visual illusions. Scientific American, 25(3s):10–13, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3s/full/scientificamericanillusions0816-
10.html;   http://www.nature.com/scientificamerican/journal/
v25/n3s/pdf/scientificamericanillusions0816-10.pdf.

Martinez-Conde:2016:ADG

[2968] Susana Martinez-Conde and Stephen L. Macknik. All deceptions great and small. Scientific American, 25(3s):14–17, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-14.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
14.pdf.

Macknik:2016:YTL

[2969] Stephen L. Macknik and Susana Martinez-Conde. Your twisted little mind. Scientific American, 25(3s):22–25, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-22.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
22.pdf.

Martinez-Conde:2016:VT

[2970] Susana Martinez-Conde and Stephen L. Macknik. Vanished without a trace. Scientific American, 25(3s):26–29, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-26.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
26.pdf.

Macknik:2016:YGO

[2971] Stephen L. Macknik and Susana Martinez-Conde. Yorick’s ghost and other afterimages. Scientific American, 25(3s):30–33, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-30.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
30.pdf.

Martinez-Conde:2016:CS

[2972] Susana Martinez-Conde and Stephen L. Macknik. Colors out of space. Scientific American, 25(3s):34–41, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-34.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
34.pdf.

Macknik:2016:AD

[2973] Stephen L. Macknik, Susana Martinez-Conde, and Ellis C. Gayles. The aviator’s dilemma. Scientific American, 25(3s):42–45, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-42.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
42.pdf.

Macknik:2016:AS

[2974] Stephen L. Macknik and Susana Martinez-Conde. Afraid of shadows. Scientific American, 25(3s):46–49, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n3s/full/scientificamericanillusions0816-
46.html;   http://www.nature.com/scientificamerican/journal/
v25/n3s/pdf/scientificamericanillusions0816-46.pdf.

Macknik:2016:NBN

[2975] Stephen L. Macknik and Susana Martinez-Conde. No brain, no pain. Scientific American, 25(3s):50–53, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n3s/full/scientificamericanillusions0816-
50.html;   http://www.nature.com/scientificamerican/journal/
v25/n3s/pdf/scientificamericanillusions0816-50.pdf.

Martinez-Conde:2016:SR

[2976] Susana Martinez-Conde, Leandro L. Di Stasi, and Stephen L. Macknik. Seeing red. Scientific American, 25(3s):54–57, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-54.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
54.pdf.

Martinez-Conde:2016:GEO

[2977] Susana Martinez-Conde and Stephen L. Macknik. Gazing at each other: What’s in a face? Scientific American, 25(3s):58–67, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-58.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
58.pdf.

Martinez-Conde:2016:EI

[2978] Susana Martinez-Conde and Stephen L. Macknik. The eyes have it. Scientific American, 25(3s):68–73, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n3s/full/scientificamericanillusions0816-
68.html;   http://www.nature.com/scientificamerican/journal/
v25/n3s/pdf/scientificamericanillusions0816-68.pdf.

Macknik:2016:IL

[2979] Stephen L. Macknik and Susana Martinez-Conde. The illusions of love. Scientific American, 25(3s):74–79, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-74.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
74.pdf.

Martinez-Conde:2016:AIA

[2980] Susana Martinez-Conde and Stephen L. Macknik. Art and illusion: Art as visual research: Kinetic illusions in op art. Scientific American, 25(3s):80–87, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3s/full/scientificamericanillusions0816-
80.html;   http://www.nature.com/scientificamerican/journal/
v25/n3s/pdf/scientificamericanillusions0816-80.pdf.

Macknik:2016:SIS

[2981] Stephen L. Macknik and Susana Martinez-Conde. Sculpting the impossible: Solid renditions of visual illusions. Scientific American, 25(3s):88– 93, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3s/full/scientificamericanillusions0816-88.html;
 http://www.nature.com/scientificamerican/journal/v25/n3s/pdf/
scientificamericanillusions0816-88.pdf.

Martinez-Conde:2016:TTT

[2982] Susana Martinez-Conde and Stephen L. Macknik. Truer than truth. Scientific American, 25(3s):94–97, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-94.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
94.pdf.

Martinez-Conde:2016:UI

[2983] Susana Martinez-Conde and Stephen L. Macknik. Urban illusions. Scientific American, 25(3s):98–101, August 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3s/full/scientificamericanillusions0816-98.html;   http://
www.nature.com/scientificamerican/journal/v25/n3s/pdf/scientificamericanillusions0816-
98.pdf.

Macknik:2016:PTA

[2984] Stephen L. Macknik, Leandro L. Di Stasi, and Susana Martinez-Conde. Perfectly timed advertising. Scientific American, 25(3s):102–105, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3s/full/scientificamericanillusions0816-102.html;
 http://www.nature.com/scientificamerican/journal/v25/n3s/pdf/
scientificamericanillusions0816-102.pdf.

Martinez-Conde:2016:FTV

[2985] Susana Martinez-Conde and Stephen L. Macknik. Food for thought: Visual illusions good enough to eat. Scientific American, 25(3s):106–112, August 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3s/full/scientificamericanillusions0816-106.html;
 http://www.nature.com/scientificamerican/journal/v25/n3s/pdf/
scientificamericanillusions0816-106.pdf.

DiChristina:2016:ENV

[2986] Mariette DiChristina. From the Editor: New views: Brains, sciences, oceans. Scientific American, 315(4):6, October 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n4/full/scientificamerican1016-6.html;   http://www.nature.
com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
6.pdf.

Anonymous:2016:Lj

[2987] Anonymous. Letters. Scientific American, 315(4):8–9, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n4/full/scientificamerican1016-8.html;   http://www.nature.
com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
8.pdf.

Schlenoff:2016:YAIj

[2988] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 315(4):86, October 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n4/full/scientificamerican1016-86.html;   http://www.nature.
com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
86.pdf.

Anonymous:2016:SASb

[2989] Anonymous. Science agenda: Science is not enough. Scientific American, 315(4):12, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
12.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-12.pdf.

Hart:2016:FOS

[2990] Carl Hart. Forum: Our senseless pot laws. Scientific American, 315(4): 13, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n4/full/scientificamerican1016-13.html;   http:
//www.nature.com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
13.pdf.

Billings:2016:ADA

[2991] Lee Billings. Advances: In the dark about dark matter. Scientific American, 315(4):14–16, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
14.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-14.pdf.

Weintraub:2016:SC

[2992] Karen Weintraub. Send in the clones. Scientific American, 315(4):17, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n4/full/scientificamerican1016-17.html;   http:
//www.nature.com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
17.pdf.

Hsu:2016:NWE

[2993] Jeremy Hsu. New weather eye in the sky. Scientific American, 315(4):18, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n4/full/scientificamerican1016-18.html;   http:
//www.nature.com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
18.pdf.

Yuhas:2016:DDW

[2994] Daisy Yuhas. Depressed? Do what you love. Scientific American, 315(4): 19, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n4/full/scientificamerican1016-19.html;   http:
//www.nature.com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
19.pdf.

Sheikh:2016:BA

[2995] Knvul Sheikh. A boost from above. Scientific American, 315(4):20, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n4/full/scientificamerican1016-20.html;   http:
//www.nature.com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
20.pdf.

Sheikh:2016:BOM

[2996] Knvul Sheikh. The breast has its own microbiome. Scientific American, 315(4):21, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
21.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-21.pdf.

Anonymous:2016:QHj

[2997] Anonymous. Quick hits. Scientific American, 315(4):22, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n4/full/scientificamerican1016-22.html;   http://www.nature.
com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
22.pdf.

Goldman:2016:TDD

[2998] Jason G. Goldman. Taking direction from a deer. Scientific American, 315(4):23, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
23.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-23.pdf.

Gannon:2016:GBI

[2999] Megan Gannon. Good-bye ice, good-bye data. Scientific American, 315(4):24, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
24.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-24.pdf.

Noonan:2016:SHB

[3000] David Noonan. The science of health: Busting blood clots. Scientific American, 315(4):26–27, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
26.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-26.pdf.

Pogue:2016:TBS

[3001] David Pogue. TechnoFiles: The bright side of Internet shaming. Scientific American, 315(4):28, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
28.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-28.pdf.

Anonymous:2016:RRRf

[3002] Anonymous. Reviews and recommendations: Recommended. Scientific American, 315(4):82, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
82.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-82.pdf.

Shermer:2016:SMM

[3003] Michael Shermer. Skeptic: Methuselah’s Moon shot. Scientific American, 315(4):84, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-

84.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-84.pdf.

Mirsky:2016:AGCb

[3004] Steve Mirsky. Anti gravity: Current events. Scientific American, 315(4): 85, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n4/full/scientificamerican1016-85.html;   http:
//www.nature.com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
85.pdf.

Fischetti:2016:GSP

[3005] Mark Fischetti. Graphic science: Pushed out. Scientific American, 315(4):88, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
88.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-88.pdf.

Deisseroth:2016:NLI

[3006] Karl Deisseroth. Neuroscience: A look inside the brain. Scientific American, 315(4):30–37, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
30.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-30.pdf.

Postberg:2016:PSU

[3007] Frank Postberg, Gabriel Tobie, and Thorsten Dambeck. Planetary science: Under the sea of Enceladus. Scientific American, 315(4):38–45, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n4/full/scientificamerican1016-38.html;   http:
//www.nature.com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
38.pdf.

Maron:2016:HRP

[3008] Dina Fine Maron. Health: The right pill for you. Scientific American, 315(4):46–51, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
46.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-46.pdf.

Anonymous:2016:SWS

[3009] Anonymous. State of the world’s science 2016: The truth brokers. Scientific American, 315(4):52–53, October 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n4/full/scientificamerican1016-52.html;   http://www.nature.
com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
52.pdf.

Seife:2016:HSS

[3010] Charles Seife. How to spin the science news. Scientific American, 315(4):54–61, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
54.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-54.pdf.

Anonymous:2016:WSB

[3011] Anonymous. What science is buzzworthy? Scientific American, 315(4):62–63, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
62.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-62.pdf.

Martinez-Conde:2016:PCS

[3012] Susana Martinez-Conde, Stephen L. Macknik, and Devin Powell. The plight of the celebrity scientist. Scientific American, 315(4):64–67, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n4/full/scientificamerican1016-64.html;   http:
//www.nature.com/scientificamerican/journal/v315/n4/pdf/scientificamerican1016-
64.pdf.

Becker:2016:WTN

[3013] Emily Becker. Weather: On the trail of El Niño. Scientific American, 315(4):68–75, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
68.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-68.pdf.

Conniff:2016:CHR

[3014] Richard Conniff. Conservation: Home off the range. Scientific American, 315(4):76–81, October 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n4/full/scientificamerican1016-
76.html;   http://www.nature.com/scientificamerican/journal/
v315/n4/pdf/scientificamerican1016-76.pdf.

Wong:2016:EHJ

[3015] Kate Wong. From the Editor: Humanity’s journey. Scientific American, 25(4s):1, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
1.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-1.pdf.

Anonymous:2016:SUW

[3016] Anonymous. The story of us: Where we came from. Scientific American, 25(4s):4–5, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
4.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-4.pdf.

Wood:2016:WF

[3017] Bernard Wood. Welcome to the family. Scientific American, 25(4s):6–11, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-6.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-6.pdf.

Wong:2016:MHa

[3018] Kate Wong. Mystery human. Scientific American, 25(4s):12–21, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-12.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-12.pdf.

deMenocal:2016:CS

[3019] Peter B. deMenocal. Climate shocks. Scientific American, 25(4s):22–27, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-22.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-22.pdf.

Stout:2016:TSA

[3020] Dietrich Stout. Tales of a stone age neuroscientist. Scientific American, 25(4s):28–35, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
28.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-28.pdf.

Marean:2016:WSS

[3021] Curtis W. Marean. When the sea saved humanity. Scientific American, 25(4s):36–43, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
36.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-36.pdf.

Hammer:2016:HH

[3022] Michael F. Hammer. Human hybrids. Scientific American, 25(4s):44–49, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-44.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-44.pdf.

Anonymous:2016:WMU

[3023] Anonymous. What makes us special. Scientific American, 25(4s):50–51, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-50.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-50.pdf.

Jablonski:2016:NT

[3024] Nina G. Jablonski. The naked truth. Scientific American, 25(4s):52–59, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-52.html;

 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-52.pdf.

Stix:2016:IF

[3025] Gary Stix. The “it” factor. Scientific American, 25(4s):60–65, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-60.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-60.pdf.

Wong:2016:NM

[3026] Kate Wong. Neandertal minds. Scientific American, 25(4s):66–73, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-66.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-66.pdf.

Edgar:2016:PT

[3027] Blake Edgar. Powers of two. Scientific American, 25(4s):74–79, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-74.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-74.pdf.

Caspari:2016:EG

[3028] Rachel Caspari. The evolution of grandparents. Scientific American, 25(4s):80–85, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
80.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-80.pdf.

Pringle:2016:OC

[3029] Heather Pringle. The origins of creativity. Scientific American, 25(4s):86–93, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
86.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-86.pdf.

Anonymous:2016:WWG

[3030] Anonymous. Where we are going. Scientific American, 25(4s):94–95, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-94.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-94.pdf.

Fischetti:2016:NP

[3031] Mark Fischetti. The networked primate. Scientific American, 25 (4s):96–99, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
96.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-96.pdf.

Hawks:2016:SEA

[3032] John Hawks. Still evolving (after all these years). Scientific American, 25(4s):100–105, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
100.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-100.pdf.

Smith:2016:SH

[3033] Cameron M. Smith. Starship humanity. Scientific American, 25 (4s):106–111, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4s/full/scientificamericanhumanity0916-
106.html;   http://www.nature.com/scientificamerican/journal/
v25/n4s/pdf/scientificamericanhumanity0916-106.pdf.

Wong:2016:PD

[3034] Kate Wong. The 1 percent difference. Scientific American, 25(4s):112, September 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4s/full/scientificamericanhumanity0916-112.html;
 http://www.nature.com/scientificamerican/journal/v25/n4s/pdf/
scientificamericanhumanity0916-112.pdf.

DiChristina:2016:ETT

[3035] Mariette DiChristina. From the Editor: Theory and truth. Scientific American, 315(5):4, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
4.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-4.pdf.

Anonymous:2016:Lk

[3036] Anonymous. Letters. Scientific American, 315(5):6–7, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n5/full/scientificamerican1116-6.html;   http://www.nature.
com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
6.pdf.

Schlenoff:2016:YAIk

[3037] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 315(5):79, November 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n5/full/scientificamerican1116-79.html;   http://www.nature.
com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
79.pdf.

Anonymous:2016:SADc

[3038] Anonymous. Science agenda: Donald Trump’s campaign for science illiteracy. Scientific American, 315(5):8, November 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n5/full/scientificamerican1116-8.html;   http://www.nature.
com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
8.pdf.

Wen:2016:FAR

[3039] Leana S. Wen and M. Cooper Lloyd. Forum: Attacking the roots of violence. Scientific American, 315(5):9, November 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n5/full/scientificamerican1116-9.html;   http://www.nature.
com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
9.pdf.

Goudarzi:2016:AWL

[3040] Sara Goudarzi. Advances: What lies beneath. Scientific American, 315(5):11–12, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
11.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-11.pdf.

Long:2016:TT

[3041] Kat Long. Tension in taxonomy. Scientific American, 315(5):13, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n5/full/scientificamerican1116-13.html;   http:
//www.nature.com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
13.pdf.

Sheng:2016:APG

[3042] Ellen Sheng. The artificial pancreas gets real. Scientific American, 315(5):14, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
14.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-14.pdf.

Garcia:2016:LET

[3043] Eduardo Garcia. Laziness: An evolutionary trait. Scientific American, 315(5):15, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
15.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-15.pdf.

Smith:2016:LL

[3044] Karl J. P. Smith. Lunar landscaping. Scientific American, 315(5):16, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n5/full/scientificamerican1116-16.html;   http:
//www.nature.com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
16.pdf.

Hsu:2016:MR

[3045] Jeremy Hsu. Mosquitoes to the rescue. Scientific American, 315(5):17, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n5/full/scientificamerican1116-17.html;   http:
//www.nature.com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
17.pdf.

Anonymous:2016:QHk

[3046] Anonymous. Quick hits. Scientific American, 315(5):18, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n5/full/scientificamerican1116-18.html;   http://www.nature.
com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
18.pdf.

Sheikh:2016:IAP

[3047] Knvul Sheikh. Introducing the acoustic prism. Scientific American, 315(5):19, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
19.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-19.pdf.

Platt:2016:HWT

[3048] John R. Platt. Home is where they make it. Scientific American, 315(5):20, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
20.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-20.pdf.

Hsu:2016:WC

[3049] Jeremy Hsu. Winds of change. Scientific American, 315(5):21, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n5/full/scientificamerican1116-21.html;   http:
//www.nature.com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
21.pdf.

Anonymous:2016:SE

[3050] Anonymous. Science in the elections. Scientific American, 315(5):22, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n5/full/scientificamerican1116-22.html;   http:

//www.nature.com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
22.pdf.

Pogue:2016:TPT

[3051] David Pogue. TechnoFiles: The problem with tech copycats. Scientific American, 315(5):23, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
23.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-23.pdf.

Jabr:2016:SHB

[3052] Ferris Jabr. The science of health: Blue light blues. Scientific American, 315(5):24–25, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
24.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-24.pdf.

Moskowitz:2016:RRR

[3053] Clara Moskowitz and Knvul Sheikh. Reviews and recommendations: Recommended. Scientific American, 315(5):76, November 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n5/full/scientificamerican1116-76.html;   http://www.nature.
com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
76.pdf.

Shermer:2016:SWG

[3054] Michael Shermer. Skeptic: Why gloom trumps glad. Scientific American, 315(5):77, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
77.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-77.pdf.

Mirsky:2016:AGM

[3055] Steve Mirsky. Anti gravity: Muddying the Waters. Scientific American, 315(5):78, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
78.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-78.pdf.

Fischetti:2016:GSM

[3056] Mark Fischetti. Graphic science: The many ways to innovate. Scientific American, 315(5):80, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
80.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-80.pdf.

Maldacena:2016:TPB

[3057] Juan Maldacena. Theoretical physics: Black holes and wormholes and the secrets of quantum spacetime. Scientific American, 315(5):26–31, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n5/full/scientificamerican1116-26.html;   http:
//www.nature.com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
26.pdf.

Belmonte:2016:BHO

[3058] Juan Carlos Izpisúa Belmonte. Biology: Human organs from animal bodies. Scientific American, 315(5):32–37, November 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n5/full/scientificamerican1116-32.html;   http://www.nature.
com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
32.pdf.

Gibbs:2016:EFU

[3059] W. Wayt Gibbs. Energy: The fusion underground. Scientific American, 315(5):38–45, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
38.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-38.pdf.

Shermer:2016:KTW

[3060] Michael Shermer, Harriet Hall, Ray Pierrehumbert, Paul Offit, and Seth Shostak. Knowledge: 5 things we know to be true. Scientific American, 315(5):46–53, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
46.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-46.pdf.

Riesch:2016:ESM

[3061] Rüdiger Riesch. Evolution: Species in the making. Scientific American, 315(5):54–61, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
54.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-54.pdf.

Nestor:2016:NGC

[3062] James Nestor. Neuroscience: Get clean or die trying. Scientific American, 315(5):62–69, November 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n5/full/scientificamerican1116-
62.html;   http://www.nature.com/scientificamerican/journal/
v315/n5/pdf/scientificamerican1116-62.pdf.

Ibbotson:2016:LLN

[3063] Paul Ibbotson and Michael Tomasello. Linguistics: Language in a new key. Scientific American, 315(5):70–75, November 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n5/full/scientificamerican1116-70.html;   http://www.nature.
com/scientificamerican/journal/v315/n5/pdf/scientificamerican1116-
70.pdf.

Gawrylewski:2016:EYS

[3064] Andrea Gawrylewski. From the Editor: The year in science. Scientific American, 25(5s):1, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
1.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-1.pdf.

Anonymous:2016:CW

[3065] Anonymous. Cosmic wonders. Scientific American, 25(5s):4–5, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5s/full/scientificamericansciencestories1216-4.
html;   http://www.nature.com/scientificamerican/journal/v25/
n5s/pdf/scientificamericansciencestories1216-4.pdf.

Anonymous:2016:MMB

[3066] Anonymous. Mysteries of mind and body. Scientific American, 25(5s):34–35, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
34.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-34.pdf.

Anonymous:2016:FT

[3067] Anonymous. Future technology. Scientific American, 25(5s):70–71, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5s/full/scientificamericansciencestories1216-70.
html;   http://www.nature.com/scientificamerican/journal/v25/
n5s/pdf/scientificamericansciencestories1216-70.pdf.

Anonymous:2016:NVL

[3068] Anonymous. New views of life on Earth. Scientific American, 25(5s):86–87, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
86.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-86.pdf.

Wong:2016:MHb

[3069] Kate Wong. Mystery human. Scientific American, 25(5s):88–97, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5s/full/scientificamericansciencestories1216-88.
html;   http://www.nature.com/scientificamerican/journal/v25/
n5s/pdf/scientificamericansciencestories1216-88.pdf.

Lemonick:2016:SSPb

[3070] Michael D. Lemonick. Space: The search for Planet X. Scientific American, 25(5s):6–13, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
6.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-6.pdf.

Libeskind:2016:AOPb

[3071] Noam I. Libeskind and R. Brent Tully. Astronomy: Our place in the Cosmos. Scientific American, 25(5s):14–21, December 2016.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n5s/full/scientificamericansciencestories1216-14.html;   http:
//www.nature.com/scientificamerican/journal/v25/n5s/pdf/scientificamericansciencestories1216-
14.pdf.

Szapudi:2016:CEPb

[3072] István Szapudi. Cosmology: The emptiest place in space. Scientific American, 25(5s):22–29, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
22.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-22.pdf.

Moskowitz:2016:ARS

[3073] Clara Moskowitz. Astrophysics: Ripples in spacetime. Scientific American, 25(5s):30–33, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
30.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-30.pdf.

Moser:2016:NWWb

[3074] May-Britt Moser and Edvard I. Moser. Neuroscience: Where am I? Where am I going? Scientific American, 25(5s):36–43, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n5s/full/scientificamericansciencestories1216-36.html;   http:
//www.nature.com/scientificamerican/journal/v25/n5s/pdf/scientificamericansciencestories1216-
36.pdf.

Weintraub:2016:FMC

[3075] Karen Weintraub. The future of medicine: The cancer defense. Scientific American, 25(5s):44–51, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
44.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-44.pdf.

Carreno:2016:NAV

[3076] Beatriz M. Carreno and Elaine R. Mardis. Novel approaches: A vaccine for cancer? Scientific American, 25(5s):50, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n5s/full/scientificamericansciencestories1216-50.html;   http:
//www.nature.com/scientificamerican/journal/v25/n5s/pdf/scientificamericansciencestories1216-
50.pdf.

Maron:2016:PHH

[3077] Dina Fine Maron. Public health: How Zika spiraled out of control. Scientific American, 25(5s):52–57, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n5s/full/scientificamericansciencestories1216-52.html;   http:
//www.nature.com/scientificamerican/journal/v25/n5s/pdf/scientificamericansciencestories1216-
52.pdf.

Gifford:2016:LL

[3078] Bill Gifford. Longevity: Living to 120. Scientific American, 25(5s):58–63, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5s/full/scientificamericansciencestories1216-58.
html;   http://www.nature.com/scientificamerican/journal/v25/
n5s/pdf/scientificamericansciencestories1216-58.pdf.

Hall:2016:BRLb

[3079] Stephen S. Hall. Biotechnology: The red line. Scientific American, 25(5s):64–69, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
64.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-64.pdf.

Lu:2016:BMLb

[3080] Timothy K. Lu and Oliver Purcell. Bioengineering: Machine life. Scientific American, 25(5s):72–77, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n5s/full/scientificamericansciencestories1216-72.html;   http:
//www.nature.com/scientificamerican/journal/v25/n5s/pdf/scientificamericansciencestories1216-
72.pdf.

Shladover:2016:TTA

[3081] Steven E. Shladover. Transportation: The truth about “self-driving” cars. Scientific American, 25(5s):78–83, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n5s/full/scientificamericansciencestories1216-78.html;   http:
//www.nature.com/scientificamerican/journal/v25/n5s/pdf/scientificamericansciencestories1216-
78.pdf.

Moskowitz:2016:PWL

[3082] Clara Moskowitz. Physics: Are we living in a computer simulation? Scientific American, 25(5s):84–85, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n5s/full/scientificamericansciencestories1216-84.html;   http:
//www.nature.com/scientificamerican/journal/v25/n5s/pdf/scientificamericansciencestories1216-
84.pdf.

Wendle:2016:SSCb

[3083] John Wendle. Sustainability: Syria’s climate refugees. Scientific American, 25(5s):98–103, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
98.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-98.pdf.

Zalasiewicz:2016:ESH

[3084] Jan Zalasiewicz. Earth science: A history in layers. Scientific American, 25(5s):104–111, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
104.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-104.pdf.

Fischetti:2016:ENC

[3085] Mark Fischetti. End note: The circle of life. Scientific American, 25(5s):112, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5s/full/scientificamericansciencestories1216-
112.html;   http://www.nature.com/scientificamerican/journal/
v25/n5s/pdf/scientificamericansciencestories1216-112.pdf.

DiChristina:2016:EIC

[3086] Mariette DiChristina. From the Editor: Ideas to change the world. Scientific American, 315(6):5, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-

5.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-5.pdf.

Anonymous:2016:Ll

[3087] Anonymous. Letters. Scientific American, 315(6):8–9, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-8.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
8.pdf.

Schlenoff:2016:YAIl

[3088] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 315(6):86, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-86.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
86.pdf.

Anonymous:2016:SABb

[3089] Anonymous. Science agenda: Bringing up baby, helping the economy. Scientific American, 315(6):10, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-10.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
10.pdf.

Valantine:2016:FSG

[3090] Hannah A. Valantine. Forum: Science has a gender problem. Scientific American, 315(6):12, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
12.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-12.pdf.

Hall:2016:AMW

[3091] Shannon Hall. Advances: The Milky Way, transformed. Scientific American, 315(6):14–16, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-

14.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-14.pdf.

Jacobson:2016:YPD

[3092] Roni Jacobson. Your pun-divided attention. Scientific American, 315(6):17, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
17.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-17.pdf.

Brouillette:2016:OO

[3093] Monique Brouillette. Organelle overhaul. Scientific American, 315(6):18, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-18a.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
18a.pdf.

Goldman:2016:BBC

[3094] Jason G. Goldman. The brainy big cats. Scientific American, 315(6):18, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-18b.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
18b.pdf.

Mandelbaum:2016:OCL

[3095] Ryan F. Mandelbaum. Obama’s conservation legacy. Scientific American, 315(6):19, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
19.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-19.pdf.

Billings:2016:HSE

[3096] Lee Billings. Hope springs eternal for easy access to water on Europa. Scientific American, 315(6):20, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-20.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
20.pdf.

Smith:2016:SP

[3097] Karl J. P. Smith. Surge protector. Scientific American, 315(6):21, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-21.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
21.pdf.

Anonymous:2016:QHl

[3098] Anonymous. Quick hits. Scientific American, 315(6):22, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-22.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
22.pdf.

Caruso:2016:TTD

[3099] Catherine Caruso. Teen takes on drought. Scientific American, 315(6):23, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-23.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
23.pdf.

Goldman:2016:OTN

[3100] Jason G. Goldman. One’s true nature. Scientific American, 315(6):24, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-24.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
24.pdf.

Anderson:2016:PUC

[3101] Mark Anderson. P.U. to Cu. Scientific American, 315(6):27, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-27.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
27.pdf.

Schmidt:2016:SHW

[3102] Charles Schmidt. The science of health: When medical tests mislead. Scientific American, 315(6):28–29, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-28.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
28.pdf.

Pogue:2016:TRF

[3103] David Pogue. TechnoFiles: Resistance is futile. Scientific American, 315(6):30, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
30.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-30.pdf.

Anonymous:2016:RRRg

[3104] Anonymous. Reviews and recommendations: Recommended. Scientific American, 315(6):82, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
82.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-82.pdf.

Shermer:2016:SBW

[3105] Michael Shermer. Skeptic: Born this way. Scientific American, 315(6):84, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-84.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
84.pdf.

Mirsky:2016:AGCc

[3106] Steve Mirsky. Anti gravity: Citrus be the place. Scientific American, 315(6):85, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
85.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-85.pdf.

Peek:2016:GSR

[3107] Katie Peek. Graphic science: Is the rise in twin births cresting? Scientific American, 315(6):88, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-88.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
88.pdf.

Anonymous:2016:IWC

[3108] Anonymous. Innovation: 2016 world changing ideas. Scientific American, 315(6):32–33, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
32.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-32.pdf.

Sneed:2016:CBB

[3109] Annie Sneed. 1. Carbon-breathing batteries. Scientific American, 315(6):34, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
34.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-34.pdf.

Sneed:2016:NAS

[3110] Annie Sneed. 2. New antibiotics from scratch. Scientific American, 315(6):35, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
35a.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-35a.pdf.

Pavlus:2016:QS

[3111] John Pavlus. 3. Quantum satellites. Scientific American, 315(6):35, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-35b.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
35b.pdf.

Pavlus:2016:IMR

[3112] John Pavlus. 4. Ingestible micro robots. Scientific American, 315(6):36–37, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-

36.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-36.pdf.

Sneed:2016:PPS

[3113] Annie Sneed. 5. Poverty-predicting software. Scientific American, 315(6):37, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
37.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-37.pdf.

Sneed:2016:CC

[3114] Annie Sneed. 6. Cool clothes. Scientific American, 315(6):38, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-38.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
38.pdf.

Sneed:2016:UVF

[3115] Annie Sneed. 7. The ultimate virus fighter. Scientific American, 315(6):39, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
39a.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-39a.pdf.

Pavlus:2016:SRS

[3116] John Pavlus. 8. Sight-reading software. Scientific American, 315(6):39, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v315/n6/full/scientificamerican1216-39b.html;   http:
//www.nature.com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
39b.pdf.

Patel:2016:PDC

[3117] Prachi Patel. 9. Paper diagnostics that cost pennies. Scientific American, 315(6):40, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
40.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-40.pdf.

Pavlus:2016:SMS

[3118] John Pavlus. 10. Super-molecules from superatoms. Scientific American, 315(6):41, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
41.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-41.pdf.

Elkins-Tanton:2016:ASS

[3119] Linda T. Elkins-Tanton. Astronomy: Solar system smashup. Scientific American, 315(6):42–49, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v315/n6/full/scientificamerican1216-
42.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-42.pdf.

Sanders:2016:VHA

[3120] Rogier W. Sanders, Ian A. Wilson, and John P. Moore. Vaccines: HIV’s Achilles’ heel. Scientific American, 315(6):50–55, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-50.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
50.pdf.

Schuur:2016:CCP

[3121] Ted Schuur. Climate change: The permafrost prediction. Scientific American, 315(6):56–61, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v315/n6/full/scientificamerican1216-
56.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-56.pdf.

dHuy:2016:AEM

[3122] Julien d’Huy. Anthropology: The evolution of myths. Scientific American, 315(6):62–69, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v315/n6/full/scientificamerican1216-
62.html;   http://www.nature.com/scientificamerican/journal/
v315/n6/pdf/scientificamerican1216-62.pdf.

Moyer:2016:ALT

[3123] Melinda Wenner Moyer. Agriculture: The looming threat of factory superbugs. Scientific American, 315(6):70–79, December 2016. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-70.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
70.pdf.

Fischman:2016:OPC

[3124] Josh Fischman. Oceanography: The plight of the celebrity scientist. Scientific American, 315(6):80–81, December 2016. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v315/
n6/full/scientificamerican1216-80.html;   http://www.nature.
com/scientificamerican/journal/v315/n6/pdf/scientificamerican1216-
80.pdf.

DiChristina:2017:EWN

[3125] Mariette DiChristina. From the Editor: What’s next for science? Scientific American, 316(1):3, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
3.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-3.pdf.

Anonymous:2017:La

[3126] Anonymous. Letters. Scientific American, 316(1):4–5, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n1/full/scientificamerican0117-4.html;   http://www.nature.
com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
4.pdf.

Schlenoff:2017:YAIa

[3127] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 316(1):71, January 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n1/full/scientificamerican0117-71.html;   http://www.nature.
com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
71.pdf.

Anonymous:2017:SALa

[3128] Anonymous. Science agenda: Let NASA take flight. Scientific American, 316(1):7, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
7.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-7.pdf.

Morris:2017:FKH

[3129] Nathaniel P. Morris. Forum: Keep hospitals weapons-free. Scientific American, 316(1):8, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
8.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-8.pdf.

Wong:2017:AWT

[3130] Kate Wong. Advances: Whose tools are these? Scientific American, 316(1):10–12, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
10.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-10.pdf.

Mandelbaum:2017:S

[3131] Ryan F. Mandelbaum. STDs on the up. Scientific American, 316(1):11, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n1/full/scientificamerican0117-11.html;   http:
//www.nature.com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
11.pdf.

Weintraub:2017:CSD

[3132] Karen Weintraub. Could a special diet replace chemotherapy? Scientific American, 316(1):14, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
14.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-14.pdf.

Pierre-Louis:2017:BB

[3133] Kendra Pierre-Louis. Bats in the Bronx. Scientific American, 316(1):15, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n1/full/scientificamerican0117-15.html;   http:
//www.nature.com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
15.pdf.

Hsu:2017:TTT

[3134] Jeremy Hsu. Telescopic tag team. Scientific American, 316(1):16, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n1/full/scientificamerican0117-16.html;   http:
//www.nature.com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
16.pdf.

Goldman:2017:RTS

[3135] Jason G. Goldman. The remarkable timing of seals. Scientific American, 316(1):17, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
17.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-17.pdf.

Zaraska:2017:VV

[3136] Marta Zaraska. Veggies with vision. Scientific American, 316(1):18, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n1/full/scientificamerican0117-18a.html;   http:
//www.nature.com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
18a.pdf.

Mandelbaum:2017:HSE

[3137] Ryan F. Mandelbaum. Hidden side effects. Scientific American, 316(1): 18, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n1/full/scientificamerican0117-18b.html;   http:
//www.nature.com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
18b.pdf.

Anonymous:2017:QHa

[3138] Anonymous. Quick hits. Scientific American, 316(1):20, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n1/full/scientificamerican0117-20.html;   http://www.nature.

com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
20.pdf.

Maron:2017:SHCa

[3139] Dina Fine Maron. The science of health: Cold comfort. Scientific American, 316(1):22–23, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
22.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-22.pdf.

Pogue:2017:TYM

[3140] David Pogue. TechnoFiles: Your e-mail password will never be safe. Scientific American, 316(1):24, January 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n1/full/scientificamerican0117-24.html;   http://www.nature.
com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
24.pdf.

Anonymous:2017:RRRa

[3141] Anonymous. Reviews and recommendations: Recommended. Scientific American, 316(1):68, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
68.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-68.pdf.

Shermer:2017:SWF

[3142] Michael Shermer. Skeptic: When facts backfire. Scientific American, 316(1):69, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
69.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-69.pdf.

Mirsky:2017:AGDa

[3143] Steve Mirsky. Anti gravity: Data deliver in the clutch. Scientific American, 316(1):70, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
70.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-70.pdf.

Fischetti:2017:GSTa

[3144] Mark Fischetti. Graphic science: Top air polluters. Scientific American, 316(1):72, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
72.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-72.pdf.

Knoblich:2017:NLB

[3145] Juergen A. Knoblich. Neuroscience: Lab-built brains. Scientific American, 316(1):26–31, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
26.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-26.pdf.

Moskowitz:2017:STS

[3146] Clara Moskowitz. Space: Tangled up in spacetime. Scientific American, 316(1):32–37, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
32.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-32.pdf.

Rubanyi:2017:CHT

[3147] Gabor Rubanyi. Cardiology: Heart therapy. Scientific American, 316(1):38–43, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
38.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-38.pdf.

Briggs:2017:CCR

[3148] Gordon Briggs and Matthias Scheutz. Computing: The case for robot disobedience. Scientific American, 316(1):44–47, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n1/full/scientificamerican0117-44.html;   http://www.nature.
com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
44.pdf.

Brusatte:2017:ETW

[3149] Stephen Brusatte. Evolution: Taking wing. Scientific American, 316(1):48–55, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
48.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-48.pdf.

Byrd:2017:CJA

[3150] Jason Byrd and Natasha Whitling. Criminal justice: Animal CSI. Scientific American, 316(1):56–63, January 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n1/full/scientificamerican0117-56.html;   http://www.nature.
com/scientificamerican/journal/v316/n1/pdf/scientificamerican0117-
56.pdf.

Schiffman:2017:QWW

[3151] Richard Schiffman. Q & A: Wildlife warrior. Scientific American, 316(1):64–67, January 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n1/full/scientificamerican0117-
64.html;   http://www.nature.com/scientificamerican/journal/
v316/n1/pdf/scientificamerican0117-64.pdf.

DiChristina:2017:EWF

[3152] Mariette DiChristina. From the Editor: From workouts to far out. Scientific American, 316(2):4, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
4.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-4.pdf.

Anonymous:2017:Lb

[3153] Anonymous. Letters. Scientific American, 316(2):5–6, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n2/full/scientificamerican0217-5.html;   http://www.nature.
com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
5.pdf.

Schlenoff:2017:YAIb

[3154] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 316(2):75, February 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n2/full/scientificamerican0217-75.html;   http://www.nature.
com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
75.pdf.

Anonymous:2017:SALb

[3155] Anonymous. Science agenda: A letter to Washington. Scientific American, 316(2):7, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
7.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-7.pdf.

Mann:2017:FCT

[3156] Michael E. Mann and Susan Joy Hassol. Forum: Climate trumps everything. Scientific American, 316(2):8, February 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n2/full/scientificamerican0217-8.html;   http://www.nature.
com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
8.pdf.

Kuchment:2017:AMM

[3157] Anna Kuchment. Advances: Man-made solutions for man-made quakes. Scientific American, 316(2):10–12, February 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n2/full/scientificamerican0217-10.html;   http://www.nature.
com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
10.pdf.

Chen:2017:LMC

[3158] Ingfei Chen. Learn Morse code, semiconsciously. Scientific American, 316(2):12–13, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
12.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-12.pdf.

Brouillette:2017:TPM

[3159] Monique Brouillette. To treat primates more humanely: Transparency. Scientific American, 316(2):14, February 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n2/full/scientificamerican0217-14.html;   http://www.nature.
com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
14.pdf.

Patel:2017:BYB

[3160] Prachi Patel. Bots in your bloodstream. Scientific American, 316(2):15, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n2/full/scientificamerican0217-15.html;   http:
//www.nature.com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
15.pdf.

Anonymous:2017:QHb

[3161] Anonymous. Quick hits. Scientific American, 316(2):16, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n2/full/scientificamerican0217-16.html;   http://www.nature.
com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
16.pdf.

Mandelbaum:2017:IR

[3162] Ryan F. Mandelbaum. Icy retreat. Scientific American, 316(2):17, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n2/full/scientificamerican0217-17.html;   http:
//www.nature.com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
17.pdf.

Gannon:2017:FF

[3163] Megan Gannon. Fruits for the frozen. Scientific American, 316(2):18, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n2/full/scientificamerican0217-18a.html;   http:
//www.nature.com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
18a.pdf.

Anonymous:2017:TMW

[3164] Anonymous. “[This] tells me we don’t just have to wring our hands about the high level of mercury in these fish. There is something we can do about it and get pretty quick results.”. Scientific American, 316(2):18, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n2/full/scientificamerican0217-18b.html;   http:
//www.nature.com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
18b.pdf.

Frazer:2017:GWA

[3165] Jennifer Frazer. Go West, allergy sufferers. Scientific American, 316(2):19, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
19.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-19.pdf.

Nuwer:2017:TSA

[3166] Rachel Nuwer. The trauma of saving animals. Scientific American, 316(2):20, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
20.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-20.pdf.

Weintraub:2017:SHS

[3167] Karen Weintraub. The science of health: Stomach upset. Scientific American, 316(2):22–23, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
22.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-22.pdf.

Pogue:2017:THS

[3168] David Pogue. TechnoFiles: How to stamp out fake news. Scientific American, 316(2):24, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
24.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-24.pdf.

Anonymous:2017:RRRb

[3169] Anonymous. Reviews and recommendations: Recommended. Scientific American, 316(2):72, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
72.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-72.pdf.

Shermer:2017:SIN

[3170] Michael Shermer. Skeptic: Imagine no universe. Scientific American, 316(2):73, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
73.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-73.pdf.

Mirsky:2017:AGN

[3171] Steve Mirsky. Anti gravity: None so blind. Scientific American, 316(2):74, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
74.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-74.pdf.

Fischetti:2017:GSN

[3172] Mark Fischetti. Graphic science: Novel math. Scientific American, 316(2):76, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
76.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-76.pdf.

Pontzer:2017:EEP

[3173] Herman Pontzer. Evolution: The exercise paradox. Scientific American, 316(2):26–31, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
26.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-26.pdf.

Ijjas:2017:CPG

[3174] Anna Ijjas, Paul J. Steinhardt, and Abraham Loeb. Cosmology: Pop goes the universe. Scientific American, 316(2):32–39, February 2017.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n2/full/scientificamerican0217-32.html;   http://www.nature.
com/scientificamerican/journal/v316/n2/pdf/scientificamerican0217-
32.pdf.

Schmale:2017:BHF

[3175] David Schmale and Shane Ross. Biology: High-flying microbes. Scientific American, 316(2):40–45, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
40.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-40.pdf.

Folger:2017:MMH

[3176] Tim Folger. Metrology: Mass hysteria. Scientific American, 316(2):46–53, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
46.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-46.pdf.

Limoli:2017:NDS

[3177] Charles L. Limoli. Neuroscience: Deep-space deal breaker. Scientific American, 316(2):54–59, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
54.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-54.pdf.

Meyer:2017:LWW

[3178] Julien Meyer. Linguistics: The whistled word. Scientific American, 316(2):60–67, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-
60.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-60.pdf.

Peplow:2017:HBM

[3179] Mark Peplow. Health: Blind medicine. Scientific American, 316(2):68–71, February 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n2/full/scientificamerican0217-

68.html;   http://www.nature.com/scientificamerican/journal/
v316/n2/pdf/scientificamerican0217-68.pdf.

DiChristina:2017:EBG

[3180] Mariette DiChristina. From the Editor: To boldly go. Scientific American, 316(3):4, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
4.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-4.pdf.

Anonymous:2017:Lc

[3181] Anonymous. Letters. Scientific American, 316(3):6–7, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n3/full/scientificamerican0317-6.html;   http://www.nature.
com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
6.pdf.

Schlenoff:2017:YAIc

[3182] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 316(3):79, March 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n3/full/scientificamerican0317-79.html;   http://www.nature.
com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
79.pdf.

Belfiore:2017:AMD

[3183] Michael Belfiore. Advances: Metal devices, in miniature. Scientific American, 316(3):16, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
16.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-16.pdf.

Biba:2017:SRM

[3184] Erin Biba. STD results in minutes. Scientific American, 316(3):18, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n3/full/scientificamerican0317-18.html;   http:

//www.nature.com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
18.pdf.

Caruso:2017:MEG

[3185] Catherine Caruso. Make Earth great again. Scientific American, 316(3): 20, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n3/full/scientificamerican0317-20.html;   http:
//www.nature.com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
20.pdf.

Dill:2017:FOE

[3186] David Dill. Forum: Our elections are not secure. Scientific American, 316(3):12, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
12.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-12.pdf.

Guterl:2017:AFO

[3187] Fred Guterl. Advances: Is fusion in our future? Scientific American, 316(3):17, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
17.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-17.pdf.

Lloyd:2017:SHG

[3188] Robin Lloyd. The science of health: Gasping for air. Scientific American, 316(3):26–27, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
26.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-26.pdf.

Mandelbaum:2017:AIT

[3189] Ryan F. Mandelbaum. Advances: It takes guts. Scientific American, 316(3):19, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
19.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-19.pdf.

Mirsky:2017:AGPa

[3190] Steve Mirsky. Anti gravity: 23 and pee. Scientific American, 316(3):78, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n3/full/scientificamerican0317-78.html;   http:
//www.nature.com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
78.pdf.

Anonymous:2017:RRRc

[3191] Anonymous. Reviews and recommendations: Recommended. Scientific American, 316(3):76, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
76.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-76.pdf.

Peek:2017:GSL

[3192] Katie Peek. Graphic science: Long live Hubble. Scientific American, 316(3):80, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
80.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-80.pdf.

Platt:2017:AHG

[3193] John R. Platt. Advances: How to get elephants to buzz off. Scientific American, 316(3):23, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
23.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-23.pdf.

Pogue:2017:TYE

[3194] David Pogue. TechnoFiles: Your echo is listening. Scientific American, 316(3):28, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
28.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-28.pdf.

Qiu:2017:AIA

[3195] Jane Qiu. Advances: Ice age tibetans. Scientific American, 316(3):14– 16, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n3/full/scientificamerican0317-14.html;   http:
//www.nature.com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
14.pdf.

Shermer:2017:SAA

[3196] Michael Shermer. Skeptic: Apocalypse AI. Scientific American, 316(3): 77, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n3/full/scientificamerican0317-77.html;   http:
//www.nature.com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
77.pdf.

Sneed:2017:AER

[3197] Annie Sneed. Advances: It’s electric — with the right mix. Scientific American, 316(3):24, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
24.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-24.pdf.

Switek:2017:FOJ

[3198] Brian Switek. Fossil octopus is a Jurassic jewel. Scientific American, 316(3):21, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
21.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-21.pdf.

Anonymous:2017:SAT

[3199] Anonymous. Science agenda: Take nukes off a short fuse. Scientific American, 316(3):10, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
10.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-10.pdf.

Anonymous:2017:AQH

[3200] Anonymous. Advances: Quick hits. Scientific American, 316(3):22, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n3/full/scientificamerican0317-22.html;   http:

//www.nature.com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
22.pdf.

Finkbeiner:2017:SNL

[3201] Ann Finkbeiner. Spaceflight: Near-light-speed mission to Alpha Centauri. Scientific American, 316(3):30–37, March 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n3/full/scientificamerican0317-30.html;   http://www.nature.
com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
30.pdf.

Marcus:2017:RH

[3202] Gary Marcus. Robotics: Am i human? Scientific American, 316(3):58– 63, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n3/full/scientificamerican0317-58.html;   http:
//www.nature.com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
58.pdf.

Noble:2017:BNB

[3203] Kimberly G. Noble. Behavioral neuroscience: Brain trust. Scientific American, 316(3):44–49, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
44.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-44.pdf.

Nordhaus:2017:ACV

[3204] Hannah Nordhaus. Agriculture: Cornboy vs. the billion-dollar bug. Scientific American, 316(3):64–71, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v316/n3/full/scientificamerican0317-
64.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-64.pdf.

Pavlus:2017:NTT

[3205] John Pavlus. The new Turing tests. Scientific American, 316(3):61–62, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n3/full/scientificamerican0317-61.html;   http:
//www.nature.com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
61.pdf.

Posey:2017:ICK

[3206] Avery D. Posey, Jr., Carl H. June, and Bruce L. Levine. Immunology: Cancer killers. Scientific American, 316(3):38–43, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n3/full/scientificamerican0317-38.html;   http://www.nature.
com/scientificamerican/journal/v316/n3/pdf/scientificamerican0317-
38.pdf.

Rochman:2017:GSB

[3207] Bonnie Rochman. Genomics: Should babies be sequenced? Scientific American, 316(3):72–75, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
72.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-72.pdf.

Vinther:2017:PTC

[3208] Jakob Vinther. Paleontology: The true colors of dinosaurs. Scientific American, 316(3):50–57, March 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n3/full/scientificamerican0317-
50.html;   http://www.nature.com/scientificamerican/journal/
v316/n3/pdf/scientificamerican0317-50.pdf.

DiChristina:2017:EKF

[3209] Mariette DiChristina. From the Editor: Keeping the future in mind. Scientific American, 316(4):4, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
4.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-4.pdf.

Anonymous:2017:Ld

[3210] Anonymous. Letters. Scientific American, 316(4):6–8, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n4/full/scientificamerican0417-6.html;   http://www.nature.
com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
6.pdf.

Schlenoff:2017:YAId

[3211] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in scientific American: 50, 100 & 150 years ago. Scientific American, 316(4):82, April 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n4/full/scientificamerican0417-82.html;   http://www.nature.
com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
82.pdf.

Anonymous:2017:SAB

[3212] Anonymous. Science agenda: A better reckoning. Scientific American, 316(4):10, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
10.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-10.pdf.

Cowtan:2017:FSS

[3213] Kevin Cowtan and Zeke Hausfather. Forum: Stay out of scientists’ e-mails. Scientific American, 316(4):12, April 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n4/full/scientificamerican0417-12.html;   http://www.nature.
com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
12.pdf.

Wright:2017:ASL

[3214] Katherine Wright. Advances: Sky is the limit. Scientific American, 316(4):14–17, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
14.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-14.pdf.

Mandelbaum:2017:TST

[3215] Ryan F. Mandelbaum. Telltale sounds of tsunamis. Scientific American, 316(4):18–19, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
18a.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-18a.pdf.

Goldman:2017:WOM

[3216] Jason G. Goldman. Which one is mom again? Scientific American, 316(4):18–19, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
18b.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-18b.pdf.

Marks:2017:QHa

[3217] Andrea Marks. Quick hits. Scientific American, 316(4):20, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n4/full/scientificamerican0417-20.html;   http://www.nature.
com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
20.pdf.

Braun:2017:CC

[3218] Ashley Braun. A crimp on crops. Scientific American, 316(4):21, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n4/full/scientificamerican0417-21.html;   http:
//www.nature.com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
21.pdf.

Drouin:2017:LL

[3219] Roger Drouin. A lift for lichens. Scientific American, 316(4):22, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n4/full/scientificamerican0417-22.html;   http:
//www.nature.com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
22.pdf.

Nguyen:2017:SG

[3220] Tien Nguyen. Screen grab. Scientific American, 316(4):23, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n4/full/scientificamerican0417-23.html;   http://www.nature.
com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
23.pdf.

Maron:2017:URF

[3221] Dina Fine Maron. Is the U.S. ready for future disease threats? Scientific American, 316(4):24, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
24.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-24.pdf.

Hsu:2017:NFF

[3222] Jeremy Hsu. NASA fights flight delays. Scientific American, 316(4):26, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n4/full/scientificamerican0417-26.html;   http:
//www.nature.com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
26.pdf.

Noonan:2017:SHE

[3223] David Noonan. The science of health: The epilepsy dilemma. Scientific American, 316(4):28–29, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
28.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-28.pdf.

Pogue:2017:TWH

[3224] David Pogue. TechnoFiles: What happened to user manuals? Scientific American, 316(4):30, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
30.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-30.pdf.

Anonymous:2017:RRRd

[3225] Anonymous. Reviews and recommendations: Recommended. Scientific American, 316(4):76, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
76.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-76.pdf.

Shermer:2017:SWTa

[3226] Michael Shermer. Skeptic: What is truth, anyway? Scientific American, 316(4):78, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-

78.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-78.pdf.

Mirsky:2017:AGFa

[3227] Steve Mirsky. Anti gravity: Floor plan. Scientific American, 316(4):80, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n4/full/scientificamerican0417-80.html;   http:
//www.nature.com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
80.pdf.

Fischetti:2017:GSTb

[3228] Mark Fischetti. Graphic science: Trillions of insects migrate. Scientific American, 316(4):84, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
84.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-84.pdf.

Kivipelto:2017:NRS

[3229] Miia Kivipelto and Krister Håkansson. Neurology: A rare success against Alzheimer’s. Scientific American, 316(4):32–37, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n4/full/scientificamerican0417-32.html;   http://www.nature.
com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
32.pdf.

Cenko:2017:AHS

[3230] S. Bradley Cenko and Neil Gehrels. Astrophysics: How to swallow a Sun. Scientific American, 316(4):38–45, April 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n4/full/scientificamerican0417-38.html;   http://www.nature.
com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
38.pdf.

Waldholz:2017:SRF

[3231] Michael Waldholz. Special report: 2017 future of medicine: Transformers. Scientific American, 316(4):46–53, April 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n4/full/scientificamerican0417-46.html;   http://www.nature.

com/scientificamerican/journal/v316/n4/pdf/scientificamerican0417-
46.pdf.

Esvelt:2017:UR

[3232] Kevin M. Esvelt. Unnatural responsibilities. Scientific American, 316(4):50–51, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
50.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-50.pdf.

Riesch:2017:BEL

[3233] Rüdiger Riesch and Martin Plath. Biology: Evolution at the limits. Scientific American, 316(4):54–59, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v316/n4/full/scientificamerican0417-
54.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-54.pdf.

Quattrociocchi:2017:SSI

[3234] Walter Quattrociocchi. Social science: Inside the echo chamber. Scientific American, 316(4):60–63, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
60.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-60.pdf.

Schmidt:2017:HDD

[3235] Charles Schmidt. Health: Don’t drink the water. Scientific American, 316(4):64–69, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
64.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-64.pdf.

Keim:2017:BWT

[3236] Paul S. Keim, David H. Walker, and Raymond A. Zilinskas. Biological weapons: Time to worry about anthrax again. Scientific American, 316(4):70–75, April 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n4/full/scientificamerican0417-
70.html;   http://www.nature.com/scientificamerican/journal/
v316/n4/pdf/scientificamerican0417-70.pdf.

DiChristina:2017:EDI

[3237] Mariette DiChristina. From the Editor: Dawn of innovation. Scientific American, 316(5):4, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
4.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-4.pdf.

Anonymous:2017:Le

[3238] Anonymous. Letters. Scientific American, 316(5):6–8, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-6.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
6.pdf.

Schlenoff:2017:YAIe

[3239] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in scientific American: 50, 100 & 150 years ago. Scientific American, 316(5):79, May 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-79.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
79.pdf.

Anonymous:2017:SAM

[3240] Anonymous. Science agenda: A March for science is not enough. Scientific American, 316(5):9, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
9.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-9.pdf.

Foley:2017:FWF

[3241] Jonathan Foley. Forum: The war on facts undermines democracy. Scientific American, 316(5):10, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
10.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-10.pdf.

Emspak:2017:ANG

[3242] Jesse Emspak. Advances: Nuclear ghosts. Scientific American, 316(5): 12–14, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n5/full/scientificamerican0517-12.html;   http:
//www.nature.com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
12.pdf.

Lloyd:2017:DM

[3243] Robin Lloyd. Dissent with modification. Scientific American, 316(5): 14, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n5/full/scientificamerican0517-14.html;   http:
//www.nature.com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
14.pdf.

Belfiore:2017:MM

[3244] Michael Belfiore. Maps on the move. Scientific American, 316(5):15, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n5/full/scientificamerican0517-15.html;   http:
//www.nature.com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
15.pdf.

Long:2017:HSD

[3245] Kat Long. Home sweet dome. Scientific American, 316(5):16, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-16.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
16.pdf.

Robitzski:2017:SH

[3246] Dan Robitzski. A spare hand. Scientific American, 316(5):17, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-17.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
17.pdf.

Laursen:2017:SF

[3247] Lucas Laursen. Soil in the forecast. Scientific American, 316(5):18, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n5/full/scientificamerican0517-18.html;   http:
//www.nature.com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
18.pdf.

Hu:2017:BBR

[3248] Jane C. Hu. Bilingual brains remember. Scientific American, 316(5): 19, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n5/full/scientificamerican0517-19.html;   http:
//www.nature.com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
19.pdf.

Marks:2017:QHb

[3249] Andrea Marks. Quick hits. Scientific American, 316(5):20, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-20.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
20.pdf.

Biba:2017:HR

[3250] Erin Biba. High robot. Scientific American, 316(5):21, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-21a.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
21a.pdf.

Lott:2017:GH

[3251] Melissa C. Lott. Green hydrogen. Scientific American, 316(5):21, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n5/full/scientificamerican0517-21b.html;   http:
//www.nature.com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
21b.pdf.

Chen:2017:MM

[3252] Angus Chen. Of meat and men. Scientific American, 316(5):22, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n5/full/scientificamerican0517-22.html;   http:

//www.nature.com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
22.pdf.

Callier:2017:SHA

[3253] Viviane Callier. The science of health: Aspirin vs. cancer. Scientific American, 316(5):24–25, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
24.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-24.pdf.

Pogue:2017:TBB

[3254] David Pogue. TechnoFiles: Building a better battery. Scientific American, 316(5):26, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
26.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-26.pdf.

Anonymous:2017:RRRe

[3255] Anonymous. Reviews and recommendations: Recommended. Scientific American, 316(5):74, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
74.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-74.pdf.

Shermer:2017:SWTb

[3256] Michael Shermer. Skeptic: On witches and terrorists. Scientific American, 316(5):77, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
77.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-77.pdf.

Mirsky:2017:AGDb

[3257] Steve Mirsky. Anti gravity: Don’t pass the weed or say “guns”. Scientific American, 316(5):78, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
78.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-78.pdf.

Peek:2017:GSE

[3258] Katie Peek. Graphic science: The elusive northwest passage. Scientific American, 316(5):80, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
80.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-80.pdf.

Wong:2017:ANO

[3259] Kate Wong. Archaeology: The new origins of technology. Scientific American, 316(5):28–35, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
28.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-28.pdf.

Cartier:2017:ASN

[3260] Kimberly Cartier and Jason T. Wright. Astrophysics: Strange news from another star. Scientific American, 316(5):36–41, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-36.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
36.pdf.

Reno:2017:EML

[3261] Philip L. Reno. Evolution: Missing links. Scientific American, 316(5): 42–47, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n5/full/scientificamerican0517-42.html;   http:
//www.nature.com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
42.pdf.

Sivaram:2017:CGW

[3262] Varun Sivaram. Climate: The global warming wild card. Scientific American, 316(5):48–53, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n5/full/scientificamerican0517-
48.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-48.pdf.

Balter:2017:NSU

[3263] Michael Balter. Neuroscience: Schizophrenia’s unyielding mysteries. Scientific American, 316(5):54–61, May 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v316/n5/full/scientificamerican0517-
54.html;   http://www.nature.com/scientificamerican/journal/
v316/n5/pdf/scientificamerican0517-54.pdf.

Fromme:2017:CSS

[3264] Petra Fromme and John C. H. Spence. Chemistry: Split-second reactions. Scientific American, 316(5):62–67, May 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-62.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
62.pdf.

Trut:2017:GHB

[3265] Lyudmila Trut and Lee Alan Dugatkin. Genetics: How to build a Dog. Scientific American, 316(5):68–73, May 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n5/full/scientificamerican0517-68.html;   http://www.nature.
com/scientificamerican/journal/v316/n5/pdf/scientificamerican0517-
68.pdf.

DiChristina:2017:ESK

[3266] Mariette DiChristina. From the Editor: From science to knowledge to hope. Scientific American, 316(6):4, June 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n6/full/scientificamerican0617-4.html;   http://www.nature.
com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
4.pdf.

Anonymous:2017:Lf

[3267] Anonymous. Letters. Scientific American, 316(6):5–6, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n6/full/scientificamerican0617-5.html;   http://www.nature.
com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
5.pdf.

Schlenoff:2017:YAIf

[3268] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in scientific American: 50, 100 & 150 years ago. Scientific American, 316(6):77, June 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n6/full/scientificamerican0617-77.html;   http://www.nature.
com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
77.pdf.

Anonymous:2017:SAS

[3269] Anonymous. Science agenda: Science without walls. Scientific American, 316(6):7, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
7.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-7.pdf.

Dijkgraaf:2017:FKI

[3270] Robbert Dijkgraaf. Forum: Knowledge is infrastructure. Scientific American, 316(6):8, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
8.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-8.pdf.

Hsu:2017:API

[3271] Jeremy Hsu. Advances: Polar ice squad. Scientific American, 316(6):10– 12, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-10.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
10.pdf.

Biba:2017:ML

[3272] Erin Biba. Missing link. Scientific American, 316(6):12, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n6/full/scientificamerican0617-12.html;   http://www.nature.
com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
12.pdf.

Kaufman:2017:SAF

[3273] Mark Kaufman. Space age firefighters. Scientific American, 316(6):14– 15, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-14a.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
14a.pdf.

Pycha:2017:PP

[3274] Anne Pycha. Probing for Parkinson’s. Scientific American, 316(6):14– 15, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-14b.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
14b.pdf.

Peplow:2017:PWR

[3275] Mark Peplow. Print, wipe, rewrite. Scientific American, 316(6):16, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-16a.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
16a.pdf.

Maron:2017:MCC

[3276] Dina Fine Maron. Menstrual cycle “on a chip”. Scientific American, 316(6):16, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
16b.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-16b.pdf.

Marks:2017:QHc

[3277] Andrea Marks. Quick hits. Scientific American, 316(6):18, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n6/full/scientificamerican0617-18.html;   http://www.nature.
com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
18.pdf.

Griswold:2017:RT

[3278] Ann Griswold. Radiation triage. Scientific American, 316(6):19, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-19.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
19.pdf.

Sneed:2017:ER

[3279] Annie Sneed. Electric Renaissance. Scientific American, 316(6):20, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-20.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
20.pdf.

Reese:2017:SHW

[3280] April Reese. Swell or high water. Scientific American, 316(6):21, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-21.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
21.pdf.

Beebe:2017:RP

[3281] Jeanette Beebe. Robo pizzaiolo. Scientific American, 316(6):22, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-22.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
22.pdf.

Weintraub:2017:SHR

[3282] Karen Weintraub. The science of health: Revenge of the super lice. Scientific American, 316(6):24–25, June 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n6/full/scientificamerican0617-24.html;   http://www.nature.
com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
24.pdf.

Pogue:2017:TID

[3283] David Pogue. TechnoFiles: The iOS 11 to-do list. Scientific American, 316(6):26, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-

26.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-26.pdf.

Anonymous:2017:RRRf

[3284] Anonymous. Reviews and recommendations: Recommended. Scientific American, 316(6):74, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
74.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-74.pdf.

Shermer:2017:SRV

[3285] Michael Shermer. Skeptic: Romance of the vanished past. Scientific American, 316(6):75, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
75.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-75.pdf.

Mirsky:2017:AGFb

[3286] Steve Mirsky. Anti gravity: Food fright! Scientific American, 316(6):76, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v316/n6/full/scientificamerican0617-76.html;   http:
//www.nature.com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
76.pdf.

Fischetti:2017:GSSa

[3287] Mark Fischetti. Graphic science: Sounds like trouble. Scientific American, 316(6):78, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
78.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-78.pdf.

Nomura:2017:CQM

[3288] Yasunori Nomura. Cosmology: The quantum multiverse. Scientific American, 316(6):28–35, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
28.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-28.pdf.

Roberts:2017:HMT

[3289] Susan B. Roberts and Sai Krupa Das. Health: The messy truth about weight loss. Scientific American, 316(6):36–41, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n6/full/scientificamerican0617-36.html;   http://www.nature.
com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
36.pdf.

Dixson:2017:MSL

[3290] Danielle L. Dixson. Marine science: Lost at sea. Scientific American, 316(6):42–45, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
42.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-42.pdf.

Petrucelli:2017:GUM

[3291] Leonard Petrucelli and Aaron D. Gitler. Genetics: Unlocking the mystery of ALS. Scientific American, 316(6):46–51, June 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v316/
n6/full/scientificamerican0617-46.html;   http://www.nature.
com/scientificamerican/journal/v316/n6/pdf/scientificamerican0617-
46.pdf.

Gies:2017:BML

[3292] Erica Gies. Biology: The meaning of lichen. Scientific American, 316(6):52–59, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
52.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-52.pdf.

Gopnik:2017:MLM

[3293] Alison Gopnik. Machine learning: Making AI more human. Scientific American, 316(6):60–65, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
60.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-60.pdf.

Zorich:2017:AGV

[3294] Zach Zorich. Archaeology: Greenland’s vanished Vikings. Scientific American, 316(6):66–73, June 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v316/n6/full/scientificamerican0617-
66.html;   http://www.nature.com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-66.pdf.

DiChristina:2017:EDC

[3295] Mariette DiChristina. From the Editor: Designing the city of tomorrow today. Scientific American, 317(1):4, July 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n1/full/scientificamerican0717-4.html;   http://www.nature.
com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
4.pdf.

Anonymous:2017:Lg

[3296] Anonymous. Letters. Scientific American, 317(1):5–8, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n1/full/scientificamerican0717-5.html;   http://www.nature.
com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
5.pdf.

Anonymous:2017:Lh

[3297] Anonymous. Letters. Scientific American, 317(1):7–8, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n1/full/scientificamerican0717-7.html;   http://www.nature.
com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
7.pdf.

Schlenoff:2017:YAIg

[3298] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in scientific American: 50, 100 & 150 years ago. Scientific American, 317(1):75, July 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n1/full/scientificamerican0717-75.html;   http://www.nature.
com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
75.pdf.

Anonymous:2017:SAW

[3299] Anonymous. Science agenda: We can’t avoid future disease outbreaks. Scientific American, 317(1):9, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
9.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-9.pdf.

Sobowale:2017:FRA

[3300] Kunmi Sobowale. Forum: Raise alcohol taxes, reduce violence. Scientific American, 317(1):10, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
10.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-10.pdf.

Posada-Swafford:2017:AAA

[3301] Angela Posada-Swafford. Advances: Amazon Atlantis. Scientific American, 317(1):12–14, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
12.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-12.pdf.

Goldman:2017:BB

[3302] Jason G. Goldman. Boogie bears. Scientific American, 317(1):14, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-14.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
14.pdf.

Marks:2017:MPT

[3303] Andrea Marks. The mouse parent trap. Scientific American, 317(1): 16, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-16a.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
16a.pdf.

Lewis:2017:SD

[3304] Tanya Lewis. The stuff of dreams. Scientific American, 317(1):16, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-16b.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
16b.pdf.

Hutson:2017:SL

[3305] Matthew Hutson. Secrets and lies. Scientific American, 317(1):17, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-17.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
17.pdf.

Goldman:2017:WW

[3306] Jason G. Goldman. Waning Woods. Scientific American, 317(1):18, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-18.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
18.pdf.

Patel:2017:HR

[3307] Prachi Patel. Hot rockets. Scientific American, 317(1):19, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n1/full/scientificamerican0717-19.html;   http://www.nature.
com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
19.pdf.

Zaraska:2017:CPH

[3308] Marta Zaraska. Can plants hear? Scientific American, 317(1):20, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-20.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
20.pdf.

Marks:2017:QHd

[3309] Andrea Marks. Quick hits. Scientific American, 317(1):22, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n1/full/scientificamerican0717-22.html;   http://www.nature.

com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
22.pdf.

Cook:2017:UT

[3310] Terri Cook. Uncharted territory. Scientific American, 317(1):23, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-23.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
23.pdf.

Caruso:2017:GRH

[3311] Catherine Caruso. Grandma’s robot helper. Scientific American, 317(1): 24, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-24.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
24.pdf.

Jabr:2017:SHP

[3312] Ferris Jabr. The science of health: Probiotics are no panacea. Scientific American, 317(1):26–27, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
26.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-26.pdf.

Pogue:2017:TCA

[3313] David Pogue. TechnoFiles: Computer-aided creativity. Scientific American, 317(1):28, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
28.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-28.pdf.

Anonymous:2017:RRRg

[3314] Anonymous. Reviews and recommendations: Recommended. Scientific American, 317(1):72, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
72.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-72.pdf.

Shermer:2017:SWY

[3315] Michael Shermer. Skeptic: Who are you? Scientific American, 317(1): 73, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-73.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
73.pdf.

Mirsky:2017:AGDc

[3316] Steve Mirsky. Anti gravity: Drilling for fossil gold. Scientific American, 317(1):74, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
74.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-74.pdf.

Fischetti:2017:GSB

[3317] Mark Fischetti and Zan Armstrong. Graphic science: The baby spike. Scientific American, 317(1):76, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n1/full/scientificamerican0717-
76.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-76.pdf.

Silva:2017:NMI

[3318] Alcino J. Silva. Neuroscience: Memory’s intricate Web. Scientific American, 317(1):30–37, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
30.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-30.pdf.

GarcxEDa-Bellido:2017:CBH

[3319] Juan García-Bellido and Sébastien Clesse. Cosmology: Black holes from the beginning of time. Scientific American, 317(1):38–43, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n1/full/scientificamerican0717-38.html;   http://www.nature.
com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
38.pdf.

McDonough:2017:SRS

[3320] William McDonough. Special report sustainable cities: How cities could save us. Scientific American, 317(1):44–48, July 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n1/full/scientificamerican0717-44.html;   http://www.nature.
com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
44.pdf.

Webber:2017:TT

[3321] Michael E. Webber. Tapping the trash. Scientific American, 317(1):48– 53, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n1/full/scientificamerican0717-48.html;   http:
//www.nature.com/scientificamerican/journal/v317/n1/pdf/scientificamerican0717-
48.pdf.

Ratti:2017:PLP

[3322] Carlo Ratti and Assaf Biderman. From parking lot to paradise. Scientific American, 317(1):54–59, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
54.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-54.pdf.

Rubino:2017:MOD

[3323] Francesco Rubino. Medicine: Operation: Diabetes. Scientific American, 317(1):60–65, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
60.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-60.pdf.

Singer:2017:BED

[3324] Thea Singer. Biology: The evolution of dance. Scientific American, 317(1):66–71, July 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n1/full/scientificamerican0717-
66.html;   http://www.nature.com/scientificamerican/journal/
v317/n1/pdf/scientificamerican0717-66.pdf.

DiChristina:2017:EB

[3325] Mariette DiChristina. From the Editor: In the beginning. Scientific American, 317(2):4, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
4.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-4.pdf.

Anonymous:2017:Li

[3326] Anonymous. Letters. Scientific American, 317(2):5–6, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n2/full/scientificamerican0817-5.html;   http://www.nature.
com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
5.pdf.

Anonymous:2017:SAN

[3327] Anonymous. Science agenda: Nuclear war should require a second opinion. Scientific American, 317(2):8, August 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n2/full/scientificamerican0817-8.html;   http://www.nature.
com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
8.pdf.

Ienca:2017:FRC

[3328] Marcello Ienca. Forum: The right to cognitive Liberty. Scientific American, 317(2):10, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
10.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-10.pdf.

Sheikh:2017:ASF

[3329] Knvul Sheikh. Advances: Saving face. Scientific American, 317(2):12– 14, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n2/full/scientificamerican0817-12.html;   http:
//www.nature.com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
12.pdf.

Biba:2017:AI

[3330] Erin Biba. Aussie invaders. Scientific American, 317(2):15, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n2/full/scientificamerican0817-15.html;   http://www.nature.
com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
15.pdf.

Nuwer:2017:FME

[3331] Rachel Nuwer. Find my elephant. Scientific American, 317(2):16, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n2/full/scientificamerican0817-16.html;   http:
//www.nature.com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
16.pdf.

Laursen:2017:EP

[3332] Lucas Laursen. Emission permission. Scientific American, 317(2):17, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n2/full/scientificamerican0817-17.html;   http:
//www.nature.com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
17.pdf.

Nemo:2017:QHa

[3333] Leslie Nemo. Quick hits. Scientific American, 317(2):18, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n2/full/scientificamerican0817-18.html;   http://www.nature.
com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
18.pdf.

Maron:2017:MBH

[3334] Dina Fine Maron. Mending a broken heart. Scientific American, 317(2): 19, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n2/full/scientificamerican0817-19.html;   http:
//www.nature.com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
19.pdf.

Greenspan:2017:AP

[3335] Jesse Greenspan. Awash in plastic. Scientific American, 317(2):20, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n2/full/scientificamerican0817-20.html;   http:
//www.nature.com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
20.pdf.

Sedacca:2017:PEW

[3336] Matthew Sedacca. Plastic-eating worms. Scientific American, 317(2):21, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n2/full/scientificamerican0817-21.html;   http:
//www.nature.com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
21.pdf.

Hall:2017:FI

[3337] Shannon Hall. Fire and ice. Scientific American, 317(2):22, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n2/full/scientificamerican0817-22.html;   http://www.nature.
com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
22.pdf.

Maron:2017:SHCb

[3338] Dina Fine Maron. The science of health: The computer will see you now. Scientific American, 317(2):24–25, August 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n2/full/scientificamerican0817-24.html;   http://www.nature.
com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
24.pdf.

Pogue:2017:TTM

[3339] David Pogue. TechnoFiles: Technology as magic. Scientific American, 317(2):26, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
26.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-26.pdf.

Anonymous:2017:RRRh

[3340] Anonymous. Reviews and recommendations: Recommended. Scientific American, 317(2):80, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-

80.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-80.pdf.

Shermer:2017:SWA

[3341] Michael Shermer. Skeptic: Are we all racists? Scientific American, 317(2):81, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
81.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-81.pdf.

Mirsky:2017:AGWa

[3342] Steve Mirsky. Anti gravity: What’s the deal? Scientific American, 317(2):82, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
82.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-82.pdf.

Schlenoff:2017:YAIh

[3343] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in scientific American: 50, 100 & 150 years ago. Scientific American, 317(2):83, August 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n2/full/scientificamerican0817-83.html;   http://www.nature.
com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
83.pdf.

Fischetti:2017:GSR

[3344] Mark Fischetti. Graphic science: Reactors reshuffled. Scientific American, 317(2):84, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
84.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-84.pdf.

VanKranendonk:2017:CLS

[3345] Martin J. Van Kranendonk, David W. Deamer, and Tara Djokic. Chemistry: Life springs. Scientific American, 317(2):28–35, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n2/full/scientificamerican0817-28.html;   http://www.nature.

com/scientificamerican/journal/v317/n2/pdf/scientificamerican0817-
28.pdf.

Vance:2017:CRV

[3346] Erik Vance. Conservation: Requiem for the vaquita. Scientific American, 317(2):36–45, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
36.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-36.pdf.

Wong:2017:POC

[3347] Kate Wong. Paleoanthropology: Our cousin neo. Scientific American, 317(2):46–47, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
46.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-46.pdf.

Tyre:2017:ERM

[3348] Peg Tyre. Education report 2017: A matter of choice. Scientific American, 317(2):48–53, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
48.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-48.pdf.

Pasachoff:2017:AGS

[3349] Jay M. Pasachoff. Astrophysics: The great solar eclipse of 2017. Scientific American, 317(2):54–61, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
54.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-54.pdf.

Fischetti:2017:YSE

[3350] Mark M. Fischetti. 1,000 years of solar eclipses. Scientific American, 317(2):62–65, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
62.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-62.pdf.

Broadfoot:2017:ABB

[3351] Marla Broadfoot. Agriculture: Building a better harvest. Scientific American, 317(2):66–73, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
66.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-66.pdf.

Fernyhough:2017:PTO

[3352] Charles Fernyhough. Psychology: Talking to ourselves. Scientific American, 317(2):74–79, August 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n2/full/scientificamerican0817-
74.html;   http://www.nature.com/scientificamerican/journal/
v317/n2/pdf/scientificamerican0817-74.pdf.

DiChristina:2017:EWE

[3353] Mariette DiChristina. From the Editor: Welcome to everybody’s issue. Scientific American, 317(3):6, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v317/n3/full/scientificamerican0917-
6.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-6.pdf.

Anonymous:2017:Lj

[3354] Anonymous. Letters. Scientific American, 317(3):7–8, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n3/full/scientificamerican0917-7.html;   http://www.nature.
com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
7.pdf.

Schlenoff:2017:YAIi

[3355] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 317(3):94, September 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n3/full/scientificamerican0917-94.html;   http://www.nature.
com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
94.pdf.

Anonymous:2017:SAE

[3356] Anonymous. Science agenda: End the assault on women’s health. Scientific American, 317(3):9, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n3/full/scientificamerican0917-
9.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-9.pdf.

Manlan:2017:FAC

[3357] Carl Manlan. Forum: Africa’s CDC can end malaria. Scientific American, 317(3):10, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
10.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-10.pdf.

Wong:2017:AOH

[3358] Kate Wong. Advances: The oldest Homo sapiens? Scientific American, 317(3):12–14, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
12.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-12.pdf.

Peck:2017:ME

[3359] Morgen Peck. A moth’s eye. Scientific American, 317(3):14, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-14.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
14.pdf.

Sedacca:2017:SS

[3360] Matthew Sedacca. Skinlike sunscreen. Scientific American, 317(3):15, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-15.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
15.pdf.

Hu:2017:BR

[3361] Jane C. Hu. Baby’s rainbow. Scientific American, 317(3):16, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-16a.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
16a.pdf.

Patel:2017:CTF

[3362] Prachi Patel. Cutting through the fog. Scientific American, 317(3):16, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-16b.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
16b.pdf.

Nemo:2017:CH

[3363] Leslie Nemo. Change of heartbeat. Scientific American, 317(3):17, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-17.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
17.pdf.

Marks:2017:LL

[3364] Andrea Marks. Living large. Scientific American, 317(3):18, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-18.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
18.pdf.

Sneed:2017:WNS

[3365] Annie Sneed. Wood is the new steel. Scientific American, 317(3):19, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-19.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
19.pdf.

Nemo:2017:QHb

[3366] Leslie Nemo. Quick hits. Scientific American, 317(3):20, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n3/full/scientificamerican0917-20.html;   http://www.nature.

com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
20.pdf.

Greenwood:2017:NOG

[3367] Veronique Greenwood. Night owl genes. Scientific American, 317(3):21, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-21.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
21.pdf.

Marks:2017:NT

[3368] Andrea Marks. Navigating by touch. Scientific American, 317(3):22, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-22.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
22.pdf.

Schmidt:2017:SHP

[3369] Charles Schmidt. The science of health: Pediatric predicament. Scientific American, 317(3):24–25, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n3/full/scientificamerican0917-
24.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-24.pdf.

Pogue:2017:TCL

[3370] David Pogue. TechnoFiles: Cut that last cord. Scientific American, 317(3):28, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
28.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-28.pdf.

Anonymous:2017:RRRi

[3371] Anonymous. Reviews and recommendations: Recommended. Scientific American, 317(3):88, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
88.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-88.pdf.

Shermer:2017:SPV

[3372] Michael Shermer. Skeptic: Postmodernism vs. science. Scientific American, 317(3):90, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
90.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-90.pdf.

Mirsky:2017:AGFc

[3373] Steve Mirsky. Anti gravity: The face of evil. Scientific American, 317(3):92, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
92.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-92.pdf.

Fischetti:2017:GSE

[3374] Mark Fischetti. Graphic science: The end. Scientific American, 317(3):96, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
96.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-96.pdf.

Anonymous:2017:WIW

[3375] Anonymous. This is not a women’s issue: Why the new science of sex & gender matters for everyone. Scientific American, 317(3):30–31, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-30.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
30.pdf.

Fine:2017:NSS

[3376] Cordelia Fine and Mark A. Elgar. The new science of sex and gender: Promiscuous men, chaste women and other gender myths. Scientific American, 317(3):32–37, September 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n3/full/scientificamerican0917-32.html;   http://www.nature.
com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
32.pdf.

Denworth:2017:TFB

[3377] Lydia Denworth. Is there a “female” brain? Scientific American, 317(3):38–43, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
38.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-38.pdf.

Olson:2017:WSG

[3378] Kristina R. Olson. When sex and gender collide. Scientific American, 317(3):44–49, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
44.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-44.pdf.

Montanez:2017:BXX

[3379] Amanda Montanez. Beyond XX and XY. Scientific American, 317(3):50–51, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
50.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-50.pdf.

Stefanick:2017:JM

[3380] Marcia L. Stefanick. Not just for men. Scientific American, 317(3):52–57, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-52.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
52.pdf.

Gold:2017:LBR

[3381] Rachel Benson Gold and Megan K. Donovan. Life before Roe. Scientific American, 317(3):58–59, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n3/full/scientificamerican0917-
58.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-58.pdf.

Cimpian:2017:BT

[3382] Andrei Cimpian and Sarah-Jane Leslie. The brilliance trap. Scientific American, 317(3):60–65, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n3/full/scientificamerican0917-
60.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-60.pdf.

Saujani:2017:GC

[3383] Reshma Saujani. Girl code. Scientific American, 317(3):66–69, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-66.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
66.pdf.

Temple-Wood:2017:BT

[3384] Emily Temple-Wood. The blogger and the trolls. Scientific American, 317(3):70–71, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
70.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-70.pdf.

Revenga:2017:WW

[3385] Ana L. Revenga and Ana Maria Munoz Boudet. Women’s work. Scientific American, 317(3):72–77, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n3/full/scientificamerican0917-
72.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-72.pdf.

Montanez:2017:MG

[3386] Amanda Montañez. Mind the gap. Scientific American, 317(3):78–79, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n3/full/scientificamerican0917-78.html;   http:
//www.nature.com/scientificamerican/journal/v317/n3/pdf/scientificamerican0917-
78.pdf.

DasGupta:2017:RMD

[3387] Monica Das Gupta. Return of the missing daughters. Scientific American, 317(3):80–85, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-

80.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-80.pdf.

Schwartz:2017:WWS

[3388] Jen Schwartz. The woman who saved the planet. Scientific American, 317(3):86–87, September 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n3/full/scientificamerican0917-
86.html;   http://www.nature.com/scientificamerican/journal/
v317/n3/pdf/scientificamerican0917-86.pdf.

DiChristina:2017:EWW

[3389] Mariette DiChristina. From the Editor: What the world needs now. Scientific American, 317(4):5, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n4/full/scientificamerican1017-
5.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-5.pdf.

Anonymous:2017:Lk

[3390] Anonymous. Letters. Scientific American, 317(4):8–9, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n4/full/scientificamerican1017-8.html;   http://www.nature.
com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
8.pdf.

Schlenoff:2017:YAIj

[3391] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 317(4):93, October 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n4/full/scientificamerican1017-93.html;   http://www.nature.
com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
93.pdf.

Anonymous:2017:SAP

[3392] Anonymous. Science agenda: Put science back in Congress. Scientific American, 317(4):10, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-

10.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-10.pdf.

Sah:2017:FFS

[3393] Sunita Sah. Forum: Forensic science must be scientific. Scientific American, 317(4):12, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
12.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-12.pdf.

Dunietz:2017:ASP

[3394] Jesse Dunietz. Advances: Space prospecting. Scientific American, 317(4):14–16, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
14.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-14.pdf.

Goldman:2017:SC

[3395] Jason G. Goldman. Scaredy-cats. Scientific American, 317(4):18–19, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n4/full/scientificamerican1017-18a.html;   http:
//www.nature.com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
18a.pdf.

Das:2017:FM

[3396] Saswato R. Das. Fossil moon. Scientific American, 317(4):18–19, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n4/full/scientificamerican1017-18b.html;   http:
//www.nature.com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
18b.pdf.

Sneed:2017:GSS

[3397] Annie Sneed. Giant shape-shifters. Scientific American, 317(4):20, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n4/full/scientificamerican1017-20.html;   http:
//www.nature.com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
20.pdf.

Waldholz:2017:GC

[3398] Michael Waldholz. Genetic compromises. Scientific American, 317(4):21, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n4/full/scientificamerican1017-21.html;   http:
//www.nature.com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
21.pdf.

Nemo:2017:QHc

[3399] Leslie Nemo. Quick hits. Scientific American, 317(4):22, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n4/full/scientificamerican1017-22.html;   http://www.nature.
com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
22.pdf.

Sedacca:2017:BB

[3400] Matthew Sedacca. Better batteries. Scientific American, 317(4):23, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n4/full/scientificamerican1017-23.html;   http:
//www.nature.com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
23.pdf.

Hsu:2017:SS

[3401] Jeremy Hsu. Sailing on sunshine. Scientific American, 317(4):24–25, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n4/full/scientificamerican1017-24.html;   http:
//www.nature.com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
24.pdf.

Nemo:2017:MG

[3402] Leslie Nemo. Moving a giant. Scientific American, 317(4):26, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n4/full/scientificamerican1017-26.html;   http:
//www.nature.com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
26.pdf.

Yuhas:2017:SHB

[3403] Daisy Yuhas. The science of health: Back to basics. Scientific American, 317(4):28–29, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
28.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-28.pdf.

Pogue:2017:TYS

[3404] David Pogue. TechnoFiles: Your security cam is watching you. Scientific American, 317(4):30, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
30.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-30.pdf.

Anonymous:2017:RRRj

[3405] Anonymous. Reviews and recommendations: Recommended. Scientific American, 317(4):86, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
86.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-86.pdf.

Shermer:2017:SSG

[3406] Michael Shermer. Skeptic: Sky gods for skeptics. Scientific American, 317(4):88, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
88.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-88.pdf.

Mirsky:2017:AGWb

[3407] Steve Mirsky. Anti gravity: What the heck is that thing? Scientific American, 317(4):90, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
90.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-90.pdf.

Fischetti:2017:GSO

[3408] Mark Fischetti. Graphic science: Opioid deaths soar. Scientific American, 317(4):96, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-

96.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-96.pdf.

Moskowitz:2017:PPN

[3409] Clara Moskowitz. Particle physics: The neutrino puzzle. Scientific American, 317(4):32–39, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
32.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-32.pdf.

Conniff:2017:CLD

[3410] Richard Conniff. Conservation: Loved to death. Scientific American, 317(4):40–45, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
40.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-40.pdf.

Erlebacher:2017:DBF

[3411] Adrian Erlebacher and Susan J. Fisher. Development: Baby’s first organ. Scientific American, 317(4):46–53, October 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n4/full/scientificamerican1017-46.html;   http://www.nature.
com/scientificamerican/journal/v317/n4/pdf/scientificamerican1017-
46.pdf.

Moyer:2017:PHJ

[3412] Melinda Wenner Moyer. Public health: Journey to gunland. Scientific American, 317(4):54–63, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
54.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-54.pdf.

Schwartz:2017:SWS

[3413] Jen Schwartz, Brooke Borel, Inga Vesper, and Lee Billings. State of the world’s science 2017: Reason on the ropes. Scientific American, 317(4):64–73, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-

64.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-64.pdf.

Fitzharris:2017:HSD

[3414] Lindsey Fitzharris. History of science: Dangerous medicine. Scientific American, 317(4):74–77, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
74.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-74.pdf.

Porco:2017:PSC

[3415] Carolyn Porco. Planetary science: Cassini at Saturn. Scientific American, 317(4):78–85, October 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n4/full/scientificamerican1017-
78.html;   http://www.nature.com/scientificamerican/journal/
v317/n4/pdf/scientificamerican1017-78.pdf.

DiChristina:2017:EPW

[3416] Mariette DiChristina. From the Editor: Peering within. Scientific American, 317(5):6, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
6.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-6.pdf.

Anonymous:2017:Ll

[3417] Anonymous. Letters. Scientific American, 317(5):8–9, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n5/full/scientificamerican1117-8.html;   http://www.nature.
com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
8.pdf.

Anonymous:2017:SAG

[3418] Anonymous. Science agenda: Get toxic chemicals out of cosmetics. Scientific American, 317(5):10, November 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n5/full/scientificamerican1117-10.html;   http://www.nature.
com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
10.pdf.

Hart:2017:FPD

[3419] Carl L. Hart. Forum: People are not dying because of opioids. Scientific American, 317(5):12, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
11.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-11.pdf.

Kwon:2017:AYB

[3420] Diana Kwon. Advances: Your brain on (actual) acid. Scientific American, 317(5):12–14, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
12.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-12.pdf.

Hutson:2017:BR

[3421] Matthew Hutson. Bad reputation. Scientific American, 317(5):15, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n5/full/scientificamerican1117-15.html;   http:
//www.nature.com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
15.pdf.

Maron:2017:CV

[3422] Dina Fine Maron. Cancer vaccine. Scientific American, 317(5):16, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n5/full/scientificamerican1117-16.html;   http:
//www.nature.com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
16.pdf.

Sheikh:2017:SSY

[3423] Knvul Sheikh. Space-saving yeast. Scientific American, 317(5):17, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n5/full/scientificamerican1117-17.html;   http:
//www.nature.com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
17.pdf.

Sneed:2017:MDT

[3424] Annie Sneed. Mosquito death traps. Scientific American, 317(5):18, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n5/full/scientificamerican1117-18.html;   http:
//www.nature.com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
18.pdf.

Nemo:2017:PEP

[3425] Leslie Nemo. Poison-eating poplars. Scientific American, 317(5):19, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n5/full/scientificamerican1117-19.html;   http:
//www.nature.com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
19.pdf.

Hall:2017:FV

[3426] Shannon Hall. Flying into volcanoes. Scientific American, 317(5):20, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n5/full/scientificamerican1117-20.html;   http:
//www.nature.com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
20.pdf.

Nemo:2017:QHd

[3427] Leslie Nemo. Quick hits. Scientific American, 317(5):21, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n5/full/scientificamerican1117-21.html;   http://www.nature.
com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
21.pdf.

Sheikh:2017:AYH

[3428] Knvul Sheikh. Antifreeze for your heart. Scientific American, 317(5):22, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n5/full/scientificamerican1117-22.html;   http:
//www.nature.com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
22.pdf.

Wallis:2017:SHH

[3429] Claudia Wallis. The science of health: Health by the numbers? Scientific American, 317(5):24, November 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n5/full/scientificamerican1117-24.html;   http://www.nature.

com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
24.pdf.

Pogue:2017:TFF

[3430] David Pogue. TechnoFiles: Fighting format rot. Scientific American, 317(5):26, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
26.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-26.pdf.

Anonymous:2017:RRRk

[3431] Anonymous. Reviews and recommendations: Recommended. Scientific American, 317(5):77, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
77.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-77.pdf.

Shermer:2017:SWS

[3432] Michael Shermer. Skeptic: What is the secret of success? Scientific American, 317(5):80, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
80.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-80.pdf.

Mirsky:2017:AGPb

[3433] Steve Mirsky. Anti gravity: Poultry problems. Scientific American, 317(5):82, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
82.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-82.pdf.

Schlenoff:2017:YAIk

[3434] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 317(5):83, November 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n5/full/scientificamerican1117-83.html;   http://www.nature.

com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
83.pdf.

Fischetti:2017:GSSb

[3435] Mark Fischetti and Jan Tulp. Graphic science: Safer by kindergarten. Scientific American, 317(5):84, November 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n5/full/scientificamerican1117-84.html;   http://www.nature.
com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
84.pdf.

Koch:2017:NHM

[3436] Christof Koch. Neuroscience: How to make a consciousness meter. Scientific American, 317(5):28–33, November 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n5/full/scientificamerican1117-28.html;   http://www.nature.
com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
28.pdf.

Olson:2017:GNB

[3437] Steve Olson. Geology: The next big bang. Scientific American, 317(5):34–41, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
34.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-34.pdf.

Conniff:2017:EGF

[3438] Richard Conniff. Ecology: Gone fishing. Scientific American, 317(5):42–47, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
42.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-42.pdf.

Gies:2017:SRG

[3439] Erica Gies. Sustainability: The radical groundwater storage test. Scientific American, 317(5):48–55, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n5/full/scientificamerican1117-

48.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-48.pdf.

Wilkinson:2017:PPM

[3440] Guy Wilkinson. Particle physics: Measuring beauty. Scientific American, 317(5):56–63, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
56.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-56.pdf.

Sauer:2017:BWS

[3441] Karin Sauer. Biochemistry: The war on slime. Scientific American, 317(5):64–69, November 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n5/full/scientificamerican1117-
64.html;   http://www.nature.com/scientificamerican/journal/
v317/n5/pdf/scientificamerican1117-64.pdf.

Scharf:2017:CZU

[3442] Caleb Scharf and Ron Miller. Cosmology: The zoomable universe. Scientific American, 317(5):70–75, November 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n5/full/scientificamerican1117-70.html;   http://www.nature.
com/scientificamerican/journal/v317/n5/pdf/scientificamerican1117-
70.pdf.

DiChristina:2017:ELS

[3443] Mariette DiChristina. From the Editor: Let’s science that. Scientific American, 317(6):4, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
4.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-4.pdf.

Anonymous:2017:Lm

[3444] Anonymous. Letters. Scientific American, 317(6):5–6, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n6/full/scientificamerican1217-5.html;   http://www.nature.
com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
5.pdf.

Schlenoff:2017:YAIl

[3445] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 317(6):92, December 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n6/full/scientificamerican1217-92.html;   http://www.nature.
com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
92.pdf.

Anonymous:2017:SAI

[3446] Anonymous. Science agenda: Ice under siege. Scientific American, 317(6):7, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
7.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-7.pdf.

Ghodrati:2017:FDD

[3447] Sahand Ghodrati. Forum: DACA’s demise is bad for our health. Scientific American, 317(6):10, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v317/n6/full/scientificamerican1217-
10.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-10.pdf.

Hall:2017:ADP

[3448] Shannon Hall. Advances: Dawn of plate tectonics. Scientific American, 317(6):12–14, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
12.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-12.pdf.

Pycha:2017:SC

[3449] Anne Pycha. Speaking in clicks. Scientific American, 317(6):15, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n6/full/scientificamerican1217-15.html;   http:
//www.nature.com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
15.pdf.

Choi:2017:QL

[3450] Charles Q. Choi. Quantum leaps. Scientific American, 317(6):16, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n6/full/scientificamerican1217-16.html;   http:
//www.nature.com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
16.pdf.

Main:2017:SW

[3451] Doug Main. Smoky wine. Scientific American, 317(6):17, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n6/full/scientificamerican1217-17.html;   http://www.nature.
com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
17.pdf.

Nuwer:2017:HSH

[3452] Rachel Nuwer. Healthy see, healthy do. Scientific American, 317(6):18, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n6/full/scientificamerican1217-18a.html;   http:
//www.nature.com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
18a.pdf.

Goldman:2017:WL

[3453] Jason G. Goldman. Wag the lizard. Scientific American, 317(6):18, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n6/full/scientificamerican1217-18b.html;   http:
//www.nature.com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
18b.pdf.

Saplakoglu:2017:QH

[3454] Yasemin Saplakoglu. Quick hits. Scientific American, 317(6):19, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n6/full/scientificamerican1217-19.html;   http:
//www.nature.com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
19.pdf.

Sutherland:2017:FH

[3455] Stephani Sutherland. Food high. Scientific American, 317(6):20–21, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n6/full/scientificamerican1217-20a.html;   http:
//www.nature.com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
20a.pdf.

Makin:2017:RRR

[3456] Simon Makin. Reprogram, restore, regenerate. Scientific American, 317(6):20–21, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
20b.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-20b.pdf.

Hsu:2017:SSD

[3457] Jeremy Hsu. Solar storm Doomsday? Scientific American, 317(6):22, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n6/full/scientificamerican1217-22.html;   http:
//www.nature.com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
22.pdf.

Wallis:2017:SHM

[3458] Claudia Wallis. The science of health: Marijuana and the teen brain. Scientific American, 317(6):25, December 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n6/full/scientificamerican1217-25.html;   http://www.nature.
com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
25.pdf.

Pogue:2017:TDS

[3459] David Pogue. TechnoFiles: The digital-subscription dilemma. Scientific American, 317(6):26, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
26.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-26.pdf.

Anonymous:2017:RRRl

[3460] Anonymous. Reviews and recommendations: Recommended. Scientific American, 317(6):86, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-

86.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-86.pdf.

Shermer:2017:SOW

[3461] Michael Shermer. Skeptic: Outlawing war. Scientific American, 317(6):88, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
88.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-88.pdf.

Mirsky:2017:AGS

[3462] Steve Mirsky. Anti gravity: Squirrels that chunk. Scientific American, 317(6):90, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
90.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-90.pdf.

Billings:2017:GSM

[3463] Lee Billings. Graphic science: Meet the neighbors. Scientific American, 317(6):94, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
94.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-94.pdf.

DiChristina:2017:SRT

[3464] Mariette DiChristina and Bernard S. Meyerson. Special report: Top 10 emerging technologies of 2017. Scientific American, 317(6):28–39, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v317/n6/full/scientificamerican1217-28.html;   http:
//www.nature.com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
28.pdf.

Stern:2017:PSP

[3465] S. Alan Stern. Planetary science: Pluto revealed. Scientific American, 317(6):40–47, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
40.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-40.pdf.

Hall:2017:GCG

[3466] Stephen S. Hall. Genetics: Could genetic engineering save the Galápagos? Scientific American, 317(6):48–57, December 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n6/full/scientificamerican1217-48.html;   http://www.nature.
com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
48.pdf.

Denworth:2017:PFY

[3467] Lydia Denworth. Psychology: I feel your pain. Scientific American, 317(6):58–63, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
58.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-58.pdf.

Nordhaus:2017:CBR

[3468] Hannah Nordhaus. Conservation: Beetle resurrection. Scientific American, 317(6):64–71, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
64.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-64.pdf.

Marvel:2017:CCC

[3469] Kate Marvel. Climate: The cloud conundrum. Scientific American, 317(6):72–77, December 2017. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v317/n6/full/scientificamerican1217-
72.html;   http://www.nature.com/scientificamerican/journal/
v317/n6/pdf/scientificamerican1217-72.pdf.

Cameron:2017:AHC

[3470] Catherine M. Cameron. Archaeology: How captives changed the world. Scientific American, 317(6):78–83, December 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v317/
n6/full/scientificamerican1217-78.html;   http://www.nature.
com/scientificamerican/journal/v317/n6/pdf/scientificamerican1217-
78.pdf.

DiChristina:2018:ES

[3471] Mariette DiChristina. From the Editor: Science is …. Scientific American, 318(1):4, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
4.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-4.pdf.

Anonymous:2018:La

[3472] Anonymous. Letters. Scientific American, 318(1):5–6, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-5.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
5.pdf.

Anonymous:2018:SASa

[3473] Anonymous. Science agenda: A school is not a military post. Scientific American, 318(1):7, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
7.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-7.pdf.

Rosenberg:2018:FWS

[3474] Andrew A. Rosenberg and Kathleen Rest. Forum: War on science agencies. Scientific American, 318(1):8, January 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-8.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
8.pdf.

Billings:2018:ACM

[3475] Lee Billings. Advances: Cosmic messengers. Scientific American, 318(1):10–12, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
10.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-10.pdf.

Saplakoglu:2018:GFW

[3476] Yasemin Saplakoglu. Gluten-free wheat? Scientific American, 318(1):14– 15, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n1/full/scientificamerican0118-14a.html;   http:
//www.nature.com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
14a.pdf.

Nuwer:2018:DW

[3477] Rachel Nuwer. The deadliest weapon. Scientific American, 318(1):14–15, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n1/full/scientificamerican0118-14b.html;   http:
//www.nature.com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
14b.pdf.

Kwon:2018:SS

[3478] Diana Kwon. Sea smarts. Scientific American, 318(1):16, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-16.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
16.pdf.

RodriguezMega:2018:PG

[3479] Emiliano Rodríguez Mega. The price of gold. Scientific American, 318(1): 18, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n1/full/scientificamerican0118-18.html;   http:
//www.nature.com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
18.pdf.

Yuhas:2018:RC

[3480] Daisy Yuhas. Ring cycle. Scientific American, 318(1):19, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-19.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
19.pdf.

Saplakoglu:2018:QHa

[3481] Yasemin Saplakoglu. Quick hits. Scientific American, 318(1):20, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n1/full/scientificamerican0118-20.html;   http:
//www.nature.com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
20.pdf.

Patel:2018:CB

[3482] Prachi Patel. Cobalt blues. Scientific American, 318(1):21, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-21.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
21.pdf.

Wallis:2018:SHM

[3483] Claudia Wallis. The science of health: The messy facts about diet and inflammation. Scientific American, 318(1):22, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-22.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
22.pdf.

Pogue:2018:THW

[3484] David Pogue. TechnoFiles: How well do films predict our tech future? Scientific American, 318(1):23, January 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-23.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
23.pdf.

Anonymous:2018:RRRa

[3485] Anonymous. Reviews and recommendations: Recommended. Scientific American, 318(1):76, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
76.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-76.pdf.

Shermer:2018:SLS

[3486] Michael Shermer. Skeptic: For the love of science. Scientific American, 318(1):77, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
77.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-77.pdf.

Mirsky:2018:AGA

[3487] Steve Mirsky. Anti gravity: Is anybody anywhere? Scientific American, 318(1):78, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
78.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-78.pdf.

Schlenoff:2018:YAIa

[3488] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 318(1):79, January 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-79.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
79.pdf.

Fischetti:2018:GSK

[3489] Mark Fischetti. Graphic science: Killer seas. Scientific American, 318(1): 80, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n1/full/scientificamerican0118-80.html;   http:
//www.nature.com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
80.pdf.

Anonymous:2018:SRF

[3490] Anonymous. Special report: The future of money. Scientific American, 318(1):24–25, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
24.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-24.pdf.

Lipton:2018:BB

[3491] Alexander Lipton and Alex ‘Sandy’ Pentland. Breaking the bank: New financial networks could stop the concentration of wealth and increased participation in the economy — but only if used with care. Scientific American, 318(1):26–31, January 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-26.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
26.pdf.

Pavlus:2018:WBC

[3492] John Pavlus. The world Bitcoin created: The first big digital currency gave us a glimpse of a new economic order — one that raises more questions than it answers. Scientific American, 318(1):32–37, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-32.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
32.pdf.

Smolenski:2018:ETU

[3493] Natalie Smolenski. The evolution of trust: The ultimate social impact of blockchain technology depends on who controls our digital identities. Scientific American, 318(1):38–41, January 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n1/full/scientificamerican0118-38.html;   http://www.nature.
com/scientificamerican/journal/v318/n1/pdf/scientificamerican0118-
38.pdf.

Albright:2018:SCW

[3494] Rebecca Albright. Sustainability: Can we save the corals? Scientific American, 318(1):42–49, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
42.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-42.pdf.

Rosenberg:2018:ASD

[3495] Leslie Rosenberg. Astrophysics: Searching for the dark. Scientific American, 318(1):50–57, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
50.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-50.pdf.

Waldholz:2018:MWA

[3496] Michael Waldholz. Medicine: War against ourselves. Scientific American, 318(1):58–63, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
58.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-58.pdf.

Russo:2018:PTW

[3497] Francine Russo. Psychology: The toxic well of loneliness. Scientific American, 318(1):64–69, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
64.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-64.pdf.

Yi:2018:EHS

[3498] Hongyu Yi. Evolution: How snakes came to slither. Scientific American, 318(1):70–75, January 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n1/full/scientificamerican0118-
70.html;   http://www.nature.com/scientificamerican/journal/
v318/n1/pdf/scientificamerican0118-70.pdf.

DiChristina:2018:EDS

[3499] Mariette DiChristina. From the Editor: Depths of space, oceans and politics. Scientific American, 318(2):3, February 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-3.html;   http://www.nature.
com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
3.pdf.

Anonymous:2018:Lb

[3500] Anonymous. Letters. Scientific American, 318(2):4–5, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-4.html;   http://www.nature.
com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
4.pdf.

Schlenoff:2018:YAIb

[3501] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 318(2):69, February 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-69.html;   http://www.nature.
com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
69.pdf.

Anonymous:2018:SAG

[3502] Anonymous. Science agenda: Go public or perish. Scientific American, 318(2):6, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
6.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-6.pdf.

Paules:2018:FUF

[3503] Catharine I. Paules and Anthony S. Fauci. Forum: A universal flu vaccine is vital. Scientific American, 318(2):8, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-8.html;   http://www.nature.
com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
8.pdf.

Sokol:2018:AEN

[3504] Joshua Sokol. Advances: The end of night. Scientific American, 318(2):10–12, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
10.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-10.pdf.

Goldman:2018:FF

[3505] Jason G. Goldman. Fruitless foragers. Scientific American, 318(2):12–13, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n2/full/scientificamerican0218-12.html;   http:
//www.nature.com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
12.pdf.

Berkowitz:2018:DR

[3506] Rachel Berkowitz. Diamond in the rough. Scientific American, 318(2):13, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n2/full/scientificamerican0218-13.html;   http:
//www.nature.com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
13.pdf.

Main:2018:SSM

[3507] Doug Main. Salmon sex moves mountains. Scientific American, 318(2):14, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
14a.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-14a.pdf.

Peek:2018:MS

[3508] Katie Peek. Moon shot. Scientific American, 318(2):14, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-14b.html;   http://www.nature.
com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
14b.pdf.

Patel:2018:RP

[3509] Prachi Patel. Rivers of plastic. Scientific American, 318(2):15–17, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n2/full/scientificamerican0218-15a.html;   http:
//www.nature.com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
15a.pdf.

Moskowitz:2018:PDL

[3510] Clara Moskowitz. Porpoise-driven life. Scientific American, 318 (2):15–17, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
15b.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-15b.pdf.

Sneed:2018:GS

[3511] Annie Sneed. Greener skies. Scientific American, 318(2):18, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n2/full/scientificamerican0218-18.html;   http:
//www.nature.com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
18.pdf.

Saplakoglu:2018:WD

[3512] Yasemin Saplakoglu. Wearable data. Scientific American, 318(2):20, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n2/full/scientificamerican0218-20.html;   http:
//www.nature.com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
20.pdf.

Saplakoglu:2018:QHb

[3513] Yasemin Saplakoglu. Quick hits. Scientific American, 318(2):21, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n2/full/scientificamerican0218-21.html;   http:
//www.nature.com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
21.pdf.

Wallis:2018:SHWa

[3514] Claudia Wallis. The science of health: Why fake operations are a good thing. Scientific American, 318(2):22, February 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-22.html;   http://www.nature.
com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
22.pdf.

Pogue:2018:TRA

[3515] David Pogue. TechnoFiles: The robotic artist problem. Scientific American, 318(2):23, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
23.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-23.pdf.

Anonymous:2018:RRRb

[3516] Anonymous. Reviews and recommendations: Recommended. Scientific American, 318(2):66, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-

66.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-66.pdf.

Shermer:2018:SAE

[3517] Michael Shermer. Skeptic: Alvy’s error and the meaning of life. Scientific American, 318(2):67, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
67.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-67.pdf.

Mirsky:2018:AGBa

[3518] Steve Mirsky. Anti gravity: Bee sides. Scientific American, 318(2):68, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n2/full/scientificamerican0218-68.html;   http:
//www.nature.com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
68.pdf.

Peek:2018:GSM

[3519] Katie Peek. Graphic science: Misled penguins. Scientific American, 318(2):72, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
72.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-72.pdf.

Natarajan:2018:AFM

[3520] Priyamvada Natarajan. Astrophysics: The first monster black holes. Scientific American, 318(2):24–29, February 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-24.html;   http://www.nature.
com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
24.pdf.

Flora:2018:TSR

[3521] Carlin Flora. Technology: Are Smartphones really destroying the adolescent brain? Scientific American, 318(2):30–37, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-30.html;   http://www.nature.

com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
30.pdf.

Seife:2018:PER

[3522] Charles Seife. Policy and ethics: [redacted]. Scientific American, 318(2):38–43, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
38.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-38.pdf.

Heffernan:2018:BTW

[3523] Olive Heffernan. Biodiversity: Troubled Waters. Scientific American, 318(2):44–49, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
44.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-44.pdf.

Fisher:2018:CST

[3524] Matthew Fisher, Joshua Knobe, Brent Strickland, and Frank C. Keil. Cognitive science: The tribalism of truth. Scientific American, 318(2):50–53, February 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n2/full/scientificamerican0218-
50.html;   http://www.nature.com/scientificamerican/journal/
v318/n2/pdf/scientificamerican0218-50.pdf.

Drescher:2018:MVD

[3525] Kristen M. Drescher and Steven Tracy. Medicine: Vanquishing diabetes. Scientific American, 318(2):54–59, February 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-54.html;   http://www.nature.
com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
54.pdf.

Overdorf:2018:CGD

[3526] Jason Overdorf. Conservation: Guardian dogs of the Mongolian steppe. Scientific American, 318(2):60–65, February 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n2/full/scientificamerican0218-60.html;   http://www.nature.

com/scientificamerican/journal/v318/n2/pdf/scientificamerican0218-
60.pdf.

DiChristina:2018:ECP

[3527] Mariette DiChristina. From the Editor: Child’s play. Scientific American, 318(3):4, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
4.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-4.pdf.

Anonymous:2018:Lc

[3528] Anonymous. Letters. Scientific American, 318(3):6–7, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n3/full/scientificamerican0318-6.html;   http://www.nature.
com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
6.pdf.

Schlenoff:2018:YAIc

[3529] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 318(3):75, March 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n3/full/scientificamerican0318-75.html;   http://www.nature.
com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
75.pdf.

Anonymous:2018:SASb

[3530] Anonymous. Science agenda: Science suffers from harassment. Scientific American, 318(3):8, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
8.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-8.pdf.

Rosenberg:2018:FTR

[3531] Andrew A. Rosenberg. Forum: Time for a reboot in Congress. Scientific American, 318(3):10, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-

10.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-10.pdf.

Jacobson:2018:AMT

[3532] Roni Jacobson. Advances: Mother tongue. Scientific American, 318(3):12–14, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
12.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-12.pdf.

Shen:2018:BE

[3533] Helen Shen. Brain enigma. Scientific American, 318(3):15, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n3/full/scientificamerican0318-15.html;   http://www.nature.
com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
15.pdf.

Harris:2018:IDD

[3534] Mark Harris. Ice-diving drones. Scientific American, 318(3):16, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n3/full/scientificamerican0318-16a.html;   http:
//www.nature.com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
16a.pdf.

Saplakoglu:2018:BTR

[3535] Yasemin Saplakoglu. Bacterial tape recorder. Scientific American, 318(3):16–17, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
16b.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-16b.pdf.

Goldman:2018:BB

[3536] Jason G. Goldman. Bird breakup. Scientific American, 318(3):18, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n3/full/scientificamerican0318-18a.html;   http:
//www.nature.com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
18a.pdf.

Anonymous:2018:FF

[3537] Anonymous. Fish frenzy. Scientific American, 318(3):18–19, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n3/full/scientificamerican0318-18b.html;   http://www.nature.
com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
18b.pdf.

Hutson:2018:IG

[3538] Matthew Hutson. Ineffective geniuses? Scientific American, 318(3):20, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n3/full/scientificamerican0318-20a.html;   http:
//www.nature.com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
20a.pdf.

Choi:2018:ISD

[3539] Charles Q. Choi. “I’m sorry, Dave”. Scientific American, 318(3):20, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n3/full/scientificamerican0318-20b.html;   http:
//www.nature.com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
20b.pdf.

Main:2018:FB

[3540] Doug Main. Fishy behemoth. Scientific American, 318(3):21, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n3/full/scientificamerican0318-21.html;   http:
//www.nature.com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
21.pdf.

Saplakoglu:2018:QHc

[3541] Yasemin Saplakoglu. Quick hits. Scientific American, 318(3):22, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n3/full/scientificamerican0318-22.html;   http:
//www.nature.com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
22.pdf.

Wallis:2018:SHP

[3542] Claudia Wallis. The science of health: A perfectly avoidable crisis. Scientific American, 318(3):24, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
24.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-24.pdf.

Pogue:2018:TVD

[3543] David Pogue. TechnoFiles: Video disorientation. Scientific American, 318(3):25, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
25.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-25.pdf.

Anonymous:2018:RRRc

[3544] Anonymous. Reviews and recommendations: Recommended. Scientific American, 318(3):72, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
72.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-72.pdf.

Shermer:2018:SF

[3545] Michael Shermer. Skeptic: Factiness. Scientific American, 318(3):73, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n3/full/scientificamerican0318-73.html;   http:
//www.nature.com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
73.pdf.

Mirsky:2018:AGSa

[3546] Steve Mirsky. Anti gravity: A staggering discovery. Scientific American, 318(3):74, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
74.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-74.pdf.

Peek:2018:GSC

[3547] Katie Peek. Graphic science: The case for daylight saving time. Scientific American, 318(3):76, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-

76.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-76.pdf.

Kwon:2018:AIS

[3548] Diana Kwon. Artificial intelligence: Self-taught robots. Scientific American, 318(3):26–31, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
26.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-26.pdf.

Winn:2018:PSS

[3549] Joshua N. Winn. Planetary science: Shadows of other worlds. Scientific American, 318(3):32–39, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
32.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-32.pdf.

Erickson:2018:EBB

[3550] Gregory M. Erickson. Evolution: The baddest bite. Scientific American, 318(3):40–45, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
40.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-40.pdf.

Dullmann:2018:CIH

[3551] Christoph E. Dullmann and Michael Block. Chemistry: Island of heavyweights. Scientific American, 318(3):46–53, March 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n3/full/scientificamerican0318-46.html;   http://www.nature.
com/scientificamerican/journal/v318/n3/pdf/scientificamerican0318-
46.pdf.

Gabrieli:2018:NLW

[3552] John Gabrieli. Neuroscience: A look within. Scientific American, 318(3):54–59, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
54.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-54.pdf.

Weintraub:2018:MMR

[3553] Karen Weintraub. Medicine: The means of reproduction. Scientific American, 318(3):60–65, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
60.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-60.pdf.

Embry:2018:ABB

[3554] Paige Embry. Agriculture: Building a backup bee. Scientific American, 318(3):66–71, March 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n3/full/scientificamerican0318-
66.html;   http://www.nature.com/scientificamerican/journal/
v318/n3/pdf/scientificamerican0318-66.pdf.

DiChristina:2018:EBBa

[3555] Mariette DiChristina. From the Editor: Big blue marble. Scientific American, 318(4):4, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
4.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-4.pdf.

Anonymous:2018:Ld

[3556] Anonymous. Letters. Scientific American, 318(4):6–7, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n4/full/scientificamerican0418-6.html;   http://www.nature.
com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
6.pdf.

Schlenoff:2018:YAId

[3557] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 318(4):79, April 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n4/full/scientificamerican0418-79.html;   http://www.nature.
com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
79.pdf.

Anonymous:2018:SAH

[3558] Anonymous. Science agenda: How to stop the epidemic of suicides. Scientific American, 318(4):10, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
10.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-10.pdf.

Elvis:2018:FAP

[3559] Martin Elvis. Forum: Has astronomy peaked? Scientific American, 318(4):11, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
11.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-11.pdf.

Greenemeier:2018:ADB

[3560] Lawrence Greenemeier. Advances: Don’t believe your eyes. Scientific American, 318(4):12–14, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
12.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-12.pdf.

Goldman:2018:BLB

[3561] Jason G. Goldman. Bonobos like bullies. Scientific American, 318(4):14, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-14.html;   http:
//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
14.pdf.

Goldman:2018:NN

[3562] Jason G. Goldman. Nature and nurture. Scientific American, 318(4):16– 17, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-16a.html;   http:
//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
16a.pdf.

Hall:2018:HR

[3563] Shannon Hall. Hot rocks. Scientific American, 318(4):16–17, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n4/full/scientificamerican0418-16b.html;   http://www.nature.
com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
16b.pdf.

Makin:2018:BB

[3564] Simon Makin. The brain’s brakes. Scientific American, 318(4):18, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-18.html;   http:
//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
18.pdf.

Saplakoglu:2018:QHd

[3565] Yasemin Saplakoglu. Quick hits. Scientific American, 318(4):19, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-19.html;   http:
//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
19.pdf.

Kornei:2018:GS

[3566] Katherine Kornei. Glacial sprint. Scientific American, 318(4):20–21, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-20a.html;   http:
//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
20a.pdf.

Nuwer:2018:DC

[3567] Rachel Nuwer. DNA clock. Scientific American, 318(4):20–21, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-20b.html;   http:
//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
20b.pdf.

Amos:2018:BI

[3568] Amy Mathews Amos. Bird, interrupted. Scientific American, 318(4):22, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-22.html;   http:

//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
22.pdf.

Perkins:2018:BB

[3569] Sid Perkins. Back to black. Scientific American, 318(4):23, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n4/full/scientificamerican0418-23a.html;   http://www.nature.
com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
23a.pdf.

Berkowitz:2018:FMS

[3570] Rachel Berkowitz. # Flu: Mining social media to predict outbreaks. Scientific American, 318(4):23, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
23b.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-23b.pdf.

Wallis:2018:SHWb

[3571] Claudia Wallis. The science of health: Why a lethal cancer is on the rise. Scientific American, 318(4):24, April 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n4/full/scientificamerican0418-24.html;   http://www.nature.
com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
24.pdf.

Pogue:2018:TAT

[3572] David Pogue. TechnoFiles: Automotive touch screens are awful. Scientific American, 318(4):25, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
25.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-25.pdf.

Anonymous:2018:RRRd

[3573] Anonymous. Reviews and recommendations: Recommended. Scientific American, 318(4):76, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
76.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-76.pdf.

Shermer:2018:SSN

[3574] Michael Shermer. Skeptic: Silent no more. Scientific American, 318(4): 77, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-77.html;   http:
//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
77.pdf.

Mirsky:2018:AGPa

[3575] Steve Mirsky. Anti gravity: Problem people. Scientific American, 318(4): 78, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n4/full/scientificamerican0418-78.html;   http:
//www.nature.com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
78.pdf.

Fischetti:2018:GSWa

[3576] Mark Fischetti. Graphic science: Where the reptiles are. Scientific American, 318(4):80, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
80.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-80.pdf.

Mitra:2018:MBP

[3577] Aditee Mitra. Marine biology: The perfect beast. Scientific American, 318(4):26–33, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
26.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-26.pdf.

Townsend:2018:MCT

[3578] Jeffrey P. Townsend. Medicine: The cancer tree. Scientific American, 318(4):34–41, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
34.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-34.pdf.

Lorimer:2018:AFN

[3579] Duncan Lorimer and Maura McLaughlin. Astronomy: Flashes in the night. Scientific American, 318(4):42–47, April 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n4/full/scientificamerican0418-42.html;   http://www.nature.
com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
42.pdf.

Francis:2018:EM

[3580] Jennifer A. Francis. Environment: Meltdown. Scientific American, 318(4):48–53, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
48.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-48.pdf.

Denworth:2018:MHP

[3581] Lydia Denworth. Mental health: Preventing suicide. Scientific American, 318(4):54–59, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
54.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-54.pdf.

Fox:2018:NBR

[3582] Douglas Fox. Neuroscience: The brain, reimagined. Scientific American, 318(4):60–67, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n4/full/scientificamerican0418-
60.html;   http://www.nature.com/scientificamerican/journal/
v318/n4/pdf/scientificamerican0418-60.pdf.

Moulton:2018:MHS

[3583] Derek E. Moulton, Alain Goriely, and Régis Chirat. Mathematics: How seashells take shape. Scientific American, 318(4):68–75, April 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n4/full/scientificamerican0418-68.html;   http://www.nature.
com/scientificamerican/journal/v318/n4/pdf/scientificamerican0418-
68.pdf.

DiChristina:2018:EOP

[3584] Mariette DiChristina. From the Editor: Our planet, ourselves. Scientific American, 318(5):4, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
4.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-4.pdf.

Anonymous:2018:Le

[3585] Anonymous. Letters. Scientific American, 318(5):6–8, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n5/full/scientificamerican0518-6.html;   http://www.nature.
com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
6.pdf.

Schlenoff:2018:YAIe

[3586] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 318(5):83, May 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n5/full/scientificamerican0518-83.html;   http://www.nature.
com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
83.pdf.

Anonymous:2018:SAE

[3587] Anonymous. Science agenda: End the war on weed. Scientific American, 318(5):10, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
10.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-10.pdf.

Einstein:2018:FSO

[3588] Lisa Einstein. Forum: The suns in our daughters. Scientific American, 318(5):12, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
12.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-12.pdf.

Vesper:2018:ABS

[3589] Inga Vesper. Advances: Batty schedules. Scientific American, 318(5):14– 17, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n5/full/scientificamerican0518-14.html;   http:

//www.nature.com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
14.pdf.

Tennenhouse:2018:WP

[3590] Erica Tennenhouse. Watchful plants. Scientific American, 318(5):16– 17, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n5/full/scientificamerican0518-16.html;   http:
//www.nature.com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
16.pdf.

Joel:2018:MB

[3591] Lucas Joel. Missing bugs. Scientific American, 318(5):18, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n5/full/scientificamerican0518-18.html;   http://www.nature.
com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
18.pdf.

Saplakoglu:2018:QHe

[3592] Yasemin Saplakoglu. Quick hits. Scientific American, 318(5):19, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n5/full/scientificamerican0518-19.html;   http:
//www.nature.com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
19.pdf.

Henderson:2018:EE

[3593] Bob Henderson. Exoplanet Everests. Scientific American, 318(5):20, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n5/full/scientificamerican0518-20.html;   http:
//www.nature.com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
20.pdf.

Saplakoglu:2018:ST

[3594] Yasemin Saplakoglu. Sweaty trees. Scientific American, 318(5):21, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n5/full/scientificamerican0518-21.html;   http:
//www.nature.com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
21.pdf.

Main:2018:OM

[3595] Doug Main. Orangutan medicine. Scientific American, 318(5):22–23, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n5/full/scientificamerican0518-22a.html;   http:
//www.nature.com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
22a.pdf.

Nuwer:2018:TT

[3596] Rachel Nuwer. Timber trafficking. Scientific American, 318(5):22–23, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n5/full/scientificamerican0518-22b.html;   http:
//www.nature.com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
22b.pdf.

Montanez:2018:VES

[3597] Amanda Montañez. Views of evolution shaped by knowledge. Scientific American, 318(5):24, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
24.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-24.pdf.

Wallis:2018:SHBa

[3598] Claudia Wallis. The science of health: The battle of the belt. Scientific American, 318(5):25, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
25.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-25.pdf.

Pogue:2018:TWH

[3599] David Pogue. TechnoFiles: Weirdest hardware product ever? Scientific American, 318(5):26, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
26.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-26.pdf.

Anonymous:2018:RRRe

[3600] Anonymous. Reviews and recommendations: Recommended. Scientific American, 318(5):78, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
78.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-78.pdf.

Shermer:2018:SYK

[3601] Michael Shermer. Skeptic: You Kant be serious. Scientific American, 318 (5):80, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n5/full/scientificamerican0518-80.html;   http:
//www.nature.com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
80.pdf.

Mirsky:2018:AGM

[3602] Steve Mirsky. Anti gravity: Mothering, wild style. Scientific American, 318(5):82, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
82.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-82.pdf.

Fischetti:2018:GSWb

[3603] Mark Fischetti. Graphic science: A wider life gap. Scientific American, 318(5):84, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
84.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-84.pdf.

Brusatte:2018:EUT

[3604] Stephen Brusatte. Evolution: The unlikely triumph of dinosaurs. Scientific American, 318(5):28–35, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
28.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-28.pdf.

Finkbeiner:2018:AMS

[3605] Ann Finkbeiner. Astronomy: Messengers from the sky. Scientific American, 318(5):36–41, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-

36.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-36.pdf.

Anonymous:2018:FME

[3606] Anonymous. The future of medicine 2018: Emerging disease in a changing world. Scientific American, 318(5):42–43, May 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n5/full/scientificamerican0518-42.html;   http://www.nature.
com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
42.pdf.

Moyer:2018:PHA

[3607] Melinda Wenner Moyer. Public health: American epidemic. Scientific American, 318(5):44–57, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
44.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-44.pdf.

Parshley:2018:ECF

[3608] Lois Parshley. Epidemiology: Catching fever. Scientific American, 318(5):58–65, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
58.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-58.pdf.

Russo:2018:POS

[3609] Francine Russo. Psychology: Our stuff, ourselves. Scientific American, 318(5):66–71, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n5/full/scientificamerican0518-
66.html;   http://www.nature.com/scientificamerican/journal/
v318/n5/pdf/scientificamerican0518-66.pdf.

Peacock:2018:NRD

[3610] Thomas Peacock and Matthew H. Alford. Natural resources: Is deepsea mining worth it? Scientific American, 318(5):72–77, May 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n5/full/scientificamerican0518-72.html;   http://www.nature.

com/scientificamerican/journal/v318/n5/pdf/scientificamerican0518-
72.pdf.

DiChristina:2018:ESS

[3611] Mariette DiChristina. From the Editor: The Sun, supercharging and the search. Scientific American, 318(6):6, June 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n6/full/scientificamerican0618-6.html;   http://www.nature.
com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
6.pdf.

Anonymous:2018:Lf

[3612] Anonymous. Letters. Scientific American, 318(6):8–9, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n6/full/scientificamerican0618-8.html;   http://www.nature.
com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
8.pdf.

Schlenoff:2018:YAIf

[3613] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 318(6):83, June 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n6/full/scientificamerican0618-83.html;   http://www.nature.
com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
83.pdf.

Anonymous:2018:SAW

[3614] Anonymous. Science agenda: Who owns the dead? Scientific American, 318(6):10, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
10.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-10.pdf.

Salovey:2018:FKC

[3615] Peter Salovey. Forum: Knowledge can be power. Scientific American, 318(6):11, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-

11.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-11.pdf.

Anonymous:2018:AW

[3616] Anonymous. Advances: Into the wildfire. Scientific American, 318(6):12– 14, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-12.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
12.pdf.

Anonymous:2018:DS

[3617] Anonymous. Deathly scent. Scientific American, 318(6):16, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n6/full/scientificamerican0618-16.html;   http://www.nature.
com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
16.pdf.

Joel:2018:SS

[3618] Lucas Joel. Suffocated seas. Scientific American, 318(6):17, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n6/full/scientificamerican0618-17.html;   http://www.nature.
com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
17.pdf.

Hsu:2018:MWR

[3619] Jeremy Hsu. Milky Way remapped. Scientific American, 318(6):18, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-18.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
18.pdf.

Smith:2018:FF

[3620] Peter Andrey Smith. Frozen fungi. Scientific American, 318(6):19, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-19.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
19.pdf.

Patel:2018:QG

[3621] Prachi Patel. Quirky graphene. Scientific American, 318(6):20, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-20.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
20.pdf.

Anonymous:2018:QH

[3622] Anonymous. Quick hits. Scientific American, 318(6):21, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n6/full/scientificamerican0618-21.html;   http://www.nature.
com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
21.pdf.

Hu:2018:BB

[3623] Jane C. Hu. Bilingual boost. Scientific American, 318(6):22, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n6/full/scientificamerican0618-22a.html;   http://www.nature.
com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
22a.pdf.

Makin:2018:CPP

[3624] Simon Makin. Couch potato persona. Scientific American, 318(6):22, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-22b.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
22b.pdf.

Wallis:2018:SHWc

[3625] Claudia Wallis. The science of health: Why we won’t miss opioids. Scientific American, 318(6):23, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
23.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-23.pdf.

Pogue:2018:TAW

[3626] David Pogue. TechnoFiles: Apps we really need. Scientific American, 318(6):24, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
24.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-24.pdf.

Anonymous:2018:RRRf

[3627] Anonymous. Reviews and recommendations: Recommended. Scientific American, 318(6):75, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
75.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-75.pdf.

Shermer:2018:SSS

[3628] Michael Shermer. Skeptic: Soul-searching. Scientific American, 318(6): 76, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-76.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
76.pdf.

Mirsky:2018:AGBb

[3629] Steve Mirsky. Anti gravity: This book stinks. Scientific American, 318(6):77, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
77.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-77.pdf.

Fischetti:2018:GSG

[3630] Mark Fischetti. Graphic science: A good life for all. Scientific American, 318(6):84, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
84.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-84.pdf.

Boyle:2018:ASL

[3631] Rebecca Boyle. Astronomy: The secret life of the Sun. Scientific American, 318(6):26–33, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-

26.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-26.pdf.

Vance:2018:NCY

[3632] Erik Vance. Neuroscience: Can you supercharge your baby? Scientific American, 318(6):34–39, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
34.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-34.pdf.

Pincott:2018:HPM

[3633] Jena Pincott. Health: A painful mystery. Scientific American, 318(6):40– 47, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-40.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
40.pdf.

Zaraska:2018:ESA

[3634] Marta Zaraska. Ecology: Shrinking animals. Scientific American, 318(6):48–52, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
48.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-48.pdf.

Mukerjee:2018:BQS

[3635] Madhusree Mukerjee. The biggest questions in science: Looking at the stars. Scientific American, 318(6):54, June 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v318/
n6/full/scientificamerican0618-54.html;   http://www.nature.
com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
54.pdf.

Musser:2018:WS

[3636] George Musser. What is spacetime? Scientific American, 318(6):55–58, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-55.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
55.pdf.

Randall:2018:WDM

[3637] Lisa Randall. What is dark matter? Scientific American, 318(6):58–59, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-58.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
58.pdf.

Koch:2018:WC

[3638] Christof Koch. What is consciousness? Scientific American, 318(6):60– 64, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-60.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
60.pdf.

Szostak:2018:HDL

[3639] Jack Szostak. How did life begin? Scientific American, 318(6):65–67, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v318/n6/full/scientificamerican0618-65.html;   http:
//www.nature.com/scientificamerican/journal/v318/n6/pdf/scientificamerican0618-
65.pdf.

Savage:2018:WLM

[3640] Neil Savage. What are the limits of manipulating nature? Scientific American, 318(6):68–72, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
68.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-68.pdf.

Gleiser:2018:HMC

[3641] Marcelo Gleiser. How much can we know? Scientific American, 318(6):72–73, June 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v318/n6/full/scientificamerican0618-
72.html;   http://www.nature.com/scientificamerican/journal/
v318/n6/pdf/scientificamerican0618-72.pdf.

DiChristina:2018:EBBb

[3642] Mariette DiChristina. From the Editor: Barriers between realities. Scientific American, 319(1):4, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
4.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-4.pdf.

Anonymous:2018:Lg

[3643] Anonymous. Letters. Scientific American, 319(1):6–7, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n1/full/scientificamerican0718-6.html;   http://www.nature.
com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
6.pdf.

Schlenoff:2018:YAIg

[3644] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 319(1):75, July 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n1/full/scientificamerican0718-75.html;   http://www.nature.
com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
75.pdf.

Anonymous:2018:SASc

[3645] Anonymous. Science agenda: Safe injection facilities save lives. Scientific American, 319(1):10, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
10.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-10.pdf.

Jackson:2018:FAM

[3646] Rob Jackson. Forum: Auto mileage rollback is a sick idea. Scientific American, 319(1):12, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
12.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-12.pdf.

Main:2018:ARS

[3647] Doug Main. Advances: Rocky secret. Scientific American, 319(1):14– 17, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n1/full/scientificamerican0718-14.html;   http:
//www.nature.com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
14.pdf.

Makin:2018:EE

[3648] Simon Makin. Eagle eye. Scientific American, 319(1):16–17, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n1/full/scientificamerican0718-16.html;   http://www.nature.
com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
16.pdf.

Patel:2018:MM

[3649] Prachi Patel. Microbe mystery. Scientific American, 319(1):18, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n1/full/scientificamerican0718-18a.html;   http:
//www.nature.com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
18a.pdf.

Kornei:2018:FB

[3650] Katherine Kornei. Fish bombs. Scientific American, 319(1):18, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n1/full/scientificamerican0718-18b.html;   http:
//www.nature.com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
18b.pdf.

Lewis:2018:TS

[3651] Tanya Lewis. Thinking slow. Scientific American, 319(1):19, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n1/full/scientificamerican0718-19.html;   http://www.nature.
com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
19.pdf.

Anonymous:2018:PSM

[3652] Anonymous. The power of science museums. Scientific American, 319(1): 20, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n1/full/scientificamerican0718-20.html;   http:

//www.nature.com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
20.pdf.

Lewis:2018:QH

[3653] Tanya Lewis. Quick hits. Scientific American, 319(1):21, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n1/full/scientificamerican0718-21.html;   http://www.nature.
com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
21.pdf.

Patel:2018:SW

[3654] Prachi Patel. Smart walls. Scientific American, 319(1):22, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n1/full/scientificamerican0718-22.html;   http://www.nature.
com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
22.pdf.

Hutson:2018:BC

[3655] Matthew Hutson. Blue’s clues. Scientific American, 319(1):23, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n1/full/scientificamerican0718-23.html;   http://www.nature.
com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
23.pdf.

Wallis:2018:SHK

[3656] Claudia Wallis. The science of health: Killer infection: Is it your genes? Scientific American, 319(1):24, July 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n1/full/scientificamerican0718-24.html;   http://www.nature.
com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
24.pdf.

Pogue:2018:TNR

[3657] David Pogue. TechnoFiles: The next round of merger mania. Scientific American, 319(1):26, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
26.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-26.pdf.

Anonymous:2018:RRRg

[3658] Anonymous. Reviews and recommendations: Recommended. Scientific American, 319(1):72, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
72.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-72.pdf.

Shermer:2018:SFM

[3659] Michael Shermer. Skeptic: The final mysterians. Scientific American, 319 (1):73, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n1/full/scientificamerican0718-73.html;   http:
//www.nature.com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
73.pdf.

Mirsky:2018:AGSb

[3660] Steve Mirsky. Anti gravity: Sex and drugs. Scientific American, 319(1): 74, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n1/full/scientificamerican0718-74.html;   http:
//www.nature.com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
74.pdf.

Fischetti:2018:GSM

[3661] Mark Fischetti. Graphic science: Mental illness overlap. Scientific American, 319(1):76, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
76.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-76.pdf.

Folger:2018:PCQ

[3662] Tim Folger. Physics: Crossing the quantum divide. Scientific American, 319(1):28–35, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
28.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-28.pdf.

Kenrick:2018:PSA

[3663] Douglas T. Kenrick, Adam B. Cohen, Steven L. Neuberg, and Robert B. Cialdini. Psychology: The science of antiscience thinking. Scientific American, 319(1):36–41, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
36.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-36.pdf.

Ungar:2018:ERP

[3664] Peter S. Ungar. Evolution: The real Paleo diet. Scientific American, 319(1):42–49, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
42.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-42.pdf.

Greenwood:2018:BCW

[3665] Veronique Greenwood. Biology: The clocks within our cells. Scientific American, 319(1):50–57, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
50.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-50.pdf.

Fischman:2018:NBB

[3666] Josh Fischman. Neuroscience: Battle of the brains. Scientific American, 319(1):58–59, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
58.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-58.pdf.

Fairley:2018:EBW

[3667] Peter Fairley. Energy: Building a weather-smart grid. Scientific American, 319(1):60–65, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n1/full/scientificamerican0718-
60.html;   http://www.nature.com/scientificamerican/journal/
v319/n1/pdf/scientificamerican0718-60.pdf.

Strickman:2018:PHB

[3668] Dan Strickman. Public health: Buzz kill. Scientific American, 319(1): 66–71, July 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n1/full/scientificamerican0718-66.html;   http:

//www.nature.com/scientificamerican/journal/v319/n1/pdf/scientificamerican0718-
66.pdf.

DiChristina:2018:ESD

[3669] Mariette DiChristina. From the Editor: A sense of discovery. Scientific American, 319(2):4, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
4.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-4.pdf.

Anonymous:2018:Lh

[3670] Anonymous. Letters. Scientific American, 319(2):6–7, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n2/full/scientificamerican0818-6.html;   http://www.nature.
com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
6.pdf.

Schlenoff:2018:YAIh

[3671] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 319(2):83, August 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n2/full/scientificamerican0818-83.html;   http://www.nature.
com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
83.pdf.

Anonymous:2018:SAP

[3672] Anonymous. Science agenda: Prepare for water day zero. Scientific American, 319(2):10, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
10.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-10.pdf.

Freeman:2018:FCC

[3673] Madison Freeman and David Yellen. Forum: Capture that carbon. Scientific American, 319(2):11, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-

11.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-11.pdf.

Hsu:2018:ACS

[3674] Jeremy Hsu. Advances: Collecting space rocks. Scientific American, 319(2):12–14, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
12.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-12.pdf.

Kornei:2018:AW

[3675] Katherine Kornei. Astro worms. Scientific American, 319(2):14, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n2/full/scientificamerican0818-14.html;   http:
//www.nature.com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
14.pdf.

Thompson:2018:TSL

[3676] Andrea Thompson. Taking stock of life. Scientific American, 319(2):16, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n2/full/scientificamerican0818-16.html;   http:
//www.nature.com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
16.pdf.

Nuwer:2018:B

[3677] Rachel Nuwer. Bloodthirsty. Scientific American, 319(2):17, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n2/full/scientificamerican0818-17.html;   http:
//www.nature.com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
17.pdf.

Anonymous:2018:FMF

[3678] Anonymous. Is the friend of my friend my double? Scientific American, 319(2):18–19, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
18a.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-18a.pdf.

Tennenhouse:2018:CH

[3679] Erica Tennenhouse. Caterpillar heartbeats. Scientific American, 319(2):18–19, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
18b.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-18b.pdf.

Amos:2018:CLH

[3680] Amy Mathews Amos. Can logging help jaguars? Scientific American, 319(2):20, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
20.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-20.pdf.

Greenemeier:2018:KR

[3681] Larry Greenemeier. I(KEA), robot. Scientific American, 319(2):22, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n2/full/scientificamerican0818-22a.html;   http:
//www.nature.com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
22a.pdf.

Makin:2018:BS

[3682] Simon Makin. Body sense. Scientific American, 319(2):22, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n2/full/scientificamerican0818-22b.html;   http://www.nature.
com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
22b.pdf.

Miller:2018:QHa

[3683] Maya Miller. Quick hits. Scientific American, 319(2):23, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n2/full/scientificamerican0818-23.html;   http://www.nature.
com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
23.pdf.

Wallis:2018:SHBb

[3684] Claudia Wallis. The science of health: Ba-boom! there goes your hearing. Scientific American, 319(2):24, August 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n2/full/scientificamerican0818-24.html;   http://www.nature.
com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
24.pdf.

Pogue:2018:TIM

[3685] David Pogue. TechnoFiles: I’ll have my AI call your AI. Scientific American, 319(2):26, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
26.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-26.pdf.

Anonymous:2018:RRRh

[3686] Anonymous. Reviews and recommendations: Recommended. Scientific American, 319(2):80, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
80.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-80.pdf.

Shermer:2018:SW

[3687] Michael Shermer. Skeptic: 23 and we. Scientific American, 319(2):81, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n2/full/scientificamerican0818-81.html;   http:
//www.nature.com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
81.pdf.

Mirsky:2018:AGTa

[3688] Steve Mirsky. Anti gravity: That’s life. Scientific American, 319(2):82, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n2/full/scientificamerican0818-82.html;   http:
//www.nature.com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
82.pdf.

Peek:2018:GSS

[3689] Katie Peek. Graphic science: Sunspot surprise. Scientific American, 319(2):84, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-

84.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-84.pdf.

Kipnis:2018:NSS

[3690] Jonathan Kipnis. Neuroscience: The seventh sense. Scientific American, 319(2):28–35, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
28.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-28.pdf.

Hossenfelder:2018:ADM

[3691] Sabine Hossenfelder and Stacy S. McGaugh. Astrophysics: Is dark matter real? Scientific American, 319(2):36–43, August 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n2/full/scientificamerican0818-36.html;   http://www.nature.
com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
36.pdf.

Schwartz:2018:CCU

[3692] Jen Schwartz. Climate change: Underwater. Scientific American, 319(2):44–55, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
44.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-44.pdf.

Piore:2018:SER

[3693] Adam Piore. Special education report 2018: Bringing Darwin back. Scientific American, 319(2):56–63, August 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n2/full/scientificamerican0818-56.html;   http://www.nature.
com/scientificamerican/journal/v319/n2/pdf/scientificamerican0818-
56.pdf.

Dixon:2018:OSM

[3694] Chris Dixon. Ocean science: Monster waves. Scientific American, 319(2):64–67, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-

64.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-64.pdf.

Ornes:2018:MAN

[3695] Stephen Ornes. Mathematics: Art by the numbers. Scientific American, 319(2):68–73, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
68.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-68.pdf.

Margonelli:2018:ETF

[3696] Lisa Margonelli. Ecology: Termites and fairy circles. Scientific American, 319(2):74–79, August 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n2/full/scientificamerican0818-
74.html;   http://www.nature.com/scientificamerican/journal/
v319/n2/pdf/scientificamerican0818-74.pdf.

DiChristina:2018:EVH

[3697] Mariette DiChristina. From the Editor: A very human story. Scientific American, 319(3):4, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
4.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-4.pdf.

Anonymous:2018:Li

[3698] Anonymous. Letters. Scientific American, 319(3):6–7, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n3/full/scientificamerican0918-6.html;   http://www.nature.
com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
6.pdf.

Schlenoff:2018:YAIi

[3699] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 319(3):103, September 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n3/full/scientificamerican0918-103.html;   http://www.nature.

com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
103.pdf.

Anonymous:2018:SAC

[3700] Anonymous. Science agenda: Clinical trials need more diversity. Scientific American, 319(3):10, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v319/n3/full/scientificamerican0918-
10.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-10.pdf.

Nebel:2018:FWS

[3701] Rebecca Nebel. Forum: Why sex matters in Alzheimer’s. Scientific American, 319(3):11, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
11.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-11.pdf.

Crow:2018:ABC

[3702] Diana Crow. Advances: Bone crushers. Scientific American, 319(3):12–15, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
12.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-12.pdf.

Hutson:2018:HL

[3703] Matthew Hutson. An honest liar. Scientific American, 319(3):14–15, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-14.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
14.pdf.

Ponchner:2018:AEF

[3704] Debbie Ponchner. Art-eating fungi. Scientific American, 319(3):16, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-16.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
16.pdf.

Redd:2018:MM

[3705] Nola Taylor Redd. Moon mystery. Scientific American, 319(3):17, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-17.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
17.pdf.

Ackerman:2018:RD

[3706] Daniel Ackerman. Rumbles in the deep. Scientific American, 319(3):18, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-18a.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
18a.pdf.

Joel:2018:DM

[3707] Lucas Joel. Death masks. Scientific American, 319(3):18, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-18b.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
18b.pdf.

Miller:2018:QHb

[3708] Maya Miller. Quick hits. Scientific American, 319(3):19, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n3/full/scientificamerican0918-19.html;   http://www.nature.
com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
19.pdf.

Laursen:2018:OYW

[3709] Lucas Laursen. Oil in your wine. Scientific American, 319(3):20–21, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-20a.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
20a.pdf.

Stringer:2018:IS

[3710] Heather Stringer. The inside scope. Scientific American, 319(3):20–21, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-20b.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
20b.pdf.

Hutson:2018:FGG

[3711] Matthew Hutson. The Facebook gender gap. Scientific American, 319(3):22, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
22.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-22.pdf.

Wallis:2018:SHCa

[3712] Claudia Wallis. The science of health: The so-called right to try. Scientific American, 319(3):24, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v319/n3/full/scientificamerican0918-
24.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-24.pdf.

Pogue:2018:TYD

[3713] David Pogue. TechnoFiles: Not your dad’s hearing aid. Scientific American, 319(3):27, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
27.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-27.pdf.

Anonymous:2018:RRRi

[3714] Anonymous. Reviews and recommendations: Recommended. Scientific American, 319(3):100, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
100.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-100.pdf.

Shermer:2018:SAF

[3715] Michael Shermer. Skeptic: Abortion facts. Scientific American, 319(3):101, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-

101.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-101.pdf.

Mirsky:2018:AGF

[3716] Steve Mirsky. Anti gravity: Furry business. Scientific American, 319(3):102, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
102.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-102.pdf.

Fischetti:2018:GSP

[3717] Mark Fischetti. Graphic science: Peak friends. Scientific American, 319(3):104, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
104.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-104.pdf.

Laland:2018:PWU

[3718] Kevin Laland. Part 1: Why us?: An evolved uniqueness. Scientific American, 319(3):32–39, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
32.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-32.pdf.

Dartnell:2018:TS

[3719] Lewis Dartnell. Techno sapiens. Scientific American, 319(3):40–41, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-40.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
40.pdf.

Suddendorf:2018:IOH

[3720] Thomas Suddendorf. Inside our heads. Scientific American, 319 (3):42–47, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
42.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-42.pdf.

Blackmore:2018:HP

[3721] Susan Blackmore. The hardest problem. Scientific American, 319(3):48–53, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
48.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-48.pdf.

Kenneally:2018:TTT

[3722] Christine Kenneally. Talking through time. Scientific American, 319(3):54–59, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
54.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-54.pdf.

Sherwood:2018:WWD

[3723] Chet C. Sherwood. Are we wired differently? Scientific American, 319(3):60–63, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
60.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-60.pdf.

Wong:2018:PUT

[3724] Kate Wong. Part 2: Us and them: Last hominin standing. Scientific American, 319(3):64–69, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v319/n3/full/scientificamerican0918-
64.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-64.pdf.

Tomasello:2018:OM

[3725] Michael Tomasello. The origins of morality. Scientific American, 319(3):70–75, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
70.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-70.pdf.

Ferguson:2018:WWF

[3726] R. Brian Ferguson. Why we fight. Scientific American, 319(3):76–81, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-76.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
76.pdf.

Schilthuizen:2018:PBU

[3727] Menno Schilthuizen. Part 3: Beyond us: Darwin in the city. Scientific American, 319(3):82–87, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v319/n3/full/scientificamerican0918-
82.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-82.pdf.

Domingos:2018:ODD

[3728] Pedro Domingos. Our digital doubles. Scientific American, 319(3):88–93, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n3/full/scientificamerican0918-88.html;   http:
//www.nature.com/scientificamerican/journal/v319/n3/pdf/scientificamerican0918-
88.pdf.

Gribbin:2018:AMW

[3729] John Gribbin. Alone in the Milky Way. Scientific American, 319(3):94–99, September 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n3/full/scientificamerican0918-
94.html;   http://www.nature.com/scientificamerican/journal/
v319/n3/pdf/scientificamerican0918-94.pdf.

DiChristina:2018:EPI

[3730] Mariette DiChristina. From the Editor: Proof of the impossible? Scientific American, 319(4):4, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
4.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-4.pdf.

Anonymous:2018:Lj

[3731] Anonymous. Letters. Scientific American, 319(4):6–8, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n4/full/scientificamerican1018-6.html;   http://www.nature.

com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
6.pdf.

Schlenoff:2018:YAIj

[3732] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 319(4):83, October 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n4/full/scientificamerican1018-83.html;   http://www.nature.
com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
83.pdf.

Anonymous:2018:SADa

[3733] Anonymous. Science agenda: Druggists shouldn’t be morality police. Scientific American, 319(4):9, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v319/n4/full/scientificamerican1018-
9.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-9.pdf.

Gogos:2018:FND

[3734] Joseph Gogos. Forum: New drugs from old. Scientific American, 319(4): 10, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-10.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
10.pdf.

Kornei:2018:ASS

[3735] Katherine Kornei. Advances: Slippery slope. Scientific American, 319(4):12–14, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
12.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-12.pdf.

Gholipour:2018:SB

[3736] Bahar Gholipour. Seeing blind. Scientific American, 319(4):14, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-14.html;   http:

//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
14.pdf.

Miller:2018:QHc

[3737] Maya Miller. Quick hits. Scientific American, 319(4):15, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n4/full/scientificamerican1018-15.html;   http://www.nature.
com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
15.pdf.

Goldman:2018:FI

[3738] Jason G. Goldman. Flock immunity. Scientific American, 319(4):16, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-16a.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
16a.pdf.

Miller:2018:BB

[3739] Maya Miller. Body balance. Scientific American, 319(4):16, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-16b.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
16b.pdf.

Makin:2018:BBC

[3740] Simon Makin. Brain bar codes. Scientific American, 319(4):17, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-17.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
17.pdf.

Maron:2018:COP

[3741] Dina Fine Maron. A conversation with the only physicist in Congress. Scientific American, 319(4):18, October 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n4/full/scientificamerican1018-18.html;   http://www.nature.
com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
18.pdf.

Joel:2018:M

[3742] Lucas Joel. Missing E.T. Scientific American, 319(4):19, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n4/full/scientificamerican1018-19.html;   http://www.nature.
com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
19.pdf.

Ackerman:2018:HI

[3743] Daniel Ackerman. Heading off injury. Scientific American, 319(4):20, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-20.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
20.pdf.

Learn:2018:TC

[3744] Joshua Rapp Learn. Tinder for cheetahs. Scientific American, 319(4):21, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-21.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
21.pdf.

Smith:2018:PR

[3745] Dana G. Smith. Postpartum relief. Scientific American, 319(4):23, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-23.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
23.pdf.

Wallis:2018:SHCb

[3746] Claudia Wallis. The science of health: Coming down from opioids. Scientific American, 319(4):24, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
24.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-24.pdf.

Pogue:2018:TJA

[3747] David Pogue. TechnoFiles: 5G is just around the corner. Scientific American, 319(4):25, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
25.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-25.pdf.

Anonymous:2018:RRRj

[3748] Anonymous. Reviews and recommendations: Recommended. Scientific American, 319(4):80, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
80.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-80.pdf.

Shermer:2018:SMC

[3749] Michael Shermer. Skeptic: A mysterious change of mind. Scientific American, 319(4):81, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
81.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-81.pdf.

Mirsky:2018:AGTb

[3750] Steve Mirsky. Anti gravity: True story. Scientific American, 319(4):82, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-82.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
82.pdf.

Fischetti:2018:GSH

[3751] Mark Fischetti. Graphic science: Honey, I shrunk the mammals. Scientific American, 319(4):84, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
84.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-84.pdf.

Cubitt:2018:MSA

[3752] Toby S. Cubitt, David Pérez-García, and Michael Wolf. Mathematics: The un(solv)able problem. Scientific American, 319(4):28–37, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-28.html;   http:

//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
28.pdf.

Borel:2018:AIC

[3753] Brooke Borel. Artificial intelligence: Clicks, lies and videotape. Scientific American, 319(4):38–43, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
38.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-38.pdf.

Vance:2018:SES

[3754] Erik Vance. Seismology: Earthquakes in the sky. Scientific American, 319(4):44–49, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
44.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-44.pdf.

Anonymous:2018:SWS

[3755] Anonymous. State of the world’s science 2018: How to fix science. Scientific American, 319(4):50–51, October 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n4/full/scientificamerican1018-50.html;   http://www.nature.
com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
50.pdf.

Ioannidis:2018:RF

[3756] John P. A. Ioannidis. Rethink funding. Scientific American, 319(4):52– 55, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-52.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
52.pdf.

Palus:2018:MRR

[3757] Shannon Palus. Make research reproducible. Scientific American, 319(4):56–59, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
56.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-56.pdf.

Moskowitz:2018:EH

[3758] Clara Moskowitz. End harassment. Scientific American, 319(4):60–61, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-60.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
60.pdf.

Boyle:2018:HYS

[3759] Rebecca Boyle. Help young scientists. Scientific American, 319(4):62–64, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-62.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
62.pdf.

Worthy:2018:BS

[3760] Graham A. J. Worthy and Cherie L. Yestrebsky. Break down silos. Scientific American, 319(4):64–67, October 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n4/full/scientificamerican1018-64.html;   http://www.nature.
com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
64.pdf.

Murray:2018:NRB

[3761] Andrew J. Murray. Neuroscience: Rabies on the brain. Scientific American, 319(4):68–73, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n4/full/scientificamerican1018-
68.html;   http://www.nature.com/scientificamerican/journal/
v319/n4/pdf/scientificamerican1018-68.pdf.

Duenas-Osorio:2018:NDW

[3762] Leonardo Dueñas-Osorio, Devika Subramanian, and Robert M. Stein. Natural disasters: This way out. Scientific American, 319(4):74–79, October 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n4/full/scientificamerican1018-74.html;   http:
//www.nature.com/scientificamerican/journal/v319/n4/pdf/scientificamerican1018-
74.pdf.

DiChristina:2018:EWW

[3763] Mariette DiChristina. From the Editor: When will we learn? Scientific American, 319(5):4, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
4.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-4.pdf.

Anonymous:2018:Lk

[3764] Anonymous. Letters. Scientific American, 319(5):5–6, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n5/full/scientificamerican1118-5.html;   http://www.nature.
com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
5.pdf.

Anonymous:2018:SADb

[3765] Anonymous. Science agenda: Dereliction of duty. Scientific American, 319(5):8, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
8.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-8.pdf.

Szalavitz:2018:FII

[3766] Maia Szalavitz. Forum: Income inequality and homicide. Scientific American, 319(5):9, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
9.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-9.pdf.

Maron:2018:AFW

[3767] Dina Fine Maron. Advances: Fixing wasting muscles. Scientific American, 319(5):10–12, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
10.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-10.pdf.

Perkins:2018:AW

[3768] Sid Perkins. Artificial wood. Scientific American, 319(5):12, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-12.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
12.pdf.

Blaszczak-Boxe:2018:FS

[3769] Agata Blaszczak-Boxe. Flooding the senses. Scientific American, 319(5):14–15, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
14a.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-14a.pdf.

Hutson:2018:LL

[3770] Matthew Hutson. Lifelong learning. Scientific American, 319(5):14–15, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-14b.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
14b.pdf.

Anonymous:2018:DAS

[3771] Anonymous. Decoding alien senses. Scientific American, 319(5):16, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-16.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
16.pdf.

Learn:2018:UA

[3772] Joshua Rapp Learn. Unwanted advances. Scientific American, 319(5):17, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-17.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
17.pdf.

Hutson:2018:TWL

[3773] Matthew Hutson. Time for women leaders. Scientific American, 319(5):18, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-

18.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-18.pdf.

Moskowitz:2018:MM

[3774] Clara Moskowitz. Measure for measure. Scientific American, 319(5):19, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-19.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
19.pdf.

Hornyak:2018:SPG

[3775] Tim Hornyak. To sleep, perchance to gene. Scientific American, 319(5):20, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
20a.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-20a.pdf.

Paliwal:2018:QH

[3776] Ankur Paliwal. Quick hits. Scientific American, 319(5):20, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-20b.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
20b.pdf.

Wallis:2018:SHWd

[3777] Claudia Wallis. The science of health: Why oral cancer threatens men. Scientific American, 319(5):22, November 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n5/full/scientificamerican1118-22.html;   http://www.nature.
com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
22.pdf.

Pogue:2018:TDM

[3778] David Pogue. TechnoFiles: Driving Mr. Pogue. Scientific American, 319(5):24, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
24.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-24.pdf.

Anonymous:2018:RRRk

[3779] Anonymous. Reviews and recommendations: Recommended. Scientific American, 319(5):80, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
80.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-80.pdf.

Shermer:2018:SFE

[3780] Michael Shermer. Skeptic: The fallacy of excluded exceptions. Scientific American, 319(5):81, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
81.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-81.pdf.

Mirsky:2018:AGPb

[3781] Steve Mirsky. Anti gravity: Pets gone wild. Scientific American, 319(5):82, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
82.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-82.pdf.

Schlenoff:2018:YAIk

[3782] Daniel C. Schlenoff. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 & 150 years ago. Scientific American, 319(5):83, November 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n5/full/scientificamerican1118-83.html;   http://www.nature.
com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
83.pdf.

Anonymous:2018:GSW

[3783] Anonymous. Graphic science: Watching North Korea. Scientific American, 319(5):84, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
84.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-84.pdf.

Paller:2018:NSL

[3784] Ken A. Paller and Delphine Oudiette. Neuroscience: Sleep learning gets real. Scientific American, 319(5):26–31, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n5/full/scientificamerican1118-26.html;   http://www.nature.
com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
26.pdf.

Marlow:2018:MTP

[3785] Jeffrey Marlow and Rogier Braakman. Microbiology: Team players. Scientific American, 319(5):32–39, November 2018. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v319/
n5/full/scientificamerican1118-32.html;   http://www.nature.
com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
32.pdf.

Coe:2018:ABT

[3786] Dan Coe. Astronomy: Back in time. Scientific American, 319(5):40–47, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-40.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
40.pdf.

Duchin:2018:MGV

[3787] Moon Duchin. Mathematics: Geometry v. gerrymandering. Scientific American, 319(5):48–53, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v319/n5/full/scientificamerican1118-
48.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-48.pdf.

Anonymous:2018:SRS

[3788] Anonymous. Special report: The science of inequality. Scientific American, 319(5):54–55, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
54.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-54.pdf.

Stiglitz:2018:RE

[3789] Joseph E. Stiglitz. A rigged economy. Scientific American, 319(5):56–61, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-56.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
56.pdf.

Sapolsky:2018:HWG

[3790] Robert M. Sapolsky. The health-wealth gap. Scientific American, 319(5):62–67, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
62.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-62.pdf.

Eubanks:2018:AB

[3791] Virginia Eubanks. Automating bias. Scientific American, 319(5):68–71, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v319/n5/full/scientificamerican1118-68.html;   http:
//www.nature.com/scientificamerican/journal/v319/n5/pdf/scientificamerican1118-
68.pdf.

Boyce:2018:ECI

[3792] James K. Boyce. The environmental cost of inequality. Scientific American, 319(5):72–77, November 2018. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v319/n5/full/scientificamerican1118-
72.html;   http://www.nature.com/scientificamerican/journal/
v319/n5/pdf/scientificamerican1118-72.pdf.

Aczel:2015:FZM

[3793] Amir D. Aczel. Finding zero: a mathematician’s odyssey to uncover the origins of numbers. Palgrave Macmillan, New York City, NY, USA, 2015. ISBN 1-137-27984-2 (hardcover). x + 242 pp. LCCN QA141.2

.A29 2015. URL http://www.netread.com/jcusers2/bk1388/842/
9781137279842/image/lgcover.9781137279842.jpg.

Ansari:2015:MR

[3794] Aziz Ansari and Eric Klinenberg. Modern romance. Penguin Press, New York, NY, USA, 2015. ISBN 1-59420-627-9 (hardcover). 277 pp. LCCN HQ801 .A595 2015.

Cobb:2015:LGSa

[3795] Matthew Cobb. Life’s greatest secret: the story of the race to crack the genetic code. Profile Books, London, UK, 2015. ISBN 1-78125-140-1 (hardcover), 1-78283-002-2 (e-book). xiv + 434 + 16 pp. LCCN QH450.2 .C63 2015.

Cobb:2015:LGSb

[3796] Matthew Cobb. Life’s greatest secret: the race to crack the genetic code. Basic Books, New York, NY, USA, 2015. ISBN 0-465-06267-9 (hardcover). xiv + 434 + 16 pp. LCCN QH447 .C62 2015.

Fontani:2015:LEP

[3797] Marco Fontani, Mariagrazia Costa, and Mary Virginia Orna. The Lost Elements: the Periodic Table’s Shadow Side. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-938334-0. xxxvii + 531 pp. LCCN QD467 .F66 2015.

Roberts:2015:GPC

[3798] Siobhan Roberts. Genius at play: the curious mathematical mind of John Horton Conway. Bloomsbury, New York, NY, USA, 2015. ISBN 1-62040-593-8. ???? pp. LCCN QA29.C66 R63 2015.